A Complete Bibliography of Publications in Linear Algebra and its Applications: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.75

Title word cross-reference

(−1, 1) [ÁAFG12]. (0, 1) [BBS12b, NP10, Ghe14a]. (2, 2, 0) [CI13, PH12]. (A, B) [PP13b]. (α, β) [HW11, HZM10]. (C, λ, μ) [dMR12]. (ℓ, m) [DFG10]. (H, m) [BOZ10]. (κ, τ) [CSZ10, CR10c]. (λ, 2) [BBS12b]. (m, s, 0) [GH13b]. (n − 3) [CGO10]. (n − 3, 2, 1) [CCGR13]. (ω) [CL12a]. (P, R) [KNS14]. (R, S_{σ}) [Tre12]. −1 [LZG14]. 0 [AKZ13, Ano12-30, CGGS13, DLMZ14, Wu10a]. 1 [Ano12-30, AHL^{+}14, CGGS13, GM14, Kal13b, LM12, Wu10a]. 1/n [CNPP12]. 1 < t < 2 [Seo14]. 2 [AIS14, AM14, AKA13, BDF11, BdlC14, BDK11, CvDKP13, CL13b, CNPP12, DoMP09, Ere13, GMT13, GG13, KY14, Rim12, Rud12, YHH12, vdH14]. 24 [KAAK11]. 2n − 3 [BCŠ10, Hil13]. 2 × 2 [CGRVC13, CGSCZ10, CM14, DW11, DMS10, JK11, KJK13, MSvW12, Yan14]. 2 × 2 × 2 [Ber13b]. 3 [BZWL13, Bre14, CILL12, CKAC14, Fri12, GOvdD14, GX12a, Kal13b, KK14, YHH12]. 3n^{2} − 2\sqrt{2}n^{3/2} − 3n [MR13]. 3n^{2} − 2\sqrt{2}n^{3/2}3n [MR14a]. 3 × 3 [Dru14, GLZ14, Sev14]. 3 × 3 × 2 [Ber13b]. 3 × 3 × 3 [BH13b]. 4 [Ban13a, BDK11, BZ12b, CK13a, FP14, NSW13, Nor14]. 4 × 4 [CJR11]. 5 [BH13b, CHY12, KRH14, Kol13, MW14a]. 5 × 5 [BAD09, DA10, Hil12a, Spe11]. 5 × n [CJR11]. 6 [DK13c, DK11, DK12a, DK13b, Kar11a]. 7 [PP13a, Zho12]. 70 [GRS^{+}11]. [1, n] [LWY10]. _{k} [HJN12]. A [AM13d, CK14, DJK12b, HM14b, LSTW13, MS14a, PPK13]. A + λB [MP14a]. A − BX − (BX)^{∗} [Tia11b]. A_{T,S}^{(2)} [SPKS12]. A^{∗}x = λAx [LCwCL11]. A^{3} = λA [Özd13]. {a_{i}}_{i=0}^{d} [Han13b]. A ∈ C^{n,n} [BGH12]. α [CM12a, Cra13, FKM13]. A ⊙ x = B ⊙ x [LdlP11]. AX + X^{∗}B = 0 [DDG^{+}13]. AXB = C [AG10]. B [AT14a, QS14, YY14c, Beh13, Sin10a]. B(X) [LJ10]. B_{0} [YY14c]. B_{D} [JM12b]. B_{ℓ_{p}} [BS14]. c [CN11c, CN13c, LL11c]. C∗ [AF12]. C^{∗} [AR12, BB14, BMGMC12, CIH11, GD11a, Kaw13, Pop12, Sha11, WW10, WD10]. C^{2} [GMV11]. C^{⋆} [Kaw12]. C_{4} [ZW12a]. A(n, 2) [Ghe14b]. D [BR14c, FMR12, Kal13a, Mit11b, NP10, Pry10]. D_{2t} [ÁAFG12]. DB [Dai11]. d × d [DHLX12]. E [BZW14]. E_{8} [KAAK11]. ℓ^{1} [WY13]. ℓ^{∞} [BEM12b]. ℓ^{p}(I) [BEM12a]. ℓ_{1} [Fou14]. ℓ_{p} [BS14]. \frac{3}{2} \sqrt{2} [LL13a]. G [CS13b, FH12a, FM13, GMS12, Mat12, AMJ14, CGMS10b, WL10a]. Γ [Lim11b, Aga14, LH10b, SR13b]. {Γ(A^{m})}_{m=1}^{∞} [CK14]. {Γ(A^{m})}_{m=1}^{∞} [PPK13]. GF(2) [MMP13a]. GL(V) [KN13a]. H [BGH12, CMRR13, FGvRR13, GEP10, QSW14, wXL14, wXZ19]. H_{2} [Kar11a]. ∞ [GL12c, LHWL12]. J [BFdP10, BLdPS10, BFdP11, BFdP12, FKR11a, FKR12, HSS10, HSS14, dlRMP12]. K [CC14, FFK11, ACM^{+}12, BM14b, BR14d, BdS10, CR10a, CH14, DCIW12a, DGMS14, DW12a, Fra13, GLS10, HL10b, LSC10, fLyH11, LWZ11, LLH10, LL14b, Nik11, SvdH11, Tre10, Zha14a, vdH13]. {K, s + 1} [LRT12, LRT13]. k − 1 [TL13b]. K_{2} × K_{n} [GK14a]. K_{3} [CPH11]. K_{4} [CPH11]. K_{4,4} [Gol13]. K_{5} [Kuz10]. L [SS12a, Fri13]. L^{2} [Car11, MZ11]. L_{p}[0, 1]\ ∪_{q>p} L_{q}[0, 1] [BFPSS12]. λ [FdCR10, MP14a]. Λ_{S} [MPT12]. LDU [CRU13, CRU14]. ⌊(n − 9)/2⌋ [Zhu12c]. log(XY) = log(X) + log(Y) [Bou10a]. LR [BM13a]. LU [Dur12, GL12b]. M [BCEM10, BCEM12, BR14c, Guo13, HLS10, HZ10, HH11a, JS11, NS11b, SHZ10, Bal12a, Bal12b, BMN13b, FJ14, Gu14, GK14c, Hua12b, MPSS10, PYZ14, TNP12, Vla12]. M_{11}(E) [KdM13]. M_{n}(Z_{k}) [MPRW11]. M_{n,m} [AAM12]. C^{m×n×l} [Fri13]. F_{q}[x] [GY13]. R [KK14]. R^{d×d} [Dai13]. R^{N} [San10, FGR13, IJ12, SA14, AMJ14]. R_{n} [Wu13b]. Z^{+} [JPS13]. A [DJK12a]. B(H) [DM11]. M [DQW13]. ⌋ [BEM12b]. ∫↕_{∈} [NT12b]. ∫↕_{∈}(C) [BE10]. GL(n; C) [MR11]. SO(n + 1) [DMMY10]. SO(n, 1) [DMMY10]. m ≤ n [SSM13]. m × n × (m − 1)n [SSM13]. μ [NT10b, dF10]. n [BE12, Dug12, Hil13, JLLY10, MR10c, Pin12, WWD13, WW13c, Wan14b, ZHZF13]. n − 3 [ACGVK14, CCGR13]. n − e [LLS12]. N_{0} [ND11, ZH11b]. n × n [Hil13, MR13, MR14a]. Ω_{n}(T) [ACM14]. P [AdFM11, BPDC14, DdF14, GTR12, MPS10, PP12b, TGS14, YY14c, CCGVO13, CCGO14, Guo10a, HL11b, JMP13, Lin10, MTS11, OM14, Wol12]. p > 0 [BFPSS12]. Pmax^{1} [GM11a]. p^{n} [Gu13, LWG13]. P_{0} [AT11, YY14c]. P_{2} [CPV10]. P′_{2} [CPV10]. ϕ_{J} [DoMP09, MPP10]. ϕ_{S} [AMP10, DMP11, GMP13]. {±1, ±i} [KP14a]. ψ [BC14a]. Q [Lee13b, AHL11, Bal10, BZ12b, Cer10, CT11b, MW14a, ND11, OdLdAK10, PHS13, Siv13, WBWH13, Yua14, Yua15, ZWL13, dFdADVJ10, BC14a, CJL13, Ern13, HWG14, IRT14, LL11a, LT11a, NT10b, VS11, VS14b, Wor13, ZYL10]. q + 1 [MM10a]. QDR [SPKS12]. QR [BV12a]. R [Kaw13, BNP11, ZWL13]. {R, s + 1, k} [CLST14]. R^{k} = I [LSTW13]. R_{−1} [CFK10a]. RA = A^{s+1}R [LSTW13]. RH_{2} [Köh14]. RH_{∞} [Köh14]. S [Drn13a, SBM11, dlCdlRMP14, QSW14]. Smax [GM11a]. σ [wH13, GEP13]. sl_{2} [BM12a]. \sqrt{2} [Sta12]. ⋆ [AKM13, DJK12b].

∑_{j=1}^{m} A^{n−j}XA^{j−1} = B [Fur10b]. T [BM14a, Blu10, ZXX13, BKMS12]. T (x) = f [HN10]. T^{∗} [LCM13]. T_{n} [GS12a]. tan θ [Nak12]. θ [GL12c]. U_{1} [YX13]. U_{q}(∫↕_{∈}) [Ter13, BCT14, Wor13]. U_{q}(sl_{2}) [BC14a, Ter14]. UT_{∞}(R) [Slo12b]. W [CL12b, KK10, PP12b, CH11, CDDY10]. W (2, 2) [CL12b]. W (n, n − 1) [KMNS12]. w_{23}(v) [AAKM14]. X [MP13b]. X + A^{T} X^{−1}A = Q [GKL11]. X = Af(X)B + C [ZLD11]. XA + AX^{T} = 0 [CGGS13, DD11a, GS13a]. XA − AX = f(X) [Bou11]. ξ [QCH11]. XM = NX [Lim11b]. xy [KLZ14a]. XY + Y X^{∗} [LLF13]. XY − Y X^{T} [CFL13a, FLC11]. yx [KLZ14a]. Z [CPR10, CPZ13, GTR12, TG13, XC13]. Z_{2} [Cen11]. ||A^{−1}||_{∞} [HZ10].

* [GD11a, XDFL10, YZ12b]. *-Lie [YZ12b]. *-modules [GD11a]. *-order [XDFL10]. *congruence [DFS14].

-adic [ZYL10]. -admissible [CS13b]. -algebra [AKM13, CL12b]. -algebras [BB14, BMGMC12, CIH11]. -alternating [BM14a]. -analogue [CJL13]. -analogues [Ern13]. -arithmetic [MPSS10]. -banded [Hua12b]. -bialgebra [Kaw12]. -bialgebras [Kaw13]. -cell [KAAK11]. -class [DJK12b, MW14a]. -claws [Ban13a]. -colored [Kal13b]. -commutative [Tre12]. -Commuting [DW12a]. -completable [HJN12]. -complete [Sin10a]. -conjecture [NT10b]. -connected [CH14]. -contractions [MS14a]. -cospectral [BZW14]. -critical [FdCR10, MPS10]. -curves [KK10]. -cycles [QSW14]. -cyclic [GLS10, LL11c, TL13b]. -decompositions [GMS12, KNS14]. -derivations [BE12, HW11, HZM10, WW13c, Wan14b, WWD13]. -diagonally [LH10b]. -digraphs [GL12c]. -dimensional [BDF11, DK13c, CK13a, CILL12, DK11, DK12a, DK13b, FP14, KRH14]. -divergence [CM12a]. -domination [LL14b]. -Drazin [CGMS10b]. -eigenvalue [AHL11]. -eigenvalues [CPZ13, QSW14, WBWH13, XC13]. -entropy [CH11]. -extensions [LCM13]. -Fibonacci [DGMS14]. -filiform [CGO10, CCGVO13, CCGO14]. -form [BDK11]. -forms [BDK11]. -free [ZW12a]. -function [SS12a]. -game [wH13]. -generalized [CMRR13]. -graded [Cen11]. -graphs [LHWL12]. -Hermitian [BFdP10, BLdPS10]. -Householder [MPT12]. -hyperreflexivity [BR14d]. -hyponormal [CDDY10]. -index [Yua15, CT11b, Yua14]. -integral [PHS13, ZWL13, dFdADVJ10]. -interpolating [BPDC14]. -invariant [PP13b]. -inverses [WL10a, SPKS12]. -involutory [fLyH11, Tre10]. -isometries [BMN13b, Gu14, Dug12]. -Jacobi [HSS10, HSS14]. -Krawtchouk [Wor13]. -letter [AM14]. -Lie [BZWL13, QCH11]. -Local [AKA13]. -Lucas [DGMS14]. -matching [Beh13]. -matrices [BK10, Dah10, Ghe14a, Zho12, Kaw13, ÁAFG12, AT11, AT14a, BCEM10, CPR10, Dai11, FH12a, FM13, FFK11, GEP10, Guo13, HLS10, HZ10, HH11a, JS11, JPS13, Mat12, ND11, SHZ10, Siv13, wXL14, wXZ19, ZH11b, ZXX13]. -matrix [BCEM12, BR14c, BGH12, Drn13a, NS11b]. -minimizations [Fou14]. -modules [AR12, AF12, Pop12, Sha11, Ter13, Ter14, WW10, WD10]. -negative [Wol12]. -Newton [JMP13]. -norm [MTS11]. -normal [BFdP10, BFdP11, BFdP12]. -numerical [CN11c, CN13c]. -observers [Blu10]. -odd [LWY10]. -operators [JLLY10]. -optimal [Mit11b, NP10]. -optimality [FMR12]. -orthogonal [VS14b, dlCdlRMP14, AMP10]. -palindromic [BM14a]. -parameters [JLLY10]. -paranormal [DJK12b]. -partial [AM13d]. -partite [Zha14a, ZWL13]. -Pascal [VS11]. -paths [QSW14]. -permanent [Cra13, dF10]. -polynomial [Lee13b, Cer10, MW14a]. -positive [MR10c, FGvRR13]. -positivity [GK14c]. -Potapov [FKR11a]. -potent [DCIW12a, CLST14, LRT12, LRT13]. -Primitivity [BM14b]. -properties [CPV10]. -property [Bal10, GTR12]. -quasiseparable-Vandermonde [BOZ10]. -Racah [BC14a, HWG14, NT10b]. -radius [OM14]. -reducible [Kar11a]. -Regular [GX12a, CSZ10, CR10a, CR10c, MM10a]. -regularity [SBM11]. -replicated [PYZ14]. -robustness [MP13b]. -root [Bal12a, Vla12, Bal12b]. -SDD [GEP13]. -seminorms [GD11a]. -separation [AHL^{+}14, vdH14]. -skew [LL11a]. -spectra [BZ12b]. -splittings [JM12b]. -spread [OdLdAK10]. -stability [Kal13a]. -stable [BR14c, BBS12b]. -Stieltjes [FKM13]. -strong [Car11]. -submanifolds [Lim11b]. -sum [AKZ13]. -summing [SR13b]. -tensors [Fri12, DQW13]. -term [BKMS12]. -tetrahedron [IRT14]. -th [TNP12]. -theory [CC14]. -Toeplitz [GG13, Rim12]. -transformations [TG13]. -trees [SvdH11, vdH13]. -tridiagonal [AMJ14]. -unconditional [LM12]. -variable [KY14]. -vectors [Aga14]. -vertices [AdFM11, DdF14]. -walk-regular [CvDKP13, DFG10]. -way [ZHZF13]. -weighing [NP10].

1 [Bar10b, Grü12, Rhe10, Zha12a]. 1-norm [LMYY11]. 15th [Ano10u]. 160pp [Gle11]. 16th [BBG^{+}13]. 17th [BFBD13].

2 [Lim13b]. 2008 [Ano10u, BBD^{+}11]. 2009 [BDK^{+}10]. 2010 [BBG^{+}13]. 2011 [BFBD13]. 2012 [Joh12]. 2nd [BBD^{+}11, Zha12a].

3-rose [LH13]. 3-transitive [MW12a].

4 [Rod12a]. 4-Regular [CLL13a]. 433 [MR14c]. 436 [CdGS20, Duk15, LT16, Vla12]. 437 [Kis15, WZ14c]. 438 [Che14b, KKL13a, LMO16, MR14a]. 439 [DWW14]. 454 [KB14a]. 455 [HS14a]. 458 [wXZ19, Yua15]. 463 [EKSV18].

7 [Gle11, Tam12]. 70th [Bai11].

978 [Bar10b, Gle11, Grü12, Lim13b, Rod12a, Tam12, Zha12a]. 978-0-387-40087-7 [Zha12a]. 978-0-387-68276-1 [Zha12a]. 978-0-521-46193-1 [Grü12]. 978-0-521-89881-2 [Lim13b]. 978-0-691-12157-4 [Gar12]. 978-0-691-12889-4 [Rod12a]. 978-0-691-14039-1 [Bar10b]. 978-0-691-14503-7 [Gle11]. 978-1-4614-1098-0 [Tam12]. 978-1-4614-1099-7 [Tam12].

A-m-Isometric [AS12a]. Abaffy

[MAGR13]. Abaffy-Broyden-Spedicato [MAGR13]. abelian [BZWL13, Car13, CS13b, DGZ13, DJ14, GL10a, Kuz10, LW12d, WY14b]. Aberth [BN13, GN13]. above [Pop13a]. Abraham [Ano13f]. Absolute [Nak10, HW14b, Kra12, LMM11, ST10a, ZHQ13]. Absolutely [SR13b, Pel14]. absorbing [Nem13]. abstract [ACG14, Cam13]. accelerate [GOSV12]. acceleration [PE13, ZY12b]. accessibility [SS10a]. according [AM14]. Accretive [Nie10, Lin13]. accretive-dissipative [Lin13]. Accurate [MM10b, KP14b]. ACI [BC12c, BC13, BHZ10, HZ11b]. ACI-matrices [BC12c, BC13, BHZ10, HZ11b]. across [PS12]. acting [BM13b, Mer10]. action [BG14b, Che14a, Fis14, WLG12]. actions [BDV12, CC14, Lim14, Rod11]. Acyclic [NS13, ACDM14, DdF13b, DdF14]. Aczél [Mos11]. adaptive [Wal11b]. addition [Fra12, Liu14b]. additional [DS13]. Additive [BS11b, xCwXL11, JH10, MD10, PIM^{+}10, CGMS10b, EN11, QCH11, Sun13]. additive-nilpotency [Sun13]. Additivity [ACG13a, BG13, Wan11a, CGMPSS14]. ADI [Jbi10]. adic [ZYL10]. Adin [Alo14]. adjacency [AFLN12, AAF^{+}12, Bap13b, BB10, CCF^{+}12, HO11, HTW13, ST10b, Wil14, XC13, YFW10]. adjoint [ADW13, BT13, CN10c, CN11d, DL14, DZ12a, HSZ12, OR12, RS14a, SS11d, SSR13, vBM13]. adjoint-commuting [CN10c, CN11d]. adjointable [FMWW12, WD10, XCS13]. adjustment [GOSV12]. admissible [CS13b]. admit [AT11, CGMJ14]. Affine [BDV12, AY13, AN13, BB13a, BV11, BB11b, Bud11, DS11, Dau12, KLP13, Lee13b, dSP12a, Wal11a]. affine-linear [BB13a]. affinity [SWA12]. after [JMP10, LSR11, tHR13]. agent [Sha14a, Zhu11b, ZY14]. agents [AIP12]. aggregation [PQ12]. Agler [BKV14]. AIC [HYF14, YiKIS12]. Aigner [WZ14a]. AINV [Raf14]. Akiyama [Rah13]. Albert [Zha12a]. Algebra [BFBD13, BPRY11, Che14b, CdGS20, DWW14, Duk15, EKSV18, HS14a, KB14a, KKL13a, Kis15, LMO16, LT16, MR14a, MR14c, Vla12, Wik12, wXZ19, Yua15, AKM13, ABK14, BM12a, BCS13a, BR11, BCT14, BC14a, BE10, BM13c, Bre14, BDV12, BM13d, aCCS14, Cas10, CGMS10b, Cen11, CL12b, Cir13, CM14, DKS10, DD10c, DLMZ14, Dub14, EvdD10, Fie11a, FKLT13, GPT14, GST13, GM11a, GM11b, GMS13, IRT14, KSS12, Kla10, Kon13, KdM13, LLR14, Lee13b, MLC^{+}10, Mar13b, MW10, MP13a, MP13b, MP14b, NT12b, Pep11, PTPL10, Ros12a, Ser13, SH13, SLS13, Tra12, TP13, Wil11, Wor13, WZ12, WZ13e, WZ14c, ZW12b, ZZ10, FdC12]. Algebrability [Nat13, BGP11, Wód14]. Algebraic [BFF^{+}11, BJ10a, Pin11, Yan10b, AJRT14, AHS10, BL12, BJ13, CKS10, Cra13, Das10a, Das13, DS12b, DGZ13, FZ13, FH14, Guo10b, GL10c, GSL11, Guo13, HM10, JLN13, KOJ11, LwCJL11, LGSC14, MZ12c, PLS14, Per14, RJH11, RR12, RMAJ10, Ser11, VS10, WKV10, WT12]. algebras [ACGVK14, AM13b, Ago14, AIS14, AAJ12, AKA13, BD12a, BZWL13, BDF11, BGP13, BdlC13, BdlC14, Ben11, BS12b, BG14b, BB14, BM13b, BS11c, BS13a, BS12e, BF14d, BMGMC12, CT14a, CGO10, CCGVO13, CCGR13, CCGO14, CK13a, CLR11, CILL12, CLOK13, CKLO13, CLOR13, CNT12, CM11a, CN10c, CL11b, CdGS12, CdGS20, CMZ10, CIH11, DDF13a, DIP13, DKS10, DH12b, DW12a, DW12b, DW13, ES13, FP14, GZH14, Gha13, GS10c, GTR12, GS10d, GS12e, HW11, Han11a, HP12a, HM14b, HW14c, HZ12a, Ika11, JV10, JQ11, JZ11, JH10, Kaw12, KO13, KRH14, Kim12, KLZ10, KLZ14a, KZ10, LMO16, LL11a, LW12c, LW13a, LLF13, LL11b, LW12d, LNT13, LMMR13a, LMMR13b, MD12a, MD12b, Mar13a, MFGD14, Mar14a, Mar14b, Mol12, Mos13, PY10, PYZ14, PR13a, Per13, QH10, QCH11, QH13, RY12a]. algebras [RRKK12, RM14, SS12b, Sev13, ŠK10, SM10b, SM12, Sze14b, Tao11, Tao13, TKLX14, TDT13, UW11, WLG11, WC12, WGL12, WWD13, WW13c, Wan14b, WX11, Wei10, Wu13b, WY14b, XW10a, XW10b, XW12, YZ13, YZ10, YZ12b, ZZ11, ZZW10, ZHQ14, ZZ12, ZXZ10]. algorithm [BM13a, BBE^{+}10, CFPP13, CILL12, GOSV12, HN10, KY13, KM14, LL13b, LdlP11, LQ11, MSP13, MAM^{+}13, RAAGAVS11, Roh11, SA10, Wal11b, WJ12, XY11, XSS13, ZCQ13]. algorithmic [MPS10]. algorithms [BSS10, CLX13, Dum13b, Fis14, GNE^{+}14, Gle11, HR11, HB12, Kam10, LMN13, Qua10, Reg13, UV13, Wil13]. alignment [YZ12a]. All-derivable [ZZW10, ZZ11, ZZ10, ZZ12, QH10]. Allan [Gar10]. allow [BDM^{+}12, OS14]. Almost [MM10a, BD13, DvDF11, Fid10, HLZ12, Peñ14, PS14b, Wor14, dSW12]. alternant [JL12]. Alternating [BE10, LKN13, BM14a, BKMS13, BC14c, FHS14a, JJKS11, MMMM10]. alternative [Bar12a]. Aluthge [BJ10b, HT10b]. ambiguous [JMS11]. among [FJ14, TL13a, VDVJT13]. Analogs [BFdP10]. analogue [CJL13, Kra13]. analogues [Ern13]. Analysis [BFRR14, Dum13b, HJ12, NT14, AS10, BHMR12, Cas13, CQYY13, Est12b, GP13a, GNE^{+}14, GMV11, Güv12, HYF14, Hua13a, HZ14, KG12a, LYS13, Lop11b, Oto12, PPZ14, PS14a, SBM11, SS13b, TH11, XSS13, Gre13]. Analytic [AL13c, Nik14, FP11, GMS13, Zha13]. analyzing [GHT11]. ancient [SS10c]. Anderson [PE13]. Ando [Lim11a]. anisotropic [KG14]. Annihilator [LP10]. Annihilator-preserving [LP10]. annihilators [WZ12, WZ14c]. Announcement [Ano12a]. Anti [YT13b, BH14b, CCL14, DH12a, KB12, SN14, XSH14]. anti-commuting [KB12]. Anti-diagonals [YT13b]. anti-magic [CCL14]. anti-norms [BH14b]. anti-reflexive [DH12a]. anti-topical [SN14]. anti-triangular [XSH14]. Antichains [Ghe14a]. antieigenvalues [GS10e, Sed11]. antieigenvectors [GS10e, Sed11]. antinegative [Per11]. antireflective [DH10b]. antiring [Tan11]. antirings [Tan10b]. antisymmetric [Jai11, Rub13]. any [AGV12, LdlP11, dSP10b, ZHQ14]. anzahl [GL14, GLW13]. apart [MAS12]. APOS [PO10]. appearing [FGQ11]. Appl [Che14b, CdGS20, DWW14, Duk15, EKSV18, HS14a, KB14a, KKL13a, Kis15, LMO16, LT16, MR14c, Vla12, Wik12, wXZ19, Yua15, MR14a]. Appl. [WZ14c]. application [Aud10b, Dah12b, DS13, DD11a, Est12b, FMR12, FNY13, FL10, GH13a, GP13a, JJKS11, LQY13, LHZH11, LZL12, LJY13, MLC^{+}10, MSvdD14, Mou12, OM10, VR12, Zha12d]. application-based [MLC^{+}10]. Applications [Grü12, MOA11, Rei11a, SRdAG10, BL13a, BFH^{+}12, BK12, CTW11, DMMY10, FFS11a, FG13b, GZX14, GLW13, GL14, HTS14, HN10, HZGY12, Jim10, KLS12, KLP12, LRT12, LZ10, LCwCL11, MGLW11, MR11, Mol11, NPP13, Nie13, PM10, Per14, Sah10, Sha13b, SYH14, ySpW12, SS13f, Tia10, Tia11b, Tia12, TT11, WFM11, ZZC14, dFBRS14, dMR12, AM13b, BPRY11, Zha12a]. Applied [BFBD13, Gar12, Rod12a, GMH14b, GMRS14, Sag13, SBM11]. approach [AFHP14, BDD13, BMS10, BOZ10, BOZ11a, BOZ11b, Böt13, BBS12a, CA10, Ern13, FDS13, Fuh10a, GHMPVP11, KMNS12, KKLY14, Liu13, MPS10, MR14b, NP13a, Nem13, PKR12, RRM11, RR12, SHZ10, Sto12, TD11, VS13, Wan14a, ZYL10, Zha13]. approaches [TmYsH11]. Approximate [Ano12a, KH13, KPRT14, BDD13, CP11, GB13, LPK14]. approximately [Wój14a]. Approximating [SCS11]. approximation [AGNS11, BV12a, BGK13, BV13b, CAV13, CN11e, DS13, DD11b, DBZZ14, LB14, fLyH11, OT10, yPjXL11, Reg13, SA10, SR12, SC10, Zha12d]. approximations [DZ12a, GO13, LS14, Ney11, Pin12]. Araki [Aud13a]. arbitrarily [Tra13]. arbitrary [BCS13a, BJZ12, Bot10b, Bot12, CSC13, Car11, FKW13, GS11a, GS12a, GS13b, Ma11, MGSW14, dSP10c, dSP12c, RS14b, WZL13, Wu10b]. Arc [Kuz10]. Arc-transitive [Kuz10]. arcs [YW11]. arise [GLS13a]. arising [BTYZ12, CL13b, DKOT12, DBZZ14, GL10c, Jim10, Pan11, RR14, SK14, Shi12a, Wik11, Wik12, Wu13a]. arithmetic [KLL11, MPSS10, ST10b, WLL14, Yam13, BT11c]. arithmetic-geometric [Yam13, BT11c]. ARMAX [KM14]. Arnold [Zha12a, LV14]. Arnoldi [LBLS12, MV12]. arrangement [CGW14]. arrangements [Buj13]. array [BH13b]. arrays [AMPT13, BH12, CC10, He11, LMMS12, OT10, ZHZF13, Ter14]. Artin [FP13]. Arveson [Far11a]. ascent [DDK14, GW14a]. aspect [CK13d, NR10]. aspects [CvDKP13]. asset [DBZZ14]. assignability [KBS13]. Assignment [LPK14, BCY12, MD13, WZ13b]. associated [AKN12, ARZ11, BBdH13, BC12a, BCF12, CH11, CN12b, CN13a, FDS13, FKR11b, FKR12, Guo13, HHMS10, He11, HKPR13, HNZ12, HHLS14, JKN14, KHG14, Kim13d, fLyH11, LLMZ12, MS14a, San10, dFBRS14]. Association [GH13a, GH13b, GZX14, WLGM11, MGLW11, MW14a]. associative [Ago14, BS12e, FP14, WC12]. assumptions [DLNN14]. Asymmetric [JW13, BRZ11, BR10]. Asymptotic [CvDKL10, FF10, FV13b, HHT13, Tre11, VS10, VV13, Dum13a, GB13, GZ12, GM11a, Jun12]. Asymptotically [Lan14]. asymptotics [BT14a, BMM12]. atoms [BBH^{+}12]. attaining [FCL10, SM13]. attainment [SP13]. attraction [Ser11, SCSS10]. augmentation [KOJ11, PQ12]. augmented [PPZ14]. August [BDK^{+}10]. authorities [BEK13]. AutoGraphiX [Ste10]. automata [KKS12]. automated [AH10]. Automorphism [ÁvW11, GST13, HK13, LMW12]. automorphisms [DH12b, GS12e, Mol11, WMZ14]. autonomous [BM10]. Average [PS14a, Sha14a, BMW10, GHMPVP11, KL13b]. averages [BJ11, PP11b]. Axiom [JS13a]. Axiomatic [LM10b]. Axiomatization [XWD13].

back [HB12]. backward [AA11, CLCL12, GJTP13, HZ14]. Bade [AGPPF12]. Bakonyi [Rod12a]. Balance [KG12a]. balanced [Cha12, GO13, Güv12, HLP12].

Balancedness [Bel14]. Ball [MM10b, ST12]. balls [FKR12]. Banach [AF12, BR11, BDFP11, BMGMC12, CGMS10b, DP10, DX13, ES13, For14, GD11a, HZ11a, Hua11a, HZGY12, HZJ12, JZ11, Kec13, LL11a, LMM11, LM12, LT13, MM13, Mos13, PR13a, QH10, SH13, SM12, WC12, YW10, ZW12b, ZHQ14]. Banachiewicz [CGMS10a]. band [Sei14]. band-dominated [Sei14]. Banded [Irv12, Ang13, Hua12b, SC13]. banks [MZ12c]. Bannai [HWG13]. Bannai/Ito [HWG13]. Bapat [Sat14]. bar [AY13]. Barabanov [Mor10]. Barbour [Nak13]. Barel [Gem10]. barrier [XSS13]. Barry [Zha12a]. Bart [Ano13d, Ano13e]. Bartholdi [MS14b, SS13a]. Bas [Lim13b]. base [CL10b, Din11, WLHL10, YW11]. based [BT10, BFS11, BB10, CH12, CFJKS13, CM12a, CNPP12, CS10b, Dah12b, FS11, GH13a, GH13b, GZX14, GH11, Ika11, IPFD13, IIK^{+}13, KOR14, KMS13, LL10d, LL14b, MLC^{+}10, MR14b, RJH11, SHZ10, DDP14, WAH13, WMZ14, wXL14, wXZ19]. Bases [BG11, ySpW11, Tan14a, BPDC14, KP13, KS11a, LM12, LLH10, Mel14, MSS14, PT14, ySpW14, YS13]. Basic [SW11, Tam12, FH10a, FH12b, FH13, FH14]. basics [BS10]. basis [AL13a, BM12a, CJR11, GN13, GK14a, MM10c, Ros12a, Wu10b]. bath [DGU14]. Baxter [Kaw13]. be [FW14b, OS14]. behavior [Dum13a, HHT13, PE13, RR11]. behaviors [NT11a, Pol12, ZHQ14]. behaviour [VV13]. being [DHLX12]. Bell [BHAP12]. Bellman [MMM13]. belong [dSP12d]. below [MQ11, dSP12a, Pop13a, TD13]. Berger [Koz14a]. Berlekamp [KY13]. Berlekamp/Massey [KY13]. Berman [Ano13f]. Bernoulli [CD14]. Bernstein [Bar10b, Aud12, MM10c, Sla10]. Best [AGNS11, BV13b, Pin12, SC10]. beta [ANF11, BRA11, DGJ10]. Bethe [RJ10, RJ11]. between [ÁvW13, AH13b, BV12b, BAD09, CFJKS13, CN10a, CN10c, DXG12, DMMY10, DdC13, HRT13, KR10, LRT13, LT11b, MZ12a, MM13, Pit11, RW10, SB11]. betweenness [Aud13b]. beyond [Suz13]. Bezout [IW13, Hua12a]. Bezoutian [Wu10b]. Bezoutians [ER13, Ros12c]. BFGS [EM12]. Bhatia [Dru12a]. Bhattacharyya [Wat13]. bi [WML13]. bi-regular [WML13]. bialgebra [Kaw12]. bialgebras [Kaw13]. bialternate [EM11]. Bias [YiKIS12, HYF14]. Bias-corrected [YiKIS12]. bicircular [Ili10b]. Bicyclic [WZY14, YFW13, HJLS11, LS12b, LZ14, LL10e, XZ13b, ZZ13]. Biderivations [DW13, Gho13]. Bidiagonal [HL11c, Hua13b]. Bijective [LT10a, Per11]. Bilinear [KLZ14a, AIS14, ABEV10, DFS12, FV13a, GK11, KKLY14, Mar13b, RO10, Sim10, Sze14a, WFM11]. Billiard [Ter14]. bimatrix [ANF11, BRA11, DGJ10]. binary [CDM12, KS13a, LM10a, Wu13a]. Binet [Kni14, Kon13]. Bini [Lim11a]. biography [BELK12]. biorthogonal [HL11e]. bipartite [Alt13, BG14a, COvdD10, CFK^{+}10b, CKST11, DC13, DC14, FTA11, GM14, Han14b, HW14c, HWG14, HLS11, HJLS11, KS13b, KS14c, LwW14, LSC11, Mey12, NP12, Row14b, SS13a, SSGL10, Ste11]. Bipartiteness [FF12]. bipartition [LY11b]. bipolar [Bha13]. biregular [BB11b, Fio12]. Birkhoff [ACDM14, AR12, CP11, CM10b, GS10a]. birthday [Bai11]. bisection [UZ14]. Biswa [BPRY11]. bisymmetric [DH10a, YWX13]. bit [ALPV14, MSP13]. bit-flipping [MSP13]. bivariate [Bal14, BL10a, Hen10]. Black [BGK13]. Blakley [Pat12a]. Block [BDG13, HSS10, RV13, BB13b, BTYZ12, BLL12, CFPP13, CH11, DW11, DMS10, DK13e, Gar13, GK12, HSS14, Hua13a, KS14b, Koz14b, KKM13, KM13b, Lin14b, LHZH11, MS11a, Mos13, NT14, Rub13, TGS14, XSH14, Yan14, Zha14b, ZBW12]. block-circulant [BTYZ12, CFPP13]. block-matrices [BLL12]. block-operator [DW11]. Block-oriented [HSS10]. blockcyclic [KN13a]. blocks [BC12b, BLL13, GC12, Zha14b]. Board [Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h]. Bohnenblust [NA13, SR13c]. Bohr [MMA11]. Bol [HP12a]. Bonan [BDK11]. Book [Gre13, Rod12a, Tam12, Zha12a]. Boolean [AK11, Bea12, Bre14, BRLS12, CK14, Kim11a, Kim13a, LaG12, LT10a, LT11b, Shi13]. border [Fri13]. both [FKR11b, FKR12]. Botha [KN13c]. Bott [ACG14]. Böttcher [Lu12]. bound [AHS10, Cha10, CGR14, CTG13, DTS11, DTL11, FCL10, Grc10, GR10, GCY14, HW14b, JMS11, Kim11a, KP12, MY14, Pat10, RJH11, RMAJ10, RL13b, Zim13]. boundary [jASZ12, jAS13, CN11a, FN11, HSZ12, vBM13]. Bounded [GMMFPSS12, lH10a, lH11b, And13, BJ10a, BHZ10, Cau11, Dom13, JRMFSS12, LT12b, LT16, Lim10, MQ11, dSP12a, Pep12, Pop13a, WY13]. boundedness [TD13]. Bounding [ZL11, AFHP14]. Bounds [ACG^{+}11, AK12a, BT14a, BMN^{+}13a, BL10a, BK12, GH11, HL10a, LWV12, LL14b, MNZ10, OdLdAK10, Pep12, Ruk14, Zhu11a, AGM14, AdF11, AdFST11, AP14, Aud10b, BB10, BHKM13, BL11, BdS10, BS13b, xCwXL11, CR10a, CGTR14, Cha12, CW10, CLS13, CFL13b, Dai11, DLS14, DHC12, DZ12c, DZ13, FTA11, FH12c, GEP10, GEP13, GJTP13, GO13, GLS13b, HJZ13, Jai11, tLyLWqW10, LCZ10, LMYY11, MZ10, MNZ12, MA10b, PJ13, Psa12, Rad10, RJ10, SK13, SBMT10, SWT13, DDP14, TC13, Wal14, WXH10a, WL10a, WA10, WZL12, XX14, XZ14, YZ12a, YWS11a, ZLW12, ZCWZ13, Zhu10b]. box [BGK13]. Boyle [Laf12]. branch [CK13b]. Brauer [HT14]. Braunschweig [BFBD13]. brickwork [Böt13]. Bridges [Kra13]. Brioschi [JL12]. Broyden [MAGR13]. Brualdi [Dru12b, HT14]. Bruce [Joh12]. Bruhat [BF14a, CDM12, Ghe14b]. Brunovsky [BBdH12]. BTTB [LZ11a]. Bunch [HLZP13]. Bürmann [Ma10a, TX12].

C [BELK12, Zha12a]. Cabri [Kla10]. cacti [BNP12, LZ12b, PAS11b, RRM11, XZ13c]. calculating [Tia12]. calculation [FS11, JT11b]. calculus [Cas13]. Callebaut [MMA12]. Cambridge [Grü12, Lim13b]. can [Cha13a, FW14b, Tra13]. Cancún [Ano10u]. CANDECOMP [ZHZF13]. CANDECOMP/PARAFAC [ZHZF13]. Canonical [BBdH13, KB12, CT11a, CKAC14, FHS11, HSZ12, HYF14, NSC13, Oto12, Rad13]. Cao [ZC14]. Carathéodory [CH11, FKR11b]. cardinality [KP13, ySpW14]. cardinals [CGMPSS14]. Carlson [DS12c, LL13b]. Cartan [RAAGAVS11, dlCdlRMP14]. Cartier [FP13]. case [BBdH12, BCŠ10, BC12b, CPH11, CT11b, Dor10, Fid10, Mat13a, PS14a, SCS11, Sko11, TZ12a]. cases [FKR11b, FKR12]. Cassels [Nie10]. Catalan [WZ14a, CMS12a, He13, SS10e]. Category [DLMZ14]. caterpillars [RMAJ10, Roj11]. Cauchy [BFdP10, FT10, Fie10, JK13, Kni14, Kon13, Sch10]. Cauchy-type [Sch10]. Cayley [AKN12, BH14a, BKV14, BBS12a, CGW14, DJ14, FFS11b, FTZ12, GL10a, HFS13, Hwa11, Hwa12, KAMS11, KM12, MSvW12, O’D14, Sch11]. cell [KAAK11]. cells [GRdS12]. Center [WZ12, WZ14c, XWD13]. center-of-gravity [XWD13]. central [CL11b, Slo12a]. Centraliser [AAK^{+}14]. centralizer [GS13b]. Centralizers [ZZC14, DGKO13, GS12a]. Centralizing [Liu14a, LW12c]. centro [Wik11, Wik12]. centro-invertible [Wik11, Wik12]. century [Abe11]. certain [AM13c, ÁNPQ12, Boj13, BS11c, COvdD10, CGW14, CJ14, DU14, FMR12, He14, Hia13, Hua13b, Jim10, LL10a, MM10a, MR12, MS12, Siv13, Sra13, XSZ13, dCF12]. Ceva [CH13]. Chain [RM14, BLLX11, CLR11, CDM12, Hun14, Kir10, Kir14, LS12c, Wu10b, dF10]. chained [HZ10]. chains [Cas13, Din11, DGU14, Ghe14b, Góm10, Hun10, MA10b, Nem13, PR10, Puł11, Rhe10, RS12c, Sku13, TL13a]. change [BBdH12, GS11b]. chaotic [VR12]. Chapman [Vse12]. character [Kra12]. Characteristic [AW13c, HP11, Lav10, VW10b, AIS14, BBC^{+}14, BMS10, BC14b, BDOvdD12, CCGR13, Cen11, EV11, FDS13, GX12b, GMT13, GW11, GWZ13, HLW14, HT10a, MMP13a, dSP10d, QY12, SS13f, WLLX11, WML13]. characteristics [MMRR12]. Characterization [ABBO11, Bal10, BC13, KNS14, KM13b, LSR11, MW10, MD12c, QH13, Tre10, Tre12, VS13, ALRV12, AF12, BV13a, BBS12b, BZ12a, CHY11, CHY12, GW13b, Han11b, Han13b, Han14b, Kho12, KK14, Kum11, LZ10, LwW14, LS12b, LHW11, LW12e, LHWL12, LH13, LHL13, LX13, NS12c, NS12b, Per12, PE13, RM10, Spe11, WZL13, XWS12]. Characterizations [BP12, BR11, CRU10, Den11, HLS10, JQ11, LRT12, LJ10, QH10, TW11a, DX13, FFK11, KR12, KK10, KK12, KK13, Ste11, WB12]. characterized [Che14a, Lu11]. characterizes [SS10a]. Characterizing [DFG10, FHRT14, QCH11, ZHQ14, LZ13, GMMFPSS12]. charges [OM12]. Charles [Gar12]. Chebyshev [Coh14, ES11, PT14, ST12]. Chemnitz [Bar13b]. chicken [RM14]. Chief [Ano11z, Bru11, Bru13, Bru14]. Chiò [Abe14]. Choi [Fur11, Ha13]. Choice [Tra13]. Cholesky [BM13a, HHT13]. chordal [Kak10, MNZ12, SM13]. chordal-structured [Kak10]. Chromatic [AAJ12, YWS11b]. circle [BR12, DW10, RL13a]. circuits [SS10b]. Circulant [CSAC10, CSAC11, CFLW13, PZVJ11, BP10, BTYZ12, CFPP13, DGMS14, Ili10a, LS12a, MP13b, MS14c, RE11, SS11b, Sbu10, SA14, TW14, VW10b, Wil14, EGR12]. circulant-Hankel [MS14c]. circulants [Mey12]. Circular [MSP11]. class [AS12b, ANP13, BT13, BK12, CT11a, Cho13, CDM12, DL14, DH12b, Dra12b, DJK12b, DJK12a, DK13e, FMR12, KL13a, Kho12, KS13a, KJK13, LRST10, LW13a, MW14a, ND11, NSC13, RR14, Sag11, ŠŠ11e, TZ13a, TmYsH11, WGL12]. Classes [MS13a, BD12b, CRSS14, Car13, Dah12a, Dra14, FM11b, FV13b, FG13c, FJ14, GS11a, GN14, JNS13, Kau12, Kus13, LRT13, LGSC14, MMRR11, NdM13, NPP13, PdFDV14, RRKK12, SCS11, Sev13, SYH14, WW13d]. Classical [CN11d, Ada14, BR14b, CN11c, CN10c, FG13c, GL14, KSS12, VS13, VS14b, WLGM11]. Classification [BS13a, BB11b, CLOK13, Cir14, Dub14, LMO16, PC10, Rom14, ŠK10, DK13a, Bud11, CK13a, CILL12, FRS14, GWH13, HWG13, wH12, INT11, KO13, KRH14, PS14b, RO10, RS12c, SSS13, Sku13, dOHKS12, AAT12a, Wag11]. Classifying [Ago14]. claws [Ban13a]. clean [BCDM13]. cleanness [TZ12b]. clique [AAJ12, DL13, GLS12, HJZ13, Lav10, ZHG13]. cliques [MR12, RM10]. Close [Car10b]. closed [BEM12b, CN12c, DX13, Fra12, HZ11a, Hua11a, Kau12, KR10, Liu14b, Net10, dSP11b, SR13c, Sha11]. closest [Köh14]. closure [Bar10a, Bar13a, DDK14]. closures [BG11]. cloth [Gar12]. cluster [Sev13]. Clustering [SCSS10]. CMV [BOZ11b, BOZ11a]. co [ACM^{+}12]. co-neighbor [ACM^{+}12]. coalesced [RW10]. coalescence [ABBO11]. Coalescing [DP12a]. Coalitional [Cec10]. cocharacter [CM14]. cocyclic [ÁAFG12]. code [LSV12]. codes [AAK^{+}14, ANP13, CNPP12, CGM^{+}10, CT14b, La 14, LSV12, PC10, TS12, Wu13a, dSC14]. Codimension [DKS13b]. coding [FFS11a]. coefficient [Bos11, HLP12, Kir14]. coefficients [BHAP12, DKOT12, HL11a, Hwa11, LTS13, LY11b, MK12, QY12, Sev13, wTmS12, wTlW13, VS14b, ZZ13]. cogenerator [BO12]. Cohen [Cam13]. coherent [HW14b]. cohomology [CT14a]. Coimbra [Ano12-30]. coincide [tHR13]. coincidences [Pel14]. cokernel [BCD10]. Colin [GB11, Gol13]. collections [DKM^{+}14]. Collocation [JK13]. color [MD12b]. colored [Kal13b]. colorings [Moh10]. column [AAM12, BBdH12, BBdH13, Bar12b, CC10, Dod13, HS12b, PP12b]. column- [PP12b]. column-finite [HS12b]. column-majorization [AAM12]. columns [PS12]. comaximal [DTL11]. combination [Kis15, dSP10b, XX12]. combinations [DCIW12a, DCIW12b, JMP13, KCID12, dSP10c, RJ11, SH13, SS10e, ZW12b]. Combinatorial [BMS10, FFK11, BBS12a, CD14, CNT12, Cha13a, Für10a, KKLY14, LaG13, Rah13]. Combined [FM11b, Blö12]. Combining [GOSV12, Reh11]. coming [WLGM11]. Comment [XLG^{+}13]. Comments [WY14a]. Common [Bou13, WW10, MM12, RSS10, SLS13]. communicability [Est12a]. communication [DS12b, RvS13]. commutant [MM12]. commutative [AKN12, AKA13, CRS14, DCIW12b, HTS14, Lak10b, Mar14b, MMA12, Seo13, ySpW14, Tan14a, Tan14b, Tre12, dO12]. commutativity [DLNN14, DK14, GN14, LHL12, LL10b, Liu14b, LSH12]. Commutator [GH12, KN10, BDF11, CFL13a, FLC11, Kha13, Lan10b, WA10]. Commutators [Bie14, CVW10, EV11, Aud10b, FCL10, GB13, HK10, KLS12, OR12]. Commute [BRZ11, HMS13, Kis15, Ogu13, XX12]. commutes [FMM13]. Commuting [DO11a, DW12a, Fra12, Fra13, XW10a, BC12a, Bou13, CN10c, CN11d, Hwa12, KSS12, KY14, KB12, LD12, Mig13, NS14, Pet10, Sar14, Siv12b, Siv12a, TZ13b, dOHKS12]. Compact [LT13, AK12b, BV13b, GP14, SM10b, TT12]. compactly [GHMPVP11]. Compactness [MN12b, DDK14]. Companion [EKSV14, EKSV18, MR11, BBE^{+}10, DDM12, Gau10, GS10d, GS12c, GS12d, LD11, MZ13, Mac13, Pat12b, DDP13, DDP14]. comparing [WNM13]. Comparison [DXG12, HTS14, BSKL13, BBM14, EvdD10, GEP10]. compensators [BO12]. competition [Kim10, Kim11a, KP12, Kim13a]. complement [BBF^{+}12, DW11, DC13, LHZH11, LZL12, Mit11a, Ney11, SvdH11]. complementarity [AS10, Bal10, CPV10, Dai11, GEP10, GEP13, JV11, SVP11, wXL14, wXZ19]. complementary [ACM14, FH10a, FH12b, FH13, FH14]. complementation [DV10, GLS13a, Tra12]. complements [Ago14, BEV13, FZW11, GS10c, HL10b, LW12b, LH10b, NY13, RTR10, Row12b, Row14b, Row14d, SC12, ZH11b, vdH13]. completable [HJN12]. Complete [HLS11, JZ14, BCS13a, CFK^{+}10b, CI10, FGS11, HTS14, HP11, Kuz10, PHS13, Raf14, Row14b, Sin10a, Zha14a, ZWL13]. Completely [BS12c, BP11, LT12b, LT16, SS13d, Aud12, BAD09, DA10, GST13, HKPR13, HH11b, Lin14a, Tao13, BSU14]. completely-Q [Tao13]. Completion [Dod13, MS10b, BCŠ10, BC12b, Dod10, DS14, JJKS11, JW13, KBS13, KKR11, LV14, SS11a]. Completions [Rod12a, BW11, BHZ10, CH11, HZ11b, HZ12b, MQ11, MQ13, Rub13, dSW12].

Complex

[DGJ10, GKL11, HS12a, JS13a, Pry10, Szö13, ANF11, BB13d, BGP13, BG12b, BW13a, Buj13, Byd10, CILL12, CK13c, FP14, GPT12, GPR13, GP13b, Ikr10, Kar11a, Kar11b, KK14, LRT13, LNT13, MZ11, MARC13, Mat13a, NV12, Ref12, SH13, dlCdlRMP14]. complexity [DS12b, KKL10, Shi13]. component [TH11]. Components [JZZ13, ABBO11, LSR11, RW10]. Componentwise [Hua13a, HZ14, Miy14, LWY14]. composite [ZXX13]. composition [LL11a]. compositions [AM14]. compound [Bap13a]. compressed [ALPV14, DPF10, KG14]. compression [MZ12b]. Compressions [ADW13, RS12b]. compressive [PS14a]. compressors [BS11b]. Computable [BdS10]. computation [BHAP12, KD12, SPKS12, TD11]. Computational [HS12c, NR10]. Computations [Gem10, HWSH13, MM10b, PIM^{+}10, PQZC11, PS14b, SE13]. Computer [CS11]. Computing [BBH^{+}12, BI13, FHS14b, KG12b, ST12, BGV12, BMS10, Fis14, LJY13, MV12, MP14a, RBP12, Roh11, Row12a, Sha14b, Uhl13, WJT13]. concave [Aud13a, FGQ11]. Concavity [Hia13, Nie11]. concept [PO10]. concerning [Bap10, BG12b, He14, LGZ14]. concise [Rho10]. condensed [BD13, Ji12a]. Condition [DDP13, AAK11, AL13a, AHAPP10, Alt13, BB13c, DH10b, Gar13, LJ11, LWY14, LS12d, XD12]. conditions [jAS13, CHK^{+}13, EGR12, FN11, FS14b, HSZ12, Hu10, JKN14, LLD13, Nak12, Sha13a, ZHZF13, vBM13]. Cone [Kus12, Nie12, AV12, BB13c, DDGH13, FW14b, GST13, Hil12a, ISYY11, Jun12, Lim14, LTX14, NN10, Pro10, Sko11, WXH10b]. Cone-theoretic [Kus12]. Cones [Sko11, AGK11, AL13b, Bar12a, BZ12a, CPH11, CFL13b, JV11, JSS13, KL13b, LT10b, LXK12, RSS10, San10, Ser11, TG13, VS10]. Conference [Ano10u, BBD^{+}11, BBG^{+}13, BFBD13]. configuration [AFLN12]. configurations [FW14b, HW14b]. confluent [KS14b]. congruent [Ikr10]. Conics [Wu13a, Mir10]. conjecture [Bap10, BBF^{+}12, CH12, CW12a, Cho13, Das11, Drn13a, Dru11, Dru12b, EGR12, GS12a, GC12, GKR13, He14, HJLS11, LYL13, LL14c, Mig13, Mit11a, NT10b, YY14a, YHY14, ZW12a, Zha13, ZC14, Dru13, Yan11]. Conjectures [Das10a, AH10, CHZ13, Das10b, Das13, HL10a, Hil14c, JNS13, Ste10]. Conjugacies [BV12b]. conjugacy [BB11b, Dau12, FG13c]. conjugation [Mat13b]. Connected [CT11b, LM10a, SBMT10, CH14, CK14, HY14, LS12b, MY14, PAS11a, ZLW12]. connectedness [UZ14]. connecting [GLP^{+}13]. connection [Bar13b, BR14b, RJH11]. connection-graph-stability [RJH11]. connections [DK13a]. connectivity [AJRT14, AHS10, BL12, Cio10a, Cio10b, Das10a, Das13, Guo10b, GSL11, KOJ11, LGSC14, LLT13, RJH11, RMAJ10, RL13b, Row14d, WKV10, WT12, YFW10]. Connes’ [HH12a]. consecutive [CY11, EWY12]. Consensus [ZY14, AIP12, Est12b, Sha14a, Zhu11b]. consequences [AJ13, DK13a, MMK13]. consistency [FDS13]. Consistent [HC14]. conspecific [AIP12]. constant [Cir14, FM11a, FP13, Koz10, LL10a, MQ14, RS12a, RS14b]. constantly [MO14]. constants [AP14, BT14a, BRZ13, BR14d, NA13, SR13c]. constrained [Byd10, CMRR13, Dai12, KBS13, yPjXL11, SA10, TH11]. constraint [HMR12, SS13b]. constraints [Cim11, DS13, Koz14b, MR10a]. constructed [BM12a, CLST14]. constructible [CL14b]. Constructing [ES14, HRW99, JM12a, MAM^{+}13, Wad14]. Construction [CN11b, HL11e, KRvS14, MS13c, PP11a, PO10, ABBO11, BW13a, CMNW14, CNPP12, HNZ12, IPFD13, JNS13, VS13]. Constructions [GOvdD12, ZXX13, CGM^{+}10]. constructive [Laf12]. contained [GL10b, KK14]. containing [BS12e, HLW10a, MAGR13, Ros12a, Ziv12]. continued [GWZ13]. continuing [Mac13]. continuity [GMMFPSS12, Pep11]. Continuous [ZWZ10, BS14, CI14, GG13, JLW11, KKS12, LT13, LTX14]. contraction [AT14b, GW10, Li10]. Contractions [Pop14, BV12b, DJK12b, DJK12a, MS14a]. Contractive [KS14a, BDS13]. contribution [Nik10a]. Contributions [BPRY11]. control [NT11a]. Controllability [FS14a, KRvS12, ST13]. controllable [BBdH12, BBdH13]. controlled [GV11]. controllers [KBS13, Per12]. conventional [Aud10a, JLW11]. converge [PPK13]. Convergence [GMV11, GL10c, Jar12, KKR11, SK14, XSS13, ZCQ13, BJ13, CQYY13, DYW14, DZ12a, FS14b, GOSV12, LKN13, Pry10, Sku13, TMSS14, Yan10b]. convergent [BGP11, CC12, LJY13, Reg13]. conversion [BSS10]. Convex [AL13b, MO14, San10, SMC11, Aud13a, Bar12a, BH11a, Dra12a, OT12, PCC12, Rez13, Wei11, Zha14c]. Convexity [Nor11, Alt13, CN13c, Hen10, MPP11, NS11c, WZ14a, WA10]. convolutional [ANP13, CNPP12, PC10, La 14]. coordinated [KRvS12, KRvS14]. copositive [DDGH13, Hil12a, Hil14b, JV10, Qi13, XY11]. coprime [TL13a]. Corach [CMS12b]. core [RDD14]. Cored [HQS13]. corepresentations [Dor10]. cores [BSST13]. Corinth [SS10c]. corollaries [Hil13]. corona [CRS14]. coronae [CT12, LL13c, MM11b]. corrected [YiKIS12]. correcting [LHG10]. correction [HYF14]. correlation [Cha14, DBZZ14, HH12a, HYF14, Oto12, SA10]. corresponding [AK11]. Corrigendum [Che14b, CdGS20, DWW14, Duk15, EKSV18, HS14a, KB14a, Kis15, LMO16, LT16, MR14a, MR14c, Wik12, WZ14c, wXZ19, Yua15]. cosine [LdS13]. cosparse [GNE^{+}14]. cospectral [BZW14]. Cospectrality [AO14]. countable [Car11]. counterexample [Dru11]. Counterexamples [AH13b, CL14a]. counterpart [FLS10]. counterparts [MO14]. Counting [AG12, GM12, Shp10]. counts [DKS13b, MW14c]. coupled [AS10, BS13a, DH10a]. coupling [Tim14, tHR13]. couplings [dSW12]. covariance [CFPP13, Güv12]. cover [Sin10c, TF10]. coverage [ACM14]. Covering [KN13a, KN13b, NCdS14, SS13a, Suz13, dSC14]. coverings [CS13b]. covers [CJ12, Kuz10, MM10a]. Coxeter [Lak10a, PS14b, Sim10]. cp [BSU14]. cp-rank [BSU14]. CPCA [TH11]. Craig [Car10b, ZY12a]. Cramer [Ji12a]. Crawford [Uhl13, WWG10]. credit [VV13, Vas14]. Criteria [Mor10, Dai13, Lim11a]. criterion [Far11a, FFG^{+}11, Gna12, KL12, LJ12]. Critical [CN12b, YHH12, AK12a, CV13, FdCR10, GLZ14, GKR13, JLW11, MPS10, PS14b]. cross [LJ12, OT10, Sav14]. cross-validation [LJ12]. crosspolytope [HLW10a]. Crouzeix [Cho13, GC12]. cryptography [Cha13b]. cube [HMP^{+}11]. cubelike [CG11]. cubic [He14, LHL13, Row14a]. Curriculum [Ano13d]. curvature [Cir14]. curve [TZ12a, XW11]. curves [ABGPSS14, Buc10, CM12b, CN12b, FP13, HK13, KK10, OZ10]. Cut [ST10b, GX12c, Row12a, WF10, Zhu10a, Zhu11a]. Cut-and-paste [ST10b]. cut-points [GX12c]. cut-vertex [Row12a]. cutsets [BMN^{+}13a]. Cvetkovic [GRS^{+}11, GRS^{+}10]. Cycle [BB11a, DF14, GK14a, LS12c]. Cycles [dOFK^{+}13, AJRT14, ABS14, Mir10, MSvdD14, Nik10b, QSW14, RRM11, Yua14, Yua15]. Cyclic [Kec13, SS10a, CN13a, CN13b, DK12b, GLS10, JPS13, KZ11, LL11c, LMT10, NCdS14, TL13b]. Cyclotomic [Gre12, BHKM13]. cylinders [KK10].

D [CNPP12]. Dai [BBGM12]. d’algébre [Cas10]. Danny [BELK12]. Darboux [CGMPSS14, DD11b, Nat13]. Darboux-like [CGMPSS14, Nat13]. data [DS13, KD12, MS13c, Wei13b]. Datta [BPRY11]. deblurring [FN11]. Decay [BB14, EY13, CSV14]. decisions [DT10]. Decomposable [FdCR10]. decomposing [dSP10b]. Decomposition [AR10, UW11, BFRR14, BH12, BL13a, Bha13, BLL12, BCMT10, CC10, CKAC14, CM10a, DoMP09, DMP11, Duk12, Duk15, GL12b, GMP13, Güv12, Kak10, Kim11b, KKB11, LKN13, Liu13, MM11a, MPP10, Pag12, Pon11, Rho10, Sar14, Seg14, SWA12, ZHZF13]. Decompositions [BKM^{+}13, AS12b, BKV14, CM11b, FRS14, GMS12, KNS14, KR12, SC13]. decreasing [CGTR14]. Dedekind [Ens10]. Dedication [Ano12b, Bai11]. Defect [GW10, Tad12, Ban13b, BDS13, Rud12, WW13d]. defective [ABBO11]. defined [DV14, Kaw12, Lak10a]. definite [AAT12a, AHAPP10, BH14a, BI13, BCS13b, BLL12, BL14, CM12a, FKM13, Fuj10, HP12b, ISYY11, KKB11, LLY11, LL13b, LLB13, Lim11b, Lim12, Lim14, MY14, Mol11, NV10, PP14, QS14, Yam13, vdW14]. definiteness [GR12, HS12c, SH10]. Definition [BPDC14, Qui11]. deflate [Mei13]. Deflating [TGM11]. deformations [DFS12, DFS14]. deformed [Han11a]. degenerate [FKR11b, FKR12, GK11, Mar13b]. degenerations [CKLO13]. degree [AW13b, BMSW10, BZ12b, CH14, Ere13, Hil13, HK13, LLS11, LL10e, LL11c, aLwW13, MAS12, SWT13, TW10a, Tan10a, YHY14]. degrees [DD10b, GM12]. Delannoy [lYnZpYH13]. delay [BCY12, Liu13, Mah11, RMT11, WZ13b]. delays [PKR12, Sha14a]. deleting [LSR11]. delta [CW11]. delta-monotone [CW11]. dendriform [BM13c]. Dennis [Bar10b, Sla10]. Dense [FHS14a, Duk12, Duk15]. densities [DGH^{+}11]. Density [MZ11, CCL14, KKL14, MRW11]. denumerable [MA10b]. dependence [Hil12b, Lu11, PS12, Wal11b]. dependent [jAS13, BRA11, DT10, GKS^{+}10, ZHZF13]. depending [DP12a]. depth [LFS12]. Derivable [Pan12a, JH10, QH10, ZZ11, ZZW10, ZZ10, ZZ12, Zho11]. derivation [AS12d, BD12a]. derivational [Pan12a]. Derivations [KLZ10, AKA13, Ben11, BS12b, BE12, BG14b, BS12e, DDF13a, DW12b, DD14, EN11, HW11, HZM10, JQ11, LL11a, LP10, LD12, LL11b, LJ10, Pet10, QH10, QCH11, QH13, SM12, WW13c, Wan14b, WX11, XW10b, XW12, YZ10, YZ12b, ZHQ14, WWD13]. derivative [HMS13, Ogu13, vBM13]. Derivatives [Jai11, San14]. derived [Bar12a, KAAK11, Vij14]. deriving [DMS13]. descent [JZZ13]. describe [LdlP11]. described [BFH^{+}12, IMA10]. Description [MMP13a, RRKK12, COvdD10, LM10b, Pol12]. descriptor [BV13a, HRT10, Jun14, Köh14, SBM11]. Design [BO12, Jim10, LLMZ12, Mah11]. designs [FMR12, Kla10, LHG10, Mit11b, NP10]. Detecting [Hen10, NV10]. Determinant [FMM13, MH13b, TT10, TW10b, ÁAFG12, CMS12a, CM12a, CH11, DHS12, DdC13, FMR12, KKL10, Vse12]. Determinantal [CM12b, CN12c, Ens10, GMT13, Abe14, BT11c, CEY14, CN11b, Dru13, GS10c, KLS12, Lin13, LSR11, NT12a, Qua10, Qua12, VS11, Wan14a, LPK14]. Determinants [GK10, SR13a, TT12, Bün14, CY11, CHZ13, DU14, EWY12, Kni14, Mat14b, RBP12, Sbu10, Sch10, SSZ13]. determination [MR14b, PR13b]. determine [DV14, RR11]. determined [ABEV10, BT14b, BG14b, BZ12b, BZW14, Gha13, HH13, JZ14, KKL13a, KKL13b, LSD14, Sto11, WFM11, WLLX11, WLG11, WS12, WLG12, WML13, WY14a]. determining [DGH^{+}10, XY11]. deterministic [NNW14]. Deveci [Hil14c]. development [LaG13]. deviation [GH11, BI12]. Devoted [BBD^{+}11]. Diagonal [PM10, BOZ11a, BOZ11b, BC12b, BC14b, BV13b, CFJKS13, Drn13b, EM12, Fri11, HZ12a, LS13c, Mol12, Pry10, Reh10, Reh11, Rub13]. Diagonalizability [KKLY14]. diagonalizable [BFK11, KB12, Lán10a, RdSP11]. Diagonalization [RE11, DFR13, Fut12, GB13]. diagonally [CHLS12, Far11b, FLH12, HZ10, LH10b, LHZH11, LZL12]. Diagonals [Lee13a, WW13a, CD13, – Dok12, Sem10, PRW11, YT13b]. diameter [CLL12, CvDKL10, HL12, HWG13, KSH12, LLS12, LL14c, WKV10, WL10b]. Diameters [LL13a]. Dias [FdC12, LPQdS10]. dibaricity [CLOR13]. dichromatic [LS11b]. Dickson [FFS11b, GN13]. dictionaries [Fou14]. Dieudonné [dlCdlRMP14, RAAGAVS11]. difference [Baj14, BM12b, BAD09, DLDV11, FLS10, LS14, MN12b, Pit11, SS11d, TX12, VS14b, Vul12, Zuo10]. differences [HKK^{+}12, RW10, ZZCW13]. different [BBM14, HRT13, Koz14b, LRT13, Tra13]. Differential [Lan10b, BM12b, BCF12, Lom11, NT11a, PLS14, Rue13, Tre11]. diffusion [KRS13]. digraph [ABS10, CK14, DK13d, GR10, HL10c, Kim10, KP12, KSH12, MS14b, MSvdD14]. digraphs [BB11a, BM14b, Bru10, BS13b, CGR14, DL13, GX12a, GL12c, HY14, Kal13b, Kim13b, LLH10, LS11b, LS12b, LYL13, Moh10, Rad10, Sch11, TC13, YW11]. dilation [Dum13a, GWW14a]. dilations [ADW13]. dimension [BCS13a, BCF12, Cau11, dHLMS13, KK14, Lop11a, MMS12, Mar10]. dimensional [BDF11, Ben14a, BS11c, BS13a, DK13a, DK13c, CK13a, CILL12, CdGS12, CdGS20, DK11, DK12a, DK13b, Dub14, FP14, GJ11, HN10, KRH14, LwCJL11, Ma11, Moj14, Sha13a, Sko11, SQ14, Suz13, Ter13, Ter14, Wój14a, WJ12]. dimensionality [Cha10]. dimensions [AJ13, CKL^{+}13, DKM^{+}14, GPT12, HH12b, Yan10b, YZ13]. dioids [BL11, HTS14]. Direct [BT13, BL14, CHZ13, GK14a, JS13b, Kak10, Kaw12, KSH12, Lee13a, PPKR12]. directed [BFRR14, BKP12, Bau12, BRZ11, KP14a, Kir10, Kir14, NSC13, RGC13, dFBRS14]. Dirichlet [PRT13, Sow13]. disc [LH10b]. Discrete [CEM14, FK13, HMR12, RY12b, AW10, BJRS11, BJ13, Cas13, Cha12, CMN10, Czo10, CN10d, CN11e, FV13b, HDPT12, HRT10, Hil12b, Jun14, LTX14, NRS12, OM10, RS14a, Sad12, VS14b, Wei13a]. discrete-time [BJ13, FV13b, HDPT12, HRT10, Jun14, Sad12]. discretization [PLS14, Vul12]. discs [CGWW13, FHM13, MH13a]. disguise [BCT14]. disjoint [BT11a, CW12b, DWXS12, LS13b, LHL14, LHGL14]. disks [Mel14, ST10b]. Displacement [ZYL10]. dissimilarity [WNM13]. dissipative [ADW13, DP12b, GO13, Lin13]. Distance [AH14, BNP11, Ili10a, NOL13, Psa12, AH13a, Ban13a, BS11a, BNP12, BNP13, BYZZ14, Cer10, CBB13, CLS13, DvDF11, DV14, DS14, Est12a, FGG10, Fio12, FS11, GH11, Gro14, HKP13, HRW99, HLW14, lH10a, Hua10, lH11b, JM12a, JZ14, KS12b, KY11, KPY11, KT10, KT12, KM13b, Lee13b, LwW14, LP12, Lim10, LC10, LHWS13, LZG14, LZ14, MAS12, NP12, NP14, NM14, SS10b, ST12, WZ13c, XZ13b, XZD14, Yu13, Zha12b, Zha14a, vDF14]. distance-biregular [Fio12]. distance-regular [Ban13a, Cer10, DvDF11, FGG10, Fio12, KY11, KPY11, Lee13b, LwW14, vDF14]. distances [BT14b, HN14, JS12]. distinct [BB13a, CGMJ14, CHLS12, FLH12, HTW13, KS13b, NS13]. Distributed [RvS13]. Distribution [PLS14, Ada14, ANF11, DGGJ11, HLP12, JPS13, MSS14, PS12, SCSS10]. distributions [Cha14, DGJ10, MW10, OM14, Sku13]. distributive [HW14a]. Dittert [CW12a]. divergence [CM12a]. divided [TX12]. Divisibility [TL13a, MM10a]. division [Bot14, KLZ14b, Liu14a, Mar10]. divisor [Bot14, TL13a, WMZ14]. divisors [CS10a, Ens10, FH10b, SDNS13]. do [Pro10, RR11, XG13]. Dodgson [Abe14]. does [BMSW11, LY11a, dSP12d, Sta12, WBHM13]. domain [Pol12]. domains [BCŠ10, BC12b, BC12c, Ens10, Hua12a, IW13, SZ14]. dominance [Mol12]. Dominant [Fie11a, BGV12, Far11b, HZ10, LH10b, LHZH11, LZL12]. dominated [Sei14]. dominating [CT11b]. domination [AHS10, Har14, LL14b, XF11, Zhu12a]. Double [OT12, AdFST11, BC14a, FG13a, God12, JL12, KSAM12, Lee13b, Mol12, PY10, WJ12, Zhu11b]. double-integrator [Zhu11b]. doubling [WCKL13]. Doubly [GS10a, AT14a, BAD09, DHS12, Fan10b, GS10b, GKR13, HP04, JP11, JLW11, JPS13, LXL^{+}14, LZL12, Mou12, MAM^{+}13, NS12c, PSW11, Sar14, Sha13a, Sha14b, XLG^{+}13]. doubly-infinite [Sha14b]. doubly-stochastic [DHS12]. Dragos [GRS^{+}10, GRS^{+}11]. Drazin [BZ13, CGMS10a, CGMS10b, CI13, DW11, DMS10, DMS13, HRT13, KCID12, MD10, Mos13, PH12, SH13, WC12, Wül13, XWS12, XSZ13, ZW12b, ZBW12, ZCC12]. Dual [CT14b, Wor13, CD10, DMMY10, GMH14b, HM14a, KAAK11, KKL13c, LH11a, LWGM10, LWGM12, LH10c, MGLW11, MSS14, RDD14, Tif11, BDK11]. Duality [BHMR12, MRS12, FP11, Pin11]. Ducci [HNZ12]. due [Seo14]. Duffin [ACG14]. Dunkl [ST10a]. dyadic [MZ12c]. Dynamic [Wil11]. dynamical [DZ12a, FGQ11, NP13a, VB10]. dynamics [JDY13, MSvdD14, Zhu11b]. Dyson [Yan10a].

e-ISBN [Tam12, Zha12a]. economics [TP13]. ed [Tam12]. Edge [CW12b, LS13b, LLT13, LHL14, BYZZ14, CL14a, Cio10a, Cio10b, EHH^{+}12, LL13c, LHGL14, SS12a, SSGL10, Sto11, TF10, WAH13]. edge-based [WAH13]. Edge-connectivity [LLT13, Cio10a, Cio10b]. Edge-disjoint [CW12b, LS13b, LHL14, LHGL14]. edges [Suz13, Zhu11a]. Edition [Bar10b, Zha12a]. Editor [Ano11z, Bru11, Bru13, Bru14]. Editor-in-Chief [Bru13, Bru14, Bru11]. Editorial [Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h]. Editors [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x]. Edmonds [KW13]. effect [Mit11b, XZD14]. Effective [ESV^{+}11]. effects [Dug11]. Efficient [BHAP12, CFPP13]. Ehrenpreis’ [Lom13]. Ehrlich [BN13, GN13]. eigendata [GP12]. Eigenfunctions [WAH13]. eigenparameter [jAS13]. eigenparameter-dependent [jAS13]. eigenproblem [GPT14]. eigenproblems [AA11, GHT11, MSS12, WZ14b]. eigenspaces [Ada14, GTV12]. eigenstructure [LdSP11]. Eigenvalue [FZ13, HKK^{+}12, JTT13, MMA11, MMRR11, Row14a, YZ12a, Zha14b, AAK11, AKM14, AHL11, BCY12, BV11, BM14a, BdFdP11, BFdP12, Bel14, Bey12, BBE^{+}10, BN13, BdS10, CW10, CQYY13, Das10b, Das11, DS11, Duk12, Duk15, Elo11, FF12, FZW11, Fan10b, FHL^{+}11, Gar13, GP12, GCY14, GM14, HHL10, HRW99, HMP12, HP04, JM12a, Jar12, JDY13, Kal13b, KP14a, KMNS12, Kol13, KZ11, KPY11, LaG12, LXL^{+}14, LCwCL11, LY11a, LW12b, LQY13, LL14a, LL10d, LHW11, LLT13, LJY14, MD13, Mou12, MAM^{+}13, MP10, MP14a, Nak10, NSW13, NS12a, NM14, NP13b, NY14, OS14, yPjXL11, PAS11a, PAS11b, Row11, SVP11, SS14, SCSS10, ŠŠ11e, Sta12, TF10, TG13, WF10, WF12, WZ13b, Wei13b, Wei13a, WJ12, XD12, XZ13c, XLG^{+}13, XE11, YZ11, Yu13, YWX13, Zhu12a, dLOdAN11, dLN13]. Eigenvalues [Alo14, CMRR13, Cio10a, Cio10b, JS13a, LLMZ12, LWY10, MR12, Moh10, MAS12, RW12, AFHP14, AGM14, AAT12a, AOTR13, BFdP11, BZZ14a, BK12, Bün14, CSZ10, CS13a, CPZ13, CFJKS13, CFK^{+}10b, CGW14, CH14, CGMJ14, CHLS12, CJ14, CW12b, DXL13, DLS14, DP12a, – Dok12, Dom13, DZ12c, DGU14, FLH12, FHRT11, GKZ11, HMTR10, HO11, Har14, HS14a, HS14b, HQS13, JPS13, KS13b, KY11, KT12, Kus13, LSC10, tLyLWqW10, LNTgW12, LS13b, LGS13, LZG14, LZ14, Lip10, LHL14, LHGL14, MVPS10, MV12, NS13, NR10, Ogu13, dSP14, Psa12, QSW14, Qui11, RR11, Roj11, Row10, SS10b, SYH14, Sra13, Ste11, TH10, Wal14, WB11, WBWH13, XC13, XZ13a, YT13b, YY14a, ZLW12, ZLL12, ZC14, ZCWZ13, ZSWB14, Zhu10b, Zhu11a, ZJ10, vDO11, JT11a, RJ11]. eigenvector [GK14b, GLS13a, MM12, RW10]. Eigenvectors [BBGM12, Bün14, BM13d, SLS13, Sri13, XZ13a, FK13]. eight [NP10]. element [Drn13b, Ney11]. element-by-element [Ney11]. Elementary [BJ10b, Buc10, DDK14, Gu14, PY10, Sin10a, BJ10c, CP12, CS10a, CDDY10, FH10b, Kec13, Kuz10, Per13, Rud12, Ruk14, SZ14, SDNS13, Yam13]. Eléments [Cas10, BR11, BPDC14, CGMS10b, Che14a, CK13c, DDGH13, LL11b, LHL12, Lim10, MS11b, WLG11, WLG12, WZ13a, Zho11, Cas10]. eleven [BSU14]. elimination [VS14a, HZ14, Ji12b]. elliptic [CEM14, DKOT12, KK12, KK13, SS11a]. elliptical [ATS12]. embedded [Mit11b]. embedding [HH12a, Wag11]. Embeddings [Pan12b, RV13]. emphasis [KN10]. enclosure [Miy13, Miy14]. enclosures [FHS14b]. encounters [RL13a]. Encyclopedia [Grü12]. Endomorphism [PYZ14, ABK14]. Endomorphisms [DGZ13, HHLS14, Mol13, OZ10, Rom14, SR13a]. energies [LS12a, XF11, Zhu12b, Zhu12c]. Energy [GRM^{+}10, KAMS11, LW12a, ACG^{+}11, ABS10, AW11, CGTR14, CFK10a, CLL13a, CLL13b, CL13a, CGR14, DXG12, DG14, DGÇ14, DLL11, FHRT14, GKZ11, GX12a, GHW^{+}13, GS11b, GMRS14, GFB14, HJL10, HLS11, HJLS11, Ili10a, LLS10, LLS11, Rad10, RGC13, RJ10, RJ11, Roj11, RL13b, SS11b, SSGL10, SW13, SRdAG10, Tia11a, TC13, VDVJT13, WL12, WZY14, YM12, ZZ13, Zha14a, ZK14]. Eneström [RL13a]. enforcement [BV13a]. Engel [Bie13]. ensembles [DGGJ13]. entangled [Fid10]. Entries [GG13, BCŠ10, Bün14, CFJKS13, CDDY10, FFJM14, NT11b, RR14, Shp10]. entropies [Ben14b, BP13, EdlPH14, IIK^{+}13]. entropy [CFPP13, CH11, DGZ13, DLLS10, EdlP14, Kim13d, MS10c, PP11b]. entropy-preserving [PP11b]. Enumerating [IJ12]. Enumeration [AM14, Car13, SS13c, GL10b]. EP [BR11, TW11a]. Equal [FFJM14, HL11c, HTW13]. Equalities [Tia10, Pop12]. Equality [Ber13a, CPH11, JM14]. equation [Bou10a, Bou11, CGGS13, DD11a, DDG^{+}13, Fuh10a, Fur10b, GS13a, GKL11, HN10, Hwa12, Kaw13, LwCJL11, Lim11b, MR10b, MR14c, Per14, SK14, ZLD11]. Equations [Ano12a, PQ10, AiS13, Ara12, AG10, Baj14, BK11a, Bie13, BM12b, BJ13, Car13, CN11e, DH10a, DH12a, DKS13b, DLDV11, Dum13a, FMWW12, FGR13, FH12c, GMH14b, GL10c, Guo13, Hla13, IW13, Jbi10, Ji12a, Jim10, JK13, KL13a, Kyr13, LV11, LV12, LLW14, LS14, LTX14, Miy13, MSP13, Mys12, PQ12, PLS14, Pit11, Pol13, Rei11b, RY12b, Roh11, Sad12, SS11d, Tre11, VS14b, WW10, WD10, WCKL13, XX14, XSZ13, Zha12d]. Equiangular [DHS10, Sin10b, Bod13, FMT12, HS12a, Szö13]. Equilibrants [JT11b]. equipped [Dor10]. equitable [Ter13]. Equivalence [Sze14a, tHR13, GPT12, GTW13, LV12, dSP10a, DDM14, Tim14]. equivalences [GHS13]. Equivalent [Hu10, CHLS12, FLH12, LLMZ12, LMT10, OM12, Van10, ZLL12]. erasure [FM12]. erasure-robust [FM12]. erasures [HS12a, LH11a, LHH13, LH10c]. Erdos [OLW14]. ergodic [Fid10, Puł11]. ergodicity [Kir14]. Erratum [Cio10a, EFN10, KKL13a, Vla12]. Error [Dai11, GEP13, GO13, Cha12, Cha10, FH12c, GEP10, GJTP13, Hua13a, HZ14, KOR14, LHG10, PLS14]. error-correcting [LHG10]. errors [AA11, CLCL12]. Essentially [ZCKS12]. estimate [KOR14]. Estimates [BR14d, Gar13, BW12, Duk12, Duk15, MM13]. Estimation [HZ10, ZJ10, BL11, GD11b, NNW14, PT13b, RvS13]. estimations [Ara12, EY13]. estimator [HLP12, LJ12]. estimators [MD12c]. Estrada [DZ11, DL11, DZ12b, DZX12, FTA11, WX14, ZZL11]. Euclidean [Bai14, Bud11, GS10c, GTR12, JV10, KT10, KT12, KM13b, LC10, Mol12, Tao11, Tao13, TKLX14]. evaluated [BW13b]. evaluating [CHZ13, KMNS12]. Evaluation [IRT14, CMS12a, Vse12]. even [ABS14, BCF12, Hil13, KS11c, OS14, PPK13, QS14, RR11, Rei11b]. eventually [MPT14]. Every [CHLS12, BFPSS12, BH13b, FLH12, GS12c, KLZ14b, TZ13b]. evil [Vse12]. Evolution [DT10, CLR11, CLOR13, LLR14, RM14]. evolutionary [CK13b]. Exact [Ber13b, Cha14, KM14]. exactly [TH10, WBWH13]. exceed [BMSW11, BZ12b, LY11a, Sta12, WBHM13, XG13]. exceptional [MW14a]. excess [FGG10, vDF14]. Exchange [JS13a]. excluded [OM12]. Existence [Fan10a, GHMPVP11, GK11, MM12, KD12]. expansion [Ma10a, WZ14b]. expansions [Moo11]. Expected [HHMS10, Kir10]. Explicit [Dod10, DS12c, Kyr13, Mir12, Rim12, Sra13, XWS12, GMV11, Koz10, RW10]. exploration [DPF10]. exponent [GKR13, GWW14b, JLW11, KSH12, PJ13]. exponential [CMN10, Fis14, Ika11, Moo11, PKR12, PT13b, VS10]. exponentials [Sha14b]. exponents [CMN10, Czo10, LT10b]. Expressing [Slo13]. expression [HZGY12, Rim12, Tia11b]. Expressions [CGMS10b, CBB13, HZ11a, TZ12a]. extendability [BB11a]. Extended [CD14, MAGR13, SN14, JR11, Kum11, Mar13a, MTS11, Pat12a, TZ12a]. Extended-valued [SN14]. Extending [Hua12a, CJ11]. Extension [Lee11, BP14, CFPP13, CH12, GH11, JMP10, Kim11b, LNT13, Ma10a, MN12a, dSP10a, tHR13]. Extensions [Fur12, YJ12, BdlC13, BdlC14, DGGJ11, FHS14a, FGH13, FG13b, Had12, JZ11, JL12, LCM13, MMA11, RS14a, SS11d, ŠK10, WGL12]. extent [dSP11b]. exterior [Ika11]. extraction [BK11b, Dur12]. extrapolated [BH14a, BJRS11]. Extremal [DGKO13, DL13, Jen13, LS12a, MZ13, MW14b, XF11, Ziv12, AHL11, BK12, CGSCZ10, DC14, DdF14, LL14d]. extremality [Ha13]. extreme [BT14a, CH14, Hil12a, LL10d, YX13]. extremum [Dax10, Tia12].

Faber [CKS10]. Faces [CM10b, PSW11]. factor [LNN14, LLF13, Shi12c, YZ12b]. Factored [Cha13a]. Factoring [DT11]. Factorization [CW11, KM11, Lak10b, Bot10b, BH13a, CRU13, CRU14, Dur12, For14, GS12d, HLZP13, Hua13a, Hua13b, LYS13, MR11, SPKS12, vdW14]. Factorizations [BBG14, KKL14, Lim12, HL11c, Hua12b, KG12b, LW13b]. Factorizing [Sow13]. factors [Jar12, Koz14b, LWY10, Sbu10, SC13]. Facts [Ber09, Sla10]. faithful [LW12d]. Fallat [Gar12]. Families [BB13d, FP13, HKPR13, KB12, Nat13, RSS10, Stu13, Wód14, Ziv12, dFdADVJ10]. family [Baj14, Cha14, GZ12, Kal13a, LX13, TH11, WJT13]. Fan [BFdP10, GMRS14, tLyLWqW10, Lin11, SRdAG10, ZCWZ13]. Fast [BSS10, HR11, Miy13, OM10, XE11, BBE^{+}10, Han13a, MS11a]. Fastest [Kir10]. fault [PP13a]. feasibility [LLW14, ZY12b]. Feedback [Jun14, SSS13, BO12, CC14, Liu13]. Ferrers [AdF11]. few [AJRT14, BG12b, FMNW14, Kla10, Mol11, dSP14]. Fibonacci [CLX13, DGMS14, LaG13]. fidelity [Kim13d]. Fiedler [BOZ11b, Grü12, Ano13y, BHMO13, BOZ11a, BCF14, BF14c, DDM12, Mac13, NdM13, Nik13, Ser13, Stu13, DDP13, DDP14]. field [AKM13, AW13c, Bie13, Bot10b, Bot12, CK13c, CJK^{+}13, EJLS11, HK13, dHLMS13, HCY10a, LHG10, LdSP11, Ma11, MMP13a, dSP10a, dSP10c, dSP10d, dSP12c, Qui11, Rad13, SS10d, Wu10b, de 13]. fields [AG12, BB13a, BBC^{+}14, BC13, FL12, GH13a, GH13b, HHLS14, KK14, Kra13, KS11b, LdS13, LT13, MZ12c, MSP13, PP13a, Per13, RS14b, SV13]. filiform [CT14a, CGO10, CLOK13, LMO16, Wu13b, CCGVO13, CCGO14]. filling [HK13]. fillings [AW10]. filter [BOZ10, Jim10, LLMZ12, Mah11, MZ12c]. Filters [GLP^{+}13]. Filtrations [Stu12]. find [CS13b, Ern13]. Finding [AS12b, BDD13, PQZC11]. Finite [Ter13, AiS13, AG12, Ada14, jASZ12, BB13a, Bar10a, Bar13a, BS11c, BS13a, BPDC14, CK13c, CKL^{+}13, DHLX12, Dai13, DKM^{+}14, Dub14, DDK14, FL12, FG13c, GH13a, GH13b, GW14a, GWW14a, GJ11, HW14a, HM14a, HS12b, HK13, HN10, HHLS14, KKS12, LHG10, LdSP11, LdS13, LX13, MB13, MSP13, Moj14, NS11a, Pol12, PRW11, Rom14, Row14c, Row14d, SR13a, Sev10, Sev13, Sha14b, Shi12a, Sim10, Sko11, SQ14, ŠŠ11e, Ter14, VW10a, Vij14, Vul12, Wój14a, Wol12, Xu14]. finite-difference [Vul12]. Finite-dimensional [Ter13, BS11c, BS13a, Dub14, HN10, Moj14, Sko11, Ter14, Wój14a]. finite-step [DHLX12]. finitely [Bar12a, Ma11, TL13a, Yan10b]. finitely-generated [Bar12a]. Finiteness [CGSCZ10, Dai13, Mor10]. Finsler [Vla12, Bal12a, Bal12b, TNP12]. First [DD10a, LJY14, LZ11b, AM14, BM10, Bos11, JMS11, NS11b, Zhu12c]. first-order [BM10]. Fischer [Dru13, Lin13]. Fisher [KS10, KS12b, KM14]. Fissioned [MW12a]. fitting [MM10c, Moh13]. five [BOZ11a, BOZ11b, NS12a, Ste11, XX14]. five-diagonal [BOZ11a, BOZ11b]. five-point [XX14]. Fixed [Ros12b, SS11c, WLL14, CLL12, Cos14, HL11a, HL12, LT11a, LT10a, LL11c, NOL13, Per14, SBMT10, Slo14, YFW10]. fixed-point [Per14]. Fixing [Lip10]. flag [Ber13a, DKM^{+}14]. Flanders [DLNN14]. flat [CSAC10, CSAC11]. flats [WLGM11]. Fliess [LZ11b]. flipping [MSP13]. flow [BOZ10, LL10d]. flows [AKZ13, AJ13]. forbidden [LS13d, Yua14, Yua15]. forcing [BBF^{+}10, CL14a, EHH^{+}12, EEH^{+}13, GB14, HCY10b, Mey12, Row12a]. form [BBdH13, BFdP12, BIT12, BDK11, CRU14, CT11a, FDS13, FP11, FHS11, GS12b, GK11, Mar13b, NSC13, Nom14, Rad13, Reh10, TX12]. formal [TZ13a]. format [BGK13]. formats [KRS13]. forms [AW13a, AIS14, BDK11, CGMS10a, CN13a, DFS12, DS10, EN11, FV13a, FGH13, GMMFPSS12, HSZ12, Hu10, JR14, KKLY14, KLZ14a, KB12, MMMM13, MPS10, NT11a, Pop10, Reh11, Sim10, Sze14a, TW14, Wil14]. formula [AS12d, Dai12, Gro14, KLS12, Kon13, Koz14a, KR10, Ma10a, MS11a, OM14, RW10, SR13c, TX12]. Formulae [BT14a]. Formulas [Ber09, Tia12, BZ13, DK13e, Ern13, GS10c, GMV11, GLW13, GL14, Kyr13, MGLW11, MAGR13, Rue13, Sla10]. four [Ben14a, BW13b, Mar10, RRM11, Tia12]. four-term [Tia12]. Fourier [Ban13b, FK13, Sav12, Xu14]. Fourth [HSZ12, jASZ12, DHS10, Lop11b]. fraction [Dur12, KY13]. fraction-free [Dur12].

Fractional [Beh13, LdS13, Mat13a, Mat14a, GMH14a]. Frame [Fut12, BHKL10, DFR13, VW10a]. framelet [MZ12c]. Frames [Wal11a, Bod13, BW13a, CKL^{+}13, DHS10, FM12, FMT12, HM14a, HG11a, Hir10, HS12a, KKL13c, KOPT13, LH11a, LHH13, LH10c, NRR^{+}11, Sin10b, Szö13]. framework [GHT11]. frameworks [AY13, AN13, Bai14]. Fréchet [San14]. Fredholm [Sei14]. free [BGP13, BP10, BdlC14, BM13c, DDF13a, Dea11, DTL11, Dur12, GS12d, KY13, Ma11, MR10b, MR14c, VB10, YW11, ZW12a]. French [Cas10]. frequencies [Koz14b]. frequency [ST12]. Frobenius [Lim13b, BT14b, CVW10, DIP13, FGH13, LN12, MZ10, NV12, Pat12a, Pit11]. Fuchsian [Kim13c]. Fuci’k [RY12b]. Fuglede [DK14]. Full [CRU14, HSS14, JP11, LMW12, BC13, GLZ14, JKV13, Sav12]. full-rank [Sav12]. function [BKV14, BV11, Boj13, BTYZ12, Car10a, CM12a, FDS13, FdC10, FHS14b, GIP12, GS12f, HLP12, LQY13, MS14b, RS14c, SS12a, SS13a, SS13c, SS13f, Tia12]. Functional [BK11a, Blu10, Ere13, BB13a, Fuh10a, Rhe10, TX12]. functionals [Ben10, GD11a, Lu11, SS12c]. Functions [Xu11, AL13b, Aud12, Aud13a, BKV14, BT11b, BGP13, BHDW12, BANP12, BB10, BEK13, BB14, BJ10a, CH11, CGMPSS14, Dra12a, FP11, FG13b, FKR11a, GHMPVP11, GKS^{+}10, GK14c, GG13, Hia13, HWH14, HH11b, IM11, JM14, JRMFSS12, JZT14, KOR14, Kia14, KKL10, MM10a, Naj13, Nak13, Nat13, OT12, PCC12, Ruk14, Sag13, San14, Sin10a, SN12, SN14, Sto11, Uch10, WLL14, Wód14, Xu14, Zha14c]. fundamental [LNN14, Lom13]. Further [DMS10, KMSC14, MH13a, PAS11a]. Furuta [BLdP12, FNY13, Wad14]. Furuta-type [Wad14]. fusion [Bod13, HM14a]. Fuzhen [Tam12]. fuzzy [MMP13b, MP13b, MS14c].

G [Loe12, Rei11a, Rhe10]. Gaddum [NY14]. gain [KG12a, Ref12]. Gallai [KW13]. game [wH13]. games [AGK11, FDS13, Jun14, Sta14, Cec10]. Gangster [FJ10a]. gap [GO13, KR10, Wol12]. gates [CLX13]. Gau [CRSS14]. Gauss [BFRR14, Ji12b]. Gaussian [ALPV14, BT14a]. gcd [KAMS11, TL13a]. gcd-graphs [KAMS11]. GCDs [BDD13]. Geary [BBGM12]. Gel’fand [Dai12]. Gel’fand-type [Dai12]. General [Jim10, AY13, BV12b, Bot10b, BGH12, CPH11, CFK10a, CMS12a, CM12b, DH10a, FFG^{+}11, JKN14, KSAM12, LHLL10, MR11, Sha13b, TT11, XW11, wXL14, wXZ19, vBM13, Beh13]. generalised [DGJ10]. Generalization [– Dok12, Mel13, AKZ13, BL13b, FFS11b, Kus12, MGLW11, MW12b]. Generalizations [FFS11a, IIK^{+}13, MS12, SN12, WX11]. Generalized [AKN12, BCEM10, Ben11, BG14b, CL12a, Dra14, Dur12, GP12, HL10c, Hun14, Hür13, Ili10b, KB14a, KB14b, KSA11, Kau12, Kim10, Kim13a, KCID12, KS11c, LL11b, MP13a, OM13, SS14, Seo13, SM10a, TX12, Wu10b, AL13b, ÁvW11, AM13d, AL13c, Ban13b, CMRR13, CJ10, CLCL12, CL11b, DH10a, DH12a, DV10, Den10, DK12b, DYW14, Dra12b, DW12b, DW13, DBZZ14, FKW13, FTZ12, FDS13, FdC10, FH12b, FH13, GP14, GL12c, He11, HOT13, HLW10b, Hu10, HZ11a, Hua11a, HP11, Hwa11, IW13, JM14, Jen13, JDY13, Kim11a, KP12, KS14b, KKM13, LW12a, Lav10, LCwCL11, LW12c, LW13a, LYS13, LMW12, LJ12, LJY13, MB13, Mar11, Mar13c, Mat12, Mel14, MS11b, Mos13, Nak10, NCI13, Pep11, Pep12, RAAGAVS11, RJ11, San10, SS13a, Siv13, SR12, SS10e, Ste13, SL14, TZ12b, TNP12, TN14, Uhl13]. generalized [WLLX11, WML13, WW13c, XW10a, XE11, Yan10a, YW10, YZ11, ZY12a, ZZC14, ZCC12]. Generalizing [CC10]. generate [AKM13, Cha13a, Kau12]. generated [Bar12a, BS11c, BS13a, GS10d, KM12, Ma11, Rez13, Yan10b]. Generating [CK13b, CN10b, Lop11a]. Generation [BFBD13, JSS13, RRZ13, Wik11, Wik12]. generators [pWlW14, WY14b]. Generic [Bat14, KBS13, Fri12, HS10, MMRR11, MMRR12]. genetic [Pan11]. genetics [LLR14]. gentle [BS10]. geodesic [Fuj11, HN14]. Geometric [Ban13a, CvDKP13, ISYY11, AGV12, BG12a, BCS13b, DMMY10, GG12, Han14a, JLLY10, KLL11, Kim13d, LLY11, Lim11a, Lim12, Seo13, Wal14, Yam13, BT11c, DD10c]. geometrical [ART13, Aud13b, BDD13, Ben14a, NS12b]. geometrically [Aud13a]. Geometry [DHC12, Fie11b, HH14, Lim13a, BH10, BCF12, Fuj10, GH13a, lH10a, Hua10, lH11b, Hua12a, HS12c, UV13, Grü12]. Germany [BFBD13]. Geronimus [DGAM14, DW10]. Gersgorin [CGWW13, MH13a, FHM13]. Gerstenhaber [de 13]. GF [AW13c]. girth [Kol13, Row11]. given [AG12, AW13b, BL12, CK13b, DL13, DZX12, FYI10, GL10b, GM12, GLS12, Hua11b, HJL10, Kir14, LS10, LZ12b, LY11b, LS11b, LJY14, MS13c, dSP11a, ŠK10, SWT13, Tan10a, WKV10, XF11, YW11, YWS11b, ZHG13, Zhu12a, Zhu12b]. global [CD11, Kra13]. GMRES [TMSS14]. GNS [Kaw13]. Gohberg [BDK^{+}10]. Goldberg [Ano13z]. Golub [Rei11a]. Gondran [GS12f, Shi11]. Good [Hua10, FG13a, NT14]. Graded [DKS10, ACGVK14, CGO10, CLOK13, Cen11, CM14, KZ10, LMO16, Mar13a, Mar14b, Per13]. gradient [JRMFSS12, LL10d, OM10]. grading [Mar14a]. Gradings [Cir13, MFGD14, Per13]. graft [XZD14]. grafting [SSGL10]. Gram [SST14, Wil13]. Gramian [HG11a]. grand [FNY13]. Graph [CS11, ACM^{+}12, AAJ12, AV12, Alt13, AH10, AH13a, AHLvdH13, AOTR13, BB13c, Bap13b, BBF^{+}12, Bau12, BOZ10, BRZ11, Boz13, BZZ14a, BW12, CMRR13, CAV13, Cer10, CTW11, CW10, CS13b, CLS13, CL13a, DGH^{+}10, Dea11, DC13, EHH^{+}12, EFN09, EFN10, ESV^{+}11, FTDA10, FH13, FL12, GX12b, GX12c, GCY14, HHMS10, HS14a, HS14b, HS10, HCY10b, HCY10a, IMA10, KR12, Kir10, Kir14, KPY11, Kuz10, LaG13, LRST10, LWV12, LFS12, LW12e, aLwW13, MYL13, MŠ10a, Mit11a, MS13c, NSW13, Nik13, OdLdAK10, PRT13, RJH11, RR12, RJ11, Roj11, Row12a, Row11, SS12a, SS13a, SS14, SM13, SRdAG10, Suz13, TW10a, TF10, Ter11, WF10, WL12, WAH13, WMZ14, Yu13, ZC14, Zhu10b, Zhu12a, Zim13]. graph-theoretic [BB13c]. Graphical [COvdD10]. Graphs [BMSW11, BL12, BCF12, CT10, DV14, Fie11b, Grü12, HTW13, HHLS14, LLS12, OS14, WKV10, WLLX11, WX14, vDO11, AO14, AFHP14, AFLN12, AJRT14, AKN12, AKZ13, AKM14, AS12c, AdFST11, AdFM11, AW11, ABS14, AH14, AT14b, AJ13, BB13a, Ban13a, BKP12, BG14a, BFK11, BHvdH11, BHMO13, BP10, Bau12, BB11a, BMSW10, Bel14, Ben14b, BFF^{+}11, BNP11, BNP13, BZ12a, BYZZ14, CvDKP13, CFG^{+}14, CR10c, CFK10a, CCF^{+}12, CM11a, CT11b, CHY11, CHY12, CFHL14, CFK^{+}10b, CLL13a, CLL13b, CGW14, CH14, CGMJ14, CL11a, CLL12, CG11, CKST11, Cio10a, CvDKL10, Cio10b, CW12b, CV13, CSAC10, CSAC11, CTG13, CS10b, DFG10, DvDF11, Das10a, Das10b, Das11, Das13, DXL13, DG14, DLS14, DC14, DO11a, DO11b, DLLS10, DL11, DLL11, DZ12b, DZX12, DZ12c, DP12b, DJ14, Est12a, EdlPH14, FF12, FWW13, FS14a, FTA11, FYI10, FN10]. graphs [FGG10, Fio12, GL10a, GZH14, GLS10, GL12a, GKZ11, GB11, GFY10, GHW^{+}13, GM12, GLS13a, GW11, GWZ13, Gu13, Guo10b, GSL11, GLS12, GLS13b, GW13b, GM14, GRM^{+}10, GS11b, HMTR10, HL10a, HL12, HLW14, HW14c, Hua10, HJL10, HLS11, HJLS11, Ili10a, JTT13, JZ14, JS12, KG12a, KP14a, KAMS11, KKL10, KS13b, Kol13, KS14c, KY11, KM12, LL13a, LW12a, LS12a, Lee13b, LwW14, LS10, LLS10, LWGM10, LSC11, LWZ11, LY11a, LWGM12, LS13b, LWG13, LTS13, LGS13, LHWS13, LZG14, LZ14, LL10e, LHW11, LL11c, LMW12, LHWL12, LZ13, LLT13, LH13, LHL13, LL14b, LL14c, LJY14, LSD14, LLT12, MWZ13, MS13a, MM10a, MR12, Mig13, MW14c, MK12, MAS12, NP12, NP14, Nik10b, Nik11, NY13, NSC13, NLL13, NS12b, Pan12b, PAS11a, PHS13, QY12, RRM11, Ref12, RM10, RL13b, RTR10, Row12b, Row14b, Row14a, Row14c, Row14d]. graphs [RW10, SS11b, SST14, SS13c, SSGL10, SW13, SBMT10, Sin10c, SK12, Sta12, Ste10, Ste11, Sto11, wTlW13, TL13b, TH10, Tif11, UZ14, Vij14, VDVJT13, WB11, WF12, WS12, WB12, WML13, WZ13c, WBWH13, WBHM13, WZY14, WY14a, WF14, WNM13, Wil14, WZL13, Wor13, XZ13b, XZ13a, XM11, XG13, YFW10, YL10, YWS11a, YWS11b, YFW13, YZF14, ZW12a, ZLW12, Zha12b, ZZ13, ZHG13, Zha14a, ZK14, ZWL13, Zhu10a, Zhu11a, Zhu12c, dFdADVJ10, dFBRS14, vDF14, vdH14]. Grassmann [BBCC13, Cen11, DKS10, Pan12b]. Grassmannians [De 11, Lim10]. gravity [XWD13]. Greedy [GNE^{+}14, GO12, Dum13b]. Greedy-like [GNE^{+}14]. green [Sev14, BTYZ12, CEM14, Mar11, Mar13c, OSZ10]. Greub [Rua12]. Gröbner [KS11a]. ground [dSP10a]. Group [RDD14, AAH13, AV12, BK11a, Bie13, CGRVC13, COvdD10, CLST14, CJL13, CK13c, CGM^{+}10, Den11, GST13, GH12, HRT13, Hil14a, HT10b, JLN13, KM12, LWGM10, LWGM12, LMW12, Ma11, MR11, Nie12, Nie13, dSP12d, Slo12a, TT12, VW10a, WZ13a, WMZ14, XSH14]. groups [Ada14, ÁvW11, BDV12, CD11, DF14, DK12b, DGZ13, FG13c, GL10a, He11, HS12b, HK13, Kim13c, KK14, KN10, KN13b, MW12a, Mol13, Rod11, Shi12a, Sin10b, Xu14, Yan10b]. growth [NA13, TZ12a, XW11]. Grüss [Coh14, GD11a, MR10c]. guessing [CFHL14]. guide [BS10]. gyroscopic [Lan13].

H [Rei11a, HQS13]. H-eigenvalues

[HQS13]. Haar [AK11]. Haar-type [AK11]. Hadamard [Aud10a, BFK11, BB13d, BR14a, CFLW13, Dra12a, EGR12, GK12, Hua11c, Kar11a, Kar11b, tLyLWqW10, Lin14b, Mit11b, NS12c, Pep12, Szö13, WZ13d, ZCWZ13]. Hadamard-type [BR14a]. Halley [Guo10a, Lin10]. Halmos [Dom10]. Hamilton [FTZ12, Hwa11, Hwa12, MSvW12]. Hamiltonian [BFS11, Lim14, LLT12, RS14a, SB12]. Hamiltonians [Han11a]. Hamiltonicity [FN10, Zho10]. Hamming [CM11a]. Handelman [Laf12]. Hankel [AHAPP10, Ang13, BH12, Boj13, BF14b, CY11, CMS12a, CHZ13, CK13d, CI10, EWY12, HR11, KS14b, MS14c, RBP12]. Hardback [Lim13b, Rod12a, Zha12a]. Hardy [AP14, Lac13]. Harm [Ano13d, Ano13e]. harmonic [Hir10, KLL11]. having [AG12, MR12, Nom14]. Haynsworth [GS10c]. heat [DGU14]. heat-bath [DGU14]. Hecke [Lee13b]. Heinz [BLdP12, KMSC14, MN12a]. Heisenberg [GS12e, MFGD14, Sze14b]. Helmert [Hür13]. hemispaces [KNS14]. Herglotz [BKV14]. Hermite [DTS11, Dra12a, GMV11, SV13]. Hermitean [LWGM10]. Hermitian [Rod12a, AAT12a, ALRV12, BH14a, BW11, BFdP10, BLdPS10, BH10, BLL13, CM12a, CGMJ14, CN11d, DP12a, Dom13, DYW14, FMWW12, FG13b, FS14b, Gar13, GG13, HFS13, HSS10, HSS14, JS13a, KM13a, LT12a, LNTgW12, SCS11, SH10, Sha14b, SM10a, Tia10, Tia11b, TCL11, VW10b, WW10, WZ14b, ZL11]. Hessenberg [BBGM12, God10, God12, HG11b, Ste13, TT11]. hexagonal [RGC13]. Hiai [Fuj10]. hierarchical [BGK13, BM13a, BPDC14, UV13]. hierarchy [GMS13]. high [BSU14, MD13, Nem13, WZ13b, ZY14]. high-order [Nem13, ZY14]. Higham [Dru13]. Higher [GW13a, LJY13, WX11, CN11a, CN12b, CLS10, HH12b, SWBS13, ST13, XW10b, ZZ11, ZZ12]. Higher-order [LJY13]. Higher-rank [GW13a]. highly [BW13a]. Hilbert [And13, AR12, AF12, BV12a, DWXS12, Dra12a, DK14, Fan12, Fie10, GN14, HKK^{+}12, HZ11a, KLP12, MMM13, Pop12, SST14, Sha11, SQ14, Tim14, WW10, WD10, tHR13, vBM13]. Hilbertian [AS12a, Gon11]. Hille [NA13, SR13c]. hitting [PR10]. Hodge [JR14]. Hoffman [BMSW11]. Hölder [FLS10, KSA11, Pat12a]. Holographic [CLX13, LMN13]. holomorphic [HWH14]. Hom [LCM13]. homeomorphism [MZ12a]. Homogeneous [EN11, PQ10, CFL13b, Tre11]. Homomorphisms [Mar10, DWW14, DW14, MW10, WW13b]. Homotopic [BI12]. Homotopy [RO10]. honor [Ano13y, GRS^{+}10, LPQdS10, Stu13]. Hopf [BH13a]. Horn [CP12]. house [BS12c]. Householder [Irv12, MPT12, dlRMP12]. Hua [Hua12a]. hubs [BEK13]. Hugo [Rod12a]. hull [Rez13, Roh11]. hulls [AM13a]. Hurwitz [BT11b, BG12b, FF10, KKB11]. hybrid [BCY12, GOSV12, Wan11b]. Hyperbolic [CN13a, GP13b, Kim13c, KK14, LYS13, NV10, Šeg10, Seg14]. hypercube [Sri13]. hypercubes [Tia11a]. hyperdeterminant [BBS12a]. hyperelliptic [OZ10]. Hyperexpansive [JJKS11]. hypergraph [AGK11]. hypergraphic [RR12]. hypergraphs [BZW14, CD12b, Duk12, Duk15, HQS13, Nik14, QSW14, XC13]. hyperinvariant [AW13c, MMP13a]. hyperplanes [De 11, LM10a, Nit10]. Hyperreflexivity [BRZ13, BR14d, PP12a]. hyperspectral [GP13a]. hypersurface [CN10b]. Hyponormal [HMT10, CDDY10]. hyponormality [KL12].

ideals [BZWL13, BDF11, BJ11, CV13, KS14a, KLZ10]. Idempotent [Ma14, WLG11, And13, Bot14, BH11a, BHMR12, KN13c, KLZ14b, LV11, PP11a, dSP10c, dSP10d, SZ14, Shi11, Sin10a, SB11, Zuo10]. Idempotents [Mat13b, ABEV10, BS12b, BS11c, BS13a, CGM12, Den11, DCIW12b, KCID12, dSP10b, SH13, Wan14b, ZW12b]. Identical [AGV12]. identifiability [BBCC13]. identified [AK11]. Identities [BS11a, KZ10, LMMS12, Abe14, BM13c, CD14, Cen11, Cha13a, Cra13, DKS10, Ere13, HP12a, Hil13, KdM13, Lan10b, Mar13b, Per13, Rah13, TW10b, dCF12]. identity [BT11c, HMP^{+}11, Hil13, Kon13, MSvW12, SS10e, WLG12]. identity-product [WLG12]. IDR [RRZ13]. if [PPK13]. Ihara [BHAP12, MM10a, SS13c]. II [LT10b, ARZ11, BF14c, CGGS13, CS10b, DK12a, EvdD10, FJMP14, HP12b, IS14, JS11, KK13, LH11a, LMMR13b, MS10b, MS10c, OT12, Siv12a, Tre10]. III [DK13b, Siv12b]. ILAS [Ano10u, BBG^{+}13, BFBD13]. ill [BJRS11, HMR12, NRS12]. ill-posed [BJRS11, HMR12, NRS12]. illustration [WA10]. illustrative [CP12]. image [FN11, GP13a]. images [Sav12]. imaginariness [BV13a]. imaginary [KS11b]. imbeddability [KM13a]. immanants [CD12a, Ye11]. Implementations [NRS12]. implicit [BBE^{+}10, DLDV11, JNS13]. Implicitly [PO11, MV12, MSS12]. implicitly-restarted [MSS12]. Implicitly-weighted [PO11]. implies [EdlP14, JN13]. imprecise [Sku13]. Impressions [WA10]. Improved [RJ10, BW12, Rua12]. Improvement [DTS11]. Improving [SR13c]. Impulse [VB10]. impulsive [LM10b]. imputation [XWD13]. In-betweenness [Aud13b]. incidence [CL13a, Dah11, DG14, DV14, RL13b, SW13, ZZ13]. Inclusion [Mel14, AAKM14, For14, FHL^{+}11, HHL10, WC11]. inclusions [MW14b]. incomplete [KD12]. increase [SC10]. increasing [Bün14]. incremental [BGV12]. indefinite [LWY14, WC11]. independence [ABK14, Dug11, FYI10, Hu10, HCY10a, LS10, TF10]. independent [RS12b, Row14c]. index [AdFST11, Bot10a, CLOR13, CFK10a, CT11b, CL10a, Das10a, Das11, DXG12, DC13, DZ11, DLDV11, DZ12b, FTA11, GW14a, GWW14a, HL10a, HW14a, HL12, Hua11b, Kim10, Kim11a, KP12, Kim13a, Kim13b, LZ12b, LL14c, MYL13, MWZ13, SBMT10, Ste10, TH10, WBHM13, XWS12, YFW13, Yua14, Yua15, ZZL11, DDM14]. index-2 [DLDV11]. Indexes [BCE^{+}10]. Indices [PdFDV14, BCdP11, Dod13, DL11, DZX12, GW10, HL10c, KHKT13, LMM11, LM12, WX14]. induced [CR10a, LMMS12]. inductive [LP14]. industrial [Zha12c]. Inequalities [BANP12, CHLW14, JMP13, MOA11, Mat14b, Shi11, ZHQ13, AP14, AK12b, Aud12, AH13b, BP14, BLdPS10, BG12a, BS12d, BH14b, Cec10, CFL13a, Coh14, CMS12b, Dra12a, Dru13, DLV13, Fur11, Fur12, GD11a, Gon11, HKK^{+}12, Hla13, KS14a, KP14b, Lac13, Lan14, Lee10, LLW14, LY13, Lin13, MA12, Maz10, MTS11, Naj13, Pat10, Pat12a, Seo13, SW11, SR13c, TG13, TKLX14, Tia10, TCL11, TPZ12, Wad14, pWlW14, Zha14b, ZH12, Zha12a]. inequality [Aud13a, BLdP12, BR14a, BR10, CMS12b, CL12c, Der13, FZ11, FT10, FLS10, FNY13, FL10, GRdS12, HOT13, IIK^{+}13, KSA11, KSAM12, KMSC14, Kia14, KPY11, Lin14b, Lu12, MMA11, MR10a, MPP11, MMM13, MR10c, Mos11, MN12a, MMA12, MMK13, Nie10, NA13, Rua12, ST10a, TmYsH11, Yam13, YJ12]. inertia [AHLvdH13, BHvdH11, BBH^{+}12, GS10c, MYL13, MWZ13, Tao11, Tia11b, YFW13, YZF14, Zha14a, dS12a, vdH13, vdH14]. inertias [BDM^{+}12, GLZ14, GOvdD14, Tia10, Tia12, YHH12]. Inexact [BGW12, YZ11, BJ13, SS13b, XE11]. inference [GD11b]. Infinite [SQ14, VS11, WY13, dFdADVJ10, Ben10, Bie14, BG11, HS12b, Per13, Pin11, Sha13a, Sha14b, Slo13, Vij14]. infinite-dimensional [Sha13a]. infinitesimal [SC12]. influence [Mac13, Nik13]. information [Gle11, Kam10, KS10, KS12b, KM14, Yan10a]. Ingram [Zha12a]. inheritance [FH10a, Pan11]. inhomogeneous [VV13]. injections [BS14]. injective [BO12, ZHQ14]. injectivity [Gna12]. Inner [MARC13, Tan14b, BKV14, BS12e, HMT10, HZGY12, Pat12a, SP13]. input [BCY12, BLLM13]. inseparable [Dor10]. Instability [KHKT13, FS11]. Integer [DJ14, BM13d, DHS12, Gre12, MRW11, MY14, MO14]. integers [Hil14a]. integrable [Sch10]. Integral [CR10c, CL13b, SS11b, Sim10, BP10, Bey12, BCŠ10, BC12b, BC12c, BD12b, CKST11, FT10, GHW^{+}13, He14, IM11, Ili10a, JK13, Kal13b, LS12a, MPS10, Mat13a, Mat14a, PHS13, SZ14, TH10, ZWL13, dFdADVJ10]. integrality [Kra12]. Integrally [GOSvdD14]. Integrating [MLC^{+}10]. integrator [Zhu11b]. intentions [NT14]. interactions [Böt13, CGGS13]. interconnection [VB10]. interior [GOSV12, LL10d, XSS13]. interlace [BH11b]. interlaced [MS13c]. Interlacing [AT14b, Kus13, AFHP14, BFdP10, FPC13, GLS13a]. internal [Gum13, Wu13a]. interplay [SB11]. interpolants [CH11]. interpolating [BPDC14]. Interpolation [AL14, HWH14, AL13b, AIL12, BO11, FKR11a, Fuh10b, KOR14, PT14, Sav14, SV13]. interpolatory [BGW12, GMV11]. interpretation [GG12]. intersection [MGLW11, SB12]. intertwining [Che13]. Interval [GPT14, Mys12, BHMR12, Dah11, Hla13, KS12a, LLD13, LLW14, LL13d, MMP13b, MP13b, MS14c, MP14b, NS11c, PM10, PKR12, Roh11, SH10]. Intervals [AG13, FZ13]. Introduction [Est12b]. Invariance [dSP10a, Bai14, BH11b]. Invariances [BHKL10]. Invariant [AFLN12, Baj14, dSC14, BK11b, BLLM13, Bou13, CHJ13, Cir14, CLHQ14, DXG12, DGÇ14, Dom10, DZ12a, FGS11, FCL10, FJ10b, Fur12, GK11, GV11, HG11a, HL11d, IJ12, Jun12, MZ14, MA12, Pro10, PP13b, RSS10, Sku13, WL12, Wój14b, ZH12]. Invariants [AW10, Bap10, Dod13, DJ14, FV13a, Für10a, Góm10, GS10b, Lop11a]. Inverse [BdP13, JS11, MZ14, NM14, SVP11, BM14a, BdFdP11, BFdP12, BT13, BCEM12, Boz13, BZ13, BZZ^{+}14b, xCwXL11, CSV14, CGMS10a, CGMS10b, CGRVC13, CI13, DS11, Den11, DW11, DMS10, DMS13, Fan10b, FHS14b, GP12, HRW99, HLS10, HH11a, HP04, JM12a, Ji12b, JDY13, JS13b, KS12a, KW12, LT12a, LXL^{+}14, LMYY11, LJY13, LSH12, MH13b, MZ10, MD10, Mos13, Mou12, MAM^{+}13, NS12a, NS11b, Nor11, PH12, Pat12b, yPjXL11, PRT13, RDD14, SS14, SH13, Wei13b, Wei13a, WJT13, WJ12, Wül13, XX14, XLG^{+}13, XSH14, Yan14, lYnZpYH13, YZ11, YWX13, ZW12b, Zha12c, ZH11b, ZCC12, vdH13]. inverse-positivity [JS13b]. Inverses [BG14a, Bün14, AM13d, ACG14, AL13c, BS14, BCEM10, COvdD10, Dra12b, Dra14, FKW13, HTS14, HF12, HZ11a, Hua11a, HZGY12, HZJ12, Hun14, KKM13, LHLL10, Mar11, Mar13c, MS11b, NCI13, Siv13, SPKS12, WC12, WL10a, XCS13, XSZ13, YW10, ZBW12, ZZC14]. Inversion [FP11, BOZ10, DV10, ER13, MS11a, MR10b, MR14c, DDP13]. inversion-free [MR10b, MR14c]. inversions [Ghe14a]. invertibility [CRS14, DCIW12a, HC14, KCID12, ZZCW13]. invertible [AK11, CHLS12, HN14, KKL14, Liu14a, TZ13b, Wik11, Wik12]. Inverting [SS10e]. investigating [TMSS14]. invincibility [FJ10a]. involution [AF12, LL11b, LHL12, LL10b, RDD14, ZZCW13]. involutions [Slo13]. involutory [Kis15, fLyH11, Tre10, XX12]. involving [BLLX11, CFL13a, Das10b, Das11, Das13, Den11, DD11d, HRT13, KPY11, Lan14, Mat14a]. irrationality [GWW14b]. Irreducibility [Had12, BB13c]. Irreducible [DDGH13, BE10, Cer10, JMS11, Kim13a, RY12a, Ser11, Ter13, Ter14, YHH12]. irregular [CH14, NLL13]. ISBN [Bar10b, Gar12, Gle11, Grü12, Lim13b, Rod12a, Tam12, Zha12a]. ISBN-13 [Bar10b, Gle11]. Ising [Böt13]. Isolation [Bea12]. Isometric [BJ10c, AS12a]. Isometries [AAH13, GP14, HN14, AM13d, BMN13b, BJ11, BJZ12, Cir14, Dug12, GW14b, GP13b, Gu14, Kho12, MS14a, Mol13, PP12a, Sar14]. isometry [BT14a, GW14a]. Isomorphisms [ÁvW13, KZ10]. Isospectral [BW12, GLP^{+}13]. isotone [NN10, NN13]. Israel [BDK^{+}10]. Issue [KPRT14, BBD^{+}11, BMS14b, FKLT13, GRS^{+}10, LPQdS10]. iterated [BV11]. iterates [CFL13b]. iteration [DYW14, HT10b, PPKR12, Rhe10, XE11, YZ11]. iterations [GIP12, HHT13]. Iterative [CAV13, FS14b, GL10c, HN10, LL13b, LJY13, NRS12, WJT13]. Ito [HWG13]. IV [BRA11, DK13c]. Iwasawa [HHT13].

J [Rod12a]. Jackknife [HYF14]. Jacobi [BP12, BdFdP11, BFdP12, BCEM12, BF14d, DD11b, HSS10, HSS14, KZ11, SB12, ŠŠ11e, SS13f, Wei13b, WJ12, Xu12]. Jacobian [BVV12, Gna12, Sun13, Yan11]. Jacobians [GS12a]. Jacobson [Wu10b, ZCC12]. James [AR12, CP11]. JB* [Ili10b]. JB*-triples [Ili10b]. Jensen [BH14b, Kia14, KLP12, KP14b, MPP11]. Johnson [Gar12, GZH14]. join [BHvdH11, BYZZ14, CMRR13, MH13b]. joined [MR12]. Joint [CN10b, Dum13a, GZ13, CN13c, DHLX12, GB13, GMV11, Koz10, Koz14a, LV11, LP12, LX13, MR14b, OM13, OM14, Pep12, Sed11]. Jordan [DWW14, BS12b, BH10, BM13c, BF14d, CT11a, CM14, CMZ10, DW14, FMM13, GS10c, GTR12, GC12, HW11, HLW10b, JV10, Ji12b, JH10, LW12d, MMMM10, Mar13b, MW10, Mol12, Mol13, NSC13, SM10b, Tao11, Tao13, TKLX14, WW13b, XW10b, ZZ11, ZH11a, ZZ10, ZZ12]. Jörg [Gre13]. Jose [LPQdS10, FdC12]. jumping [DKOT12].

Kaczmarz [NT14, PPKR12]. Kadanoff [BBGM12]. Kadison [BR10, DH12b, WY14b, YJ12]. Kakeya [RL13a]. Kamvar [Gle11]. Kapranov [Shi12b]. Karaduman [Hil14c]. Karcher [BI13, LY13]. Kato [BLdP12, AS12d, CL12a]. Kaufman [HLZP13]. Keller [YT13a]. Kellogg [Joh12]. kernel [BKV14, BCD10, CD12c, Dur12, OM10, Oto12, SST14, SWA12, Wor14]. Kernels [HG11a, Che13, Hun14, Kec13, Lom11, Rud12]. Kerov [GH12, Slo12a]. Khinchin [FT10]. kind [Xu12]. Kippenhahn [GW13a]. Kirchhoff [BCE^{+}10, DXG12, DC13]. Kittaneh [Dru12a]. KKR [BTYZ12]. Klein [GS12e]. Kleiner [SY12]. Kleinian [KK14, Yan10b]. Ko [OLW14]. Kohn [YM12]. Krawtchouk [NT12b, Wor13]. Krein [AGPPF12, PT13a]. Kronecker [BJRS11, CSV14, Dod13, HFS13, HTS14, HF12, OZ10, Tad12]. Kruskal [BH13b, Der13, Rho10]. Krylov [BFS11, CK13d, DZ12a, JR11, Jbi10, MSS12, RRZ13, Sad12, SE13, Sto12, Xu11, Gre13]. Krylov-type [SE13]. Kwong [Naj13]. Ky [Lin11, GMRS14, SRdAG10].

labeled [WNM13]. Lagrange

[Ma10a, TX12]. Laguerre [Lan14]. Lanczos [BFS11, PPZ14]. Laplace [MW10, TW10b]. Laplacian [LT11a, AFHP14, ACG^{+}11, ACM^{+}12, AKM14, AOTR13, AT14b, BS11a, Bap13a, BMSW11, Bel14, BL10b, BRZ11, Boz13, BZ12a, CGTR14, CFK10a, CT10, CW10, CJ12, CT12, CTG13, CS10b, Das10b, Das11, DXG12, DXL13, DGÇ14, DLS14, DL11, DZ12c, Est12b, FF12, FTDA10, FYI10, FHRT11, FHRT14, GK14b, GLS13a, GLS12, GLS13b, GW13b, GCY14, HMTR10, HL10a, Har14, HL11a, HL12, HJZ13, HT10a, HY14, HQS13, LSC10, LS10, LSC11, LWZ11, LW12b, LZ12b, LTS13, LGS13, LY11b, LL10c, LL11c, LW12e, LLT13, LH13, LL14b, LL14c, LSD14, MK12, NSW13, NP14, NLL13, QSW14, QY12, RJ10, RW10, SW13, Sri13, Suz13, TW10a, wTmS12, wTlW13, Ter11, VDVJT13, WB11, WL12, WF12, WBHM13, WAH13, XZ13b, XM11, YY14a, YFW10, YL10, YWS11a, YWS11b, ZZ13, ZHG13, Zho10, ZSWB14, Zhu10a, Zhu10b, dLOdAN11, dLN13, vDO11, vDF14]. Laplacian-eigenvector [RW10]. Laplacian-energy-like [DXG12, DGÇ14, WL12]. Laplacianness [Hu10]. Laplacians [AH13a, Bau12, BFF^{+}11]. Large [AM13c, BGP13, BFdP12, GM14, LRV12, LSV12, dSP12a, AW11, BHMO13, FHRT14, GG13, Jbi10, KKM13, LHZH11, MAS12, dSP11a, dSP11b, dSP12b, Sad12, Wan11b, WCKL13, WZ14b, ZL11]. Large-scale [LRV12, KKM13, Sad12, WCKL13]. largest [AHL11, ABK14, CFK^{+}10b, CW10, CQYY13, Das10b, Das11, Kol13, KY11, KPY11, LY11a, LGS13, LGSC14, LHW11, Nik11, PJ13, QSW14, Sta12, WZY14, ZLW12, ZCQ13, Zhu12c, dLN13]. last [LWGM10, LWGM12]. Lattice [NN13, HW14a, IPFD13, Maz10, PY10, Suz13]. Lattice-like [NN13]. Lattices [CMZ10, CSC13, Kub13]. Laurent [He11, Lak10b]. law [DD11c, MTS11]. laws [CIH11, SR13a]. Lawson [Seo14]. layout [LRST10]. LB [Nom14]. LCM [TL13a]. LCMs [BDD13]. LDU [BH12, Kim11b]. leading [JN13, KS11a]. learning [MLC^{+}10, YZ12a]. Least [LB14, MR10a, Bel14, Byd10, DS13, Dum13b, FF12, FZW11, GJTP13, GCY14, HMR12, Ji12a, Kol13, Kyr13, LJ11, LW12b, LWY14, LZ11a, LZG14, LZ14, LXK12, MM10c, MR13, MR14a, MS11b, Miy14, PO11, dSP12d, PAS11a, PAS11b, PPKR12, TF10, WF10, WF12, WBWH13, XZ13c, XZ13a, Yu13, Zhu12a, LKN13, SS10c]. least-squares [GJTP13, HMR12, Ji12a, LXK12, LKN13, SS10c]. Ledermann [Hür13]. Left [CL12b, Raf14, MVPS10]. Left-looking [Raf14]. Left-symmetric [CL12b]. Lehman [Shi12a]. Leibniz [LMO16, AM13b, BCS13a, CGO10, CCGR13, CCGO14, CK13a, CILL12, CLOK13, CKLO13, KRH14, MD12a, MSD13, RRKK12, TX12]. Leja [KOR14]. lemma [BLL12, ZCC12]. Lemmens [Lim13b]. length [CCGO14, CK13b, CDM12, DF14, Nik10b, Rud12]. Lengths [CL11b]. Leonard [BM12a, GWH13, Han11b, Han13b, Han14b, HWG13, HWG14, wH12, NT12b, Nom14, Wor13]. less [LZG14, LHW11]. letter [AM14, HMS13]. letters [AM14]. level [Blö12, KO13, KW12]. levels [MR12]. Levi [BdlC13]. Levine [Sat11]. Lewin [LYL13]. Lexicographical [WL10b]. Li [Lim11a, Dru12b]. Lie [CdGS20, AIS14, BD12a, BZWL13, BM12a, BDF11, BdlC13, BdlC14, Ben11, BE12, BCS13a, BM13b, BS12e, BDV12, CT14a, CNT12, CdGS12, CLHQ14, DDF13a, DD10b, DW12b, GST13, Han11a, HT10b, JQ11, LL11b, LW12d, LCM13, LNT13, LJ10, MD12b, Mar14a, MSvW12, NT12b, Pop13b, QCH11, QH13, ŠK10, SM12, TT12, WGL12, WW13c, Wan14b, Wu13b, XW12, YZ10, YZ12b]. Lie-orthogonal [Pop13b]. Lie-solvable [MSvW12]. Lieb [Aud13a, GK14c]. Liesen [Gre13]. like [CGMPSS14, DXG12, DGÇ14, DF14, GNE^{+}14, GTR12, Nat13, NN13, PT13b, WL12]. Lim [Seo14]. Limit [BHDW12, BMSW11, OM14]. Limited [EM12]. Limited-memory [EM12]. Limiting [PS12]. limits [SC12, VW10b]. Line [CKAC14, RJ11, Roj11, ZSWB14, GKZ11, GW13b, GRM^{+}10, LW12a, LL10a, LFS12, MWZ13, MV12, Vij14, WS12, WY14a, WF14]. Lineability [CGMPSS14, BG13]. linéaire [Cas10]. Linear [Ano12a, AJ13, BEM12a, BDK11, BFBD13, Bra10, Che14a, Che14b, CdGS20, CD12a, DS13, DS12b, DD10c, DWW14, Duk15, EKSV18, Gem10, HS14a, HQ13, KSB12, KB14a, KKL13a, KPRT14, LMO16, LT16, LPS13, Lu11, MR14a, MR14c, NT11a, PQ10, Rue13, SHZ10, SA14, Vla12, Wik12, WZ14c, XW11, YWX13, Yua15, Zho11, AS10, Ara12, AMJ14, ADW13, BB13a, BV12a, BEM12b, BLLX11, Bal10, BBCC13, BV11, BGP13, BM10, Ben10, BS12d, BLLM13, BJRS11, Bou10b, CDP10, Car13, CC14, Cas10, Cec10, CM13, CPV10, CFHL14, CA10, CW11, CL11b, Cos14, CT14b, CMN10, Czo10, CN10d, CN11e, Dai11, Dai12, DLDV11, DX13, EY13, FFS11a, FGQ11, FFS11b, FGR13, FS14b, Fur11, FRS14, GZX14, GEP10, GEP13, GD11a, GRdS12, GMH14a, GMH14b, GK11, GHT11, GJTP13]. linear [GD11b, Hla13, HM10, HG10, HG12, HN10, HZ11a, Hua11a, HZJ12, Irv12, JV10, JV11, Ji12a, JMP13, Jun14, KH13, KSAM12, KRS13, KRvS12, KRvS14, Kis15, Kla10, KN13b, KN13c, KKM13, KB12, Lan13, Li12, LW13a, LLD13, LLW14, LT10a, LHZH11, LJ12, LS14, LMMR13a, LMMR13b, LM10b, LdlP11, LL13d, MMS12, Mah11, MD13, MZ12a, MR11, MLC^{+}10, MW14b, MO14, MSP13, MR10c, Mys12, NP13a, NRS12, OLW14, PQ12, dSP10b, dSP10c, dSP10e, dSP12b, dSP12d, Per11, PKR12, PPKR12, PTPL10, PE13, Pro10, RAY14, RdSP11, RS14a, RŠ10, Roh11, RS12c, SSS13, SB12, SVP11, SW11, SS11d, SH13, SLS13, SS10e, TZ13b, TmYsH11, Tao13, Tra12, Tre11, TP13, Wal11b, WGL12, XDFL10, XX12, wXL14, wXZ19, YT13a, YW10, YJ12, ZW12b, dOFK^{+}13, dlP11, dSC14, BPRY11]. Linear-quadratic [NT11a, Jun14]. Linearity [BEV13]. linearization [MP10]. linearizations [AA11, AAK11, BCF14, BF14c, DDM12, HMP12]. linearly [AA14, Bün14, CK14, CGM^{+}10, Gna12, RS12b, ZHZF13]. lines [BH10, DS10, Lee13c]. link [BCF12, Pit11]. Liouville [jAS13, KZ11]. Lipschitz [BJ10a, GV11, JRMFSS12, Koz10, Lán10a, PP13b, Rod11, Rod12b]. Lipschitzian [JV11]. List [Ano11a, Ano12c, Ano13e, Ano13f]. Lists [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, CL14b, ES14]. lists [JNS13, KS13a]. lit [wH13]. lit-only [wH13]. Littlewood [AW10, AP14, Lac13]. LLL [LQ11]. LMI [PKR12]. loadings [ZHZF13]. Local [Cos14, BS11b, Ben13, Bou10b, CD11, JZZ13, KY11, Kra13, TZ12b, Tra12, AKA13, FP11]. local-global [Kra13]. localization [Elo11, GPR13, LL14a, RŠ10]. localizations [DTL11]. localize [SYH14]. Locating [JT11a]. location [JTT13, Wil11]. Loewner [Aud13c]. Loewy [Ano13-27]. Log [WZ14a, Alt13, CM12a, Fur12]. Log-convexity [WZ14a, Alt13]. log-determinant [CM12a]. logarithms [BLdPS10, Chi13, LNN14]. logistic [YiKIS12]. lollipop [GSL11, GW13b, WS12, WY14a]. lonesums [KKL13a, KKL13b]. long [BLLX11, CBB13, WZL13]. looking [Raf14]. loop [Bap10, GS10b, YW11]. loop-free [YW11]. loopy [GM12]. Lorentz [KN10, TD13]. loss [HLP12]. Low [BGV12, KRS13, AJ13, CAV13, DBZZ14, MZ12b, Sad12, GPT12]. Low-rank [BGV12, KRS13, MZ12b, Sad12]. Lower [AP14, MNZ10, MNZ12, Pat10, PJ13, Rad10, Cha10, CFL13b, Grc10, GR10, GCY14, LLH10, LMYY11, RJH11, Slo12a, WXH10a, Zim13]. lowering [BC14a]. Lowest [LTX14]. Lowest-rank [LTX14]. Löwner [Han13a, MN12a]. LP [Pin11]. LTI [Per12]. LU [Hua13a]. Lucas [DGMS14]. Lur’e [Rei11b]. Lyapunov [FH10b, GTR12, GKS^{+}10, Jbi10, LXK12, LTX14, PJ13, Sad12]. Lyapunov-like [GTR12]. Lyapunov-type [LXK12].

M [Gar12, Gem10, AS12a, Rhe10].

Macaulay [BDD14]. magic

[CMNW14, CCL14, dHLMS13, Hun10, LLN^{+}12, LGZ14, Nor12, Nor14]. Mahalanobis [GH11]. Main [CSZ10, LS13c, TH10]. Majorization [BEM12b, BD12b, CPK11, Dah10, For14, KT12, MOA11, Nie11, Uch10, Zha12a, AK12a, AAM12, AMJ14, BEM12a, Dug11, FZ11, FW14a, Fur12, KP14b, LL14d, Nie13, SA10, SA14, TKLX14, TPZ12, Zha14b]. majorizations [Nie12]. make [DTL11]. manifold [YZ12a]. Manin [FP13]. manpower [DT10]. many [DLNN14, Mat14a, MO14, TL13a]. map [GK11, Ha13, HNZ12, Lin14a, dSP12d]. Maple [GS12b, Kla10]. mapping [AA14, JSS13, Sko11]. mappings [BP11, CW11, CN10c, CN11d, FFS11a, FRS14, HW11, Han14a, KN13c, LLF13, RS12c, Wój14a, XW10a, Zho11, ZXZ10, dOFK^{+}13]. Maps [ABSV12, Cos11, GN14, HLW10b, HH11b, JG10, LP12, MS10c, WLG12, ZH11a, ABEV10, BS12d, Bou10b, BMS14a, CC12, Che14a, CFL13b, CLS10, DD10b, DHKQ13, DW12a, Fra12, Fra13, FH10b, Fur11, Gna12, GZ13, HKPR13, HQ13, Ika11, JH10, KS14a, KSAM12, LT12b, LT16, LP10, LW12c, LW13a, LHL12, LT10a, LL10b, LT13, Liu14a, Liu14b, MR10c, Nie13, Pan12a, RO10, Sun13, Wan11a, WFM11, YT13a, YJ12, dlP11]. Marcus [LL12]. Marcus-Minc [LL12]. Marix [BW11]. market [Vas14]. Markov [CK13b, Cas13, DT10, DGU14, Góm10, HB12, Hun10, Hun14, Kir10, Kir14, MA10b, Nem13, PR10, Puł11, Rhe10, Sku13, VV13, Vas14]. Markovian [Hun14, Koz14a]. Marshall [Zha12a]. Martingale [Dah12a]. Massey [KY13]. Mastronardi [Gem10]. matchgates [LMN13]. matching [Beh13, GL12a, HL11a, KW13, LT11a, SS11c, Ste10, wTmS12]. matchings [GX12b, LSC10, TS12]. mathematical [BELK12]. Mathematics [Ber09, Gar12, Grü12, Lim13b, Rod12a, Sla10, FdC12, Bar10b]. Mathias [Lim11a]. matrice [BW12]. Matrices [Bar12b, Bra10, FJ11, Fie11b, Gem10, HMS13, KKL13a, KKL13b, LSTW13, Rei11a, SM13, AMP10, AiS13, AG12, AFLN12, AA14, AG13, AM13a, AAF^{+}12, AKM13, AGM14, AB12, AHAPP10, Ali12, ANP13, Alo14, ÁNPQ12, ÁAFG12, AdF11, ART13, ALRV12, Ang13, AT11, AT14a, AK11, Aud10a, Aud13c, BBG14, BT14a, BB13b, BT11a, BB13c, Ban13b, BS11a, BG14a, BP12, BKM^{+}13, BB13d, BH12, BMW10, BSK12, BSKL13, BFdP10, BLdPS10, BFdP11, BdFdP11, BFdP12, BT13, BdP13, BDH^{+}12, BOZ10, BOZ11a, BANP12, BCEM10, BCEM12, Ben10, BM13a, Ben14a, BB10, BB14, BS12c, BP11, BDG13, BD13, Bie13, Bie14, BBE^{+}10, BI13, BH10, BBGM12, BTYZ12, BSU14, BCS13b, BC12c, BC13, BC14b, Bot10a, Bot12, Bou13, BLL12, BLL13, BL14, BRZ13, Bre14, BRLS12, BHZ10, BK10, BD12b]. matrices [BFH^{+}12, BKMS13, BC14c, BBS12b, BLS14, BZ13, BK12, Buj13, Bün14, BAD09, BS13b, BW13b, CRS14, CRSS14, CHJ13, CD14, CRU10, CRU13, CRU14, CSV14, CH13, Car10a, CT11a, CN12a, CFPP13, CR10b, CK13b, CAV13, CGM11, CGMS10a, CGRVC13, COvdD10, Cau11, CCF^{+}12, CS10a, Cha14, Cha13a, CPR10, CTW11, CM12a, CL10a, CGMJ14, CL10b, CG13, Che14b, CJ11, CK13d, CN13a, CN13b, Cho13, CN11d, CGSCZ10, CD13, Cir13, CM14, CNPP12, CD12a, Coh14, CDM12, CP10, Cos11, CFLW13, CL12c, Dah10, Dah11, Dai11, DHLX12, DoMP09, DL14, Dax10, DLNN14, DH10a, DMP11, DD11b, DK12b, DP12a, DHS12, DGMS14, DFS12, DFS14, DS14, – Dok12, DKOT12, DHKQ13, DO11a, DO11b, Dom10, Dom13, DA10, DT11, DdF13b, DdF14, DBZZ14, DdC13, DGMS10, DHS10, ER13, Elo11, EvdD10]. matrices [Ens10, ES11, Est12b, EGR12, Fan10b, FM11a, FFJM14, FGS11, FFG^{+}11, Far11b, FKW13, FJ10a, FdC10, Fie10, FH10a, FM11b, Fie11a, FH12a, FH12b, FM13, FH13, Fie13, FHS14a, FH14, FMR12, FFK11, Fou14, FGvRR13, Fra12, Fra13, FW14b, Fuj10, FJ10b, FJ14, Fut12, FHS11, GMS12, GEP10, GOSvdD14, Gar13, Gau10, GTW13, Ghe14a, God12, Gol13, GS10b, GL12b, GGK^{+}13, Gre12, GS10d, GS12d, GZ12, GR12, GKR13, Guo13, GY13, GS12f, GG13, HO11, HC14, HFS13, HSS10, HSS14, HTS14, HRW99, HF12, HR11, HMT10, HP12b, HL11b, Hil14b, HM10, HS12b, HB12, HN14, lH10a, HLS10, HZ10, Hua10, HCY10a, HL10b, HC10, HZ11b, lH11b, HH11a, Hua11c, HL11d, Hua12a, HLZ12, HZ12b, Hua12b, HLZP13, Hua13a, HTW13, Hua13b, HHLS14, Hür13, HP04]. matrices [HP11, Hwa12, IMA10, Ikr10, ISYY11, JM12a, JP11, JM12b, JZT14, JKV13, JLW11, JT11b, JS11, JK11, JPS13, JS13b, Kak10, KSB12, KS12a, KMNS12, KH13, Kar11a, Kar11b, KM13a, KSS12, Kau12, KKL14, Kis15, KS10, KN13a, KJK13, KW12, KT10, KT12, KM13b, KS11b, KS11c, Laf12, LS13a, Lav10, LRT12, LRT13, LLY11, LV11, LL12, LV12, LL13b, Lee13a, Lee13c, LHLL10, tLyLWqW10, fLyH11, LNTgW12, LL14a, LT10a, LD11, LT11b, Lim11b, Lim12, Lim14, LC10, Lin13, LCZ10, LHZH11, LMYY11, LS12c, LZL12, LS13d, LJY13, Liu14a, Liu14b, LMT10, LM10c, MMS12, MVPS10, Mac13, MM12, MM10b, MARC13, Mat14b, MS10b, Mat12, MH13b, MRW11, MPT14, MY14, MQ11, MQ13, MQ14, MMRR11, MMRR12, MPP10, MPRW11, MPT12, Mer12, Mer10, MSvW12, Mit11b, MPSS10, MMP13b, MS13c, Mos13, Mou12, MAM^{+}13]. matrices [MP13b, MS14c, MP14b, NS13, NS12a, ND11, NPP13, NS14, Nik11, NY13, NT11b, Nor12, NV12, NS12c, O’D14, Ogu13, OSZ10, OM12, iO12, Özd13, Pan11, PP11a, PM10, dSP10a, dSP10c, dSP10d, dSP10e, dSP11a, dSP11b, dSP12a, dSP12b, dSP12c, dSP14, yPjXL11, Per11, Pop10, Pop14, PV12, PJ13, Pry10, PSW11, Qui11, RR14, RY12a, RMT11, RMP14, RW12, RdSP11, RR11, RSS10, Ros12a, Rub13, Sag13, SZ14, SS12b, Sbu10, SB12, SS14, Ser11, SCSS10, SS13b, SST14, Sev10, Sev14, SSMS14, SHZ10, SH10, Sha14b, SYH14, SS13d, SS13e, Shi12a, Shi12b, Shi12c, Shp10, SLS13, Siv13, Slo13, Slo14, SC12, SDNS13, Sow13, Spe11, Sra13, ŠŠ11e, SS13f, SS10e, Stu13, Sun13, Szö13, Tad12, TT10, Tan10b, TL13a, TX12, TMSS14, TW14, DDP13, DDP14]. matrices [TW10b, Tia10, TW11a, TCL11, Tre10, Tre12, TW11b, Tsa11, VS10, VU14, Van10, VR12, VS11, Voy13, WLHL10, WXH10b, WXH10a, WFM11, WW13a, WW13d, WY13, Wil14, WMZ14, Wu10a, WJ12, XX14, XZ13a, XZ14, XY11, Xu11, Xu12, XX12, XLG^{+}13, wXL14, XSH14, wXZ19, Yam13, YT13b, Yan14, YX13, YS13, YWX13, ZWZ10, ZYL10, ZLL12, Zha14b, Zha14a, ZLH^{+}14, ZZ10, Zha12c, Zho12, ZL11, ZH11b, ZBW12, ZCWZ13, ZJ10, Zuo10, ZXX13, dCF12, dFBRS14, dHM11, de 13, dlCdlRMP14, dlRMP12, Gar10, JMW11, Kaw13, Gar12, BOZ11b, Grü12]. Matricial [JS13a, CH11, FKR11b, FKM13, tHR13]. Matrix [jASZ12, jAS13, Bar10b, Ber09, BC12b, Bot14, CC12, DW10, Gem10, GB13, HH13, HNZ12, HJ12, JS13a, Koz14b, MPT14, MZ12c, PQZC11, Ros12c, Sla10, UW11, AAK11, AAKM14, AKM13, AAJ12, AAT12a, AAT12b, ABBO11, ÁvW11, ÁvW13, Ano12-30, AH13a, AW10, AHL^{+}14, AH13b, AKA13, BPA^{+}11, BLLX11, BH14a, BT10, Bap13b, Bap13a, BDD14, Bat14, BM14a, BS12a, BHDW12, BMS10, BR14b, Ben13, BCEM12, BCD10, BL13a, BEK13, BIT12, BMR11, BK11b, BG12a, BR14c, BCŠ10, Bou10a, Bou11, BMM12, BCDM13, BGH12, BDOvdD12, BKMS12, BSST13, CC10, CRS14, Cas13, CLST14, CJR11, CGGS13, Cha13b, CFJKS13, CH11, CL13b, CK13d, CHLS12, CK14, CN10c, CL11b, CD11, CJ14, Cim11, CI14, Dah12b, Dah12a, Dai13, DIP13, Dau12, DD10a, DDM12, DTS11, DH10a, DH12a, DW11, DF14, DKS10]. matrix [DKS13b, Dod10, Dod13, DS14, DGKO13, DMS10, DMS13, Drn13a, Drn13b, Dru14, DW12b, DW13, DWW14, DW14, DZ13, DLV13, EKSV14, EKSV18, EM11, Ern13, FP13, FLH12, FDS13, FdC10, FMM13, Fis14, FHL^{+}11, FP11, FG13b, FKR12, FH12c, FHS14b, Fuh10b, FLS10, FL10, FRS14, GEP13, GHMPVP11, GPT12, GS13a, GW14a, GWW14a, GN13, Ger12, Gho13, GP13a, Grc10, GIP12, Gro14, GTV12, GS10d, GS12c, Guo10a, GKL11, GM11b, Güv12, HH12a, Han11a, Hia13, HW14c, Hu10, Hwa11, IM11, IPFD13, JM14, Jbi10, Ji12b, Jim10, JN13, Jun12, KL13a, Kak10, KOR14, KHKT13, KSA11, Kaw12, KR12, Kha13, KLS12, Kim11a, Kim11b, KLL11, Kim13a, Kis15, KS10, KS12b, KM14, KS14b, KKR11, KZ11, KLZ14a, KLZ14b, Kus13, Kyr13, LS13a, LT12a, LNN14, LV14, LAL11, LL12]. matrix [LV12, LS11a, LZ12a, LW12c, LW13a, LY13, Lin10, LS13c, LH10b, LSR11, LLMZ12, LW13b, LX13, LNT13, LS12d, Lop11b, LSH12, MZ13, MMMM10, MMMM13, MM12, MSP11, MZ12a, Mar10, MR13, MR14a, MM13, Mat13a, Mat14a, Mei13, Mel13, MS12, MR10b, MR14c, Moo11, MN12b, NR10, NM14, NS11b, NV10, O’D14, OT12, Pál13, PIM^{+}10, PCC12, PPK13, Pat12a, Pat12b, dSP10b, Per13, PS12, Pin12, PRW11, Pop13a, Psa12, Rah13, Rei11b, Rim12, Rod11, Rod12b, Sah10, ST10b, SS11a, SBM11, SK14, SK13, Shi13, SA10, SC13, SR12, Sta14, Ste13, TD13, Tan11, TZ12b, TZ13a, TmYsH11, TTZ13, DDM14, Tia11b, Tia12, TGM11, TT11, TDT13, Uhl13, VS13, WWG10, WLG11, WW13b, WW13c, Wat13, Wei13b, WJT13, Wei10, WCKL13, Wik11, Wik12, WZL12, WY14b, Wül13, XW10a, XD12, Xu11]. matrix [XX12, wXL14, wXZ19, lYnZpYH13, YZ12a, YZ13, Zha12d, ZLD11, ZXZ10, dF10, vdW14, KY13, Tam12, Rod12a]. matrix-based [Dah12b, IPFD13]. matrix-convex [PCC12]. matrix-monotone [PCC12]. matrix-valued [KKR11, Tia12]. matroids [LM10a, TN14]. Matsaev [Dru11]. Max [EvdD10, FH14, Ser11, AGNS11, BM13d, Cec10, Fie11a, GPT14, GM11a, GM11b, GMS13, KSS12, Mer10, MP13a, Mys12, MP14b, Nit10, NS11c, Pep11, SW11, Ser13, WLL14]. Max-algebra [EvdD10]. Max-algebraic [Ser11]. max-linear [SW11]. max-min [GPT14]. max-plus [AGNS11, BM13d, Mer10, Mys12, MP14b]. max-weighted [WLL14]. Maxima [Yua14, Yua15]. Maximal [BCFP12, Fan12, FMR12, LT10b, RRM11, Sev14, ÁAFG12, ADW13, BNP13, CT14a, CT10, CT11b, CVW10, CDM12, DF14, DdF13b, HLS11, HJLS11, LLS11, SS11b, SBMT10, TW10a, WX14, BDK11]. maximality [HZ12a]. Maximization [Tia11b]. Maximizing [YFW10, WKV10]. Maximum [CKL^{+}13, EdlP14, HS10, NY13, Sin10c, VDVJT13, BZ12b, CCGO14, Car11, CFPP13, CH11, CH14, CCL14, CL10b, DZX12, EHH^{+}12, GL12a, Ghe14b, LLS11, LQY13, LL11c, LHL13, Sav14, SWT13]. maximum-volume [Sav14]. maxitive [Pon11]. may [SC10]. MDP [ANP13]. MDS [CNPP12, La 14]. MDS-convolutional [La 14]. mean [AGK11, BP11, BG12a, BI13, Góm10, JLLY10, KLL11, Kim13d, LL13b, NS11b, WLL14, Yam13, BT11c]. Means [CM12a, Aud13b, AH13b, BN13, BCS13b, BH14b, DLV13, Fuj11, Han14a, HP12b, HKK^{+}12, KSAM12, KLL11, LLY11, LL12, Lim11a, Lim12, LY13, LP14, Nak13, Pál13, PP14, Seo13, Seo14, ŠK10]. measure [GN14, KKR11, Maz10, WNM13]. measured [KD12]. measurements [ALPV14, FMNW14, Jen13, KG14]. measures [DW10, KS12b, Pon11, DGH^{+}11]. median [GM14]. mediant [HNZ12]. meeting [Ano12-30]. Meini [Lim11a]. members [BRA11]. Memorial [BMS14b]. Memoriam [Tsa12, BDK^{+}10, Lin11]. memory [EM12]. Mena [Kra13]. Mertens [Car10a]. mesh [GS12b]. metabelian [DDF13a]. method [BCY12, BFS11, BVV12, Bey12, BN13, BTYZ12, BH13a, BJRS11, CDP10, CHZ13, Dah12b, FS11, GS13b, GN13, GO12, Gul11, Guo10a, HSS14, JR11, LS11a, LKN13, Lin10, LL10d, LJY13, MV12, MSS12, MR10b, MR14c, MP14a, NT14, RJH11, Sad12, Sha14b, Tra13, wXL14, wXZ19]. Methods [Gre13, HRW99, JM12a, BV12a, BGV12, Blö12, BFF^{+}11, BJ13, CAV13, CQYY13, DS12b, DYW14, FS14b, GOSV12, GL10c, HSS10, Jar12, Jbi10, Ji12b, LB14, NRS12, Nik14, PT13b, RRZ13, SE13, SK14, SWA12, Wan11b, WJT13]. metric [BT10, GB13, GO13, ISYY11, Lim13a, SB11, Wol12, WC11]. metrics [HP12b, HKPR13, Kum11, LRST10, TNP12]. Metropolis [Din11]. Metzler [VS10]. Meurant [Rei11a]. México [Ano10u]. Michael [BMS14b, Tsa12]. microscopic [FGQ11]. midpoint [Lim13a]. might [PPK13]. migration [VV13, Vas14]. Mihály [Rod12a]. Miki [Tsa12]. Milman [AGPPF12]. min [Cec10, GPT14, Nit10, NS11c]. Minc [LL12]. Minimal [AHAPP10, HLW10a, Hil14b, KS11a, LW12d, Lop11a, Zhu12b, ALRV12, BL10b, BJ11, Dod13, GLZ14, HJL10, Lip10, WLGM11, ZK14]. minimality [KRvS14]. Minimax [HLP12]. minimization [IW13, LNN14, LLB13, TmYsH11, Tia11b, Zha12d]. minimizations [Fou14]. Minimizing [O’D14, YM12, LLS12]. Minimum [CH11, CJ12, GK14a, Hog10, MD13, MNZ10, PSW11, SK12, TG13, AAF^{+}12, AHL^{+}14, BBF^{+}10, BBF^{+}12, BKM^{+}13, BMN^{+}13a, BL12, BMS14a, DGH^{+}10, Dea11, EHH^{+}12, FCL10, FL12, Gar13, GHW^{+}13, HHMS10, HK13, HCY10b, HCY10a, JMS11, Ji12a, Kir14, Kyr13, LS10, MS13a, MGSW14, Mit11a, MNZ12, NS11a, SM13, SHS12, SvdH11, YL10, YW11, Zim13]. minimum-norm [Ji12a]. Minisymposium [Ano13y]. Miniversal [DFS12, DFS14]. Minkowski [BdFdP11, BH14b, Lán10a]. minor [CN12a, TN14]. minors [BS11a, Bap13a, BR14c, CN12a, CJR11, JN13, KMNS12, OTdDv12, JMW11]. Minoux [GS12f, Shi11]. Miroslav [Grü12, Ano13y, Nik13, Stu13]. Mirsky [– Dok12]. missing [MŠ10a]. misspecification [HYF14]. Mixed [DD11c, FPC13, DKM^{+}14, GP12, GD11b, Güv12, Lan14, LWY14, LJY14, Nem13, Wei13b]. Mixed-type [DD11c]. mixing [Kir10, Puł11]. mixture [CR10b]. MLE [TZ12a]. mobility [JP11]. Mock [IT11]. mode [ZHZF13]. model [BGW12, Blö12, Böt13, DT10, GNE^{+}14, GD11b, Güv12, HYF14, JDY13, KD12, LwCJL11, LBLS12, PS12, SS11c, TP13, VV13, XW11]. modeling [FGQ11]. modelled [LHH13]. Modelling [BL11]. models [ATS12, Bal12a, Bal12b, Bos11, CR10b, FH10c, LZ11b, PTPL10, SBM11, SK14, TZ12a, Vla12, YiKIS12]. modes [DD10c]. modification [LL14a]. modified [DMS13, LMT10, wXL14, wXZ19]. Modular [Sze14b, WZ13a]. module [DTL11, LT12b, LT16, NCdS14]. modules [AR12, AF12, BO12, aCCS14, Cer10, Dub14, GD11a, IRT14, MLC^{+}10, OZ10, PYZ14, Pop12, RAY14, Sha11, Ter13, Ter14, Wag11, WW10, WD10, WZ12, WZ13e, WZ14c, dSC14]. moduli [BMS14a, FP14]. modulo [Gu13, LWG13]. modulus [wXL14, wXZ19]. modulus-based [wXL14, wXZ19]. moment [DW10, Lop11b, PLL12]. Moments [Rei11a, Rod12a, AW13b, BW11, CLL12, MW14c, Nem13, SWBS13, WL10b]. monoid [ABK14, Bre14, Car13]. monomial [KS11a]. monomials [JMP13]. monotone [Aud12, BHDW12, CW11, HKPR13, JP11, Kum11, OT12, PCC12, San14, Uch10]. Monotonically [Reg13]. monotonicity [Aud13b, FJMP14, JMP12, MS12].

Monotony [BM12b]. Monov [CL14b]. Moore [Boz13, xCwXL11, HF12, HZ11a, Ji12b, KS12a, MŠ10a, MZ10, Nor11, Pat12b, RDD14, WJT13, XCS13, Yan14, ZZCW13]. Morrison [MS11a]. Morse [AJ13]. Moshe [Ano13z]. most [BH13b, KY11, LL13a]. motions [AY13, AN13]. Motzkin [CY11]. mouth [HH12b]. Multi [Blö12, LLY11, BCY12, CC10, GMS13, Sha14a, SB11, Zhu11b, ZY14]. multi-agent [Sha14a, Zhu11b, ZY14]. multi-input [BCY12]. Multi-level [Blö12]. multi-metric [SB11]. multi-objective [GMS13]. multi-step [BCY12]. Multi-variable [LLY11]. multi-way [CC10]. Multidimensional [KLP12, Lee13c, Xu14, NT11a, OT10]. multifilar [EFN09, EFN10]. multigrid [BFF^{+}11]. multilinear [BW13b, FGH13, FKLT13, GMMFPSS12, Kon13, Moh13, NA13, Pel14, SR13b, SR13c, FdC12]. multinomial [YiKIS12]. multipartite [CHLW14, JZ14, PHS13]. Multiple [DGAM14, DKS13a, FPC13, KSH12, LNTgW12, Lip10, MP14a, Psa12, Row10, Sha14a, XD12, ZY12b]. multiple-sets [ZY12b]. multiplication [HH13, Kau12, MR13, MR14a, Sow13]. multiplications [BLLX11, RO10]. Multiplicative [BE12, CLS10, JK11, WW13c, ZXZ10, Bal10, xCwXL11, GO12, Wan11a, UW11]. multiplicativity [LS12a]. multiplicities [BZZ14a, OS14]. multiplicity [AKM14, AGV12, GS12d, JNS13, KS13a, Row11, Row14a]. multiplier [Aud13c, CPK11]. multipliers [LP10]. multiplying [O’D14]. Multispherical [KT10, KM13b]. multivariable [PP14]. multivariate [AS14, BDD13, DU14, GH11, Han14a, LW13b, Wal11b]. Musings [Moh13]. Mutation [Sev13, Sev10]. mutually [Kis15, XX12]. Mysteries [NSW13].

N [BPRY11, Gem10]. Nanjing [BBD^{+}11]. NASC [BBD^{+}11]. Natural [MRW11]. Naturally [ACGVK14, CGO10, CLOK13, LMO16]. near [MV12]. nearest [ABBO11, NR10]. Nearly [SSMS14, Bod13, CFK^{+}10b]. necessarily [Pop14]. Necessary [CHK^{+}13, LLD13, Sha13a, FS14b, LS12d]. needed [DTL11]. negacyclic [La 14]. negative [Aud10a, BV13a, CR10b, FKM13, KS13b, MYL13, MWZ13, PT13a, Wol12, ZCKS12]. neighbor [ACM^{+}12, Böt13]. neighbourhood [LL13c]. nest [Gha13, JH10, OTdDv12, QH10, SM12, ZZ11, ZZW10, ZHQ14]. nested [AdFST11]. network [BCE^{+}10, FFS11a, SB11, TN14]. networks [BFRR14, Car11, Est12b, GPR13, Gle11, HWSH13, Kam10]. Neumann [Tsa12, BD12a, BMS14b, DLLS10, LZ10, LLF13, PRT13, QH13, SS10a, SSS13, YZ12b]. Nevanlinna [AL13b]. Neville [HZ14]. news [ZCKS12]. Newton [BH13a, BJ13, CA10, FS11, Guo10a, Jar12, JMP10, JMP13, SK14, SV11]. Newton-based [FS11]. Newton-type [CA10]. nice [CHZ13]. NIEP [CL14b, ES14]. Nijienhuis [LCM13]. Nil [BCDM13]. Nil-clean [BCDM13]. nilindex [ACGVK14, CCGR13]. nilpotency [Sun13, Tan10b]. Nilpotent [BDF11, Hua11b, BDH^{+}12, BdlC13, BdlC14, BVV12, CLOR13, CdGS12, CdGS20, FP14, GS13b, Kha13, KLZ14b, MMS12, NS14, OR12, Tan11, WGL12, de 13, BVV12]. nilpotent-centralizer [GS13b]. Nilpotent-Jacobian [BVV12]. nilradical [CLOK13, LMO16, ŠK10]. Nineteenth [Abe11]. nnd [NCI13]. no [KD12, MD13, Ziv12]. nodes [LWV12]. Nomura [CM11a]. Non [JRMFSS12, MMA12, ALPV14, Aud10a, BH14a, BP10, BT13, BCS13a, CR10b, CKST11, Cos14, DL14, FKM13, GK11, KS13b, LLH10, LYL13, MMP13a, Net10, dSP10e, dSP12b, Per14, PT13a, Qui11, Seo13, WZ14b, YW11, dO12]. non-bipartite [CKST11]. non-closed [Net10]. Non-commutative [MMA12, Seo13, dO12]. non-degenerate [GK11]. non-fixed [Cos14]. non-Gaussian [ALPV14]. non-Hermitian [BH14a, WZ14b]. non-hyperinvariant [MMP13a]. non-Lie [BCS13a]. Non-Lipschitz [JRMFSS12]. non-negative [Aud10a, CR10b, FKM13, KS13b, PT13a]. non-powerful [LLH10, LYL13, YW11]. non-self-adjoint [BT13, DL14]. non-singular [dSP10e]. non-singularity [dSP12b]. non-square-free [BP10]. non-symmetric [Per14]. non-zero [Qui11]. nonabsolutely [BGP11]. nonautonomous [AS12c]. nonbinary [Sri13]. noncentral [Hu10]. Noncommutative [ARZ11]. Noncommuting [DO11b]. nondefective [XD12]. nondegenerate [FKR11b, FKR12]. nonderogatory [Bot10b, FHS11, LS13a]. Nonexistence [MW14a]. nonhomogeneous [Puł11]. nonincreasing [DHKQ13, PT13b]. Nonlinear [BMS14a, FGR13, LN12, LLF13, LHL12, RS14c, SS12c, YZ10, YZ12b, BD12a, Bey12, FZ13, GHT11, Hil12b, Jar12, RY12b, WCKL13, Xu14, Lim13b]. Nonnegative [CN12a, FJ11, Gar12, MS11b, PRW11, Siv13, SDNS13, AG13, Aga14, Ben14a, BR14c, BK10, BAD09, BSST13, BS13b, CDP10, CS10a, CPZ13, CQYY13, CD13, CKAC14, CL12c, Drn13b, DZ13, Fri11, FGH13, GG12, GP13a, GL12b, Sem10, GGK^{+}13, GKR13, Hua11c, Hua13a, JLW11, KG12b, KSS12, Laf12, Lav10, LC10, NS12a, NT11b, PV12, Qi13, Sah10, Ser11, Spe11, VR12, Voy13, XZ14, ZCQ13]. Nonnegativity [HRT10, HRT13, PP11a]. nonpositive [CRU13, CRU14]. nonpositivity [HC10]. nonpowerful [JMS11]. nonreal [PR13b]. Nonsingular [BC12c, DdF14, HK13, HLZ12, BHZ10, CRU13, CP10, DQW13, FLH12, FJ14, HC10, LLN^{+}12, LS13d, LJY14, NPP13, SYH14]. Nonsingularity [Zuo10, BB13b, LS12c]. nonsymmetric [GL10c, KW12, LwCJL11, WXH10b, XD12]. nontrivial [Dod13, ZZ11, Zho11]. nonzero [CP10, GOSvdD14, HTW13, MB13, MZ14, PR13b, YHH12]. nonzeros [MGSW14]. Nordhaus [NY14]. Nordhaus-Gaddum [NY14]. Norm [BG12a, MM13, Pop12, Aud10b, BT14b, BJ11, CVW10, CFL13a, FCL10, FLC11, Gon11, Hia13, HMR12, Hua11c, Ji12a, KS14a, Kyr13, LMM11, LMYY11, MD13, MA12, MZ10, MTS11, NNW14, RR11, SP13, SR12, TGS14]. norm- [TGS14]. normal [AM13a, BFdP10, BFdP11, BFdP12, BDG13, BD13, BR11, Buj13, CH13, Chi13, CI10, Ger12, HLP12, KM13a, MSS14, yPjXL11, Sed11, SS13d, SS13e, Slo12b, Sze14a, Van10, WW13d]. normalizable [GOSvdD14]. Normalized [Bau12, AT14b, Boz13, CFK10a, CJ12, CL13a, LGSC14, vDO11]. normally [CM13]. normed [Wój14a]. norms [Aud13c, BH14b, BLS14, CGSCZ10, Fur12, GK14c, HK10, Lac13, Lan14, Lee10, LT12b, LT16, MW14b, Mor10, NY13, Pat10, ZH12]. Note [Ano11z, Ano12-28, BC14c, GL10b, HJL10, KP13, LLS10, LHL13, LHGL14, Rhe10, BI12, Ben14b, BC14b, BZ13, BZZ14a, CG13, Che14b, DA10, DD11d, DLLS10, EFN09, EFN10, FWW13, Fan10b, FH12b, FLC11, GOvdD14, GLS13b, Had13, Har14, HZM10, HZ12a, HCY10a, Ikr10, JM12a, Li10, LS12b, Lom13, MB13, Mat12, Mig13, NdM13, Sah10, Tao11, Wal14, WB11, XLG^{+}13, Yan11, dHM11]. Notes [Far11b, Fie10, Ha13]. notion [Han11b, Han14b, TN14]. notions [BBM14]. novel [Ma10a]. November [BBD^{+}11]. nowhere [BS14, FFS11a]. NSC [BBD^{+}11]. null [BDD14]. Nullity [BH11b, Bri13, CL11a, EHH^{+}12, FWW13, GFY10, GX12c, HS10, LFS12, Sin10c, TL13b]. nullspace [JKN14]. nullspace-type [JKN14]. Number [AiS13, BPA^{+}11, AG12, AKM13, AAJ12, AHAPP10, AdFM11, AHS10, BB13a, Bea12, BLL13, CRSS14, CS13a, CL14a, CFHL14, DTL11, Dom13, DH10b, DL13, DdF13b, DdF14, EHH^{+}12, FYI10, FG13c, Gar13, GB11, GLS12, HL11a, HJZ13, Hir10, HTW13, HJL10, LS10, LT11a, LZ12b, LS11b, LHL13, LYL13, LL14b, LJY14, MGSW14, MZ12c, NS13, NOL13, Row12a, Sin10c, SWT13, Ste10, TF10, Uhl13, Wik11, Wik12, XF11, YW11, YWS11b, ZHG13, Zhu12a]. numbers [AAK11, AL13a, BHKM13, CY11, CMS12a, EWY12, GWW14b, He13, KL12, LaG13, Lak10a, LJ11, LWY14, MGLW11, DDP13, WWG10, WZ14a, XD12, ZC14]. numeric [PS14b]. Numerical [Ali12, BS12a, CGWW13, DGH^{+}11, Gau10, Gle11, HMP^{+}11, Kam10, KI10, TW11b, Tsa11, AM13a, CG13, Che14b, CN10b, CN11a, CN11c, CN12b, CN13c, CP11, CLS10, DPF10, DD11d, GW13a, GTW13, GW14b, GP14, GZ13, HH12a, HB12, LMM11, LM12, Lee13a, LP12, LPS13, PLS14, PT13a, PGM^{+}11, VU14, WW13a, WC11, LCwCL11]. Numerically [FM12]. Nussbaum [Lim13b].

objective [GMS13]. objects [NP10]. oblique [ACG14, CM10a, Hua11a, SS10d, WL10a]. observability [KRvS12]. observable [BCdP11]. observations [Fie13, RvS13]. observers [Blu10]. occasion [Bai11]. occurrences [AM14]. October [BDK^{+}10]. octonion [RO10]. odd [CMNW14, GW11, GWZ13, HWG13, KS11c, LLN^{+}12, LGZ14, LWY10, Rim12, Yua14, Yua15]. off [BC14b, CFJKS13]. off-diagonal [BC14b, CFJKS13]. old [ES14, RTR10]. Olkin [Zha12a]. One [ALPV14, CRS14, GS12b, ABEV10, BBdH12, BBdH13, Bat14, BR14c, CT11a, CT11b, DHLX12, FGvRR13, Fri11, KO13, KY11, Lee13b, LFS12, LT11b, LX13, MMRR11, MMRR12, dSP12d, Pet10, PS14b, RW12, Sag11, Sav12, Slo12b, Suz13, WFM11, WMZ14]. One-bit [ALPV14]. one-dimensional [Suz13]. one-dominating-vertex [CT11b]. One-peak [GS12b, PS14b]. ones [TZ13b]. only [Aga14, CGMJ14, Dod13, wH13, RR11, WZ13d]. onto [Cha10, NN10]. open [HH12b, LGZ14, Zha12c]. operation [CMRR13]. operations [NN13, SL14]. Operator [CMS12b, Gon11, Kia14, Kim13d, Mos11, SP13, Seo14, AR10, AG10, ACG13a, Aud13b, BC14a, Böt13, BH14b, Chi13, CDDY10, CI14, DW11, DD11d, DX13, ET13, FMWW12, Fur10b, HC14, Han14a, HZ12a, HN10, IIK^{+}13, JG10, JR14, KS14a, KSAM12, KJK13, KW12, KLP12, MPP11, MTS11, MMK13, Nak13, OM10, PP14, San14, Uch10, WW10, WD10, WLL14, Wój14b, XSZ13, XSH14, ZWZ10, ZHQ14, Zha14c, dSC14]. Operators [Bra10, AS12a, AR10, And13, ACG13b, AK12b, BV12a, Bar10a, Bar13a, BS12a, BCD10, BR11, BC12a, BJ10a, BJ10b, BJ10c, BV13b, BB11b, Bud11, CRS14, Cam13, CL12a, CEM14, Che13, Den10, DWXS12, DCIW12a, DD11c, Dra12a, DP10, DJK12b, DDK14, DK14, DK13e, ES13, FJ10a, For14, GS10a, GN14, GP14, Sem10, Gu14, GS10e, HG11a, HKK^{+}12, HN14, HLW10b, HZ11a, Hua11a, HZJ12, JLLY10, KL12, KI10, KRS13, Kec13, KR10, KB12, Lac13, Lán10a, LP12, LCM13, LMMR13a, LMMR13b, Lom11, Lu11, MM13, Mat13a, Mat14a, MPP11, Mol11, MMM13, MN12b, Nie10, PP14, PY10, Pel14, Pep12, PT13a, Pop13b, Pro10, Rez13, RS12b, Ros12b, Rud12, ST10a, Sed11, Sei14, SR13b, Sha13a, Sha11, SM10a, SM10b, TD13, Tim14, VU14, XDFL10, XCS13, YW10, ZH11a]. operators [ZHQ13, dOHKS12]. Oppenheim [Lin14b]. Optimal [JZT14, LH11a, LJ12, LH10c, TDT13, Wil13, CGM^{+}10, EY13, GOSV12, GIP12, HCY10a, LHH13, LLD13, LL13d, MZ10, Mit11b, NP10, NA13, Peñ14, Sta14]. optimality [FMR12]. optimisation [GMS13]. Optimization [BH14a, MO14]. Optimizing [Gol13, LS11a]. optimum [CLL13a, GX12a]. options [PZVJ11]. Orbit [RMP14, Bar10a, Bar13a, Dau12, MZ12a, Rez13]. orbits [Baj14, CN12c, DD11a, LPS13, TT12]. Order [Bra10, FJ10b, Mol11, jASZ12, BP10, BM10, BL12, Blö12, Bos11, Bre14, BF14a, CR10a, CMNW14, CK13c, CL13b, CFL13b, CKAC14, CDM12, CIH11, DD10a, DTL11, DD11c, DM11, FK13, GOvdD14, Ghe14b, HSZ12, Hir10, HL10b, Kar11a, Kar11b, KM11, LLN^{+}12, LBLS12, LHZH11, LJY13, LGZ14, LS14, LZ11b, MD13, NS12a, Nem13, Nor14, QS14, Rim12, SS11b, SS11d, SBMT10, Slo14, ST13, Tra13, WZ13b, XDFL10, ZY14]. order- [Nor14]. order-preserving [CFL13b]. ordered [Car13]. Ordering [AJRT14, HL11a, WT12, JP11, iO12, RMAJ10, WL10b]. orderings [Nie12]. orders [BJZ12, FJ14, KS11b, YHH12]. Ordinary [CM14]. organization [Zha12c]. organizing [Gle11, Kam10]. orientations [CEY14, Tia11a]. Oriented [Zhu12c, CLL13a, CLL13b, GX12b, GHW^{+}13, HSS10, RR12, WZY14, XG13]. origin [LSR11]. Orlicz [AM13c, ARZ11]. Orthogonal [CMS12a, Moo11, AAH13, BR14b, BSS10, BMM12, BR12, CM11b, Dax10, DW10, FPC13, GW11, GWZ13, Hür13, KN10, LRV12, LAL11, LWGM12, LNT13, MPRW11, Mer12, MNZ12, NM14, O’D14, dSP12d, Pop13b, RBP12, RL13a, ySpW12, TT11, VS13, VS14b, Wei10, dF10, dlCdlRMP14, AMP10]. Orthogonality [Gro14, AR12, CP11, HH13]. orthogonalization [OM10]. Oscillation [Hil12b]. oscillators [VR12]. Ostrowski [Had12, Had13, HT14, Tao11]. outer [Dra12b, HZJ12]. outerplanar [Sin10c, SK12]. output [BO12]. Oxford [Bar10b, Gar12].

P [DdF13b]. P-vertices [DdF13b]. p.p [LZ12a]. Padé [DD11b, GIP12]. PageRank [GPR13, TD11]. pages [Rei11a]. pair [BBdH12, BBdH13, Bar10a, BCdP11, BC12a, DHLX12, DKS13b, Hwa12, NT10a, dOHKS12]. Pairs [Cal12, FdC10, AW10, BM12a, Bar13a, BK11b, BC14a, CGSCZ10, DS14, FRS14, God10, God12, GTV12, Han11b, Han13b, Han14b, HG11b, HWG14, INT11, IS14, JG10, Lim10, NT10b, NT12b, Nom14, Pet10, dCF12]. Pairwise [LD12, Tra13, ACM^{+}12, EvdD10]. Pálfia [Seo14]. Palindromic [IM11, BM14a, GN13, LCwCL11]. Pall [BFdP10]. Paperback [Bar10b, Tam12]. paperfolding [GWW14b]. Parabolic [HS12b]. paraboloids [KK12, KK13]. paracomplete [CM13]. paracontractions [Moj14]. PARAFAC [ZHZF13]. parallelogram [MTS11]. parallelotopes [GK10]. Parameter [GD11b, PT13b, BR14b, Hil12b, HMP12, Kar11b, MP10, VW10b]. parameterization [CR10b]. parameterized [Gna12, KLL11, YWX13]. parameters [BBF^{+}10, DP12a, DZX12, HHMS10, Han13b, HB12, HL11c, JLLY10, Zhu12b]. Parametric [He13]. parametrization [BO12, Dau12]. parametrized [Fuj10, GHT11]. paranormal [DJK12b]. Paratransitive [LMMR13a, LMMR13b]. Parity [MS13b]. Parseval [Kon13]. part [Li10, AAT12a]. Parter [FdC14b]. Partial [BP13, HZ12b, MQ11, MQ13, MQ14, RMT11, WZ13b, AM13d, BFRR14, BCY12, BO11, CFLW13, DM11, Fan10a, GW14a, GW14b, Kho12, MD13, PLS14, PP12a, Rub13, SvdH11, Zha12c, vdH13]. particular [BBdH12, BDM^{+}12, GOvdD14, ŠŠ11e]. partite [Zha14a, ZWL13]. partition [FdC14a, HM10, NS11a]. partitioned [BLL13, CGMS10a, CTW11, XCS13]. partitioning [Dru14]. partitions [CD10, Car13, GL10b, MS13b]. Pascal [DK12b, Kau12, Kim11b, VS11]. passage [NS11b]. paste [ST10b]. Pate [Pat12a]. Path [Est12b, Fuj11, HM14b, Nak13, SS14, Sin10c]. paths [AS12c, Gum13, Nik10b, QSW14]. pattern [AT11, AHL^{+}14, AM14, BDH^{+}12, CFJKS13, CL10b, DT11, HL11d, JMS11, LSR11, MGSW14, PP11a, WLHL10, YS13, ZLH^{+}14]. Patterns [BKMS13, BVV12, BDM^{+}12, BFH^{+}12, CP10, EKSV14, EKSV18, GS11a, GLZ14, GS12a, GS13b, GOSvdD14, GOvdD14, GK14b, GB14, GOvdD12, GS12f, HJN12, HLS10, Hua11b, MZ14, Ma14, Mit11b, OTdDv12, YHH12, dS12a]. Paved [NT14]. payoff [AGK11]. peak [GS12b, PS14b]. Peano [ABGPSS14]. peculiar [RR14]. pedagogical [Kla10]. Pellet [Mel13]. pencil [Dod13, GHMPVP11, XD12]. Pencils [RS12a, RS14b, Bat14, BIT12, BCF14, BF14c, DKS13b, Dod10, DS14, FRS14, LLB13, MSP11, Rei11b, SBM11]. pendant [GFY10, LWZ11, LZ12b, LJY14, Suz13, XZ13c]. pendent [BNP11, BNP13, HJL10, NOL13]. Penrose [Boz13, xCwXL11, HF12, HZ11a, Ji12b, KS12a, MZ10, Nor11, Pat12b, RDD14, WJT13, XCS13, Yan14, ZZCW13]. pentadiagonal [Elo11]. Perfect [BP10, CG11, Beh13, LSC10, wTmS12]. Perfectness [FPC13]. perform [CHK^{+}13]. Periodic [AS12c, BM10, Li12, ZLH^{+}14, BdFdP11, BT13, Cal12, Fid10, SB12, Tsa11, Xu12, YM12]. peripheral [ZH11a]. permanent [Cra13, DdC13, Zha13, dF10]. permanental [AdF11, LZ13]. permanents [Brä12, CW12a, FH12b, PSW11]. permissible [Buj13]. permutahedra [Dah10]. Permutation [DF14, LS13c, RR14, SMC11]. Permutation-like [DF14]. permutations [AK11, BF14b, ST12]. Perron [Lim13b, Czo10, FGH13, FJMP14, JMP12, LN12, NV12, Pit11, ZH11b]. personalized [GPR13, Gle11, Kam10]. Perturbation [BV12a, BK11b, CM13, HZGY12, LNTgW12, MMRR12, AL13c, BI12, BBdH12, BBdH13, xCwXL11, CS10a, DD10a, DX13, Guo10b, Jai11, LS11a, LYS13, Lip10, MMRR11, MZ10, MA10b, Vul12, WF12, WL10a, YW10]. Perturbations [CGM11, CL14b, HZ11a, Bat14, BCdP11, BHKL10, BdS10, CL12a, DvDF11, FGR13, FGvRR13, Hua11a, HZJ12, MMRR11, MMRR12, RW12, Rod12b, SWA12]. perturbed [BDG13, BR14c, CGMS10b, GC12, GTV12, HH11a, Mat14b, PPZ14]. Pete [Bai11]. Petz [Fuj10]. Pfaffian [Buc10, TT10]. Pfaffians [Ika11]. Phase [FMNW14, Byd10, Gul11]. phenomenon [BL11, RR14]. Pick [AL13b, CH11]. Piecewise [FGQ11, CA10]. piling [OT12]. Pinkus [Gar10]. Pisa [BBG^{+}13]. pivot [BH11b, Bri13, KB14a, KB14b, Mit11b]. pivoted [LQ11]. pivoting [VS14a, PQZ13, Raf14]. placement [BO12, RMT11]. planar [BHMO13]. plane [Buc10, Buj13, CM12b, HK13, MZ11, dlP11]. planes [Dau12]. planning [DT10]. player [Jun14]. plus [AGNS11, BH12, BM13d, Mer10, Mys12, MP14b]. Poincaré [BFdP10]. point [ACG14, CLCL12, CI14, DYW14, DD14, GOSV12, JH10, LWGM10, LWGM12, Per14, SS13b, Ter13, Wan11b, XX14, XSS13, ZWZ10]. point-stabilizer [LWGM10]. points [AK12a, AR10, BS11d, CN13b, GX12c, Li12, Pan12a, QH10, WLL14, Wu13a, ZZ11, ZZW10, ZZ10, ZZ12, ZXZ10]. Poisson [Mar13a, XX14]. Polar [CM10a, AGK11, CD10, CM11b, DoMP09, DMP11, GMP13, LNN14, LWGM10, LWGM12, MGLW11, MPP10, WLGM11, Wor13]. pole [BO12, KBS13, RMT11]. Poloni [Lim11a]. Pólya [Seo13]. polygon [VW10b]. polyhedral [LT10b, MS11b]. Polynomial [AM13a, LMM11, LM12, MM10c, Mar13b, AA11, AB12, ABSV12, BMS10, BN13, BC14b, BO11, BH11b, BHAP12, Cer10, CL13b, DTL11, FH10c, Gna12, GX12b, GS12c, HB12, HP11, JKV13, KL13a, KL12, KH13, KdM13, KW13, Lee13b, LD12, LW13b, LM10c, MZ13, MW14a, Mel14, Moo11, PQZC11, Pet10, PT14, RO10, SK14, Sun13, TW14, VS14b, Wu10b, ZYL10, dCF12, vdW14]. polynomial-Vandermonde [ZYL10]. Polynomials [NT12a, AAK11, AK12a, Aga14, AAT12a, AAT12b, AL14, Bal14, BM14a, BR14b, BK11b, BG12b, BL10a, BSS10, BH13a, BMM12, BR12, BW13b, CMS12a, CKS10, CN10a, CN11b, CL11b, Cim11, CD12c, DD10a, DDM12, DTS11, DW10, DD10b, ES11, FPC13, GMMFPSS12, GW13a, GN13, GS12d, GM11b, He14, HLW14, Hen10, HT10a, Hwa11, IM11, Kal13a, KHKT13, KH13, LT12a, Lan14, Lav10, LL10a, LZ13, MMMM10, MMMM13, MZ12a, MZ11, Mei13, Mel13, NV10, NT12b, Psa12, QY12, Qua12, RBP12, RL13a, Sha10a, Sim10, TTZ13, DDP14, DDM14, TGM11, TT11, VS13, WLLX11, WML13, WZ13d, dF10, dO12]. polytope [CGSCZ10, CM10b, KAAK11, PSW11]. Polytopes [Dah11, ACDM14, Bar12a, Beh13, BG12b, BW13a, Dah12a]. polyvectors [De 11]. Poncelet [Mir10, Mir12]. Pontryagin [CD12c, Wor14, dSW12]. pooling [LHG10]. population [LLR14, MSvdD14, RM14]. porism [Mir12]. Porta [CMS12b]. portfolio [DBZZ14]. Portuguese [FdC12]. posed [BJRS11, HMR12, NRS12]. poset [Kha13]. posets [Dor10, GS12b, PS14b, SY12, Sim10]. position [AY13, FFG^{+}11]. Positive [AG10, BKV14, BLL13, Dru14, EEH^{+}13, Fri11, Fur10b, Gar10, LV14, MYL13, Pop10, SH10, WD10, AHAPP10, AL13b, ACG13b, BH14a, BMW10, BS12c, BP11, BS12d, BR14c, BI13, BSU14, BCS13b, BLL12, BL14, BAD09, CRU10, Cen11, CM12a, DA10, DP10, FFJM14, FJ10a, FV13b, FGvRR13, FW14b, Fuj10, Fur11, GS12b, GD11a, Gna12, GST13, GR12, HP12b, HKPR13, HJN12, HN14, HS12c, ISYY11, Kak10, KSAM12, KS11c, Lac13, LLY11, LL13b, LT11a, LW13a, LLB13, Lim11b, Lim12, Lim14, MWZ13, MPS10, MW14b, Mat14b, MPT14, MY14, Mol11, MR10c, OTdDv12, PP14, dSP10d, Pep12, PV12, QS14, Ros12b, SSMS14, Tan10a, Voy13, WXH10b, Yam13, YFW13, YJ12, ZCKS12, Zha14b, Zim13, vdW14]. positive-definite [CM12a]. Positive-kernel [BKV14]. Positivity [BMR11, BP14, FF10, GK14c, HH11a, JN13, JS13b, KS12a, Kus12, Peñ14]. Positivstellensätze [Cim11, SS12b]. post [BDV12]. post-Lie [BDV12]. posteriori [Cha12]. Potapov [FKR11a, FKR12]. potent [CLST14, DCIW12a, LRT12, LRT13, Rom14, SR13a]. Potentially [BVV12, GOvdD12]. Power [GW14a, AH13b, Für10a, GW14b, HQS13, JMS11, LY13, PP12a, SS11b, Seo14, SHS12, TL13a, WF14]. powerful [LLH10, LYL13, YW11, ZLH^{+}14]. powers [BLdPS10, Bie14, CL13b, DXL13, GW10, GKR13, Jai11, JLW11, Pat10, Rim12, SS10b, WWG10, WZ13d, HOT13]. pp [Bar10b, Gar12, Grü12, Lim13b, Rod12a, Tam12, Zha12a]. PPT [Lin14a]. pre [BM13c, Pop12]. pre-Hilbert [Pop12]. pre-Jordan [BM13c]. preconditioned [Jbi10]. preconditioner [CJ10, DKOT12, GOSV12, Raf14]. preconditioners [JZT14, LZ11a, PZVJ11, TDT13]. Preconditioning [BTYZ12, PIM^{+}10, SS13b, Tre11, TD11, Wan11b]. predetermined [MVPS10]. predictable [KS11a]. Preface [ACGN10, ACT14, Ano10u, Ano11-27, Ano12-29, Ano12-30, BBD^{+}11, BBSMT13, BBG^{+}13, BFBD13, BBD^{+}13, BGL^{+}11, CFPT12, EGLR14, FKLT13, KPRT14, BFH^{+}12]. Preliminary [Ara12]. Preprocessing [PQ10, PQZC11, PQZ13]. Prescribed [Kra12, Bar12b, BCdP11, BMSW10, BCŠ10, BC12b, BC14b, Fri11, LL10e, MS13c, NR10, Psa12, RS12a, RS14b, SDNS13, WL10b, YT13b]. presence [LS13a]. preserve [Cos11, FdC10]. preserved [DV10]. preserver [BMGMC12, RŠ10]. Preservers [BSK12, Ben13, FHL^{+}11, HHL10, AMJ14, BEM12a, BEM12b, BSKL13, CD12a, Cos14, CFLY12, CLHQ14, KSB12, LPS13, LT11b, dSP10e, dSP12b, Per11, RdSP11, SA14]. preserves [Hua12b]. preserving [AA11, ABSV12, BB13c, BCS13b, Bou10b, BMS14a, CFL13b, CLS10, CL14b, DD10b, GN14, HLW10b, HQ13, lH10a, lH11b, HH11b, JG10, LL12, LP10, LP12, LLF13, LHL12, Lim10, LL10b, Liu14b, MS10c, PP11b, Rod12b, Ros12b, TGM11, ZH11a]. Press [Bar10b, Gar12, Gle11, Grü12, Lim13b, Rei11a, Rod12a]. Pretty [FG13a]. pricing [PZVJ11]. prime [DF14, DW14, FK13, Hir10, LL11b, LHL12, LL10b, QCH11, SS11b, Sbu10]. Primitive [AB12, CL10b, YW11, CL10a, HL10c, Kim10, Kim11a, KP12, Kim13b, LLH10, LYL13, PPK13, WLHL10, YHY14]. primitivity [Sch11, BM14b]. Princeton [Bar10b, Gar12, Gle11, Rei11a, Rod12a]. Principal [JS13a, BBC^{+}14, BFdP11, BR14c, Bri13, BDOvdD12, FG13b, Guo10a, JN13, KB14a, KB14b, LS12d, OTdDv12, Sev13, TH11]. Principally [SS13e]. principle [Car11, Kra13, Lom13, MW12b, Puł11]. Principles [Gre13, CPZ13, GJ11, LLB13]. probabilities [Ber13b, VV13]. Probability [CR10b, LHH13, DD11d, GY13, OM14]. problem [AW13a, BBdH13, BV11, BM14a, BdFdP11, BCEM12, BBGM12, BCŠ10, Byd10, CPV10, CI10, CI13, DS11, Dod10, DS12c, DS14, Fan10b, FKR11a, GS10a, GEP13, GP12, HH12a, HRW99, HMP12, HLS11, HP04, JJKS11, JM12a, JDY13, LXL^{+}14, LCwCL11, LJ11, LWY14, LGZ14, Mou12, MAM^{+}13, MP10, NS12a, NM14, NS11b, Nik10a, OM10, PH12, yPjXL11, PRT13, SS11a, SS14, SS11c, SR12, ŠŠ11e, SMC11, TD11, VS10, Vul12, Wei13b, WJ12, XLG^{+}13, YWX13, Zha12c, ZY12b, vdH13, LPK14]. Problems [MNZ10, AA14, AS10, jASZ12, jAS13, ACG14, Bal10, BBF^{+}10, BFdP12, BT13, BdP13, Bey12, BN13, BC12b, BJRS11, Brä12, BMGMC12, CRS14, Dai11, DS13, DKOT12, DYW14, GMS12, GEP10, GJTP13, HMR12, Hog10, HWH14, Jar12, JV11, JRMFSS12, JW13, KKR11, KZ11, LT12a, LRST10, fLyH11, LZ11a, LL10d, LXK12, MS10b, MR10a, Miy14, Nak10, NRS12, NY14, Pin11, PPKR12, PE13, RŠ10, SVP11, SHZ10, SB11, Stu12, TmYsH11, Wan11b, Wei13a, wXL14, wXZ19, XE11, XSS13, YZ11]. procedure [LBLS12, SV11]. procedures [Pál13]. Proceedings [Ano10u, BBG^{+}13, BFBD13, Ano12-30]. process [BFS11, FFS11b, GMS13, PPZ14, VV13, Vas14]. processes [KS10, MAGR13]. Procrustes [fLyH11]. produce [Tra13]. Product [PGM^{+}11, AR10, Ali12, Aud10a, BM13c, BZZ^{+}14b, Cal12, Cha13b, CG13, Che14b, CJK^{+}13, CT14b, Dug12, EM11, FF10, Gha13, GK14a, GK12, HFS13, HTS14, HMT10, HLW10b, HP11, KSH12, KY14, tLyLWqW10, LPS13, LLF13, LD11, Lin14b, LH10b, Pat12a, dSP11b, RAAGAVS11, SP13, Sha13b, Sha11, Tad12, WLG12, ZCWZ13]. product-set [Cal12]. Products [ACG13b, BMN13b, Bot10a, CM11b, SZ14, AM13b, BG14b, CSV14, CLHQ14, DKS10, DD11c, DD14, Ens10, FHL^{+}11, FJ14, GKZ11, HM14b, Hua11c, JG10, Kim12, KN13a, KN13c, KLZ14a, Koz14b, Kub13, LW12e, MARC13, Pep12, PV12, QCH11, SCSS10, Slo13, Slo14, Tan14b, Voy13, ZH11a, ZZCW13]. Professor [Lin11]. profiles [CEY14]. programming [AHAPP10, GS12b, GMH14a, LLD13, LL13d]. programs [Sag11]. project [NN10]. projection [Cha10, Hua11a, LRV12, LB14, NN10, NN13, PPKR12]. Projections [BJ11, ACG13b, ACG14, BS10, CSC13, CM10a, CM11b, Ili10b, JG10, SS10d, WL10a, ZZCW13]. Projective [BH10, BEV13, De 11, DTL11, lH11b, LSV12, Mer10]. projectors [HRT13]. Prony [PT13b]. Prony-like [PT13b]. Proof [Das11, Das13, JS13a, ZW12a, BL13a, CP12, Cho13, Dom10, FGG10, GS12a, Hil14c, Kak10, KN13c, Kon13, LYL13, Rho10, Yam13, ZY12a, vdW14]. proofs [CHZ13, Dra14, FF10, Sat11, Sat14]. proper [MS12, TG13]. Properties [ANF11, CFLY12, DCIW12b, HM14a, ÁNPQ12, ACG13a, BHDW12, BCD10, BB14, BJ10c, BLS14, CPV10, CJ11, Dax10, Den10, DJK12b, DGMS10, Far11b, FPC13, FH10a, FH13, FJ10b, GB13, GM11a, GM11b, HG11b, HL11d, Hun10, Hun14, ISYY11, Jun12, KKL13c, Kus13, LaG12, LZ13, MŠ10a, Nor12, Ref12, Rod11, Rod12b, ST10b, SSZ13, SSR13, Sto11, Tre10, Tre12, WC11, Zho12]. property [Aud13b, Bai14, Bal10, Ben10, CL12a, CGSCZ10, DK14, GTR12, GLS13a, JV10, JV11, KP14a, KS11a, Lán10a, LNN14, Mor10, Mou12, NP13b, PP12b, Puł11, Slo12b, Tao13, TGS14, AGPPF12, LV14, UW11]. protocol [AIP12]. PSD [LMT10]. Pseudo [Fio12, WC12, BdP13, Kni14, Wik11, Wik12]. pseudo-determinants [Kni14]. Pseudo-distance-regularized [Fio12]. pseudo-random [Wik11, Wik12]. pseudo-symmetric [BdP13]. pseudographs [LL14d]. Pseudospectra [RR11, AGV12]. pseudospectral [ABBO11]. pseudospectrum [CFLY12]. Pták [Ser13]. PU [XJ10]. Publications [Ano13e, Ano13f]. Pure [BFBD13, HQ13, SK14]. Putnam [DK14]. pyramid [BPDC14]. Python [GS12b].

Q [Tao13]. QR [BBE^{+}10, LYS13]. QRacah [wH12]. QTT [KRS13, Sav12]. QTT-rank-one [Sav12]. Quadratic [GKS^{+}10, KD12, Pol13, BCY12, DS11, Den10, EY13, GS10a, GLP^{+}13, GS12b, Gre12, HMP12, Hu10, JV10, Jun14, KS11b, LT12a, MPS10, MR10a, Mei13, MP10, NT11a, dSP12c, Pop10, Sha13a, Tia12, TGM11, XSZ13, Zha12d]. Quadrature [BB10, Rei11a, BFRR14]. quadric [GS12e]. quadrics [Baj14, BS11d]. quadriculated [ST10b]. quantized [Han11a]. Quantum [Dug11, Wei11, BC14a, CHLW14, GZ13, HWSH13, Jen13, KLS12, SL14, Wor13]. quantum-trace [KLS12]. Quasi [aCCS14, CRU13, MS14a, Wu13b, BM10, BFdP12, BIT12, BDFP11, CRU14, DS14, DJK12a, ES13, ET13, Fuj11, GLS10, HTS14, HF12, Kim12, iO12, PLL12, WLHL10, WLL14, XM11, LBLS12]. Quasi- [CRU13, GLS10, CRU14]. quasi-arithmetic [WLL14]. Quasi-Arnoldi [LBLS12]. quasi-Banach [BDFP11]. quasi-class [DJK12a]. quasi-inverses [HTS14]. Quasi-isometries [MS14a]. quasi-Jacobi [BFdP12]. quasi-Kronecker [HF12]. quasi-linear [BM10]. quasi-means [Fuj11]. quasi-primitive [WLHL10]. quasi-regular [DS14]. quasi-thin [Kim12]. quasi-tree [XM11]. quasi-trees [PLL12]. quasi-triangular [ES13]. quasi-Weierstraß [BIT12]. Quasi-Whittaker [aCCS14]. quasideterminants [Abe11]. Quasioptimality [Sav14]. quasiseparable [BOZ10, BOZ11a, BOZ11b, OM12]. quaternary [BE10]. quaternion [Kyr13]. Quaternionic [Kim13c, KAAK11]. question [Cha13b, CL14a, Dru12a, Qui11]. questions [CL14b, Shp10]. quiver [Stu12]. Quivers [Sev10, GS12b, Lop11a]. quotient [BO12, Bot14, LL10d]. quotients [Dax10].

R [Gar12, Gem10]. R. [Vse12]. Racah [BC14a, GWH13, HWG14, NT10b]. racetrack [SS10c]. Radial [Sri13, CGR14, MZ11]. radii [CLS10, FKR11b, LSC11, LL11c, ST13, XZ13b, XG13, YWS11a, YWS11b]. radius [Alt13, Aud10a, BMSW10, BMSW11, BL10b, BNP11, BNP12, BNP13, BC14c, BS13b, CT10, CLS13, CKST11, CvDKL10, CTG13, DHLX12, Dai12, DP10, DZ13, Dum13a, FYI10, FN10, GLS10, GL12a, GP14, GR10, GMV11, GLS12, GL12c, GLS13b, HJZ13, HY14, Hua11c, KI10, Koz10, Koz14a, LLS12, LL13a, LS10, LT11a, LWZ11, LWV12, LP12, LS11b, LCZ10, LL10e, LX13, aLwW13, LLT12, MR14b, MP13a, NP12, Nik10b, NOL13, NLL13, OM13, OM14, Pep11, Pep12, SWT13, WXH10a, WB12, WZ13c, XZD14, XZ14, YFW10, ZW12a, Zha12b, ZHG13, Zho10, Zhu10a]. radix [Dum13a]. radix-rational [Dum13a]. Rado [CP12, OLW14]. Rado-Horn [CP12]. ramified [MM10a]. Ramsey [ZC14]. Randić [CFK10a, GMRS14, GFB14]. Random [Bod13, LAL11, Bos11, CLL13b, DLLS10, DLL11, PS12, Wik11, Wik12]. randomization [PQ12]. Randomized [PQ10, PQZ13, GO12, NT14]. range [Ali12, BT11a, CG13, Che14b, CN10b, CN11c, CN13c, DWXS12, DDK14, DGH^{+}11, For14, GH11, HH12a, KI10, LPS13, NRS12, PT13a, PGM^{+}11, Sha11, SL14, VU14, WC11]. ranges [ACG13a, BS12a, CGWW13, CN11a, CN12b, CP11, CLS10, DD11d, Gau10, GW13a, GTW13, GW14b, GZ13, HMP^{+}11, Lee13a, TW11b, Tsa11, WW13a]. Rank [BCS13b, DP12b, FGvRR13, IW13, LT11b, LX13, MNZ10, iO12, AAF^{+}12, ABEV10, AHL^{+}14, BGV12, BT10, Bal14, BBF^{+}10, BBF^{+}12, BKM^{+}13, BBC^{+}14, Bar10a, Bar13a, Bat14, BMN^{+}13a, Bea12, BDG13, BR14c, BSU14, BC13, BdS10, BH13b, BRLS12, BHZ10, BDOvdD12, BKMS12, BZZ^{+}14b, CT14a, CRU14, CR10b, CAV13, Cau11, CHY11, CHY12, CN11a, CN12b, CLS10, DHLX12, DoMP09, DGH^{+}10, Dea11, DS14, DBZZ14, DGMS10, EHH^{+}12, FM11a, FP13, FdC14a, Fra13, FL12, Fri13, GW13a, GG12, GS12f, HHMS10, Hog10, HCY10b, HCY10a, HZ11b, HTW13, IMA10, JKV13, KG12b, KRS13, Kim11a, Lee13b, LT10a, LC10, LTX14, MAGR13, MS13a, MR13, MR14a, MZ12b, MQ13, MQ14, MMRR11, MMRR12, Mit11a, MNZ12, dSP11a, dSP12a, Per11, Pin12, RW12, Sad12, Sag11, Sav12, SM13, SHS12, Shi11, Shi12c, SvdH11]. rank [SK12, SC10, Tia11b, Tra13, TDT13, WFM11, WZL13, WMZ14, Zha12d, Zim13]. rank- [BdS10, Fra13]. rank-1 [SC10]. Rank-one [LX13, Bat14, DHLX12, Sag11]. Rank-preserving [BCS13b]. Rank-width [iO12]. Ranking [BEK13, Dah12b, Tra13]. Ranks [BL13b, BSK12, BSKL13, Ber13b, BBM14, CHLW14, Fri12, JZ11, KSB12, MQ11, Shi12b, Shi13, SSM13, Tia12]. Raphael [Ano13-27]. rate [CNPP12, EY13]. rates [GL10c]. ratio [EJLS11]. Rational [BKV14, Duk12, Duk15, Lom11, AL13b, AIL12, Baj14, BT11b, Boj13, Bre14, Dum13a, MR10b, MR14c, Rad13]. rationals [HNZ12]. ratios [BRA11]. Ray [LS12c, GS11a, JMS11, LSR11, LS13d, MGSW14, ZLH^{+}14]. Rayleigh [LL10d, SS12c]. rays [BF14b, Hil12a, MZ11]. rd [CKAC14]. Re [NCI13]. Re-nnd [NCI13]. reachability [BLLM13]. reachable [SS10a]. real [Aga14, AKM13, AAT12a, Bal14, BGP13, BH13a, BDOvdD12, Bün14, Cau11, CGSCZ10, Coh14, Dai13, FGvRR13, Gre12, He14, Ikr10, Kal13a, Lav10, MS10b, MSS12, PR13b, RdSP11, Sag13, SYH14, WZ13d, Wód14, XZ13a]. real-nonreal [PR13b]. realisable [ES14]. realizability [CL14b, DHLX12]. realization [OZ10]. Realizations [BCdP11, BKV14]. realized [FW14b]. reals [KN10]. Rearrangement [TCL11]. reasoning [DD10c]. receptance [RMT11]. Recht [CMS12b]. reciprocal [DKOT12, KP14a, LaG12, NP13b]. reciprocals [GIP12]. reciprocity [SR13a]. Reconstructing [JMW11]. reconstruction [GHMPVP11, MRS12, Xu12]. recovered [Lee13c]. recoveries [JKN14]. recovery [DPF10, NNW14, PS14a]. recruitment [DT10]. rectangular [CRU10, CRU14, Coh14, DDM12, GY13, MRW11, ZCQ13]. Recurrence [LM10c, VS14b, lY12, Góm10, Lin10]. Recursion [JR11, BHAP12]. reduced [Hwa12, TW10a]. reducibility [GTW13]. reducible [Gau10, Kar11a, ZLH^{+}14]. reducing [DFS14, LHZH11, MAGR13]. Reduction [CN11c, KS11b, BGW12, BDG13, Blö12, Cha10, KMS13, LS13a, LRV12, LBLS12, MAGR13]. reductions [BW12]. redundant [Fou14]. refined [BDM^{+}12, GLZ14, GOvdD14, YHH12]. refinement [BK11b]. refinements [KMSC14]. reflection [MV12]. Reflexive [KS14c, DH12a, RRM11]. Reflexivity [Rud12, BJ10a, HW14a]. Regarding [Wal11b]. region [CKAC14, LSR11]. regression [ATS12, HLP12, LJ12, YiKIS12]. Regular [CD10, CLL13a, CCL14, GX12a, Han14a, KS13b, Kol13, Kub13, LS14, QSW14, Row12b, AAJ12, AKA13, Ban13a, BP12, Bat14, BIT12, BZ12a, CvDKP13, CSZ10, CR10a, CR10c, Cer10, CMNW14, CS13b, Cio10a, Cio10b, CW12b, CP10, DFG10, DvDF11, DS14, ES13, FGG10, Fio12, FP11, FG13c, GWW14b, GM14, HM10, HL10b, HLZ12, Hua12b, HLZP13, Hua13b, KS14c, KY11, KPY11, LZ10, LLN^{+}12, Lee13b, LwW14, LLS10, LGZ14, MM10a, MSP11, MW14c, NS12c, NS12b, RTR10, SS10a, SSS13, SS13a, Ste11, WML13, vDF14]. regularity [CN10d, EdlP14, GZ12, SBM11]. regularization [HRT10, LRV12, SS10c]. regularized [Fio12, XSS13, LB14, LKN13]. Regularizing [FRS14]. regulators [EY13]. related [Ano12-30, BG12a, CD14, Car10a, CMS12b, Dah11, Dru14, Duk12, Duk15, GMS12, HW11, HP12b, JRMFSS12, KSAM12, KP14b, LRST10, Maz10, MPSS10, MTS11, Naj13, Nak13, Rah13, RO10, Sei14, SHZ10, Tim14, TT11, ZYL10]. relating [Ste10]. relation [CFJKS13, DdC13, GB13, JV10, LSH12, Yan10a]. Relations [LRT13, AK12a, ADW13, BR12, CM13, DW10, JR11, KP14b, LM10c, Mar11, Mar13c, Tim14, Wój14a, lY12]. Relationship [CN10a]. Relationships [HRT13, MS12]. Relative [BMM12, Sha10a, IIK^{+}13, Koz14b, MS10c, Nak10, Nie11]. Relaxed [FFS11b, Nak12, SK14]. Remark [SK13]. Remarks [Lu12, PHS13, Rod11, dOHKS12, YY14c]. removal [MNZ12]. removed [LWV12]. Rényi [IIK^{+}13]. repetition [BCF14, BF14c]. replicated [PYZ14]. replication [AT14b]. report [BFH^{+}12]. Representation [BY11, OM12, YZ13, Ben14a, BD13, BBS12a, CM12b, CN11b, DMS10, DKM^{+}14, Irv12, KAAK11, Kyr13, Qua12, Sag13]. Representations [DW11, XCS13, XSH14, jASZ12, jAS13, BY11, BE10, Buc10, CN12c, Den11, DD10c, Dor10, Han11a, IM11, Kaw13, LW12d, LZ11b, Ma11, Net10, NT12a, Qua10, RMP14, Ros12c, SY12, Sze14b, VS11, XSZ13, Yan14, GMT13]. represented [GN13]. Reprint [BOZ11b, Mar13c]. Reproducing [Wor14, Che13, CD12c, SST14]. require [BDM^{+}12, GOvdD14]. Required [CP10]. residual [CI14, KOR14, Lin10, ZWZ10]. residuations [GMH14b]. Resistance [BYZZ14, BS11a, ESV^{+}11]. Resolution [Ste10]. resolvent [BMW10, EFN09, EFN10, JZZ13, KL13b, Sah10]. Resource [Cec10, Wil11]. resources [Cec10]. respect [FZ11, ISYY11, KKR11, Wu10b]. respective [Köh14]. restarted [MV12, MSS12]. Restricted [Oto12, BT14a, Cha10, NRS12, Shp10]. restrictions [Wój14b]. result [Cam13]. Resultant [ER13, Rue13]. resultants [BL10a]. resulted [XX14]. resulting [Cha10]. Results [Tam12, AT14a, BG12b, CW12a, CvDKL10, CI13, DMS10, GL12c, LJ11, LKN13, LW13b, MH13a, MK12, MD10, Mos13, Naj13, PAS11a, RTR10]. Retaining [GR12]. retrieval [FMNW14]. Reverse [CIH11, BLdP12, DD11c, KSA11]. Reversible [GP13b]. reversion [Boj13]. Review [Bar10b, Gar10, Gar12, Gle11, Gre13, Grü12, Lim13b, Rei11a, Rod12a, Sla10, Tam12, Zha12a, Lan13]. reviewers [Ano12c, Ano11a]. revisited [CTW11, EM11, FHM13, Qua10]. Rheinboldt [Rua12]. Riccati [BJ13, GL10c, Guo13, LwCJL11, Per14]. Richardson [AW10, CDP10]. ridge [LJ12]. Riemannian [HP12b, Yam13]. right [KLZ10, LT12b, LT16]. rigid [BCS13a, Hen10]. rigidity [AN13, LV14]. ring [CGRVC13, DW14, KLZ14b, Pet10, TZ12b]. rings [Ada14, AKN12, ÁvW11, ÁvW13, AW10, BE12, BCDM13, CRS14, DWW14, DW14, Ere13, Gho13, Gre12, LaG12, Lak10b, LZ10, LD12, LZ12a, LHL12, LL10b, Liu14a, Mar10, MSvW12, RDD14, SS10a, SSS13, TZ12b, TZ13a, Wan11a, WC12, WW13b, ZZCW13]. Riordan [AMPT13, BH12, CJ11, CJL13, CK13c, He11, JLN13, LM10c, LMMS12, WZ14a]. RIPless [KG14]. risk [VV13, Vas14]. Ritz [Buj13, CH13]. Robust [KKB11, SBM11, FM12]. Robustness [MS14c, NRR^{+}11, CKL^{+}13, Fou14, MMP13b, MP13b, MP14b]. Roger [Lim13b]. Roichman [Alo14]. ROM [Hür13]. root [Bal12a, Bal12b, DHS10, FHS14b, FJMP14, Guo10a, JMP12, KAAK11, Lin10, PQZC11, TNP12, Vla12]. root-finding [PQZC11]. rooted [RM10]. Roots [GS12d, Abe11, BLS14, DKS13a, GS12b, HMP^{+}11, He14, HL11b, MPT14, MR12, DDP14]. rose [LH13]. rotation [GTV12]. Rotfel’d [Lee10, Lee11]. Rothblum [Loe12, BMS14b]. roulette [CN12b]. routing [BL11]. Row [AAM12, Bar12b, CC10, Dod10]. rows [AG12, HTW13, PS12]. Roy [Pat12a]. rule [BB10, Ji12a]. rule-based [BB10]. rules [PR10, SB11]. Ryser [EGR12].

S [Bar10b, Sla10]. Saddle

[ACG14, CLCL12, DYW14, SS13b, Wan11b]. saddle-point [DYW14, Wan11b]. Said [MM10b]. Sakamoto [Car10b, ZY12a]. Salem [Lak10a]. same [HZ11b, HZ12b, MQ13, SS13c]. sampling [DPF10, GHMPVP11]. sandpile [AV12]. satellite [FNY13]. satisfy [dCF12]. satisfying [Özd13]. Scalable [KOPT13]. scalar [AR10, Cim11, RAAGAVS11]. scale [KKM13, LRV12, LHZH11, Sad12, WCKL13]. scaled [LJ11, Sra13]. scaling [Fri11, JT11b, Ser13]. scattered [DS13]. Schatten [MTS11]. Schauder [Per14]. schemes [GH13a, GH13b, GZX14, GMV11, Kim12, MGLW11, MW12a, MW14a, WLGM11, ZY12b]. Scherk [dlCdlRMP14]. Schmidt [AS14, Wil13]. Schmidt-Spitzer [AS14]. Schneider [Tao11]. school [FdC12]. Schreier [FP13]. Schröder [EWY12, lYnZpYH13]. Schrödinger [WZ14c, aCCS14, DLMZ14, Dub14, WZ12, WZ13e]. Schur [Aud13c, BFS11, CGMS10a, DV10, DW11, FKR11b, GS10c, GLS13a, GK12, HMS13, HL10b, KL12, KY14, LH10b, LHZH11, LZL12, Ney11, Seg14, SC12, Sto12, Tim14, tHR13]. Schur-type [BFS11]. Schwarz [GD11a, GO12]. Science [CS11]. Scrambling [Kim13b, CL10a, HL10c]. SDD [GEP13]. Search [MŠ10a, CKAC14, Gle11, Kam10, WZ14b]. secant [BL13b]. secants [BBCC13]. Second [Bar10b, AHL11, Bos11, Böt13, CCL14, Das10b, Das11, Kal13b, Kol13, KY11, KPY11, LY11a, LBLS12, LGS13, LHW11, LS14, SS11d, Sta12, Tam12, WZY14, ZLW12, dLN13, vBM13]. second-neighbor [Böt13]. second-order [LBLS12, LS14, SS11d]. section [Sha14b]. Seidel [DHS10]. selecting [HYF14, YiKIS12]. selection [dHM11]. Self [OR12, RS14a, SS11d, ADW13, BT13, DL14, DZ12a, Gle11, HSZ12, Kam10, MSS14, SSR13, Tif11, vBM13, BDK11]. Self-adjoint [OR12, RS14a, SS11d, ADW13, DZ12a, HSZ12, SSR13, vBM13]. self-dual [MSS14, Tif11, BDK11]. self-organizing [Gle11, Kam10]. selfadjoint [Dra12a, MMRR12, RS12b]. Semi [LW12c, Vas14, AS12a, BCS13b, BLL12, BL14, CKAC14, CT14b, DYW14, FKR11b, GL10a, Gon11, Hun10, KN13b, LLB13, Lop11a, VV13, ZH11a, vdW14]. semi-Cayley [GL10a]. Semi-centralizing [LW12c]. semi-convergence [DYW14]. semi-definite [BCS13b, BLL12, BL14, LLB13, vdW14]. semi-groups [KN13b]. semi-Hilbertian [AS12a, Gon11]. semi-invariants [Lop11a]. semi-linear [CT14b]. semi-magic [Hun10]. Semi-Markov [Vas14, VV13]. semi-nonnegative [CKAC14]. semi-radii [FKR11b]. semi-symmetric [CKAC14]. semi-triple [ZH11a]. semicrossed [DD14]. semidefinite [AHAPP10, Bal10, BMW10, BMN^{+}13a, CPV10, Dea11, EEH^{+}13, FJ10a, FS14b, Fur10b, Kak10, LV14, Mit11a, MNZ12, Net10, Pop10, Sag11, Sag13, SM13, SvdH11, WXH10b, Zha14b, Zim13]. semidefiniteness [Dru14]. semifields [Sin10a]. Semigroup [Mer10, Tan11]. Semigroups [Sem10, BPA^{+}11, BMR11, CD13, Jun12, KLP13, Mar10, OR12, PRW11, Pop13a, SLS13]. semilinear [KP13, ySpW11, ySpW12, ySpW14, dOHKS12, dOFK^{+}13]. semimodules [AGNS11, BH11a, LT10a, Sin10a, SN12, SN14, Tan14a, Tan14b]. seminorms [GD11a]. semipositive [JT11b].

Semiregular [BL10b]. semiring

[FKW13, JP11, SB11, Tan14b]. semirings [AB12, BHMR12, DO11a, DO11b, KP13, KSB12, Per11, SS11c, Shi11, ySpW11, ySpW14, Tan14a]. Semiseparable [Gem10]. semisimple [FG13c, HT10b]. Semismoothness [LQY13]. semispaces [NS11c]. semitransitively [BM13b]. sense [JP11]. sensing [ALPV14, KG14, PS14a]. Sensitivity [Cas13]. Sep [Gle11]. separability [HQ13]. separable [AS12b, Nie12, tHR13]. separate [DU14]. separating [LV11, LT13]. Separation [Bar12a, AHL^{+}14, BH11a, CJ14, LH10b, NS11c, vdH14]. separators [NdM13]. sequence [AW13b, BBC^{+}14, BMSW10, BDS13, BDFP11, BCFP12, BDOvdD12, CCGR13, CGTR14, CH12, CK14, DP10, Lac13, aLwW13, MN12b, PPK13, Pep12, TD13, Tan10a]. sequences [BV12a, Ben10, BHKL10, CC12, CL13b, CPK11, Dum13a, FKR11b, FKR12, FKM13, HNZ12, JJKS11, JMP13, LL10e, LM10c, SCS11, Sev14, SV13, VS14b, Wan14a, lY12, dF10]. sequential [Dug11]. Series [Gar12, Rod12a, Zha12a, dMR12, BGP11, Boj13, He11, KM14, Lak10b, Slo12a]. set [AHLvdH13, BS14, BB13c, BHvdH11, Cal12, FdC10, JZZ13, Kim13b, LT11a, Lop11a, LdlP11, MS13b, Mar14a, Roh11, Tan10a, XWD13]. Sets [FdC14b, PV12, BBH^{+}12, BJ10a, BH11a, Cal12, CSZ10, CR10c, CP11, CLHQ14, Dau12, FHL^{+}11, GLZ14, HT14, HHL10, HRT13, HCY10b, JSS13, Lim13a, MH13b, Mey12, MS11b, NN13, Net10, Pep12, PR13a, Row14c, Sin10b, VW10b, WLHL10, YHH12, ZY12b, ZSWB14, Ziv12, dS12a, dSC14, vdH14]. setting [GJ11, Gul11]. seven [BSU14]. Several [GL14, ZY12b, CHZ13, FJ10b, Pro10, SV11, Zha14c]. Seysen [Maz10]. shadows [DGH^{+}11, GZ13]. Sham [YM12]. Shannon [IIK^{+}13]. shape [Cha14, GL10b, HG10, IS14, NT10a]. shapes [HJN12]. Shapley [FDS13]. sharable [Cec10]. share [Pro10]. Sharp [BS13b, CW10, CLS13, DZ13, HJZ13, NCdS14, Pel14, XX14, XZ14, YWS11a, ZLW12, AHS10, CTG13, DFR13, Der13, HG11b, INT11, KHG14, Lan14, RMAJ10, RL13b]. sharpened [Alt13]. Sharpening [Reh10]. sharply [MW12a]. Shaun [Gar12]. Sheffer [He11, Wan14a, lY12]. Sheffer-type [He11]. Sherman [MS11a]. shift [CN13a, CN13b, GTW13, HG11a, Mei13, TW11b, Tsa11, VU14, WW13a]. shift-and-deflate [Mei13]. shift-invariant [HG11a]. shifts [KY14]. Shorrocks [JP11]. Short [Mir10, BL13a, KN13c]. Shortest [Voy13]. shrinking [MD12c]. sided [CRS14, LYS13, PPZ14, Šeg10]. Sign [AT11, BDM^{+}12, Hua12b, OTdDv12, AHL^{+}14, BP12, BDH^{+}12, BKMS13, BC14c, CFJKS13, CL10b, CGSCZ10, CP10, FHS14a, FHS14b, GLZ14, GS12a, GOvdD14, GK14b, GB14, GIP12, GOvdD12, HL10b, Hua11b, HLZ12, HLZP13, Hua13b, JMS11, KKB11, MMRR12, PP11a, PR13b, WLHL10, YS13, dS12a]. sign-definite [KKB11]. sign-matrices [CGSCZ10]. sign-patterns [dS12a]. Signal [BOZ10, BO12]. Signature [HW14c, HLZP13, O’D14, Sin10b, WF14]. signed [AHLvdH13, AT14b, Bel14, FWW13, GKZ11, LLH10, LYL13, Vij14, YW11, ZBW12]. Signless [BZ12a, YWS11b, Zho10, ACG^{+}11, ACM^{+}12, AOTR13, BMSW11, CT10, CW10, CT12, CTG13, CS10b, Das10b, Das11, DLS14, FF12, GK14b, GW13b, GCY14, HL10a, HL12, HJZ13, HY14, LS10, LWZ11, LW12b, LZ12b, LTS13, LL10c, LL11c, LLT13, LL14b, LL14c, LSD14, MK12, NLL13, TW10a, WB11, WF12, XZ13b, YY14a, YFW10, YWS11a, ZZ13, ZHG13, Zhu10a, dLOdAN11, dLN13]. Siler [CPH11]. Silva [FdC12, LPQdS10]. similar [GS12c, HLZP13]. similarity [Bar10a, Bar13a, CAV13, CD11, CLHQ14, Far11a, FGS11, FFG^{+}11, FJ14, FHS11, Ger12, LNT13, dSP10a, YY14b]. similarly [Nie12]. similitudes [GS12e]. Simple [ATS12, CBB13, GJTP13, WZ13e, Dub14, FGG10, KZ10, MB13, RR11, ZY12a]. Simplification [dO12]. simulation [Hür13, LAL11]. Simultaneous [BLLM13, GHS13, LS13a, LV11, LV12, MM11a, Ara12, Bar10a, Bar13a, Gna12, dSP10a]. sine [LdS13]. Singer [DH12b, WY14b]. single [BR14b, MM12]. Singular [AK12b, BHZ10, CN13b, DS10, HK10, BFRR14, BGV12, BCEM12, BR14c, CQYY13, DD10a, DW11, DYW14, DdF13b, DdC13, DLV13, Dur12, GH13a, GH13b, GZX14, HSZ12, JK13, KN13b, Liu14a, LS14, MB13, MM11a, ND11, Nik11, OLW14, dSP10e, RR14, Siv13, Vul12, Wül13, XSS13, ZCQ13, ZJ10, ZH12]. singularity [dSP12b]. Sivasubramanian [Sat14]. Six [CdGS12, CdGS20, LGSC14, DK13a]. Six-dimensional [CdGS12, CdGS20, DK13a]. size [BL12, Cir13, NS11a, ST12, SBMT10]. skeleton [ACDM14]. sketching [NNW14]. Skew [ABS14, CCF^{+}12, DKS13b, MMMM13, AW13a, ABS10, CLL13a, CLL13b, CD12a, CLHQ14, DYW14, FM11a, GX12a, GHW^{+}13, IMA10, LL11a, LL11b, LHL12, MS13a, NS12c, iO12, Özd13, Sev10, Sev14, Sha14b, TT10, Tia11a, WZY14, XG13, YT13b, Yan10a, Zhu12c, de 13]. Skew-adjacency [CCF^{+}12]. skew-energy [WZY14]. skew-Hermitian [DYW14, Sha14b]. Skew-symmetric [DKS13b, MMMM13, AW13a, CD12a, FM11a, IMA10, MSS12, iO12, Özd13, Sev10]. skew-symmetrizable [Sev14]. skews [Hil13]. slack [GGK^{+}13]. Slant [GS10e, Sed11]. Slepian [FZ11]. slice [Fri11]. sliding [Koz14b]. small [BLL13, Cal12, DGH^{+}10, DGZ13, DD10b, FK13, JSS13, Kol13, NS13, Shi12c, WB12, YFW13]. smallest [GK14b, GLS12, Kal13b, KPY11, ZHG13, ZJ10, dLOdAN11]. Smith [FP11, MMMM13, RRM11, Sad12, TW14, Wil14]. Smooth [LM10b]. Smooth/impulsive [LM10b]. snub [KAAK11]. Sobolev [KKLY14, RBP12]. Sobolev-type [KKLY14]. SOC [PCC12]. SOC-convex [PCC12]. SOC-monotone [PCC12]. Solution [Ano12a, Byd10, Dru12b, HJLS11, KPRT14, LwCJL11, CQYY13, CI10, DS11, DD11a, DDG^{+}13, DKS13b, Dod10, DS12c, GKL11, HMR12, HLS11, Ji12a, Jim10, KMS13, KD12, LCwCL11, LHZH11, LdlP11, MS11a, PLS14, Per14, Roh11, SMC11, VS10, pWlW14, Zha12c, ZLD11]. Solutions [Brä12, AiS13, AG10, BLLX11, BM10, BM12b, DH12a, FMWW12, FH12c, Fur10b, Kyr13, LV11, LV12, LLD13, LL13d, LTX14, Mir12, Miy13, Miy14, Sag11, SW11, WW10, WD10, XSS13]. Solvability [LLW14, XSZ13, Bie13, Hla13]. Solvable [SM10b, WGL12, CLOK13, CKLO13, CM13, KRH14, LMO16, MSvW12, ŠK10]. solve [Bou11]. solved [AA11]. Solvents [LT12a]. solver [PPKR12]. solves [BGW12]. Solving [BN13, MZ12b, MSP13, PQ12, SB11, TmYsH11, WCKL13, Bey12, HN10, KKM13, SK14, SHZ10, XX14, ZY12b]. Some [AT14a, BDH^{+}12, BCD10, BS12d, CPZ13, CW12a, CGM^{+}10, Cra13, CLHQ14, CI13, Dai13, DD14, FH10a, FH13, Fie13, GL12c, HG11b, Hun10, JLN13, JL12, KK12, KK13, tLyLWqW10, LJ11, LKN13, LCZ10, LSD14, MA12, Maz10, MK12, MS14b, Naj13, NPP13, SYH14, Shp10, Sta12, Sto11, TKLX14, TPZ12, Wad14, WXH10a, YW10, YY14c, Zha14c, ZBW12, AM13a, AdFM11, BP14, Bar13a, BM12b, BRZ13, BH11a, BZ12a, BZW14, CRSS14, Cau11, CHZ13, CFL13a, DH12b, Duk12, Duk15, Fan12, Fis14, Fra12, Gha13, GL10c, HLW14, HG10, HG12, HL11c, JNS13, KI10, KKR11, KJK13, KP14b, Kus13, Kyr13, LLD13, LY13, LW12e, LHL13, MZ11, MR12, MN12b, Nat13, RRKK12, RMAJ10, RM10, Seo13, SHZ10, SSZ13, wTlW13, Uch10, VU14, WW10, WFM11, WW13d, Wód14, Wój14b, YT13a, Zha12b, Zho11, ZCWZ13, ZXZ10, ZH12]. Some [DGGJ11, SBMT10]. Somos [CH12]. Somos-4 [CH12]. SOR [LMT10]. Sorensen [BELK12]. Space [Bra10, And13, BV12a, Bai14, BR11, BFK^{+}13, BEV13, BPDC14, DK13a, DK13c, CD12c, Dai13, DK11, DK12a, DK13b, DV14, DWXS12, DK14, FP14, GH13b, GN14, HKK^{+}12, HN14, dHLMS13, KLP12, LSV12, MARC13, Mat13b, Mer10, MMM13, PO10, dSP11b, dSP12b, PGM^{+}11, QH10, Ros12b, Tim14, pWlW14, WC11, vBM13]. Spaceability [BS14, BDFP11, BCFP12, CGMPSS14, RS14c]. spaceable [BFPSS12]. Spaces [Özd13, dSP14, Qui11, AS12a, AM13c, ARZ11, ABGPSS14, BG13, BDD14, BdFdP11, BDFP11, BCFP12, BCF12, BRZ13, Bud11, CD10, Cau11, CM13, Cir14, De 11, DW10, Dra12a, DP10, DX13, Dub14, Fan12, FM11a, FFS11b, For14, GZX14, Gon11, GLW13, GL14, HMT10, HZ11a, Hua11a, HZGY12, HZJ12, KP13, Kec13, Köh14, Lac13, Lán10a, LMM11, LM12, LHG10, LT11b, LT13, MMS12, MZ12a, MZ11, MM13, MN12b, OLW14, dSP12a, Pep12, Per11, PT13a, RAAGAVS11, RS14c, SP13, SST14, ySpW11, ySpW12, ySpW14, SM12, TD13, Wal11a, WLGM11, WY13, Wój14a, Wol12, Wor14, YW10, ZHQ14, de 13, dSW12, tHR13]. spanned [LT10a, RY12a]. Spanning [GS10b, Bap10, CW12b, Góm10, LS13b, LHL14, LHGL14]. Sparse [Ano12a, Dum13b, GP13a, KR12, KPRT14, PT14, CCL14, MSP13, NNW14, Rue13, Wan11b]. Sparsity [KKL13c, DT11, JKN14, MZ13, Zho12]. Special [BM13c, GRS^{+}10, HP12a, LPQdS10, Stu13, AAH13, BDH^{+}12, DGMS10, FM11b, FV13b, FKLT13, HLS10, Nor12, Nor14, PdFDV14, RR14, TZ12a, TD13, Xu12, BBD^{+}11, KPRT14]. specified [Nik10b]. Spectra [Bal12a, Bal12b, Bru10, CD12b, DL14, DK13d, JMP10, LL13c, MM11b, Nor14, Vla12, ABS14, AH14, BV12a, Ben13, Ben14a, Bou10b, BZ12b, BZW14, BW12, CTW11, CL12c, CS11, Ili10a, JZ14, LaG13, Laf12, LS14, LSD14, NP14, WS12, WY14a]. Spectral [ÁNPQ12, AS10, AW13b, Aud10a, BV13a, CR10a, FN10, GM11b, HY14, JS12, Kal13b, LS11b, LHW11, aLwW13, Lop11b, LLT12, MW14c, Ref12, RM10, SS13b, SSR13, DDM14, UZ14, WB12, WC11, Alt13, AH10, BT13, BdP13, BMSW10, BMSW11, BR14b, BL10b, BNP11, BNP12, BNP13, BH13a, BC14c, BZ12a, BS13b, CT10, CLS13, CLL12, CKST11, CvDKL10, CTG13, CS10b, DHLX12, Dai12, Dai13, DD10a, DP10, DZ13, Dum13a, FYI10, FZ13, FGG10, FG13b, Für10a, FJ10b, GLS10, GL12a, GR10, GMV11, GLS12, GL12c, GLS13b, GW13b, Güv12, HJZ13, Hil12b, HH11b, Hua11c, JZZ13, KG12a, Koz10, Koz14a, KKL13c, LLS12, LL13a, LS10, LT11a, LSC11, LWZ11, LWV12, LS12b, LCZ10, LL10e, LL11c, LW12e, LHWL12, LH13, LHL13, LX13, MR14b, MS13c, Mou12, MP13a, NP12, Nik10b, Nik13, NOL13, NLL13, OM13, OM14]. spectral [PLL12, Pep11, Pep12, PR13a, PS12, PRT13, PS14b, SWA12, SBMT10, SWT13, SR12, Ste11, WXH10a, WZ13c, WL10b, XZ13b, XZD14, XZ14, XG13, XE11, YFW10, YWS11a, YWS11b, ZW12a, Zha12b, ZHG13, Zho10, Zhu10a, vDF14, CHK^{+}13]. Spectrally [Bot10b, ES13, GS11a, GS12a, GS13b, MGSW14]. Spectrum [DHKQ13, KM12, Suz13, ACM^{+}12, AR10, Ang13, jASZ12, AT14b, BS11b, CJ12, Cos11, Cos14, CT12, CI14, Drn13b, FTDA10, GL10a, HLW10b, HZ12b, Kra12, LV11, LHWS13, LL14b, MSP11, SDNS13, TW10a, Ter11, ZWZ10, ZH11a, ZL11, vBM13, RY12b]. Spedicato [MAGR13]. speed [Zhu11b]. Spheres [PP13a, KK10, KK12, KK13]. Spherical [DGGJ13]. spill [KD12, MD13]. spill-over [KD12, MD13]. Spin [PP13a]. Spitzer [AS14]. splines [ACG14]. split [JZ11, MD12a, MD12b, MSD13, Ros12c, ZY12b]. splitting [DYW14, MZ12c, Rad13, wXL14, wXZ19]. splittings [JM12b, MS12]. spread [CL14a, Drn13b, EHH^{+}12, LL10c, OdLdAK10, SK13, WZL12, XM11, YL10]. Springer [Tam12, Zha12a]. square [BP10, Bot12, BLS14, Che14a, DD10a, DF14, FHS14b, GS12d, Lav10, MARC13, Nor12]. square-free [GS12d]. square-zero [Bot12, Che14a]. squared [Sad12]. squarefree [Hil14a]. squares [BW11, Byd10, CMNW14, CCL14, DS13, Dum13b, GJTP13, HMR12, dHLMS13, Ji12a, Kyr13, LLN^{+}12, LJ11, LWY14, LZ11a, LGZ14, LXK12, MM10c, MR10a, Miy14, Nor14, PO11, PPKR12, Wu10a, LB14, LKN13, SS10c, Rod12a]. SSM [JV10]. SSM-property [JV10]. stabilisation [LS12d]. Stability [CN11e, Fou14, Kal13a, Lan13, Vul12, BV12a, BG12b, Dai12, Fis14, FV13b, GV11, GM11a, HDPT12, HKP13, HRT10, KKB11, LS11a, Mor10, OM13, PM10, PKR12, PP13b, RJH11]. stabilizability [FV13b]. stabilization [Liu13]. stabilizer [LWGM10, LWGM12, Ye11]. Stabilizers [GS12e]. stabilizing [BO12, GKL11, Per12]. Stable [BLLX11, JZ11, BR14c, BBS12b, DX13, GOvdD12, HZJ12, Köh14]. stably [DP12b]. Standard [AAT12b, ySpW12, BM12a, GH11, HH14, Hil13]. Star [DM11, RTR10, Row14d, JR14, PRT13, Row12b, Row14b, Row14c, ZSWB14]. Starlike [BZ12b, SHS12]. stars [FG13a]. State [Liu13, ARZ11, BP10, BLLM13, CFG^{+}14, CG11, FG13a, GKS^{+}10, RvS13]. state-dependent [GKS^{+}10]. states [AS12b, CHLW14, HQ13, MS10c, SS10a]. static [Per12]. stationarity [Bos11]. stationary [FS14b, Gul11]. statistical [KS12b]. statistically [CC12]. Statistics [Zha12a]. Stein [Ara12, Fuh10a]. Steiner [FMT12]. stencil [XX14]. step [BCY12, DHLX12]. Stewart [Bai11]. Stiefel [FV13a]. Stieltjes [FKM13, PRT13]. Stirling [CD14]. Stochastic [Ben10, Pan11, DHS12, Fan10b, GS10a, GS10b, Gul11, HL11b, Hun10, HP04, JP11, LL12, LXL^{+}14, LL14a, MZ12b, Mou12, MAM^{+}13, PP11b, PSW11, SK14, Vas14, XLG^{+}13]. stochasticity [Sha13a]. Stokes [OM10]. Stopping [Lim11a]. Størmer [HOT13]. straightforward [vdW14]. Strakos [Gre13]. strategies [CKAC14, Jun14, KM11, Sta14]. strategy [Raf14]. Stratification [JKV13]. streaming [NNW14]. Strict [BJZ12, CN13c, Cim11, Mol12, Alt13, Peñ14]. strictly [CRU10, HLZ12, LZL12]. strings [PRT13]. Strong [HDPT12, LL10b, Liu14b, LL13d, TZ12b, Bar10a, Bar13a, BCF14, BF14c, Car11, CPR10, HKP13, Hla13, LLD13, LHL12, BDK11, LV14]. Strongly [BMGMC12, Tif11, ÁvW13, Cam13, HMT10, HY14, LS12b, NS12b, PY10, Row12b]. Structural [CJ11, Zho12, DIP13, JDY13, LZ12a]. structurally [KBS13]. Structure [BR12, Bre14, Cer10, DGU14, ET13, Fuj10, LL12, AA11, ART13, BD12a, BBdH12, BRA11, BJRS11, CKS10, EFN09, EFN10, HG12, Hua12b, JLN13, JK11, KRS13, KW13, LS13d, MD12b, Mar14b, Moj14, Nit10, OT12, PGM^{+}11, RE11, Rod12b, TGM11, ZYL10]. Structured [AAK11, AA14, BCF14, BF14c, CLCL12, AA11, AAT12b, Bat14, BdP13, BDG13, DKOT12, DT11, HF12, HL11d, Kak10, MMRR11, MMRR12, ST13, LBLS12]. Structures [DV10, GW14b, AM13c, BOZ10, BE10, BDV12, CNT12, CL12b, JKN14, KM13b, MW14a, MMMM10, RMT11]. student [DD10c]. Study [CNT12, KZ11]. Sturm [jAS13, KZ11, Qua10, SB12]. sub [JS13b, Nie13]. sub- [Nie13]. sub-direct [JS13b]. subadditivity [BP13].

Subalgebras

[GS10d, BZWL13, Sed11, ŠK10]. subclasses [MS12, Siv13]. subconstituent [LWGM10, LWGM12]. Subconstituents [GW11, GWZ13, Gu13]. subdiagonals [CGM11]. subdivision [BYZZ14, GMV11, Gum13, LL13c]. subdivision-edge [BYZZ14, LL13c]. subdivision-vertex [BYZZ14, LL13c]. subeigenvectors [BM13d]. subgraph [CR10a, MW14c]. Subgraphs [Ter11, AAJ12]. Subgroup [FW14a, GH12]. subgroups [CJ14, HS12b, Slo12a, Slo12b]. subject [YFW10]. sublinear [NdM13]. submanifolds [Lim11b]. Submatrices [JS13a, BFdP11, FG13b, LS12d, LSH12]. Submatrix [FJMP14, JMP12, BC14b]. Submodular [FG13b]. submodules [NCdS14]. subnormality [JJKS11, KY14, TT11]. Suborbits [LWGM10, LWGM12]. subpencils [RS12a, RS14b]. subpolynomial [SR13c]. Subresultants [DKS13a]. subscalarity [KJK13]. subschemes [CM12b]. subsemigroups [Tan11]. subsemilattices [ABK14]. subset [Dai13, dHM11]. Subsets [Wód14, AM13c, CMRR13, Fra12, Liu14b, RS14c]. Subspace [Gre13, BEM12b, BT10, BT14b, Bou13, Cha10, DZ12a, FP11, JR11, MSS12, NP13a, NS11a, PY10, dSP11a, RRZ13, Wój14b, XE11, YZ11]. Subspaces [DGMS10, AW13c, BGV12, BFK^{+}13, Cir14, Dom10, DKM^{+}14, GH13b, GV11, Gro14, GS12e, GLW13, HW14a, HG11a, IJ12, Irv12, MMP13a, MD12c, PP13b, SSR13, WZ14b, BDK11]. substitutions [BK11a]. substructuring [Blö12]. Subtracting [SC10]. successions [MS13b]. such [KY11, LSTW13, MP14a]. sufficiency [XW11]. Sufficient [EGR12, CHK^{+}13, LLD13, LS12d, Sha13a]. Sum [PR10, DDM14, AFHP14, AKZ13, AM14, AOTR13, BZ13, CSV14, DXL13, DZ12c, FHRT11, HMTR10, HTS14, Kaw12, KLZ14b, MPRW11, Mer12, Nik11, TT10, TT12, TZ13b, Zuo10]. summable [dMR12]. Summary [DK13a]. summing [Cam13, Pel14, SR13b]. Sums [Bot12, dSP12c, Rod12a, dSW12, BW11, Bar12b, Ben13, BL14, BG11, CY11, EWY12, Fri11, HS14a, HS14b, JS13b, Kra12, Lee13a, OM12, dSP10d, PT13b]. sumset [Cal12]. super [DKS10, Ros12b]. super-operators [Ros12b]. superadditive [Nie13]. superalgebras [LW12d, LCM13, MSD13]. Superfast [KMS13]. supergraphs [MNZ12]. superoptimal [CJ10]. superpositions [BMN^{+}13a]. superquadratic [Kia14]. superregular [ANP13, CNPP12]. Supplements [HK13]. support [FK13, Wei11]. supported [GHMPVP11]. supremum [XDFL10]. surgery [CKL^{+}13, NRR^{+}11]. Surjections [Lim10, lH10a, lH11b]. Surjective [Bou10b]. surjectivity [BMS14a]. survey [AH10, AH14, Sei14, YS13]. survival [VV13]. Suslin [JR14]. SVD [Šeg10, Sto12]. switched [FV13b]. switching [Dai12, GKS^{+}10, OM13, ZY14]. Sylvester [FH12c, FH10b, IW13, KS10, LV11, Lim11b, LSH12, Miy13, Qua10]. Symbolic [SPKS12, AMPT13, PS14b, dO12]. symbols [BS12a, BCD10, CRS14, KL12, SR13a]. Symmetric [BCMT10, Car10a, VS14a, JS13a, Lim11b, Qi13, Qua12, AiS13, AW13a, Aga14, ART13, Bal12a, Bal12b, BH12, BSK12, BSKL13, BdP13, BCEM10, BCEM12, BM13a, BBM14, BI13, BW13a, BDOvdD12, BK12, Bün14, CFJKS13, CL10a, CL12b, CL10b, CN11d, CD12a, CKAC14, DKS13b, Elo11, ES11, Fan10b, FM11a, GPT12, GKL11, HLZP13, HP04, IMA10, JV11, JDY13, KG12b, KM12, KL13b, LXL^{+}14, LQY13, LNT13, LS14, LXK12, LTX14, MMMM13, Mar13b, MY14, MSS12, iO12, Özd13, Per14, QS14, RV13, Reg13, Rub13, Ruk14, Sag13, SS13b, Sev10, Sha10a, SS11d, Spe11, Sta14, TT10, TG13, Vla12, WLHL10, XZ13a, XLG^{+}13, YT13b, YW11, Zho12, dS12a, GMT13]. symmetric/skew [MSS12]. symmetric/skew-symmetric [MSS12]. symmetries [AL14, EV11, fLyH11, MFGD14, Tre10]. symmetrizability [KKLY14]. symmetrizable [CGM12, Sev14]. symmetrization [Pál13]. symmetrized [Ber13a, HH14, Pat10]. symmetry [BLS14, MZ12c, Sko11, VW10a]. Symplectic [MVPS10, BFS11, DK13a, DK13c, DK11, DK12a, DK13b, GH13a, GH13b, Gu13, GLW13, Hil12b, LHG10, LWG13, dSP12d]. synchronization [VR12]. Synthetic [FN11]. system [DH12a, FMWW12, FZ13, KHG14, KAAK11, Köh14, LdlP11, MD13, RMT11, SHZ10, WD10, pWlW14]. Systems [BFK^{+}13, PQ10, AW13a, BLLX11, Baj14, BBCC13, Bap10, BV11, BM10, BV13a, BH10, BTYZ12, BLLM13, CDP10, Car13, CC14, CA10, CLCL12, CMN10, Czo10, CN10d, Dai12, DZ12a, Dur12, EM12, ET13, FGQ11, FV13b, FS14b, GLP^{+}13, GRdS12, GS10b, GHT11, GKS^{+}10, HDPT12, HRT10, Hil12b, Hla13, IT11, KH13, KRvS12, KRvS14, KMS13, KKM13, Lan13, LBLS12, LHZH11, Liu13, LM10b, Mah11, MS11a, MZ12a, MRS12, MZ12b, Miy14, MSP13, Mys12, NP13a, OM13, PQ12, Per12, PKR12, RGC13, RS14a, SSS13, Sch10, SB12, SVP11, SW11, SS11c, Sha14a, ST13, Tre11, VB10, WZ13b, Wei10, Wor13, YM12, ZCKS12, Zhu11b, ZY14, Gem10, KPRT14]. Syzygies [Din11]. syzygy [KMS13]. Szechtman [Qui11]. Szeged [FTDA10]. Szego [Seo13].

tail [Han11b, Han14b]. talk [Stu13]. Tallini [HK13]. tangential [Fuh10b]. Tanigawa [Rah13]. Tao [MW12b]. TCP [SWBS13]. TD [IS14, Nom14]. TD-pairs [IS14]. teaching [MLC^{+}10, PTPL10, TP13]. technique [Mei13, Rhe10, RŠ10, Row12a].

Techniques [DGH^{+}10, IPFD13, KY14, Tam12]. tennis [Dah12b]. tensegrity [AN13]. Tensor [Dug12, KS10, Nem13, SC13, BCMT10, CS13a, DKS10, Fri11, HM14b, HWSH13, Jai11, KRS13, Kub13, LPS13, LKN13, LQY13, MZ12b, Pat10, PGM^{+}11, QS14, Reg13, Rho10, SE13, SC10, ZCQ13]. Tensored [FH10c]. Tensors [Bra10, TS12, Bal12a, Bal12b, BGK13, BBCC13, Ber13a, Ber13b, BBM14, BZZ^{+}14b, BL13b, CPZ13, CQYY13, CKAC14, DQW13, FdCR10, Fri12, FKLT13, Fri13, HH14, KM11, Qi13, RV13, RE11, Sav14, Sha13b, SSZ13, SQ14, SSM13, UV13, Vla12, XC13, YHY14, YY14c, YY14b]. term [BSK12, BSKL13, BKMS12, FdC14a, KSB12, Tia12]. terms [CGMS10a, HJZ13, Hun14, MW14c, RL13b]. Terwilliger [GZH14, Kim12]. tessellations [CSAC10, CSAC11]. test [Ara12, Hua13b, Peñ14]. testing [Wal11b]. tetrahedron [IRT14]. Tetris [CHK^{+}13]. th [Guo10a, HL11b, LSC10, LLH10, Lin10, TNP12]. their [AK11, BEM12b, BSKL13, BZ12b, BZW14, CEM14, CTW11, Dor10, FKW13, GS12b, GKZ11, GB13, GIP12, GLW13, GL14, HMS13, HK13, HQS13, HHLS14, JZ14, KKL13a, KKL13b, Lee13c, LH10b, LHZH11, LSD14, Lom11, MGLW11, MMMM13, MR12, MS12, MSS14, Mit11b, Mou12, NY13, Ogu13, QSW14, Qui11, ySpW12, Tia12, VS10, WLLX11, WS12, WML13, WY14a, Wód14, Wu13a, ZHQ14, Zha14b, ZZC14, Ziv12]. theorem [AS14, BH13b, Bri13, CH13, Car10b, CP12, DLNN14, – Dok12, Dom10, Elo11, FTZ12, Fid10, FGG10, FGH13, Gum13, GK12, GMRS14, HH12b, Han13a, Hua12a, Hwa11, KN13c, Kra13, KW13, Laf12, LL14d, Mel13, MW12b, Nak12, OLW14, PR13a, Per14, Pit11, Pon11, Rho10, RL13a, RAAGAVS11, SY12, SSGL10, SRdAG10, Tao11, Wag11, Wat13, ZY12a, de 13, vDF14, Had12, Had13, Lee11, DDM14]. Theorems [CTW11, LHZH11, Bar12a, BFdP10, Hil12b, Nak10, Pin11, Sat11, Sat14, YW10, dlCdlRMP14]. theoretic [BB13c, Böt13, FH13, Kus12]. theories [MLC^{+}10]. Theory [Ber09, BFH^{+}12, MOA11, Sla10, Tam12, Zha12a, Ano12-30, AH10, BV12a, BPA^{+}11, Beh13, BS10, BBS12a, CC14, CS10b, DD10a, DD11a, DD11d, GL10c, Ika11, JZZ13, KZ11, LN12, MMRR11, MMRR12, Nik13, NV12, PO10, RR12, RL13a, SB12, Sei14, SRdAG10, TN14, Zha12d, HB12, Lim13b]. Thin [God10, God12, Shi12a, Cer10, Kim12]. Third [Bra10, KM11, SHS12]. Third-Order [Bra10, KM11]. Thirring [Aud13a]. Thompson [ISYY11, Lim13a]. Three [Cha14, Dor10, HLW14, Kar11b, BS11d, CKST11, Cir13, DP12a, DD11c, Dra14, KS13b, LZG14, MAS12, dSP10b, WBWH13, vDO11]. Three-by-three [Cha14]. Three-equipped [Dor10]. Three-parameter [Kar11b]. threshold [Bap13b, JTT13, VDVJT13]. thresholded [GR12]. tight [Bod13, BW13a, DHS10, FMT12, HS12a, Sin10b, Szö13]. tightly [BS13a]. Tikhonov [LRV12]. tilting [PYZ14]. Time [Pol12, BCY12, BLLM13, BJ13, DT10, DZ12a, FV13b, HDPT12, HRT10, Jun14, Kir10, KM14, Liu13, Mah11, PKR12, RMT11, Sad12, Sha14a, WZ13b]. time-delay [Mah11]. Time-domain [Pol12]. time-invariant [DZ12a]. time-varying [Liu13, PKR12, Sha14a]. times [BRZ11, Góm10, PR10]. Tits [GS12b]. TN [JW13]. Toeplitz [BH12, BS12a, BCD10, BBGM12, BF14b, Bün14, CRS14, CGM11, DK13e, Elo11, Gar13, GG13, HR11, KL12, KMS13, KW12, LMYY11, LJY13, MS11a, RR14, Rim12, SCSS10]. tomographic [PS14a]. TOP [Sha10b]. Topical [SN12, Sin10a, SN14]. topics [Ano12-30]. Topological [Bud11, RS12c, ABGPSS14, FRS14, JZZ13, TN14, AJ13]. topologies [Sha14a, ZY14]. tori [CSAC10, CSAC11]. torsion [Wag11]. Total [HC10, LB14, BP14, Dum13b, Kus12, LJ11, PO11, Peñ14, RMAJ10, YM12, SS10c]. Totally [BK10, FJ11, Gar10, HJN12, HWG14, AG13, CRU10, CRU13, CRU14, CDP10, FFJM14, GL12b, Hua13a, KS11c, Gar12]. tournament [BRLS12]. tournaments [BF14a, NS12c]. TP [HJN12, JW13, PS14b]. TP-critical [PS14b]. Trace [Aud12, BLdPS10, DK13e, LLB13, FL10, HH13, Hia13, KLS12, KK14, Lu11, MY14, MSvW12, Ros12b, Spe11, WLHL10]. trace-preserving [Ros12b]. traces [FF10, KHG14, WXH10a, WZ13a]. track [Han13a]. Tracts [Lim13b]. traditional [SLS13]. traffic [FGQ11]. training [DT10]. transfer [BKV14, BP10, CFG^{+}14, CG11, FG13a]. transfer-function [BKV14]. transferring [Wój14a]. transform [BH14a, Boj13, BJ10b, Bri13, DH10b, FK13, HFS13, KB14a, KB14b, O’D14]. transformation [DMMY10, Lak10a, Li12, LLMZ12, TZ13b, XZD14, XE11, ZLL12]. Transformations [FKM13, BB13a, Bal10, Buc10, DK13d, DD11b, DGAM14, DFS14, GTR12, HG10, HG12, JV10, RAY14, Sim10, wTlW13, Tao13, TG13, TGM11, VW10b]. Transformed [KS12b]. transforms [LdSP11, LdS13, Xu14]. transition [BANP12]. transitive [Kuz10, MW12a]. translation [GS12b, JMP10]. transmission [Wei13a]. transport [GL10c, LwCJL11]. transpositions [KM12]. transversals [AGK11, Fan10a]. treatment [AMPT13]. tree [Bap10, CGTR14, DT11, FHRT11, Góm10, GS10b, LLS11, MR14b, SHS12, XM11]. tree-based [MR14b]. tree-structured [DT11]. trees [AJRT14, AW13b, BL10b, BZ12b, CJ12, CW12b, DZ11, FZW11, FHRT14, GFY10, GLS13a, HL11a, HT10a, wH13, JNS13, KS13a, LSC10, LT11a, LWZ11, LW12b, LS13b, LGSC14, LY11b, LZ14, LHL14, LHGL14, MWZ13, NP13b, NOL13, PLL12, PdFDV14, RJ10, RJ11, Row10, SvdH11, SWT13, Tan10a, wTmS12, WT12, WF14, WL10b, XF11, ZZL11, Zhu12b, vdH13, JT11a]. tri [EN11]. tri-additive [EN11]. triangle [Dea11, MPSS10, PY10]. triangle-free [Dea11]. triangles [He13, LHL13]. Triangular [DGMS14, Kaw12, ÁvW11, ÁvW13, BBG14, BG14a, Ben11, BE12, Bie13, Cir13, CM14, CI14, DYW14, DW12a, DWW14, DW14, ES13, Ere13, FFG^{+}11, FdC10, Gho13, HC14, HW11, JQ11, LMYY11, MW12a, VS11, Wan11a, WWD13, WW13b, WX11, WMZ14, XW10b, XW12, XSH14, YZ13, YZ10, ZZ10, ZZ12]. triangularizability [YT13a]. triangularizable [dSP12c, RY12a]. Triangularizing [TTZ13]. trichotomy [MSP11]. tricyclic [CL11a, GL12a, LWZ11, LY11a]. Tridiagonal [BC14a, NT10b, NT11b, ÁNPQ12, AMJ14, BT13, BDH^{+}12, BDG13, BC12a, BK12, CM10b, GG13, HH11a, INT11, IT11, KHG14, LHLL10, MS11a, NT10a, Qua10, Rim12, SC12, Van10, Wül13, dF10, dS12a]. tridiagonalization [PPZ14]. trigonometric [CN11b, LdSP11, LLMZ12, Sra13, ZLL12]. trilinear [DS10]. triple [Mol13, Reh10, Reh11, SM12, XW12, ZH11a]. triples [AAT12b, BM12a, GWH13, HWG13, HWG14, wH12, Siv12b, Siv12a, Ili10b]. Tripotency [Kis15, XX12]. tripotent [Kis15, XX12]. tripotents [BY11]. trivectors [DK13a, DK13c, DK11, DK12a, DK13b]. Tropical [AGM14, AGK11, BS11d, DHS12, GS12f, dlP11, Cas10, CJR11, GMH14a, GMH14b, JK11, KNS14, LdlP11, Shi12b, Shi12c, SLS13, Wag11, Wil11]. tropicale [Cas10]. truncated [DK13e, Sto12]. truncation [Cha12, DGMS14]. trust [CKAC14]. Tsallis [IIK^{+}13]. TSVD [BJRS11]. TT-cross [OT10]. Tucker [BGK13]. tuples [BDS13]. Two [AH13a, BC12a, BSST13, wH13, IPFD13, LB14, LHG10, LYS13, Ma11, Šeg10, Ste11, TT11, AKM14, AR10, AIL12, AK11, BB13b, Bot10a, Bot12, BS10, BS11c, BS13a, Brä12, BZ13, CGMJ14, CN10a, HO11, Hil14c, HMP12, Jun14, Kis15, KW12, KKM13, KR10, LLS11, LwCJL11, LSR11, Lop11a, LMT10, MW14a, Mar10, MH13b, MP10, NR10, NS12a, PPZ14, dSP10c, dSP12c, Pet10, PSW11, SW11, SH13, TZ12a, TH10, TZ13b, WBWH13, XX12, ZW12b, Ziv12, Zuo10, FF10]. two-by-two [BB13b, KKM13]. two-cyclic [LMT10]. Two-dimensional [Ma11, LwCJL11]. two-level [KW12]. Two-lit [wH13]. two-parameter [HMP12, MP10]. two-player [Jun14]. Two-sided [LYS13, Šeg10, PPZ14]. two-variable [AIL12]. Tykhonov [SS10c]. type [AiS13, AAT12a, AK11, BR14a, BRA11, BdlC14, BFS11, BC14a, CL12a, CA10, Dai12, DD11c, Dra12a, FPC13, GH13b, GWH13, GD11a, He11, HM14b, HWG13, HWG14, wH12, IS14, JKN14, KKLY14, Lee10, Lin13, Lin14b, LXK12, MFGD14, NY14, NT10b, Pag12, Pat12a, Rhe10, ST10a, SE13, SS10c, Sch10, Seo13, Sev10, Sev13, Tao11, TPZ12, Wad14, Wik11, Wik12, Wil13, Wol12, Wor13, CGGS13]. Typical [SSM13, Bal14, Ber13b, Fri12].

UK [Lim13b]. umbral [Ern13]. Unbounded [And13, CGM12, For14, GMMFPSS12, Pop14]. uncertain [Liu13, Sha14a]. Uncertainty [Yan10a, GJ11, MW12b]. unconditional [LM12]. underapproximation [GP13a]. underdetermined [Miy14]. Undirected [CGMJ14, CFHL14, FS14a]. Uni [ZHZF13].

Uni-mode [ZHZF13]. unicellular [FGS11]. Unicyclic [AW11, AKM14, BMSW10, CLL12, DC14, DZ12b, FWW13, FYI10, GHW^{+}13, HL12, HJL10, HLS11, Kal13b, KP14a, LFS12, LTS13, LZ14, SSGL10, YL10, YZF14, Zhu12c]. Unified [AM13b, JKN14]. Uniform [LMYY11, CD12b, JZZ13, Nik14, QSW14, XC13]. uniformly [MD12c]. unimodular [GY13, KBS13, MRW11]. union [Ziv12]. unions [Dau12]. Unipotent [CHJ13, Bot10a]. uniquely [KKL13a, KKL13b, Lee13c]. Uniqueness [LL10a, Der13, Mor10, Sta14, Wei13a, ZHZF13, Dra14]. unit [AG12, BR12, DW10, MPS10, Ref12, RL13a]. Unital [KLP13, BS12b, BG14b, Wan14b]. unitarily [FCL10, Fur12, Ikr10, MA12, Van10, ZH12]. Unitary [GPT12, Ger12, LNT13, AR10, BFK^{+}13, Bud11, CLHQ14, Far11a, FGS11, FFG^{+}11, FHS11, GHS13, KAMS11, LNN14, Li10, LWGM10, LPS13, LMW12, Mol13, Pop10, SY12, Ste13, Tad12, TMSS14, GTW13]. unitriangular [Bie14]. Univ [Gle11]. univariate [BL10a]. Universal [CFG^{+}14, HO11, AAF^{+}12, AN13, LV14]. universally [HCY10a]. Universitext [Tam12]. University [Bar10b, Gar12, Grü12, Lim13b, Rei11a, Rod12a]. Unordered [KS13a]. unreduced [Wül13]. unstructured [RW12]. updates [EM12]. Updating [SWA12, JDY13, KD12]. upon [CFJKS13]. Upper [BHKM13, CFL13b, DLS14, DZ12c, WZL12, AHS10, BBG14, CR10a, CGTR14, CLS13, Cir13, CM14, CGR14, CTG13, CI14, DWW14, DW14, FFG^{+}11, FTA11, FdC10, Gho13, GLS13b, HC14, JMS11, PJ13, RMAJ10, RL13b, SK13, SWT13, TC13, WW13b, WMZ14, ZLW12, ZZ10, Zhu10b]. upset [BRLS12]. Uriel [BMS14b, Loe12]. Use [PTPL10, VS10]. Using [MS11a, TP13, AHAPP10, AHS10, BBH^{+}12, BEK13, BRZ11, BTYZ12, BHAP12, CNT12, GS12d, Han11b, Han13b, Han14b, Kim11a, KZ11, MSP13, OM12, PS14b, RMT11, RW10, SPKS12, TMSS14, UV13, VS13, WXH10a].

validation [LJ12]. valuation [AW10]. value [jASZ12, AK12b, BFRR14, BHMO13, BMSW11, CQYY13, DLV13, FDS13, FdC10, MM11a, NdM13, ST10a, XWD13, ZCQ13, ZJ10, ZHQ13]. valued [BJ10a, CD12c, KKR11, Nie13, SN14, Tia12, vBM13]. values [AHL11, AM14, Buj13, CH13, CN12b, CJK^{+}13, EJLS11, HHMS10, HK10, Kra12, MB13, MP14a, Nik11, SS10d, ZH12]. Vandebril [Gem10]. Vandermonde [BOZ10, DU14, God12, KS14b, MM10b, ZYL10]. vanishing [KLZ14a]. variable [AIL12, KY14, LLY11]. variables [BW13b, DU14, Gna12, HYF14, Mat14a, Pet10, SV11, YiKIS12, Zha14c, dO12]. Variance [Aud10b, Fie13, Güv12]. Variances [CEY14]. variate [ANF11, DGJ10, Mat13a]. variation [Kal13b]. variational [CPZ13, TmYsH11]. variations [Mel13]. varieties [BL13b, NS14, Siv12b, Siv12a]. variety [Ber13a]. various [BBC^{+}14, FJ14]. VARMAX [KS10]. varying [Liu13, PKR12, Sha14a]. Vector [FM11a, ABGPSS14, BG13, BJ10a, BCF12, DK13c, CD12c, De 11, DK11, DK12a, DK13b, DV14, FW14b, KM14, Nie13, PO10, PP13a, Pol13, RR14, Wal11a]. vector-valued [Nie13]. vectors [Aga14, Cos14, Dah11, Fan12, HL11e, Nie12, Sav12, ySpW12]. Verdière [GB11, Gol13]. verification [JNS13]. Verified [FH12c]. Verma [WZ14c, WZ12]. Vershik [GH12, Slo12a]. version [KSA11, Kon13, Laf12, MP13a, NS11c, Raf14]. versions [GO12]. versus [Aud10a, BI12, Bea12, PQZ13]. Vertex [EHH^{+}12, BNP13, BYZZ14, CMRR13, CT11b, CJ12, LL13c, Row12a, TF10]. vertical [MV12]. vertices [ACM^{+}12, AdFM11, BNP11, DdF13b, DdF14, FdC14b, Gum13, HJL10, LLS11, LWZ11, LZ12b, LJY14, MAS12, NOL13, SWT13, WF10, XZ13c, Zhu10a, Zhu11a]. very [FMNW14]. via [ABBO11, Alt13, BFRR14, BY11, BO12, BBS12a, Car13, Cas13, Che13, CK13d, CGSCZ10, DK12b, Fur12, GRdS12, JJKS11, KL12, KKB11, LRV12, LS12a, LM10c, MW12a, MW10, NM14, Qua10, Rad13, RRM11]. view [ATS12, HM10, JKN14, Ter13]. viewpoint [PO10]. viii [Grü12]. Vitae [Ano13d]. Vol [Lim13b, Grü12]. Volume [Gem10, Sav14]. volumes [GK10]. Voronoi [GRdS12]. vs [PCC12].

W [Zha12a]. Walk [EdlPH14, Ben14b, CvDKP13, CBB13, DFG10, EdlP14]. Walks [MSvdD14]. Wang [Koz14a]. Waring [AW13a]. wavelet [HL11e]. way [CC10, ZHZF13]. ways [Wad14]. Weak [GMH14b, Hla13, KP14b]. weakly [HZ10]. Wedderburn [MAGR13]. Weibull [Fou14]. Weierstraß [BIT12]. weighing [KMNS12, NP10]. weight [AG12, Dub14, LT11a, LSV12, Tan10a, WZ13e]. Weighted [GTW13, JLLY10, LP14, Pál13, PP14, BKP12, CN13a, CN13b, EWY12, GX12b, GHW^{+}13, GS11b, KP14a, KKS12, KKL10, KLL11, KY14, KW13, LLY11, LT11a, MR12, MS14b, NP13b, NSC13, PO11, RM10, Tan10a, TW11a, TW11b, Tsa11, VU14, WW13a, WLL14, WY13, XCS13, Yam13, YZF14]. weighted-EP [TW11a]. weights [KP14a, Stu12, Tsa11]. Weil [AiS13]. Weitzenböck [DDF13a]. Welch [DHC12]. well [iO12]. well-quasi-ordering [iO12]. Wenzel [Lu12]. Weyl [DV14, FKR11b, FKR12, Nak10]. Wg [MD10]. Which [GGK^{+}13, CFK^{+}10b, DHLX12, FdC14b, Gu14, Ogu13]. Whitney [FV13a, Wag11]. Whittaker [aCCS14]. whose [BMSW11, BZ12b, CK14, DU14, HZ11b, HZ12b, HTW13, JG10, LY11a, LSR11, LHW11, MQ11, MQ13, SS14, Sta12, WBHM13, Wu10a, XG13, vDO11]. width [iO12]. Wielandt [HMS13, PR13a]. Wiener [BH13a]. Wigner [Yan10a]. Williams [ST10a]. Wishart [BRA11, DGGJ11, MW10]. without [ABS14, Ban13a, BNP13, VS14a, LMN13, MPP11, NT12a, Nik10b, PV12, Qui11]. Woerdeman [Rod12a]. Wold [Pag12, Sar14]. Wold-type [Pag12]. Woodbury [MS11a]. Words [Ros12a, LSV12]. work [HK13, Mac13]. workshop [BFH^{+}12]. wreath [Kim12]. Wu [CRSS14].

X [BF14b]. X-rays [BF14b]. xi [Lim13b, Rei11a]. xii [Rod12a]. xv [Gar12]. xvii [Tam12]. xxvii [Zha12a]. xxxix [Bar10b].

Yanase [Yan10a]. Yang [Kaw13]. Yuan [ZC14].

Zarankiewicz [Nik10a]. Zdenek [Gre13]. Zero [BBF^{+}10, GWW14a, Gha13, GB14, Mey12, PR13b, BG14b, Bot12, CT11a, CL14a, Che14a, CL10b, Drn13b, EHH^{+}12, EEH^{+}13, FFS11a, GOSvdD14, HLS10, HCY10b, HL11c, HL11d, JH10, KH13, KLZ14a, MZ14, Ma14, QCH11, Qui11, Row12a, Spe11, WLHL10, WMZ14, YHH12]. Zero-dilation [GWW14a]. zerodivisor [WMZ14]. Zero-nonzero [PR13b, GOSvdD14, MZ14, YHH12]. zero-one [CT11a]. zero-pattern [HL11d]. zerosymmetric [CL10b, WLHL10]. zeros [Aga14, ABSV12, CN10a, DD10b, FPC13, Hil14b, Mel14, SS11c, WZ13d]. zerosumfree [KP13, ySpW11]. zeta [MM10a, MS14b, SS13a, SS13c, Sto11]. Zhang [Tam12]. Zinbiel [ACGVK14, CCGVO13]. zonotopes [GK10, HLW10a].

References

Adhikari:2011:BES

[AA11] Bibhas Adhikari and Rafikul Alam. On backward errors of structured polynomial eigenproblems solved by structure preserving linearizations. Linear Algebra and its Applications, 434(9):1989–2017, May 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adhikari:2014:SMP

[AA14] Bibhas Adhikari and Rafikul Alam. Structured mapping problems for linearly structured matrices. Linear Algebra and its Applications, 444(??):132–145, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                                     //www.sciencedirect.com/
                                                     science/article/pii/S0024379513007118.

Ahmadi:2012:MRU

[AAF^{+}12] B. Ahmadi, F. Alinaghipour, Shaun M. Fallat, Yi-Zheng Fan, K. Meagher, and S. Nasserasr. The minimum rank of universal adjacency matrices. Linear Algebra and its Applications, 437 (8):2064–2076, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                                     //www.sciencedirect.com/
                                                     science/article/pii/S0024379512004247.

Alvarez:2012:MDC

[ÁAFG12] V. Álvarez, J. A. Armario, M. D. Frau, and F. Gudiel. The maximal determinant of cocyclic (−1, 1)-matrices over D_{2t}. Linear Algebra and its Applications, 436(4): 858–873, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100423X.

Abe:2013:ISO

[AAH13] Toshikazu Abe, Shigeki Akiyama, and Osamu Hatori. Isometries of the special orthogonal group. Linear Algebra and its Applications, 439 (1):174–188, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001912.

Akbari:2012:CNC

[AAJ12] S. Akbari, M. Aryapoor, and M. Jamaali. Chromatic number and clique number of subgraphs of regular graph of matrix algebras. Linear Algebra and its Applications, 436(7):2419–2424, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006537.

Adhikari:2011:SEC

[AAK11] Bibhas Adhikari, Rafikul Alam, and Daniel Kressner. Structured eigenvalue condition numbers and linearizations for matrix polynomials. Linear Algebra and its Applications, 435(9):2193–2221, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alahmadi:2014:CC

[AAK^{+}14] Adel Alahmadi, Shefa Alamoudi, Suat Karadeniz, Bahattin Yildiz, Cheryl Praeger, and Patrick Solé. Centraliser codes. Linear Algebra and its Applications, 463(??):68–77, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005680.

Ahmadi:2014:IM

[AAKM14] M. H. Ahmadi, N. Akhlaghinia, G. B. Khosrovshahi, and Ch. Maysoori. On the inclusion matrix w_{23}(v). Linear Algebra and its Applications, 461(??):42–50, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004005.

Armandnejad:2012:RCM

[AAM12] A. Armandnejad, F. Akbarzadeh, and Z. Mohammadi. Row and columnmajorization on M_{n,m}. Linear Algebra and its Applications, 437(3):1025–1032, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002637.

Al-Ammari:2012:HMP

[AAT12a] Maha Al-Ammari and Françoise Tisseur. Hermitian matrix polynomials with real eigenvalues of definite type. Part I: Classification. Linear Algebra and its Applications, 436(10):3954–3973, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951000443X.

Al-Ammari:2012:STS

[AAT12b] Maha Al-Ammari and Françoise Tisseur. Standard triples of structured matrix polynomials. Linear Algebra and its Applications, 437(3): 817–834, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002510.

Akiyama:2012:PMP

[AB12] Shigeki Akiyama and Horst Brunotte. Primitive matrices over polynomial semirings. Linear Algebra and its Applications, 436(9):3568– 3596, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008652.

Alam:2011:CCN

[ABBO11] Rafikul Alam, Shreemayee Bora, Ralph Byers, and Michael L. Overton. Characterization and construction of the nearest defective matrix via coalescence of pseudospectral components. Linear Algebra and its Applications, 435(3):494–513, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abeles:2011:NCR

[Abe11] Francine F. Abeles. Nineteenth century roots of quasideterminants. Linear Algebra and its Applications, 435(5):1019–1024, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abeles:2014:CDD

[Abe14] Francine F. Abeles. Chiò’s and Dodgson’s determinantal identities. Linear Algebra and its Applications, 454(??):130–137, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002249.

Alaminos:2010:BMD

[ABEV10] J. Alaminos, M. Brešar, J. Extremera, and A. R. Villena. On bilinear maps determined by rank one idempotents. Linear Algebra and its Applications, 432(2–3):738–743, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Albuquerque:2014:PCT

[ABGPSS14] N. Albuquerque, L. Bernal-González, D. Pellegrino, and J. B. Seoane-Sepúlveda. Peano curves on topological vector spaces. Linear Algebra and its Applications, 460(??):81–96, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004790.

Araujo:2014:LSE

[ABK14] João Araújo, Wolfram Bentz, and Janusz Konieczny. The largest subsemilattices of the endomorphism monoid of an independence algebra. Linear Algebra and its Applications, 458(??):60–79, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003619.

Adiga:2010:SED

[ABS10] C. Adiga, R. Balakrishnan, and Wasin So. The skew energy of a digraph. Linear Algebra and its Applications, 432(7):1825–1835, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anuradha:2014:SSG

[ABS14] A. Anuradha, R. Balakrishnan, and Wasin So. Skew spectra of graphs without even cycles. Linear Algebra and its Applications, 444 (??):67–80, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007647.

Alaminos:2012:MPZ

[ABSV12] J. Alaminos, M. Bresar, S. Spenko, and A. R. Villena. Maps preserving zeros of a polynomial. Linear Algebra and its Applications, 436(7):2504–2512, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006975.

Cai:2014:QWM

[aCCS14] Yan an Cai, Yongsheng Cheng, and Ran Shen. Quasi-Whittaker modules for the Schrödinger algebra. Linear Algebra and its Applications, 463(??):16–32, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005734.

Abreu:2014:SAB

[ACDM14] Nair Abreu, Liliana Costa, Geir Dahl, and Enide Martins. The skeleton of acyclic Birkhoff polytopes. Linear Algebra and its Applications, 457(??):29–48, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003127.

Abreu:2011:BSL

[ACG^{+}11] Nair Abreu, Domingos M. Cardoso, Ivan Gutman, Enide A. Martins, and María Robbiano. Bounds for the signless Laplacian energy. Linear Algebra and its Applications, 435(10):2365–2374, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arias:2013:APO

[ACG13a] M. Laura Arias, Gustavo Corach, and M. Celeste Gonzalez. Additivity properties of operator ranges. Linear Algebra and its Applications, 439(11):3581–3590, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300579X.

Arias:2013:PPP

[ACG13b] M. Laura Arias, Gustavo Corach, and M. Celeste Gonzalez. Products of projections and positive operators. Linear Algebra and its Applications, 439(7):1730–1741, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003340.

Arias:2014:SPP

[ACG14] M. Laura Arias, Gustavo Corach, and M. Celeste Gonzalez. Saddle point problems, Bott–Duffin inverses, abstract splines and oblique projections. Linear Algebra and its Applications, 457(??): 61–75, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002869.

Abreu:2010:P

[ACGN10] Nair Abreu, Dragos Cvetković, Ivan Gutman, and Vladimir Nikiforov. Preface. Linear Algebra and its Applications, 432(9):2161–2162, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adashev:2014:NGZ

[ACGVK14] J. Q. Adashev, L. M. Camacho, S. Gómez-Vidal, and I. A. Karimjanov. Naturally graded Zinbiel algebras with nilindex n − 3. Linear Algebra and its Applications, 443(??):86–104, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007192.

Abreu:2012:LSL

[ACM^{+}12] Nair M. M. Abreu, Domingos M. Cardoso, Enide A. Martins, Maria Robbiano, and B. San Martiń. On the Laplacian and signless Laplacian spectrum of a graph with k pairwise co-neighbor vertices. Linear Algebra and its Applications, 437(9):2308– 2316, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003655.

Abreu:2014:CCI

[ACM14] Nair Abreu, Liliana Costa, and Enide Andrade Martins. On complementary coverage of Ω_{n}(T). Linear Algebra and its Applications, 442(??): 135–144, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005107.

Abreu:2014:P

[ACT14] Nair Abreu, Domingos Cardoso, and Vilmar Trevisan.

Preface. Linear Algebra and its Applications, 442 (??):1, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005545.

Adam:2014:DEC

[Ada14]  Michael Adam.  On the

distribution of eigenspaces in classical groups over finite rings. Linear Algebra and its Applications, 443(??): 50–65, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005727.

Alzer:2011:NPB

[AdF11] H. Alzer and C. M. da Fonseca. New permanental bounds for Ferrers matrices. Linear Algebra and its Applications, 435(11):2813–2827, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andelic:2011:NVS

[AdFM11] Milica An– delić, C. M. da Fonseca, and Ricardo Mamede. On the number of P -vertices of some graphs. Linear Algebra and its Applications, 434(2):514–525, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andelic:2011:BID

[AdFST11] M. An– delić, C. M. da Fonseca, S. K. Simić, and D. V. Tošić. On bounds for the index of double nested graphs. Linear Algebra and its Applications, 435(10):2475–2490, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Azizov:2013:CMD

[ADW13] T. Ya. Azizov, A. Dijksma, and G. Wanjala. Compressions of maximal dissipative and self-adjoint linear relations and of dilations. Linear Algebra and its Applications, 439(3):771–792, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002462.

Asadi:2012:CHM

[AF12] M. B. Asadi and M. Frank. A characterization of Hilbert C∗-modules as Banach modules with involution. Linear Algebra and its Applications, 437(2):722–725, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001954.

Abiad:2014:IAB

[AFHP14] A. Abiad, M. A. Fiol, W. H. Haemers, and G. Perarnau. An interlacing approach for bounding the sum of Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 448(??):11–21, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000688.

Abreu:2012:IAM

[AFLN12] M. Abreu, M. J. Funk, D. Labbate, and V. Napolitano. Invariant adjacency matrices of configuration graphs. Linear Algebra and its Applications, 437(8):2026–2037, October 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200420X.

Arias:2010:PSO

[AG10] M. Laura Arias and M. Celeste Gonzalez. Positive solutions to operator equations AXB = C. Linear Algebra and its Applications, 433(6): 1194–1202, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abdel-Ghaffar:2012:CMF

[AG12] Khaled A. S. Abdel-Ghaffar. Counting matrices over finite fields having a given number of rows of unit weight. Linear Algebra and its Applications, 436(7):2665–2669, April 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006458.

Adm:2013:ITN

[AG13] Mohammad Adm and Jürgen Garloff. Intervals of totally nonnegative matrices. Linear Algebra and its Applications, 439(12):3796–3806, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006411.

Agapito:2014:SPO

[Aga14] José Agapito. On symmetric polynomials with only real zeros and nonnegative γ-vectors. Linear Algebra and its Applications, 451 (??):260–289, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001621.

Allamigeon:2011:TPC

[AGK11] Xavier Allamigeon, Stéphane Gaubert, and Ricardo D. Katz. Tropical polar cones, hypergraph transversals, and mean payoff games. Linear Algebra and its Applications, 435(7):1549–1574, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akian:2014:TBE

[AGM14] Marianne Akian, Stéphane Gaubert, and Andrea Marchesini. Tropical bounds for eigenvalues of matrices. Linear Algebra and its Applications, 446(??):281–303, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300829X.

Akian:2011:BAM

[AGNS11] Marianne Akian, Stéphane Gaubert, Viorel Nitica, and Ivan Singer. Best approximation in max-plus semimodules. Linear Algebra and its Applications, 435(12):3261– 3296, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agore:2014:CCA

[Ago14] A. L. Agore. Classifying complements for associative algebras. Linear Algebra and its Applications, 446 (??):345–355, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000305.

Armario:2012:KMP

[AGPPF12] R. Armario, F. J. García-Pacheco, and F. J. Pérez-Fernández. On the Krein– Milman Property and the Bade Property. Linear Algebra and its Applications, 436(5):1489–1502, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100601X.

Armentia:2012:IPG

[AGV12] Gorka Armentia, Juan-Miguel Gracia, and Francisco E. Velasco. Identical pseudospectra of any geometric multiplicity. Linear Algebra and its Applications, 436(6): 1683–1688, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511000395.

Aouchiche:2010:SAC

[AH10] M. Aouchiche and P. Hansen. A survey of automated conjectures in spectral graph theory. Linear Algebra and its Applications, 432(9):2293–2322, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aouchiche:2013:TLD

[AH13a] Mustapha Aouchiche and Pierre Hansen. Two Laplacians for the distance matrix of a graph. Linear Algebra and its Applications, 439(1):21–33, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001614.

Audenaert:2013:MIB

[AH13b] Koenraad M. R. Audenaert and Fumio Hiai. On matrix inequalities between the power means: Counterexamples. Linear Algebra and its Applications, 439(5):1590– 1604, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002747.

Aouchiche:2014:DSG

[AH14] Mustapha Aouchiche and Pierre Hansen. Distance spectra of graphs: a survey. Linear Algebra and its Applications, 458(??):301–386, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003759.

Al-Homidan:2010:MCN

[AHAPP10] Suliman Al-Homidan, Mohammad M. Alshahrani, Cosmin G. Petra, and Florian A. Potra. Minimal condition number for positive definite Hankel matrices using semidefinite programming. Linear Algebra and its Applications, 433(6):1101– 1109, November 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aouchiche:2011:EVS

[AHL11] Mustapha Aouchiche, Pierre Hansen, and Claire Lucas. On the extremal values of the second largest Q-eigenvalue. Linear Algebra and its Applications, 435(10):2591–2606, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arav:2014:MRS

[AHL^{+}14] Marina Arav, Frank J. Hall, Zhongshan Li, Hein van der Holst, Lihua Zhang, and Wenyan Zhou. The minimum rank of a sign pattern matrix with a 1-separation. Linear Algebra and its Applications, 448(??):205–216, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000378.

Arav:2013:ISS

[AHLvdH13] Marina Arav, Frank J. Hall, Zhongshan Li, and Hein van der Holst. The inertia set of a signed graph. Linear Algebra and its Applications, 439(5):1506–1529, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300308X.

Aouchiche:2010:SUB

[AHS10] M. Aouchiche, P. Hansen, and D. Stevanović. A sharp upper bound on algebraic connectivity using domination number. Linear Algebra and its Applications, 432(11):2879–2893, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:2012:TVR

[AIL12] A. C. Antoulas, A. C. Ionita, and S. Lefteriu. On twovariable rational interpolation. Linear Algebra and its Applications, 436(8):2889– 2915, April 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005295.

Abaid:2012:CPC

[AIP12] Nicole Abaid, Irina Igel, and Maurizio Porfiri. On the consensus protocol of conspecific agents. Linear Algebra and its Applications, 437(1):221–235, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000948.

Abara:2013:NSE

[AiS13] Ma. Nerissa M. Abara and Ken ichi Shinoda. Number of solutions of equations of Weil type on finite symmetric matrices. Linear Algebra and its Applications, 438(11):4322–4334, June 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001031.

Ahmadi:2014:LAB

[AIS14] S. Ruhallah Ahmadi, Mohammad A. Izadi, and Fernando Szechtman. Lie algebras and bilinear forms in characteristic 2. Linear Algebra and its Applications, 448 (??):299–314, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000329.

Ayala:2013:LFM

[AJ13] Víctor Ayala and Ivan Jirón. Linear flows and Morse graphs: Topological consequences in low dimensions. Linear Algebra and its Applications, 439(8):2177–2194, October 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003960.

Abreu:2014:OTG

[AJRT14] Nair Abreu, Claudia Marcela Justel, Oscar Rojo, and Vilmar Trevisan. Ordering trees and graphs with few cycles by algebraic connectivity. Linear Algebra and its Applications, 458(??):429–453, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003814.

Atreas:2011:BIM

[AK11] Nikolaos Atreas and Costas Karanikas. Boolean invertible matrices identified from two permutations and their corresponding Haar-type matrices. Linear Algebra and its Applications, 435(1):95–105, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adm:2012:BMR

[AK12a] Mohammad Adm and Fuad Kittaneh. Bounds and majorization relations for the critical points of polynomials. Linear Algebra and its Applications, 436(7):2494– 2503, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007221.

Audeh:2012:SVI

[AK12b] Wasim Audeh and Fuad Kittaneh. Singular value inequalities for compact operators. Linear Algebra and its Applications, 437(10):2516– 2522, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004958.

Ayupov:2013:LDM

[AKA13] Shavkat Ayupov, Karimbergen Kudaybergenov, and Amir Alauadinov. 2-Local derivations on matrix algebras over commutative regular algebras. Linear Algebra and its Applications, 439 (5):1294–1311, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002759.

Aiura:2013:NMG

[AKM13] Daishi Aiura, Naonori Kakimura, and Kazuo Murota. On the number of matrices to generate a matrix ⋆-algebra over the real field. Linear Algebra and its Applications, 438 (3):1252–1266, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006465.

Akbari:2014:MLE

[AKM14] Saieed Akbari, Dariush Kiani, and Maryam Mirzakhah. The multiplicity of Laplacian eigenvalue two in unicyclic graphs. Linear Algebra and its Applications, 445 (??):18–28, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007209.

Afkhami:2012:GCG

[AKN12] Mojgan Afkhami, Kazem Khashyarmanesh, and Khosro Nafar. Generalized Cayley graphs associated to commutative rings. Linear Algebra and its Applications, 437(3):1040–1049, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002480.

Akbari:2013:GSF

[AKZ13] S. Akbari, M. Kano, and S. Zare. A generalization of 0-sum flows in graphs. Linear Algebra and its Applications, 438(9):3629–3634, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000517.

Ahues:2013:CNB

[AL13a] Mario Ahues and Balmohan V. Limaye. On condition numbers of a basis. Linear Algebra and its Applications, 439(11):3359–3377, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005703.

Alpay:2013:CCG

[AL13b] Daniel Alpay and Izchak Lewkowicz. Convex cones of generalized positive rational functions and the Nevanlinna–Pick interpolation. Linear Algebra and its Applications, 438(10): 3949–3966, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000870.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Avrachenkov:2013:APG

[AL13c] Konstantin E. Avrachenkov and Jean B. Lasserre. Analytic perturbation of generalized inverses. Linear Algebra and its Applications, 438 (4):1793–1813, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007348.

Alpay:2014:IPS

[AL14] Daniel Alpay and Izchak Lewkowicz. Interpolation by polynomials with symmetries. Linear Algebra and its Applications, 456(??): 64–81, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007677.

Alizadeh:2012:NRP

[Ali12]  R. Alizadeh.  Numerical

range and product of matrices. Linear Algebra and its Applications, 437(6):1422– 1425, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002984.

Alon:2014:EAR

[Alo14] Gil Alon. Eigenvalues of the Adin–Roichman matrices. Linear Algebra and its Applications, 450(??):280– 292, June 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001566.

Ai:2014:OBC

[ALPV14] Albert Ai, Alex Lapanowski, Yaniv Plan, and Roman Vershynin. One-bit compressed sensing with non-Gaussian measurements. Linear Algebra and its Applications, 441(??):222–239, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002450.

Andruchow:2012:CMH

[ALRV12] E. Andruchow, G. Larotonda, L. Recht, and A. Varela. A characterization of minimal Hermitian matrices. Linear Algebra and its Applications, 436(7):2366–2374, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006264.

Altenberg:2013:SCS

[Alt13] Lee Altenberg. A sharpened condition for strict log-convexity of the spectral radius via the bipartite graph. Linear Algebra and its Applications, 438 (9):3702–3718, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000578.

Liu:2013:SRD

[aLwW13] Chia an Liu and Chih wen Weng. Spectral radius and degree sequence of a graph. Linear Algebra and its Applications, 438(8):3511–3515, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000025.

Afshin:2013:PNH

[AM13a] Hamid Reza Afshin and Mohammad Ali Mehrjoofard. Polynomial numerical hulls of some normal matrices. Linear Algebra and its Applications, 439(8):2453–2467, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004242.

Agore:2013:UPL

[AM13b] A. L. Agore and G. Militaru. Unified products for Leibniz algebras. Applications. Linear Algebra and its Applications, 439(9):2609–2633, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004722.

Akbarbaglu:2013:LSC

[AM13c] I. Akbarbaglu and S. Maghsoudi. Large structures in certain subsets of Orlicz spaces. Linear Algebra and its Applications, 438 (11):4363–4373, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300102X.

Arias:2013:PIG

[AM13d] M. Laura Arias and Mostafa Mbekhta. A-partial isometries and generalized inverses. Linear Algebra and its Applications, 439(5):1286– 1293, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002735.

Asakly:2014:ECA

[AM14] Walaa Asakly and Toufik Mansour. Enumeration of compositions according to the sum of the values of the first letters of the occurrences of a 2-letter pattern. Linear Algebra and its Applications, 449(??):43–59, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000950.

Armandnejad:2014:LPG

[AMJ14] Ali Armandnejad, Saeedeh Mohtashami, and Mina Jamshidi. On linear preservers of gtridiagonal majorization on R^{n}. Linear Algebra and its Applications, 459(??): 145–153, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004194.

Abara:2010:OM

[AMP10] Ma. Nerissa M. Abara, Dennis I. Merino, and Agnes T. Paras. ϕ_{S}-Orthogonal matrices. Linear Algebra and its Applications, 432(11):2834– 2846, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agapito:2013:STR

[AMPT13] José Agapito, Ângela Mestre, Pasquale Petrullo, and Maria M. Torres. A symbolic treatment of Riordan arrays. Linear Algebra and its Applications, 439(7):1700–1715, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003194.

Alfakih:2013:AMU

[AN13] A. Y. Alfakih and Viet-Hang Nguyen. On affine motions and universal rigidity of tensegrity frameworks. Linear Algebra and its Applications, 439(10):3134–3147, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005156.

Ando:2013:UBI

[And13] Tsuyoshi Ando. Unbounded or bounded idempotent operators in Hilbert space. Linear Algebra and its Applications, 438(10):3769–3775, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004976.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Arashi:2011:PCB

[ANF11] M. Arashi, Daya K. Nagar, and Z. Farshidian Far. Properties of the complex bimatrix variate beta distribution. Linear Algebra and its Applications, 434(9):2018–2029, May 1, 2011. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Angerer:2013:SBH

[Ang13] Wolfgang P. Angerer. On the spectrum of banded Hankel matrices. Linear Algebra and its Applications, 439 (5):1496–1505, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003078.

Anonymous:2010:LEa

[Ano10a] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(1):ii–iii, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEb

[Ano10b] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(2–3):ii–iii, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEc

[Ano10c] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(4):ii–iii, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEd

[Ano10d] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(5):ii–iii, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEe

[Ano10e] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(6):ii–iii, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEf

[Ano10f] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(7):ii–iii, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEg

[Ano10g] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(8):ii–iii, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEh

[Ano10h] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(9):ii–iii, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEi

[Ano10i] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(10):ii–iii, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEj

[Ano10j] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(11):ii–iii, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEk

[Ano10k] Anonymous. Lists of Editors. Linear Algebra and its Applications, 432(12):ii–iii, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEl

[Ano10l] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(1):ii–iii, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEm

[Ano10m] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(2):ii–iii, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEn

[Ano10n] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(3):ii–iii, September 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEo

[Ano10o] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(4):ii–iii, October 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEp

[Ano10p] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(5):ii–iii, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEq

[Ano10q] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(6):ii–iii, November 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEr

[Ano10r] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(7):ii–iii, December 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEs

[Ano10s] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(8–10):ii–iii, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:LEt

[Ano10t] Anonymous. Lists of Editors. Linear Algebra and its Applications, 433(11–12): ii–iii, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2010:PIC

[Ano10u]  Anonymous.  Preface to

the 15th ILAS Conference Proceedings Cancún, México 2008. Linear Algebra and its Applications, 432(8):1865– 1866, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LR

[Ano11a] Anonymous. List of Reviewers. Linear Algebra and its Applications, 434(1):iv– viii, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEa

[Ano11b] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(1):ii–iii, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEb

[Ano11c] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(2):ii–iii, January 15, 2011. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEc

[Ano11d] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(3):ii–iii, February 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEd

[Ano11e] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(4):ii–iii, February 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEe

[Ano11f] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(5):ii–iii, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEf

[Ano11g] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(6):ii–iii, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEg

[Ano11h] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(7):ii–iii, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEh

[Ano11i] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(8):ii–iii, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEi

[Ano11j] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(9):ii–iii, May 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEj

[Ano11k] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(10):ii–iii, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEk

[Ano11l] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(11):ii–iii, June 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEl

[Ano11m] Anonymous. Lists of Editors. Linear Algebra and its Applications, 434(12):ii–iii, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEm

[Ano11n] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(1):ii–iii, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEn

[Ano11o] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(2):ii–iii, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEo

[Ano11p] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(3):ii–iii, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEp

[Ano11q] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(4):ii–iii, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEq

[Ano11r] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(5):ii–iii, September 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEr

[Ano11s] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(6):ii–iii, September 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEs

[Ano11t] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(7):ii–iii, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEt

[Ano11u] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(8):ii–iii, October 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEu

[Ano11v] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(9):ii–iii, November 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEv

[Ano11w] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(10):ii–iii, November 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEw

[Ano11x] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(11):ii–iii, December 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:LEx

[Ano11y] Anonymous. Lists of Editors. Linear Algebra and its Applications, 435(12):ii–iii, December 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:NEC

[Ano11z] Anonymous. Note from Editor in Chief. Linear Algebra and its Applications, 434(1): xv–xix, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2011:P

[Ano11-27] Anonymous. Preface. Linear Algebra and its Applications, 434(7):1611–1612, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2012:ASA

[Ano12a] Anonymous. Announcement: Sparse approximate solution of linear [equations]. Linear Algebra and its Applications, 436(4):??, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Anonymous:2012:D

[Ano12b]  Anonymous.  Dedication.

Linear Algebra and its Applications, 436(6):1541–1544, March 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000080.

Anonymous:2012:LR

[Ano12c] Anonymous. List of reviewers. Linear Algebra and its Applications, 436(1):??, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006896.

Anonymous:2012:LEa

[Ano12d] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(1):ii–iii, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005581.

Anonymous:2012:LEb

[Ano12e] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(2):ii–iii, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006665.

Anonymous:2012:LEc

[Ano12f] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(3):ii–iii, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100677X.

Anonymous:2012:LEd

[Ano12g] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(4):ii–iii, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008007.

Anonymous:2012:LEe

[Ano12h] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(5):ii–iii, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008494.

Anonymous:2012:LEf

[Ano12i] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(6):ii–iii, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000328.

Anonymous:2012:LEg

[Ano12j] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(7):ii–iii, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000663.

Anonymous:2012:LEh

[Ano12k] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(8):ii–iii, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001048.

Anonymous:2012:LEi

[Ano12l] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(9):ii–iii, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001267.

Anonymous:2012:LEj

[Ano12m] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(10):ii–iii, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001498.

Anonymous:2012:LEk

[Ano12n] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(11):ii–iii, June 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001887.

Anonymous:2012:LEl

[Ano12o] Anonymous. Lists of Editors. Linear Algebra and its Applications, 436(12):ii–iii, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002200.

Anonymous:2012:LEm

[Ano12p] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(1):ii–iii, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002376.

Anonymous:2012:LEn

[Ano12q] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(2):ii–iii, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002807.

Anonymous:2012:LEo

[Ano12r] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(3):ii–iii, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Anonymous:2012:LEp

[Ano12s] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(4):ii–iii, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200345X.

Anonymous:2012:LEq

[Ano12t] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(5):ii–iii, September 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003722.

Anonymous:2012:LEr

[Ano12u] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(6):ii– iii, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004016.

Anonymous:2012:LEs

[Ano12v] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(7):ii–iii, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004569.

Anonymous:2012:LEt

[Ano12w] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(8):ii–iii, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004818.

Anonymous:2012:LEu

[Ano12x] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(9):ii–iii, November 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005204.

Anonymous:2012:LEv

[Ano12y] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(10): ii–iii, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005435.

Anonymous:2012:LEw

[Ano12z] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(11):ii– iii, December 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Anonymous:2012:LEx

[Ano12-27] Anonymous. Lists of Editors. Linear Algebra and its Applications, 437(12): ii–iii, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006283.

Anonymous:2012:NE

[Ano12-28] Anonymous. Note from Editors. Linear Algebra and its Applications, 436(1):xi– xvi, January 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006884.

Anonymous:2012:P

[Ano12-29] Anonymous. Preface. Linear Algebra and its Applications, 436(10):3793–3800, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001450.

Anonymous:2012:PPC

[Ano12-30] Anonymous. Preface to the proceedings of the Coimbra meeting on 0-1 matrix theory and related topics. Linear Algebra and its Applications, 436(4):789–790, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006835.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Linear Algebra and its Applications, 439(7):ii–iii, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004400.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Linear Algebra and its Applications, 439(8):ii–iii, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004837.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Linear Algebra and its Applications, 439(9):ii–iii, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379513005260.

Anonymous:2013:HBC

[Ano13d] Anonymous. Harm Bart curriculum vitae. Linear Algebra and its Applications, 439 (3):513–514, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002425.

Anonymous:2013:HBL

[Ano13e] Anonymous. Harm Bart list of publications. Linear Algebra and its Applications, 439 (3):515–519, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002413.

Anonymous:2013:LPA

[Ano13f] Anonymous. List of publications (Abraham Berman). Linear Algebra and its Applications, 438(10):3724–3734, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512007525.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Anonymous:2013:LEa

[Ano13g] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(1):ii–iii, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200688X.

Anonymous:2013:LEb

[Ano13h] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(2):ii–iii, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007069.

Anonymous:2013:LEc

[Ano13i] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(3):ii–iii, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007926.

Anonymous:2013:LEd

[Ano13j] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(4):ii–iii, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008087.

Anonymous:2013:LEe

[Ano13k] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(5):ii–iii, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008543.

Anonymous:2013:LEf

[Ano13l] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(6):ii–iii, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000244.

Anonymous:2013:LEg

[Ano13m] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(7):ii–iii, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000633.

Anonymous:2013:LEh

[Ano13n] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(8):ii–iii, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001183.

Anonymous:2013:LEi

[Ano13o] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(9):ii–iii, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001390.

Anonymous:2013:LEj

[Ano13p] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(10):ii–iii, May 15, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001699.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Anonymous:2013:LEk

[Ano13q] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(11):ii–iii, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001997.

Anonymous:2013:LEl

[Ano13r] Anonymous. Lists of Editors. Linear Algebra and its Applications, 438(12):ii–iii, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002310.

Anonymous:2013:LEm

[Ano13s] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(1):ii–iii, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002577.

Anonymous:2013:LEn

[Ano13t] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(2):ii–iii, July 15, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002899.

Anonymous:2013:LEo

[Ano13u] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(3):ii–iii, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003248.

Anonymous:2013:LEp

[Ano13v] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(4):ii–iii, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003479.

Anonymous:2013:LEq

[Ano13w] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(5):ii–iii, September 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003753.

Anonymous:2013:LEr

[Ano13x] Anonymous. Lists of Editors. Linear Algebra and its Applications, 439(6):ii– iii, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004023.

Anonymous:2013:MHM

[Ano13y] Anonymous. Minisymposium in honor of Miroslav Fiedler. Linear Algebra and its Applications, 439(4):809, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002231.

Anonymous:2013:MG

[Ano13z] Anonymous. Moshe Goldberg. Linear Algebra and its Applications, 438(10): 3735–3738, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007537.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Anonymous:2013:RL

[Ano13-27] Anonymous. Raphael Loewy. Linear Algebra and its Applications, 438(10):3739–3744, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512007549.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Linear Algebra and its Applications, 455(??):ii–iii, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003292.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Linear Algebra and its Applications, 456(??):ii–iii, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514003486.

Anonymous:2014:EBc

[Ano14c] Anonymous. Editorial Board. Linear Algebra and its Applications, 457(??):ii–iii, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514004273.

Anonymous:2014:EBd

[Ano14d] Anonymous. Editorial Board. Linear Algebra and its Applications, 458(??):ii–iii, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004455.

Anonymous:2014:EBe

[Ano14e] Anonymous. Editorial Board. Linear Algebra and its Applications, 459(??):ii–iii, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005151.

Anonymous:2014:EBf

[Ano14f] Anonymous. Editorial Board. Linear Algebra and its Applications, 460(??):ii–iii, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514005345.

Anonymous:2014:EBg

[Ano14g] Anonymous. Editorial Board. Linear Algebra and its Applications, 461(??):ii–iii, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514005564.

Anonymous:2014:EBh

[Ano14h] Anonymous. Editorial Board. Linear Algebra and its Applications, 463(??):ii–iii, December 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514006211.

Almeida:2013:NCS

[ANP13] P. Almeida, D. Napp, and R. Pinto. A new class of superregular matrices and MDP convolutional codes. Linear Algebra and its Applications, 439(7):2145–2157, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004126.

Alvarez-Nodarse:2012:SPC

[ÁNPQ12] R. Álvarez-Nodarse, J. Petronilho, and N. R. Quintero. Spectral properties of certain tridiagonal matrices. Linear Algebra and its Applications, 436(3):682–698, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005854.

Abdollahi:2014:CG

[AO14] Alireza Abdollahi and Mohammad Reza Oboudi. Cospectrality of graphs. Linear Algebra and its Applications, 451 (??):169–181, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001384.

Ashraf:2013:SSL

[AOTR13] F. Ashraf, G. R. Omidi, and B. Tayfeh-Rezaie. On the sum of signless Laplacian eigenvalues of a graph. Linear Algebra and its Applications, 438(11):4539–4546, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000852.

Araujo:2014:LBC

[AP14] Gustavo Araújo and Daniel Pellegrino. Lower bounds for the constants of the Hardy– Littlewood inequalities. Linear Algebra and its Applications, 463(??):10–15, December 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005722.

Albeverio:2010:DSO

[AR10] Sergio Albeverio and Slavik Rabanovich. Decomposition of a scalar operator into a product of unitary operators with two points in spectrum. Linear Algebra and its Applications, 433(6):1127–1137, November 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arambasic:2012:BJO

[AR12] Ljiljana Arambasić and Rajna Rajić. The Birkhoff–James orthogonality in Hilbert C^{∗}- modules. Linear Algebra and its Applications, 437(7): 1913–1929, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003631.

Arashi:2012:PTS

[Ara12]  M. Arashi.  Preliminary

test and Stein estimations in simultaneous linear equations. Linear Algebra and its Applications, 436(5):1195– 1211, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005775.

Andreani:2013:GSS

[ART13] Roberto Andreani, Marcos Raydan, and Pablo Tarazaga. On the geometrical structure of symmetric matrices. Linear Algebra and its Applications, 438(3):1201–1214, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006118.

Al-Rashed:2011:NOS

[ARZ11] Maryam H. A. Al-Rashed and Boguslaw Zegarliński. Noncommutative Orlicz spaces associated to a state II. Linear Algebra and its Applications, 435(12):2999–3013, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amri:2010:SAC

[AS10] Amine Amri and Alberto Seeger. Spectral analysis of coupled linear complementarity problems. Linear Algebra and its Applications, 432(10): 2507–2523, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahmed:2012:MIO

[AS12a] Ould Ahmed Mahmoud Sid Ahmed and Adel Saddi. A-m-Isometric operators in semi-Hilbertian spaces. Linear Algebra and its Applications, 436(10):3930–3942, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379510004751.

Alfsen:2012:FDC

[AS12b] Erik Alfsen and Fred Shultz. Finding decompositions of a class of separable states. Linear Algebra and its Applications, 437(10):2613–2629, November 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004648.

Alves:2012:PPN

[AS12c] João Ferreira Alves and Luís Silva. Periodic paths on nonautonomous graphs. Linear Algebra and its Applications, 437(3):1003–1015, August 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002625.

Ames:2012:NDF

[AS12d] Brendan P. W. Ames and Hristo S. Sendov. A new derivation of a formula by Kato. Linear Algebra and its Applications, 436(3):722– 730, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100574X.

Alexandersson:2014:AMS

[AS14] Per Alexandersson and Boris Shapiro. Around a multivariate Schmidt-Spitzer theorem. Linear Algebra and its Applications, 446(??):356– 368, April 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000226.

Araujo:2011:SPM

[AT11] C. Mendes Araújo and Juan R. Torregrosa. Sign pattern matrices that admit P_{0}- matrices. Linear Algebra and its Applications, 435(8):2046– 2053, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2014:SRM

[AT14a] C. Mendes Araújo and Juan R. Torregrosa. Some results on B-matrices and doubly B-matrices. Linear Algebra and its Applications, 459(??):101–120, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004170.

Atay:2014:SNL

[AT14b] Fatihcan M. Atay and Hande Tunçel. On the spectrum of the normalized Laplacian for signed graphs: Interlacing, contraction, and replication. Linear Algebra and its Applications, 442(??): 165–177, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005211.

Arashi:2012:SRV

[ATS12] M. Arashi, S. M. M. Tabatabaey, and H. Soleimani. Simple regression in view of elliptical models. Linear Algebra and its Applications, 437(7): 1675–1691, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003588.

Audenaert:2010:SRH

[Aud10a] Koenraad M. R. Audenaert. Spectral radius of Hadamard product versus conventional product for non-negative matrices. Linear Algebra and its Applications, 432(1):366– 368, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2010:VBA

[Aud10b] Koenraad M. R. Audenaert. Variance bounds, with an application to norm bounds for commutators. Linear Algebra and its Applications, 432(5): 1126–1143, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2012:TIC

[Aud12] Koenraad M. R. Audenaert. Trace inequalities for completely monotone functions and Bernstein functions. Linear Algebra and its Applications, 437(2):601–611, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002339.

Audenaert:2013:ALT

[Aud13a] Koenraad M. R. Audenaert. An Araki–Lieb–Thirring inequality for geometrically concave and geometrically convex functions. Linear Algebra and its Applications, 438(8):3454–3462, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000463.

Audenaert:2013:BGM

[Aud13b] Koenraad M. R. Audenaert. In-betweenness, a geometrical monotonicity property for operator means. Linear Algebra and its Applications, 438 (4):1769–1778, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001868.

Audenaert:2013:SMN

[Aud13c] Koenraad M. R. Audenaert. Schur multiplier norms for Loewner matrices. Linear Algebra and its Applications, 439(9):2598–2608, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300503X.

Alfaro:2012:SGC

[AV12] Carlos A. Alfaro and Carlos E. Valencia. On the sandpile group of the cone of a graph. Linear Algebra and its Applications, 436(5): 1154–1176, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100543X.

Anh:2011:AGG

[ÁvW11] P. N. Ánh and L. van Wyk. Automorphism groups of generalized triangular matrix rings. Linear Algebra and its Applications, 434(4):1018– 1026, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anh:2013:IBS

[ÁvW13] P. N. Ánh and L. van Wyk. Isomorphisms between strongly triangular matrix rings. Linear Algebra and its Applications, 438(11): 4374–4381, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300092X.

Appleby:2010:IMP

[AW10] Glenn D. Appleby and Tamsen Whitehead. Invariants of matrix pairs over discrete valuation rings and Littlewood– Richardson fillings. Linear Algebra and its Applications, 432(5):1277–1298, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andriantiana:2011:UGL

[AW11] Eric Ould Dadah Andriantiana and Stephan Wagner. Unicyclic graphs with large energy. Linear Algebra and its Applications, 435 (6):1399–1414, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abo:2013:WPS

[AW13a] Hirotachi Abo and Jia Wan. On Waring’s problem for systems of skew-symmetric forms. Linear Algebra and its Applications, 439(8):2330– 2349, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004515.

Andriantiana:2013:SMT

[AW13b] Eric Ould Dadah Andriantiana and Stephan Wagner. Spectral moments of trees with given degree sequence. Linear Algebra and its Applications, 439(12):3980– 4002, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006393.

Astuti:2013:CHS

[AW13c] Pudji Astuti and Harald K. Wimmer. Characteristic and hyperinvariant subspaces over the field GF (2). Linear Algebra and its Applications, 438 (4):1551–1563, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

Alfakih:2013:AMB

[AY13] A. Y. Alfakih and Yinyu Ye. On affine motions and bar frameworks in general position. Linear Algebra and its Applications, 438 (1):31–36, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006556.

Burer:2009:DBD

[BAD09] Samuel Burer, Kurt M. Anstreicher, and Mirjam Dür. The difference between 5 × 5 doubly nonnegative and completely positive matrices. Linear Algebra and its Applications, 431(9):1539–1552, October 1, 2009. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See note [DA10].

Bai:2011:DPS

[Bai11] Zhaojun Bai. Dedication to Pete Stewart on the occasion of his 70th birthday. Linear Algebra and its Applications, 435(3):421, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baird:2014:IPF

[Bai14] Paul Baird. An invariance property for frameworks in Euclidean space. Linear Algebra and its Applications, 440(??):243–265, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006745.

Bajo:2014:IQO

[Baj14]  Ignacio Bajo.  Invariant

quadrics and orbits for a family of rational systems of difference equations. Linear Algebra and its Applications, 449(??):500–511, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001104.

Balaji:2010:CPM

[Bal10] R. Balaji. Characterization of Q-property for multiplicative transformations in semidefinite linear complementarity problems. Linear Algebra and its Applications, 432(11): 2754–2763, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balan:2012:SSTa

[Bal12a] Vladimir Balan. Spectra of symmetric tensors and m-root Finsler models. Linear Algebra and its Applications, 436(1):152–162, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Balan:2012:SSTb

[Bal12b] Vladimir Balan. Spectra of symmetric tensors and m-root Finsler models. Linear Algebra and its Applications, 436(5):1061–1071, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004964.

See erratum [Vla12].

Ballico:2014:TRR

[Bal14] Edoardo Ballico. On the typical rank of real bivariate polynomials. Linear Algebra and its Applications, 452 (??):263–269, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400192X.

Bang:2013:GDR

[Ban13a] S. Bang. Geometric distanceregular graphs without 4claws. Linear Algebra and its Applications, 438(1):37– 46, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005629.

Banica:2013:DGF

[Ban13b] Teodor Banica. The defect of generalized Fourier matrices. Linear Algebra and its Applications, 438(9): 3667–3688, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000736.

Ben-Ari:2012:IFT

[BANP12] Iddo Ben-Ari, Michael Neumann, and Olga Pryporova. Inequalities for functions of transition matrices. Linear Algebra and its Applications, 436(2):335–348, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003648.

Bapat:2010:CCS

[Bap10] R. B. Bapat. On a conjecture concerning spanning tree invariants and loop systems. Linear Algebra and its Applications, 433(8–10):1642–1645, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2013:MCM

[Bap13a] R. B. Bapat. On minors of the compound matrix of a Laplacian. Linear Algebra and its Applications, 439 (11):3378–3386, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Bapat:2013:AMT

[Bap13b] R. B. Bapat. On the adjacency matrix of a threshold graph. Linear Algebra and its Applications, 439 (10):3008–3015, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005041.

Barria:2010:SCS

[Bar10a] José Barría. On the strong closure of the simultaneous similarity orbit of a pair of finite rank operators. Linear Algebra and its Applications, 432(8):1873–1877, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2010:RMM

[Bar10b]  Harm Bart.  Review of

Matrix Mathematics, Second Edition by Dennis S. Bernstein, Princeton University Press, Princeton and Oxford (2009). xxxix + 1139 pp., Paperback, ISBN-13 978-0-691-14039-1. Linear Algebra and its Applications, 433(11– 12):2272–2274, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bartl:2012:STC

[Bar12a] David Bartl. Separation theorems for convex polytopes and finitely-generated cones derived from theorems of the alternative. Linear Algebra and its Applications, 436 (9):3784–3789, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100824X.

Barvinok:2012:MPR

[Bar12b] Alexander Barvinok. Matrices with prescribed row and column sums. Linear Algebra and its Applications, 436(4):820–844, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510005859.

Barria:2013:SCS

[Bar13a] José Barría. The strong closure of the simultaneous similarity orbit of some pairs of finite rank operators. Linear Algebra and its Applications, 438(6):2770–2776, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008233.

Bart:2013:CC

[Bar13b] Harm Bart. The Chemnitz connection. Linear Algebra and its Applications, 439 (3):520–523, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001092.

Batzke:2014:GRO

[Bat14] Leonhard Batzke. Generic rank-one perturbations of structured regular matrix pencils. Linear Algebra and its Applications, 458(??): 638–670, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004108.

Bauer:2012:NGL

[Bau12] Frank Bauer. Normalized graph Laplacians for directed graphs. Linear Algebra and its Applications, 436(11):4193–4222, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000845.

Benzi:2010:QRB

[BB10] Michele Benzi and Paola Boito. Quadrature rule-based bounds for functions of adjacency matrices. Linear Algebra and its Applications, 433(3):637–652, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2011:CEG

[BB11a] LeRoy B. Beasley and David E. Brown. Cycle extendability in graphs and digraphs. Linear Algebra and its Applications, 435(7):1513–1519, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Budnitska:2011:CAO

[BB11b] Tetiana Budnitska and Nadiya Budnitska. Classification of affine operators up to biregular conjugacy. Linear Algebra and its Applications, 434 (5):1195–1199, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bach:2013:NDF

[BB13a] Eric Bach and Andrew Bridy. On the number of distinct functional graphs of affinelinear transformations over finite fields. Linear Algebra and its Applications, 439 (5):1312–1320, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002760.

Bai:2013:NBT

[BB13b] Zheng-Jian Bai and Zhong-Zhi Bai. On nonsingularity of block two-by-two matrices. Linear Algebra and its Applications, 439(8): 2388–2404, October 15, 2013.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003923.

Banaji:2013:GTC

[BB13c] Murad Banaji and Andrew Burbanks. A graph-theoretic condition for irreducibility of a set of cone preserving matrices. Linear Algebra and its Applications, 438(11):4103–4113, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000918.

BarroseSa:2013:FCH

[BB13d] Nuno Barros e Sá and Ingemar Bengtsson. Families of complex Hadamard matrices. Linear Algebra and its Applications, 438(7): 2929–2957, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007598.

Benzi:2014:DPF

[BB14] Michele Benzi and Paola Boito. Decay properties for functions of matrices over C^{∗}-algebras. Linear Algebra and its Applications, 456 (??):174–198, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300726X.

Barrett:2014:PRC

[BBC^{+}14] Wayne Barrett, Steve Butler, Minerva Catral, Shaun M. Fallat, H. Tracy Hall, Leslie Hogben, P. van den Driessche, and Michael Young. The principal rank characteristic sequence over various fields. Linear Algebra and its Applications, 459(??): 222–236, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004145.

Ballico:2013:GSI

[BBCC13] Edoardo Ballico, Alessandra Bernardi, Maria Virginia Catalisano, and Luca Chiantini. Grassmann secants, identifiability, and linear systems of tensors. Linear Algebra and its Applications, 438(1):121–135, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006076.

Bai:2011:PSI

[BBD^{+}11] Zhong-Zhi Bai, Michele Benzi, Iain S. Duff, Andreas Frommer, and Zhong-Ci Shi. Preface [Special issue: Devoted to the 2nd NASC 08 Conference in Nanjing (NSC), November 2–5, 2008]. Linear Algebra and its Applications, 434(11): 2223–2224, June 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2013:P

[BBD^{+}13] Albrecht Böttcher, Richard Brualdi, Harry Dym, Rien Kaashoek, and Andre Ran. Preface. Linear Algebra and its Applications, 439(3): 511–512, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002449.

Baragana:2012:CBS

[BBdH12] Itziar Baragaña, M. Asunción Beitia, and Inmaculada de Hoyos. The change of the Brunovsky structure of a controllable pair under one column perturbation in a particular case. Linear Algebra and its Applications, 436(6):1572– 1588, March 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005398.

Baragana:2013:CFA

[BBdH13] Itziar Baragaña, M. Asunción Beitia, and Inmaculada de Hoyos. Canonical form associated with the problem of perturbation of one column of a controllable pair. Linear Algebra and its Applications, 438 (4):1587–1626, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006604.

Bini:2010:FIQ

[BBE^{+}10] D. A. Bini, P. Boito, Y. Eidelman, L. Gemignani, and I. Gohberg. A fast implicit QR eigenvalue algorithm for companion matrices. Linear Algebra and its Applications, 432 (8):2006–2031, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2010:ZFP

[BBF^{+}10] Francesco Barioli, Wayne Barrett, Shaun M. Fallat, H. Tracy Hall, Leslie Hogben, Bryan Shader, P. van den Driessche, and Hein van der Holst. Zero forcing parameters and minimum rank problems. Linear Algebra and its Applications, 433(2):401– 411, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2012:GCC

[BBF^{+}12] Francesco Barioli, Wayne Barrett, Shaun M. Fallat, H. Tracy Hall, Leslie Hogben, and Hein van der Holst. On the graph complement conjecture for minimum rank. Linear Algebra and its Applications, 436(12):

4373–4391, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510006713.

Bini:2013:PIC

[BBG^{+}13] Dario A. Bini, Albrecht Böttcher, Luca Gemignani, Leslie Hogben, and Françoise Tisseur. Preface to the 16th ILAS Conference Proceedings, Pisa 2010. Linear Algebra and its Applications, 438 (4):1495–1496, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200047X.

Bachman:2014:FUT

[BBG14] Dale Bachman, Nicholas R. Baeth, and James Gossell. Factorizations of upper triangular matrices. Linear Algebra and its Applications, 450 (??):138–157, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001050.

Bogoya:2012:EHT

[BBGM12] J. M. Bogoya, A. Böttcher, S. M. Grudsky, and E. A. Maksimenko. Eigenvectors of Hessenberg Toeplitz matrices and a problem by Dai, Geary, and Kadanoff. Linear Algebra and its Applications, 436(9):3480–3492, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008299.

Barrett:2012:CIS

[BBH^{+}12] Wayne Barrett, Steve Butler, H. Tracy Hall, John Sinkovic, Wasin So, Colin Starr, and Amy Yielding. Computing inertia sets using atoms. Linear Algebra and its Applications, 436(12):4489–4502, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006094.

Bernardi:2014:CDN

[BBM14] Alessandra Bernardi, Jérôme Brachat, and Bernard Mourrain. A comparison of different notions of ranks of symmetric tensors. Linear Algebra and its Applications, 460(??):205–230, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400487X.

Bremner:2012:CHC

[BBS12a] Murray R. Bremner, Mikelis G. Bickis, and Mohsen Soltanifar. Cayley’s hyperdeterminant: a combinatorial approach via representation theory. Linear Algebra and its Applications, 437(1):94–112, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200119X.

Bruen:2012:CSM

[BBS12b] Aiden A. Bruen, Trevor C. Bruen, and Robert Silverman. A characterization of (λ, 2)- stable (0, 1) matrices. Linear Algebra and its Applications, 436(4):814–819, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002667.

Barrett:2013:P

[BBSMT13] Wayne Barrett, Richard A. Brualdi, Naomi Shaked-Monderer, and Eitan Tadmor. Preface. Linear Algebra and its Applications, 438(10):3723, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007550.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Bockting-Conrad:2012:TCO

[BC12a] Sarah Bockting-Conrad. Two commuting operators associated with a tridiagonal pair. Linear Algebra and its Applications, 437(1):242–270, July 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001371.

Borobia:2012:MCP

[BC12b] Alberto Borobia and Roberto Canogar. Matrix completion problems over integral domains: The case with a diagonal of prescribed blocks. Linear Algebra and its Applications, 436(1):222–236, January 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100499X.

Borobia:2012:NAM

[BC12c] Alberto Borobia and Roberto Canogar. Nonsingular ACImatrices over integral domains. Linear Algebra and its Applications, 436(11): 4311–4316, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000766.

Borobia:2013:CFR

[BC13] Alberto Borobia and Roberto Canogar. Characterization of full rank ACI-matrices over fields. Linear Algebra and its Applications, 439(12):3752– 3762, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Bockting-Conrad:2014:TPR

[BC14a] Sarah Bockting-Conrad. Tridiagonal pairs of q-Racah type, the double lowering operator ψ, and the quantum algebra U_{q}(∫↕_{∈}). Linear Algebra and its Applications, 445(??):256–279, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008057.

Borobia:2014:NMP

[BC14b] Alberto Borobia and Roberto Canogar. A note on matrices with prescribed off-diagonal submatrix and characteristic polynomial. Linear Algebra and its Applications, 458(??):99–107, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400370X.

Brualdi:2014:NSR

[BC14c] Richard A. Brualdi and Joshua Cooper. Note on the spectral radius of alternating sign matrices. Linear Algebra and its Applications, 442(??):99–105, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005016.

Benhida:2010:SPK

[BCD10] C. Benhida, M. C. Câmara, and C. Diogo. Some properties of the kernel and the cokernel of Toeplitz operators with matrix symbols. Linear Algebra and its Applications, 432(1):307–317, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Breaz:2013:NCM

[BCDM13] S. Breaz, G. Calugareanu, P. Danchev, and T. Micu. Nilclean matrix rings. Linear Algebra and its Applications, 439(10):3115–3119, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005417.

Beitia:2011:RPO

[BCdP11] M. A. Beitia, A. Compta, I. Hoyos de, and M. Peña. Realizations of perturbations of an observable pair with prescribed indices. Linear Algebra and its Applications, 434 (5):1325–1335, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendito:2010:KIN

[BCE^{+}10] E. Bendito, A. Carmona, A. M. Encinas, J. M. Gesto, and M. Mitjana. Kirchhoff Indexes of a network. Linear Algebra and its Applications, 432(9):2278–2292, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendito:2010:GIS

[BCEM10] E. Bendito, A. Carmona, A. M. Encinas, and M. Mitjana. Generalized inverses of symmetric M-matrices. Linear Algebra and its Applications, 432(9):2438–2454, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendito:2012:MIP

[BCEM12] E. Bendito, A. Carmona, A. M. Encinas, and M. Mitjana. The M-matrix inverse problem for singular and symmetric Jacobi matrices. Linear Algebra and its Applications, 436(5): 1090–1098, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004940.

Boza:2012:GAV

[BCF12] Luis Boza, Alfonso Carriazo, and Luis M. Fernández. Graphs associated with vector spaces of even dimension: a link with differential geometry. Linear Algebra and its Applications, 437(1):60–76, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001206.

Bueno:2014:SSLa

[BCF14] M. I. Bueno, K. Curlett, and S. Furtado. Structured strong linearizations from Fiedler pencils with repetition I. Linear Algebra and its Applications, 460(??): 51–80, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400490X.

Botelho:2012:MSS

[BCFP12] Geraldo Botelho, Daniel Cariello, Vinićius V. Fávaro, and Daniel Pellegrino. Maximal spaceability in sequence spaces. Linear Algebra and its Applications, 437(12):2978– 2985, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005149.

Brachat:2010:STD

[BCMT10] Jerome Brachat, Pierre Comon, Bernard Mourrain, and Elias Tsigaridas. Symmetric tensor decomposition. Linear Algebra and its Applications, 433 (11–12):1851–1872, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2010:MCP

[BCŠ10] Alberto Borobia, Roberto Canogar, and Helena Šmigoc. A matrix completion problem over integral domains: the case with 2n − 3 prescribed entries. Linear Algebra and its Applications, 433(3):606– 617, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bermudez:2013:CRL

[BCS13a] J. M. Ancochea Bermúdez and R. Campoamor-Stursberg. On a complete rigid Leibniz non-Lie algebra in arbitrary dimension. Linear Algebra and its Applications, 438(8):3397–3407, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000542.

Bonnabel:2013:RPG

[BCS13b] Silvère Bonnabel, Anne Collard, and Rodolphe Sepulchre. Rank-preserving geometric means of positive semidefinite matrices. Linear Algebra and its Applications, 438(8):3202–3216, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008646.

Bockting-Conrad:2014:AID

[BCT14] Sarah Bockting-Conrad and Paul Terwilliger. The algebra U_{q}(∫↕_{∈}) in disguise. Linear Algebra and its Applications, 459(??):548–585, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004728.

Bai:2012:MSH

[BCY12] Zheng-Jian Bai, Mei-Xiang Chen, and Jin-Ku Yang. A multi-step hybrid method for multi-input partial quadratic eigenvalue assignment with time delay. Linear Algebra and its Applications, 437(7): 1658–1669, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003345.

Bai:2012:SNL

[BD12a] Zhaofang Bai and Shuanping Du. The structure of nonlinear Lie derivation on von Neumann algebras. Linear Algebra and its Applications, 436(7):2701–2708, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007506.

Brualdi:2012:MCI

[BD12b] Richard A. Brualdi and Geir Dahl. Majorization classes of integral matrices. Linear Algebra and its Applications, 436(4):802–813, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510006129.

Bevilacqua:2013:CRA

[BD13] Roberto Bevilacqua and Gianna M. Del Corso. A condensed representation of almost normal matrices. Linear Algebra and its Applications, 438(11):4408–4425, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001055.

Batselier:2013:GAF

[BDD13] Kim Batselier, Philippe Dreesen, and Bart De Moor. A geometrical approach to finding multivariate approximate LCMs and GCDs. Linear Algebra and its Applications, 438 (9):3618–3628, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300044X.

Batselier:2014:NSM

[BDD14] Kim Batselier, Philippe Dreesen, and Bart De Moor. On the null spaces of the Macaulay matrix. Linear Algebra and its Applications, 460(??): 259–289, November 1, 2014.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004868.

Bartolone:2011:NLA

[BDF11] C. Bartolone, A. Di Bartolo, and G. Falcone. Nilpotent Lie algebras with 2-dimensional commutator ideals. Linear Algebra and its Applications, 434(3):650–656, February 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2011:IEP

[BdFdP11] Natália Bebiano, C. M. da Fonseca, and João da Providência. An inverse eigenvalue problem for periodic Jacobi matrices in Minkowski spaces. Linear Algebra and its Applications, 435(8):2033–2045, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botelho:2011:SBQ

[BDFP11] G. Botelho, D. Diniz, V. V. Fávaro, and D. Pellegrino. Spaceability in Banach and quasi-Banach sequence spaces. Linear Algebra and its Applications, 434(5):1255–1260, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bevilacqua:2013:BTR

[BDG13] Roberto Bevilacqua, Gianna M. Del Corso, and Luca Gemignani. Block tridiagonal reduction of perturbed normal and rank structured matrices. Linear Algebra and its Applications, 439(11): 3505–3517, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005892.

Behn:2012:SNT

[BDH^{+}12] Antonio Behn, Kenneth R. Driessel, Irvin Roy Hentzel, Kent A. Vander Velden, and James Wilson. Some nilpotent, tridiagonal matrices with a special sign pattern. Linear Algebra and its Applications, 436(12): 4446–4450, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005027.

Bart:2010:MIG

[BDK^{+}10] Harm Bart, Harry Dym, Rien Kaashoek, Peter Lancaster, Alexander Markus, and Leiba Rodman. In memoriam: Israel Gohberg, August 23, 1928–October 12, 2009. Linear Algebra and its Applications, 433(5):877–892, October 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bilge:2011:MLS

[BDK11] Ayşe Hümeyra Bilge, Tekin Dereli, and Şahin Koçak. Maximal Linear Subspaces of Strong Self-Dual 2-forms and the Bonan 4-form. Linear Algebra and its Applications, 434(5):1200–1214, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benito:2013:LEN

[BdlC13] Pilar Benito and Daniel dela Concepción. On Levi extensions of nilpotent Lie algebras. Linear Algebra and its Applications, 439(5):1441– 1457, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003030.

Benito:2014:EFN

[BdlC14] Pilar Benito and Daniel dela Concepción. On extensions of free nilpotent Lie algebras of type 2. Linear Algebra and its Applications, 457 (??):363–382, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003425.

Bodine:2012:SPR

[BDM^{+}12] E. Bodine, L. Deaett, J. J. McDonald, D. D. Olesky, and P. van den Driessche. Sign patterns that require or allow particular refined inertias. Linear Algebra and its Applications, 437(9):2228– 2242, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003667.

Brualdi:2012:PRC

[BDOvdD12] R. A. Brualdi, L. Deaett, D. D. Olesky, and P. van den Driessche. The principal rank characteristic sequence of a real symmetric matrix. Linear Algebra and its Applications, 436(7):2137–2155, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007567.

Bebiano:2013:ISP

[BdP13] Natália Bebiano and João da Providência. Inverse spectral problems for structured pseudo-symmetric matrices. Linear Algebra and its Applications, 438(10): 4062–4074, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005642.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Brandts:2010:CEB

[BdS10] Jan H. Brandts and Ricardo Reis da Silva. Computable eigenvalue bounds for rank-k perturbations. Linear Algebra and its Applications, 432(12):3100–3116, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhattacharyya:2013:DSC

[BDS13] Tirthankar Bhattacharyya, Bata Krishna Das, and Santanu Sarkar. The defect sequence for contractive tuples. Linear Algebra and its Applications, 438(1):315– 330, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006003.

Burde:2012:AAL

[BDV12] Dietrich Burde, Karel Dekimpe, and Kim Vercammen. Affine actions on Lie groups and post-Lie algebra structures. Linear Algebra and its Applications, 437(5):1250–1263, September 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002716.

Bremner:2010:AQA

[BE10] Murray R. Bremner and Hader A. Elgendy. Alternating quaternary algebra structures on irreducible representations of ∫↕_{∈}(C). Linear Algebra and its Applications, 433(8–10):1686–1705, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benkovic:2012:MLD

[BE12] Dominik Benkovic and Daniel Eremita. Multiplicative Lie n-derivations of triangular rings. Linear Algebra and its Applications, 436(11): 4223–4240, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000869.

Beasley:2012:INV

[Bea12] LeRoy B. Beasley. Isolation number versus Boolean rank. Linear Algebra and its Applications, 436(9):3469– 3474, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008305.

Behrend:2013:FPM

[Beh13] Roger E. Behrend. Fractional perfect b-matching polytopes I: General theory. Linear Algebra and its Applications, 439(12):3822–3858, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006204.

Benzi:2013:RHA

[BEK13] Michele Benzi, Ernesto Estrada, and Christine Klymko. Ranking hubs and authorities using matrix functions. Linear Algebra and its Applications, 438(5):2447–2474, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200732X.

Belardo:2014:BLE

[Bel14] Francesco Belardo. Balancedness and the least eigenvalue of Laplacian of signed graphs. Linear Algebra and its Applications, 446(??):133– 147, April 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000160.

Benner:2012:MBD

[BELK12] Peter Benner, Mark Embree, Richard B. Lehoucq, and C. T. Kelley. A mathematical biography of Danny C. Sorensen. Linear Algebra and its Applications, 436 (8):2717–2724, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000961.

Bahrami:2012:LPM

[BEM12a] F. Bahrami, A. Bayati Eshkaftaki, and S. M. Manjegani.

Linear preservers of majorization on ℓ^{p}(I). Linear Algebra and its Applications, 436 (9):3177–3195, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007130.

Bahrami:2012:MCL

[BEM12b] F. Bahrami, A. Bayati Eshkaftaki, and S. M. Manjegani. Majorization on ℓ^{∞} and on its closed linear subspace ⌋, and their linear preservers. Linear Algebra and its Applications, 437(9):2340–2358, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004363.

Beneduci:2010:SMP

[Ben10] Roberto Beneduci. Stochastic matrices and a property of the infinite sequences of linear functionals. Linear Algebra and its Applications, 433(6): 1224–1239, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benkovic:2011:GLD

[Ben11] Dominik Benkovič. Generalized Lie derivations on triangular algebras. Linear Algebra and its Applications, 434(6): 1532–1544, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendaoud:2013:PLS

[Ben13] Mohamed Bendaoud. Preservers of local spectra of matrix sums. Linear Algebra and its Applications, 438(5):2500–2507, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007586.

Benvenuti:2014:GRS

[Ben14a] Luca Benvenuti. A geometrical representation of the spectra of four dimensional nonnegative matrices. Linear Algebra and its Applications, 445(??):162–180, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008276.

Benzi:2014:NWE

[Ben14b] Michele Benzi. A note on walk entropies in graphs. Linear Algebra and its Applications, 445(??):395–399, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008227.

Bernstein:2009:MMT

[Ber09] Dennis S. Bernstein. Matrix Mathematics: Theory, Facts, and Formulas. Princeton University Press, Princeton, NJ, USA, second edition, 2009. ISBN 0-691-13287-9 (hardcover), 0-691-14039-1 (paperback), 1-4008-3334-5 (e-book). xlii + 1139 pp. LCCN QA188 .B475 2009; QA188 .B475X 2009 (LC). URL http://

www.jstor.org/stable/10.
2307/j.ctt7t833;   http:
//www.loc.gov/catdir/toc/
ecip0826/2008036257.html.

Berget:2013:EST

[Ber13a] Andrew Berget. Equality of symmetrized tensors and the flag variety. Linear Algebra and its Applications, 438(2):658–662, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002606.

Bergqvist:2013:EPT

[Ber13b] Göran Bergqvist. Exact probabilities for typical ranks of 2 × 2 × 2 and 3 × 3 × 2 tensors. Linear Algebra and its Applications, 438(2): 663–667, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001753.

Braun:2013:LCP

[BEV13] Michael Braun, Tuvi Etzion, and Alexander Vardy. Linearity and complements in projective space. Linear Algebra and its Applications, 438 (1):57–70, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006209.

Beyn:2012:IMS

[Bey12] Wolf-Jürgen Beyn. An integral method for solving nonlinear eigenvalue problems. Linear Algebra and its Applications, 436(10): 3839–3863, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511002540.

Brualdi:2014:BOT

[BF14a] Richard A. Brualdi and Eliseu Fritscher. Bruhat order of tournaments. Linear Algebra and its Applications, 458(??): 261–279, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003681.

Brualdi:2014:HTX

[BF14b] Richard A. Brualdi and Eliseu Fritscher. Hankel and Toeplitz X-rays of permutations. Linear Algebra and its Applications, 449 (??):350–380, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001049.

Bueno:2014:SSLb

[BF14c] M. I. Bueno and S. Furtado. Structured strong linearizations from Fiedler pencils with repetition II. Linear Algebra and its Applications, 463 (??):282–321, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005886.

Burde:2014:JJA

[BF14d] Dietrich Burde and Alice Fialowski. Jacobi–Jordan algebras. Linear Algebra and its Applications, 459(??): 586–594, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004844.

Bollhofer:2013:PIC

[BFBD13] Matthias Bollhöfer, Heike Faßbender, Ravi Bapat, and Froilán M. Dopico. Preface to the 17th ILAS Conference ‘Pure and Applied Linear Algebra: the New Generation’ Proceedings, Braunschweig, Germany, 2011. Linear Algebra and its Applications, 439(4):807–808, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300222X.

Bebiano:2010:ACP

[BFdP10] N. Bebiano, S. Furtado, and J. da Providência. Analogs of Cauchy–Poincaré and Fan– Pall interlacing theorems for J-Hermitian and J-normal matrices. Linear Algebra and its Applications, 433(1):80– 90, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2011:EPS

[BFdP11] N. Bebiano, S. Furtado, and J. da Providência. On the eigenvalues of principal submatrices of J-normal matrices. Linear Algebra and its Applications, 435(12):3101– 3114, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2012:LQJ

[BFdP12] N. Bebiano, S. Furtado, and J. da Providência. Large a quasi-Jacobi form for Jnormal matrices and inverse eigenvalue problems. Linear Algebra and its Applications, 436(6):1739–1753, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004666.

Bolten:2011:AMM

[BFF^{+}11] Matthias Bolten, Stephanie Friedhoff, Andreas Frommer, Matthias Heming, and Karsten Kahl. Algebraic multigrid methods for Laplacians of graphs. Linear Algebra and its Applications, 434 (11):2225–2243, June 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2012:TAM

[BFH^{+}12] Richard Brualdi, Shaun Fallat, Leslie Hogben, Bryan Shader, and P. van den Driessche. Theory and applications of matrices described by patterns: Preface and workshop report. Linear Algebra and its Applications, 436(12):4349–4351, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000894.

Barik:2011:HDG

[BFK11] S. Barik, S. Fallat, and S. Kirkland. On Hadamard diagonalizable graphs. Linear Algebra and its Applications, 435(8):1885–1902, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bondarenko:2013:SSU

[BFK^{+}13] Vitalij M. Bondarenko, Vyacheslav Futorny, Tatiana Klimchuk, Vladimir V. Sergeichuk, and Kostyantyn Yusenko. Systems of subspaces of a unitary space. Linear Algebra and its Applications, 438(5):2561–2573, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007689.

Botelho:2012:SE

[BFPSS12] G. Botelho, V. V. Fávaro, D. Pellegrino, and J. B. Seoane-Sepúlveda. L_{p}[0, 1]\∪_{q>p} L_{q}[0, 1] is spaceable for every p > 0. Linear Algebra and its Applications, 436 (9):2963–2965, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008640.

Baglama:2014:ADN

[BFRR14] J. Baglama, C. Fenu, L. Reichel, and G. Rodriguez. Analysis of directed networks via partial singular value decomposition and Gauss quadrature. Linear Algebra and its Applications, 456(??):93– 121, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003097.

Benner:2011:HKS

[BFS11] Peter Benner, Heike Faßbender, and Martin Stoll. A Hamiltonian Krylov–Schurtype method based on the symplectic Lanczos process. Linear Algebra and its Applications, 435(3):578–600, August 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bruhn:2011:BCU

[BG11] Henning Bruhn and Agelos Georgakopoulos. Bases and closures under infinite sums. Linear Algebra and its Applications, 435(8):2007–2018, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2012:NIR

[BG12a] Rajendra Bhatia and Priyanka Grover. Norm inequalities related to the matrix geometric mean. Linear Algebra and its Applications, 437 (2):726–733, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001814.

Bialas:2012:FRC

[BG12b] Stanislaw Bialas and Michal Góra. A few results concerning the Hurwitz stability of polytopes of complex polynomials. Linear Algebra and its Applications, 436(5): 1177–1188, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005878.

Bartoszewicz:2013:ALV

[BG13] Artur Bartoszewicz and Szymon Glab. Additivity and lineability in vector spaces. Linear Algebra and its Applications, 439(7):2123–2130, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003959.

Bapat:2014:ITM

[BG14a] R. B. Bapat and E. Ghorbani. Inverses of triangular matrices and bipartite graphs. Linear Algebra and its Applications, 447(??):68–73, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001857.

Benkovic:2014:GDU

[BG14b] Dominik Benkovic and Mateja Grasic. Generalized derivations on unital algebras determined by action on zero products. Linear Algebra and its Applications, 445(??): 347–368, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008082.

Bru:2012:GM

[BGH12] Rafael Bru, Isabel Giménez, and Apostolos Hadjidimos. Is A ∈ C^{n,n} a general Hmatrix? Linear Algebra and its Applications, 436(2): 364–380, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002126.

Ballani:2013:BBA

[BGK13] Jonas Ballani, Lars Grasedyck, and Melanie Kluge. Black box approximation of tensors in hierarchical Tucker format. Linear Algebra and its Applications, 438(2): 639–657, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100591X.

Butkovic:2011:P

[BGL^{+}11] Peter Butkovic, Alexander Guterman, Jean Jacques Loiseau, William M. McEneaney, and Edouard Wagneur. Preface. Linear Algebra and its Applications, 435(7): 1491–1493, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bartoszewicz:2011:ANC

[BGP11] Artur Bartoszewicz, Szymon Głab̧, and Tadeusz Poreda.

On algebrability of nonabsolutely convergent series. Linear Algebra and its Applications, 435(5):1025–1028, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bartoszewicz:2013:LFL

[BGP13] Artur Bartoszewicz, Szymon Glab̧, and Adam Paszkiewicz. Large free linear algebras of real and complex functions. Linear Algebra and its Applications, 438(9):3689– 3701, May 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300075X.

Baker:2012:LRI

[BGV12] C. G. Baker, K. A. Gallivan, and P. Van Dooren. Lowrank incremental methods for computing dominant singular subspaces. Linear Algebra and its Applications, 436(8):2866–2888, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005301.

Beattie:2012:ISI

[BGW12] Christopher Beattie, Serkan Gugercin, and Sarah Wyatt. Inexact solves in interpolatory model reduction. Linear Algebra and its Applications, 436(8):2916–2943, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005271.

Blunck:2010:PLJ

[BH10] Andrea Blunck and Hans Havlicek. Projective lines over Jordan systems and geometry of Hermitian matrices. Linear Algebra and its Applications, 433(3):672–680, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Briec:2011:SCS

[BH11a] Walter Briec and Charles Horvath. On the separation of convex sets in some idempotent semimodules. Linear Algebra and its Applications, 435(7):1542–1548, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brijder:2011:NIP

[BH11b] Robert Brijder and Hendrik Jan Hoogeboom. Nullity invariance for pivot and the interlace polynomial. Linear Algebra and its Applications, 435(2):277–288, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barry:2012:RAL

[BH12] P. Barry and A. Hennessy. Riordan arrays and the LDU decomposition of symmetric Toeplitz plus Hankel matrices. Linear Algebra and its Applications, 437(6):1380– 1393, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200300X.

Bottcher:2013:WHS

[BH13a] Albrecht Böttcher and Martin Halwass. Wiener–Hopf and spectral factorization of real polynomials by Newton’s method. Linear Algebra and its Applications, 438(12):4760–4805, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001511.

Bremner:2013:KTE

[BH13b] Murray R. Bremner and Jiaxiong Hu. On Kruskal’s theorem that every 3 × 3 × 3 array has rank at most 5. Linear Algebra and its Applications, 439(2):401–421, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002164.

Bai:2014:OEC

[BH14a] Zhong-Zhi Bai and Apostolos Hadjidimos. Optimization of extrapolated Cayley transform with non-Hermitian positive definite matrix. Linear Algebra and its Applications, 463(??):322–339, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005655.

Bourin:2014:JMI

[BH14b] Jean-Christophe Bourin and Fumio Hiai. Jensen and Minkowski inequalities for operator means and antinorms. Linear Algebra and its Applications, 456(??): 22–53, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003218.

Bhatia:2013:BD

[Bha13] Rajendra Bhatia. The bipolar decomposition. Linear Algebra and its Applications, 439 (10):3031–3037, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005612.

Bulo:2012:ECI

[BHAP12] Samuel Rota Bulò, Edwin R. Hancock, Furqan Aziz, and Marcello Pelillo. Efficient computation of Ihara coefficients using the Bell polynomial recursion. Linear Algebra and its Applications, 436(5):1436–1441, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006008.

Behrndt:2012:LPM

[BHDW12] Jussi Behrndt, Seppo Hassi, Henk De Snoo, and Rudi Wietsma. Limit properties of monotone matrix functions. Linear Algebra and its Applications, 436(5):935– 953, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004290.

Bishop:2010:IFS

[BHKL10] Shannon Bishop, Christopher Heil, Yoo Young Koo, and Jae Kun Lim. Invariances of frame sequences under perturbations. Linear Algebra and its Applications, 432(6):1501– 1514, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Betsumiya:2013:UBC

[BHKM13] Koichi Betsumiya, Mitsugu Hirasaka, Takao Komatsu, and Akihiro Munemasa. Upper bounds on cyclotomic numbers. Linear Algebra and its Applications, 438(1): 111–120, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Barriere:2013:FVL

[BHMO13] Lali Barrière, Clemens Huemer, Dieter Mitsche, and David Orden. On the Fiedler value of large planar graphs. Linear Algebra and its Applications, 439(7): 2070–2084, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003868.

Brunsch:2012:DIA

[BHMR12] Thomas Brunsch, Laurent Hardouin, Carlos Andrey Maia, and Jörg Raisch. Duality and interval analysis over idempotent semirings. Linear Algebra and its Applications, 437(10):2436–2454, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004739.

Barrett:2011:ISJ

[BHvdH11] Wayne Barrett, H. Tracy Hall, and Hein van der Holst. The inertia set of the join of graphs. Linear Algebra and its Applications, 434(10): 2197–2203, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2010:SNB

[BHZ10] Richard A. Brualdi, Zejun Huang, and Xingzhi Zhan. Singular, nonsingular, and bounded rank completions of ACI-matrices. Linear Algebra and its Applications, 433(7): 1452–1462, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bahri:2012:NHD

[BI12] Y. Bahri and A. Ilahi. A note on Homotopic Deviation versus perturbation. Linear Algebra and its Applications, 436(7):2353–2358, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007026.

Bini:2013:CKM

[BI13] Dario A. Bini and Bruno Iannazzo. Computing the Karcher mean of symmetric positive definite matrices. Linear Algebra and its Applications, 438(4):1700– 1710, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006616.

Bier:2013:SEE

[Bie13] Agnieszka Bier. On solvability of Engel equations in the group of triangular matrices over a field. Linear Algebra and its Applications, 438(5):2320–2330, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007197.

Bier:2014:CPI

[Bie14] Agnieszka Bier. Commutators and powers of infinite unitriangular matrices. Linear Algebra and its Applications, 457(??):162–178, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003115.

Berger:2012:QWF

[BIT12] Thomas Berger, Achim Ilchmann, and Stephan Trenn. The quasi-Weierstraß form for regular matrix pencils. Linear Algebra and its Applications, 436(10):4052–4069, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379509006776.

Botelho:2010:ARS

[BJ10a] Fernanda Botelho and James Jamison. Algebraic reflexivity of sets of bounded operators on vector valued Lipschitz functions. Linear Algebra and its Applications, 432 (12):3337–3342, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botelho:2010:EOA

[BJ10b] Fernanda Botelho and James Jamison. Elementary operators and the Aluthge transform. Linear Algebra and its Applications, 432(1):275– 282, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botelho:2010:IPE

[BJ10c] Fernanda Botelho and James Jamison. Isometric properties of elementary operators. Linear Algebra and its Applications, 432(1):357–365, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botelho:2011:PAI

[BJ11] Fernanda Botelho and James Jamison. Projections as averages of isometries on minimal norm ideals. Linear Algebra and its Applications, 435 (6):1344–1355, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bouhamidi:2013:CIN

[BJ13] Abderrahman Bouhamidi and Khalide Jbilou. On the convergence of inexact Newton methods for discretetime algebraic Riccati equations. Linear Algebra and its Applications, 439(7):2057– 2069, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300387X.

Bouhamidi:2011:ETM

[BJRS11] A. Bouhamidi, K. Jbilou, L. Reichel, and H. Sadok. An extrapolated TSVD method for linear discrete ill-posed problems with Kronecker structure. Linear Algebra and its Applications, 434(7):1677– 1688, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botelho:2012:SIA

[BJZ12] Fernanda Botelho, James Jamison, and Bentuo Zheng. Strict isometries of arbitrary orders. Linear Algebra and its Applications, 436 (9):3303–3314, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007658.

Brualdi:2010:TNM

[BK10] Richard A. Brualdi and Steve Kirkland. Totally nonnegative (0,1)-matrices. Linear Algebra and its Applications, 432(7): 1650–1662, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bessenyei:2011:FEG

[BK11a] Mihály Bessenyei and Csaba G. Kézi. Functional equations and group substitutions. Linear Algebra and its Applications, 434(6):1525– 1531, March 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Betcke:2011:PER

[BK11b] Timo Betcke and Daniel Kressner. Perturbation, extraction and refinement of invariant pairs for matrix polynomials. Linear Algebra and its Applications, 435(3):514– 536, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buchholzer:2012:BEE

[BK12] Hannes Buchholzer and Christian Kanzow. Bounds for the extremal eigenvalues of a class of symmetric tridiagonal matrices with applications. Linear Algebra and its Applications, 436(7):1837–1849, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007075.

Barrett:2013:DMR

[BKM^{+}13] Wayne Barrett, Mark Kempton, Nicole Malloy, Curtis Nelson, William Sexton, and John Sinkovic. Decompositions of minimum rank matrices. Linear Algebra and its Applications, 438(10): 3913–3948, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000067.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Brualdi:2012:TRM

[BKMS12] Richard A. Brualdi, Kathleen P. Kiernan, Seth A. Meyer, and Michael W. Schroeder. On the t-term rank of a matrix. Linear Algebra and its Applications, 436(6):1632–1643, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004368.

Brualdi:2013:PAS

[BKMS13] Richard A. Brualdi, Kathleen P. Kiernan, Seth A. Meyer, and Michael W. Schroeder. Patterns of alternating sign matrices. Linear Algebra and its Applications, 438(10):3967–3990, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002078.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Bapat:2012:WDG

[BKP12] R. B. Bapat, D. Kalita, and S. Pati. On weighted directed graphs. Linear Algebra and its Applications, 436 (1):99–111, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004848.

Ball:2014:RCI

[BKV14] Joseph A. Ball and Dmitry S. Kaliuzhnyi-Verbovetskyi. Rational Cayley inner Herglotz– Agler functions: Positivekernel decompositions and transfer-function realizations. Linear Algebra and its Applications, 456(??):138–156, September 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006435.

Bistritz:2010:BRU

[BL10a] Yuval Bistritz and Alexander Lifshitz. Bounds for resultants of univariate and bivariate polynomials. Linear Algebra and its Applications, 432 (8):1995–2005, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biyikoglu:2010:STM

[BL10b] Türker Bıyıkoḡlu and Josef Leydold. Semiregular trees with minimal Laplacian spectral radius. Linear Algebra and its Applications, 432(9): 2335–2341, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boutin:2011:MRP

[BL11] Olivier Boutin and Anne L’Anton. Modelling routing phenomenon with bounds estimation in dioids. Linear Algebra and its Applications, 435(7):1520–1541, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biyikoglu:2012:GGO

[BL12] Türker Biyikoglu and Josef Leydold. Graphs of given order and size and minimum algebraic connectivity. Linear Algebra and its Applications, 436(7):2067–2077, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006628.

Benitez:2013:SPM

[BL13a] Julio Benítez and Xiaoji Liu. A short proof of a matrix decomposition with applications. Linear Algebra and its Applications, 438(3): 1398–1414, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006994.

Buczynski:2013:RTG

[BL13b] Jaroslaw Buczyński and J. M. Landsberg. Ranks of tensors and a generalization of secant varieties. Linear Algebra and its Applications, 438(2):668–689, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003308.

Bourin:2014:DSP

[BL14] Jean-Christophe Bourin and Eun-Young Lee. Direct sums of positive semi-definite matrices. Linear Algebra and its Applications, 463(??):273– 281, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005977.

Bebiano:2012:RHK

[BLdP12] N. Bebiano, R. Lemos, and J. da Providência. On a reverse Heinz–Kato–Furuta inequality. Linear Algebra and its Applications, 437(7): 1892–1905, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003576.

Bebiano:2010:TIL

[BLdPS10] N. Bebiano, R. Lemos, J. da Providência, and G. Soares. Trace inequalities for logarithms and powers of J-Hermitian matrices. Linear Algebra and its Applications, 432(12):3172–3182, July 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2012:DLP

[BLL12] Jean-Christophe Bourin, Eun-Young Lee, and Minghua Lin. On a decomposition lemma for positive semi-definite block-matrices. Linear Algebra and its Applications, 437(7):1906–1912, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200314X.

Bourin:2013:PMP

[BLL13] Jean-Christophe Bourin, Eun-Young Lee, and Minghua Lin. Positive matrices partitioned into a small number of Hermitian blocks. Linear Algebra and its Applications, 438(5):2591–2598, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200763X.

Bonilla:2013:SSI

[BLLM13] M. Bonilla, G. Lebret, J. J. Loiseau, and M. Malabre. Simultaneous state and input reachability for linear time invariant systems. Linear Algebra and its Applications, 439 (5):1415–1440, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003029.

Bai:2011:SSL

[BLLX11] Zhaojun Bai, Cherung Lee, Ren-Cang Li, and Shufang Xu. Stable solutions of linear systems involving long chain of matrix multiplications. Linear Algebra and its Applications, 435(3):659– 673, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blomeling:2012:MLS

[Blö12]  Frank Blömeling.  Multi-

level substructuring combined with model order reduction methods. Linear Algebra and its Applications, 436(10): 3864–3882, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001741.

Brunnock:2014:SRN

[BLS14] R. Brunnock, M. C. Lettington, and K. M. Schmidt. On square roots and norms of matrices with symmetry properties. Linear Algebra and its Applications, 459(??): 175–207, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004236.

Blumthaler:2010:FO

[Blu10] Ingrid Blumthaler. Functional T -observers. Linear Algebra and its Applications, 432(6): 1560–1577, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bayat:2010:PSF

[BM10] M. Bayat and B. Mehri. Periodic solutions of first-order autonomous quasi-linear systems. Linear Algebra and its Applications, 432(2–3):485– 492, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balmaceda:2012:LTL

[BM12a] Jose Maria P. Balmaceda and Jryl P. Maralit. Leonard triples from Leonard pairs constructed from the standard basis of the Lie algebra sl_{2}. Linear Algebra and its Applications, 437(7): 1961–1977, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200328X.

Bobok:2012:MSS

[BM12b] Jozef Bobok and Ivo Marek. Monotony of solutions of some difference and differential equations. Linear Algebra and its Applications, 436(10):3919–3929, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379510004799.

Benner:2013:CAS

[BM13a] Peter Benner and Thomas Mach. The LR Cholesky algorithm for symmetric hierarchical matrices. Linear Algebra and its Applications, 439(4):1150–1166, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001845.

Bernik:2013:LAA

[BM13b] Janez Bernik and Mitja Mastnak. Lie algebras acting semitransitively. Linear Algebra and its Applications, 438(6):2777–2792, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008385.

Bremner:2013:SIP

[BM13c] Murray R. Bremner and Sara Madariaga. Special identities for the pre-Jordan product in the free dendriform algebra. Linear Algebra and its Applications, 439(2):435–454, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001936.

Butkovic:2013:IES

[BM13d] Peter Butkovic and Marie MacCaig. On integer eigenvectors and subeigenvectors in the max-plus algebra. Linear Algebra and its Applications, 438(8):3408–3424, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000037.

Batzke:2014:IEP

[BM14a] Leonhard Batzke and Christian Mehl. On the inverse eigenvalue problem for T -alternating and T - palindromic matrix polynomials. Linear Algebra and its Applications, 452(??):172– 191, July 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001815.

Beasley:2014:PD

[BM14b] LeRoy B. Beasley and Sarah Mousley. k-Primitivity of digraphs. Linear Algebra and its Applications, 449 (??):512–519, May 15, 2014.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001062.

Burgos:2012:SPP

[BMGMC12] M. Burgos, A. C. Márquez-García, and A. Morales-Campoy. Strongly preserver problems in Banach algebras and C^{∗}-algebras. Linear Algebra and its Applications, 437 (5):1183–1193, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002959.

Branquinho:2012:RAO

[BMM12] A. Branquinho, F. Marcellán, and A. Mendes. Relative asymptotics for orthogonal matrix polynomials. Linear Algebra and its Applications, 437(7):1458–1481, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003084.

Beagley:2013:BMS

[BMN^{+}13a] Jonathan Beagley, Lon H. Mitchell, Sivaram K. Narayan, Eileen Radzwion, Sara P. Rimer, Rachael Tomasino, Jennifer Wolfe, and Andrew M. Zimmer. Bounds for minimum semidefinite rank from superpositions and cutsets. Linear Algebra and its Applications, 438(10): 4041–4061, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005381.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Bermudez:2013:PI

[BMN13b] Teresa Bermúdez, Antonio Martinón, and Juan Agustín Noda. Products of misometries. Linear Algebra and its Applications, 438 (1):80–86, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200537X.

Bernik:2011:PMS

[BMR11] Janez Bernik, Mitja Mastnak, and Heydar Radjavi. Positivity and matrix semigroups. Linear Algebra and its Applications, 434(3):801– 812, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belardo:2010:CAC

[BMS10] Francesco Belardo, Enzo M. Li Marzi, and Slobodan K. Simić. Combinatorial approach for computing the characteristic polynomial of a matrix. Linear Algebra and its Applications, 433(8–10):1513– 1523, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourhim:2014:NMP

[BMS14a] A. Bourhim, J. Mashreghi, and A. Stepanyan. Nonlinear maps preserving the minimum and surjectivity moduli. Linear Algebra and its Applications, 463(??):171– 189, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005746.

Brualdi:2014:MIM

[BMS14b] Richard A. Brualdi, Volker Mehrmann, and Peter Semrl. Memorial issue for Michael Neumann and Uriel Rothblum. Linear Algebra and its Applications, 447(??):1, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000901.

Belardo:2010:SRU

[BMSW10] Francesco Belardo, Enzo M. Li Marzi, Slobodan K. Simić, and Jianfeng Wang. On the spectral radius of unicyclic graphs with prescribed degree sequence. Linear Algebra and its Applications, 432(9):2323– 2334, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-

3795 (print), 1873-1856 (electronic).

Belardo:2011:GWS

[BMSW11] Francesco Belardo, Enzo M. Li Marzi, Slobodan K. Simić, and Jianfeng Wang. Graphs whose signless Laplacian spectral radius does not exceed the Hoffman limit value. Linear Algebra and its Applications, 435(11):2913– 2920, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bauschke:2010:RAP

[BMW10] Heinz H. Bauschke, Sarah M. Moffat, and Xianfu Wang. The resolvent average for positive semidefinite matrices. Linear Algebra and its Applications, 432(7):1757– 1771, March 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2013:SPE

[BN13] Dario A. Bini and Vanni Noferini. Solving polynomial eigenvalue problems by means of the Ehrlich–Aberth method. Linear Algebra and its Applications, 439(4): 1130–1149, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001559.

Bose:2011:DSR

[BNP11] Surya Sekhar Bose, Milan Nath, and Somnath Paul. Distance spectral radius of graphs with r pendent vertices. Linear Algebra and its Applications, 435(11):2828– 2836, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bose:2012:DSR

[BNP12] Surya Sekhar Bose, Milan Nath, and Somnath Paul. On the distance spectral radius of cacti. Linear Algebra and its Applications, 437 (9):2128–2141, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004387.

Bose:2013:MDS

[BNP13] Surya Sekhar Bose, Milan Nath, and Somnath Paul. On the maximal distance spectral radius of graphs without a pendent vertex. Linear Algebra and its Applications, 438(11):4260–4278, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000815.

Brambilla:2011:PPI

[BO11] Maria Chiara Brambilla and Giorgio Ottaviani. On partial polynomial interpolation.

Linear Algebra and its Applications, 435(6):1415–1445, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blumthaler:2012:DPP

[BO12] Ingrid Blumthaler and Ulrich Oberst. Design, parametrization, and pole placement of stabilizing output feedback compensators via injective cogenerator quotient signal modules. Linear Algebra and its Applications, 436 (5):963–1000, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004216.

Bodmann:2013:RFF

[Bod13] Bernhard G. Bodmann. Random fusion frames are nearly equiangular and tight. Linear Algebra and its Applications, 439(5):1401–1414, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002863.

Bojicic:2013:HTS

[Boj13]  Radica Bojicić.  Hankel

transform of a series reversion of a certain rational function. Linear Algebra and its Applications, 438 (11):4237–4248, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000864.

Boshnakov:2011:FSO

[Bos11] Georgi N. Boshnakov. On first and second order stationarity of random coefficient models. Linear Algebra and its Applications, 434(2):415–423, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:2010:PTU

[Bot10a] J. D. Botha. Products of two unipotent matrices of index 2. Linear Algebra and its Applications, 433(7):1447– 1451, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:2010:SAN

[Bot10b] J. D. Botha. Spectrally arbitrary, nonderogatory factorization over a general field. Linear Algebra and its Applications, 433(1):1–11, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:2012:STS

[Bot12] J. D. Botha. Sums of two square-zero matrices over an arbitrary field. Linear Algebra and its Applications, 436(3):516–524, February 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005234.

Bottcher:2013:OTA

[Böt13] Albrecht Böttcher. An operator theoretic approach to the brickwork Ising model with second-neighbor interactions. Linear Algebra and its Applications, 439(3): 675–685, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007392.

Botha:2014:MDI

[Bot14] J. D. Botha. Matrix division with an idempotent divisor or quotient. Linear Algebra and its Applications, 446(??):71–90, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000032.

Bourgeois:2010:ME

[Bou10a] Gerald Bourgeois. The matrix equation log(XY) = log(X)+ log(Y). Linear Algebra and its Applications, 432(8):1878– 1884, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourhim:2010:SLM

[Bou10b] Abdellatif Bourhim. Surjective linear maps preserving local spectra. Linear Algebra and its Applications, 432 (1):383–393, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourgeois:2011:HSM

[Bou11] Gerald Bourgeois. How to solve the matrix equation XA − AX = f(X). Linear Algebra and its Applications, 434(3):657–668, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourgeois:2013:CIS

[Bou13] Gérald Bourgeois. Common invariant subspace and commuting matrices. Linear Algebra and its Applications, 438(7):3030–3038, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008488.

Bella:2010:SFG

[BOZ10] T. Bella, V. Olshevsky, and P. Zhlobich. Signal flow graph approach to inversion of (H, m)-quasiseparable-Vandermonde matrices and new filter structures. Linear Algebra and its Applications, 432(8):2032–2051, April 1, 2010. CODEN LAAPAW.

Bella:2011:QAF

[BOZ11a] T. Bella, V. Olshevsky, and P. Zhlobich. A quasiseparable approach to five-diagonal CMV and Fiedler matrices. Linear Algebra and its Applications, 434(4):957–976, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bella:2011:RQA

[BOZ11b] T. Bella, V. Olshevsky, and P. Zhlobich. Reprint of: A quasiseparable approach to five-diagonal CMV and Fiedler matrices. Linear Algebra and its Applications, 434 (7):1773–1792, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bozzo:2013:MPI

[Boz13] Enrico Bozzo. The Moore– Penrose inverse of the normalized graph Laplacian. Linear Algebra and its Applications, 439(10):3038–3043, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005582.

Basic:2010:PST

[BP10] Milan Bašić and Marko D. Petković. Perfect state transfer in integral circulant graphs of non-square-free order. Linear Algebra and its Applications, 433(1):149–163, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Besenyei:2011:CPM

[BP11] Ádám Besenyei and Dénes Petz. Completely positive mappings and mean matrices. Linear Algebra and its Applications, 435(5):984–997, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barreras:2012:CJS

[BP12] A. Barreras and J. M. Pena. Characterizations of Jacobi sign regular matrices. Linear Algebra and its Applications, 436(2):381–388, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001005.

Besenyei:2013:PSE

[BP13] Ádám Besenyei and Dénes Petz. Partial subadditivity of entropies. Linear Algebra and its Applications, 439 (10):3297–3305, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002437.

Barreras:2014:EST

[BP14] A. Barreras and J. M. Peña. On the extension of some total positivity inequalities. Linear Algebra and its Applications, 448(??):153–167, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000421.

Baeth:2011:NTM

[BPA^{+}11] Nicholas Baeth, Vadim Ponomarenko, Donald Adams, Rene Ardila, David Hannasch, Audra Kosh, Hanah McCarthy, and Ryan Rosenbaum. Number theory of matrix semigroups. Linear Algebra and its Applications, 434(3):694–711, February 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bravo:2014:DPI

[BPDC14] Cedric M. A. Ayala Bravo, Renato Pavanello, Philippe R. B. Devloo, and Jorge L. D. Calle. Definition of a P - interpolating space of hierarchical bases of finite elements on the pyramid. Linear Algebra and its Applications, 460(??):174–204, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004832.

Brualdi:2011:CBD

[BPRY11] Richard A. Brualdi, Robert J. Plemmons, Lothar Reichel, and Qiang Ye. Contributions of Biswa N. Datta to Linear Algebra and its Applications. Linear Algebra and its Applications, 434(7):1613–1614, April 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2010:AKI

[BR10] Jean-Christophe Bourin and Éric Ricard. An asymmetric Kadison’s inequality. Linear Algebra and its Applications, 433(3):499–510, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boasso:2011:CEN

[BR11] Enrico Boasso and Vladimir Rakocevic. Characterizations of EP and normal Banach algebra elements and Banach space operators. Linear Algebra and its Applications, 435 (2):342–353, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Branquinho:2012:SRO

[BR12] A. Branquinho and M. N. Rebocho. Structure relations for orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 436(11):4296–4310, June 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001139.

Bekjan:2014:HTI

[BR14a] Turdebek N. Bekjan and Madi Raikhan. An Hadamardtype inequality. Linear Algebra and its Applications, 443(??):228–234, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007179.

Bella:2014:SCM

[BR14b] Tom Bella and Jenna Reis. The spectral connection matrix for classical orthogonal polynomials of a single parameter. Linear Algebra and its Applications, 458(??):161–182, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003668.

Bierkens:2014:SMP

[BR14c] Joris Bierkens and André Ran. A singular M-matrix perturbed by a nonnegative rank one matrix has positive principal minors; is it D-stable? Linear Algebra and its Applications, 457 (??):191–208, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003139.

Bracic:2014:EHC

[BR14d] Janko Bracic and Tina Rudolf. Estimates of khyperreflexivity constants. Linear Algebra and its Applications, 458(??):47–59, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514003371.

Braman:2010:TOT

[Bra10] Karen Braman. Third-order tensors as linear operators on a space of matrices. Linear Algebra and its Applications, 433(7):1241–1253, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bekker:2011:WRD

[BRA11] A. Bekker, J. J. J. Roux, and M. Arashi. Wishart ratios with dependent structure: New members of the bimatrix beta type IV. Linear Algebra and its Applications, 435(12):3243–3260, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Branden:2012:STP

[Brä12] Petter Brändén. Solutions to two problems on permanents. Linear Algebra and its Applications, 436 (1):53–58, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100471X.

Bremner:2014:SRM

[Bre14] Murray R. Bremner. Structure of the rational monoid algebra for Boolean matrices of order 3. Linear Algebra and its Applications, 449 (??):381–401, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001025.

Brijder:2013:NTP

[Bri13] Robert Brijder. The nullity theorem for principal pivot transform. Linear Algebra and its Applications, 439(11): 3638–3642, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005697.

Brown:2012:BRU

[BRLS12] David E. Brown, Scott Roy, J. Richard Lundgren, and Daluss J. Siewert. Boolean rank of upset tournament matrices. Linear Algebra and its Applications, 436(9): 3239–3246, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007397.

Brualdi:2010:SD

[Bru10] Richard A. Brualdi. Spectra of digraphs. Linear Algebra and its Applications, 432 (9):2181–2213, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2011:EC

[Bru11] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 434(4):849–853, February 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2013:EC

[Bru13] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 438(1):1–6, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006544.

Brualdi:2014:EC

[Bru14] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 440(??):360–362, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006290.

Boley:2011:CTD

[BRZ11] Daniel Boley, Gyan Ranjan, and Zhi-Li Zhang. Commute times for a directed graph using an asymmetric Laplacian. Linear Algebra and its Applications, 435(2):224–242, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bracic:2013:HCS

[BRZ13] Janko Bracic, Viktória Rozborová, and Michal Zajac. Hyperreflexivity constants of some spaces of matrices. Linear Algebra and its Applications, 439(5):1340–1349, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002796.

Bottcher:2010:GGB

[BS10] A. Böttcher and I. M. Spitkovsky. A gentle guide to the basics of two projections theory. Linear Algebra and its Applications, 432(6):1412– 1459, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2011:IML

[BS11a] R. B. Bapat and Sivaramakrishnan Sivasubramanian. Identities for minors of the Laplacian, resistance and distance matrices. Linear Algebra and its Applications, 435(6):1479–1489, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendaoud:2011:ALS

[BS11b] M. Bendaoud and M. Sarih. Additive local spectrum compressors. Linear Algebra and its Applications, 435(6):1473– 1478, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2011:CFD

[BS11c] A. Böttcher and I. M. Spitkovsky. On certain finitedimensional algebras generated by two idempotents. Linear Algebra and its Applications, 435(8):1823–1836, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brodsky:2011:TQT

[BS11d] Sarah B. Brodsky and Bernd Sturmfels. Tropical quadrics through three points. Linear Algebra and its Applications, 435(7):1778–1785, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2012:NRT

[BS12a] Natalia Bebiano and Ilya M. Spitkovsky. Numerical ranges of Toeplitz operators with matrix symbols. Linear Algebra and its Applications, 436(6):1721–1726, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100468X.

Benkovic:2012:JDU

[BS12b] Dominik Benkovic and Nejc Sirovnik. Jordan derivations of unital algebras with idempotents. Linear Algebra and its Applications, 437 (9):2271–2284, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004442.

Berman:2012:CPH

[BS12c] Abraham Berman and Dafna Shasha. Completely positive house matrices. Linear Algebra and its Applications, 436(1):12–26, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004563.

Bhatia:2012:SIP

[BS12d] Rajendra Bhatia and Rajesh Sharma. Some inequalities for positive linear maps. Linear Algebra and its Applications, 436(6):1562–1571, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951000501X.

Bresar:2012:LAA

[BS12e] Matej Bresar and Spela Spenko. On Lie and associative algebras containing inner derivations. Linear Algebra and its Applications, 437 (2):648–658, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002108.

Bottcher:2013:CFD

[BS13a] A. Böttcher and I. M. Spitkovsky. Classification of the finite-dimensional algebras generated by two tightly coupled idempotents. Linear Algebra and its Applications, 439(3):538–551, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005617.

Butler:2013:SBS

[BS13b] Brian K. Butler and Paul H. Siegel. Sharp bounds on the spectral radius of nonnegative matrices and digraphs. Linear Algebra and its Applications, 439(5):1468–1478, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003054.

Balcerzak:2014:SSC

[BS14] Marek Balcerzak and Filip Strobin. Spaceability of the set of continuous injections from B_{ℓ_{p}} into ℓ_{p} with nowhere continuous inverses. Linear Algebra and its Applications, 450(??):76–82, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001098.

Beasley:2012:PTR

[BSK12] LeRoy B. Beasley, Seok-Zun Song, and Kyung-Tae Kang. Preservers of term ranks of symmetric matrices. Linear Algebra and its Applications, 436(6):1727–1738, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004678.

Beasley:2013:CTR

[BSKL13] LeRoy B. Beasley, Seok-Zun Song, Kyung-Tae Kang, and Sang-Gu Lee. A comparison of term ranks of symmetric matrices and their preservers. Linear Algebra and its Applications, 438(10): 3745–3754, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100262X.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Bostan:2010:FCA

[BSS10] Alin Bostan, Bruno Salvy, and Éric Schost. Fast conversion algorithms for orthogonal polynomials. Linear Algebra and its Applications, 432 (1):249–258, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2013:TCN

[BSST13] Peter Butkovic, Hans Schneider, Sergei Sergeev, and Bit-Shun Tam. Two cores of a nonnegative matrix. Linear Algebra and its Applications, 439(7):1929–1954, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003716.

Bomze:2014:SEC

[BSU14] Immanuel M. Bomze, Werner Schachinger, and Reinhard Ullrich. From seven to eleven: Completely positive matrices with high cp-rank. Linear Algebra and its Applications, 459(??):208–221, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003917.

Baksalary:2010:SMB

[BT10] Oskar Maria Baksalary and Götz Trenkler. On a subspace metric based on matrix rank. Linear Algebra and its Applications, 432(6):1475– 1491, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2011:DRM

[BT11a] Oskar Maria Baksalary and Götz Trenkler. On disjoint range matrices. Linear Algebra and its Applications, 435 (6):1222–1240, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barkovsky:2011:HRF

[BT11b] Yury Barkovsky and Mikhail Tyaglov. Hurwitz rational functions. Linear Algebra and its Applications, 435(8):1845– 1856, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bayat:2011:AGM

[BT11c] M. Bayat and H. Teimoori. Arithmetic-Geometric Mean determinantal identity. Linear Algebra and its Applications, 435(11):2936–2941, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2013:DIS

[BT13] Natália Bebiano and Mikhail Tyaglov. Direct and inverse spectral problems for a class of non-self-adjoint periodic tridiagonal matrices. Linear Algebra and its Applications, 439 (11):3490–3504, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300582X.

Bah:2014:BRI

[BT14a] Bubacarr Bah and Jared Tanner. Bounds of restricted isometry constants in extreme asymptotics: Formulae for Gaussian matrices. Linear Algebra and its Applications, 441(??):88–109, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008312.

Baksalary:2014:SDD

[BT14b] Oskar Maria Baksalary and Götz Trenkler. On subspace distances determined by the Frobenius norm. Linear Algebra and its Applications, 448(??):245–263, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000354.

Bolten:2012:PSA

[BTYZ12] Matthias Bolten, Alexander Thiess, Irad Yavneh, and Rudolf Zeller. Preconditioning systems arising from the KKR Green function method using block-circulant matrices. Linear Algebra and its Applications, 436(2):

436–446, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004241.

Buckley:2010:ETP

[Buc10]  Anita Buckley.  Elemen-

tary transformations of Pfaffian representations of plane curves. Linear Algebra and its Applications, 433(4):758– 780, October 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Budnitska:2011:TCA

[Bud11] Tetiana Budnitska. Topological classification of affine operators on unitary and Euclidean spaces. Linear Algebra and its Applications, 434(2): 582–592, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bujanovic:2013:PAR

[Buj13]  Zvonimir Bujanović.  On

the permissible arrangements of Ritz values for normal matrices in the complex plane. Linear Algebra and its Applications, 438(12): 4606–4624, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001328.

Bunger:2014:IDE

[Bün14] F. Bünger. Inverses, determinants, eigenvalues, and eigenvectors of real symmetric Toeplitz matrices with linearly increasing entries. Linear Algebra and its Applications, 459(??):595–619, October 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400473X.

Barnsley:2011:EPL

[BV11] Michael Barnsley and Andrew Vince. The eigenvalue problem for linear and affine iterated function systems. Linear Algebra and its Applications, 435(12):3124–3138, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Badawy:2012:PTA

[BV12a] M. Badawy and E. Van Vleck. Perturbation theory for the approximation of stability spectra by QR methods for sequences of linear operators on a Hilbert space. Linear Algebra and its Applications, 437(1):37–59, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001176.

Barreira:2012:CBG

[BV12b] Luis Barreira and Claudia Valls. Conjugacies between general contractions. Linear Algebra and its Applications, 436(9):3087–3098, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007208.

Benner:2013:SCE

[BV13a] Peter Benner and Matthias Voigt. Spectral characterization and enforcement of negative imaginariness for descriptor systems. Linear Algebra and its Applications, 439(4):1104–1129, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000451.

Bottazzi:2013:BAD

[BV13b] Tamara Bottazzi and Alejandro Varela. Best approximation by diagonal compact operators. Linear Algebra and its Applications, 439 (10):3044–3056, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005399.

Bergsma:2012:PNP

[BVV12] Hannah Bergsma, Kevin N. Vander Meulen, and Adam Van Tuyl. Potentially nilpotent patterns and the Nilpotent-Jacobian method. Linear Algebra and its Applications, 436(12):4433–4445, June 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004228.

Bakonyi:2011:MCM

[BW11] Mihaly Bakonyi and Hugo J. Woerdeman. Marix completions, moments, and sums of Hermitian squares. Princeton series in applied mathematics. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-12889-8 (hardcover). ???? pp. LCCN ????

Bunimovich:2012:IGR

[BW12] L. A. Bunimovich and B. Z. Webb. Isospectral graph reductions and improved estimates of matrices’ spectra. Linear Algebra and its Applications, 437(7):1429– 1457, October 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003072.

Broome:2013:CHS

[BW13a] Helen Broome and Shayne Waldron. On the construction of highly symmetric tight frames and complex polytopes. Linear Algebra and its Applications, 439(12):4135– 4151, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006228.

Buzinski:2013:MPF

[BW13b] David Buzinski and Robin Winstanley. On multilinear polynomials in four variables evaluated on matrices. Linear Algebra and its Applications, 439(9):2712–2719, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005028.

Bikchentaev:2011:RTR

[BY11] Airat M. Bikchentaev and Rinat S. Yakushev. Representation of tripotents and representations via tripotents. Linear Algebra and its Applications, 435(9):2156–2165, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bydder:2010:SCL

[Byd10] Mark Bydder. Solution of a complex least squares problem with constrained phase. Linear Algebra and its Applications, 433(11–12):1719– 1721, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bu:2014:RDS

[BYZZ14] Changjiang Bu, Bo Yan, Xiuqing Zhou, and Jiang Zhou. Resistance distance in subdivision-vertex join and subdivision-edge join of graphs. Linear Algebra and its Applications, 458(??): 454–462, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003838.

Bu:2012:SLS

[BZ12a] Changjiang Bu and Jiang Zhou. Signless Laplacian spectral characterization of the cones over some regular graphs. Linear Algebra and its Applications, 436 (9):3634–3641, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000055.

Bu:2012:STW

[BZ12b] Changjiang Bu and Jiang Zhou. Starlike trees whose maximum degree exceed 4 are determined by their Qspectra. Linear Algebra and its Applications, 436(1): 143–151, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004770.

Bu:2013:NFD

[BZ13] Changjiang Bu and Chundi Zhang. A note on the formulas for the Drazin inverse of the sum of two matrices. Linear Algebra and its Applications, 439(3):565–576, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000888.

Bu:2014:CHS

[BZW14] Changjiang Bu, Jiang Zhou, and Yimin Wei. E-cospectral hypergraphs and some hypergraphs determined by their spectra. Linear Algebra and its Applications, 459(??): 397–403, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004704.

Bai:2013:ILA

[BZWL13] RuiPu Bai, Lihong Zhang, Yong Wu, and Zhenheng Li. On 3-Lie algebras with abelian ideals and subalgebras. Linear Algebra and its Applications, 438(5):2072– 2082, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007847.

Bu:2014:NMG

[BZZ14a] Changjiang Bu, Xu Zhang, and Jiang Zhou. A note on the multiplicities of graph eigenvalues. Linear Algebra and its Applications, 442 (??):69–74, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005004.

Bu:2014:IRP

[BZZ^{+}14b] Changjiang Bu, Xu Zhang, Jiang Zhou, Wenzhe Wang, and Yimin Wei. The inverse, rank and product of tensors. Linear Algebra and its Applications, 446 (??):269–280, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008239.

Chen:2010:NTA

[CA10] Jinhai Chen and Ravi P. Agarwal. On Newton-type approach for piecewise linear systems. Linear Algebra and its Applications, 433(7):1463– 1471, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Caldeira:2012:PSS

[Cal12]  Cristina Caldeira.  Pairs

of sets with small sumset and small periodic productset. Linear Algebra and its Applications, 436(6):1689– 1700, March 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004708.

Campos:2013:ARC

[Cam13] Jamilson R. Campos. An abstract result on Cohen strongly summing operators. Linear Algebra and its Applications, 439(12):4047–4055, December 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006186.

Cardinal:2010:SMR

[Car10a] Jean-Paul Cardinal. Symmetric matrices related to the Mertens function. Linear Algebra and its Applications, 432(1):161–172, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carrieu:2010:CCS

[Car10b] Hervé Carrieu. Close to the Craig–Sakamoto’s theorem. Linear Algebra and its Applications, 432(2–3):777– 779, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cardanobile:2011:SMP

[Car11] Stefano Cardanobile. The L^{2}- strong maximum principle on arbitrary countable networks. Linear Algebra and its Applications, 435(6):1315–1325, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carriegos:2013:ECL

[Car13] Miguel V. Carriegos. Enumeration of classes of linear systems via equations and via partitions in an ordered abelian monoid. Linear Algebra and its Applications, 438 (3):1132–1148, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006714.

Castella:2010:EAL

[Cas10] Dominique Castella. Eléments d’algébre linéaire tropicale. (French) [Elements of tropical linear algebra]. Linear Algebra and its Applications, 432 (6):1460–1474, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Caswell:2013:SAD

[Cas13] Hal Caswell. Sensitivity analysis of discrete Markov chains via matrix calculus. Linear Algebra and its Applications, 438(4):1727–1745, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006185.

Causin:2011:DSR

[Cau11] Andrea Causin. On the dimension of some real spaces of bounded rank matrices. Linear Algebra and its Applications, 434(2):501–506, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cason:2013:IML

[CAV13] T. P. Cason, P. A. Absil, and P. Van Dooren. Iterative methods for low rank approximation of graph similarity matrices. Linear Algebra and its Applications, 438 (4):1863–1882, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007841.

Chebotarev:2013:SEL

[CBB13] Pavel Chebotarev, R. B. Bapat, and R. Balaji. Simple expressions for the long walk distance. Linear Algebra and its Applications, 439(4):893–898, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005927.

Caiafa:2010:GCR

[CC10] Cesar F. Caiafa and Andrzej Cichocki. Generalizing the column–row matrix decomposition to multi-way arrays. Linear Algebra and its Applications, 433(3):557– 573, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2012:MMS

[CC12] Chi-Tung Chang and Chang-Pao Chen. Matrix maps of statistically convergent sequences. Linear Algebra and its Applications, 437 (12):2896–2909, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005678.

Carriegos:2014:TFA

[CC14] Miguel V. Carriegos and Ángel Luis Muñoz Castañeda. On the K-theory of feedback actions on linear systems. Linear Algebra and its Applications, 440(??): 233–242, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006836.

Cavers:2012:SAM

[CCF^{+}12] M. Cavers, S. M. Cioaba, S. Fallat, D. A. Gregory, W. H. Haemers, S. J. Kirkland, J. J. McDonald, and M. Tsatsomeros. Skew-adjacency matrices of graphs. Linear Algebra and its Applications, 436(12): 4512–4529, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000833.

Camacho:2014:FLA

[CCGO14] L. M. Camacho, E. M. Cañete, J. R. Gómez, and B. A. Omirov. p-Filiform Leibniz algebras of maximum length. Linear Algebra and its Applications, 450 (??):316–333, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001347.

Camacho:2013:LAN

[CCGR13] L. M. Camacho, E. M. Cañete, J. R. Gómez, and Sh. B. Redjepov. Leibniz algebras of nilindex n − 3 with characteristic sequence (n − 3, 2, 1). Linear Algebra and its Applications, 438 (4):1832–1851, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004496.

Camacho:2013:FZA

[CCGVO13] L. M. Camacho, E. M. Cañete, S. Gómez-Vidal, and B. A. Omirov. p-Filiform Zinbiel algebras. Linear Algebra and its Applications, 438(7):2958–2972, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008373.

Chen:2014:RSA

[CCL14] Kejun Chen, Guangzhou Chen, and Wen Li. Regular sparse anti-magic squares with the second maximum density. Linear Algebra and its Applications, 457(??):12– 28, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002882.

Cardinali:2010:RPD

[CD10] Ilaria Cardinali and Bart De Bruyn. Regular partitions of dual polar spaces. Linear Algebra and its Applications, 432(2–3):744–769, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cigler:2011:LGS

[CD11] Grega Cigler and Roman Drnovšek. From local to global similarity of matrix groups. Linear Algebra and its Applications, 435(6):1285– 1295, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:2012:LPI

[CD12a] M. Purificação Coelho and M. Antónia Duffner. Linear preservers of immanants on skew-symmetric matrices. Linear Algebra and its Applications, 436(7):2536– 2553, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007889.

Cooper:2012:SUH

[CD12b] Joshua Cooper and Aaron Dutle. Spectra of uniform hypergraphs. Linear Algebra and its Applications, 436 (9):3268–3292, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007610.

Curgus:2012:RKP

[CD12c] B. ’Curgus and A. Dijksma. On the reproducing kernel of a Pontryagin space of vector valued polynomials. Linear Algebra and its Applications, 436(5):1312–1343, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005787.

Cigler:2013:SMN

[CD13] Grega Cigler and Roman Drnovsek. On semigroups of matrices with nonnegative diagonals. Linear Algebra and its Applications, 438(1):626–633, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006593.

Can:2014:EBS

[CD14] Mümün Can and M. Cihat Dagli. Extended Bernoulli and Stirling matrices and related combinatorial identities. Linear Algebra and its Applications, 444(??):114– 131, March 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007696.

Cho:2010:EOH

[CDDY10] M. Chō, S. V. Djordjević, B. P. Duggal, and T. Yamazaki. On an elementary operator with w-hyponormal operator entries. Linear Algebra and its Applications, 433 (11–12):2070–2079, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cicalo:2012:SDN

[CdGS12] Serena Cicalò, Willem A. de Graaf, and Csaba Schneider. Six-dimensional nilpotent Lie algebras. Linear Algebra and its Applications, 436(1):163–189, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004861.

See [CdGS20].

Cicalo:2020:CSD

[CdGS20] Serena Cicalò, Willem A. de Graaf, and Csaba Schneider. Corrigendum to “Sixdimensional nilpotent Lie algebras” [Linear Algebra Appl. 436 (2012) 163–189]. Linear Algebra and its Applications, 604(??):507–508, November 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379520303189.

See [CdGS12].

Conflitti:2012:MLC

[CDM12] Alessandro Conflitti, C. M. Da Fonseca, and Ricardo Mamede. The maximal length of a chain in the Bruhat order for a class of binary matrices. Linear Algebra and its Applications, 436(3): 753–757, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100588X.

Carnicer:2010:RMT

[CDP10] J. M. Carnicer, J. Delgado, and J. M. Peña. Richardson method and totally nonnegative linear systems. Linear Algebra and its Applications, 433(11–12):2010–2017, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cechlarova:2010:MML

[Cec10] Katarína Cechlárová. On max–min linear inequalities and Coalitional Resource Games with sharable resources. Linear Algebra and its Applications, 433(1):127– 135, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carmona:2014:DEO

[CEM14] A. Carmona, A. M. Encinas, and M. Mitjana. Discrete elliptic operators and their Green operators. Linear Algebra and its Applications, 442(??):115–134, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004680.

Centrone:2011:GIG

[Cen11] Lucio Centrone. Z_{2}-graded identities of the Grassmann algebra in positive characteristic. Linear Algebra and its Applications, 435(12):3297– 3313, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cerzo:2010:STI

[Cer10] Diana R. Cerzo. Structure of thin irreducible modules of a Q-polynomial distanceregular graph. Linear Algebra and its Applications, 433(8– 10):1573–1613, December 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2014:VDP

[CEY14] Huilan Chang, Sen-Peng Eu, and Pei-Lan Yen. Variances and determinantal profiles of orientations. Linear Algebra and its Applications, 457 (??):209–223, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003188.

Cameron:2014:UST

[CFG^{+}14] Stephen Cameron, Shannon Fehrenbach, Leah Granger, Oliver Hennigh, Sunrose Shrestha, and Christino Tamon. Universal state transfer on graphs. Linear Algebra and its Applications, 455(??):115–142, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002845.

Chang:2014:LGN

[CFHL14] Gerard Jennhwa Chang, Keqin Feng, Liang-Hao Huang, and Mei Lu. The linear guessing number of undirected graphs. Linear Algebra and its Applications, 449 (??):119–131, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000779.

Charles:2013:RBD

[CFJKS13] Zachary B. Charles, Miriam Farber, Charles R. Johnson, and Lee Kennedy-Shaffer. The relation between the diagonal entries and the eigenvalues of a symmetric matrix, based upon the sign pattern of its off-diagonal entries. Linear Algebra and its Applications, 438(3):1427–1445, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006799.

Cavers:2010:NLE

[CFK10a] Michael Cavers, Shaun Fallat, and Steve Kirkland. On the normalized Laplacian energy and general Randić index R_{−1} of graphs. Linear Algebra and its Applications, 433(1): 172–190, July 15, 2010. CO-

DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2010:LEB

[CFK^{+}10b] Yi-Fan Chen, Hung-Lin Fu, In-Jae Kim, Eryn Stehr, and Brendon Watts. On the largest eigenvalues of bipartite graphs which are nearly complete. Linear Algebra and its Applications, 432(2– 3):606–614, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2013:SNI

[CFL13a] Che-Man Cheng, Kin-Sio Fong, and Weng-Fai Lei. On some norm inequalities involving the commutator and XY − Y X^{T}. Linear Algebra and its Applications, 438(6):2793–2807, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007860.

Chodrow:2013:ULB

[CFL13b] Philip Chodrow, Cole Franks, and Brian Lins. Upper and lower bounds for the iterates of order-preserving homogeneous maps on cones. Linear Algebra and its Applications, 439(4):999–1005, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007379.

Craigen:2013:CPH

[CFLW13] R. Craigen, G. Faucher, R. Low, and T. Wares. Circulant partial Hadamard matrices. Linear Algebra and its Applications, 439(11):3307– 3317, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005533.

Cui:2012:PPP

[CFLY12] Jianlian Cui, Virginia Forstall, Chi-Kwong Li, and Vincent Yannello. Properties and preservers of the pseudospectrum. Linear Algebra and its Applications, 436(2): 316–325, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511002680.

Carli:2013:EAM

[CFPP13] F. P. Carli, A. Ferrante, M. Pavon, and G. Picci. An efficient algorithm for maximum entropy extension of block-circulant covariance matrices. Linear Algebra and its Applications, 439(8): 2309–2329, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Cvetkovic:2012:P

[CFPT12] Ljiljana Cvetković, Andreas Frommer, Juan Manuel Peña, and Michael Tsatsomeros. Preface. Linear Algebra and its Applications, 436(2): 263–264, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004393.

Cheung:2011:PST

[CG11] Wang-Chi Cheung and Chris Godsil. Perfect state transfer in cubelike graphs. Linear Algebra and its Applications, 435(10):2468–2474, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2013:NNR

[CG13] Che-Man Cheng and Yuan Gao. A note on numerical range and product of matrices. Linear Algebra and its Applications, 438(7): 3139–3143, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000189.

See corrigendum [Che14b].

Chan:2013:MEI

[CGGS13] Alice Zhuo-Yu Chan, Luis Alberto Garcia German, Stephan Ramon Garcia, and Amy L. Shoemaker. On the matrix equation XA + AX^{T} = 0, II: Type 0-1 interactions. Linear Algebra and its Applications, 439 (12):3934–3944, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006034.

See also Part I [GS13a].

Couselo:2010:SCL

[CGM^{+}10] Elena Couselo, Santos González, Victor Markov, Consuelo Martínez, and Alexander Nechaev. Some constructions of linearly optimal group codes. Linear Algebra and its Applications, 433(2):356– 364, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Castillo:2011:PST

[CGM11] K. Castillo, L. Garza, and F. Marcellán. Perturbations on the subdiagonals of Toeplitz matrices. Linear Algebra and its Applications, 434(6):1563–1579, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Corach:2012:USI

[CGM12] Gustavo Corach, M. Celeste Gonzalez, and Alejandra Maestripieri. Unbounded symmetrizable idempotents.

Linear Algebra and its Applications, 437(2):659–674, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200211X.

Chen:2014:UGH

[CGMJ14] Zhao Chen, Matthew Grimm, Paul McMichael, and Charles R. Johnson. Undirected graphs of Hermitian matrices that admit only two distinct eigenvalues. Linear Algebra and its Applications, 458(??): 403–428, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000895.

Ciesielski:2014:LSA

[CGMPSS14] Krzysztof Chris Ciesielski, José L. Gámez-Merino, Daniel Pellegrino, and Juan B. Seoane-Sepúlveda. Lineability, spaceability, and additivity cardinals for Darbouxlike functions. Linear Algebra and its Applications, 440(??):307–317, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300668X.

Castro-Gonzalez:2010:DIP

[CGMS10a] N. Castro-González and M. F. Martínez-Serrano. Drazin inverse of partitioned matrices in terms of Banachiewicz– Schur forms. Linear Algebra and its Applications, 432(7): 1691–1702, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Castro-Gonzalez:2010:EDI

[CGMS10b] N. Castro-González and M. F. Martínez-Serrano. Expressions for the g-Drazin inverse of additive perturbed elements in a Banach algebra. Linear Algebra and its Applications, 432(8):1885–1895, April 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Camacho:2010:NGF

[CGO10] L. M. Camacho, J. R. Gómez, and B. A. Omirov. Naturally graded (n − 3)-filiform Leibniz algebras. Linear Algebra and its Applications, 433(2): 433–446, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cruz:2014:UBE

[CGR14] Roberto Cruz, Hernán Giraldo, and Juan Rada. An upper bound for the energy of radial digraphs. Linear Algebra and its Applications, 442(??):75–81, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004631.

Castro-Gonzalez:2013:GIM

[CGRVC13] N. Castro-González, J. Robles, and J. Y. Vélez-Cerrada. The group inverse of 2 × 2 matrices over a ring. Linear Algebra and its Applications, 438(9):3600–3609, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000384.

Cicone:2010:FPP

[CGSCZ10] Antonio Cicone, Nicola Guglielmi, Stefano Serra-Capizzano, and Marino Zennaro. Finiteness property of pairs of 2×2 signmatrices via real extremal polytope norms. Linear Algebra and its Applications, 432 (2–3):796–816, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carmona:2014:DSU

[CGTR14] Juan Carmona, Ivan Gutman, Nelda Jaque Tamblay, and María Robbiano. A decreasing sequence of upper bounds for the Laplacian energy of a tree. Linear Algebra and its Applications, 446(??):304–313, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000317.

Chen:2014:ECC

[CGW14] Bai Fan Chen, Ebrahim Ghorbani, and Kok Bin Wong. On the eigenvalues of certain Cayley graphs and arrangement graphs. Linear Algebra and its Applications, 444(??):246–253, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007738.

Chang:2013:NRG

[CGWW13] Chi-Tung Chang, Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Numerical ranges and Gersgorin discs. Linear Algebra and its Applications, 438(3):1170– 1192, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006660.

Chen:2011:MEI

[CH11] Gong-Ning Chen and Yong-Jian Hu. Minimum wentropy interpolants for matricial Carathéodory functions and maximum determinant completions of associated block Pick matrix. Linear Algebra and its Applications, 434(8):1851–1878, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2012:CBS

[CH12] Xiang-Ke Chang and Xing-Biao Hu. A conjecture based on Somos-4 sequence and its extension. Linear Algebra and its Applications, 436(11):4285–4295, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000808.

Carden:2013:RVN

[CH13] Russell Carden and Derek J. Hansen. Ritz values of normal matrices and Ceva’s theorem. Linear Algebra and its Applications, 438(11): 4114–4129, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000177.

Chen:2014:EEM

[CH14] Xiaodan Chen and Yaoping Hou. The extreme eigenvalues and maximum degree of k-connected irregular graphs. Linear Algebra and its Applications, 463(??): 33–44, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005643.

Chandola:2010:LBE

[Cha10] Himanshu Chandola. A lower bound on the error in dimensionality reduction resulting from projection onto a restricted subspace. Linear Algebra and its Applications, 433(11–12):2147–2151, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chahlaoui:2012:PEB

[Cha12] Y. Chahlaoui. A posteriori error bounds for discrete balanced truncation. Linear Algebra and its Applications, 436(8):2744–2763, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005374.

Chamberland:2013:FMC

[Cha13a] Marc Chamberland. Factored matrices can generate combinatorial identities. Linear Algebra and its Applications, 438(4):1667–1677, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006136.

Chang:2013:MPQ

[Cha13b] Mei-Chu Chang. On a matrix product question in cryptography. Linear Algebra and its Applications, 439(7): 1742–1748, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Chai:2014:TTC

[Cha14] Kian Ming A. Chai. Threeby-three correlation matrices: its exact shape and a family of distributions. Linear Algebra and its Applications, 458(??):589–604, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400408X.

Chen:2013:IOR

[Che13]  Li Chen.  On intertwin-

ing operators via reproducing kernels. Linear Algebra and its Applications, 438 (9):3661–3666, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000505.

Chen:2014:LMC

[Che14a]  Hung-Yuan Chen.  Lin-

ear maps characterized by the action on square-zero elements. Linear Algebra and its Applications, 450 (??):243–249, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001311.

Cheng:2014:CSN

[Che14b] Che-Man Cheng. Corrigendum to “A note on numerical range and product of matrices” [Linear Algebra Appl. 438 (7) (2013) 3139–3143]. Linear Algebra and its Applications, 459(??): 622–624, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004054.

See [CG13].

Chiumiento:2013:NOL

[Chi13] Eduardo Chiumiento. On normal operator logarithms. Linear Algebra and its Applications, 439(2):455–462, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002218.

Can:2013:UIM

[CHJ13] Mahir Bilen Can, Roger Howe, and Michael Joyce. Unipotent invariant matrices. Linear Algebra and its Applications, 439(1):196–210, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001535.

Casazza:2013:NSC

[CHK^{+}13] Peter G. Casazza, Andreas Heinecke, Keri Kornelson, Yang Wang, and Zhengfang Zhou. Necessary and sufficient conditions to perform Spectral Tetris. Linear Algebra and its Applications, 438(5):2239–2255, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007604.

Choi:2012:EIM

[CHLS12] Man-Duen Choi, Zejun Huang, Chi-Kwong Li, and Nung-Sing Sze. Every invertible matrix is diagonally equivalent to a matrix with distinct eigenvalues. Linear Algebra and its Applications, 436 (9):3773–3776, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008263.

Cadney:2014:IRM

[CHLW14] Josh Cadney, Marcus Huber, Noah Linden, and Andreas Winter. Inequalities for the ranks of multipartite quantum states. Linear Algebra and its Applications, 452(??):153–171, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001797.

Choi:2013:PCC

[Cho13] Daeshik Choi. A proof of Crouzeix’s conjecture for a class of matrices. Linear Algebra and its Applications, 438(8):3247–3257, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000475.

Chang:2011:CGR

[CHY11] Gerard J. Chang, Liang-Hao Huang, and Hong-Gwa Yeh. A characterization of graphs with rank 4. Linear Algebra and its Applications, 434 (8):1793–1798, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2012:CGR

[CHY12] Gerard J. Chang, Liang-Hao Huang, and Hong-Gwa Yeh. A characterization of graphs with rank 5. Linear Algebra and its Applications, 436(11):4241–4250, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000857.

Chang:2013:DME

[CHZ13] Xiang-Ke Chang, Xing-Biao Hu, and Ying-Nan Zhang. A direct method for evaluating some nice Hankel determinants and proofs of several conjectures. Linear Algebra and its Applications, 438(5):2523–2541, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200780X.

Chugunov:2010:CSN

[CI10] V. N. Chugunov and Kh. D. Ikramov. A complete solution of the normal Hankel problem. Linear Algebra and its Applications, 432(12):3210–3230, July 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic-Ilic:2013:SRI

[CI13] Dragana S. Cvetković-Ilić. Some results on the (2, 2, 0) Drazin inverse problem. Linear Algebra and its Applications, 438(12):4726–4741, June 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001043.

Cvetkovic-Ilic:2014:PRC

[CI14] Dragana S. Cvetković-Ilić. The point, residual and continuous spectrum of an upper triangular operator matrix. Linear Algebra and its Applications, 459(??): 357–367, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004595.

Cvetkovic-Ilic:2011:ROL

[CIH11] Dragana S. Cvetković-Ilić and Robin Harte. Reverse order laws in C^{∗}-algebras. Linear Algebra and its Applications, 434(5):1388–1394, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Casas:2012:ACD

[CILL12] J. M. Casas, M. A. Insua, M. Ladra, and S. Ladra. An algorithm for the classification of 3-dimensional complex Leibniz algebras. Linear Algebra and its Applications, 436(9):3747–3756, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008251.

Cimpric:2011:SPM

[Cim11] J. Cimprič. Strict Positivstellensätze for matrix polynomials with scalar constraints. Linear Algebra and its Applications, 434(8):1879–1883, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cioab:2010:EEE

[Cio10a] Sebastian M. Cioab. Erratum to “Eigenvalues and edge-connectivity of regular graphs”. Linear Algebra and its Applications, 432(9):2455, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [Cio10b].

Cioaba:2010:EEC

[Cio10b] Sebastian M. Cioabă. Eigenvalues and edge-connectivity of regular graphs. Linear Algebra and its Applications, 432(1):458–470, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See erratum [Cio10a].

Cirrito:2013:GAU

[Cir13] Alessio Cirrito. Gradings on the algebra of upper triangular matrices of size three. Linear Algebra and its Applications, 438(11): 4520–4538, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000979.

Cirici:2014:CIS

[Cir14] Joana Cirici. Classification of isometries of spaces of constant curvature and invariant subspaces. Linear Algebra and its Applications, 450 (??):250–279, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001396.

Chen:2010:GSP

[CJ10] Jian-Long Chen and Xiao-Qing Jin. The generalized superoptimal preconditioner. Linear Algebra and its Applications, 432(1):203–217, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2011:SPR

[CJ11] Gi-Sang Cheon and Sung-Tae Jin. Structural properties of Riordan matrices and extending the matrices. Linear Algebra and its Applications, 435(8):2019–2032, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2012:MVC

[CJ12] Hao Chen and Jürgen Jost. Minimum vertex covers and the spectrum of the normalized Laplacian on trees. Linear Algebra and its Applications, 437(4):1089–1101, August 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002698.

Cigler:2014:SEC

[CJ14] Grega Cigler and Marjan Jerman. On separation of eigenvalues by certain matrix subgroups. Linear Algebra and its Applications, 440(??):213–217, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Corey:2013:PFV

[CJK^{+}13] Daniel Corey, Charles R. Johnson, Ryan Kirk, Brian Lins, and Ilya Spitkovsky. The product field of values. Linear Algebra and its Applications, 438(5):2155–2173, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007422.

Cheon:2013:ARG

[CJL13] Gi-Sang Cheon, Ji-Hwan Jung, and Yongdo Lim. A q-analogue of the Riordan group. Linear Algebra and its Applications, 439(12):4119– 4129, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006459.

Chan:2011:MMT

[CJR11] Melody Chan, Anders Jensen, and Elena Rubei. The 4 × 4 minors of a 5 × n matrix are a tropical basis. Linear Algebra and its Applications, 435(7):1598–1611, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Canete:2013:CDL

[CK13a] Elisa M. Cañete and Abror Kh. Khudoyberdiyev. The classification of 4-dimensional Leibniz algebras. Linear Algebra and its Applications, 439 (1):273–288, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300181X.

Casanellas:2013:GME

[CK13b] Marta Casanellas and Anna M. Kedzierska. Generating Markov evolutionary matrices for a given branch length. Linear Algebra and its Applications, 438(5):2484– 2499, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007884.

Cheon:2013:EFO

[CK13c] Gi-Sang Cheon and Hana Kim. The elements of finite order in the Riordan group over the complex field. Linear Algebra and its Applications, 439(12):4032–4046, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005855.

Cheon:2013:NAH

[CK13d] Gi-Sang Cheon and Hana Kim. A new aspect of Hankel matrices via Krylov matrix. Linear Algebra and its Applications, 438(1):

361–373, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200599X.

Choi:2014:MSB

[CK14] Jihoon Choi and Suh-Ryung Kim. On the matrix sequence {Γ(A^{m})}_{m=1}^{∞} for a Boolean matrix A whose digraph is linearly connected. Linear Algebra and its Applications, 450(??):56–75, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001256.

Coloigner:2014:LST

[CKAC14] Julie Coloigner, Ahmad Karfoul, Laurent Albera, and Pierre Comon. Line search and trust region strategies for canonical decomposition of semi-nonnegative semisymmetric 3 rd order tensors. Linear Algebra and its Applications, 450(??):334– 374, June 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000639.

Copenhaver:2013:MRS

[CKL^{+}13] Martin S. Copenhaver, Yeon Hyang Kim, Cortney Logan, Kyanne Mayfield, Sivaram K. Narayan, and Jonathan Sheperd. Maximum robustness and surgery of frames in finite dimensions. Linear Algebra and its Applications, 439(5):1330– 1339, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002784.

Casas:2013:DSL

[CKLO13] J. M. Casas, A. Kh. Khudoyberdiyev, M. Ladra, and B. A. Omirov. On the degenerations of solvable Leibniz algebras. Linear Algebra and its Applications, 439(2):472–487, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002243.

Cheon:2010:ASF

[CKS10] Gi-Sang Cheon, Hana Kim, and Louis W. Shapiro. An algebraic structure for Faber polynomials. Linear Algebra and its Applications, 433(6): 1170–1179, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chung:2011:NBI

[CKST11] Taeyoung Chung, Jack Koolen, Yoshio Sano, and Tetsuji Taniguchi. The non-bipartite integral graphs with spectral radius three. Linear Algebra and its Applications, 435 (10):2544–2559, November 15,

2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2010:SIS

[CL10a] Shexi Chen and Bolian Liu. The scrambling index of symmetric primitive matrices. Linear Algebra and its Applications, 433(6):1110–1126, November 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2010:PZS

[CL10b] Bo Cheng and Bolian Liu. Primitive zero-symmetric sign pattern matrices with the maximum base. Linear Algebra and its Applications, 433 (2):365–379, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2011:NTG

[CL11a] Bo Cheng and Bolian Liu. On the nullity of tricyclic graphs. Linear Algebra and its Applications, 434(8):1799–1810, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chuang:2011:LCL

[CL11b] Chen-Lian Chuang and Tsiu-Kwen Lee. Lengths of central linear generalized polynomials in matrix algebras. Linear Algebra and its Applications, 435(12):3206–3211, December 15, 2011. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2012:GKT

[CL12a] Xiaohong Cao and Aifang Liu. Generalized Kato type operators and property (ω) under perturbations. Linear Algebra and its Applications, 436(7):2231–2239, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007245.

Chen:2012:LSA

[CL12b] Hongjia Chen and Junbo Li. Left-symmetric algebra structures on the W - algebra W (2, 2). Linear Algebra and its Applications, 437(7):1821–1834, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200362X.

Cronin:2012:ISN

[CL12c] Anthony G. Cronin and Thomas J. Laffey. An inequality for the spectra of nonnegative matrices. Linear Algebra and its Applications, 436(9):3225–3238, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007488.

Cheng:2013:NIE

[CL13a] Bo Cheng and Bolian Liu. The normalized incidence energy of a graph. Linear Algebra and its Applications, 438(11):4510–4519, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000499.

Cheon:2013:IPS

[CL13b] Gi-Sang Cheon and Yongdo Lim. Integral polynomial sequences arising from matrix powers of order 2. Linear Algebra and its Applications, 438(1):269–287, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005940.

Chang:2014:CES

[CL14a] Gerard Jennhwa Chang and Jephian Chin-Hung Lin. Counterexamples to an edge spread question for zero forcing number. Linear Algebra and its Applications, 446 (??):192–195, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000275.

Cronin:2014:PCL

[CL14b] Anthony G. Cronin and Thomas J. Laffey. Perturbations on constructible lists preserving realizability in the NIEP and questions of Monov. Linear Algebra and its Applications, 445(??): 206–222, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300801X.

Chen:2012:SBE

[CLCL12] Xiao Shan Chen, Wen Li, Xiaojun Chen, and Jun Liu. Structured backward errors for generalized saddle point systems. Linear Algebra and its Applications, 436(9): 3109–3119, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007063.

Cui:2014:SUS

[CLHQ14] Jianlian Cui, Qiting Li, Jinchuan Hou, and Xiaofei Qi. Some unitary similarity invariant sets preservers of skew Lie products. Linear Algebra and its Applications, 457(??):76–92, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002894.

Cheng:2012:SMU

[CLL12] Bo Cheng, Bolian Liu, and Jianxi Liu. On the spectral moments of unicyclic graphs with fixed diameter. Linear Algebra and its Applications, 437(4):1123–1131, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002881.

Chen:2013:ROG

[CLL13a] Xiaolin Chen, Xueliang Li, and Huishu Lian. 4-regular oriented graphs with optimum skew energy. Linear Algebra and its Applications, 439 (10):2948–2960, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004977.

Chen:2013:SER

[CLL13b] Xiaolin Chen, Xueliang Li, and Huishu Lian. The skew energy of random oriented graphs. Linear Algebra and its Applications, 438(11):4547–4556, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001080.

Casas:2013:CSL

[CLOK13] J. M. Casas, M. Ladra, B. A. Omirov, and I. A. Karimjanov. Classification of solvable Leibniz algebras with naturally graded filiform nilradical. Linear Algebra and its Applications, 438(7):2973–3000, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008300.

See corrigendum [LMO16].

Casas:2013:NID

[CLOR13] J. M. Casas, M. Ladra, B. A. Omirov, and U. A. Rozikov. On nilpotent index and dibaricity of evolution algebras. Linear Algebra and its Applications, 439(1):90–105, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001894.

Casas:2011:CEA

[CLR11] J. M. Casas, M. Ladra, and U. A. Rozikov. A chain of evolution algebras. Linear Algebra and its Applications, 435 (4):852–870, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Clark:2010:MMP

[CLS10] Sean Clark, Chi-Kwong Li, and Nung-Sing Sze. Multiplicative maps preserving the higher rank numerical ranges and radii. Linear Algebra and its Applications, 432(11): 2729–2738, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2013:SUB

[CLS13] Yingying Chen, Huiqiu Lin, and Jinlong Shu. Sharp upper bounds on the distance spectral radius of a graph. Linear Algebra and its Applications, 439(9):2659–2666, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004746.

Catral:2014:MGC

[CLST14] Minerva Catral, Leila Lebtahi, Jeffrey Stuart, and Néstor Thome. On a matrix group constructed from an {R, s + 1, k}-potent matrix. Linear Algebra and its Applications, 461(??):200– 210, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005278.

Cai:2013:HAF

[CLX13] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms by Fibonacci gates. Linear Algebra and its Applications, 438(2):690–707, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001625.

Corach:2010:PDO

[CM10a] G. Corach and A. Maestripieri. Polar decomposition of oblique projections. Linear Algebra and its Applications, 433(3):511–519, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:2010:FFT

[CM10b] Liliana Costa and Enide Andrade Martins. Faces of faces of the tridiagonal Birkhoff polytope. Linear Algebra and its Applications, 432(6):1384– 1404, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:2011:HGN

[CM11a] Ada Chan and Akihiro Munemasa. Hamming graphs in Nomura algebras. Linear Algebra and its Applications, 435(2): 330–341, July 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Corach:2011:POP

[CM11b] G. Corach and A. Maestripieri. Products of orthogonal projections and polar decompositions. Linear Algebra and its Applications, 434(6):1594– 1609, March 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chebbi:2012:MHP

[CM12a] Zeineb Chebbi and Maher Moakher. Means of Hermitian positive-definite matrices based on the log-determinant α-divergence function. Linear Algebra and its Applications, 436(7):1872–1889, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100783X.

Chiantini:2012:DRS

[CM12b] Luca Chiantini and Juan Migliore. Determinantal representation and subschemes of general plane curves. Linear Algebra and its Applications, 436(5):1001–1013, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004587.

Chafai:2013:PNS

[CM13] Ezzeddine Chafai and Maher Mnif. Perturbation of normally solvable linear relations in paracomplete spaces. Linear Algebra and its Applications, 439(7):1875–1885, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003613.

Cirrito:2014:OGC

[CM14] Alessio Cirrito and Fabrizio Martino. Ordinary and graded cocharacter of the Jordan algebra of 2 × 2 upper triangular matrices. Linear Algebra and its Applications, 451(??):246–259, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001554.

Czornik:2010:EED

[CMN10] Adam Czornik, Piotr Mokry, and Aleksander Nawrat. On the exponential exponents of discrete linear systems. Linear Algebra and its Applications, 433(4):867–875, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:2014:CRM

[CMNW14] C.-Y. Jean Chan, Meera G. Mainkar, Sivaram K. Narayan, and Jordan D. Webster. A construction of regular magic squares of odd order. Linear Algebra and its Applications, 457(??):293–302, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003231.

Cardoso:2013:EHG

[CMRR13] Domingos M. Cardoso, Enide A. Martins, Maria Robbiano, and Oscar Rojo. Eigenvalues of a H-generalized join graph operation constrained by vertex subsets. Linear Algebra and its Applications, 438(8):3278–3290, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008439.

Chammam:2012:OPC

[CMS12a] Wathek Chammam, Francisco Marcellán, and Ridha Sfaxi. Orthogonal polynomials, Catalan numbers, and a general Hankel determinant evaluation. Linear Algebra and its Applications, 436(7):2105–2116, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100704X.

Conde:2012:OIR

[CMS12b] Cristian Conde, Mohammad Sal Moslehian, and Ameur Seddik. Operator inequalities related to the Corach–Porta–Recht inequality. Linear Algebra and its Applications, 436(9): 3008–3017, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006410.

Covas:2010:LJA

[CMZ10] Ricardo Covas, João Tiago Mexia, and Roman Zmyślony. Lattices of Jordan algebras. Linear Algebra and its Applications, 432(10):2679–2690, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheung:2010:RBZ

[CN10a] W. S. Cheung and T. W. Ng. Relationship between the zeros of two polynomials. Linear Algebra and its Applications, 432(1):107–115, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2010:JNR

[CN10b] Mao-Ting Chien and Hiroshi Nakazato. Joint numerical range and its generating hypersurface. Linear Algebra and its Applications, 432(1): 173–179, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chooi:2010:CAC

[CN10c] Wai Leong Chooi and Wei Shean Ng. On classical adjointcommuting mappings between matrix algebras. Linear Algebra and its Applications, 432 (10):2589–2599, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Czornik:2010:RDL

[CN10d] Adam Czornik and Aleksander Nawrat. On the regularity of discrete linear systems. Linear Algebra and its Applications, 432(11):2745– 2753, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2011:BHR

[CN11a] Mao-Ting Chien and Hiroshi Nakazato. The boundary of higher rank numerical ranges. Linear Algebra and its Applications, 435(11):2971–2985, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2011:CDR

[CN11b] Mao-Ting Chien and Hiroshi Nakazato. Construction of determinantal representation of trigonometric polynomials. Linear Algebra and its Applications, 435(6):1277–1284, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2011:RNR

[CN11c] Mao-Ting Chien and Hiroshi Nakazato. Reduction of the c-numerical range to the classical numerical range. Linear Algebra and its Applications, 434(3):615–624, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chooi:2011:CAC

[CN11d] Wai Leong Chooi and Wei Shean Ng. Classical adjointcommuting mappings on Hermitian and symmetric matrices. Linear Algebra and its Applications, 435(2):202–223, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Czornik:2011:SLA

[CN11e] Adam Czornik and Aleksander Nawrat. Stability by the linear approximation for discrete equations. Linear Algebra and its Applications, 435(4):742–750, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cardon:2012:NMM

[CN12a] David A. Cardon and Pace P. Nielsen. Nonnegative minors of minor matrices. Linear Algebra and its Applications, 436(7):2187–2200, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006902.

Chien:2012:CVH

[CN12b] Mao-Ting Chien and Hiroshi Nakazato. Critical values for higher rank numerical ranges associated with roulette curves. Linear Algebra and its Applications, 437 (9):2117–2127, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004417.

Chien:2012:DRC

[CN12c] Mao-Ting Chien and Hiroshi Nakazato. Determinantal representations of closed orbits. Linear Algebra and its Applications, 437(3):992–1002, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002613.

Chien:2013:HFA

[CN13a] Mao-Ting Chien and Hiroshi Nakazato. Hyperbolic forms associated with cyclic weighted shift matrices. Linear Algebra and its Applications, 439(11):3541– 3554, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005740.

Chien:2013:SPC

[CN13b] Mao-Ting Chien and Hiroshi Nakazato. Singular points of cyclic weighted shift matrices. Linear Algebra and its Applications, 439(12):4090– 4100, December 15, 2013.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006319.

Chien:2013:SCJ

[CN13c] Mao-Ting Chien and Hiroshi Nakazato. Strict convexity of the joint c-numerical range. Linear Algebra and its Applications, 438(3):1305– 1321, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006672.

Climent:2012:CMC

[CNPP12] Joan-Josep Climent, Diego Napp, Carmen Perea, and Raquel Pinto. A construction of MDS 2D convolutional codes of rate 1/n based on superregular matrices. Linear Algebra and its Applications, 437(3):766–780, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002042.

Ceballos:2012:SLA

[CNT12] Manuel Ceballos, Juan Núñez, and Ángel F. Tenorio. Study of Lie algebras by using combinatorial structures. Linear Algebra and its Applications, 436(2):349–363, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

Cohen:2014:CGI

[Coh14] Joel E. Cohen. Chebyshev and Grüss inequalities for real rectangular matrices. Linear Algebra and its Applications, 447(??):133–138, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001869.

Costara:2011:MMP

[Cos11] C. Costara. Maps on matrices that preserve the spectrum. Linear Algebra and its Applications, 435(11):2674–2680, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costara:2014:LSL

[Cos14] Constantin Costara. Local spectrum linear preservers at non-fixed vectors. Linear Algebra and its Applications, 457(??):154–161, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400322X.

Catral:2010:GDG

[COvdD10] M. Catral, D. D. Olesky, and P. van den Driessche. Graphical description of group inverses of certain bipartite matrices. Linear Algebra and its Applications, 432(1):36– 52, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cortes:2010:RNP

[CP10] V. Cortes and J. M. Peña. Required nonzero patterns for nonsingular sign regular matrices. Linear Algebra and its Applications, 432(8):1990– 1994, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chorianopoulos:2011:BJA

[CP11] Christos Chorianopoulos and Panayiotis Psarrakos. Birkhoff– James approximate orthogonality sets and numerical ranges. Linear Algebra and its Applications, 434(9):2089– 2108, May 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Casazza:2012:EIP

[CP12] Peter G. Casazza and Jesse Peterson. An elementary, illustrative proof of the Rado-Horn theorem. Linear Algebra and its Applications, 437 (10):2523–2537, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004521.

Carlson:2011:ESC

[CPH11] David H. Carlson, Russell T.

Potter, and Richard D. Hill. The equality of the Siler cones K_{3} and K_{4}, the general case. Linear Algebra and its Applications, 434(3):723– 729, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Church:2011:MMS

[CPK11] Amber Church, Rajesh Pereira, and David Kribs. Majorization and multiplier sequences. Linear Algebra and its Applications, 435(9):2132–2139, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chandrashekaran:2010:SM

[CPR10] A. Chandrashekaran, T. Parthasarathy, and G. Ravindran. On strong Z-matrices. Linear Algebra and its Applications, 432(4): 964–969, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chandrashekaran:2010:PSL

[CPV10] A. Chandrashekaran, T. Parthasarathy, and V. Vetrivel. On the P′_{2} and P_{2}-properties in the semidefinite linear complementarity problem. Linear Algebra and its Applications, 432(1):134–143, January 1, 2010. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2013:SVP

[CPZ13] K. C. Chang, K. J. Pearson, and Tan Zhang. Some variational principles for Zeigenvalues of nonnegative tensors. Linear Algebra and its Applications, 438(11): 4166–4182, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001146.

Chen:2013:SML

[CQYY13] Zhongming Chen, Liqun Qi, Qingzhi Yang, and Yuning Yang. The solution methods for the largest eigenvalue (singular value) of nonnegative tensors and convergence analysis. Linear Algebra and its Applications, 439 (12):3713–3733, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005831.

Cardoso:2010:SUB

[CR10a] Domingos M. Cardoso and Peter Rowlinson. Spectral upper bounds for the order of a k-regular induced subgraph. Linear Algebra and its Applications, 433(5):1031– 1037, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-

Carlini:2010:PMN

[CR10b] Enrico Carlini and Fabio Rapallo. Probability matrices, non-negative rank, and parameterization of mixture models. Linear Algebra and its Applications, 433(2):424– 432, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carvalho:2010:IGR

[CR10c] Paula Carvalho and Paula Rama. Integral graphs and (κ, τ)-regular sets. Linear Algebra and its Applications, 432(9):2409–2417, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Crane:2013:SAI

[Cra13] Harry Crane. Some algebraic identities for the αpermanent. Linear Algebra and its Applications, 439 (11):3445–3459, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005843.

Camara:2014:OSI

[CRS14] M. C. Câmara, L. Rodman, and I. M. Spitkovsky. One sided invertibility of matrices over commutative rings, corona problems, and Toeplitz operators with matrix symbols. Linear Algebra and its Applications, 459(??):58–82, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004078.

Camenga:2014:GWN

[CRSS14] Kristin A. Camenga, Patrick X. Rault, Tsvetanka Sendova, and Ilya M. Spitkovsky. On the Gau–Wu number for some classes of matrices. Linear Algebra and its Applications, 444(??):254–262, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007830.

Canto:2010:CRT

[CRU10] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Characterizations of rectangular totally and strictly totally positive matrices. Linear Algebra and its Applications, 432(10): 2623–2633, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Canto:2013:QFN

[CRU13] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Quasi-LDU factorization of nonsingular totally nonpositive matrices. Linear Algebra and its Applications, 439(4):836–851, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004454.

Canto:2014:FRF

[CRU14] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Full rank factorization in quasi-LDU form of totally nonpositive rectangular matrices. Linear Algebra and its Applications, 440(??):61–82, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300699X.

Ccapa:2010:EDP

[CS10a] Javier Ccapa and Ricardo L. Soto. On elementary divisors perturbation of nonnegative matrices. Linear Algebra and its Applications, 432(2– 3):546–555, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2010:TST

[CS10b] Dragoš Cvetković and Slobodan K. Simić. Towards a spectral theory of graphs based on the signless Laplacian, II. Linear Algebra and its Applications, 432(9):2257–2272, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2011:GSC

[CS11] Dragoš Cvetković and Slobodan Simić. Graph spectra in Computer Science. Linear Algebra and its Applications, 434(6):1545–1562, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cartwright:2013:NET

[CS13a] Dustin Cartwright and Bernd Sturmfels. The number of eigenvalues of a tensor. Linear Algebra and its Applications, 438(2):942–952, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004629.

Chen:2013:HFG

[CS13b] Haimiao Chen and Hao Shen. How to find G-admissible abelian regular coverings of a graph? Linear Algebra and its Applications, 438(8): 3303–3320, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000426.

Costa:2010:CGT

[CSAC10] S. I. R. Costa, J. E. Strapasson, M. M. S. Alves, and T. B. Carlos. Circulant graphs and tessellations on flat tori. Linear Algebra and its Applications, 432(1):369–382, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:2011:CGT

[CSAC11] S. I. R. Costa, J. E. Strapasson, M. M. S. Alves, and T. B. Carlos. Circulant graphs and tessellations on flat tori. Linear Algebra and its Applications, 434(8):1811–1823, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Campello:2013:PAL

[CSC13] Antonio Campello, João Strapasson, and Sueli I. R. Costa. On projections of arbitrary lattices. Linear Algebra and its Applications, 439(9):2577– 2583, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004734.

Canuto:2014:DIM

[CSV14] C. Canuto, V. Simoncini, and M. Verani. On the decay of the inverse of matrices that are sum of Kronecker products. Linear Algebra and its Applications, 452(??):21–39, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001736.

Cardoso:2010:MER

[CSZ10] Domingos M. Cardoso, Irene Sciriha, and Cheryl Zerafa. Main eigenvalues and (κ, τ)- regular sets. Linear Algebra and its Applications, 432 (9):2399–2408, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2010:GMS

[CT10] Ting-Jung Chang and Bit-Shun Tam. Graphs with maximal signless Laplacian spectral radius. Linear Algebra and its Applications, 432(7): 1708–1733, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cardon:2011:JCF

[CT11a] David A. Cardon and Bradford Tuckfield. The Jordan canonical form for a class of zero-one matrices. Linear Algebra and its Applications, 435(11):2942–2954, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2011:CGM

[CT11b] Ting-Chung Chang and Bit-Shun Tam. Connected graphs with maximal Q-index: The one-dominating-vertex case. Linear Algebra and its Applications, 435(10):2451–2461, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-

Cui:2012:SSL

[CT12] Shu-Yu Cui and Gui-Xian Tian. The spectrum and the signless Laplacian spectrum of coronae. Linear Algebra and its Applications, 437(7):1692–1703, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003813.

Cagliero:2014:CFL

[CT14a] Leandro Cagliero and Paulo Tirao. The cohomology of filiform Lie algebras of maximal rank. Linear Algebra and its Applications, 455(??): 143–167, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400278X.

Cuitino:2014:DCP

[CT14b] Luis Felipe Tapia Cuitiño and Andrea Luigi Tironi. Dual codes of product semi-linear codes. Linear Algebra and its Applications, 457(??):114– 153, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002912.

Cui:2013:SUB

[CTG13] Shu-Yu Cui, Gui-Xian Tian, and Jing-Jing Guo. A sharp upper bound on the signless Laplacian spectral radius of graphs. Linear Algebra and its Applications, 439 (8):2442–2447, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300414X.

Chang:2011:TPM

[CTW11] Ting-Chung Chang, Bit-Shun Tam, and Shu-Hui Wu. Theorems on partitioned matrices revisited and their applications to graph spectra. Linear Algebra and its Applications, 434(2):559–581, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Corrales:2013:CIG

[CV13] Hugo Corrales and Carlos E. Valencia. On the critical ideals of graphs. Linear Algebra and its Applications, 439 (12):3870–3892, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006307.

Cioaba:2010:ARS

[CvDKL10] Sebastian M. Cioabă, Edwin R. van Dam, Jack H. Koolen, and Jae-Ho Lee.

Asymptotic results on the spectral radius and the diameter of graphs. Linear Algebra and its Applications, 432(2– 3):722–737, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Camara:2013:GAW

[CvDKP13] Marc Cámara, Edwin R. van Dam, Jack H. Koolen, and Jongyook Park. Geometric aspects of 2-walk-regular graphs. Linear Algebra and its Applications, 439 (9):2692–2710, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004795.

Cheng:2010:CMF

[CVW10] Che-Man Cheng, Seak-Weng Vong, and David Wenzel. Commutators with maximal Frobenius norm. Linear Algebra and its Applications, 432 (1):292–306, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2010:SBL

[CW10] Yanqing Chen and Ligong Wang. Sharp bounds for the largest eigenvalue of the signless Laplacian of a graph. Linear Algebra and its Applications, 433(5):908–913, October 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2011:FDM

[CW11] Changhao Chen and Shengyou Wen. Factorization of deltamonotone linear mappings. Linear Algebra and its Applications, 435(8):2087–2095, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2012:SRT

[CW12a] Gi-Sang Cheon and Ian M. Wanless. Some results towards the Dittert conjecture on permanents. Linear Algebra and its Applications, 436(4):791–801, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379510004490.

Cioaba:2012:EDS

[CW12b] Sebastian M. Cioaba and Wiseley Wong. Edge-disjoint spanning trees and eigenvalues of regular graphs. Linear Algebra and its Applications, 437(2):630–647, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002133.

Cameron:2011:HDS

[CY11] Naiomi T. Cameron and Andrew C. M. Yip. Hankel determinants of sums of consecutive Motzkin numbers. Linear Algebra and its Applications, 434(3):712–722, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Czornik:2010:PED

[Czo10] Adam Czornik. On the Perron exponents of discrete linear systems. Linear Algebra and its Applications, 432(1):394– 401, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dong:2010:NCP

[DA10] Hongbo Dong and Kurt Anstreicher. A note on “5 × 5 completely positive matrices”. Linear Algebra and its Applications, 433(5):1001– 1004, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [BAD09].

Dahl:2010:MPM

[Dah10]  Geir Dahl.  Majorization

permutahedra and (0,1)- matrices. Linear Algebra and its Applications, 432(12): 3265–3271, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:2011:PRI

[Dah11] Geir Dahl. Polytopes related to interval vectors and incidence matrices. Linear Algebra and its Applications, 435 (11):2955–2960, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:2012:MMC

[Dah12a] Geir Dahl. Martingale matrix classes and polytopes. Linear Algebra and its Applications, 437(7):1722–1733, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003291.

Dahl:2012:MBR

[Dah12b] Geir Dahl. A matrix-based ranking method with application to tennis. Linear Algebra and its Applications, 437(1):26–36, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001218.

Dai:2011:EBL

[Dai11] Ping-Fan Dai. Error bounds for linear complementarity problems of DB-matrices. Linear Algebra and its Applications, 434(3):830–840, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dai:2012:GTS

[Dai12] Xiongping Dai. A Gel’fandtype spectral radius formula and stability of linear constrained switching systems. Linear Algebra and its Applications, 436(5): 1099–1113, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005416.

Dai:2013:SCS

[Dai13] Xiongping Dai. Some criteria for spectral finiteness of a finite subset of the real matrix space R^{d×d}. Linear Algebra and its Applications, 438(6):2717–2727, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007409.

Das:2010:CIA

[Das10a] Kinkar Ch. Das. Conjectures on index and algebraic connectivity of graphs. Linear Algebra and its Applications, 433(8–10):1666–1673, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2010:CIS

[Das10b] Kinkar Ch. Das. On conjectures involving second largest signless Laplacian eigenvalue of graphs. Linear Algebra and its Applications, 432(11): 3018–3029, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2011:PCI

[Das11] Kinkar Ch. Das. Proof of conjecture involving the second largest signless Laplacian eigenvalue and the index of graphs. Linear Algebra and its Applications, 435 (10):2420–2424, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2013:PCI

[Das13] Kinkar Ch. Das. Proof of conjectures involving algebraic connectivity of graphs. Linear Algebra and its Applications, 438(8):3291–3302, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000050.

Daugulis:2012:PMC

[Dau12] Peteris Daugulis. A parametrization of matrix conjugacy orbit sets as unions of affine planes. Linear Algebra and its Applications, 436(3):709– 721, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005726.

Dax:2010:EPO

[Dax10] Achiya Dax. On extremum properties of orthogonal quotients matrices. Linear Algebra and its Applications, 432 (5):1234–1257, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duan:2014:GLR

[DBZZ14] Xuefeng Duan, Jianchao Bai, Maojun Zhang, and Xinjun Zhang. On the generalized low rank approximation of the correlation matrices arising in the asset portfolio. Linear Algebra and its Applications, 461(??):1–17, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004765.

Deng:2013:KIC

[DC13] Qingying Deng and Haiyan Chen. On the Kirchhoff index of the complement of a bipartite graph. Linear Algebra and its Applications, 439(1):167–173, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001924.

Deng:2014:EBU

[DC14] Qingying Deng and Haiyan Chen. On extremal bipartite unicyclic graphs. Linear Algebra and its Applications, 444(??):89–99, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007763.

daCruz:2012:PMS

[dCF12] Henrique F. da Cruz and Rosário Fernandes. On pairs of matrices that satisfy certain polynomial identities. Linear Algebra and its Applications, 436(6):1589–1605, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006598.

Deng:2012:ICP

[DCIW12a] Chunyuan Deng, Dragana S. Cvetković-Ilić, and Yimin Wei. On invertibility of combinations of k-potent operators. Linear Algebra and its Applications, 437(1):376– 387, July 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001413.

Deng:2012:PCC

[DCIW12b] Chunyuan Deng, Dragana S. Cvetković-Ilić, and Yimin Wei. Properties of the combinations of commutative idempotents. Linear Algebra and its Applications, 436(1): 202–221, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

DeTeran:2010:FOS

[DD10a] Fernando De Terán and Froilán M. Dopico. First order spectral perturbation theory of square singular matrix polynomials. Linear Algebra and its Applications, 432(4): 892–910, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dinh:2010:MPZ

[DD10b] Toan Duc Dinh and Michael Donzella. On maps preserving zeros of Lie polynomials of small degrees. Linear Algebra and its Applications, 432(2– 3):493–498, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dogan-Dunlap:2010:LAS

[DD10c] Hamide Dogan-Dunlap. Linear algebra students’ modes of reasoning: Geometric representations. Linear Algebra and its Applications, 432(8): 2141–2159, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeTeran:2011:SEA

[DD11a] Fernando De Terán and Froilán M. Dopico. The solution of the equation XA + AX^{T} = 0 and its application to the theory of orbits. Linear Algebra and its Applications, 434(1):44–67, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Derevyagin:2011:DTJ

[DD11b] Maxim Derevyagin and Vladimir Derkach. Darboux transformations of Jacobi matrices and Padé approximation. Linear Algebra and its Applications, 435(12):3056–3084, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dincic:2011:MTR

[DD11c] Nebojsa C. Dincić and Dragan S. Djordjević. Mixed-type reverse order law for products of three operators. Linear Algebra and its Applications, 435(11):2658–2673, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dou:2011:NOP

[DD11d] Yan-Ni Dou and Hong-Ke Du. A note on operator probability theory involving numerical ranges. Linear Algebra and its Applications, 435 (12):3233–3242, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duchsherer:2014:SPD

[DD14] Melissa Duchsherer and Benton L. Duncan. Some point derivations for semicrossed products. Linear Algebra and its Applications, 459(??): 237–247, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004613.

Duffner:2013:RBD

[DdC13] M. Antónia Duffner and Henrique F. da Cruz. A relation between the determinant and the permanent on singular matrices. Linear Algebra and its Applications, 438 (9):3654–3660, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000165.

Dangovski:2013:WDF

[DDF13a] Rumen Dangovski, Vesselin Drensky, and Sehmus Findik. Weitzenböck derivations of free metabelian Lie algebras. Linear Algebra and its Applications, 439(10):3279– 3296, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003157.

Du:2013:SAM

[DdF13b] Zhibin Du and C. M. da Fonseca. The singular acyclic matrices with maximal number of P-vertices. Linear Algebra and its Applications, 438(5):2274–2279, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007331.

Du:2014:NAM

[DdF14] Zhibin Du and C. M. da Fonseca. Nonsingular acyclic matrices with an extremal number of P -vertices. Linear Algebra and its Applications, 442(??):2–19, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000591.

DeTeran:2013:SE

[DDG^{+}13] Fernando De Terán, Froilán M. Dopico, Nathan Guillery, Daniel Montealegre, and Nicolás Reyes. The solution of the equation AX + X^{∗}B = 0. Linear Algebra and its Applications, 438 (7):2817–2860, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008191.

Dickinson:2013:IEC

[DDGH13] Peter J. C. Dickinson, Mirjam Dür, Luuk Gijben, and Roland Hildebrand. Irreducible elements of the copositive cone. Linear Algebra and its Applications, 439 (6):1605–1626, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007240.

Duggal:2014:EOF

[DDK14] B. P. Duggal, S. V. Djordjević, and C. S. Kubrusly. Elementary operators, finite ascent, range closure and compactness. Linear Algebra and its Applications, 449 (??):334–349, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000871.

DeTeran:2012:FCL

[DDM12] Fernando De Terán, Froilán M. Dopico, and D. Steven Mackey. Fiedler companion linearizations for rectangular matrix polynomials. Linear Algebra and its Applications, 437(3):957–991, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002595.

Teran:2014:SEM

[DDM14] Fernando De Terán, Froilán M. Dopico, and D. Steven Mackey. Spectral equivalence of matrix polynomials and the Index Sum Theorem. Linear Algebra and its Applications, 459(??):264–333, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400456X.

Teran:2013:CNI

[DDP13] Fernando De Terán, Froilán M. Dopico, and Javier Pérez. Condition numbers for inversion of Fiedler companion matrices. Linear Algebra and its Applications, 439(4):944–981, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006970.

Teran:2014:NBR

[DDP14] Fernando De Terán, Froilán M. Dopico, and Javier Pérez. New bounds for roots of polynomials based on Fiedler companion matrices. Linear Algebra and its Applications, 451 (??):197–230, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001578.

DeBruyn:2011:PVS

[De 11] Bart De Bruyn. On polyvectors of vector spaces and hyperplanes of projective Grassmannians. Linear Algebra and its Applications, 435(5):1055– 1084, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deSeguinsPazzis:2013:GTS

[de 13] Clément de Seguins Pazzis. On Gerstenhaber’s theorem for spaces of nilpotent matrices over a skew field. Linear Algebra and its Applications, 438(11):4426–4438, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000992.

Deaett:2011:MSR

[Dea11] Louis Deaett. The minimum semidefinite rank of a triangle-free graph. Linear Algebra and its Applications, 434(8):1945–1955, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2010:PGQ

[Den10] Chun Yuan Deng. On properties of generalized quadratic operators. Linear Algebra and its Applications, 432(4):847– 856, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2011:CRG

[Den11] Chun Yuan Deng. Characterizations and representations of the group inverse involving idempotents. Linear Algebra and its Applications, 434(4): 1067–1079, February 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Derksen:2013:KUI

[Der13] Harm Derksen. Kruskal’s uniqueness inequality is sharp. Linear Algebra and its Applications, 438(2):708–712, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004873.

daFonseca:2010:PTM

[dF10] C. M. da Fonseca. The μ-permanent of a tridiagonal matrix, orthogonal polynomials, and chain sequences. Linear Algebra and its Applications, 432(5):1258–1266, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2014:PLM

[DF14] Guodong Deng and Yun Fan. Permutation-like matrix groups with a maximal cycle of prime square length. Linear Algebra and its Applications, 450(??):44–55, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400127X.

deFreitas:2014:MAD

[dFBRS14] Maria Aguieiras A. de Freitas, Andréa Soares Bonifácio, Maria Robbiano, and Bernardo San Martín. On matrices associated to directed graphs and applications. Linear Algebra and its Applications, 442(??):156–164, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004564.

deFreitas:2010:IFI

[dFdADVJ10] Maria Aguieiras A. de Freitas, Nair M. M. de Abreu, Renata R. Del-Vecchio, and Samuel Jurkiewicz. Infinite families of Q-integral graphs. Linear Algebra and its Applications, 432(9):2352–2360, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dalfo:2010:CWR

[DFG10] C. Dalfó, M. A. Fiol, and E. Garriga. Characterizing (ℓ, m)-walk-regular graphs. Linear Algebra and its Applications, 433(11–12):1821– 1826, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Denman:2013:SFD

[DFR13] Richard T. Denman, Fumiko Futamura, and Kendall C. Richards. On sharp frame diagonalization. Linear Algebra and its Applications, 438(5):2210–2224, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007446.

Dmytryshyn:2012:MDM

[DFS12] Andrii R. Dmytryshyn, Vyacheslav Futorny, and Vladimir V. Sergeichuk. Miniversal deformations of matrices of bilinear forms. Linear Algebra and its Applications, 436 (7):2670–2700, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007518.

Dmytryshyn:2014:MDM

[DFS14] Andrii Dmytryshyn, Vyacheslav Futorny, and Vladimir V. Sergeichuk. Miniversal deformations of matrices under *congruence and reducing transformations. Linear Algebra and its Applications, 446(??):388–420, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000342.

Das:2014:IEG

[DG14] Kinkar Ch. Das and Ivan Gutman. On incidence energy of graphs. Linear Algebra and its Applications, 446 (??):329–344, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008343.

Derevyagin:2014:MGT

[DGAM14] M. Derevyagin, J. C. García-Ardila, and F. Marcellán. Multiple Geronimus transformations. Linear Algebra and its Applications, 454(??): 158–183, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002493.

Das:2014:LEL

[DGÇ14] Kinkar Ch. Das, Ivan Gutman, and A. Sinan Çevik. On the Laplacian-energy-like invariant. Linear Algebra and its Applications, 442(??): 58–68, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003133.

Diaz-Garcia:2011:WDS

[DGGJ11] José A. Díaz-García and Ramón Gutiérrez-Jáimez. On Wishart distribution: Some extensions. Linear Algebra and its Applications, 435 (6):1296–1310, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Diaz-Garcia:2013:SE

[DGGJ13] José A. Diáz-Garciá and Ramón Gutiérrez-Jáimez. Spherical ensembles. Linear Algebra and its Applications, 438(8):3174–3201, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000396.

DeLoss:2010:TDM

[DGH^{+}10] Laura DeLoss, Jason Grout, Leslie Hogben, Tracy McKay, Jason Smith, and Geoff Tims. Techniques for determining the minimum rank of a small graph. Linear Algebra and its Applications, 432(11):2995– 3001, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dunkl:2011:NSM

[DGH^{+}11] Charles F. Dunkl, Piotr Gawron, John A. Holbrook, Zbigniew Puchała, and Karol Życzkowski. Numerical shadows: Measures and densities on the numerical range. Linear Algebra and its Applications, 434(9):2042–2080, May 1, 2011. CODEN LAAPAW.

Diaz-Garcia:2010:CBV

[DGJ10] José A. Díaz-García and Ramón Gutiérrez Jáimez. Complex bimatrix variate generalised beta distributions. Linear Algebra and its Applications, 432(2–3):571–582, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolinar:2013:EMC

[DGKO13] Gregor Dolinar, Alexander Guterman, Bojan Kuzma, and Polona Oblak. Extremal matrix centralizers. Linear Algebra and its Applications, 438(7):2904–2910, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008658.

Dumas:2010:SMS

[DGMS10] Jean-Guillaume Dumas, Rod Gow, Gary McGuire, and John Sheekey. Subspaces of matrices with special rank properties. Linear Algebra and its Applications, 433(1): 191–202, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dixon:2014:TTF

[DGMS14] John Dixon, Michael Goldenberg, Ben Mathes, and Justin Sukiennik. Triangular truncation of k-Fibonacci and k-Lucas circulant matrices. Linear Algebra and its Applications, 440(??): 177–187, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005818.

Dyer:2014:SEH

[DGU14] Martin Dyer, Catherine Greenhill, and Mario Ullrich. Structure and eigenvalues of heatbath Markov chains. Linear Algebra and its Applications, 454(??):57–71, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002328.

Dikranjan:2013:EAG

[DGZ13] D. Dikranjan, K. Gong, and P. Zanardo. Endomorphisms of abelian groups with small algebraic entropy. Linear Algebra and its Applications, 439(7):1894–1904, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003637.

Dehghan:2010:GCM

[DH10a] Mehdi Dehghan and Masoud Hajarian. The general coupled matrix equations over generalized bisymmetric matrices.

Linear Algebra and its Applications, 432(6):1531–1552, March 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Donatelli:2010:CNA

[DH10b] Marco Donatelli and Martin Hanke. On the condition number of the antireflective transform. Linear Algebra and its Applications, 432(7):1772– 1784, March 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dehghan:2012:GRA

[DH12a] Mehdi Dehghan and Masoud Hajarian. On the generalized reflexive and anti-reflexive solutions to a system of matrix equations. Linear Algebra and its Applications, 437 (11):2793–2812, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005083.

Dong:2012:SAC

[DH12b] Aiju Dong and Chengjun Hou. On some automorphisms of a class of Kadison– Singer algebras. Linear Algebra and its Applications, 436(7):2037–2053, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006288.

Datta:2012:GWB

[DHC12] S. Datta, S. Howard, and D. Cochran. Geometry of the Welch bounds. Linear Algebra and its Applications, 437 (10):2455–2470, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004405.

Dolinar:2013:SNM

[DHKQ13] Gregor Dolinar, Jinchuan Hou, Bojan Kuzma, and Xiaofei Qi. Spectrum nonincreasing maps on matrices. Linear Algebra and its Applications, 438(8):3504– 3510, April 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000347.

Hou:2013:DSM

[dHLMS13] Xiang dong Hou, Ana G. Lecuona, Gary L. Mullen, and James A. Sellers. On the dimension of the space of magic squares over a field. Linear Algebra and its Applications, 438(8):3463–3475, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008695.

Dai:2012:FSR

[DHLX12] Xiongping Dai, Yu Huang, Jun Liu, and Mingqing Xiao.

The finite-step realizability of the joint spectral radius of a pair of d × d matrices one of which being rank-one. Linear Algebra and its Applications, 437(7):1548–1561, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003564.

deHoog:2011:NSS

[dHM11] F. R. de Hoog and R. M. M. Mattheij. A note on subset selection for matrices. Linear Algebra and its Applications, 434(8):1845–1850, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duncan:2010:ETF

[DHS10] David M. Duncan, Thomas R. Hoffman, and James P. Solazzo. Equiangular tight frames and fourth root Seidel matrices. Linear Algebra and its Applications, 432(11): 2816–2823, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dinitz:2012:TDI

[DHS12] Thomas Dinitz, Matthew Hartman, and Jenya Soprunova. Tropical determinant of integer doublystochastic matrices. Linear Algebra and its Applications, 436(5):1212–1227, March 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005738.

Dinwoodie:2011:SMB

[Din11] Ian H. Dinwoodie. Syzygies for Metropolis base chains. Linear Algebra and its Applications, 434(10):2176–2186, May 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dascalescu:2013:FSM

[DIP13] S. Dascalescu, M. C. Iovanov, and S. Predut. Frobenius structural matrix algebras. Linear Algebra and its Applications, 439(10):3166– 3172, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005624.

Ducey:2014:IIA

[DJ14] Joshua E. Ducey and Deelan M. Jalil. Integer invariants of abelian Cayley graphs. Linear Algebra and its Applications, 445(??):316–325, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008021.

Duggal:2012:QCC

[DJK12a] B. P. Duggal, I. H. Jeon, and I. H. Kim. On quasiclass A contractions. Linear Algebra and its Applications, 436(9):3562–3567, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008627.

Duggal:2012:PCP

[DJK12b] Bhagwati P. Duggal, In Ho Jeon, and In Hyoun Kim. On ⋆-paranormal contractions and properties for ⋆-class A operators. Linear Algebra and its Applications, 436 (5):954–962, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004381.

DeBruyn:2011:TDS

[DK11] B. De Bruyn and M. Kwiatkowski. On the trivectors of a 6dimensional symplectic vector space. Linear Algebra and its Applications, 435(2):289–306, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2012:TDS

[DK12a] B. De Bruyn and M. Kwiatkowski. On the trivectors of a 6dimensional symplectic vector space. II. Linear Algebra and its Applications, 437 (5):1215–1233, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200290X.

Devecic:2012:CGP

[DK12b] Ömür Devecić and Erdal Karaduman. The cyclic groups via the Pascal matrices and the generalized Pascal matrices. Linear Algebra and its Applications, 437 (10):2538–2545, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004727.

Bruyn:2013:CTS

[DK13a] B. De Bruyn and M. Kwiatkowski. The classification of the trivectors of a six-dimensional symplectic space: Summary, consequences and connections. Linear Algebra and its Applications, 438(8):3516– 3529, April 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000487.

DeBruyn:2013:TDS

[DK13b] B. De Bruyn and M. Kwiatkowski. On the trivectors of a 6dimensional symplectic vector space. III. Linear Algebra and its Applications, 438(1):374–398, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Bruyn:2013:TDS

[DK13c] B. De Bruyn and M. Kwiatkowski. On the trivectors of a 6dimensional symplectic vector space. IV. Linear Algebra and its Applications, 438(5):2405–2429, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007501.

Deng:2013:SDT

[DK13d] Aiping Deng and Alexander Kelmans. Spectra of digraph transformations. Linear Algebra and its Applications, 439(1):106–132, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001882.

Dym:2013:TFC

[DK13e] Harry Dym and David P. Kimsey. Trace formulas for a class of truncated block Toeplitz operators. Linear Algebra and its Applications, 439(10):3070–3099, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004989.

Duggal:2014:PFC

[DK14] B. P. Duggal and C. S. Kubrusly. A Putnam– Fuglede commutativity property for Hilbert space operators. Linear Algebra and its Applications, 458(??): 108–115, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003772.

Draper:2014:FRF

[DKM^{+}14] Bruce Draper, Michael Kirby, Justin Marks, Tim Marrinan, and Chris Peterson. A flag representation for finite collections of subspaces of mixed dimensions. Linear Algebra and its Applications, 451 (??):15–32, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001669.

Dolgov:2012:RPS

[DKOT12] Sergey Dolgov, Boris N. Khoromskij, Ivan Oseledets, and Eugene Tyrtyshnikov. A reciprocal preconditioner for structured matrices arising from elliptic problems with jumping coefficients. Linear Algebra and its Applications, 436(9):2980–3007, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

DiVincenzo:2010:GIT

[DKS10] Onofrio Mario Di Vincenzo, Plamen Koshlukov, and Ednei Aparecido Santulo Jr. Graded identities for tensor products of matrix (super)algebras over the Grassmann algebra. Linear Algebra and its Applications, 432(2– 3):780–795, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DAndrea:2013:SMR

[DKS13a] Carlos D’Andrea, Teresa Krick, and Agnes Szanto. Subresultants in multiple roots. Linear Algebra and its Applications, 438(5):1969– 1989, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007811.

Dmytryshyn:2013:SSM

[DKS13b] Andrii Dmytryshyn, Bo Kågström, and Vladimir V. Sergeichuk. Skew-symmetric matrix pencils: Codimension counts and the solution of a pair of matrix equations. Linear Algebra and its Applications, 438(8):3375–3396, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008324.

Du:2011:ELE

[DL11] Zhibin Du and Zhongzhu Liu. On the Estrada and Laplacian Estrada indices of graphs. Linear Algebra and its Applications, 435(8):2065–2076, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2013:EDG

[DL13] S. W. Drury and Huiqiu Lin. Extremal digraphs with given clique number. Linear Algebra and its Applications, 439 (2):328–345, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300219X.

Davies:2014:SCN

[DL14] E. B. Davies and Michael Levitin. Spectra of a class of nonself-adjoint matrices. Linear Algebra and its Applications, 448(??):55–84, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000433.

delaCruz:2014:CDS

[dlCdlRMP14] Ralph John de la Cruz, Kennett L. de la Rosa, Dennis I. Merino, and Agnes T. Paras. The Cartan–Dieudonné–Scherk theorems for complex Sorthogonal matrices. Linear Algebra and its Applications, 458(??):251–260, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003723.

Du:2011:ILI

[DLDV11] Nguyen Huu Du, Le Cong Loi, Trinh Khanh Duy, and Vu Tien Viet. On index-2 linear implicit difference equations. Linear Algebra and its Applications, 434(2):394– 414, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2011:ERG

[DLL11] Wenxue Du, Xueliang Li, and Yiyang Li. The energy of random graphs. Linear Algebra and its Applications, 435 (10):2334–2346, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2010:NNE

[DLLS10] Wenxue Du, Xueliang Li, Yiyang Li, and Simone Severini. A note on the von Neumann entropy of random graphs. Linear Algebra and its Applications, 433 (11–12):1722–1725, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dubsky:2014:CIS

[DLMZ14] Brendan Dubsky, Rencai Lü, Volodymyr Mazorchuk, and Kaiming Zhao. Category 0 for the Schrödinger algebra. Linear Algebra and its Applications, 460(??): 17–50, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004807.

deLima:2013:SLE

[dLN13] Leonardo Silva de Lima and Vladimir Nikiforov. On the second largest eigenvalue of the signless Laplacian. Linear Algebra and its Applications, 438(3):1215–1222, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006155.

DeTeran:2014:FTM

[DLNN14] Fernando De Terán, Ross A. Lippert, Yuji Nakatsukasa, and Vanni Noferini. Flanders’ theorem for many matrices under commutativity assumptions. Linear Algebra and its Applications, 443 (??):120–138, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007258.

deLima:2011:SES

[dLOdAN11] Leonardo Silva de Lima, Carla Silva Oliveira, Nair Maria Maia de Abreu, and Vladimir Nikiforov. The smallest eigenvalue of the signless Laplacian. Linear Algebra and its Applications, 435(10):2570–2584, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

delaPuente:2011:TLM

[dlP11] M. J. de la Puente. Tropical linear maps on the plane. Linear Algebra and its Applications, 435(7):1681–1710, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

delaRosa:2012:HM

[dlRMP12] Kennett L. de la Rosa, Dennis I. Merino, and Agnes T. Paras. The J–Householder matrices. Linear Algebra and its Applications, 436(5): 1189–1194, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005684.

Das:2014:UBS

[DLS14] Kinkar Ch. Das, Muhuo Liu, and Haiying Shan. Upper bounds on the (signless) Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 459(??):

334–341, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004686.

Dumitru:2013:SVI

[DLV13] Raluca Dumitru, Rachel Levanger, and Bogdan Visinescu. On singular value inequalities for matrix means. Linear Algebra and its Applications, 439(8):2405–2410, October 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004655.

Dolinar:2011:SPO

[DM11] Gregor Dolinar and Janko Marovt. Star partial order on B(H). Linear Algebra and its Applications, 434(1):319– 326, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dohi:2010:DTB

[DMMY10] Ryota Dohi, Yoichi Maeda, Masahiro Mori, and Hirokazu Yoshida. A dual transformation between SO(n + 1) and SO(n, 1) and its geometric applications. Linear Algebra and its Applications, 432(2– 3):770–776, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DelaCruz:2011:PDM

[DMP11] Ralph John De la Cruz, Dennis I. Merino, and Agnes T. Paras. The ϕ_{S} polar decomposition of matrices. Linear Algebra and its Applications, 434(1):4–13, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deMalafosse:2012:SSA

[dMR12] Bruno de Malafosse and Vladimir Rakocević. Series summable (C, λ, μ) and applications. Linear Algebra and its Applications, 436(11): 4089–4100, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007476.

Dopazo:2010:FRR

[DMS10] E. Dopazo and M. F. Martínez-Serrano. Further results on the representation of the Drazin inverse of a 2 × 2 block matrix. Linear Algebra and its Applications, 432(8): 1896–1904, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dopazo:2013:DDI

[DMS13] E. Dopazo and M. F. Martínez-Serrano. On deriving the Drazin inverse of a modified matrix. Linear Algebra and its Applications, 438 (4):1678–1687, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004721.

Dolzan:2011:CGM

[DO11a] David Dolžan and Polona Oblak. Commuting graphs of matrices over semirings. Linear Algebra and its Applications, 435(7):1657–1665, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolzan:2011:NGM

[DO11b] David Dolžan and Polona Oblak. Noncommuting graphs of matrices over semirings. Linear Algebra and its Applications, 435(7):1649–1656, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deOliveira:2012:SSP

[dO12] Maurício C. de Oliveira. Simplification of symbolic polynomials on non-commutative variables. Linear Algebra and its Applications, 437(7): 1734–1748, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003679.

Dodig:2010:ESR

[Dod10] Marija Dodig. Explicit solution of the row completion problem for matrix pencils. Linear Algebra and its Applications, 432(5):1299–1309, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2013:CMP

[Dod13] Marija Dodig. Completion up to a matrix pencil with column minimal indices as the only nontrivial Kronecker invariants. Linear Algebra and its Applications, 438(8): 3155–3173, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000414.

deOliveira:2013:CLS

[dOFK^{+}13] Debora Duarte de Oliveira, Vyacheslav Futorny, Tatiana Klimchuk, Dmitry Kovalenko, and Vladimir V. Sergeichuk. Cycles of linear and semilinear mappings. Linear Algebra and its Applications, 438(8):3442–3453, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000104.

deOliveira:2012:RCP

[dOHKS12] Debora Duarte de Oliveira, Roger A. Horn, Tatiana Klimchuk, and Vladimir V. Sergeichuk. Remarks on the classification of a pair of commuting semilinear operators. Linear Algebra and its Applications, 436(9):3362–3372, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007725.

Domanov:2010:ISM

[Dom10]  Ignat Domanov.  On in-

variant subspaces of matrices: a new proof of a theorem of Halmos. Linear Algebra and its Applications, 433 (11–12):2255–2256, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Domokos:2013:HMB

[Dom13] M. Domokos. Hermitian matrices with a bounded number of eigenvalues. Linear Algebra and its Applications, 439(12):3964–3979, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006083.

Dato-on:2009:PDM

[DoMP09] Jolle E. Dato-on, Dennis I. Merino, and Agnes T. Paras. The ϕ_{J} polar decomposition of matrices with rank 2. Linear Algebra and its Applications, 430(2–3):756–761, January 15, 2009. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See related work [MPP10].

Dorado:2010:TEP

[Dor10] Ivon Dorado. Three-equipped posets and their representations and corepresentations (inseparable case). Linear Algebra and its Applications, 433(11–12):1827–1850, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drnovsek:2010:SRP

[DP10] Roman Drnovšek and Aljoša Peperko. On the spectral radius of positive operators on Banach sequence spaces. Linear Algebra and its Applications, 433(1):241–247, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dieci:2012:HMD

[DP12a] Luca Dieci and Alessandro Pugliese. Hermitian matrices depending on three parameters: Coalescing eigenvalues. Linear Algebra and its Applications, 436(11): 4120–4142, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000572.

Duarte:2012:RSD

[DP12b] Pedro Duarte and Telmo Peixe. Rank of stably dissipative graphs. Linear Algebra and its Applications, 437 (10):2573–2586, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004508.

Dossal:2010:NEC

[DPF10] Charles Dossal, Gabriel Peyré, and Jalal Fadili. A numerical exploration of compressed sampling recovery. Linear Algebra and its Applications, 432(7):1663–1679, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:2013:TNT

[DQW13] Weiyang Ding, Liqun Qi, and Yimin Wei. M-tensors and nonsingular M-tensors. Linear Algebra and its Applications, 439(10):3264–3278, November 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005570.

Dragomir:2012:HHT

[Dra12a] S. S. Dragomir. Hermite– Hadamard’s type inequalities for convex functions of selfadjoint operators in Hilbert spaces. Linear Algebra and its Applications, 436(5):1503–1515, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006574.

Drazin:2012:COG

[Dra12b]  Michael P. Drazin.  A

class of outer generalized inverses. Linear Algebra and its Applications, 436(7): 1909–1923, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006367.

Drazin:2014:GIU

[Dra14] Michael P. Drazin. Generalized inverses: Uniqueness proofs and three new classes. Linear Algebra and its Applications, 449(??):402–416, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001013.

Drnovsek:2013:MC

[Drn13a] Roman Drnovsek. On the S-matrix conjecture. Linear Algebra and its Applications, 439(11):3555–3560, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005673.

Drnovsek:2013:SSN

[Drn13b] Roman Drnovsek. The spread of the spectrum of a nonnegative matrix with a zero diagonal element. Linear Algebra and its Applications, 439 (8):2381–2387, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004539.

Drury:2011:CCM

[Dru11] S. W. Drury. A counterexample to a conjecture of Matsaev. Linear Algebra and its Applications, 435(2):323–329, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2012:QBK

[Dru12a] S. W. Drury. On a question of Bhatia and Kittaneh. Linear Algebra and its Applications, 437(7):1955–1960, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003278.

Drury:2012:SCB

[Dru12b] S. W. Drury. Solution of the conjecture of Brualdi and Li. Linear Algebra and its Applications, 436(9): 3392–3399, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Drury:2013:FDI

[Dru13] S. W. Drury. Fischer determinantal inequalities and Higham’s Conjecture. Linear Algebra and its Applications, 439(10):3129–3133, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005454.

Drury:2014:PSI

[Dru14] S. W. Drury. Positive semidefiniteness of a 3 × 3 matrix related to partitioning. Linear Algebra and its Applications, 446(??):369–376, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000172.

Draisma:2010:SLT

[DS10] Jan Draisma and Ron Shaw. Singular lines of trilinear forms. Linear Algebra and its Applications, 433(3):690– 697, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Datta:2011:SAQ

[DS11] Biswa Nath Datta and Vadim Sokolov. A solution of the affine quadratic inverse eigenvalue problem. Linear Algebra and its Applications, 434(7):

1745–1760, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deSa:2012:ISS

[dS12a] Eduardo Marques de Sá. The inertia sets of symmetric tridiagonal sign-patterns. Linear Algebra and its Applications, 436(6):1754–1762, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004654.

Deaett:2012:LAM

[DS12b] Louis Deaett and Venkatesh Srinivasan. Linear algebraic methods in communication complexity. Linear Algebra and its Applications, 436(12): 4459–4472, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005209.

Dodig:2012:ESC

[DS12c] Marija Dodig and Marko Stosić. Explicit solution of the Carlson problem. Linear Algebra and its Applications, 436(1):190–201, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004952.

Damm:2013:LLS

[DS13] Tobias Damm and Dominik Stahl. Linear least squares problems with additional constraints and an application to scattered data approximation. Linear Algebra and its Applications, 439(4): 933–943, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006398.

Dodig:2014:RDP

[DS14] Marija Dodig and Marko Stosić. The rank distance problem for pairs of matrices and a completion of quasiregular matrix pencils. Linear Algebra and its Applications, 457(??):313–347, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003206.

dosSantos:2014:ISU

[dSC14] O. J. N. T. N. dos Santos and E. L. Monte Carmelo. Invariant sets under linear operator and covering codes over modules. Linear Algebra and its Applications, 444 (??):42–52, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007726.

Pazzis:2010:ISS

[dSP10a] Clément de Seguins Pazzis. Invariance of simultaneous similarity and equivalence of matrices under extension of the ground field. Linear Algebra and its Applications, 433(3):618–624, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2010:DML

[dSP10b] Clément de Seguins Pazzis. On decomposing any matrix as a linear combination of three idempotents. Linear Algebra and its Applications, 433(4):843–855, October 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2010:LCT

[dSP10c] Clément de Seguins Pazzis. On linear combinations of two idempotent matrices over an arbitrary field. Linear Algebra and its Applications, 433(3): 625–636, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2010:SIM

[dSP10d] Clément de Seguins Pazzis. On sums of idempotent matrices over a field of positive characteristic. Linear Algebra and its Applications, 433 (4):856–866, October 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2010:SLP

[dSP10e] Clément de Seguins Pazzis. The singular linear preservers of non-singular matrices. Linear Algebra and its Applications, 433(2):483–490, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2011:MGR

[dSP11a] Clément de Seguins Pazzis. On the matrices of given rank in a large subspace. Linear Algebra and its Applications, 435(1):147–151, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2011:WEL

[dSP11b] Clément de Seguins Pazzis. To what extent is a large space of matrices not closed under product? Linear Algebra and its Applications, 435(11): 2708–2721, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pazzis:2012:LAS

[dSP12a] Clément de Seguins Pazzis. Large affine spaces of matrices with rank bounded below. Linear Algebra and its Applications, 437(2):499–518, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002066.

Pazzis:2012:LPN

[dSP12b] Clément de Seguins Pazzis. The linear preservers of nonsingularity in a large space of matrices. Linear Algebra and its Applications, 436 (9):3507–3530, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008639.

Pazzis:2012:STT

[dSP12c] Clément de Seguins Pazzis. Sums of two triangularizable quadratic matrices over an arbitrary field. Linear Algebra and its Applications, 436 (9):3293–3302, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007695.

Pazzis:2012:WDL

[dSP12d] Clément de Seguins Pazzis. When does a linear map belong to at least one orthogonal or symplectic group? Linear Algebra and its Applications, 436(5):1385–1405, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006197.

Pazzis:2014:SMF

[dSP14] Clément de Seguins Pazzis. Spaces of matrices with few eigenvalues. Linear Algebra and its Applications, 449 (??):210–311, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000822.

deSnoo:2012:SCC

[dSW12] Henk de Snoo and Harald Woracek. Sums, couplings, and completions of almost Pontryagin spaces. Linear Algebra and its Applications, 437(2):559–580, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002030.

Dimitriou:2010:ETD

[DT10] V. A. Dimitriou and N. Tsantas. Evolution of a time dependent Markov model for training and recruitment decisions in manpower planning. Linear Algebra and its Applications, 433(11–12):1950– 1972, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Druinsky:2011:FMT

[DT11] Alex Druinsky and Sivan Toledo. Factoring matrices with a tree-structured sparsity pattern. Linear Algebra and its Applications, 435 (5):1099–1110, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Diaz–Toca:2011:PBN

[DTL11] Gema M. Diaz-Toca and Henri Lombardi. A polynomial bound on the number of comaximal localizations needed in order to make free a projective module. Linear Algebra and its Applications, 435(2):354–360, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Defez:2011:IBH

[DTS11] Emilio Defez, Michael M. Tung, and Jorge Sastre. Improvement on the bound of Hermite matrix polynomials. Linear Algebra and its Applications, 434(8):1910–1919, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeMarchi:2014:CMV

[DU14] Stefano De Marchi and Konstantin Usevich. On certain multivariate Vandermonde determinants whose variables separate. Linear Algebra and its Applications, 449(??):17–27, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Dubsky:2014:CSW

[Dub14] Brendan Dubsky. Classification of simple weight modules with finite-dimensional weight spaces over the Schrödinger algebra. Linear Algebra and its Applications, 443(??): 204–214, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007131.

Duggal:2011:QES

[Dug11] B. P. Duggal. Quantum effects, sequential independence and majorization. Linear Algebra and its Applications, 435(12):3014–3023, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2012:TPI

[Dug12] B. P. Duggal. Tensor product of n-isometries. Linear Algebra and its Applications, 437(1):307–318, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001619.

Dukes:2012:RDD

[Duk12] Peter Dukes. Rational decomposition of dense hypergraphs and some related eigenvalue estimates. Linear Algebra and its Applications, 436(9):3736–3746, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000626.

See corrigendum [Duk15].

Dukes:2015:CSD

[Duk15] Peter J. Dukes. Corrigendum to “Rational decomposition of dense hypergraphs and some related eigenvalue estimates” [Linear Algebra Appl. 436 (9) (2012) 3736–3746]. Linear Algebra and its Applications, 467(??):267–269, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514007587.

See [Duk12].

Dumas:2013:JSR

[Dum13a] Philippe Dumas. Joint spectral radius, dilation equations, and asymptotic behavior of radix-rational sequences. Linear Algebra and its Applications, 438(5): 2107–2126, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007239.

Dumitrescu:2013:STL

[Dum13b] Bogdan Dumitrescu. Sparse total least squares: Analysis and greedy algorithms. Linear Algebra and its Applications, 438(6):2661–2674, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007628.

Dureisseix:2012:GFF

[Dur12] David Dureisseix. Generalized fraction-free LU factorization for singular systems with kernel extraction. Linear Algebra and its Applications, 436(1):27–40, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004617.

Delvaux:2010:SPG

[DV10] Steven Delvaux and Marc Van Barel. Structures preserved by generalized inversion and Schur complementation. Linear Algebra and its Applications, 432(2–3):817–836, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeSchepper:2014:GDW

[DV14] Anneleen De Schepper and Hendrik Van Maldeghem. Graphs, defined by Weyl distance or incidence, that determine a vector space. Linear Algebra and its Applications, 449(??):435–464, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000810.

Dalfo:2011:PAD

[DvDF11] C. Dalfó, E. R. van Dam, and M. A. Fiol. On perturbations of almost distanceregular graphs. Linear Algebra and its Applications, 435 (10):2626–2638, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dette:2010:MMU

[DW10] Holger Dette and Jens Wagener. Matrix measures on the unit circle, moment spaces, orthogonal polynomials and the Geronimus relations. Linear Algebra and its Applications, 432(7):1609– 1626, March 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2011:RDI

[DW11] Chunyuan Deng and Yimin Wei. Representations for the Drazin inverse of 2 × 2 blockoperator matrix with singular Schur complement. Linear Algebra and its Applications, 435(11):2766–2783, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2012:CMT

[DW12a] Yiqiu Du and Yu Wang. kcommuting maps on triangular algebras. Linear Algebra and its Applications, 436(5):1367–1375, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006070.

Du:2012:LDG

[DW12b] Yiqiu Du and Yu Wang. Lie derivations of generalized matrix algebras. Linear Algebra and its Applications, 437 (11):2719–2726, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200448X.

Du:2013:BGM

[DW13] Yiqiu Du and Yu Wang. Biderivations of generalized matrix algebras. Linear Algebra and its Applications, 438(11):4483–4499, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001353.

Du:2014:JHU

[DW14] Yiqiu Du and Yu Wang. Jordan homomorphisms of upper triangular matrix rings over a prime ring. Linear Algebra and its Applications, 458(??):197–206, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003620.

Du:2014:CSH

[DWW14] Yiqiu Du, Yao Wang, and Yu Wang. Corrigendum to “Jordan homomorphisms of upper triangular matrix rings” [Linear Algebra Appl. 439(12) (2013) 4063–4069]. Linear Algebra and its Applications, 452(??):345–350, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001888.

See [WW13b].

Deng:2012:DRO

[DWXS12] Chunyuan Deng, Yimin Wei, Qingxiang Xu, and Chuanning Song. On disjoint range operators in a Hilbert space. Linear Algebra and its Applications, 437(9):2366– 2385, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004326.

Du:2013:CSP

[DX13] Fapeng Du and Yifeng Xue. The characterizations of the stable perturbation of a closed operator by a linear operator in Banach spaces. Linear Algebra and its Applications, 438(5):2046–2053, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200746X.

Das:2012:CBK

[DXG12] Kinkar Ch. Das, Kexiang Xu, and Ivan Gutman. Comparison between Kirchhoff index and the Laplacian-energy-like invariant. Linear Algebra and its Applications, 436 (9):3661–3671, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000481.

Das:2013:SPL

[DXL13] Kinkar Ch. Das, Kexiang Xu, and Muhuo Liu. On sum of powers of the Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 439 (11):3561–3575, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006046.

Dou:2014:SCG

[DYW14] Yan Dou, Ai-Li Yang, and Yu-Jiang Wu. On semiconvergence of generalized skew-Hermitian triangular splitting iteration methods for singular saddle-point problems. Linear Algebra and its Applications, 459(??): 493–510, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004637.

Du:2011:EIT

[DZ11] Zhibin Du and Bo Zhou. The Estrada index of trees. Linear Algebra and its Applications, 435(10):2462–2467, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Druskin:2012:CKS

[DZ12a] Vladimir Druskin and Mikhail Zaslavsky. On convergence of Krylov subspace approximations of time-invariant self-adjoint dynamical systems. Linear Algebra and its Applications, 436(10): 3883–3903, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100173X.

Du:2012:EIU

[DZ12b] Zhibin Du and Bo Zhou. The Estrada index of unicyclic graphs. Linear Algebra and its Applications, 436 (9):3149–3159, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007178.

Du:2012:UBS

[DZ12c] Zhibin Du and Bo Zhou. Upper bounds for the sum of Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 436(9): 3672–3683, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000559.

Duan:2013:SBS

[DZ13] Xing Duan and Bo Zhou. Sharp bounds on the spectral radius of a nonnegative matrix. Linear Algebra and its Applications, 439 (10):2961–2970, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005405.

Du:2012:MEI

[DZX12] Zhibin Du, Bo Zhou, and Rundan Xing. On maximum Estrada indices of graphs with given parameters. Linear Algebra and its Applications, 436(9):3767–3772, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008706.

Estrada:2014:MWE

[EdlP14] Ernesto Estrada and José A. de la Peña. Maximum walk entropy implies walk regularity. Linear Algebra and its Applications, 458(??): 542–547, October 1, 2014.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003991.

Estrada:2014:WEG

[EdlPH14] Ernesto Estrada, José A. de la Peña, and Naomichi Hatano. Walk entropies in graphs. Linear Algebra and its Applications, 443(??): 235–244, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007064.

Ekstrand:2013:PSZ

[EEH^{+}13] Jason Ekstrand, Craig Erickson, H. Tracy Hall, Diana Hay, Leslie Hogben, Ryan Johnson, Nicole Kingsley, Steven Osborne, Travis Peters, Jolie Roat, Arianne Ross, Darren D. Row, Nathan Warnberg, and Michael Young. Positive semidefinite zero forcing. Linear Algebra and its Applications, 439(7):1862–1874, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003625.

Ejov:2009:NGR

[EFN09] Vladimir Ejov, Shmuel Friedland, and Giang T. Nguyen. A note on the graph’s resolvent and the multifilar structure. Linear Algebra and its Applications, 431(8):1367– 1379, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See erratum [EFN10].

Ejov:2010:ENG

[EFN10] Vladimir Ejov, Shmuel Friedland, and Giang T. Nguyen. Erratum to “A note on the graph’s resolvent and the multifilar structure”. Linear Algebra and its Applications, 432 (9):2456, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [EFN09].

Ernst:2014:P

[EGLR14] Oliver G. Ernst, Chun-Hua Guo, Jörg Liesen, and Leiba Rodman. Preface. Linear Algebra and its Applications, 456(??):1–2, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000366.

Euler:2012:SCC

[EGR12] Reinhardt Euler, Luis H. Gallardo, and Olivier Rahavandrainy. Sufficient conditions for a conjecture of Ryser about Hadamard Circulant matrices. Linear Algebra and its Applications, 437 (12):2877–2886, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005630.

Edholm:2012:VES

[EHH^{+}12] Christina J. Edholm, Leslie Hogben, My Huynh, Joshua LaGrange, and Darren D. Row. Vertex and edge spread of zero forcing number, maximum nullity, and minimum rank of a graph. Linear Algebra and its Applications, 436(12):4352–4372, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379510005392.

Einstein:2011:RFV

[EJLS11] Eduard Einstein, Charles R. Johnson, Brian Lins, and Ilya Spitkovsky. The ratio field of values. Linear Algebra and its Applications, 434(4):1119– 1136, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eastman:2014:CMP

[EKSV14] B. Eastman, I.-J. Kim, B. L. Shader, and K. N. Vander Meulen. Companion matrix patterns. Linear Algebra and its Applications, 463 (??):255–272, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/

science/article/pii/S0024379514005953.

See corrigendum [EKSV18].

Eastman:2018:CSC

[EKSV18] B. Eastman, I.-J. Kim, B. L. Shader, and K. N. Vander Meulen. Corrigendum to “Companion matrix patterns” [Linear Algebra Appl. 463 (2014) 255–272]. Linear Algebra and its Applications, 538(??):225–227, February 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379517305967.

See [EKSV14].

Elouafi:2011:ELT

[Elo11] Mohamed Elouafi. An eigenvalue localization theorem for pentadiagonal symmetric Toeplitz matrices. Linear Algebra and its Applications, 435(11):2986–2998, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2011:BMP

[EM11] L. Elsner and V. Monov. The bialternate matrix product revisited. Linear Algebra and its Applications, 434(4):1058– 1066, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Erway:2012:LMB

[EM12] Jennifer B. Erway and Roummel F. Marcia. Limitedmemory BFGS systems with diagonal updates. Linear Algebra and its Applications, 437(1):333–344, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001358.

Ebanks:2011:HTA

[EN11] Bruce Ebanks and C. T. Ng. Homogeneous tri-additive forms and derivations. Linear Algebra and its Applications, 435(11):2731–2755, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ensenbach:2010:DDP

[Ens10] Marc Ensenbach. Determinantal divisors of products of matrices over Dedekind domains. Linear Algebra and its Applications, 432(11): 2739–2744, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ehrhardt:2013:RMI

[ER13] Torsten Ehrhardt and Karla Rost. Resultant matrices and inversion of Bezoutians. Linear Algebra and its Applications, 439(3):621–639, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006453.

Eremita:2013:FID

[Ere13] Daniel Eremita. Functional identities of degree 2 in triangular rings. Linear Algebra and its Applications, 438(1):584–597, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005873.

Ernst:2013:UAF

[Ern13] Thomas Ernst. An umbral approach to find q-analogues of matrix formulas. Linear Algebra and its Applications, 439(4):1167–1182, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002139.

Erdmann:2011:CPS

[ES11] Karin Erdmann and Sibylle Schroll. Chebyshev polynomials on symmetric matrices. Linear Algebra and its Applications, 434(12):2475–2496, June 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ehrhardt:2013:BAQ

[ES13] Torsten Ehrhardt and Bernd Silbermann. Banach algebras of quasi-triangular operators are spectrally regular. Linear Algebra and its Applications, 439(3):577–583, August 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006192.

Ellard:2014:CNR

[ES14] Richard Ellard and Helena Smigoc. Constructing new realisable lists from old in the NIEP. Linear Algebra and its Applications, 440(??): 218–232, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006708.

Estrada:2012:CDG

[Est12a]  Ernesto Estrada.  The

communicability distance in graphs. Linear Algebra and its Applications, 436(11): 4317–4328, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200081X.

Estrada:2012:PLM

[Est12b] Ernesto Estrada. Path Laplacian matrices: Introduction and application to the analysis of consensus in networks. Linear Algebra and its Applications, 436(9):3373– 3391, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007750.

Ellens:2011:EGR

[ESV^{+}11] W. Ellens, F. M. Spieksma, P. Van Mieghem, A. Jamakovic, and R. E. Kooij. Effective graph resistance. Linear Algebra and its Applications, 435(10):2491–2506, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Esslamzadeh:2013:SQO

[ET13] G. H. Esslamzadeh and F. Taleghani. Structure of quasi operator systems. Linear Algebra and its Applications, 438(3):1372–1392, February 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006817.

Ellers:2011:CSC

[EV11] Erich W. Ellers and Oliver Villa. Commutators of symmetries in characteristic 2. Linear Algebra and its Applications, 434(6):1588–1593, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2010:MAP

[EvdD10] L. Elsner and P. van den Driessche. Max-algebra and pairwise comparison matrices, II. Linear Algebra and its Applications, 432(4):927– 935, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eu:2012:HDS

[EWY12] Sen-Peng Eu, Tsai-Lien Wong, and Pei-Lan Yen. Hankel determinants of sums of consecutive weighted Schröder numbers. Linear Algebra and its Applications, 437(9): 2285–2299, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003862.

Estevez:2013:DRE

[EY13] Daniel Estévez and Dmitry V. Yakubovich. Decay rate estimations for linear quadratic optimal regulators. Linear Algebra and its Applications, 439(11):3332–3358, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005442.

Fanai:2010:EPT

[Fan10a] Hamid-Reza Fanaï. Existence of partial transversals. Linear Algebra and its Applications, 432(10):2608–2614, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2010:NIE

[Fan10b] Maozhong Fang. A note on the inverse eigenvalue problem for symmetric doubly stochastic matrices. Linear Algebra and its Applications, 432(11):2925–2927, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See original paper [HP04] and comments [Fan10b, XLG^{+}13].

Fang:2012:MVS

[Fan12] Li Fang. Maximal vectors in some Hilbert spaces. Linear Algebra and its Applications, 437(4):1102–1108, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002893.

Farenick:2011:ACU

[Far11a] Douglas Farenick. Arveson’s criterion for unitary similarity. Linear Algebra and its Applications, 435(4):769– 777, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farid:2011:NMD

[Far11b] F. O. Farid. Notes on matrices with diagonally dominant properties. Linear Algebra and its Applications, 435 (11):2793–2812, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fong:2010:AUI

[FCL10] Kin-Sio Fong, Che-Man Cheng, and Io-Kei Lok. Another unitarily invariant norm attaining the minimum norm bound for commutators. Linear Algebra and its Applications, 433(11–12):1793–1797, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2010:PMP

[FdC10] Rosário Fernandes and Henrique F. da Cruz. Pairs of matrices that preserve the value of a generalized matrix function on the set of the upper triangular matrices. Linear Algebra and its Applications, 433(7):1336–1347, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2012:MAJ

[FdC12] Rosário Fernandes and Henrique F. da Cruz. The Multilinear Algebra of José Dias da Silva and the Portuguese school of mathematics. Linear Algebra and its Applications, 436(6):1545–1561, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511002503.

Fernandes:2014:TRP

[FdC14a] Rosário Fernandes and Henrique F. da Cruz. On the term rank partition. Linear Algebra and its Applications, 458(??):134–148, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003693.

Fernandes:2014:SPV

[FdC14b] Rosário Fernandes and Henrique F. da Cruz. Sets of Parter vertices which are Parter sets. Linear Algebra and its Applications, 448 (??):37–54, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400069X.

Fernandes:2010:DCT

[FdCR10] Rosário Fernandes, Henrique F. da Cruz, and Fátima Rodrigues. Decomposable λcritical tensors. Linear Algebra and its Applications, 433 (1):297–317, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2013:MAA

[FDS13] Yuan Feng, Theo S. H. Driessen, and Georg Still. A matrix approach to associated consistency of the Shapley value for games in generalized characteristic function form. Linear Algebra and its Applications, 438(11): 4279–4295, June 1, 2013.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000876.

Fleischhack:2010:APH

[FF10] Christian Fleischhack and Shmuel Friedland. Asymptotic positivity of Hurwitz product traces: Two proofs. Linear Algebra and its Applications, 432(6):1363–1383, March 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2012:BLE

[FF12] Shaun Fallat and Yi-Zheng Fan. Bipartiteness and the least eigenvalue of signless Laplacian of graphs. Linear Algebra and its Applications, 436(9):3254–3267, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007580.

Farenick:2011:CUSb

[FFG^{+}11] Douglas Farenick, Vyacheslav Futorny, Tatiana G. Gerasimova, Vladimir V. Sergeichuk, and Nadya Shvai. A criterion for unitary similarity of upper triangular matrices in general position. Linear Algebra and its Applications, 435(6):1356–1369, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farber:2014:EET

[FFJM14] Miriam Farber, Mitchell Faulk, Charles R. Johnson, and Evan Marzion. Equal entries in totally positive matrices. Linear Algebra and its Applications, 454 (??):91–106, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002213.

Foniok:2011:CCM

[FFK11] Jan Foniok, Komei Fukuda, and Lorenz Klaus. Combinatorial characterizations of K-matrices. Linear Algebra and its Applications, 434(1): 68–80, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2011:GAN

[FFS11a] Min Fang, Gui-Liang Feng, and Jia-Yu Shao. Generalizations and applications of the nowhere zero linear mappings in network coding. Linear Algebra and its Applications, 434(3):841–848, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Felipe:2011:RLS

[FFS11b] Raúl Felipe and Raúl Felipe-Sosa. Relaxed linear spaces and a generalization of the Cayley–Dickson process. Linear Algebra and its Applications, 434(2):424–442, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2013:PGS

[FG13a] Xiaoxia Fan and Chris Godsil. Pretty good state transfer on double stars. Linear Algebra and its Applications, 438(5):2346–2358, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007161.

Friedland:2013:SSF

[FG13b] S. Friedland and S. Gaubert. Submodular spectral functions of principal submatrices of a Hermitian matrix, extensions and applications. Linear Algebra and its Applications, 438(10):3872–3884, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007646.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Fulman:2013:NRS

[FG13c] Jason Fulman and Robert Guralnick. The number of regular semisimple conjugacy classes in the finite classical groups. Linear Algebra and its Applications, 439 (2):488–503, July 15, 2013.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002267.

Fiol:2010:SPS

[FGG10] M. A. Fiol, S. Gago, and E. Garriga. A simple proof of the spectral excess theorem for distance-regular graphs. Linear Algebra and its Applications, 432(9):2418–2422, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2013:PFT

[FGH13] S. Friedland, S. Gaubert, and L. Han. Perron–Frobenius theorem for nonnegative multilinear forms and extensions. Linear Algebra and its Applications, 438(2):738– 749, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001765.

Farhi:2011:PLC

[FGQ11] N. Farhi, M. Goursat, and J.- P. Quadrat. Piecewise linear concave dynamical systems appearing in the microscopic traffic modeling. Linear Algebra and its Applications, 435(7):1711–1735, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Forrester:2013:NPL

[FGR13] W. Phillip Forrester, Edgar E. Galindo, and Adolfo J. Rumbos. Nonlinear perturbations of linear equations in R^{n}. Linear Algebra and its Applications, 439(10):3100– 3114, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005491.

Farenick:2011:CUSa

[FGS11] Douglas Farenick, Tatiana G. Gerasimova, and Nadya Shvai. A complete unitary similarity invariant for unicellular matrices. Linear Algebra and its Applications, 435(2): 409–419, July 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fourie:2013:ROP

[FGvRR13] J. H. Fourie, G. J. Groenewald, D. B. Janse van Rensburg, and A. C. M. Ran. Rank one perturbations of Hpositive real matrices. Linear Algebra and its Applications, 439(3):653–674, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002723.

Fiedler:2010:SIP

[FH10a] Miroslav Fiedler and Frank J. Hall. Some inheritance properties for complementary basic matrices. Linear Algebra and its Applications, 433 (11–12):2060–2069, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2010:EDS

[FH10b] P. A. Fuhrmann and U. Helmke. On the elementary divisors of the Sylvester and Lyapunov maps. Linear Algebra and its Applications, 432(10): 2572–2588, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2010:TPM

[FH10c] P. A. Fuhrmann and U. Helmke. Tensored polynomial models. Linear Algebra and its Applications, 432(2–3):678– 721, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2012:M

[FH12a] Miroslav Fiedler and Frank J. Hall. G-matrices. Linear Algebra and its Applications, 436(3):731–741, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005453.

Fiedler:2012:NPG

[FH12b] Miroslav Fiedler and Frank J. Hall. A note on permanents and generalized complementary basic matrices. Linear Algebra and its Applications, 436(9):3553–3561, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008664.

Frommer:2012:VEB

[FH12c] Andreas Frommer and Behnam Hashemi. Verified error bounds for solutions of Sylvester matrix equations. Linear Algebra and its Applications, 436(2):405–420, January 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510006300.

Fiedler:2013:SGT

[FH13] Miroslav Fiedler and Frank J. Hall. Some graph theoretic properties of generalized complementary basic matrices. Linear Algebra and its Applications, 438(8): 3365–3374, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000153.

Fiedler:2014:MAC

[FH14] Miroslav Fiedler and Frank J. Hall. Max algebraic complementary basic matrices. Linear Algebra and its Applications, 457(??):287–292, September 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003103.

Forstall:2011:PEI

[FHL^{+}11] Virginia Forstall, Aaron Herman, Chi-Kwong Li, Nung-Sing Sze, and Vincent Yannello. Preservers of eigenvalue inclusion sets of matrix products. Linear Algebra and its Applications, 434(1):285– 293, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2013:GDR

[FHM13] Miroslav Fiedler, Frank J. Hall, and Rachid Marsli. Geršgorin discs revisited. Linear Algebra and its Applications, 438(1):598–603, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005861.

Fritscher:2011:SLE

[FHRT11] Eliseu Fritscher, Carlos Hoppen, Israel Rocha, and Vilmar Trevisan. On the sum of the Laplacian eigenvalues of a tree. Linear Algebra and its Applications, 435(2):371–399, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritscher:2014:CTL

[FHRT14] Eliseu Fritscher, Carlos Hoppen, Israel Rocha, and Vilmar Trevisan. Characterizing trees with large Laplacian energy. Linear Algebra and its Applications, 442 (??):20–49, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000906.

Futorny:2011:CFN

[FHS11] Vyacheslav Futorny, Roger A. Horn, and Vladimir V. Sergeichuk. A canonical form for nonderogatory matrices under unitary similarity. Linear Algebra and its Applications, 435(4):830–841, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2014:DAS

[FHS14a] Miroslav Fiedler, Frank J. Hall, and Mikhail Stroev. Dense alternating sign matrices and extensions. Linear Algebra and its Applications, 444(??):219–226, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007799.

Frommer:2014:CEI

[FHS14b] Andreas Frommer, Behnam Hashemi, and Thomas Sablik. Computing enclosures for the inverse square root and the sign function of a matrix. Linear Algebra and its Applications, 456(??): 199–213, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007969.

Fidaleo:2010:EET

[Fid10] Francesco Fidaleo. The entangled ergodic theorem in the almost periodic case. Linear Algebra and its Applications, 432(2–3):526–535, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2010:NHC

[Fie10] Miroslav Fiedler. Notes on Hilbert and Cauchy matrices. Linear Algebra and its Applications, 432(1):351–356, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2011:DMM

[Fie11a] Miroslav Fiedler. Dominant matrices and max algebra. Linear Algebra and its Applications, 434(4):1189–1194, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2011:MGG

[Fie11b] Miroslav Fiedler. Matrices and Graphs in Geometry, volume 139 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-46193-6. viii + 197 pp. LCCN QA447 .F45

2011. URL http://assets.
cambridge.org/97805214/
61931/cover/9780521461931.
jpg;   http://catdir.loc.
gov/catdir/enhancements/
fy1101/2010046601-b.html;
 http://catdir.loc.gov/
catdir/enhancements/fy1101/
2010046601-d.html;   http:
//catdir.loc.gov/catdir/
enhancements/fy1101/2010046601-
t.html.

Fiedler:2013:SOV

[Fie13] Miroslav Fiedler. Some observations on variance matrices. Linear Algebra and its Applications, 439(2):504–509, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002516.

Fiol:2012:PDR

[Fio12]  M. A. Fiol.  Pseudo-

distance-regularized graphs are distance-regular or distancebiregular. Linear Algebra and its Applications, 437 (12):2973–2977, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Fischer:2014:SSA

[Fis14] Thomas M. Fischer. On the stability of some algorithms for computing the action of the matrix exponential. Linear Algebra and its Applications, 443(??):1–20, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007003.

Ferguson:2010:GOI

[FJ10a] Timothy Ferguson and Charles Johnson. Gangster operators and invincibility of positive semidefinite matrices. Linear Algebra and its Applications, 433(11–12):2096–2110, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2010:OIS

[FJ10b] Susana Furtado and Charles R. Johnson. Order invariant spectral properties for several matrices. Linear Algebra and its Applications, 432(8):1950– 1960, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2011:TNM

[FJ11] Shaun M. Fallat and Charles R. Johnson. Totally Nonnegative Matrices. Princeton Series in Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-12157-5 (hardcover), 1-4008-3901-7 (e-book). xv + 248 pp. LCCN QA188 .F35 2011. URL http://

www.jstor.org/stable/10.
2307/j.ctt7scff.

Furtado:2014:SCA

[FJ14] Susana Furtado and Charles R. Johnson. On the similarity classes among products of m nonsingular matrices in various orders. Linear Algebra and its Applications, 450 (??):217–242, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400161X.

Furtado:2014:SMP

[FJMP14] Susana Furtado, C. R. Johnson, C. Marijuán, and M. Pisonero. Submatrix monotonicity of the Perron root, II. Linear Algebra and its Applications, 458(??): 679–688, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004169.

Fendler:2013:DFT

[FK13] Gero Fendler and Norbert Kaiblinger. Discrete Fourier transform of prime order: Eigenvectors with small support. Linear Algebra and its Applications, 438(1):288– 302, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005939.

Friedland:2013:PSI

[FKLT13] Shmuel Friedland, Tamara Kolda, Lek-Heng Lim, and Eugene Tyrtyshnikov. Preface to the special issue on tensors and multilinear algebra. Linear Algebra and its Applications, 438(2):635– 638, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006763.

Fritzsche:2013:TMS

[FKM13] Bernd Fritzsche, Bernd Kirstein, and Conrad Mädler. Transformations of matricial α-Stieltjes non-negative definite sequences. Linear Algebra and its Applications, 439 (12):3893–3933, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006216.

Fritzsche:2011:IPP

[FKR11a] Bernd Fritzsche, Bernd Kirstein, and Uwe Raabe. On an interpolation problem for J-Potapov functions. Linear Algebra and its Applications, 434(3):741–784, February 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2011:SWS

[FKR11b] Bernd Fritzsche, Bernd Kirstein, and Uwe Raabe. On the sequences of the Weyl semiradii associated with matricial Carathéodory and Schur sequences in both nondegenerate and degenerate cases. Linear Algebra and its Applications, 435(4):778–803, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2012:WMB

[FKR12] Bernd Fritzsche, Bernd Kirstein, and Uwe Raabe. On the Weyl matrix balls associated with J–Potapov sequences in both nondegenerate and degenerate cases. Linear Algebra and its Applications, 436(5):1028–1060, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004757.

Farid:2013:MAS

[FKW13] F. O. Farid, Israr Ali Khan, and Qing-Wen Wang. On matrices over an arbitrary semiring and their generalized inverses. Linear Algebra and its Applications, 439(7):2085–2105, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300390X.

Furuichi:2010:MTI

[FL10] Shigeru Furuichi and Minghua Lin. A matrix trace inequality and its application. Linear Algebra and its Applications, 433(7):1324–1328, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2012:MRG

[FL12] Shmuel Friedland and Raphael Loewy. On the minimum rank of a graph over finite fields. Linear Algebra and its Applications, 436(6): 1710–1720, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004915.

Fong:2011:NNC

[FLC11] Kin-Sio Fong, Io-Kei Lok, and Che-Man Cheng. A note on the norm of the commutator and the norm of XY − Y X^{T}. Linear Algebra and its Applications, 435(6):1193–1201, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2012:ENM

[FLH12] Xin-Lei Feng, Zhongshan Li, and Ting-Zhu Huang. Is every nonsingular matrix diagonally equivalent to a matrix with all distinct eigenvalues? Linear Algebra and its Applications, 436(1):120–125, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004812.

Fujii:2010:DCM

[FLS10] Masatoshi Fujii, Eun-Young Lee, and Yuki Seo. A difference counterpart to a matrix Hölder inequality. Linear Algebra and its Applications, 432(10):2565–2571, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:PPA

[fLyH11] Jiao fen Li and Xi yan Hu. Procrustes problems and associated approximation problems for matrices with kinvolutory symmetries. Linear Algebra and its Applications, 434(3):820–829, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fania:2011:VSS

[FM11a] Maria Lucia Fania and Emilia Mezzetti. Vector spaces of skew-symmetric matrices of constant rank. Linear Algebra and its Applications, 434(12): 2383–2403, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2011:CMS

[FM11b] Miroslav Fiedler and Thomas L. Markham. Combined matrices in special classes of matrices. Linear Algebra and its Applications, 435(8):1945– 1955, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fickus:2012:NER

[FM12] Matthew Fickus and Dustin G. Mixon. Numerically erasurerobust frames. Linear Algebra and its Applications, 437 (6):1394–1407, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003102.

Fiedler:2013:MM

[FM13] Miroslav Fiedler and Thomas L. Markham. More on Gmatrices. Linear Algebra and its Applications, 438(1): 231–241, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006179.

Ferrer:2013:DMC

[FMM13] Josep Ferrer, David Mingueza, and M. Eulalia Montoro. Determinant of a matrix that commutes with a Jordan matrix. Linear Algebra and its Applications, 439(12):3945– 3954, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006447.

Fickus:2014:PRV

[FMNW14] Matthew Fickus, Dustin G. Mixon, Aaron A. Nelson, and Yang Wang. Phase retrieval from very few measurements. Linear Algebra and its Applications, 449(??):475–499, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000780.

Filipiak:2012:MDC

[FMR12] K. Filipiak, A. Markiewicz, and R. Rózański. Maximal determinant over a certain class of matrices and its application to D-optimality of designs. Linear Algebra and its Applications, 436(4): 874–887, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003417.

Fickus:2012:SET

[FMT12] Matthew Fickus, Dustin G. Mixon, and Janet C. Tremain. Steiner equiangular tight frames. Linear Algebra and its Applications, 436(5): 1014–1027, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004769.

Farid:2012:HSS

[FMWW12] F. O. Farid, M. S. Moslehian, Qing-Wen Wang, and Zhong-Cheng Wu. On the Hermitian solutions to a system of adjointable operator equations. Linear Algebra and its Applications, 437(7): 1854–1891, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003643.

Fiedler:2010:SRH

[FN10] Miroslav Fiedler and Vladimir Nikiforov. Spectral radius and Hamiltonicity of graphs. Linear Algebra and its Applications, 432(9):2170–2173, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2011:SBC

[FN11] Ying Wai (Daniel) Fan and James G. Nagy. Synthetic boundary conditions for image deblurring. Linear Algebra and its Applications, 434 (11):2244–2268, June 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2013:SGF

[FNY13] Masatoshi Fujii, Ritsuo Nakamoto, and Keisuke Yonezawa. A satellite of the grand Furuta inequality and its application. Linear Algebra and its Applications, 438(4): 1580–1586, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002734.

Forough:2014:MRI

[For14] M. Forough. Majorization, range inclusion, and factorization for unbounded operators on Banach spaces. Linear Algebra and its Applications, 449(??):60–67, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001001.

Foucart:2014:SRI

[Fou14]  Simon Foucart.  Stabil-

ity and robustness of ℓ_{1}- minimizations with Weibull matrices and redundant dictionaries. Linear Algebra and its Applications, 441 (??):4–21, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007008.

Franchi:2011:IRA

[FP11] Massimo Franchi and Paolo Paruolo. Inversion of regular analytic matrix functions: Local Smith form and subspace duality. Linear Algebra and its Applications, 435 (11):2896–2912, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farnell:2013:FAS

[FP13] Shawn Farnell and Rachel Pries. Families of Artin– Schreier curves with Cartier– Manin matrix of constant rank. Linear Algebra and its Applications, 439(7):2158– 2166, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004114.

Fialowski:2014:MSD

[FP14] Alice Fialowski and Michael Penkava. The moduli space of 4-dimensional nilpotent complex associative algebras. Linear Algebra and its Applications, 457(??):408–427, September 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002948.

Fidalgo:2013:MTM

[FPC13] U. Fidalgo, S. Medina Peralta, and J. Mínguez Ceniceros. Mixed type multiple orthogonal polynomials: Perfectness and interlacing properties of zeros. Linear Algebra and its Applications, 438 (3):1229–1239, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006581.

Franca:2012:CMS

[Fra12] Willian Franca. Commuting maps on some subsets of matrices that are not closed under addition. Linear Algebra and its Applications, 437 (1):388–391, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001620.

Franca:2013:CMR

[Fra13] Willian Franca. Commuting maps on rank-k matrices. Linear Algebra and its Applications, 438(6):2813– 2815, March 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200818X.

Friedland:2011:PDS

[Fri11] Shmuel Friedland. Positive diagonal scaling of a nonnegative tensor to one with prescribed slice sums. Linear Algebra and its Applications, 434(7):1615–1619, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2012:GTR

[Fri12] Shmuel Friedland. On the generic and typical ranks of 3-tensors. Linear Algebra and its Applications, 436(3): 478–497, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004137.

Friedland:2013:TBR

[Fri13] Shmuel Friedland. On tensors of border rank l in C^{m×n×l}. Linear Algebra and its Applications, 438(2): 713–737, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004186.

Futorny:2014:RDM

[FRS14] Vyacheslav Futorny, Tetiana Rybalkina, and Vladimir V. Sergeichuk. Regularizing decompositions for matrix pencils and a topological classification of pairs of linear mappings. Linear Algebra and its Applications, 450 (??):121–137, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001323.

Freitag:2011:NBM

[FS11] M. A. Freitag and A. Spence. A Newton-based method for the calculation of the distance to instability. Linear Algebra and its Applications, 435 (12):3189–3205, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farrugia:2014:CUG

[FS14a] Alexander Farrugia and Irene Sciriha. Controllability of undirected graphs. Linear Algebra and its Applications, 454(??):138–157, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400247X.

Frommer:2014:NCC

[FS14b] Andreas Frommer and Daniel B. Szyld. On necessary conditions for convergence of stationary iterative methods for Hermitian semidefinite linear systems. Linear Algebra and its Applications, 453 (??):192–201, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Feng:2010:CKI

[FT10] Bao Qi Feng and Andrew Tonge. A Cauchy–Khinchin integral inequality. Linear Algebra and its Applications, 433(5):1024–1030, October 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fath-Tabar:2011:NUB

[FTA11] G. H. Fath-Tabar and A. R. Ashrafi. New upper bounds for Estrada index of bipartite graphs. Linear Algebra and its Applications, 435 (10):2607–2611, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fath-Tabar:2010:SLS

[FTDA10] Gholam Hossein Fath-Tabar, Tomislav Došlić, and Ali Reza Ashrafi. On the Szeged and the Laplacian Szeged spectrum of a graph. Linear Algebra and its Applications, 433(3):662–671, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2012:GCH

[FTZ12] L. G. Feng, H. J. Tan, and K. M. Zhao. A generalized Cayley–Hamilton theorem. Linear Algebra and its Applications, 436(7):2440– 2445, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008329.

Fuhrmann:2010:FAS

[Fuh10a] P. A. Fuhrmann. A functional approach to the Stein equation. Linear Algebra and its Applications, 432(12): 3031–3071, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2010:TMI

[Fuh10b] P. A. Fuhrmann. On tangential matrix interpolation. Linear Algebra and its Applications, 433(11–12):2018–2059, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2010:SHP

[Fuj10] Jun Ichi Fujii. Structure of Hiai–Petz parametrized geometry for positive definite matrices. Linear Algebra and its Applications, 432(1):318– 326, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2011:PQM

[Fuj11] Jun Ichi Fujii. Path of quasimeans as a geodesic. Linear Algebra and its Applications, 434(2):542–558, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furer:2010:PCS

[Für10a] Martin Fürer. On the power of combinatorial and spectral invariants. Linear Algebra and its Applications, 432(9):2373– 2380, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2010:PSS

[Fur10b]  Takayuki Furuta.  Posi-

tive semidefinite solutions of the operator equation ∑_{j=1}^{m} A^{n−j}XA^{j−1} = B. Linear Algebra and its Applications, 432(4):949–955, February 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2011:ACI

[Fur11] Takayuki Furuta. Around Choi inequalities for positive linear maps. Linear Algebra and its Applications, 434(1): 14–17, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2012:EIU

[Fur12]  Takayuki Furuta.  Exten-

sions of inequalities for unitarily invariant norms via log majorization. Linear Algebra and its Applications, 436 (9):3463–3468, May 1, 2012.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008342.

Futamura:2012:FDM

[Fut12] Fumiko Futamura. Frame diagonalization of matrices. Linear Algebra and its Applications, 436(9):3201– 3214, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100735X.

First:2013:SWI

[FV13a] Uriya A. First and Uzi Vishne. Stiefel–Whitney invariants for bilinear forms. Linear Algebra and its Applications, 439(7):1905–1917, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003662.

Fornasini:2013:ASS

[FV13b] Ettore Fornasini and Maria Elena Valcher. Asymptotic stability and stabilizability of special classes of discrete-time positive switched systems. Linear Algebra and its Applications, 438(4):1814–1831, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006112.

Francis:2014:SM

[FW14a] Andrew R. Francis and Henry P. Wynn. Subgroup majorization. Linear Algebra and its Applications, 444 (??):53–66, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007805.

Frenkel:2014:VCC

[FW14b] Péter E. Frenkel and Mihály Weiner. On vector configurations that can be realized in the cone of positive matrices. Linear Algebra and its Applications, 459(??):465–474, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004674.

Fan:2013:NNU

[FWW13] Yi-Zheng Fan, Yue Wang, and Yi Wang. A note on the nullity of unicyclic signed graphs. Linear Algebra and its Applications, 438 (3):1193–1200, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006519.

Feng:2010:LSR

[FYI10] Lihua Feng, Guihai Yu, and Aleksandar Ilić. The Laplacian spectral radius for unicyclic graphs with given independence number. Linear Algebra and its Applications, 433(5):934–944, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2011:SIR

[FZ11] Longxiang Fang and Xinsheng Zhang. Slepian’s inequality with respect to majorization. Linear Algebra and its Applications, 434(4):1107–1118, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2013:ESI

[FZ13] Wenying Feng and Guang Zhang. Eigenvalue and spectral intervals for a nonlinear algebraic system. Linear Algebra and its Applications, 439(1):1–20, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001651.

Fan:2011:LEC

[FZW11] Yi-Zheng Fan, Fei-Fei Zhang, and Yi Wang. The least eigenvalue of the complements of trees. Linear Algebra and its Applications, 435(9):2150– 2155, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garloff:2010:RTP

[Gar10] Jürgen Garloff. Review of Totally Positive Matrices by Allan Pinkus. Linear Algebra and its Applications, 433(5): 1052–1053, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garloff:2012:RBN

[Gar12] Jürgen Garloff. Review of Totally Nonnegative Matrices by Shaun M. Fallat and Charles R. Johnson, Princeton University Press, Princeton and Oxford (2011), xv + 248 pp., Princeton Series in Applied Mathematics, ISBN 978-0-691-12157-4 cloth. Linear Algebra and its Applications, 436(9):3790–3792, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008238.

See [FJ11].

Garoni:2013:EME

[Gar13]  Carlo Garoni.  Estimates

for the minimum eigenvalue and the condition number of Hermitian (block) Toeplitz matrices. Linear Algebra and its Applications, 439(3): 707–728, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006234.

Gau:2010:NRR

[Gau10] Hwa-Long Gau. Numerical ranges of reducible companion matrices. Linear Algebra and its Applications, 432(5):1310– 1321, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:2011:CVN

[GB11] Felix Goldberg and Abraham Berman. On the Colin de Verdière number of graphs. Linear Algebra and its Applications, 434(7):1656–1662, April 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Glashoff:2013:MCT

[GB13] Klaus Glashoff and Michael M. Bronstein. Matrix commutators: their asymptotic metric properties and relation to approximate joint diagonalization. Linear Algebra and its Applications, 439 (8):2503–2513, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004643.

Goldberg:2014:ZFS

[GB14] Felix Goldberg and Abraham Berman. Zero forcing for sign patterns. Linear Algebra and its Applications, 447(??):56–67, April 15, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007982.

Greenbaum:2012:CCP

[GC12] Anne Greenbaum and Daeshik Choi. Crouzeix’s conjecture and perturbed Jordan blocks. Linear Algebra and its Applications, 436 (7):2342–2352, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007853.

Guo:2014:LBL

[GCY14] Shu-Guang Guo, Yong-Gao Chen, and Guanglong Yu. A lower bound on the least signless Laplacian eigenvalue of a graph. Linear Algebra and its Applications, 448 (??):217–221, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000330.

Ghazanfari:2011:SGT

[GD11a] Amir G. Ghazanfari and Sever S. Dragomir. Schwarz and Grüss type inequalities for C^{∗}-seminorms and positive linear functionals on Banach *-modules. Linear Algebra and its Applications, 434(4):944–956, February 15, 2011. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Gumedze:2011:PEI

[GD11b] F. N. Gumedze and T. T. Dunne. Parameter estimation and inference in the linear mixed model. Linear Algebra and its Applications, 435(8): 1920–1944, October 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gemignani:2010:MCS

[Gem10] L. Gemignani. Matrix Computations and Semiseparable Matrices. Volume I: Linear Systems, R. Vandebril, M. Van Barel, N. Mastronardi. Linear Algebra and its Applications, 432(1):483–484, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Esnaola:2010:CEB

[GEP10] M. García-Esnaola and J. M. Peña. A comparison of error bounds for linear complementarity problems of Hmatrices. Linear Algebra and its Applications, 433(5):956– 964, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Esnaola:2013:EBL

[GEP13] M. García-Esnaola and J. M. Peña. Error bounds for the linear complementarity problem with a Σ-SDD matrix. Linear Algebra and its Applications, 438(3):1339– 1346, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006830.

Gerasimova:2012:USN

[Ger12] Tatiana G. Gerasimova. Unitary similarity to a normal matrix. Linear Algebra and its Applications, 436 (9):3777–3783, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008275.

Gutman:2014:RE

[GFB14] Ivan Gutman, Boris Furtula, and S. Burcu Bozkurt. On Randić energy. Linear Algebra and its Applications, 442(??):50–57, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003984.

Gong:2010:NGP

[GFY10] Shi-Cai Gong, Yi-Zheng Fan, and Zhi-Xiang Yin. On the nullity of graphs with pendant trees. Linear Algebra and its Applications, 433(7):1374– 1380, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gillis:2012:GIN

[GG12] Nicolas Gillis and François Glineur. On the geometric interpretation of the nonnegative rank. Linear Algebra and its Applications, 437 (11):2685–2712, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005022.

Gutierrez-Gutierrez:2013:ECF

[GG13] Jesús Gutiérrez-Gutiérrez. Entries of continuous functions of large Hermitian tridiagonal 2-Toeplitz matrices. Linear Algebra and its Applications, 439(1):34–54, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002097.

Gouveia:2013:WNM

[GGK^{+}13] João Gouveia, Roland Grappe, Volker Kaibel, Kanstantsin Pashkovich, Richard Z. Robinson, and Rekha R. Thomas. Which nonnegative matrices are slack matrices? Linear Algebra and its Applications, 439(10):2921–2933, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005089.

Gath:2011:BME

[GH11] E. G. Gath and K. Hayes. Bounds for a multivariate extension of range over standard deviation based on the Mahalanobis distance. Linear Algebra and its Applications, 435(6):1267–1276, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gupta:2012:CSV

[GH12] Chander K. Gupta and Waldemar Holubowski. Commutator subgroup of Vershik– Kerov group. Linear Algebra and its Applications, 436(11):4279–4284, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000985.

Gao:2013:ASBa

[GH13a] You Gao and Yifan He. Association schemes based on singular symplectic geometry over finite fields and its application. Linear Algebra and its Applications, 438(1): 549–558, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005897.

Gao:2013:ASBb

[GH13b] You Gao and Yifan He. Association schemes based on the subspaces of type (m, s, 0) in singular symplectic space over finite fields. Linear Algebra and its Applications, 439 (11):3435–3444, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005752.

Ghahramani:2013:ZPD

[Gha13] Hoger Ghahramani. Zero product determined some nest algebras. Linear Algebra and its Applications, 438(1): 303–314, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005903.

Ghebleh:2014:AIM

[Ghe14a] M. Ghebleh. Antichains of (0, 1)-matrices through inversions. Linear Algebra and its Applications, 458(??): 503–511, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003875.

Ghebleh:2014:MCB

[Ghe14b] M. Ghebleh. On maximum chains in the Bruhat order of A(n, 2). Linear Algebra and its Applications, 446 (??):377–387, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Garcia:2011:MPA

[GHMPVP11] A. G. García, M. A. Hernández-Medina, G. Pérez-Villalón, and A. Portal. A matrix pencil approach to the existence of compactly supported reconstruction functions in average sampling. Linear Algebra and its Applications, 435 (11):2837–2859, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghosseiri:2013:BUT

[Gho13] Nader M. Ghosseiri. On biderivations of upper triangular matrix rings. Linear Algebra and its Applications, 438(1):250–260, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005988.

Gerasimova:2013:SUE

[GHS13] Tatiana G. Gerasimova, Roger A. Horn, and Vladimir V. Sergeichuk. Simultaneous unitary equivalences. Linear Algebra and its Applications, 438(10):3829–3835, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006963.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Grammont:2011:FAN

[GHT11] Laurence Grammont, Nicholas J. Higham, and Françoise Tisseur. A framework for analyzing nonlinear eigenproblems and parametrized linear systems. Linear Algebra and its Applications, 435(3):623– 640, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gong:2013:IWO

[GHW^{+}13] Shi-Cai Gong, Yao-Ping Hou, Ching-Wah Woo, Guang-Hui Xu, and Xiao-Ling Shen. On the integral weighted oriented unicyclic graphs with minimum skew energy. Linear Algebra and its Applications, 439(1):262–272, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001493.

Greco:2012:PIM

[GIP12] Federico Greco, Bruno Iannazzo, and Federico Poloni. The Padé iterations for the matrix sign function and their reciprocals are optimal. Linear Algebra and its Applications, 436(3):472– 477, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Ghobber:2011:UPF

[GJ11] Saifallah Ghobber and Philippe Jaming. On uncertainty principles in the finite dimensional setting. Linear Algebra and its Applications, 435(4):751– 768, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gratton:2013:SBE

[GJTP13] Serge Gratton, Pavel Jiránek, and David Titley-Peloquin. Simple backward error bounds for linear least-squares problems. Linear Algebra and its Applications, 439(1):78– 89, July 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001900.

Gover:2010:DVP

[GK10] Eugene Gover and Nishan Krikorian. Determinants and the volumes of parallelotopes and zonotopes. Linear Algebra and its Applications, 433 (1):28–40, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gongopadhyay:2011:EIN

[GK11] Krishnendu Gongopadhyay and Ravi S. Kulkarni. On the existence of an invariant non-degenerate bilinear form under a linear map. Linear Algebra and its Applications, 434(1):89–103, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gunther:2012:STB

[GK12] M. Günther and L. Klotz. Schur’s theorem for a block Hadamard product. Linear Algebra and its Applications, 437(3):948–956, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002650.

Ghareghani:2014:MCB

[GK14a] N. Ghareghani and G. B. Khosrovshahi. Minimum cycle basis of direct product of K_{2} × K_{n}. Linear Algebra and its Applications, 458(??):671–678, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400411X.

Goldberg:2014:SPS

[GK14b] Felix Goldberg and Steve Kirkland. On the sign patterns of the smallest signless Laplacian eigenvector. Linear Algebra and its Applications, 443(??):66–85, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Gunther:2014:LFP

[GK14c] M. Günther and L. Klotz. Lieb functions and m-positivity of norms. Linear Algebra and its Applications, 456(??): 54–63, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003595.

Guo:2011:CSS

[GKL11] Chun-Hua Guo, Yueh-Cheng Kuo, and Wen-Wei Lin. Complex symmetric stabilizing solution of the matrix equation X + A^{T} X^{−1}A = Q. Linear Algebra and its Applications, 435(6):1187–1192, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guillot:2013:CEC

[GKR13] Dominique Guillot, Apoorva Khare, and Bala Rajaratnam. The critical exponent conjecture for powers of doubly nonnegative matrices. Linear Algebra and its Applications, 439(8):2422–2427, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004308.

Griggs:2010:QLF

[GKS^{+}10] Wynita M. Griggs, Christopher K. King, Robert N. Shorten, Oliver Mason, and Kai Wulff. Quadratic Lyapunov functions for systems with state-dependent switching. Linear Algebra and its Applications, 433(1):52– 63, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Germina:2011:PLG

[GKZ11] K. A. Germina, Shahul Hameed K., and Thomas Zaslavsky. On products and line graphs of signed graphs, their eigenvalues and energy. Linear Algebra and its Applications, 435(10):2432–2450, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2010:SSC

[GL10a] Xing Gao and Yanfeng Luo. The spectrum of semi-Cayley graphs over abelian groups. Linear Algebra and its Applications, 432(11):2974–2983, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gessel:2010:NEP

[GL10b] Ira M. Gessel and Nicholas Loehr. Note on enumeration of partitions contained in a given shape. Linear Algebra and its Applications, 432(2– 3):583–585, January 15, 2010.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2010:CRS

[GL10c] Chun-Hua Guo and Wen-Wei Lin. Convergence rates of some iterative methods for nonsymmetric algebraic Riccati equations arising in transport theory. Linear Algebra and its Applications, 432 (1):283–291, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geng:2012:SRT

[GL12a] Xianya Geng and Shuchao Li. On the spectral radius of tricyclic graphs with a maximum matching. Linear Algebra and its Applications, 436(10):4043–4051, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379509006788.

Goodearl:2012:DTN

[GL12b] K. R. Goodearl and T. H. Lenagan. LU decomposition of totally nonnegative matrices. Linear Algebra and its Applications, 436(7): 2554–2566, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007464.

Guo:2012:SRS

[GL12c] Guangquan Guo and Juan Liu. Some results on the spectral radius of generalized ∞ and θ-digraphs. Linear Algebra and its Applications, 437 (9):2200–2208, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004302.

Guo:2014:SAF

[GL14] Jun Guo and Fenggao Li. Several anzahl formulas of classical spaces and their applications. Linear Algebra and its Applications, 453 (??):154–173, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002286.

Gleich:2011:RNA

[Gle11] David F. Gleich. Review of: “Numerical algorithms for personalized search in self-organizing information networks” by Sep Kamvar, Princeton Univ. Press, 2010, 160pp., ISBN-13: 978-0-691-14503-7. Linear Algebra and its Applications, 435(4):908– 909, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garvey:2013:FCI

[GLP^{+}13] S. D. Garvey, P. Lancaster, A. A. Popov, U. Prells, and I. Zaballa. Filters connecting isospectral quadratic systems. Linear Algebra and its Applications, 438(4):1497– 1516, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002643.

Geng:2010:SRQ

[GLS10] Xianya Geng, Shuchao Li, and Slobodan K. Simić. On the spectral radius of quasi-k-cyclic graphs. Linear Algebra and its Applications, 433(8– 10):1561–1572, December 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2012:SLS

[GLS12] Ji-Ming Guo, Jianxi Li, and Wai Chee Shiu. The smallest Laplacian spectral radius of graphs with a given clique number. Linear Algebra and its Applications, 437(4):1109–1122, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002923.

Griffing:2013:EIP

[GLS13a] Alexander R. Griffing, Benjamin R. Lynch, and Eric A.

Stone. An eigenvector interlacing property of graphs that arise from trees by Schur complementation of the Laplacian. Linear Algebra and its Applications, 438(3): 1078–1094, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007021.

Guo:2013:NUB

[GLS13b] Ji-Ming Guo, Jianxi Li, and Wai Chee Shiu. A note on the upper bounds for the Laplacian spectral radius of graphs. Linear Algebra and its Applications, 439(6):1657– 1661, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003108.

Guo:2013:AFS

[GLW13] Jun Guo, Fenggao Li, and Kaishun Wang. Anzahl formulas of subspaces in symplectic spaces and their applications. Linear Algebra and its Applications, 438(8): 3321–3335, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008476.

Gao:2014:MCS

[GLZ14] Wei Gao, Zhongshan Li, and Lihua Zhang. The minimal critical sets of refined inertias for 3 × 3 full sign patterns. Linear Algebra and its Applications, 458(??): 183–196, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003747.

Gursoy:2011:PAS

[GM11a] Buket Benek Gursoy and Oliver Mason. Pmax^{1} and Smax properties and asymptotic stability in the max algebra. Linear Algebra and its Applications, 435(5):1008– 1018, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gursoy:2011:SPM

[GM11b] Buket Benek Gursoy and Oliver Mason. Spectral properties of matrix polynomials in the max algebra. Linear Algebra and its Applications, 435(7):1626–1636, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Greenhill:2012:CLG

[GM12] Catherine Greenhill and Brendan D. McKay. Counting loopy graphs with given degrees. Linear Algebra and its Applications, 436(4): 901–926, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100276X.

Guo:2014:LRB

[GM14] Krystal Guo and Bojan Mohar. Large regular bipartite graphs with median eigenvalue 1. Linear Algebra and its Applications, 449 (??):68–75, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000834.

Goncalves:2014:TFL

[GMH14a] Vinicius Mariano Gonçalves, Carlos Andrey Maia, and Laurent Hardouin. On tropical fractional linear programming. Linear Algebra and its Applications, 459(??): 384–396, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400439X.

Goncalves:2014:WDR

[GMH14b] Vinicius Mariano Gonçalves, Carlos Andrey Maia, and Laurent Hardouin. Weak dual residuations applied to tropical linear equations. Linear Algebra and its Applications, 445(??):69–84, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Gamez-Merino:2012:BUP

[GMMFPSS12] José L. Gámez-Merino, Gustavo A. Muñoz-Fernández, Daniel Pellegrino, and Juan B. Seoane-Sepúlveda. Bounded and unbounded polynomials and multilinear forms: Characterizing continuity. Linear Algebra and its Applications, 436(1):237–242, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005003.

Granario:2013:PD

[GMP13] Daryl Q. Granario, Dennis I. Merino, and Agnes T. Paras. The ϕ_{S} polar decomposition. Linear Algebra and its Applications, 438 (1):609–620, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200585X.

Gutman:2014:KFT

[GMRS14] Ivan Gutman, Enide A. Martins, María Robbiano, and Bernardo San Martín. Ky Fan theorem applied to Randić energy. Linear Algebra and its Applications, 459(??):23–42, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004200.

Ganikhodzhaev:2012:DMR

[GMS12] Rasul Ganikhodzhaev, Farrukh Mukhamedov, and Mansoor Saburov. Gdecompositions of matrices and related problems I. Linear Algebra and its Applications, 436(5):1344–1366, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005933.

Gursoy:2013:AHP

[GMS13] Buket Benek Gursoy, Oliver Mason, and Sergeĭ Sergeev. The analytic hierarchy process, max algebra and multiobjective optimisation. Linear Algebra and its Applications, 438(7):2911–2928, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008257.

Grenet:2013:SDR

[GMT13] Bruno Grenet, Thierry Monteil, and Stéphan Thomassé. Symmetric Determinantal Representations in characteristic 2. Linear Algebra and its Applications, 439(5):1364– 1381, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Guglielmi:2011:CAH

[GMV11] Nicola Guglielmi, Carla Manni, and Davide Vitale. Convergence analysis of C^{2} Hermite interpolatory subdivision schemes by explicit joint spectral radius formulas. Linear Algebra and its Applications, 434(4):884–902, February 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gemignani:2013:EAM

[GN13] Luca Gemignani and Vanni Noferini. The Ehrlich–Aberth method for palindromic matrix polynomials represented in the Dickson basis. Linear Algebra and its Applications, 438(4):1645–1666, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007324.

Geher:2014:MCH

[GN14] György Pál Gehér and Gergő Nagy. Maps on classes of Hilbert space operators preserving measure of commutativity. Linear Algebra and its Applications, 463(??):205– 227, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005709.

Gnacadja:2012:JCS

[Gna12] Gilles Gnacadja. A Jacobian criterion for the simultaneous injectivity on positive variables of linearly parameterized polynomial maps. Linear Algebra and its Applications, 437(2):612–622, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002145.

Giryes:2014:GLA

[GNE^{+}14] R. Giryes, S. Nam, M. Elad, R. Gribonval, and M. E. Davies. Greedy-like algorithms for the cosparse analysis model. Linear Algebra and its Applications, 441 (??):22–60, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001870.

Griebel:2012:GRV

[GO12] Michael Griebel and Peter Oswald. Greedy and randomized versions of the multiplicative Schwarz method. Linear Algebra and its Applications, 437(7):1596–1610, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003540.

Guiver:2013:EBG

[GO13] Chris Guiver and Mark R. Opmeer. Error bounds in the gap metric for dissipative balanced approximations. Linear Algebra and its Applications, 439(12):3659–3698, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005880.

Godjali:2010:THP

[God10] Ali Godjali. Thin Hessenberg pairs. Linear Algebra and its Applications, 432(12): 3231–3249, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godjali:2012:THP

[God12] Ali Godjali. Thin Hessenberg pairs and double Vandermonde matrices. Linear Algebra and its Applications, 436(9):3018–3060, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006446.

Goldberg:2013:OCV

[Gol13] Felix Goldberg. Optimizing Colin de Verdière matrices of K_{4,4}. Linear Algebra and its Applications, 438(10):4090–4101, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006702.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Gomez:2010:MRT

[Góm10] Ricardo Gómez. On mean recurrence times of Markov chains and spanning tree invariants. Linear Algebra and its Applications, 433(11– 12):1714–1718, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:2011:ONI

[Gon11] María Celeste Gonzalez. Operator norm inequalities in semi-Hilbertian spaces. Linear Algebra and its Applications, 434(1):370–378, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghidini:2012:CHP

[GOSV12] Carla T. L. S. Ghidini, A. R. L. Oliveira, Jair. Silva, and M. I. Velazco. Combining a hybrid preconditioner and a optimal adjustment algorithm to accelerate the convergence of interior point methods. Linear Algebra and its Applications, 436(5):1267–1284, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Garnett:2014:INM

[GOSvdD14] C. Garnett, D. D. Olesky, B. L. Shader, and P. van den Driessche. Integrally normalizable matrices and zerononzero patterns. Linear Algebra and its Applications, 449(??):132–153, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000718.

Grundy:2012:CPS

[GOvdD12] D. A. Grundy, D. D. Olesky, and P. van den Driessche. Constructions for potentially stable sign patterns. Linear Algebra and its Applications, 436(12):4473–4488, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005921.

Garnett:2014:NSP

[GOvdD14] C. Garnett, D. D. Olesky, and P. van den Driessche. A note on sign patterns of order 3 that require particular refined inertias. Linear Algebra and its Applications, 450(??):293–300, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001372.

Ghanbari:2012:GIE

[GP12] K. Ghanbari and F. Parvizpour. Generalized inverse eigenvalue problem with mixed eigendata. Linear Algebra and its Applications, 437(8): 2056–2063, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003825.

Gillis:2013:SNM

[GP13a] Nicolas Gillis and Robert J. Plemmons. Sparse nonnegative matrix underapproximation and its application to hyperspectral image analysis. Linear Algebra and its Applications, 438(10): 3991–4007, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003096.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Gongopadhyay:2013:RCH

[GP13b] Krishnendu Gongopadhyay and John R. Parker. Reversible complex hyperbolic isometries. Linear Algebra and its Applications, 438(6): 2728–2739, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008361.

Goncalves:2014:IGN

[GP14] Maria Inez Cardoso Gonçalves and Vladimir G. Pestov. Isometries of a generalized numerical radius on compact operators. Linear Algebra and its Applications, 450 (??):301–315, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001074.

Garcia:2013:LPP

[GPR13] E. García, F. Pedroche, and M. Romance. On the localization of the personalized PageRank of complex networks. Linear Algebra and its Applications, 439(3): 640–652, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007835.

Garcia:2012:UEC

[GPT12] Stephan Ramon Garcia, Daniel E. Poore, and James E. Tener. Unitary equivalence to a complex symmetric matrix: Low dimensions. Linear Algebra and its Applications, 437 (1):271–284, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000936.

Gavalec:2014:IEM

[GPT14] Martin Gavalec, Ján Plavka, and Hana Tomásková. Interval eigenproblem in maxmin algebra. Linear Algebra and its Applications, 440 (??):24–33, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006691.

Gudino:2010:LBS

[GR10] E. Gudiño and J. Rada. A lower bound for the spectral radius of a digraph. Linear Algebra and its Applications, 433(1):233–240, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guillot:2012:RPD

[GR12] Dominique Guillot and Bala Rajaratnam. Retaining positive definiteness in thresholded matrices. Linear Algebra and its Applications, 436(11):4143–4160, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000614.

Grcar:2010:MLB

[Grc10] Joseph F. Grcar. A matrix lower bound. Linear Algebra and its Applications, 433(1): 203–220, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-

Goberna:2012:VCL

[GRdS12] M. A. Goberna, M. M. L. Rodríguez, and V. N. Vera de Serio. Voronoi cells via linear inequality systems. Linear Algebra and its Applications, 436(7):2169–2186, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008330.

Greaves:2012:CMR

[Gre12] Gary Greaves. Cyclotomic matrices over real quadratic integer rings. Linear Algebra and its Applications, 437 (9):2252–2261, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004338.

Greenbaum:2013:BRB

[Gre13] Anne Greenbaum. Book review: Krylov Subspace Methods: Principles and Analysis, Jörg Liesen, Zdenek Strakos (2013). Linear Algebra and its Applications, 439 (9):2711, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004266.

Gutman:2010:ELG

[GRM^{+}10] Ivan Gutman, María Robbiano, Enide Andrade Martins, Domingos M. Cardoso, Luis Medina, and Oscar Rojo. Energy of line graphs. Linear Algebra and its Applications, 433(7):1312–1323, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grover:2014:OMS

[Gro14] Priyanka Grover. Orthogonality to matrix subspaces, and a distance formula. Linear Algebra and its Applications, 445(??):280–288, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007787.

Gutman:2010:SIH

[GRS^{+}10] Ivan Gutman, Peter Rowlinson, Slobodan K. Simic, Dragan Stevanovic, and Edwin R. van Dam. Special issue in honor of Dragos Cvetković. Linear Algebra and its Applications, 432(10):2727, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutman:2011:DC

[GRS^{+}11] Ivan Gutman, Peter Rowlinson, Slobodan K. Simić, Dragan Stevanović, and Edwin R. van Dam. Dragos Cvetkovic,

70. Linear Algebra and its Applications, 435(10):2321–2333, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grunbaum:2012:RBG

[Grü12] Branko Grünbaum. Review of Matrices and Graphs in Geometry. Encyclopedia of Mathematics and its Applications, vol. 139 by Miroslav Fiedler, Cambridge University Press (2011). viii + 197 pp., ISBN: 978-0-521-46193-1. Linear Algebra and its Applications, 436(7): 2709–2710, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006252.

Ganikhodzhaev:2010:DSQ

[GS10a] Rasul Ganikhodzhaev and Farruh Shahidi. Doubly stochastic quadratic operators and Birkhoff’s problem. Linear Algebra and its Applications, 432(1):24–35, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gomez:2010:STI

[GS10b] Ricardo Gómez and José Miguel Salazar. Spanning tree invariants, loop systems and doubly stochastic matrices. Linear Algebra and its Applications, 432(2–3):556–565, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2010:SCS

[GS10c] M. Seetharama Gowda and Roman Sznajder. Schur complements, Schur determinantal and Haynsworth inertia formulas in Euclidean Jordan algebras. Linear Algebra and its Applications, 432(6):1553– 1559, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guersenzvaig:2010:SMA

[GS10d] N. H. Guersenzvaig and Fernando Szechtman. Subalgebras of matrix algebras generated by companion matrices. Linear Algebra and its Applications, 432(10):2691–2700, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gustafson:2010:SAS

[GS10e] Karl Gustafson and Morteza Seddighin. Slant antieigenvalues and slant antieigenvectors of operators. Linear Algebra and its Applications, 432(5): 1348–1362, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2011:NCS

[GS11a] Yubin Gao and Yanling Shao. New classes of spectrally arbitrary ray patterns. Linear Algebra and its Applications, 434(10):2140–2148, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutman:2011:ECW

[GS11b] Ivan Gutman and Jia-Yu Shao. The energy change of weighted graphs. Linear Algebra and its Applications, 435 (10):2425–2431, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garnett:2012:PCC

[GS12a] Colin Garnett and Bryan L. Shader. A proof of the T_{n} conjecture: Centralizers, Jacobians and spectrally arbitrary sign patterns. Linear Algebra and its Applications, 436(12):4451–4458, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005015.

Gasiorek:2012:OPP

[GS12b] Marcin Gaşiorek and Daniel Simson. One-peak posets with positive quadratic Tits form, their mesh translation quivers of roots, and programming in Maple and Python. Linear Algebra and its Applications, 436(7):2240–2272, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007555.

Guersenzvaig:2012:EMS

[GS12c] N. H. Guersenzvaig and Fernando Szechtman. Is every matrix similar to a polynomial in a companion matrix? Linear Algebra and its Applications, 437(7): 1611–1627, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003382.

Guersenzvaig:2012:RMS

[GS12d] N. H. Guersenzvaig and Fernando Szechtman. Roots multiplicity and square-free factorization of polynomials using companion matrices. Linear Algebra and its Applications, 436(9):3160–3164, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007154.

Gulde:2012:SSU

[GS12e] Michael Gulde and Markus Stroppel. Stabilizers of subspaces under similitudes of the Klein quadric, and automorphisms of Heisenberg algebras. Linear Algebra and its Applications, 437(4): 1132–1161, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002492.

Guterman:2012:TPM

[GS12f] A. E. Guterman and Ya. N. Shitov. Tropical patterns of matrices and the Gondran– Minoux rank function. Linear Algebra and its Applications, 437(7):1793–1811, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003047.

Garcia:2013:MEI

[GS13a] Stephan Ramon Garcia and Amy L. Shoemaker. On the matrix equation XA + AX^{T} = 0. Linear Algebra and its Applications, 438(6): 2740–2746, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007719.

See also Part II [CGGS13].

Garnett:2013:NCM

[GS13b] Colin Garnett and Bryan L. Shader. The nilpotentcentralizer method for spectrally arbitrary patterns. Linear Algebra and its Applications, 438(10):3836–3850, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379511006951.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Guo:2011:ACL

[GSL11] Ji-Ming Guo, Wai Chee Shiu, and Jianxi Li. The algebraic connectivity of lollipop graphs. Linear Algebra and its Applications, 434(10): 2204–2210, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2013:AGC

[GST13] M. Seetharama Gowda, Roman Sznajder, and Jiyuan Tao. The automorphism group of a completely positive cone and its Lie algebra. Linear Algebra and its Applications, 438(10): 3862–3871, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006987.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Gowda:2012:PLL

[GTR12] M. Seetharama Gowda, Jiyuan Tao, and G. Ravindran. On the P -property of Z and Lyapunov-like transformations on Euclidean Jordan algebras. Linear Algebra and its Applications, 436 (7):2201–2209, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007166.

Grubisic:2012:REP

[GTV12] Luka Grubisić, Ninoslav Truhar, and Kresimir Veselić. The rotation of eigenspaces of perturbed matrix pairs. Linear Algebra and its Applications, 436(11):4161–4178, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512000900.

Gau:2013:WSM

[GTW13] Hwa-Long Gau, Ming-Cheng Tsai, and Han-Chun Wang. Weighted shift matrices: Unitary equivalence, reducibility and numerical ranges. Linear Algebra and its Applications, 438(1):498–513, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006428.

Gu:2013:SSG

[Gu13] Zhenhua Gu. Subconstituents of symplectic graphs modulo p^{n}. Linear Algebra and its Applications, 439(5):1321– 1329, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002772.

Gu:2014:EOW

[Gu14] Caixing Gu. Elementary operators which are m-isometries.

Linear Algebra and its Applications, 451(??):49–64, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001645.

Gulinsky:2011:SSS

[Gul11] O. Gulinsky. On stochastic setting of stationary phase method. Linear Algebra and its Applications, 435(7):1575– 1584, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gumbrell:2013:STV

[Gum13] Lee Gumbrell. A subdivision theorem for vertices not on internal paths. Linear Algebra and its Applications, 439 (10):2790–2794, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005569.

Guo:2010:NMH

[Guo10a] Chun-Hua Guo. On Newton’s method and Halley’s method for the principal p th root of a matrix. Linear Algebra and its Applications, 432(8):1905– 1922, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2010:ACG

[Guo10b] Ji-Ming Guo. The algebraic connectivity of graphs under perturbation. Linear Algebra and its Applications, 433(6): 1148–1153, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2013:ARE

[Guo13] Chun-Hua Guo. On algebraic Riccati equations associated with M-matrices. Linear Algebra and its Applications, 439(10):2800–2814, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300517X.

Guven:2012:SDC

[Güv12] Bilgehan Güven. The spectral decomposition of a covariance matrix for the balanced mixed analysis of variance model. Linear Algebra and its Applications, 436 (9):3337–3346, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007713.

Gracia:2011:LSC

[GV11] Juan-Miguel Gracia and Francisco E. Velasco. Lipschitz stability of controlled invariant subspaces. Linear Algebra and its Applications, 434(4):1137– 1162, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2010:DIP

[GW10] Hwa-Long Gau and Pei Yuan Wu. Defect indices of powers of a contraction. Linear Algebra and its Applications, 432 (11):2824–2833, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gu:2011:SOG

[GW11] Zhenhua Gu and Zhe-Xian Wan. Subconstituents of orthogonal graphs of odd characteristic. Linear Algebra and its Applications, 434(12): 2430–2447, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2013:HRN

[GW13a] Hwa-Long Gau and Pei Yuan Wu. Higher-rank numerical ranges and Kippenhahn polynomials. Linear Algebra and its Applications, 438 (7):3054–3061, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008221.

Guo:2013:SLS

[GW13b] Guangquan Guo and Guoping Wang. On the (signless)

Laplacian spectral characterization of the line graphs of lollipop graphs. Linear Algebra and its Applications, 438(12):4595–4605, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008701.

Gau:2014:PPI

[GW14a] Hwa-Long Gau and Pei Yuan Wu. Power partial isometry index and ascent of a finite matrix. Linear Algebra and its Applications, 459(??):136–144, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004388.

Gau:2014:SNR

[GW14b] Hwa-Long Gau and Pei Yuan Wu. Structures and numerical ranges of power partial isometries. Linear Algebra and its Applications, 440(??):325–341, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007106.

Gao:2013:CLT

[GWH13] Suogang Gao, Yan Wang, and Bo Hou. The classification of Leonard triples of Racah type. Linear Algebra and its Applications, 439(7):1834–1861, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003418.

Gau:2014:ZDI

[GWW14a] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Zero-dilation index of a finite matrix. Linear Algebra and its Applications, 440(??): 111–124, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006769.

Guo:2014:IER

[GWW14b] Ying-Jun Guo, Zhi-Xiong Wen, and Wen Wu. On the irrationality exponent of the regular paperfolding numbers. Linear Algebra and its Applications, 446 (??):237–264, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008318.

Gu:2013:SOG

[GWZ13] Zhenhua Gu, Zhe-Xian Wan, and Kai Zhou. Subconstituents of orthogonal graphs of odd characteristic — continued. Linear Algebra and its Applications, 439(10):2861– 2877, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005090.

Gong:2012:RDO

[GX12a] Shi-Cai Gong and Guang-Hui Xu. 3-regular digraphs with optimum skew energy. Linear Algebra and its Applications, 436(3):465–471, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003065.

Gong:2012:CPM

[GX12b] Shi-Cai Gong and Guang-Hui Xu. The characteristic polynomial and the matchings polynomial of a weighted oriented graph. Linear Algebra and its Applications, 436 (9):3597–3607, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000031.

Gong:2012:NGC

[GX12c] Shi-Cai Gong and Guang-Hui Xu. On the nullity of a graph with cut-points. Linear Algebra and its Applications, 436(1):135–142, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004897.

Guo:2013:PRU

[GY13] Xiangqian Guo and Guangyu Yang. The probability of rectangular unimodular matrices over F_{q}[x]. Linear Algebra and its Applications, 438(6):2675–2682, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008038.

Guglielmi:2012:ARF

[GZ12] N. Guglielmi and M. Zennaro. On the asymptotic regularity of a family of matrices. Linear Algebra and its Applications, 436(7):2093–2104, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007609.

Gutkin:2013:JNR

[GZ13] Eugene Gutkin and Karol Zyczkowski. Joint numerical ranges, quantum maps, and joint numerical shadows. Linear Algebra and its Applications, 438(5):2394– 2404, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007732.

Gao:2014:TAJ

[GZH14] Suogang Gao, Liwei Zhang, and Bo Hou. The Terwilliger algebras of Johnson graphs. Linear Algebra and its Applications, 443 (??):164–183, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007167.

Gao:2014:ASB

[GZX14] You Gao, Xiaojuan Zhang, and Yanyan Xue. Association schemes based on singular linear spaces and its applications. Linear Algebra and its Applications, 453 (??):125–140, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001931.

Ha:2013:NEC

[Ha13] Kil-Chan Ha. Notes on extremality of the Choi map. Linear Algebra and its Applications, 439(10):3156–3165, November 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005661.

Hadjidimos:2012:IEO

[Had12] Apostolos Hadjidimos. Irreducibility and extensions of Ostrowski’s Theorem. Linear Algebra and its Applications, 436(7):2156–2168, April 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007786.

Hadjidimos:2013:NOT

[Had13] Apostolos Hadjidimos. A note on Ostrowski’s Theorem. Linear Algebra and its Applications, 439(12):3785– 3795, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006289.

Hanna:2011:MRD

[Han11a] L. A-M. Hanna. On matrix representations of deformed Lie algebras for quantized Hamiltonians. Linear Algebra and its Applications, 434(2):507–513, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanson:2011:CLP

[Han11b] Edward Hanson. A characterization of Leonard pairs using the notion of a tail. Linear Algebra and its Applications, 435(11):2961–2970, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2013:FTL

[Han13a] Frank Hansen. The fast track to Löwner’s theorem. Linear Algebra and its Applications, 438(11):4557–4571, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000840.

Hanson:2013:CLP

[Han13b] Edward Hanson. A characterization of Leonard pairs using the parameters {a_{i}}_{i=0}^{d}. Linear Algebra and its Applications, 438(5):2289–2305, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007513.

Hansen:2014:ROM

[Han14a] Frank Hansen. Regular operator mappings and multivariate geometric means. Linear Algebra and its Applications, 461(??):123–138, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004819.

Hanson:2014:CBL

[Han14b] Edward Hanson. A characterization of bipartite Leonard pairs using the notion of a tail. Linear Algebra and its Applications, 452(??):46–67, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001414.

Har:2014:NLE

[Har14] Jan Har. A note on Laplacian eigenvalues and domination. Linear Algebra and its Applications, 449(??):115–118, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000925.

Holzel:2012:PMM

[HB12] Matthew S. Holzel and Dennis S. Bernstein. From polynomial matrices to Markov parameters and back: Theory and numerical algorithms. Linear Algebra and its Applications, 437(3): 783–808, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001966.

Huang:2010:TNN

[HC10] Rong Huang and Delin Chu. Total nonpositivity of nonsingular matrices. Linear Algebra and its Applications, 432(11): 2931–2941, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hai:2014:CIU

[HC14] Guojun Hai and Alatancang Chen. Consistent invertibility of upper triangular operator matrices. Linear Algebra and its Applications, 455 (??):22–31, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002535.

Huang:2010:NUO

[HCY10a] Liang-Hao Huang, Gerard J. Chang, and Hong-Gwa Yeh. A note on universally optimal matrices and field independence of the minimum rank of a graph. Linear Algebra and its Applications, 433(3):585– 594, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:MRZ

[HCY10b] Liang-Hao Huang, Gerard J. Chang, and Hong-Gwa Yeh. On minimum rank and zero forcing sets of a graph. Linear Algebra and its Applications, 432(11):2961–2973, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Halikias:2012:SSD

[HDPT12] G. Halikias, L. Dritsas, A. Pantelous, and V. Tsoulkas. Strong stability of discretetime systems. Linear Algebra and its Applications, 436 (7):1890–1908, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006586.

He:2011:RAA

[He11] Tian-Xiao He. Riordan arrays associated with Laurent series and generalized Sheffertype groups. Linear Algebra and its Applications, 435 (6):1241–1256, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

He:2013:PCN

[He13] Tian-Xiao He. Parametric Catalan numbers and Catalan triangles. Linear Algebra and its Applications, 438 (3):1467–1484, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006982.

He:2014:CCI

[He14]  Junhua He.  On a con-

jecture concerning integral real roots of certain cubic polynomials. Linear Algebra and its Applications, 446 (??):265–268, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008379.

Henrion:2010:DRC

[Hen10] Didier Henrion. Detecting rigid convexity of bivariate polynomials. Linear Algebra and its Applications, 432(5): 1218–1233, February 15, 2010.

Heersink:2012:MPI

[HF12] Daniel K. Heersink and Reinhard Furrer. On Moore– Penrose inverses of quasi-Kronecker structured matrices. Linear Algebra and its Applications, 436(3):561– 570, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005210.

Hardy:2013:CTK

[HFS13] Yorick Hardy, Ajda Fosner, and Willi-Hans Steeb. Cayley transform and the Kronecker product of Hermitian matrices. Linear Algebra and its Applications, 439 (12):4023–4031, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006149.

Hou:2010:SSL

[HG10] Bo Hou and Suogang Gao. The shape of some linear transformations. Linear Algebra and its Applications, 433 (11–12):2088–2095, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hemmat:2011:KGO

[HG11a] A. Askari Hemmat and J.-P. Gabardo. Kernels of Gramian operators for frames in shiftinvariant subspaces. Linear Algebra and its Applications, 435(5):911–922, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2011:SPS

[HG11b] Bo Hou and Suogang Gao. Some properties of sharp Hessenberg pairs. Linear Algebra and its Applications, 435(8): 1987–1996, October 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2012:SSL

[HG12] Bo Hou and Suogang Gao. The structure of some linear transformations. Linear Algebra and its Applications, 437(9):2110–2116, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004211.

Huang:2011:IPP

[HH11a] Jie Huang and Ting-Zhu Huang. The inverse positivity of perturbed tridiagonal Mmatrices. Linear Algebra and its Applications, 434(1):131– 143, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:MCP

[HH11b] Li Huang and Jinchuan Hou. Maps completely preserving spectral functions. Linear Algebra and its Applications, 435(11):2756–2765, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadwin:2012:CNR

[HH12a] Don Hadwin and Deguang Han. The correlation numerical range of a matrix and Connes’ embedding problem. Linear Algebra and its Applications, 436(9):3139– 3148, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007142.

Hajja:2012:OMT

[HH12b] Mowaffaq Hajja and Mostafa Hayajneh. The open mouth theorem in higher dimensions. Linear Algebra and its Applications, 437(3):1057– 1069, August 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002121.

Heunen:2013:MMD

[HH13] Chris Heunen and Clare Horsman. Matrix multiplication is determined by orthogonality and trace. Linear Algebra and its Applications, 439 (12):4130–4134, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006071.

Harmon:2014:GSS

[HH14] Hank G. Harmon and Randall R. Holmes. Geometry of standard symmetrized tensors. Linear Algebra and its Applications, 460(??): 136–150, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004959.

Hartman:2010:PEI

[HHL10] Jim Hartman, Aaron Herman, and Chi-Kwong Li. Preservers of eigenvalue inclusion sets. Linear Algebra and its Applications, 433(5):1038– 1051, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2014:GAM

[HHLS14] Li-Ping Huang, Zejun Huang, Chi-Kwong Li, and Nung-Sing Sze. Graphs associated with matrices over finite fields and their endomorphisms. Linear Algebra and its Applications, 447(??):2–25, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Hall:2010:EVP

[HHMS10] H. Tracy Hall, Leslie Hogben, Ryan Martin, and Bryan Shader. Expected values of parameters associated with the minimum rank of a graph. Linear Algebra and its Applications, 433(1):101–117, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holmes:2013:ABI

[HHT13] Randall R. Holmes, Huajun Huang, and Tin-Yau Tam. Asymptotic behavior of Iwasawa and Cholesky iterations. Linear Algebra and its Applications, 438(10): 3755–3768, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003636.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Hiai:2013:CCM

[Hia13] Fumio Hiai. Concavity of certain matrix trace and norm functions. Linear Algebra and its Applications, 439(5):1568– 1589, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002826.

Hildebrand:2012:ERC

[Hil12a] Roland Hildebrand. The extreme rays of the 5 × 5 copositive cone. Linear Algebra and its Applications, 437(7): 1538–1547, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002935.

Hilscher:2012:OTD

[Hil12b] Roman Simon Hilscher. Oscillation theorems for discrete symplectic systems with nonlinear dependence in spectral parameter. Linear Algebra and its Applications, 437 (12):2922–2960, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200496X.

Hill:2013:IDS

[Hil13] Jordan Dale Hill. An identity of degree 2n − 3 for the n×n skews, n even, and corollaries for standard identities. Linear Algebra and its Applications, 439(2):310–327, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002140.

Hilberdink:2014:GSI

[Hil14a] Titus Hilberdink. The group of squarefree integers. Linear Algebra and its Applications, 457(??):383–399, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003395.

Hildebrand:2014:MZC

[Hil14b] Roland Hildebrand. Minimal zeros of copositive matrices. Linear Algebra and its Applications, 459(??): 154–174, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004418.

Hiller:2014:PTC

[Hil14c] Josh Hiller. A proof of two conjectures of Deveci and Karaduman. Linear Algebra and its Applications, 446 (??):163–165, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008355.

Hirn:2010:NHF

[Hir10] Matthew Hirn. The number of harmonic frames of prime order. Linear Algebra and its Applications, 432(5):1105– 1125, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2012:MA

[HJ12] Roger A. Horn and Charles R. (Charles Royal) Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, second edition, 2012. ISBN 0-521-83940-8 (hardcover), 0-521-54823-3 (paperback), 1-283-74139-3, 1-139-77904-4, 1-139-77600-2 (e-book), 1-139-02041-2 (ebook). xviii + 643 pp. LCCN QA188 .H66 2012.

Huo:2010:NUG

[HJL10] Bofeng Huo, Shengjin Ji, and Xueliang Li. Note on unicyclic graphs with given number of pendent vertices and minimal energy. Linear Algebra and its Applications, 433(7):1381– 1387, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huo:2011:SCM

[HJLS11] Bofeng Huo, Shengjin Ji, Xueliang Li, and Yongtang Shi. Solution to a conjecture on the maximal energy of bipartite bicyclic graphs. Linear Algebra and its Applications, 435(4):804–810, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoff:2012:TPS

[HJN12] D. Hoff, C. R. Johnson, and S. Nasserasr. Totally positive shapes and TP_{k}-completable patterns. Linear Algebra and its Applications, 436(12): 4412–4422, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004320.

He:2013:SBS

[HJZ13] Bian He, Ya-Lei Jin, and Xiao-Dong Zhang. Sharp bounds for the signless Laplacian spectral radius in terms of clique number. Linear Algebra and its Applications, 438(10):3851–3861, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007014.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Hirzallah:2010:SVN

[HK10] Omar Hirzallah and Fuad Kittaneh. Singular values, norms, and commutators. Linear Algebra and its Applications, 432(5):1322–1336, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Homma:2013:NPF

[HK13] Masaaki Homma and Seon Jeong Kim. Nonsingular plane filling curves of minimum degree over a finite field and their automorphism groups: Supplements to a work of Tallini. Linear Algebra and its Applications, 438(3): 969–985, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006568.

Hirzallah:2012:EID

[HKK^{+}12] Omar Hirzallah, Fuad Kittaneh, Mario Krnić, Neda Lovricević, and Josip Pecarić. Eigenvalue inequalities for differences of means of Hilbert space operators. Linear Algebra and its Applications, 436(5):1516–1527, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006215.

Halikias:2013:DSS

[HKP13] G. D. Halikias, N. Karcanias, and A. Papageorgiou. The distance to strong stability. Linear Algebra and its Applications, 439(10):2721– 2735, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005521.

Hiai:2013:FCP

[HKPR13] Fumio Hiai, Hideki Kosaki, Dénes Petz, and Mary Beth Ruskai. Families of completely positive maps associated with monotone metrics. Linear Algebra and its Applications, 439(7):1749–1791, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003388.

Hansen:2010:BCS

[HL10a] Pierre Hansen and Claire Lucas. Bounds and conjectures for the signless Laplacian index of graphs. Linear Algebra and its Applications, 432(12): 3319–3336, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:SCS

[HL10b] Rong Huang and Jianzhou Liu. On Schur complements of sign regular matrices of order k. Linear Algebra and its Applications, 433(1):143–148, July 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:GSI

[HL10c] Yufei Huang and Bolian Liu. Generalized scrambling indices of a primitive digraph. Linear Algebra and its Applications, 433(11–12):1798– 1808, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

He:2011:OTF

[HL11a] Shushan He and Shuchao Li. Ordering of trees with fixed matching number by the Laplacian coefficients. Linear Algebra and its Applications, 435(5):1171–1186, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2011:RSM

[HL11b] Nicholas J. Higham and Lijing Lin. On p th roots of stochastic matrices. Linear Algebra and its Applications, 435(3): 448–463, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:BFS

[HL11c] Rong Huang and Jianzhou Liu. Bidiagonal factorizations with some parameters equal to zero. Linear Algebra and its Applications, 434(3):730– 740, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:ZPI

[HL11d] Rong Huang and Jianzhou Liu. On zero-pattern invariant properties of structured matrices. Linear Algebra and its Applications, 435(4):871– 883, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2011:CBW

[HL11e] Seok Yoon Hwang and Jeong Yeon Lee. Construction of biorthogonal wavelet vectors. Linear Algebra and its Applications, 434(4):1171– 1188, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

He:2012:SLI

[HL12] Shushan He and Shuchao Li. On the signless Laplacian index of unicyclic graphs with fixed diameter. Linear Algebra and its Applications, 436(1):252–261, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005131.

Hladik:2013:WSS

[Hla13] Milan Hladík. Weak and strong solvability of interval linear systems of equations and inequalities. Linear Algebra and its Applications, 438(11):4156–4165, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001134.

Hu:2012:MER

[HLP12] Guikai Hu, Qingguo Li, and Ping Peng. Minimax estimator of regression coefficient in normal distribution under balanced loss function. Linear Algebra and its Applications, 436(5):1228–1237, March 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005945.

Huang:2010:CIM

[HLS10] Rong Huang, Jianzhou Liu, and Nung-Sing Sze. Characterizations of inverse Mmatrices with special zero patterns. Linear Algebra and its Applications, 433(5):994– 1000, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huo:2011:CSP

[HLS11] Bofeng Huo, Xueliang Li, and Yongtang Shi. Complete solution to a problem on the maximal energy of unicyclic bipartite graphs. Linear Algebra and its Applications, 434 (5):1370–1377, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Henk:2010:MZC

[HLW10a] Martin Henk, Eva Linke, and Jörg M. Wills. Minimal zonotopes containing the crosspolytope. Linear Algebra and its Applications, 432(11): 2942–2952, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2010:MPS

[HLW10b] Jinchuan Hou, Chi-Kwong Li, and Ngai-Ching Wong. Maps preserving the spectrum of generalized Jordan product of operators. Linear Algebra and its Applications, 432(4):1049– 1069, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

He:2014:TDC

[HLW14] Changxiang He, Shiqiong Liu, and Baofeng Wu. Three distance characteristic polynomials of some graphs. Linear Algebra and its Applications, 452(??):281–291, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400189X.

Huang:2012:NAS

[HLZ12] Rong Huang, Jianzhou Liu, and Li Zhu. Nonsingular almost strictly sign regular matrices. Linear Algebra and its Applications, 436(11):4179–4192, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000602.

Huang:2013:BKF

[HLZP13] Rong Huang, Jianzhou Liu, Li Zhu, and Jinsong Pan. The Bunch–Kaufman factorization of symmetric matrices signature similar to sign regular matrices. Linear Algebra and its Applications, 439 (5):1458–1467, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003042.

Hogben:2010:LAV

[HM10] Leslie Hogben and Jillian McLeod. A linear algebraic view of partition regular matrices. Linear Algebra and its Applications, 433 (11–12):1809–1820, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heineken:2014:PFD

[HM14a] Sigrid B. Heineken and Patricia M. Morillas. Properties of finite dual fusion frames. Linear Algebra and its Applications, 453(??):1–27, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002225.

Hille:2014:TPP

[HM14b] Lutz Hille and Jürgen Müller. On tensor products of path algebras of type A. Linear Algebra and its Applications, 448(??):222–244, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514000652.

Harris:2011:NRC

[HMP^{+}11] Thomas Ryan Harris, Michael Mazzella, Linda J. Patton, David Renfrew, and Ilya M. Spitkovsky. Numerical ranges of cube roots of the identity. Linear Algebra and its Applications, 435(11):2639–2657, December 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hochstenbach:2012:LQT

[HMP12] Michiel E. Hochstenbach, Andrej Muhic, and Bor Plestenjak. On linearizations of the quadratic two-parameter eigenvalue problem. Linear Algebra and its Applications, 436(8):2725–2743, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005386.

Hochstenbach:2012:DIP

[HMR12] M. E. Hochstenbach, N. McNinch, and L. Reichel. Discrete ill-posed least-squares problems with a solution norm constraint. Linear Algebra and its Applications, 436(10): 3801–3818, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005404.

Holtz:2013:MCT

[HMS13] Olga Holtz, Volker Mehrmann, and Hans Schneider. Matrices that commute with their derivative. On a letter from Schur to Wielandt. Linear Algebra and its Applications, 438(5):2574–2590, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007203.

Henn:2010:HSH

[HMT10] Mark-Alexander Henn, Christian Mehl, and Carsten Trunk. Hyponormal and strongly hyponormal matrices in inner product spaces. Linear Algebra and its Applications, 433(6):1055–1076, November 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haemers:2010:SLE

[HMTR10] W. H. Haemers, A. Mohammadian, and B. Tayfeh-Rezaie. On the sum of Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 432(9):2214–2221, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:IAS

[HN10] Jianguo Huang and Liwei Nong. An iterative algorithm for solving a finitedimensional linear operator equation T (x) = f with applications. Linear Algebra and its Applications, 432(5):1176– 1188, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Honma:2014:IGD

[HN14] Shigekazu Honma and Tatsuya Nogawa. Isometries of the geodesic distances for the space of invertible positive operators and matrices. Linear Algebra and its Applications, 444(??):152–164, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007751.

Hogenson:2012:MSA

[HNZ12] Kirsten Hogenson, Shannon Negaard, and Ryan J. Zerr. Matrix sequences associated with the Ducci map and the mediant construction of the rationals. Linear Algebra and its Applications, 437 (1):285–293, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001401.

Haemers:2011:UAM

[HO11] W. H. Haemers and G. R. Omidi. Universal adjacency matrices with two eigenvalues. Linear Algebra and its Applications, 435(10):2520–2529, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2010:MRP

[Hog10] Leslie Hogben. Minimum rank problems. Linear Algebra and its Applications, 432(8):1961– 1974, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoa:2013:GPS

[HOT13] Dinh Trung Hoa, Hiroyuki Osaka, and Ho Minh Toan. On generalized Powers–Størmer’s inequality. Linear Algebra and its Applications, 438(1): 242–249, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006167.

Hwang:2004:IEP

[HP04] Suk-Geun Hwang and Sung-Soo Pyo. The inverse eigenvalue problem for symmetric doubly stochastic matrices. Linear Algebra and its Applications, 379(1):77–83, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See comments [Fan10b, XLG^{+}13].

Hwang:2011:CPG

[HP11] Suk-Geun Hwang and Jin-Woo Park. Characteristic polynomial of a generalized complete product of matrices. Linear Algebra and its Applications, 434(5):1362–1369, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hentzel:2012:SIB

[HP12a] Irvin R. Hentzel and Luiz A. Peresi. Special identities for Bol algebras. Linear Algebra and its Applications, 436 (7):2315–2330, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006549.

Hiai:2012:RMP

[HP12b] Fumio Hiai and Dénes Petz. Riemannian metrics on positive definite matrices related to means. II. Linear Algebra and its Applications, 436 (7):2117–2136, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007269.

Hou:2013:LMP

[HQ13] Jinchuan Hou and Xiaofei Qi. Linear maps preserving separability of pure states. Linear Algebra and its Applications, 439(5):1245–1257, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002504.

Hu:2013:CHP

[HQS13] Shenglong Hu, Liqun Qi, and Jia-Yu Shao. Cored hypergraphs, power hypergraphs and their Laplacian H-eigenvalues. Linear Algebra and its Applications, 439 (10):2980–2998, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005429.

Heinig:2011:FAT

[HR11] Georg Heinig and Karla Rost. Fast algorithms for Toeplitz and Hankel matrices. Linear Algebra and its Applications, 435(1):1–59, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Herrero:2010:NSR

[HRT10] A. Herrero, A. Ramírez, and N. Thome. Nonnegativity, stability, and regularization of discrete-time descriptor systems. Linear Algebra and its Applications, 432(4):837– 846, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Herrero:2013:RBD

[HRT13] Alicia Herrero, Francisco J. Ramírez, and Néstor Thome. Relationships between different sets involving group and Drazin projectors and nonnegativity. Linear Algebra and its Applications, 438(4): 1688–1699, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006124.

Hayden:1999:MCD

[HRW99] Thomas L. Hayden, Robert Reams, and James Wells. Methods for constructing distance matrices and the inverse eigenvalue problem. Linear Algebra and its Applications, 295(1–3):97–112, July 01, 1999. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
cas/tree/store/laa/sub/
1999/295/1-3/6412.pdf;
http://www.elsevier.com/
cgi-bin/cas/tree/store/
laa/cas_sub/browse/browse.
cgi?year=1999&volume=295&
issue=1-3&aid=6412. See

note [JM12a].

Hogben:2010:MGN

[HS10] Leslie Hogben and Bryan Shader. Maximum generic nullity of a graph. Linear Algebra and its Applications, 432(4):857–866, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffman:2012:CET

[HS12a] Thomas R. Hoffman and James P. Solazzo. Complex equiangular tight frames and erasures. Linear Algebra and its Applications, 437 (2):549–558, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000882.

Holubowski:2012:PSG

[HS12b] Waldemar Holubowski and Roksana Slowik. Parabolic subgroups of groups of column-finite infinite matrices. Linear Algebra and its Applications, 437(2):519– 524, July 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001668.

Huhtanen:2012:CGP

[HS12c] Marko Huhtanen and Otto Seiskari. Computational geometry of positive definiteness. Linear Algebra and its Applications, 437(7): 1562–1578, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003333.

Harrell:2014:CSS

[HS14a] Evans M. Harrell II and Joachim Stubbe. Corrigendum to “On sums of graph eigenvalues” [Linear Algebra Appl. 455 (2014) 168–186]. Linear Algebra and its Applications, 458(??): 699–700, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003851.

See [HS14b].

Harrell:2014:SGE

[HS14b] Evans M. Harrell II and Joachim Stubbe. On sums of graph eigenvalues. Linear Algebra and its Applications, 455(??):168–186, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002791.

See corrigendum [HS14a].

Hari:2010:BOJ

[HSS10] Vjeran Hari, Sanja Singer, and Saša Singer. Blockoriented J-Jacobi methods for Hermitian matrices. Linear Algebra and its Applications, 433(8–10):1491–1512, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hari:2014:FBJ

[HSS14] Vjeran Hari, Sanja Singer, and Sasa Singer. Full block J-Jacobi method for Hermitian matrices. Linear Algebra and its Applications, 444(??):1–27, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007623.

Hao:2012:FOC

[HSZ12] Xiaoling Hao, Jiong Sun, and Anton Zettl. Fourth order canonical forms of singular self-adjoint boundary conditions. Linear Algebra and its Applications, 437(3): 899–916, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002534.

Heydari:2010:CLP

[HT10a] Abbas Heydari and Bijan Taeri. On the characteristic and Laplacian polynomials of trees. Linear Algebra and its Applications, 432(2– 3):661–669, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:AIS

[HT10b] Huajun Huang and Tin-Yau Tam. Aluthge iteration in semisimple Lie group. Linear Algebra and its Applications, 432(12):3250–3257, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2014:BOB

[HT14] A. Hadjidimos and M. Tzoumas. On Brauer–Ostrowski and Brualdi sets. Linear Algebra and its Applications, 449 (??):175–193, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000962.

Hashemi:2014:CQI

[HTS14] Behnam Hashemi, Hanieh Tavakolipour, and Mahsa Nasrollahi Shirazi. Comparison of the quasi-inverses of the Kronecker sum and product of matrices over complete commutative dioids with applications. Linear Algebra and its Applications, 448 (??):22–36, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000573.

Huang:2013:GWA

[HTW13] Liang-Hao Huang, Bit-Shun Tam, and Shu-Hui Wu. Graphs whose adjacency matrices have rank equal to the number of distinct nonzero rows. Linear Algebra and its Applications, 438(10): 4008–4040, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004752.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Hu:2010:ECN

[Hu10] Jianhua Hu. Equivalent conditions for noncentral generalized Laplacianness and independence of matrix quadratic forms. Linear Algebra and its Applications, 433(4):796– 809, October 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2010:GDG

[Hua10] Li-Ping Huang. Good distance graphs and the geometry of matrices. Linear Algebra and its Applications, 433 (1):221–232, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:POP

[Hua11a] Qianglian Huang. On perturbations for oblique projection generalized inverses of closed linear operators in Banach spaces. Linear Algebra and its Applications, 434(12): 2468–2474, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:NSP

[Hua11b] Zejun Huang. Nilpotent sign patterns of a given index. Linear Algebra and its Applications, 434(4):880–883, February 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:SRS

[Hua11c] Zejun Huang. On the spectral radius and the spectral norm of Hadamard products of nonnegative matrices. Linear Algebra and its Applications, 434(2):457–462, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2012:EHT

[Hua12a] Li-Ping Huang. Extending Hua’s theorem on the geometry of matrices to Bézout domains. Linear Algebra and its Applications, 436(7): 2446–2473, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006343.

Huang:2012:SSP

[Hua12b] Rong Huang. Sign structure preserves for m-banded factorizations of sign regular matrices. Linear Algebra and its Applications, 436 (7):1990–2000, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006380.

Huang:2013:CEA

[Hua13a] Rong Huang. Componentwise error analysis for the block LU factorization of totally nonnegative matrices. Linear Algebra and its Applications, 439(10):2888–2900, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005181.

Huang:2013:TBF

[Hua13b] Rong Huang. A test and bidiagonal factorization for certain sign regular matrices. Linear Algebra and its Applications, 438(3):1240– 1251, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005964.

Hunter:2010:SSP

[Hun10] Jeffrey J. Hunter. Some stochastic properties of “semimagic” and “magic” Markov chains. Linear Algebra and its Applications, 433(5):893– 907, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hunter:2014:GIM

[Hun14] Jeffrey J. Hunter. Generalized inverses of Markovian kernels in terms of properties of the Markov chain. Linear Algebra and its Applications, 447(??):38–55, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005557.

Hurlimann:2013:GHL

[Hür13] Werner Hürlimann. Generalized Helmert–Ledermann orthogonal matrices and ROM simulation. Linear Algebra and its Applications, 439(7): 1716–1729, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003364.

Han:2011:JDT

[HW11] Dong Han and Feng Wei. Jordan (α, β)-derivations on triangular algebras and related mappings. Linear Algebra and its Applications, 434(1):259– 284, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Harrison:2014:RIF

[HW14a] K. J. Harrison and J. A. Ward. The reflexivity index of a finite distributive lattice of subspaces. Linear Algebra and its Applications, 455(??):73–81, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002523.

Hobart:2014:ABC

[HW14b] Sylvia A. Hobart and Jason Williford. The absolute bound for coherent configurations. Linear Algebra and its Applications, 440 (??):50–60, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006800.

Holguin:2014:SMA

[HW14c] Valeria Aguirre Holguín and Piotr J. Wojciechowski. Signature matrix algebras and bipartite graphs. Linear Algebra and its Applications, 451 (??):97–106, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001608.

Hwang:2011:GCH

[Hwa11] Suk-Geun Hwang. A generalized Cayley–Hamilton theorem for polynomials with matrix coefficients. Linear Algebra and its Applications, 434(2):475–479, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2012:RCH

[Hwa12] Suk-Geun Hwang. The reduced Cayley–Hamilton equation for a pair of commuting matrices. Linear Algebra and its Applications, 436 (7):2078–2083, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100766X.

Hou:2013:CLT

[HWG13] Bo Hou, Lijuan Wang, and Suogang Gao. The classification of Leonard triples that have Bannai/Ito type and odd diameter. Linear Algebra and its Applications, 439 (9):2667–2691, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004552.

Hou:2014:TBL

[HWG14] Bo Hou, Jing Wang, and Suogang Gao. Totally bipartite Leonard pairs and totally bipartite Leonard triples of q-Racah type. Linear Algebra and its Applications, 448 (??):168–204, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000676.

Hsu:2014:IPH

[HWH14] Ming-Hsiu Hsu, Lih-Chung Wang, and Zhen He. Interpolation problems for holomorphic functions. Linear Algebra and its Applications, 452 (??):270–280, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001906.

Huckle:2013:CQT

[HWSH13] T. Huckle, K. Waldherr, and T. Schulte-Herbrüggen. Computations in quantum tensor networks. Linear Algebra and its Applications, 438(2):750–781, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008366.

Hong:2014:SRS

[HY14] Wenxi Hong and Lihua You. Spectral radius and signless Laplacian spectral radius of strongly connected digraphs. Linear Algebra and its Applications, 457(??):93– 113, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002870.

Hashiyama:2014:JBC

[HYF14] Yusuke Hashiyama, Hirokazu Yanagihara, and Yasunori Fujikoshi. Jackknife bias correction of the AIC for selecting variables in canonical correlation analysis under model misspecification. Linear Algebra and its Applications, 455(??):82–106, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002547.

Huang:2010:EWC

[HZ10] Ting-Zhu Huang and Yan Zhu. Estimation of ||A^{−1}||_{∞} for weakly chained diagonally dominant M-matrices. Linear Algebra and its Applications, 432(2–3):670–677, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:PEG

[HZ11a] Qianglian Huang and Wenxiao Zhai. Perturbations and expressions for generalized inverses in Banach spaces and Moore–Penrose inverses in Hilbert spaces of closed linear operators. Linear Algebra and its Applications, 435(1): 117–127, July 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:AMA

[HZ11b] Zejun Huang and Xingzhi Zhan. ACI-matrices all of whose completions have the same rank. Linear Algebra and its Applications, 434(8): 1956–1967, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2012:NDM

[HZ12a] Chengjun Hou and Haiyan Zhang. A note on the diagonal maximality of operator algebras. Linear Algebra and its Applications, 436 (7):2406–2418, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006392.

Huang:2012:PMA

[HZ12b] Zejun Huang and Xingzhi Zhan. Partial matrices all of whose completions have the same spectrum. Linear Algebra and its Applications, 436(9):3061–3064, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006434.

Huang:2014:CBE

[HZ14] Rong Huang and Li Zhu. Componentwise backward error analysis of Neville elimination. Linear Algebra and its Applications, 451 (??):33–48, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400158X.

Huang:2012:PEI

[HZGY12] Qianglian Huang, Lanping Zhu, Wanhui Geng, and Jiena Yu. Perturbation and expression for inner inverses in Banach spaces and its applications. Linear Algebra and its Applications, 436 (9):3721–3735, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000535.

Huang:2012:SPO

[HZJ12] Qianglian Huang, Lanping Zhu, and Yueyu Jiang. On stable perturbations for outer inverses of linear operators in Banach spaces. Linear Algebra and its Applications, 437(7):1942–1954, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003412.

Hou:2010:ND

[HZM10] Chengjun Hou, Wenmin Zhang, and Qing Meng. A note on (α, β)-derivations. Linear Algebra and its Applications, 432(10):2600–2607, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Isa:2013:GOS

[IIK^{+}13] Hiroshi Isa, Masatoshi Ito, Eizaburo Kamei, Hiroaki Tohyama, and Masayuki Watanabe. Generalizations of operator Shannon inequality based on Tsallis and Rényi relative entropies. Linear Algebra and its Applications, 439 (10):3148–3155, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005648.

Ide:2012:EIS

[IJ12] Josh Ide and Lenny Jones. Enumerating invariant subspaces of R^{n}. Linear Algebra and its Applications, 437(7):1845–1853, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002674.

Ikai:2011:TPB

[Ika11] Hisatoshi Ikai. On the theory of Pfaffians based on exponential maps in exterior algebras. Linear Algebra and its Applications, 434(4):1094– 1106, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:2010:NCM

[Ikr10] Khakim D. Ikramov. A note on complex matrices that are unitarily congruent to real matrices. Linear Algebra and its Applications, 433(4):838– 842, October 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ilic:2010:DSD

[Ili10a] Aleksandar Ilić. Distance spectra and distance energy of integral circulant graphs. Linear Algebra and its Applications, 433(5):1005–1014, October 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ilisevic:2010:GBP

[Ili10b] Dijana Ilišević. Generalized bicircular projections on JB*- triples. Linear Algebra and its Applications, 432(5):1267– 1276, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Iannazzo:2011:PMP

[IM11] Bruno Iannazzo and Beatrice Meini. Palindromic matrix polynomials, matrix functions and integral representations. Linear Algebra and its Applications, 434(1):174–184, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

IIRGMR:2010:MRS

[IMA10] IMA-ISU research group on minimum rank. Minimum rank of skew-symmetric matrices described by a graph. Linear Algebra and its Applications, 432(10):2457–2472, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2011:CST

[INT11] Tatsuro Ito, Kazumasa Nomura, and Paul Terwilliger. A classification of sharp tridiagonal pairs. Linear Algebra and its Applications, 435(8):1857– 1884, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oum:2012:RWW

[iO12] Sang il Oum. Rank-width and well-quasi-ordering of skew-symmetric or symmetric matrices. Linear Algebra and its Applications, 436 (7):2008–2036, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100663X.

Interlando:2013:TMB

[IPFD13] J. Carmelo Interlando, Trajano Pires da Nóbrega Neto, André Luiz Flores, and José Othon Dantas Lopes. Two matrix-based lattice construction techniques. Linear Algebra and its Applications, 438(7):3001–3010, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007616.

Ito:2014:EMT

[IRT14] Tatsuro Ito, Hjalmar Rosengren, and Paul Terwilliger. Evaluation modules for the q-tetrahedron algebra. Linear Algebra and its Applications, 451(??):107–168, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Irving:2012:BHR

[Irv12] G. Irving. Banded Householder representation of linear subspaces. Linear Algebra and its Applications, 436 (9):3196–3200, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007294.

Ito:2014:TPT

[IS14] Tatsuro Ito and Jugo Sato. TD-pairs of type II with shape 1, 2, …, 2, 1. Linear Algebra and its Applications, 461(??): 51–91, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400500X.

Ito:2011:GPP

[ISYY11] Masatoshi Ito, Yuki Seo, Takeaki Yamazaki, and Masahiro Yanagida. Geometric properties of positive definite matrices cone with respect to the Thompson metric. Linear Algebra and its Applications, 435(8):2054–2064, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2011:MTS

[IT11] Tatsuro Ito and Paul Terwilliger. Mock tridiagonal systems. Linear Algebra and its Applications, 435(8):1997– 2006, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2013:RMG

[IW13] Naoharu Ito and Harald K. Wimmer. Rank minimization of generalized Sylvester equations over Bezout domains. Linear Algebra and its Applications, 439(3):592– 599, August 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200804X.

Jain:2011:DAT

[Jai11] Tanvi Jain. Derivatives for antisymmetric tensor powers and perturbation bounds. Linear Algebra and its Applications, 435(5):1111–1121, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jarlebring:2012:CFN

[Jar12] Elias Jarlebring. Convergence factors of Newton methods for nonlinear eigenvalue problems. Linear Algebra and its Applications, 436(10): 3943–3953, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Ao:2013:MRS

[jAS13] Ji jun Ao and Jiong Sun. Matrix representations of Sturm–Liouville problems with eigenparameter-dependent boundary conditions. Linear Algebra and its Applications, 438(5):2359–2365, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007288.

Ao:2012:MRF

[jASZ12] Ji jun Ao, Jiong Sun, and Anton Zettl. Matrix representations of fourth order boundary value problems with finite spectrum. Linear Algebra and its Applications, 436(7):2359–2365, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006914.

Jbilou:2010:APK

[Jbi10] K. Jbilou. ADI preconditioned Krylov methods for large Lyapunov matrix equations. Linear Algebra and its Applications, 432(10):2473– 2485, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2013:SGI

[JDY13] Jiashang Jiang, Hua Dai, and Yongxin Yuan. A symmetric generalized inverse eigenvalue problem in structural dynamics model updating. Linear Algebra and its Applications, 439(5):1350–1363, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002838.

Jencova:2013:EGQ

[Jen13] Anna Jencová. Extremal generalized quantum measurements. Linear Algebra and its Applications, 439(12):4070– 4079, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006253.

Ji:2010:MPO

[JG10] Guoxing Ji and Yaling Gao. Maps preserving operator pairs whose products are projections. Linear Algebra and its Applications, 433(7):1348– 1364, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiao:2010:AMD

[JH10] Meiyan Jiao and Jinchuan Hou. Additive maps derivable or Jordan derivable at zero point on nest algebras.

Linear Algebra and its Applications, 432(11):2984–2994, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2012:CCR

[Ji12a] Jun Ji. A condensed Cramer’s rule for the minimum-norm least-squares solution of linear equations. Linear Algebra and its Applications, 437 (9):2173–2178, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004478.

Ji:2012:GJE

[Ji12b] Jun Ji. Gauss–Jordan elimination methods for the Moore–Penrose inverse of a matrix. Linear Algebra and its Applications, 437(7): 1835–1844, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003795.

Jimenez:2010:GSC

[Jim10] M. Elena Domínguez Jiménez. General solution of certain matrix equations arising in filter design applications. Linear Algebra and its Applications, 432(8):2077–2088, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jablonski:2011:HCP

[JJKS11] Zenon J. Jabłoński, Il Bong Jung, Jung Ah Kwak, and Jan Stochel. Hyperexpansive completion problem via alternating sequences; an application to subnormality. Linear Algebra and its Applications, 434(12):2497–2526, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2011:MST

[JK11] Marianne Johnson and Mark Kambites. Multiplicative structure of 2 × 2 tropical matrices. Linear Algebra and its Applications, 435(7):1612– 1625, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Junghanns:2013:CCS

[JK13] Peter Junghanns and Robert Kaiser. Collocation for Cauchy singular integral equations. Linear Algebra and its Applications, 439(3): 729–770, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006751.

Juditsky:2014:UVN

[JKN14] Anatoli Juditsky, Fatma Kilinç Karzan, and Arkadi Nemirovski. On a unified view of nullspace-type conditions for recoveries associated with general sparsity structures. Linear Algebra and its Applications, 441(??): 124–151, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300476X.

Johansson:2013:SFR

[JKV13] Stefan Johansson, Bo Kågström, and Paul Van Dooren. Stratification of full rank polynomial matrices. Linear Algebra and its Applications, 439(4):1062–1090, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008683.

Johnson:2012:SEB

[JL12] Warren P. Johnson and Emil Lalov. Some extensions of Brioschi’s double alternant. Linear Algebra and its Applications, 436(7):1825– 1836, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007099.

Jung:2010:WGM

[JLLY10] Changdo Jung, Hosoo Lee, Yongdo Lim, and Takeaki Yamazaki. Weighted geometric mean of n-operators with n-parameters. Linear Algebra and its Applications, 432 (6):1515–1530, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jean-Louis:2013:SAS

[JLN13] Candice Jean-Louis and Asamoah Nkwanta. Some algebraic structure of the Riordan group. Linear Algebra and its Applications, 438(5):2018– 2035, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007562.

Johnson:2011:CEC

[JLW11] Charles R. Johnson, Brian Lins, and Olivia Walch. The critical exponent for continuous conventional powers of doubly nonnegative matrices. Linear Algebra and its Applications, 435(9):2175–2182, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jaklic:2012:NMC

[JM12a] Gasper Jaklic and Jolanda Modic. A note on “Methods for constructing distance matrices and the inverse eigenvalue problem”. Linear Algebra and its Applications, 437 (11):2781–2792, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005162.

See [HRW99].

Jena:2012:SM

[JM12b] Litismita Jena and Debasisha Mishra. B_{D}-splittings of matrices. Linear Algebra and its Applications, 437(4): 1162–1173, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200273X.

Jafari:2014:EGM

[JM14] M. H. Jafari and A. R. Madadi. On the equality of generalized matrix functions. Linear Algebra and its Applications, 456(??): 16–21, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003035.

Johnson:2010:SNA

[JMP10] C. R. Johnson, C. Marijuán, and M. Pisonero. Spectra that are Newton after extension or translation. Linear Algebra and its Applications, 433(8–10):1623–1641, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2012:SMP

[JMP12] C. R. Johnson, C. Marijuán, and M. Pisonero. Submatrix monotonicity of the Perron root. Linear Algebra and its Applications, 437(10):2429– 2435, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005101.

Johnson:2013:ILC

[JMP13] C. R. Johnson, C. Marijuán, and M. Pisonero. Inequalities for linear combinations of monomials in p-Newton sequences. Linear Algebra and its Applications, 439(7): 2038–2056, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003893.

Jeon:2011:MUB

[JMS11] Jong Sam Jeon, Judith J. Mc-Donald, and Jeffrey L. Stuart. The minimum upper bound on the first ambiguous power of an irreducible, nonpowerful ray or sign pattern. Linear Algebra and its Applications, 435(5):1147–1156, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2011:RMM

[JMW11] Charles R. Johnson, Joshua J. Mollner, and Ashlyn M. Winkler. Reconstructing Matrices from Minors. Linear Algebra and its Applications, 434(7):

1733–1744, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2013:WPL

[JN13] Charles R. Johnson and Sivaram K. Narayan. When the positivity of the leading principal minors implies the positivity of all principal minors of a matrix. Linear Algebra and its Applications, 439 (10):2934–2947, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005168.

Johnson:2013:ICM

[JNS13] Charles R. Johnson, Jonathan Nuckols, and Calum Spicer. The implicit construction of multiplicity lists for classes of trees and verification of some conjectures. Linear Algebra and its Applications, 438(5):1990–2003, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008014.

Johnson:2012:BK

[Joh12] Charles R. Johnson. Bruce Kellogg 1930–2012. Linear Algebra and its Applications, 437(11):2683–2684, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200451X.

Jankovic:2011:FOS

[JP11] Branka Janković and Endre Pap. Full ordering in the Shorrocks mobility sense of the semiring of monotone doubly stochastic matrices. Linear Algebra and its Applications, 435(7):1585–1597, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2013:DED

[JPS13] Charles R. Johnson, Zachary Price, and Ilya M. Spitkovsky. The distribution of eigenvalues of doubly cyclic Z^{+}- matrices. Linear Algebra and its Applications, 439(11): 3576–3580, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005776.

Ji:2011:CLD

[JQ11] Peisheng Ji and Weiqing Qi. Characterizations of Lie derivations of triangular algebras. Linear Algebra and its Applications, 435(5):1137– 1146, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jagels:2011:RRE

[JR11] Carl Jagels and Lothar Reichel. Recursion relations for the extended Krylov subspace method. Linear Algebra and its Applications, 434(7):1716– 1732, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jose:2014:SFH

[JR14] Selby Jose and Ravi A. Rao. Suslin forms and the Hodge star operator. Linear Algebra and its Applications, 452 (??):328–344, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001839.

Jimenez-Rodriguez:2012:NLF

[JRMFSS12] P. Jiménez-Rodríguez, G. A. Muñoz-Fernández, and J. B. Seoane-Sepúlveda. Non-Lipschitz functions with bounded gradient and related problems. Linear Algebra and its Applications, 437(4):1174– 1181, August 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002741.

Johnson:2011:IMI

[JS11] Charles R. Johnson and Ronald L. Smith. Inverse Mmatrices, II. Linear Algebra and its Applications, 435(5): 953–983, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jovanovic:2012:SDG

[JS12] Irena Jovanović and Zoran Stanić. Spectral distances of graphs. Linear Algebra and its Applications, 436(5): 1425–1435, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006021.

Johnson:2013:MPS

[JS13a] Charles R. Johnson and Helena Smigoc. A matricial proof of the symmetric exchange axiom for eigenvalues of principal submatrices of a complex Hermitian matrix. Linear Algebra and its Applications, 438 (9):3719–3722, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000554.

Jose:2013:IPS

[JS13b] Shani Jose and K. C. Sivakumar. On inverse-positivity of sub-direct sums of matrices. Linear Algebra and its Applications, 439(6):1670– 1677, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003121.

Johnston:2013:GMC

[JSS13] Nathaniel Johnston, Lukasz Skowronek, and Erling Størmer. Generation of mapping cones from small sets. Linear Algebra and its Applications, 438(7):3062–3075, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008464.

Jacobs:2011:LET

[JT11a] David P. Jacobs and Vilmar Trevisan. Locating the Eigenvalues of Trees. Linear Algebra and its Applications, 434 (1):81–88, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2011:ESM

[JT11b] Charles R. Johnson and Zheng Tong. Equilibrants, semipositive matrices, calculation and scaling. Linear Algebra and its Applications, 434(7):1638–1647, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jacobs:2013:ELT

[JTT13] David P. Jacobs, Vilmar Trevisan, and Fernando Tura. Eigenvalue location in threshold graphs. Linear Algebra and its Applications, 439 (10):2762–2773, November 15, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300493X.

Jungers:2012:APM

[Jun12] Raphaël M. Jungers. On asymptotic properties of matrix semigroups with an invariant cone. Linear Algebra and its Applications, 437 (5):1205–1214, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002704.

Jungers:2014:FSD

[Jun14]  Marc Jungers.  Feedback

strategies for discrete-time linear-quadratic two-player descriptor games. Linear Algebra and its Applications, 440(??):1–23, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006952.

Jeyaraman:2010:JQS

[JV10] I. Jeyaraman and V. Vetrivel. Jordan quadratic SSM-property and its relation to copositive linear transformations on Euclidean Jordan algebras. Linear Algebra and its Applications, 433(2):390–400, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jeyaraman:2011:LPL

[JV11] I. Jeyaraman and V. Vetrivel. On the Lipschitzian property in linear complementarity problems over symmetric cones. Linear Algebra and its Applications, 435(4):842– 851, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2013:ATT

[JW13] Charles R. Johnson and Zhen Wei. Asymmetric TP and TN completion problems. Linear Algebra and its Applications, 438(5):2127–2135, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007872.

Ji:2011:SRS

[JZ11] You Qing Ji and Yuan Hang Zhang. Stable ranks of split extensions of Banach algebras. Linear Algebra and its Applications, 434(10): 2149–2157, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jin:2014:CMG

[JZ14] Ya-Lei Jin and Xiao-Dong Zhang. Complete multipartite graphs are determined by their distance spectra. Linear Algebra and its Applications, 448(??):285–291, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000585.

Jin:2014:OPF

[JZT14] Xiao-Qing Jin, Zhi Zhao, and Sik-Chung Tam. Optimal preconditioners for functions of matrices. Linear Algebra and its Applications, 457 (??):224–243, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003140.

Jiang:2013:CTU

[JZZ13] Qiaofen Jiang, Huaijie Zhong, and Shifang Zhang. Components of topological uniform descent resolvent set and local spectral theory. Linear Algebra and its Applications, 438(3):1149–1158, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200701X.

Koca:2011:QRS

[KAAK11] Mehmet Koca, Mudhahir Al-Ajmi, and Nazife Ozdes Koca. Quaternionic representation of snub 24-cell and its dual polytope derived from E_{8} root system. Linear Algebra and its Applications, 434(4):977– 989, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-

Kakimura:2010:DPM

[Kak10] Naonori Kakimura. A direct proof for the matrix decomposition of chordal-structured positive semidefinite matrices. Linear Algebra and its Applications, 433(4):819–823, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalinina:2013:SDS

[Kal13a] Elizabeth A. Kalinina. Stability and D-stability of the family of real polynomials. Linear Algebra and its Applications, 438(6):2635–2650, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008208.

Kalita:2013:SIV

[Kal13b] D. Kalita. Spectral integral variation and unicyclic 3colored digraphs with second smallest eigenvalue 1. Linear Algebra and its Applications, 439(1):55–65, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001584.

Kamvar:2010:NAP

[Kam10] Sep Kamvar. Numerical algorithms for personalized search in self-organizing information networks. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14503-2 (hardcover). xiv + 139 pp. LCCN ZA4460 .K36 2010.

Kiani:2011:EUC

[KAMS11] Dariush Kiani, Mohsen Molla Haji Aghaei, Yotsanan Meemark, and Borworn Suntornpoch. Energy of unitary Cayley graphs and gcd-graphs. Linear Algebra and its Applications, 435(6):1336–1343, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karlsson:2011:RCH

[Kar11a]  Bengt R. Karlsson.  H_{2}-

reducible complex Hadamard matrices of order 6. Linear Algebra and its Applications, 434(1):239–246, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karlsson:2011:TPC

[Kar11b] Bengt R. Karlsson. Threeparameter complex Hadamard matrices of order 6. Linear Algebra and its Applications, 434(1):247–258, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kautsky:2012:GPM

[Kau12] Jaroslav Kautsky. Generalized Pascal matrices generate classes closed under multiplication. Linear Algebra and its Applications, 437 (12):2887–2895, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005174.

Kawamura:2012:TBD

[Kaw12] Katsunori Kawamura. Triangular C^{⋆}-bialgebra defined as the direct sum of matrix algebras. Linear Algebra and its Applications, 436 (7):2638–2652, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006355.

Kawamura:2013:MYB

[Kaw13] Katsunori Kawamura. R-Matrices and the Yang– Baxter equation on GNS representations on C^{∗}-bialgebras. Linear Algebra and its Applications, 438(1):573–583, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006040.

Kumbasar:2012:CFF

[KB12] Yalçin Kumbasar and Ayse Hümeyra Bilge. Canonical forms for families of anti-commuting diagonalizable linear operators. Linear Algebra and its Applications, 436(1):79– 85, January 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004836.

Kannan:2014:CSP

[KB14a] M. Rajesh Kannan and R. B. Bapat. Corrigendum to “Generalized principal pivot transform” [Linear Algebra Appl. 454 (2014) 49–56]. Linear Algebra and its Applications, 459(??):620–621, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004789.

See [KB14b].

Kannan:2014:GPP

[KB14b] M. Rajesh Kannan and R. B. Bapat. Generalized principal pivot transform. Linear Algebra and its Applications, 454(??):49–56, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002298.

See corrigendum [KB14a].

Kalaimani:2013:GPA

[KBS13] Rachel K. Kalaimani, Madhu N. Belur, and Sivaramakrishnan Sivasubramanian. Generic pole assignability, structurally constrained controllers and unimodular completion. Linear Algebra and its Applications, 439(12):4003–4022, December 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300623X.

Koliha:2012:GDI

[KCID12] J. J. Koliha, Dragana Cvetković-Ilić, and Chunyuan Deng. Generalized Drazin invertibility of combinations of idempotents. Linear Algebra and its Applications, 437 (9):2317–2324, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004351.

Kuo:2012:QMU

[KD12] Yueh-Cheng Kuo and Biswa N. Datta. Quadratic model updating with no spill-over and incomplete measured data: Existence and computation of solution. Linear Algebra and its Applications, 436 (7):2480–2493, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007622.

Koshlukov:2013:PIA

[KdM13] Plamen Koshlukov and Thiago Castilho de Mello. On the polynomial identities of the algebra M_{11}(E). Linear Algebra and its Applications, 438(11):4469–4482, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000967.

Keckic:2013:CKE

[Kec13] Dragoljub J. Keckić. Cyclic kernels of elementary operators on Banach spaces. Linear Algebra and its Applications, 438(5):2628–2633, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007653.

K:2012:BGG

[KG12a] Shahul Hameed K. and K. A. Germina. Balance in gain graphs — a spectral analysis. Linear Algebra and its Applications, 436(5):1114– 1121, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005179.

Kalofolias:2012:CSN

[KG12b] V. Kalofolias and E. Gallopoulos. Computing symmetric nonnegative rank factorizations. Linear Algebra and its Applications, 436(2): 421–435, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Kueng:2014:RCS

[KG14] R. Kueng and D. Gross. RIPless compressed sensing from anisotropic measurements. Linear Algebra and its Applications, 441(??): 110–123, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002802.

Karcanias:2013:AZP

[KH13] Nicos Karcanias and George Halikias. Approximate zero polynomials of polynomial matrices and linear systems. Linear Algebra and its Applications, 439(4):1091–1103, August 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000141.

Khatami:2013:PNC

[Kha13] Leila Khatami. The poset of the nilpotent commutator of a nilpotent matrix. Linear Algebra and its Applications, 439(12):3763–3776, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006162.

Kang:2014:TAS

[KHG14] Na Kang, Bo Hou, and Suogang Gao. The traces associated with a sharp tridiagonal system. Linear Algebra and its Applications, 457(??):1–11, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002924.

Kapitula:2013:IIM

[KHKT13] Todd Kapitula, Elizabeth Hibma, Hwa-Pyeong Kim, and Jonathan Timkovich. Instability indices for matrix polynomials. Linear Algebra and its Applications, 439 (11):3412–3434, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300548X.

Khosravi:2012:CCP

[Kho12] Maryam Khosravi. A characterization of the class of partial isometries. Linear Algebra and its Applications, 437 (5):1300–1304, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003011.

Karaev:2010:NRN

[KI10] M. T. Karaev and N. Sh. Iskenderov. Numerical range and numerical radius for some operators. Linear Algebra and its Applications, 432(12): 3149–3158, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kian:2014:OJI

[Kia14]  Mohsen Kian.  Operator

Jensen inequality for superquadratic functions. Linear Algebra and its Applications, 456(??):82–87, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951200866X.

Kim:2010:GCI

[Kim10] Hwa Kyung Kim. Generalized competition index of a primitive digraph. Linear Algebra and its Applications, 433(1): 72–79, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2011:BGC

[Kim11a] Hwa Kyung Kim. A bound on the generalized competition index of a primitive matrix using Boolean rank. Linear Algebra and its Applications, 435(9):2166–2174, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2011:LDE

[Kim11b] Ik-Pyo Kim. LDU decomposition of an extension matrix of the Pascal matrix. Linear Algebra and its Applications, 434(10):2187–2196, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2012:TAW

[Kim12] Kijung Kim. Terwilliger algebras of wreath products by quasi-thin schemes. Linear Algebra and its Applications, 437(11):2773–2780, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005095.

Kim:2013:GCI

[Kim13a] Hwa Kyung Kim. Generalized competition index of an irreducible Boolean matrix. Linear Algebra and its Applications, 438(6):2747– 2756, March 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007707.

Kim:2013:SIS

[Kim13b] Hwa Kyung Kim. Scrambling index set of primitive digraphs. Linear Algebra and its Applications, 439(7): 1886–1893, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003649.

Kim:2013:QHF

[Kim13c] Joonhyung Kim. Quaternionic hyperbolic Fuchsian groups. Linear Algebra and its Applications, 438(9): 3610–3617, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000931.

Kim:2013:OEF

[Kim13d] Sejong Kim. Operator entropy and fidelity associated with the geometric mean. Linear Algebra and its Applications, 438(5):2475– 2483, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007720.

Kirkland:2010:FET

[Kir10] Steve Kirkland. Fastest expected time to mixing for a Markov chain on a directed graph. Linear Algebra and its Applications, 433 (11–12):1988–1996, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2014:MCE

[Kir14] Steve Kirkland. The minimum coefficient of ergodicity for a Markov chain with a given directed graph. Linear Algebra and its Applications, 447(??):139–154, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004357.

Kisi:2015:CTL

[Kis15] Emre Kisi. Corrigendum to “Tripotency of a linear combination of two involutory matrices and a tripotent matrix that mutually commute” [Linear Algebra Appl. 437

(9) (2012) 2091–2109]. Linear Algebra and its Applications, 477(??):211–212, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379515001536.

See [XX12].

Ko:2013:SSC

[KJK13] Eungil Ko, Sungeun Jung, and Yoenha Kim. On subscalarity of some 2 × 2 class A operator matrices. Linear Algebra and its Applications, 438(3):1322–1338, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006611.

Kim:2010:NCS

[KK10] Dong-Soo Kim and Young Ho Kim. New characterizations of spheres, cylinders and W -curves. Linear Algebra and its Applications, 432(11): 3002–3006, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2012:SCS

[KK12] Dong-Soo Kim and Young Ho Kim. Some characterizations of spheres and elliptic paraboloids. Linear Algebra and its Applications, 437(1):113–120, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001437.

Kim:2013:SCS

[KK13] Dong-Soo Kim and Young Ho Kim. Some characterizations of spheres and elliptic paraboloids II. Linear Algebra and its Applications, 438 (3):1356–1364, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006489.

Kim:2014:CCH

[KK14] Joonhyung Kim and Sungwoon Kim. A characterization of complex hyperbolic Kleinian groups in dimension 3 with trace fields contained in R. Linear Algebra and its Applications, 455(??):107–114, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400281X.

Knap:2011:RHS

[KKB11] M. J. Knap, L. H. Keel, and S. P. Bhattacharyya. Robust Hurwitz stability via signdefinite decomposition. Linear Algebra and its Applications, 434(7):1663–1676, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2010:WCD

[KKL10] Dongseok Kim, Young Soo Kwon, and Jaeun Lee. The weighted complexity and the determinant functions of graphs. Linear Algebra and its Applications, 433(2):348– 355, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2013:ESU

[KKL13a] Hyun Kwang Kim, Denis S. Krotov, and Joon Yop Lee. Erratum to “Matrices uniquely determined by their lonesums” [Linear Algebra Appl. 438(7) (2013) 3107– 3123]. Linear Algebra and its Applications, 439(1):289, July 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002127.

See [KKL13b].

Kim:2013:MUD

[KKL13b] Hyun Kwang Kim, Denis S. Krotov, and Joon Yop Lee. Matrices uniquely determined by their lonesums. Linear Algebra and its Applications, 438(7):3107–3123, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008348.

See erratum [KKL13a].

Krahmer:2013:SSP

[KKL13c] Felix Krahmer, Gitta Kutyniok, and Jakob Lemvig. Sparsity and spectral properties of dual frames. Linear Algebra and its Applications, 439(4):982–998, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007264.

Kim:2014:FID

[KKL14] Sejong Kim, Sungwoon Kim, and Hosoo Lee. Factorizations of invertible density matrices. Linear Algebra and its Applications, 463 (??):190–204, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005990.

Kim:2014:DSS

[KKLY14] H. K. Kim, K. H. Kwon, L. L. Littlejohn, and G. J. Yoon. Diagonalizability and symmetrizability of Sobolevtype bilinear forms: a combinatorial approach. Linear Algebra and its Applications, 460(??):111–124, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004881.

Kucera:2013:LSG

[KKM13] R. Kucera, T. Kozubek, and A. Markopoulos. On largescale generalized inverses in solving two-by-two block linear systems. Linear Algebra and its Applications, 438 (7):3011–3029, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007410.

Klotz:2011:CMR

[KKR11] Lutz Klotz, Peter Kunkel, and David Rudolph. Convergence in measure with respect to a matrix-valued measure and some matrix completion problems. Linear Algebra and its Applications, 434(4):990– 999, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-

Kari:2012:CWF

[KKS12] Jarkko Kari, Alexandr Kazda, and Paula Steinby. On continuous weighted finite automata. Linear Algebra and its Applications, 436(7): 1791–1824, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007282.

Kang:2012:CHT

[KL12] Dong-O Kang and Woo Young Lee. A criterion on the hyponormality of Toeplitz operators with polynomial symbols via Schur numbers. Linear Algebra and its Applications, 436(9):3608–3617, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200002X.

Kaashoek:2013:CMP

[KL13a] M. A. Kaashoek and L. Lerer. On a class of matrix polynomial equations. Linear Algebra and its Applications, 439 (3):613–620, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006441.

Kum:2013:RAS

[KL13b] Sangho Kum and Yongdo Lim. The resolvent average on symmetric cones. Linear Algebra and its Applications, 438(3):1159–1169, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006696.

Klasa:2010:FPD

[Kla10] J. Klasa. A few pedagogical designs in linear algebra with Cabri and Maple. Linear Algebra and its Applications, 432(8):2100–2111, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2011:MGM

[KLL11] Sejong Kim, Jimmie Lawson, and Yongdo Lim. The matrix geometric mean of parameterized, weighted arithmetic and harmonic means. Linear Algebra and its Applications, 435(9):2114–2131, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krnic:2012:MJO

[KLP12] Mario Krnić, Neda Lovricević, and Josip Pecarić. Multidimensional Jensen’s operator on a Hilbert space and applications. Linear Algebra and its Applications, 436 (7):2583–2596, April 1, 2012.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100646X.

Kribs:2013:UAS

[KLP13] David W. Kribs, Jeremy Levick, and Rajesh Pereira. Unital affine semigroups. Linear Algebra and its Applications, 439(7):2106–2113, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003935.

Khurana:2012:QTD

[KLS12] Dinesh Khurana, T. Y. Lam, and Noam Shomron. A quantum-trace determinantal formula for matrix commutators, and applications. Linear Algebra and its Applications, 436(7):2380–2397, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005222.

Kosan:2010:DRI

[KLZ10] M. Tamer Koşan, Tsiu-Kwen Lee, and Yiqiang Zhou. Derivations and right ideals of algebras. Linear Algebra and its Applications, 432(11): 2773–2781, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kosan:2014:BFM

[KLZ14a] M. Tamer Kosan, Tsiu-Kwen Lee, and Yiqiang Zhou. Bilinear forms on matrix algebras vanishing on zero products of xy and yx. Linear Algebra and its Applications, 453(??):110–124, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001967.

Kosan:2014:WEM

[KLZ14b] M. Tamer Kosan, Tsiu-Kwen Lee, and Yiqiang Zhou. When is every matrix over a division ring a sum of an idempotent and a nilpotent? Linear Algebra and its Applications, 450(??):7–12, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001268.

Kilmer:2011:FST

[KM11] Misha E. Kilmer and Carla D. Martin. Factorization strategies for third-order tensors. Linear Algebra and its Applications, 435(3):641–658, August 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krakovski:2012:SCG

[KM12] Roi Krakovski and Bojan Mohar. Spectrum of Cayley graphs on the symmetric group generated by transpositions. Linear Algebra and its Applications, 437(3): 1033–1039, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002662.

Katsouleas:2013:IHN

[KM13a] Georgios Katsouleas and John Maroulas. The imbeddability for Hermitian and normal matrices. Linear Algebra and its Applications, 439 (3):552–564, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002085.

Kurata:2013:CMB

[KM13b] Hiroshi Kurata and Shun Matsuura. Characterization of multispherical and block structures of Euclidean distance matrices. Linear Algebra and its Applications, 439 (10):3177–3183, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005636.

Klein:2014:AEF

[KM14] André Klein and Guy Mélard. An algorithm for the exact Fisher information matrix of vector ARMAX time series. Linear Algebra and its Applications, 446(??):1–24, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008392.

Karapiperi:2012:EAE

[KMNS12] Anna Karapiperi, Marilena Mitrouli, Michael G. Neubauer, and Jennifer Seberry. An eigenvalue approach evaluating minors for weighing matrices W (n, n − 1). Linear Algebra and its Applications, 436(7):2054–2066, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007270.

Khalil:2013:SST

[KMS13] Houssam Khalil, Bernard Mourrain, and Michelle Schatzman. Superfast solution of Toeplitz systems based on syzygy reduction. Linear Algebra and its Applications, 438(9):3563–3575, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000773.

Kaur:2014:FRH

[KMSC14] Rupinderjit Kaur, Mohammad Sal Moslehian, Mandeep Singh, and Cristian Conde. Further refinements of the Heinz inequality. Linear Algebra and its Applications, 447 (??):26–37, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000748.

Knuppel:2010:CGO

[KN10] Frieder Knüppel and Klaus Nielsen. Commutator groups of orthogonal groups over the reals with emphasis on Lorentz groups. Linear Algebra and its Applications, 433 (11–12):2111–2121, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knuppel:2013:CIP

[KN13a] Frieder Knüppel and Klaus Nielsen. Covering GL(V) and products of blockcyclic matrices. Linear Algebra and its Applications, 438(5): 2508–2519, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007690.

Knuppel:2013:CSL

[KN13b] Frieder Knüppel and Klaus Nielsen. Covering singular linear semi-groups. Linear Algebra and its Applications, 438(7):3039–3053, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008440.

Knuppel:2013:SPB

[KN13c] Frieder Knüppel and Klaus Nielsen. A short proof of Botha’s theorem on products of idempotent linear mappings. Linear Algebra and its Applications, 438(5): 2520–2522, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007306.

Knill:2014:CBP

[Kni14] Oliver Knill. Cauchy–Binet for pseudo-determinants. Linear Algebra and its Applications, 459(??):522–547, October 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004625.

Katz:2014:CTH

[KNS14] Ricardo D. Katz, Viorel Nitica, and Sergei Sergeev. Characterization of tropical hemispaces by (P, R)- decompositions. Linear Algebra and its Applications, 440(??):131–163, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006630.

Khudoyberdiyev:2013:CAL

[KO13] A. Kh. Khudoyberdiyev and B. A. Omirov. The classification of algebras of level one. Linear Algebra and its Applications, 439(11):3460– 3463, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005764.

Kohler:2014:CSD

[Köh14] M. Köhler. On the closest stable descriptor system in the respective spaces RH_{2} and RH_{∞}. Linear Algebra and its Applications, 443(??): 34–49, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300709X.

Kirkland:2011:ACA

[KOJ11] Steve Kirkland, Carla Silva Oliveira, and Claudia Marcela Justel. On algebraic connectivity augmentation. Linear Algebra and its Applications, 435(10):2347–2356, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koledin:2013:RGG

[Kol13] Tamara Koledin. Regular graphs with girth at least 5 and small second largest eigenvalue. Linear Algebra and its Applications, 439 (5):1229–1244, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002498.

Konstantopoulos:2013:MAP

[Kon13] Takis Konstantopoulos. A multilinear algebra proof of the Cauchy–Binet formula and a multilinear version of Parseval’s identity. Linear Algebra and its Applications, 439(9):2651–2658, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004606.

Kutyniok:2013:SF

[KOPT13] Gitta Kutyniok, Kasso A. Okoudjou, Friedrich Philipp, and Elizabeth K. Tuley. Scalable frames. Linear Algebra and its Applications, 438(5): 2225–2238, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007768.

Kandolf:2014:RBE

[KOR14] Peter Kandolf, Alexander Ostermann, and Stefan Rainer. A residual based error estimate for Leja interpolation of matrix functions. Linear Algebra and its Applications, 456(??):157–173, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Kozyakin:2010:ELC

[Koz10] Victor Kozyakin. An explicit Lipschitz constant for the joint spectral radius. Linear Algebra and its Applications, 433(1):12–18, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kozyakin:2014:BWF

[Koz14a] Victor Kozyakin. The Berger– Wang formula for the Markovian joint spectral radius. Linear Algebra and its Applications, 448(??):315–328, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000408.

Kozyakin:2014:MPC

[Koz14b]  Victor Kozyakin.  Matrix

products with constraints on the sliding block relative frequencies of different factors. Linear Algebra and its Applications, 457(??):244– 260, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002961.

Kim:2012:BGC

[KP12] Hwa Kyung Kim and Sung Gi Park. A bound of generalized competition index of a primitive digraph. Linear Algebra and its Applications, 436(1):86–98, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004903.

Kanan:2013:NCB

[KP13] Asmaa M. Kanan and Zoran Z. Petrović. Note on cardinality of bases in semilinear spaces over zerosumfree semirings. Linear Algebra and its Applications, 439 (10):2795–2799, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004965.

Kalita:2014:REP

[KP14a] Debajit Kalita and Sukanta Pati. A reciprocal eigenvalue property for unicyclic weighted directed graphs with weights from {±1, ±i}. Linear Algebra and its Applications, 449(??):417–434, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001037.

Krnic:2014:MAW

[KP14b] Mario Krnić and Josip Pecarić. More accurate weak majorization relations for the Jensen and some related inequalities. Linear Algebra and its Applications, 458(??): 573–588, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004066.

Kutyniok:2014:PSI

[KPRT14] Gitta Kutyniok, Allan Pinkus, Holger Rauhut, and Vladimir Temlyakov. Preface to the Special Issue on Sparse Approximate Solution of Linear Systems. Linear Algebra and its Applications, 441 (??):1–3, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004187.

Koolen:2011:IIS

[KPY11] Jack H. Koolen, Jongyook Park, and Hyonju Yu. An inequality involving the second largest and smallest eigenvalue of a distance-regular graph. Linear Algebra and its Applications, 434(12):2404– 2412, June 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kulkarni:2010:FGB

[KR10] S. H. Kulkarni and G. Ramesh. A formula for gap between two closed operators. Linear Algebra and its Applications, 432(11):3012–3017, June 1, 2010. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-

1856 (electronic).

Khare:2012:SMD

[KR12] Kshitij Khare and Bala Rajaratnam. Sparse matrix decompositions and graph characterizations. Linear Algebra and its Applications, 437 (3):932–947, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002583.

Krakovski:2012:PAV

[Kra12] Roi Krakovski. Prescribed absolute values, character sums and spectrum integrality. Linear Algebra and its Applications, 437(8):2077– 2089, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004120.

Krakovski:2013:ABM

[Kra13] Roi Krakovski. An analogue of Bridges and Mena’s theorem for local fields and a local-global principle. Linear Algebra and its Applications, 438(1):342–347, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006027.

Khudoyberdiyev:2014:CDS

[KRH14] A. Kh. Khudoyberdiyev, I. S. Rakhimov, and Sh. K. Said Husain. On classification of 5-dimensional solvable Leibniz algebras. Linear Algebra and its Applications, 457 (??):428–454, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003255.

Kazeev:2013:LRT

[KRS13] Vladimir Kazeev, Oleg Reichmann, and Christoph Schwab. Low-rank tensor structure of linear diffusion operators in the TT and QTT formats. Linear Algebra and its Applications, 438(11): 4204–4221, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300058X.

Kempker:2012:COC

[KRvS12] Pia L. Kempker, André C. M. Ran, and Jan H. van Schuppen. Controllability and observability of coordinated linear systems. Linear Algebra and its Applications, 437 (1):121–167, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000924.

Kempker:2014:CMC

[KRvS14] Pia L. Kempker, André C. M. Ran, and Jan H. van Schuppen. Construction and minimality of coordinated linear systems. Linear Algebra and its Applications, 452 (??):202–236, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001840.

Klein:2010:TSM

[KS10] André Klein and Peter Spreij. Tensor Sylvester matrices and the Fisher information matrix of VARMAX processes. Linear Algebra and its Applications, 432(8):1975–1989, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuijper:2011:MGB

[KS11a] M. Kuijper and K. Schindelar. Minimal Gröbner bases and the predictable leading monomial property. Linear Algebra and its Applications, 434 (1):104–116, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kures:2011:RMO

[KS11b] Miroslav Kureš and Ladislav Skula. Reduction of matrices over orders of imaginary quadratic fields. Linear Algebra and its Applications, 435(8):1903–1919, October 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kushel:2011:GEO

[KS11c] O. Y. Kushel and P. Sharma. Generalized even and odd totally positive matrices. Linear Algebra and its Applications, 435(11):2722–2730, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kannan:2012:MPI

[KS12a] M. Rajesh Kannan and K. C. Sivakumar. Moore–Penrose inverse positivity of interval matrices. Linear Algebra and its Applications, 436(3):571–578, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005246.

Klein:2012:TSD

[KS12b] André Klein and Peter Spreij. Transformed statistical distance measures and the Fisher information matrix. Linear Algebra and its Applications, 437(2):692–712, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200184X.

Kim:2013:UML

[KS13a] In-Jae Kim and Bryan L. Shader. Unordered multiplicity lists of a class of binary trees. Linear Algebra and its Applications, 438(10): 3781–3788, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005180.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Koledin:2013:RBG

[KS13b] Tamara Koledin and Zoran Stanić. Regular bipartite graphs with three distinct non-negative eigenvalues. Linear Algebra and its Applications, 438(8):3336– 3349, April 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000360.

Kapil:2014:CMO

[KS14a] Yogesh Kapil and Mandeep Singh. Contractive maps on operator ideals and norm inequalities. Linear Algebra and its Applications, 459(??): 475–492, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004650.

Klein:2014:BHG

[KS14b] André Klein and Peter Spreij. A block Hankel generalized confluent Vandermonde matrix. Linear Algebra and its Applications, 455(??):32– 72, August 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002808.

Koledin:2014:RBR

[KS14c] Tamara Koledin and Zoran Stanić. Reflexive bipartite regular graphs. Linear Algebra and its Applications, 442(??):145–155, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004710.

Kaur:2011:GMV

[KSA11] Rupinderjit Kaur, Mandeep Singh, and Jaspal Singh Aujla. Generalized matrix version of reverse Hölder inequality. Linear Algebra and its Applications, 434(3):636– 640, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaur:2012:GDI

[KSAM12] Rupinderjit Kaur, Mandeep Singh, Jaspal Singh Aujla, and M. S. Moslehian. A general double inequality related to operator means and positive linear maps. Linear Algebra and its Applications, 437(3):1016–1024, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002601.

Kang:2012:LPT

[KSB12] Kyung-Tae Kang, Seok-Zun Song, and LeRoy B. Beasley. Linear preservers of term ranks of matrices over semirings. Linear Algebra and its Applications, 436(7):1850– 1862, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006306.

Kim:2012:EDD

[KSH12] Byeong Moon Kim, Byung Chul Song, and Woonjae Hwang. The exponent of a digraph and the diameter of its multiple direct product. Linear Algebra and its Applications, 437(10):2601–2612, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004922.

Katz:2012:CMM

[KSS12] Ricardo D. Katz, Hans Schneider, and Sergeĭ Sergeev. On commuting matrices in max algebra and in classical nonnegative algebra. Linear Algebra and its Applications, 436(2):276–292, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510004350.

Kurata:2010:MED

[KT10] Hiroshi Kurata and Pablo Tarazaga. Multispherical Euclidean distance matrices. Linear Algebra and its Applications, 433(3):534–546, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kurata:2012:MEE

[KT12] Hiroshi Kurata and Pablo Tarazaga. Majorization for the eigenvalues of Euclidean distance matrices. Linear Algebra and its Applications, 436(5):1473–1481, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006057.

Kubrusly:2013:RLT

[Kub13]  C. S. Kubrusly.  Regu-

lar lattices of tensor products. Linear Algebra and its Applications, 438(1):428– 435, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005393.

Kumagai:2011:CEM

[Kum11] Wataru Kumagai. A characterization of extended monotone metrics. Linear Algebra and its Applications, 434 (1):224–231, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kushel:2012:CTG

[Kus12] O. Y. Kushel. Cone-theoretic generalization of total positivity. Linear Algebra and its Applications, 436(3):537– 560, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005143.

Kushel:2013:IPE

[Kus13] O. Y. Kushel. Interlacing properties of the eigenvalues of some matrix classes. Linear Algebra and its Applications, 439(8):2293–2308, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300431X.

Kuzman:2010:ATE

[Kuz10]  Boštjan Kuzman.  Arc-

transitive elementary abelian covers of the complete graph K_{5}. Linear Algebra and its Applications, 433(11–12):

1909–1921, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koyuncu:2012:INT

[KW12] Selcuk Koyuncu and Hugo J. Woerdeman. The inverse of nonsymmetric two-level Toeplitz operator matrices. Linear Algebra and its Applications, 437(9):2142–2158, November 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003370.

Ku:2013:GES

[KW13] Cheng Yeaw Ku and Kok Bin Wong. Gallai–Edmonds structure theorem for weighted matching polynomial. Linear Algebra and its Applications, 439(11):3387–3411, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005806.

Koolen:2011:DRG

[KY11] Jack H. Koolen and Hyonju Yu. The distance-regular graphs such that all of its second largest local eigenvalues are at most one. Linear Algebra and its Applications, 435(10):2507–2519, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaltofen:2013:FFM

[KY13] Erich Kaltofen and George Yuhasz. A fraction free Matrix Berlekamp/Massey algorithm. Linear Algebra and its Applications, 439(9):2515– 2526, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004151.

Kim:2014:SPT

[KY14] Jaewoong Kim and Jasang Yoon. Schur product techniques for the subnormality of commuting 2-variable weighted shifts. Linear Algebra and its Applications, 453 (??):174–191, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002274.

Kyrchei:2013:ERF

[Kyr13] Ivan Kyrchei. Explicit representation formulas for the minimum norm least squares solutions of some quaternion matrix equations. Linear Algebra and its Applications, 438(1):136–152, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200612X.

Koshlukov:2010:IIG

[KZ10] Plamen Koshlukov and Mikhail Zaicev. Identities and isomorphisms of graded simple algebras. Linear Algebra and its Applications, 432(12):3141– 3148, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kong:2011:SJC

[KZ11] Qingkai Kong and Anton Zettl. The study of Jacobi and cyclic Jacobi matrix eigenvalue problems using Sturm– Liouville theory. Linear Algebra and its Applications, 434 (7):1648–1655, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

LaGuardia:2014:NMC

[La 14] Giuliano G. La Guardia. On negacyclic MDS-convolutional codes. Linear Algebra and its Applications, 448(??):85– 96, May 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000627.

Lacruz:2013:HLI

[Lac13]  Miguel Lacruz.  Hardy–

Littlewood inequalities for norms of positive operators on sequence spaces. Linear Algebra and its Applications, 438(1):153–156, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006052.

Laffey:2012:CVB

[Laf12] Thomas J. Laffey. A constructive version of the Boyle– Handelman theorem on the spectra of nonnegative matrices. Linear Algebra and its Applications, 436(6):1701– 1709, March 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006562.

LaGrange:2012:BRR

[LaG12] John D. LaGrange. Boolean rings and reciprocal eigenvalue properties. Linear Algebra and its Applications, 436(7):1863–1871, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007543.

LaGrange:2013:CDF

[LaG13] John D. LaGrange. A combinatorial development of Fibonacci numbers in graph spectra. Linear Algebra and its Applications, 438(11): 4335–4347, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Lakatos:2010:SND

[Lak10a] Piroska Lakatos. Salem numbers defined by Coxeter transformation. Linear Algebra and its Applications, 432(1):144– 154, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lakos:2010:FLS

[Lak10b] Gyula Lakos. Factorization of Laurent series over commutative rings. Linear Algebra and its Applications, 432(1):338– 346, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ledermann:2011:ROM

[LAL11] Walter Ledermann, Carol Alexander, and Daniel Ledermann. Random orthogonal matrix simulation. Linear Algebra and its Applications, 434(6):1444–1467, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Langi:2010:DOM

[Lán10a] Zsolt Lángi. On diagonalizable operators in Minkowski spaces with the Lipschitz property. Linear Algebra and its Applications, 433(11– 12):2161–2167, December 30, 2010. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Lanski:2010:DCI

[Lan10b] Charles Lanski. Differential commutator identities. Linear Algebra and its Applications, 433(6):1212–1223, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lancaster:2013:SLG

[Lan13] Peter Lancaster. Stability of linear gyroscopic systems: a review. Linear Algebra and its Applications, 439(3): 686–706, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300013X.

Langenau:2014:ASI

[Lan14] Holger Langenau. Asymptotically sharp inequalities for polynomials involving mixed Laguerre norms. Linear Algebra and its Applications, 458(??):116–127, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003784.

Lavallee:2010:CPN

[Lav10] Sylvain Lavallée. Characteristic polynomials of nonnegative real square matrices and generalized clique polynomials. Linear Algebra and its Applications, 433(1):118–126, July 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2014:TPM

[LB14] Geunseop Lee and Jesse L. Barlow. Two projection methods for Regularized Total Least Squares approximation. Linear Algebra and its Applications, 461(??): 18–41, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004960.

Li:2012:SQA

[LBLS12] Yung-Ta Li, Zhaojun Bai, Wen-Wei Lin, and Yangfeng Su. A Structured Quasi-Arnoldi procedure for model order reduction of secondorder systems. Linear Algebra and its Applications, 436(8):2780–2794, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005350.

Lin:2010:NRE

[LC10] Matthew M. Lin and Moody T. Chu. On the nonnegative rank of Euclidean distance matrices. Linear Algebra and its Applications, 433(3):681– 689, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2013:HNO

[LCM13] Yan Liu, Liangyun Chen, and Yao Ma. Hom–Nijienhuis operators and T^{∗}-extensions of hom–Lie superalgebras. Linear Algebra and its Applications, 439(7):2131–2144, October 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003947.

Li:2011:PGE

[LCwCL11] Tiexiang Li, Chun-Yueh Chiang, Eric King wah Chu, and Wen-Wei Lin. The palindromic generalized eigenvalue problem A^{∗}x = λAx: Numerical solution and applications. Linear Algebra and its Applications, 434(11):2269–2284, June 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2010:SNB

[LCZ10] Qingbing Liu, Guoliang Chen, and Linlin Zhao. Some new bounds on the spectral radius of matrices. Linear Algebra and its Applications, 432(4): 936–948, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2011:PCM

[LD11] Arthur Lim and Jialing Dai. On product of companion matrices. Linear Algebra and its Applications, 435(11): 2921–2935, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2012:PCD

[LD12] Jiantao Li and Xiankun Du. Pairwise commuting derivations of polynomial rings. Linear Algebra and its Applications, 436(7):2375–2379, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007002.

Lorenzo:2011:ADS

[LdlP11] E. Lorenzo and M. J. de la Puente. An algorithm to describe the solution set of any tropical linear system A⊙x = B ⊙ x. Linear Algebra and its Applications, 435(4):884– 901, August 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lima:2013:FCS

[LdS13] J. B. Lima and R. M. Campello de Souza. Fractional cosine and sine transforms over finite fields. Linear Algebra and its Applications, 438(8): 3217–3230, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000086.

Lima:2011:EFF

[LdSP11] J. B. Lima, R. M. Campello de Souza, and D. Panario. The eigenstructure of finite field trigonometric transforms. Linear Algebra and its Applications, 435(8):1956– 1971, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2010:RTI

[Lee10] Eun-Young Lee. Rotfel’d type inequalities for norms. Linear Algebra and its Applications, 433(3):580–584, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2011:ERT

[Lee11] Eun-Young Lee. Extension of Rotfel’d Theorem. Linear Algebra and its Applications, 435(4):735–741, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2013:DNR

[Lee13a] Hsin-Yi Lee. Diagonals and numerical ranges of direct sums of matrices. Linear Algebra and its Applications, 439 (9):2584–2597, November 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004709.

Lee:2013:PDR

[Lee13b] Jae-Ho Lee. Q-polynomial distance-regular graphs and a double affine Hecke algebra of rank one. Linear Algebra and its Applications, 439 (10):3184–3240, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005144.

Lee:2013:MMU

[Lee13c] Joon Yop Lee. Multidimensional matrices uniquely recovered by their lines. Linear Algebra and its Applications, 438(5):2430–2437, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007367.

Li:2012:NLG

[LFS12] Hong-Hai Li, Yi-Zheng Fan, and Li Su. On the nullity of the line graph of unicyclic graph with depth one. Linear Algebra and its Applications, 437(8):2038–2055, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004144.

Li:2013:SLL

[LGS13] Jianxi Li, Ji-Ming Guo, and Wai Chee Shiu. On the second largest Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 438(5):2438–2446, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200777X.

Li:2014:SCT

[LGSC14] Jianxi Li, Ji-Ming Guo, Wai Chee Shiu, and An Chang. Six classes of trees with largest normalized algebraic connectivity. Linear Algebra and its Applications, 452 (??):318–327, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001748.

Liu:2014:OPC

[LGZ14] Lele Liu, Zhenlin Gao, and Weiping Zhao. On an open problem concerning regular magic squares of odd order. Linear Algebra and its Applications, 459(??):1– 12, October 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004157.

Huang:2010:BDP

[lH10a] Wen ling Huang. Bounded distance preserving surjections in the geometry of matrices. Linear Algebra and its Applications, 433(11– 12):1973–1987, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2010:SCD

[LH10b] Jianzhou Liu and Zejun Huang. The Schur complements of γ-diagonally and product γ-diagonally dominant matrix and their disc separation. Linear Algebra and its Applications, 432(4): 1090–1104, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lopez:2010:ODF

[LH10c] Jerry Lopez and Deguang Han. Optimal dual frames for erasures. Linear Algebra and its Applications, 432(1):471– 482, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leng:2011:ODF

[LH11a] Jinsong Leng and Deguang Han. Optimal dual frames for erasures II. Linear Algebra and its Applications, 435 (6):1464–1472, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2011:BDP

[lH11b] Wen ling Huang. Bounded distance preserving surjections in the projective geometry of matrices. Linear Algebra and its Applications, 435(1):175–185, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2013:LSC

[LH13] Fenjin Liu and Qiongxiang Huang. Laplacian spectral characterization of 3rose graphs. Linear Algebra and its Applications, 439 (10):2914–2920, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004928.

Li:2010:TEC

[LHG10] Zengti Li, Tayuan Huang, and Suogang Gao. Two error-correcting pooling designs from symplectic spaces over a finite field. Linear Algebra and its Applications, 433(6):1138–1147, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2014:NED

[LHGL14] Qinghai Liu, Yanmei Hong, Xiaofeng Gu, and Hong-Jian Lai. Note on edgedisjoint spanning trees and eigenvalues. Linear Algebra and its Applications, 458(??): 128–133, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003644.

Leng:2013:PMO

[LHH13] Jinsong Leng, Deguang Han, and Tingzhu Huang. Probability modelled optimal frames for erasures. Linear Algebra and its Applications, 438(11):4222–4236, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000827.

Liau:2012:NSC

[LHL12] Pao-Kuei Liau, Wei-Lu Huang, and Cheng-Kai Liu. Nonlinear strong commutativity preserving maps on skew elements of prime rings with involution. Linear Algebra and its Applications, 436(9): 3099–3108, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007087.

Liu:2013:NSC

[LHL13] Fenjin Liu, Qiongxiang Huang, and Hong-Jian Lai. Note on the spectral characterization of some cubic graphs with maximum number of triangles. Linear Algebra and its Applications, 438(3): 1393–1397, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006805.

Liu:2014:EDS

[LHL14] Qinghai Liu, Yanmei Hong, and Hong-Jian Lai. Edgedisjoint spanning trees and eigenvalues. Linear Algebra and its Applications, 444 (??):146–151, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007775.

Li:2010:IGT

[LHLL10] Hou-Biao Li, Ting-Zhu Huang, Xing-Ping Liu, and Hong Li. On the inverses of general tridiagonal matrices. Linear Algebra and its Applications, 433(5):965–983, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2011:SCG

[LHW11] Fenjin Liu, Qiongxiang Huang, and Jianfeng Wang. Spectral characterization of graphs whose second largest eigenvalue is less than 1. Linear Algebra and its Applications, 434(2):381–393, January 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2012:SCG

[LHWL12] Fenjin Liu, Qiongxiang Huang, Jianfeng Wang, and Qinghai Liu. The spectral characterization of ∞-graphs. Linear Algebra and its Applications, 437(7):1482–1502, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200287X.

Lin:2013:DSG

[LHWS13] Huiqiu Lin, Yuan Hong, Jianfeng Wang, and Jinlong Shu. On the distance spectrum of graphs. Linear Algebra and its Applications, 439(6):1662– 1669, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002814.

Liu:2011:TSC

[LHZH11] Jianzhou Liu, Zhuohong Huang, Li Zhu, and Zejun Huang. Theorems on Schur complement of block diagonally dominant matrices and their application in reducing the order for the solution of large scale linear systems. Linear Algebra and its Applications, 435(12):3085–3100, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2010:NUP

[Li10] Chi-Kwong Li. A note on the unitary part of a contraction. Linear Algebra and its Applications, 433(8–10):1533– 1535, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2012:PPL

[Li12]  Hua-Chieh Li.  Periodic

points of a linear transformation. Linear Algebra and its Applications, 437(10):2489– 2497, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004703.

Lim:2010:SGP

[Lim10] Ming-Huat Lim. Surjections on Grassmannians preserving pairs of elements with bounded distance. Linear Algebra and its Applications, 432(7):1703–1707, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2011:SCA

[Lim11a] Yongdo Lim. Stopping criteria for the Ando–Li–Mathias and Bini–Meini–Poloni geometric means. Linear Algebra and its Applications, 434(8):1884– 1892, April 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2011:SSP

[Lim11b]  Yongdo Lim.  Symmetric

Γ-submanifolds of positive definite matrices and the Sylvester equation XM = NX. Linear Algebra and its Applications, 435(9):2285– 2295, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2012:FGM

[Lim12] Yongdo Lim. Factorizations and geometric means of positive definite matrices. Linear Algebra and its Applications, 437(9):2159–2172, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004430.

Lim:2013:GMS

[Lim13a] Yongdo Lim. Geometry of midpoint sets for Thompson’s metric. Linear Algebra and its Applications, 439 (1):211–227, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300195X.

Lim:2013:RNP

[Lim13b] Yongdo Lim. Review of Nonlinear Perron–Frobenius Theory by Bas Lemmens and Roger Nussbaum, Cambridge Tracts in Mathematics, Vol.

189. Cambridge University Press, Cambridge, UK(2012). xi + 323 pp, Hardback, ISBN: 978-0-521-89881-2. Linear Algebra and its Applications, 438(8):3560–3561, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000062.

Lim:2014:HAC

[Lim14] Yongdo Lim. Hamiltonian actions on the cone of positive definite matrices. Linear Algebra and its Applications, 460(??):1–16, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004546.

Lin:2010:RRH

[Lin10] Minghua Lin. A residual recurrence for Halley’s method for the matrix p th root. Linear Algebra and its Applications, 432(11):2928–2930, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2011:MPK

[Lin11] Bor-Luh Lin. In memoriam: Professor Ky Fan (1914– 2010). Linear Algebra and its Applications, 434(1):1–3, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2013:FTD

[Lin13] Minghua Lin. Fischer type determinantal inequalities for accretive-dissipative matrices. Linear Algebra and its Applications, 438(6):2808– 2812, March 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200821X.

Lin:2014:CPM

[Lin14a] Minghua Lin. A completely PPT map. Linear Algebra and its Applications, 459(??): 404–410, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004911.

Lin:2014:OTI

[Lin14b] Minghua Lin. An Oppenheim type inequality for a block Hadamard product. Linear Algebra and its Applications, 452(??):1–6, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001694.

Lippert:2010:FME

[Lip10] Ross A. Lippert. Fixing multiple eigenvalues by a minimal perturbation. Linear Algebra and its Applications, 432(7): 1785–1817, March 15, 2010.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2013:SFS

[Liu13] Pin-Lin Liu. State feedback stabilization of timevarying delay uncertain systems: a delay decomposition approach. Linear Algebra and its Applications, 438(5):2188–2209, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007185.

Liu:2014:CMI

[Liu14a] Cheng-Kai Liu. Centralizing maps on invertible or singular matrices over division rings. Linear Algebra and its Applications, 440(??): 318–324, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006356.

Liu:2014:SCP

[Liu14b] Cheng-Kai Liu. Strong commutativity preserving maps on subsets of matrices that are not closed under addition. Linear Algebra and its Applications, 458(??): 280–290, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Lu:2010:CLD

[LJ10] Fangyan Lu and Wu Jing. Characterizations of Lie derivations of B(X). Linear Algebra and its Applications, 432(1): 89–99, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:SRC

[LJ11] Bingyu Li and Zhongxiao Jia. Some results on condition numbers of the scaled total least squares problem. Linear Algebra and its Applications, 435(3):674–686, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2012:OGR

[LJ12] Xu-Qing Liu and Hong-Yan Jiang. Optimal generalized ridge estimator under the generalized crossvalidation criterion in linear regression. Linear Algebra and its Applications, 436(5):1238–1245, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100615X.

Liu:2013:HOC

[LJY13] Xiaoji Liu, Hongwei Jin, and Yaoming Yu. Higher-order convergent iterative method for computing the generalized inverse and its application to Toeplitz matrices. Linear Algebra and its Applications, 439(6):1635–1650, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003169.

Liu:2014:FEN

[LJY14] Ruifang Liu, Huicai Jia, and Jinjiang Yuan. First eigenvalue of nonsingular mixed graphs with given number of pendant vertices. Linear Algebra and its Applications, 453(??):28–43, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002201.

Li:2013:SCR

[LKN13] Na Li, Stefan Kindermann, and Carmeliza Navasca. Some convergence results on the Regularized Alternating Least-Squares method for tensor decomposition. Linear Algebra and its Applications, 438(2):796–812, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007828.

Lebl:2010:UCP

[LL10a] Jiří Lebl and Daniel Lichtblau. Uniqueness of certain polynomials constant on a line. Linear Algebra and its Applications, 433(4):824– 837, October 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2010:SCP

[LL10b] Jer-Shyong Lin and Cheng-Kai Liu. Strong commutativity preserving maps in prime rings with involution. Linear Algebra and its Applications, 432(1):14–23, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2010:SLS

[LL10c] Muhuo Liu and Bolian Liu. The signless Laplacian spread. Linear Algebra and its Applications, 432(2–3):505–514, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2010:MBR

[LL10d] Shu-Tian Liu and Xin-Long Luo. A method based on Rayleigh quotient gradient flow for extreme and interior eigenvalue problems. Linear Algebra and its Applications, 432(7):1851–1863, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2010:SRB

[LL10e] Yingluan Liu and Bolian Liu. The spectral radius of bicyclic graphs with prescribed degree sequences. Linear Algebra and its Applications, 433(5):1015– 1023, October 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2011:CSD

[LL11a] Pjek-Hwee Lee and Cheng-Kai Liu. On the composition of q-skew derivations in Banach algebras. Linear Algebra and its Applications, 434(12):2413–2429, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liau:2011:GLD

[LL11b] Pao-Kuei Liau and Cheng-Kai Liu. Generalized Lie derivations on skew elements of prime algebras with involution. Linear Algebra and its Applications, 435(1):67– 76, July 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2011:SRS

[LL11c] Muhuo Liu and Bolian Liu. On the spectral radii and the signless Laplacian spectral radii of c-cyclic graphs with fixed maximum degree. Linear Algebra and its Applications, 435(12):3045–3055, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2012:SPM

[LL12] Hosoo Lee and Yongdo Lim. Structure preserving matrix means on the Marcus-Minc stochastic matrices. Linear Algebra and its Applications, 437(10):2397–2407, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004715.

Lan:2013:DGS

[LL13a] Jingfen Lan and Linyuan Lu. Diameters of graphs with spectral radius at most \frac{3}{2} \sqrt{2}. Linear Algebra and its Applications, 438(11): 4382–4407, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000530.

Lee:2013:CIM

[LL13b] Hosoo Lee and Yongdo Lim. Carlson’s iterative mean algorithm of positive definite matrices. Linear Algebra and its Applications, 439(4): 1183–1201, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002486.

Liu:2013:SSV

[LL13c] Xiaogang Liu and Pengli Lu. Spectra of subdivision-vertex and subdivision-edge neighbourhood coronae. Linear Algebra and its Applications, 438(8):3547–3559, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000207.

Luo:2013:SOS

[LL13d] Jiajia Luo and Wei Li. Strong optimal solutions of interval linear programming. Linear Algebra and its Applications, 439(8):2479–2493, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004217.

Li:2014:MEL

[LL14a] Chaoqian Li and Yaotang Li. A modification of eigenvalue localization for stochastic matrices. Linear Algebra and its Applications, 460 (??):231–241, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004820.

Liu:2014:BSL

[LL14b] Huiqing Liu and Mei Lu. Bounds of signless Laplacian spectrum of graphs based on the k-domination number. Linear Algebra and its Applications, 440(??):83– 89, January 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300640X.

Liu:2014:CDS

[LL14c] Huiqing Liu and Mei Lu. A conjecture on the diameter and signless Laplacian index of graphs. Linear Algebra and its Applications, 450 (??):158–174, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001359.

Liu:2014:MTE

[LL14d] Muhuo Liu and Bolian Liu. The majorization theorem of extremal pseudographs. Linear Algebra and its Applications, 459(??):13–22, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004121.

Liang:2013:TMP

[LLB13] Xin Liang, Ren-Cang Li, and Zhaojun Bai. Trace minimization principles for positive semi-definite pencils. Linear Algebra and its Applications, 438(7):3085–3106, April 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008415.

Li:2013:NSC

[LLD13] Wei Li, Jiajia Luo, and Chongyang Deng. Necessary and sufficient conditions of some strong optimal solutions to the interval linear programming. Linear Algebra and its Applications, 439 (10):3241–3255, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005120.

Li:2013:NMP

[LLF13] Changjing Li, Fangyan Lu, and Xiaochun Fang. Nonlinear mappings preserving product XY + Y X^{∗} on factor von Neumann algebras. Linear Algebra and its Applications, 438(5):2339– 2345, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007252.

Liang:2010:LBP

[LLH10] Chaohua Liang, Bolian Liu, and Yufei Huang. The k th lower bases of primitive non-powerful signed digraphs. Linear Algebra and its Applications, 432(7):1680– 1690, March 15, 2010. CO-

Liu:2012:EET

[LLMZ12] Yiying Liu, Zhiping Lin, Giuseppe Molteni, and Dongye Zhang. Eigenvalues and equivalent transformation of a trigonometric matrix associated with filter design. Linear Algebra and its Applications, 437(12):2961–2972, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004946.

Lee:2012:NRM

[LLN^{+}12] Michael Z. Lee, Elizabeth Love, Sivaram K. Narayan, Elizabeth Wascher, and Jordan D. Webster. On nonsingular regular magic squares of odd order. Linear Algebra and its Applications, 437 (6):1346–1355, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002686.

Labra:2014:EAP

[LLR14] A. Labra, M. Ladra, and U. A. Rozikov. An evolution algebra in population genetics. Linear Algebra and its Applications, 457(??):348–362, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003383.

Li:2010:NER

[LLS10] Xueliang Li, Yiyang Li, and Yongtang Shi. Note on the energy of regular graphs. Linear Algebra and its Applications, 432(5):1144–1146, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:MET

[LLS11] Jing Li, Xueliang Li, and Yongtang Shi. On the maximal energy tree with two maximum degree vertices. Linear Algebra and its Applications, 435(9):2272–2284, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lan:2012:GDM

[LLS12] Jingfen Lan, Linyuan Lu, and Lingsheng Shi. Graphs with diameter n − e minimizing the spectral radius. Linear Algebra and its Applications, 437(11):2823–2850, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004429.

Lu:2012:SRH

[LLT12] Mei Lu, Huiqing Liu, and Feng Tian. Spectral radius and Hamiltonian graphs.

Linear Algebra and its Applications, 437(7):1670–1674, October 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003837.

Liu:2013:ECS

[LLT13] Huiqing Liu, Mei Lu, and Feng Tian. Edge-connectivity and (signless) Laplacian eigenvalue of graphs. Linear Algebra and its Applications, 439(12):3777–3784, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300637X.

Li:2014:SFI

[LLW14] Haohao Li, Jiajia Luo, and Qin Wang. Solvability and feasibility of interval linear equations and inequalities. Linear Algebra and its Applications, 463(??): 78–94, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005710.

Lee:2011:MVW

[LLY11] Hosoo Lee, Yongdo Lim, and Takeaki Yamazaki. Multivariable weighted geometric means of positive definite matrices. Linear Algebra and its Applications, 435(2):307–322, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lemos:2010:CHB

[LM10a] Manoel Lemos and T. R. B. Melo. Connected hyperplanes in binary matroids. Linear Algebra and its Applications, 432(1):259–274, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2010:SIL

[LM10b] Vakhtang Lomadze and Hasan Mahmood. Smooth/impulsive linear systems: Axiomatic description. Linear Algebra and its Applications, 433(11– 12):1997–2009, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luzon:2010:RRP

[LM10c] Ana Luzón and Manuel A. Morón. Recurrence relations for polynomial sequences via Riordan matrices. Linear Algebra and its Applications, 433(7):1422–1446, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2012:PNI

[LM12] Han Ju Lee and Miguel Martiń. Polynomial numerical indices of Banach spaces with 1-unconditional bases. Linear Algebra and its Applications, 437(8):2001–2008, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004193.

Lee:2011:PNI

[LMM11] Han Ju Lee, Miguel Martín, and Javier Merí. Polynomial numerical indices of Banach spaces with absolute norm. Linear Algebra and its Applications, 435(2):400–408, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Livshits:2013:PALa

[LMMR13a] L. Livshits, G. MacDonald, L. W. Marcoux, and H. Radjavi. Paratransitive algebras of linear operators. Linear Algebra and its Applications, 439(7):1955–1973, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003698.

Livshits:2013:PALb

[LMMR13b] L. Livshits, G. MacDonald, L. W. Marcoux, and H. Radjavi. Paratransitive algebras of linear operators II. Linear Algebra and its Applications, 439(7):1974–1989, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003704.

Luzon:2012:IIR

[LMMS12] Ana Luzón, Donatella Merlini, Manuel A. Morón, and Renzo Sprugnoli. Identities induced by Riordan arrays. Linear Algebra and its Applications, 436(3):631– 647, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005805.

Landsberg:2013:HAM

[LMN13] J. M. Landsberg, Jason Morton, and Serguei Norine. Holographic algorithms without matchgates. Linear Algebra and its Applications, 438(2):782–795, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000584.

Ladra:2016:CSC

[LMO16] M. Ladra, K. K. Masutova, and B. A. Omirov. Corrigendum to “Classification of solvable Leibniz algebras with naturally graded filiform nilradical” [Linear Algebra Appl. 438 (7) (2013) 2973–3000]. Linear Algebra and its Applications, 507(??): 513–517, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/

science/article/pii/S0024379516302749.

See [CLOK13].

Louka:2010:MPE

[LMT10] M. A. Louka, N. M. Missirlis, and F. I. Tzaferis. Is modified PSD equivalent to modified SOR for two-cyclic matrices? Linear Algebra and its Applications, 432(11):2798–2815, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2012:FAG

[LMW12] Wen Liu, Changli Ma, and Kaishun Wang. Full automorphism group of generalized unitary graphs. Linear Algebra and its Applications, 437(2):684–691, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001796.

Liu:2011:UBN

[LMYY11] X. Liu, S. McKee, J. Y. Yuan, and Y. X. Yuan. Uniform bounds on the 1-norm of the inverse of lower triangular Toeplitz matrices. Linear Algebra and its Applications, 435(5):1157–1170, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lemmens:2012:NPF

[LN12] Bas Lemmens and Roger D. Nussbaum. Nonlinear Perron– Frobenius theory, volume 189 of Cambridge tracts in mathematics. Cambridge University Press, Cambridge, UK, 2012. ISBN 0-521-89881-1 (hardback). xii + 323 pp. LCCN QA188 .L456

2012. URL http://assets.
cambridge.org/97805218/
98812/cover/9780521898812.
jpg.

Lankeit:2014:MML

[LNN14] Johannes Lankeit, Patrizio Neff, and Yuji Nakatsukasa. The minimization of matrix logarithms: On a fundamental property of the unitary polar factor. Linear Algebra and its Applications, 449 (??):28–42, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000792.

Liu:2013:USC

[LNT13] Xuhua Liu, Brice M. Nguelifack, and Tin-Yau Tam. Unitary similarity to a complex symmetric matrix and its extension to orthogonal symmetric Lie algebras. Linear Algebra and its Applications, 438(10):3789–3796, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005696.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Li:2012:PME

[LNTgW12] Ren-Cang Li, Yuji Nakatsukasa, Ninoslav Truhar, and Wei guo Wang. Perturbation of multiple eigenvalues of Hermitian matrices. Linear Algebra and its Applications, 437(1):202–213, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001164.

Loewy:2012:UGR

[Loe12] Raphael Loewy. Uriel G. Rothblum (1947–2012). Linear Algebra and its Applications, 437(12):2997–3009, December 15, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005368.

Lomadze:2011:RDO

[Lom11] Vakhtang Lomadze. Rational differential operators and their kernels. Linear Algebra and its Applications, 435 (11):2870–2888, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2013:NEF

[Lom13] Vakhtang Lomadze. A note on Ehrenpreis’ fundamental principle. Linear Algebra and its Applications, 438(5): 2083–2089, March 1, 2013.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007483.

Lopatin:2011:MGS

[Lop11a] A. A. Lopatin. Minimal generating set for semi-invariants of quivers of dimension two. Linear Algebra and its Applications, 434(8):1920–1944, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loperfido:2011:SAF

[Lop11b] Nicola Loperfido. Spectral analysis of the fourth moment matrix. Linear Algebra and its Applications, 435(8):1837– 1844, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2010:APM

[LP10] Jiankui Li and Zhidong Pan. Annihilator-preserving maps, multipliers, and derivations. Linear Algebra and its Applications, 432(1):5–13, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2012:MPJ

[LP12] Chi-Kwong Li and Edward Poon. Maps preserving the joint numerical radius distance of operators. Linear Algebra and its Applications, 437(5):1194–1204, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002947.

Lim:2014:WIM

[LP14] Yongdo Lim and Miklós Pálfia. Weighted inductive means. Linear Algebra and its Applications, 453 (??):59–83, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001943.

Leventides:2014:ADA

[LPK14] John Leventides, George Petroulakis, and Nicos Karcanias. The approximate Determinantal Assignment Problem. Linear Algebra and its Applications, 461(??):139– 162, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004571.

Laffey:2010:SIH

[LPQdS10] Thomas J. Laffey, Thomas H. Pate, Joao F. Queiró, and Eduardo Marques de Sá. Special issue in honor of Jose Dias da Silva. Linear Algebra and its Applications, 432 (10):2726, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2013:LPT

[LPS13] Chi-Kwong Li, Yiu-Tung Poonl, and Nung-Sing Sze. Linear preservers of tensor product of unitary orbits, and product numerical range. Linear Algebra and its Applications, 438(10): 3797–3803, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005830.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Luk:2011:PLA

[LQ11] Franklin T. Luk and Sanzheng Qiao. A pivoted LLL algorithm. Linear Algebra and its Applications, 434(11): 2296–2307, June 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2013:SME

[LQY13] Guoyin Li, Liqun Qi, and Gaohang Yu. Semismoothness of the maximum eigenvalue function of a symmetric tensor and its application. Linear Algebra and its Applications, 438(2):813–833, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007440.

Letchford:2010:CMR

[LRST10] Adam N. Letchford, Gerhard Reinelt, Hanna Seitz, and Dirk Oliver Theis. On a class of metrics related to graph layout problems. Linear Algebra and its Applications, 433 (11–12):1760–1777, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lebtahi:2012:CPM

[LRT12] Leila Lebtahi, Oscar Romero, and Néstor Thome. Characterizations of {K, s + 1}- potent matrices and applications. Linear Algebra and its Applications, 436(2): 293–306, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510006105.

Lebtahi:2013:RBI

[LRT13] Leila Lebtahi, Óscar Romero, and Néstor Thome. Relations between {K, s+1}-potent matrices and different classes of complex matrices. Linear Algebra and its Applications, 438(4):1517–1531, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007439.

Lampe:2012:LST

[LRV12] Jörg Lampe, Lothar Reichel, and Heinrich Voss. Largescale Tikhonov regularization via reduction by orthogonal projection. Linear Algebra and its Applications, 436(8): 2845–2865, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005313.

Li:2010:MSL

[LS10] Ruilin Li and Jinsong Shi. The minimum signless Laplacian spectral radius of graphs with given independence number. Linear Algebra and its Applications, 433(8–10):1614– 1622, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:PMO

[LS11a] Nian Li and Joseph M. Steiner. A perturbation method for optimizing matrix stability. Linear Algebra and its Applications, 434(3):641– 649, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2011:SRD

[LS11b] Huiqiu Lin and Jinlong Shu. Spectral radius of digraphs with given dichromatic number. Linear Algebra and its Applications, 434(12):2462– 2467, June 15, 2011. CO-DEN LAAPAW. ISSN 0024-

Le:2012:EEI

[LS12a] T. A. Le and J. W. Sander. Extremal energies of integral circulant graphs via multiplicativity. Linear Algebra and its Applications, 437 (6):1408–1421, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002765.

Lin:2012:NSC

[LS12b] Huiqiu Lin and Jinlong Shu. A note on the spectral characterization of strongly connected bicyclic digraphs. Linear Algebra and its Applications, 436(7):2524–2530, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006513.

Liu:2012:RNC

[LS12c] Yue Liu and Hai-Ying Shan. Ray nonsingularity of cycle chain matrices. Linear Algebra and its Applications, 437 (12):2910–2921, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005575.

Locatelli:2012:NSC

[LS12d] Arturo Locatelli and Nicola Schiavoni. A necessary and sufficient condition for the stabilisation of a matrix and its principal submatrices. Linear Algebra and its Applications, 436(7):2311–2314, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007233.

Laffey:2013:SRM

[LS13a] Thomas J. Laffey and Helena Smigoc. Simultaneous reduction of matrices in the presence of a nonderogatory matrix. Linear Algebra and its Applications, 438(10): 3885–3890, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007762.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Li:2013:EDS

[LS13b] Guojun Li and Lingsheng Shi. Edge-disjoint spanning trees and eigenvalues of graphs. Linear Algebra and its Applications, 439(10):2784– 2789, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005600.

Lindner:2013:MDP

[LS13c] Marko Lindner and Gilbert Strang. The main diagonal of a permutation matrix. Linear Algebra and its Applications, 439(3):524–537, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002467.

Liu:2013:FSR

[LS13d] Yue Liu and Hai-Ying Shan. A forbidden structure of ray nonsingular matrices. Linear Algebra and its Applications, 439(8):2367–2380, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004667.

Liu:2014:RAS

[LS14] Yan Liu and Yuming Shi. Regular approximations of spectra of singular secondorder symmetric linear difference equations. Linear Algebra and its Applications, 444(??):183–210, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007222.

Li:2010:LET

[LSC10] Jianxi Li, Wai Chee Shiu, and An Chang. On the k th Laplacian eigenvalues of trees with perfect matchings. Linear Algebra and its Applications, 432(4):1036– 1041, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:LSRb

[LSC11] Jianxi Li, Wai Chee Shiu, and Wai Hong Chan. On the Laplacian spectral radii of bipartite graphs. Linear Algebra and its Applications, 435(9): 2183–2192, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2014:SGD

[LSD14] Muhuo Liu, Haiying Shan, and Kinkar Ch. Das. Some graphs determined by their (signless) Laplacian spectra. Linear Algebra and its Applications, 449(??):154–165, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000949.

Luo:2012:CRS

[LSH12] Yousong Luo, Uwe Schwerdtfeger, and Robin Hill. A commutativity relation for submatrices of the inverse of a Sylvester matrix. Linear Algebra and its Applications, 437 (10):2560–2572, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Liu:2011:CRP

[LSR11] Yue Liu, Jia-Yu Shao, and Ling-Zhi Ren. Characterization of ray pattern matrix whose determinantal region has two components after deleting the origin. Linear Algebra and its Applications, 435(12):3139–3150, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lebtahi:2013:MSI

[LSTW13] Leila Lebtahi, Jeffrey Stuart, Néstor Thome, and James R. Weaver. Matrices A such that RA = A^{s+1}R when R^{k} = I. Linear Algebra and its Applications, 439(4): 1017–1023, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007641.

Limbupasiriporn:2012:LWC

[LSV12] J. Limbupasiriporn, L. Storme, and P. Vandendriessche. Large weight code words in projective space codes. Linear Algebra and its Applications, 437(3):809–816, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002558.

Lim:2010:BLM

[LT10a] Ming-Huat Lim and Sin-Chee Tan. Bijective linear maps on semimodules spanned by Boolean matrices of fixed rank. Linear Algebra and its Applications, 433(7):1365– 1373, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:2010:MEP

[LT10b] Raphael Loewy and Bit-Shun Tam. Maximal exponents of polyhedral cones (II). Linear Algebra and its Applications, 432(11):2861–2878, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2011:LSRa

[LT11a] Shuchao Li and Yi Tian. On the (Laplacian) spectral radius of weighted trees with fixed matching number q and a positive weight set. Linear Algebra and its Applications, 435(6):1202–1212, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2011:ROP

[LT11b] Ming-Huat Lim and Sin-Chee Tan. Rank one preservers between spaces of Boolean matrices. Linear Algebra and its Applications, 434(2):526– 541, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-

Lancaster:2012:HQM

[LT12a] Peter Lancaster and Françoise Tisseur. Hermitian quadratic matrix polynomials: Solvents and inverse problems. Linear Algebra and its Applications, 436(10):4017–4026, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510003435.

Levene:2012:CBN

[LT12b] Rupert H. Levene and Richard M. Timoney. Completely bounded norms of right module maps. Linear Algebra and its Applications, 436(5):1406–1424, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006203.

See corrigendum [LT16].

Lin:2013:CSM

[LT13] Y.-F. Lin and I. G. Todorov. Compact separating maps on continuous fields of Banach spaces. Linear Algebra and its Applications, 439 (5):1221–1228, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002255.

Levene:2016:CSC

[LT16] Rupert H. Levene and Richard M. Timoney. Corrigendum to “Completely bounded norms of right module maps” [Linear Algebra Appl. 436 (5) (2012) 1406–1424]. Linear Algebra and its Applications, 505(??):387–389, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379516301781.

See [LT12b].

Li:2013:SLC

[LTS13] Hong-Hai Li, Bit-Shun Tam, and Li Su. On the signless Laplacian coefficients of unicyclic graphs. Linear Algebra and its Applications, 439(7):2008–2028, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003844.

Luo:2014:LRS

[LTX14] Ziyan Luo, Jiyuan Tao, and Naihua Xiu. Lowest-rank solutions of continuous and discrete Lyapunov equations over symmetric cone. Linear Algebra and its Applications, 452(??):68–88, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001724.

Lu:2011:LDO

[Lu11] Fangyan Lu. Linear dependence of operators characterized by trace functionals. Linear Algebra and its Applications, 434(1):343–355, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2012:RBW

[Lu12]  Zhiqin Lu.  Remarks on

the Böttcher–Wenzel inequality. Linear Algebra and its Applications, 436(7):2531– 2535, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006495.

Lee:2011:SSS

[LV11] Sang-Gu Lee and Quoc-Phong Vu. Simultaneous solutions of Sylvester equations and idempotent matrices separating the joint spectrum. Linear Algebra and its Applications, 435(9):2097– 2109, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2012:SSM

[LV12] Sang-Gu Lee and Quoc-Phong Vu. Simultaneous solutions of matrix equations and simultaneous equivalence of matrices. Linear Algebra and its Applications, 437 (9):2325–2339, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200434X.

Laurent:2014:PSM

[LV14] M. Laurent and A. Varvitsiotis. Positive semidefinite matrix completion, universal rigidity and the Strong Arnold Property. Linear Algebra and its Applications, 452 (??):292–317, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001591.

Lang:2012:EGL

[LW12a] Weiwei Lang and Ligong Wang. Energy of generalized line graphs. Linear Algebra and its Applications, 437 (9):2386–2396, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003850.

Li:2012:LES

[LW12b] Shuchao Li and Shujing Wang. The least eigenvalue of the signless Laplacian of the complements of trees. Linear Algebra and its Applications, 436(7):2398–2405, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Li:2012:SCM

[LW12c] Yanbo Li and Feng Wei. Semicentralizing maps of generalized matrix algebras. Linear Algebra and its Applications, 436(5):1122–1153, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100526X.

Liu:2012:MFR

[LW12d] Wende Liu and Shujuan Wang. Minimal faithful representations of abelian Jordan algebras and Lie superalgebras. Linear Algebra and its Applications, 437 (5):1293–1299, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003059.

Liu:2012:LSC

[LW12e] Xiaogang Liu and Suijie Wang. Laplacian spectral characterization of some graph products. Linear Algebra and its Applications, 437(7):1749–1759, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003400.

Li:2013:CGP

[LW13a] Xin Li and Wei Wu. A class of generalized positive linear maps on matrix algebras. Linear Algebra and its Applications, 439(10):2844– 2860, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005430.

Liu:2013:NRM

[LW13b] Jinwang Liu and Mingsheng Wang. New results on multivariate polynomial matrix factorizations. Linear Algebra and its Applications, 438 (1):87–95, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006222.

Li:2011:SNA

[LwCJL11] Tiexiang Li, Eric King wah Chu, Jong Juang, and Wen-Wei Lin. Solution of a nonsymmetric algebraic Riccati equation from a twodimensional transport model. Linear Algebra and its Applications, 434(1):201–214, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2013:MSG

[LWG13] Fenggao Li, Kaishun Wang, and Jun Guo. More on symplectic graphs modulo p^{n}. Linear Algebra and its Applications, 438(6):2651– 2660, March 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004740.

Li:2010:SPS

[LWGM10] Fenggao Li, Kaishun Wang, Jun Guo, and Jianmin Ma. Suborbits of a point-stabilizer in the unitary group on the last subconstituent of Hermitean dual polar graphs. Linear Algebra and its Applications, 433(2):333–341, August 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2012:SPS

[LWGM12] Fenggao Li, Kaishun Wang, Jun Guo, and Jianmin Ma. Suborbits of a point stabilizer in the orthogonal group on the last subconstituent of orthogonal dual polar graphs. Linear Algebra and its Applications, 436(5):1297–1311, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005817.

Li:2012:BSR

[LWV12] Cong Li, Huijuan Wang, and Piet Van Mieghem. Bounds for the spectral radius of a graph when nodes are removed. Linear Algebra and its Applications, 437 (1):319–323, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001693.

Lee:2014:CBD

[LwW14] Guang-Siang Lee and Chih wen Weng. A characterization of bipartite distanceregular graphs. Linear Algebra and its Applications, 446 (??):91–103, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300832X.

Lu:2010:EOF

[LWY10] Hongliang Lu, Zefang Wu, and Xu Yang. Eigenvalues and [1, n]-odd factors. Linear Algebra and its Applications, 433(4):750–757, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2014:MCC

[LWY14] Hanyu Li, Shaoxin Wang, and Hu Yang. On mixed and componentwise condition numbers for indefinite least squares problem. Linear Algebra and its Applications, 448(??):104–129, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Li:2011:SLS

[LWZ11] Ke Li, Ligong Wang, and Guopeng Zhao. The signless Laplacian spectral radius of tricyclic graphs and trees with k pendant vertices. Linear Algebra and its Applications, 435(4):811–822, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2013:ROC

[LX13] Jun Liu and Mingqing Xiao. Rank-one characterization of joint spectral radius of finite matrix family. Linear Algebra and its Applications, 438(8):3258–3277, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000190.

Luo:2012:LTL

[LXK12] Ziyan Luo, Naihua Xiu, and Lingchen Kong. Lyapunovtype least-squares problems over symmetric cones. Linear Algebra and its Applications, 437(10):2498–2515, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004685.

Lei:2014:SDS

[LXL^{+}14] Ying-Jie Lei, Wei-Ru Xu, Yong Lu, Yan-Ru Niu, and Xian-Ming Gu. On the symmetric doubly stochastic inverse eigenvalue problem. Linear Algebra and its Applications, 445(??): 181–205, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008033.

Li:2011:TGW

[LY11a] Shuchao Li and Huangxu Yang. On tricyclic graphs whose second largest eigenvalue does not exceed 1. Linear Algebra and its Applications, 434(10):2211–2221, May 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2011:LCT

[LY11b] Weiqi Lin and Weigen Yan. Laplacian coefficients of trees with a given bipartition. Linear Algebra and its Applications, 435(1):152–162, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2012:RRS

[lY12] Sheng liang Yang. Recurrence relations for the Sheffer sequences. Linear Algebra and its Applications, 437 (12):2986–2996, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005563.

Lim:2013:SIM

[LY13] Yongdo Lim and Takeaki Yamazaki. On some inequalities for the matrix power and Karcher means. Linear Algebra and its Applications, 438 (3):1293–1304, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006477.

Liu:2013:PCL

[LYL13] Muhuo Liu, Lihua You, and Bolian Liu. The proof of a conjecture on the Lewin number of primitive non-powerful signed digraphs. Linear Algebra and its Applications, 438(5):2366–2377, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007744.

Yang:2013:SMI

[lYnZpYH13] Sheng liang Yang, Sai nan Zheng, Shao peng Yuan, and Tian-Xiao He. Schröder matrix as inverse of Delannoy matrix. Linear Algebra and its Applications, 439 (11):3605–3614, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006150.

Li:2013:TSG

[LYS13] Hanyu Li, Hu Yang, and Hua Shao. Two-sided generalized hyperbolic QR factorization and its perturbation analysis. Linear Algebra and its Applications, 438 (3):1267–1292, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006842.

Lee:2010:CNR

[LZ10] Tsiu-Kwen Lee and Yiqiang Zhou. A characterization of von Neumann regular rings and applications. Linear Algebra and its Applications, 433 (8–10):1536–1540, December 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2011:BPB

[LZ11a] Fu-Rong Lin and De-Cai Zhang. BTTB preconditioners for BTTB least squares problems. Linear Algebra and its Applications, 434(11): 2285–2295, June 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2011:FOR

[LZ11b] Vakhtang Lomadze and M. Khurram Zafar. First order representations of Fliess models. Linear Algebra and its Applications, 434(4):1027–1057, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2012:PPS

[LZ12a] Chunna Li and Yiqiang Zhou. On p.p. structural matrix rings. Linear Algebra and its Applications, 436(9):3692– 3700, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000547.

Li:2012:SLI

[LZ12b] Shuchao Li and Minjie Zhang. On the signless Laplacian index of cacti with a given number of pendant vertices. Linear Algebra and its Applications, 436(12): 4400–4411, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003016.

Liu:2013:CPP

[LZ13] Shunyi Liu and Heping Zhang. On the characterizing properties of the permanental polynomials of graphs. Linear Algebra and its Applications, 438(1):157–172, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006507.

Lin:2014:LDE

[LZ14] Hongying Lin and Bo Zhou. On least distance eigenvalues of trees, unicyclic graphs and bicyclic graphs. Linear Algebra and its Applications, 443(??):153–163, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300712X.

Lin:2014:GLT

[LZG14] Huiqiu Lin, Mingqing Zhai, and Shicai Gong. On graphs with at least three distance eigenvalues less than −1. Linear Algebra and its Applications, 458(??):548–558, October 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004091.

Liu:2012:SCS

[LZL12] Jianzhou Liu, Juan Zhang, and Yu Liu. The Schur complement of strictly doubly diagonally dominant matrices and its application. Linear Algebra and its Applications, 437(1):168–183, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Ma:2010:NEL

[Ma10a] Xinrong Ma. A novel extension of the Lagrange– Bürmann expansion formula. Linear Algebra and its Applications, 433(11–12):2152– 2160, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mouhoubi:2010:NPB

[MA10b] Zahir Mouhoubi and Djamil Aïssani. New perturbation bounds for denumerable Markov chains. Linear Algebra and its Applications, 432(7):1627–1649, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ma:2011:TDR

[Ma11] Jing Ma. Two-dimensional representations of finitely generated free group over an arbitrary field. Linear Algebra and its Applications, 434(12): 2456–2461, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matharu:2012:SIU

[MA12] Jagjit Singh Matharu and Jaspal Singh Aujla. Some inequalities for unitarily invariant norm. Linear Algebra and its Applications, 436(6):1623–1631, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510004210.

Ma:2014:IZP

[Ma14] Chao Ma. Idempotent zero patterns. Linear Algebra and its Applications, 449 (??):465–474, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001086.

Mackey:2013:CIF

[Mac13] D. Steven Mackey. The continuing influence of Fiedler’s work on companion matrices. Linear Algebra and its Applications, 439(4):810– 817, August 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007361.

Mahdavi-Amiri:2013:ERR

[MAGR13] Nezam Mahdavi-Amiri and Effat Golpar-Raboky. Extended rank reduction formulas containing Wedderburn and Abaffy-Broyden-Spedicato rank reducing processes. Linear Algebra and its Applications, 439(11):3318– 3331, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Mahmoud:2011:NFD

[Mah11] Magdi S. Mahmoud. New filter design for linear timedelay systems. Linear Algebra and its Applications, 434(4): 1080–1093, February 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mourad:2013:ACD

[MAM^{+}13] Bassam Mourad, Hassan Abbas, Ayman Mourad, Ahmad Ghaddar, and Issam Kaddoura. An algorithm for constructing doubly stochastic matrices for the inverse eigenvalue problem. Linear Algebra and its Applications, 439 (5):1382–1400, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002851.

Marovt:2010:HMS

[Mar10] Janko Marovt. Homomorphisms of matrix semigroups over division rings from dimension two to four. Linear Algebra and its Applications, 432(6):1595–1607, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mary:2011:GIG

[Mar11] X. Mary. On generalized inverses and Green’s relations.

Linear Algebra and its Applications, 434(8):1836–1844, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martin:2013:EGP

[Mar13a] Antonio J. Calderón Martín. On extended graded Poisson algebras. Linear Algebra and its Applications, 439(4):879–892, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005125.

Martino:2013:PIJ

[Mar13b] Fabrizio Martino. Polynomial identities for the Jordan algebra of a degenerate symmetric bilinear form. Linear Algebra and its Applications, 439(12):4080–4089, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006320.

Mary:2013:RGI

[Mar13c] X. Mary. Reprint of: On generalized inverses and Green’s relations. Linear Algebra and its Applications, 438(4): 1532–1540, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008294.

Martin:2014:LAS

[Mar14a] Antonio J. Calderón Martín. Lie algebras with a set grading. Linear Algebra and its Applications, 452(??):7– 20, July 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400175X.

Martin:2014:SGC

[Mar14b] Antonio J. Calderón Martín. On the structure of graded commutative algebras. Linear Algebra and its Applications, 447(??):110–118, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001602.

Martinez-Avendano:2013:IPS

[MARC13] Rubén A. Martínez-Avendaño and Josué I. Rios-Cangas. Inner products on the space of complex square matrices. Linear Algebra and its Applications, 439(11):3620– 3637, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005867.

Mohar:2012:EGV

[MAS12] Bojan Mohar, Azhvan Sheikh Ahmady, and Rayman Preet Singh. Eigenvalues of graphs with vertices of large degree at distance three apart. Linear Algebra and its Applications, 436(11):4342–4347, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000997.

Matsuura:2012:NGM

[Mat12] Masaya Matsuura. A note on generalized G-matrices. Linear Algebra and its Applications, 436(9):3475–3479, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008287.

Mathai:2013:FIO

[Mat13a] A. M. Mathai. Fractional integral operators in the complex matrix variate case. Linear Algebra and its Applications, 439(10):2901–2913, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005223.

Matvejchuk:2013:ISC

[Mat13b] Marjan Matvejchuk. Idempotents in a space with conjugation. Linear Algebra and its Applications, 438 (1):71–79, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Mathai:2014:FIO

[Mat14a] A. M. Mathai. Fractional integral operators involving many matrix variables. Linear Algebra and its Applications, 446(??):196–215, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000196.

Matic:2014:IDP

[Mat14b]  Ivan Matic.  Inequalities

with determinants of perturbed positive matrices. Linear Algebra and its Applications, 449(??):166–174, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000937.

Maze:2010:SIR

[Maz10] Gérard Maze. Some inequalities related to the Seysen measure of a lattice. Linear Algebra and its Applications, 433 (8–10):1659–1665, December 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ma:2013:NSN

[MB13] Wei Ma and Zheng-Jian Bai. A note on simple nonzero finite generalized singular values. Linear Algebra and its Applications, 438 (8):3425–3441, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000359.

Mosic:2010:ARW

[MD10] Dijana Mosić and Dragan S. Djordjević. Additive results for the Wg–Drazin inverse. Linear Algebra and its Applications, 432(11):2847–2860, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martin:2012:SLA

[MD12a] Antonio J. Calderón Martín and José M. Sánchez Delgado. On split Leibniz algebras. Linear Algebra and its Applications, 436(6): 1651–1663, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001807.

Martin:2012:SSL

[MD12b] Antonio J. Calderón Martín and José M. Sánchez Delgado. On the structure of split Lie color algebras. Linear Algebra and its Applications, 436(2):307–315, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001030.

Mom:2012:CEU

[MD12c] Alain Mom and Pierre Druilhet. Characterization of estimators uniformly shrinking on subspaces. Linear Algebra and its Applications, 437(7):1771–1778, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003138.

Mao:2013:MNP

[MD13] Xiaobin Mao and Hua Dai. Minimum norm partial eigenvalue assignment of high order linear system with no spill-over. Linear Algebra and its Applications, 438(5): 2136–2154, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007793.

Meini:2013:SDT

[Mei13] Beatrice Meini. A ‘shiftand-deflate’ technique for quadratic matrix polynomials. Linear Algebra and its Applications, 438(4):1946– 1961, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007816.

Melman:2013:GVP

[Mel13] A. Melman. Generalization and variations of Pellet’s theorem for matrix polynomials. Linear Algebra and its Applications, 439(5):1550– 1567, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003145.

Melman:2014:IDP

[Mel14] A. Melman. Inclusion disks for polynomial zeros in generalized bases. Linear Algebra and its Applications, 445(??):326–346, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008240.

Merlet:2010:SMA

[Mer10] Glenn Merlet. Semigroup of matrices acting on the maxplus projective space. Linear Algebra and its Applications, 432(8):1923–1935, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merino:2012:SOM

[Mer12] Dennis I. Merino. The sum of orthogonal matrices. Linear Algebra and its Applications, 436(7):1960–1968, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007051.

Meyer:2012:ZFS

[Mey12] Seth A. Meyer. Zero forcing sets and bipartite circulants. Linear Algebra and its Applications, 436(4): 888–900, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006550.

Martin:2014:GSH

[MFGD14] Antonio Jesús Calderón Martín, Cristina Draper Fontanals, Cándido Martín González, and José María Sánchez Delgado. Gradings and symmetries on Heisenberg type algebras. Linear Algebra and its Applications, 458(??): 463–502, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003905.

Ma:2011:GFI

[MGLW11] Jianmin Ma, Jun Guo, Fenggao Li, and Kaishun Wang. A generalization of the formulas for intersection numbers of dual polar association schemes and their applications. Linear Algebra and its Applications, 434(5):1272– 1284, March 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mei:2014:MNN

[MGSW14] Yinzhen Mei, Yubin Gao, Yanling Shao, and Peng Wang. The minimum number of nonzeros in a spectrally arbitrary ray pattern. Linear Algebra and its Applications, 453(??):99–109, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002195.

Marsli:2013:FRG

[MH13a] Rachid Marsli and Frank J. Hall. Further results on Gersgorin Discs. Linear Algebra and its Applications, 439 (1):189–195, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001523.

Mattila:2013:DIJ

[MH13b] Mika Mattila and Pentti Haukkanen. Determinant and inverse of join matrices on two sets. Linear Algebra and its Applications, 438(10):3891–3904, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008317.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Miguel:2013:NCA

[Mig13] C. Miguel. A note on a conjecture about commuting graphs. Linear Algebra and its Applications, 438(12): 4750–4756, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001341.

Mirman:2010:SCP

[Mir10] Boris Mirman. Short cycles of Poncelet’s conics. Linear Algebra and its Applications, 432(10):2543–2564, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mirman:2012:ESP

[Mir12] Boris Mirman. Explicit solutions to Poncelet’s porism. Linear Algebra and its Applications, 436(9):3531–3552, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379511008603.

Mitchell:2011:GCC

[Mit11a] Lon H. Mitchell. On the graph complement conjecture for minimum semidefinite rank. Linear Algebra and its Applications, 435(6):1311–1314, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitrouli:2011:ODE

[Mit11b] M. Mitrouli. D-optimal designs embedded in Hadamard matrices and their effect on the pivot patterns. Linear Algebra and its Applications, 434(7):1761–1772, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miyajima:2013:FES

[Miy13] Shinya Miyajima. Fast enclosure for solutions of Sylvester equations. Linear Algebra and its Applications, 439(4): 856–878, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005058.

Miyajima:2014:CES

[Miy14] Shinya Miyajima. Componentwise enclosure for solutions of least squares problems and underdetermined systems. Linear Algebra and its Applications, 444 (??):28–41, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007829.

Mirzakhah:2012:SRS

[MK12] Maryam Mirzakhah and Dariush Kiani. Some results on signless Laplacian coefficients of graphs. Linear Algebra and its Applications, 437 (9):2243–2251, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003849.

Martin:2010:ILT

[MLC^{+}10] William Martin, Sergio Loch, Laurel Cooley, Scott Dexter, and Draga Vidakovic. Integrating learning theories and application-based modules in teaching linear algebra. Linear Algebra and its Applications, 432(8):2089–2099, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Malmskog:2010:ADI

[MM10a] Beth Malmskog and Michelle Manes. “almost divisibility” in the Ihara zeta functions of certain ramified covers of q+1regular graphs. Linear Algebra and its Applications, 432 (10):2486–2506, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marco:2010:ACS

[MM10b] Ana Marco and José-Javier Martínez. Accurate computations with Said–Ball– Vandermonde matrices. Linear Algebra and its Applications, 432(11):2894–2908, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marco:2010:PLS

[MM10c] Ana Marco and José-Javier Martínez. Polynomial least squares fitting in the Bernstein basis. Linear Algebra and its Applications, 433(7): 1254–1264, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maehara:2011:SSV

[MM11a] Takanori Maehara and Kazuo Murota. Simultaneous singular value decomposition. Linear Algebra and its Applications, 435(1):106–116, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McLeman:2011:SC

[MM11b] Cam McLeman and Erin Mc-Nicholas. Spectra of coronae. Linear Algebra and its Applications, 435(5):998– 1007, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Magret:2012:ECE

[MM12] M. Dolors Magret and M. Eulalia Montoro. On the existence of a common eigenvector for all matrices in the commutant of a single matrix. Linear Algebra and its Applications, 437(5):1285–1292, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Mastylo:2013:NEM

[MM13] Mieczyslaw Mastylo and Pawel Mleczko. Norm estimates for matrix operators between Banach spaces. Linear Algebra and its Applications, 438(3):986–1001, February 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006659.

Matharu:2011:EEB

[MMA11] Jagjit Singh Matharu, Mohammad Sal Moslehian, and Jaspal Singh Aujla. Eigenvalue extensions of Bohr’s inequality. Linear Algebra and its Applications, 435(2):270– 276, July 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moslehian:2012:NCC

[MMA12] Mohammad Sal Moslehian, Jagjit Singh Matharu, and Jaspal Singh Aujla. Noncommutative Callebaut inequality. Linear Algebra and its Applications, 436(9): 3347–3353, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007671.

Moslehian:2013:OIC

[MMK13] M. S. Moslehian, J. Mićić, and M. Kian. An operator inequality and its consequences. Linear Algebra and its Applications, 439(3):584–591, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006039.

Morassaei:2013:BIH

[MMM13] A. Morassaei, F. Mirzapour, and M. S. Moslehian. Bellman inequality for Hilbert space operators. Linear Algebra and its Applications, 438(10):3776–3780, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004927.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Mackey:2010:JSA

[MMMM10] D. Steven Mackey, Niloufer Mackey, Christian Mehl, and Volker Mehrmann. Jordan structures of alternating matrix polynomials. Linear Algebra and its Applications, 432(4):867–891, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mackey:2013:SSM

[MMMM13] D. Steven Mackey, Niloufer Mackey, Christian Mehl, and Volker Mehrmann. Skewsymmetric matrix polynomials and their Smith forms. Linear Algebra and its Applications, 438(12):4625–4653, June 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001110.

Mingueza:2013:DCN

[MMP13a] David Mingueza, M. Eulàlia Montoro, and Juan R. Pacha. Description of characteristic non-hyperinvariant subspaces over the field GF(2). Linear Algebra and its Applications, 439(12):3734–3745, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006599.

Molnarova:2013:RIF

[MMP13b] M. Molnárová, H. Mysková, and J. Plavka. The robustness of interval fuzzy matrices. Linear Algebra and its Applications, 438(8): 3350–3364, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000074.

Mehl:2011:EPT

[MMRR11] Christian Mehl, Volker Mehrmann, André C. M. Ran, and Leiba Rodman. Eigenvalue perturbation theory of classes of structured matrices under generic structured rank one perturbations. Linear Algebra and its Applications, 435(3): 687–716, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehl:2012:PTS

[MMRR12] Christian Mehl, Volker Mehrmann, André C. M. Ran, and Leiba Rodman. Perturbation theory of selfadjoint matrices and sign characteristics under generic structured rank one perturbations. Linear Algebra and its Applications, 436(10):4027–4042, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379510001850.

MacDonald:2012:DLS

[MMS12] G. W. MacDonald, J. A. Mac-Dougall, and L. G. Sweet. On the dimension of linear spaces of nilpotent matrices. Linear Algebra and its Applications, 436(7):2210–2230, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007257.

Moslehian:2012:ELH

[MN12a] Mohammad Sal Moslehian and Hamed Najafi. An extension of the Löwner–Heinz inequality. Linear Algebra and its Applications, 437(9): 2359–2365, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004132.

Mursaleen:2012:CMO

[MN12b] M. Mursaleen and Abdullah K. Noman. Compactness of matrix operators on some new difference sequence spaces. Linear Algebra and its Applications, 436 (1):41–52, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004630.

Mitchell:2010:LBM

[MNZ10] Lon H. Mitchell, Sivaram K. Narayan, and Andrew M. Zimmer. Lower bounds in minimum rank problems. Linear Algebra and its Applications, 432(1):430–440, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mitchell:2012:LBM

[MNZ12] Lon H. Mitchell, Sivaram K. Narayan, and Andrew M. Zimmer. Lower bounds for minimum semidefinite rank from orthogonal removal and chordal supergraphs. Linear Algebra and its Applications, 436(3):525–536, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005167.

Melamed:2014:CIO

[MO14] Michal Melamed and Shmuel Onn. Convex integer optimization by constantly many linear counterparts. Linear Algebra and its Applications, 447(??):88–109, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000251.

Marshall:2011:ITM

[MOA11] Albert W. Marshall, Ingram Olkin, and Barry C. Arnold. Inequalities: Theory of Majorization and Its Applications, volume ?? of Springer Series in Statistics. Springer Science+Business Media, LLC, New York, NY, USA, second edition, 2011. CODEN ???? ISBN 0-387-40087-7, 0-387-68276-7 (e-book). ISSN 0172-7397. xxvii + 909 pp. LCCN QA295 .M37 2011; QA295 .M42 2011. URL http:/

/link.springer.com/book/
10.1007/978-0-387-68276-
1;   http://www.loc.gov/
catdir/enhancements/fy1114/
2010931704-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1114/2010931704-d.html;
 http://www.loc.gov/catdir/

enhancements/fy1114/2010931704-
t.html.

Mohar:2010:ECD

[Moh10] Bojan Mohar. Eigenvalues and colorings of digraphs. Linear Algebra and its Applications, 432(9):2273–2277, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mohlenkamp:2013:MMF

[Moh13] Martin J. Mohlenkamp. Musings on multilinear fitting. Linear Algebra and its Applications, 438(2):834–852, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003247.

Mojskerc:2014:SFD

[Moj14]  Blaz Mojskerc.  On the

structure of finite-dimensional paracontractions. Linear Algebra and its Applications, 446(??):148–162, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400024X.

Molnar:2011:OAP

[Mol11] Lajos Molnár. Order automorphisms on positive definite operators and a few applications. Linear Algebra and its Applications, 434(10): 2158–2169, May 15, 2011.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moldovan:2012:SDD

[Mol12] Melania M. Moldovan. Strict double diagonal dominance in Euclidean Jordan algebras. Linear Algebra and its Applications, 436(7):2513– 2523, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007531.

Molnar:2013:JTE

[Mol13] Lajos Molnár. Jordan triple endomorphisms and isometries of unitary groups. Linear Algebra and its Applications, 439(11):3518–3531, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005685.

Moore:2011:OPE

[Moo11] Gerald Moore. Orthogonal polynomial expansions for the matrix exponential. Linear Algebra and its Applications, 435(3):537–559, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Morris:2010:CSF

[Mor10] Ian D. Morris. Criteria for the stability of the finiteness property and for the uniqueness of Barabanov norms. Linear Algebra and its Applications, 433(7):1301–1311, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moslehian:2011:OAI

[Mos11] Mohammad Sal Moslehian. Operator Aczél inequality. Linear Algebra and its Applications, 434(8):1981–1987, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mosic:2013:MRG

[Mos13]  Dijana Mosić.  More re-

sults on generalized Drazin inverse of block matrices in Banach algebras. Linear Algebra and its Applications, 439 (8):2468–2478, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004576.

Mourad:2012:SPD

[Mou12]  Bassam Mourad.  On a

spectral property of doubly stochastic matrices and its application to their inverse eigenvalue problem. Linear Algebra and its Applications, 436(9):3400–3412, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007774.

Muhic:2010:QTP

[MP10] Andrej Muhič and Bor Plestenjak. On the quadratic two-parameter eigenvalue problem and its linearization. Linear Algebra and its Applications, 432(10):2529–2542, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Muller:2013:GSR

[MP13a] Vladimir Müller and Aljosa Peperko. Generalized spectral radius and its max algebra version. Linear Algebra and its Applications, 439(4):1006–1016, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007380.

Myskova:2013:RIC

[MP13b] H. Mysková and J. Plavka. X-robustness of interval circulant matrices in fuzzy algebra. Linear Algebra and its Applications, 438(6):2757– 2769, March 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008336.

Muhic:2014:MCA

[MP14a] Andrej Muhic and Bor Plestenjak. A method for computing all values λ such that A + λB has a multiple eigenvalue. Linear Algebra and its Applications, 440(??):345–359, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006344.

Myskova:2014:RIM

[MP14b] H. Mysková and J. Plavka. The robustness of interval matrices in max-plus algebra. Linear Algebra and its Applications, 445(??):85– 102, March 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008069.

Merino:2010:PDM

[MPP10] Dennis I. Merino, Agnes T. Paras, and Diane Christine P. Pelejo. On the ϕ_{J} polar decomposition of matrices. Linear Algebra and its Applications, 432(5):1165–1175, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See related work [DoMP09].

Micic:2011:JIO

[MPP11] Jadranka Mićić, Zlatko Pavić, and Josip Pečarić. Jensen’s inequality for operators without operator convexity. Linear Algebra and its Applications, 434(5):1228–1237, March 1, 2011. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic).

Merino:2011:SOM

[MPRW11] Dennis I. Merino, Agnes T. Paras, Edgar Reyes, and Gary Walls. The sum of orthogonal matrices in M_{n}(Z_{k}). Linear Algebra and its Applications, 434(10):2170–2175, May 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marczak:2010:CIQ

[MPS10] Grzegorz Marczak, Agnieszka Polak, and Daniel Simson. P -critical integral quadratic forms and positive unit forms: an algorithmic approach. Linear Algebra and its Applications, 433(11–12):1873–1888, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moghaddamfar:2010:MRA

[MPSS10] A. R. Moghaddamfar, S. M. H. Pooya, S. Navid Salehy, and S. Nima Salehy. On the matrices related to the marithmetic triangle. Linear Algebra and its Applications, 432(1):53–69, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merino:2012:HM

[MPT12] Dennis I. Merino, Agnes T. Paras, and Terrence Erard D. Teh. The Λ_{S}-Householder matrices. Linear Algebra and its Applications, 436(7): 2653–2664, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100694X.

McDonald:2014:MRE

[MPT14] Judith J. McDonald, Pietro Paparella, and Michael J. Tsatsomeros. Matrix roots of eventually positive matrices. Linear Algebra and its Applications, 456(??):122– 137, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007210.

McTigue:2011:PMW

[MQ11] James McTigue and Rachel Quinlan. Partial matrices whose completions have ranks bounded below. Linear Algebra and its Applications, 435 (9):2259–2271, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McTigue:2013:PMW

[MQ13] James McTigue and Rachel Quinlan. Partial matrices whose completions all have the same rank. Linear Algebra and its Applications, 438(1):348–360, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005587.

McTigue:2014:PMC

[MQ14] James McTigue and Rachel Quinlan. Partial matrices of constant rank. Linear Algebra and its Applications, 446 (??):177–191, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008288.

Mead:2010:LSP

[MR10a] Jodi L. Mead and Rosemary A. Renaut. Least squares problems with inequality constraints as quadratic constraints. Linear Algebra and its Applications, 432(8): 1936–1949, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Monsalve:2010:NIF

[MR10b] Marlliny Monsalve and Marcos Raydan. A new inversionfree method for a rational matrix equation. Linear Algebra and its Applications, 433 (1):64–71, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [MR14c].

Moslehian:2010:GIP

[MR10c] Mohammad Sal Moslehian and Rajna Rajić. A Grüss inequality for n-positive linear maps. Linear Algebra and its Applications, 433(8–10):1555– 1560, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marrero:2011:CFG

[MR11] J. Abderramán Marrero and M. Rachidi. Companion factorization in the general linear group GL(n; C) and applications. Linear Algebra and its Applications, 434(5):1261– 1271, March 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Medina:2012:ECW

[MR12] Luis Medina and Oscar Rojo. Eigenvalues of certain weighted graphs joined at their roots having cliques at some levels. Linear Algebra and its Applications, 437 (3):878–898, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002546.

Massarenti:2013:RIM

[MR13] Alex Massarenti and Emanuele Raviolo. The rank of n×n matrix multiplication is at least 3n^{2} − 2\sqrt{2}n^{3/2} − 3n. Linear Algebra and its Applications, 438(11):4500–4509, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000955.

See corrigendum [MR14a].

Massarenti:2014:CSR

[MR14a] Alex Massarenti and Emanuele Raviolo. Corrigendum to “The rank of n × n matrix multiplication is at least 3n^{2} − 2\sqrt{2}n^{3/2}3n” [Linear Algebra Appl. 438 (11) (2013) 4500–4509]. Linear Algebra and its Applications, 445(??): 369–371, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008070.

See [MR13].

Moller:2014:TBA

[MR14b] Claudia Möller and Ulrich Reif. A tree-based approach to joint spectral radius determination. Linear Algebra and its Applications, 463 (??):154–170, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005412.

Monsalve:2014:CSN

[MR14c] Marlliny Monsalve and Marcos Raydan. Corrigendum to “A new inversion-free method for a rational matrix equation” [Linear Algebra Appl. 433 (1) (2010) 64–71]. Linear Algebra and its Applications, 448(??):343–344, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000883.

See [MR10b].

Massey:2012:DRS

[MRS12] Pedro G. Massey, Mariano A. Ruiz, and Demetrio Stojanoff. Duality in reconstruction systems. Linear Algebra and its Applications, 436(3):447–464, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100142X.

Maze:2011:NDR

[MRW11] Gérard Maze, Joachim Rosenthal, and Urs Wagner. Natural density of rectangular unimodular integer matrices. Linear Algebra and its Applications, 434(5):1319–1324, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Macaj:2010:SPM

[MŠ10a] Martin Mačaj and Jozef Širáň. Search for properties of the missing Moore graph. Linear Algebra and its Applications, 432(9):2381–2398, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matos:2010:CPR

[MS10b] Isabel T. Matos and Fernando C. Silva. Completion problems for real matrices, II. Linear Algebra and its Applications, 432(1):180–202, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2010:MSP

[MS10c] Lajos Molnár and Patrícia Szokol. Maps on states preserving the relative entropy II. Linear Algebra and its Applications, 432(12):3343–3350, July 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Malyshev:2011:USM

[MS11a] Alexander Malyshev and Miloud Sadkane. Using the Sherman–Morrison–Woodbury inversion formula for a fast solution of tridiagonal block Toeplitz systems. Linear Algebra and its Applications, 435(11):2693–2707, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mishra:2011:NGI

[MS11b] Debasisha Mishra and K. C. Sivakumar. Nonnegative generalized inverses and least elements of polyhedral sets. Linear Algebra and its Applications, 434(12):2448–2455, June 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mishra:2012:GMM

[MS12] D. Mishra and K. C. Sivakumar. Generalizations of matrix monotonicity and their relationships with certain subclasses of proper splittings. Linear Algebra and its Applications, 436(7):2604– 2614, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007592.

Mallik:2013:CGM

[MS13a] Sudipta Mallik and Bryan L. Shader. Classes of graphs with minimum skew rank 4. Linear Algebra and its Applications, 439(11):3643–3657, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005788.

Mansour:2013:PSS

[MS13b] Toufik Mansour and Mark Shattuck. Parity successions in set partitions. Linear Algebra and its Applications, 439 (9):2642–2650, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004758.

Monfared:2013:CMG

[MS13c] Keivan Hassani Monfared and Bryan L. Shader. Construction of matrices with a given graph and prescribed interlaced spectral data. Linear Algebra and its Applications, 438(11):4348–4358, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001006.

Mbekhta:2014:QIA

[MS14a] Mostafa Mbekhta and Laurian Suciu. Quasi-isometries associated to A-contractions. Linear Algebra and its Applications, 459(??):430–453, October 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004662.

Mizuno:2014:SWB

[MS14b] Hirobumi Mizuno and Iwao Sato. Some weighted Bartholdi zeta function of a digraph. Linear Algebra and its Applications, 445(??):1–17, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007817.

Myskova:2014:RFI

[MS14c] H. Mysková and L. Stefanský. Robustness of fuzzy interval circulant-Hankel matrices. Linear Algebra and its Applications, 444(??):165– 182, March 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007714.

Martin:2013:SLS

[MSD13] Antonio J. Calderón Martiń and José M. Sánchez-Delgado. On split Leibniz superalgebras. Linear Algebra and its Applications, 438(12): 4709–4725, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000797.

Malyshev:2011:CTS

[MSP11] Alexander Malyshev, Miloud Sadkane, and Olivier Pourquier. Circular trichotomy of the spectrum of regular matrix pencils. Linear Algebra and its Applications, 435(3):717– 733, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mofrad:2013:SSL

[MSP13] Asieh A. Mofrad, M.-R. Sadeghi, and D. Panario. Solving sparse linear systems of equations over finite fields using bit-flipping algorithm. Linear Algebra and its Applications, 439(7): 1815–1824, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003431.

Mehrmann:2012:IRK

[MSS12] V. Mehrmann, C. Schröder, and V. Simoncini. An implicitly-restarted Krylov subspace method for real symmetric/skew-symmetric eigenproblems. Linear Algebra and its Applications, 436(10):4070–4087, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379509005898.

Mishra:2014:SDN

[MSS14] Alok Mishra, Rajendra Kumar Sharma, and Wagish Shukla. On the self-dual normal bases and their distribution. Linear Algebra and its Applications, 457(??): 179–190, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003164.

Moon:2014:WCD

[MSvdD14] J. W. Moon, Zhisheng Shuai, and P. van den Driessche. Walks and cycles on a digraph with application to population dynamics. Linear Algebra and its Applications, 451(??):182–196, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Meyer:2012:CHT

[MSvW12] Johan Meyer, Jenő Szigeti, and Leon van Wyk. A Cayley–Hamilton trace identity for 2 × 2 matrices over Lie-solvable rings. Linear Algebra and its Applications, 436(7):2578–2582, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100752X.

Moslehian:2011:SNI

[MTS11] Mohammad Sal Moslehian, Masaru Tominaga, and Kichi-Suke Saito. Schatten pnorm inequalities related to an extended operator parallelogram law. Linear Algebra and its Applications, 435 (4):823–829, August 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meerbergen:2012:RIR

[MV12] Karl Meerbergen and Raf Vandebril. A reflection on the implicitly restarted Arnoldi method for computing eigenvalues near a vertical line. Linear Algebra and its Applications, 436(8): 2828–2844, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005325.

Macias-Virgos:2010:SMP

[MVPS10] E. Macías-Virgós and M. J. Pereira-Sáez. Symplectic matrices with predetermined left eigenvalues. Linear Algebra and its Applications, 432(1): 347–350, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Masaro:2010:CWL

[MW10] Joe Masaro and Chi Song Wong. Characterization of Wishart–Laplace distributions via Jordan algebra homomorphisms. Linear Algebra and its Applications, 432 (6):1578–1594, March 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ma:2012:FTS

[MW12a] Jianmin Ma and Kaishun Wang. Fissioned triangular schemes via sharply 3transitive groups. Linear Algebra and its Applications, 436(7):2618–2629, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007701.

Murty:2012:UPG

[MW12b] M. Ram Murty and Junho Peter Whang. The uncertainty principle and a generalization of a theorem of Tao. Linear Algebra and its Applications, 437(1):214–220, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001395.

Ma:2014:NEC

[MW14a] Jianmin Ma and Kaishun Wang. Nonexistence of exceptional 5-class association schemes with two Qpolynomial structures. Linear Algebra and its Applications, 440(??):278–285, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006988.

Mason:2014:ENP

[MW14b] Oliver Mason and Fabian Wirth. Extremal norms for positive linear inclusions. Linear Algebra and its Applications, 444(??):100–113, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007180.

Minchenko:2014:SMR

[MW14c] Marsha Minchenko and Ian M. Wanless. Spectral moments of regular graphs in terms of subgraph counts. Linear Algebra and its Applications, 446(??):166–176, April 1, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008367.

Ma:2013:PNIb

[MWZ13] Xiaobin Ma, Dein Wong, and Min Zhu. The positive and the negative inertia index of line graphs of trees. Linear Algebra and its Applications, 439(10):3120–3128, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005387.

McKee:2014:TBP

[MY14] James McKee and Pavlo Yatsyna. A trace bound for positive definite connected integer symmetric matrices. Linear Algebra and its Applications, 444(??):227–230, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007702.

Ma:2013:PNIa

[MYL13] Haicheng Ma, Wenhua Yang, and Shenggang Li. Positive and negative inertia index of a graph. Linear Algebra and its Applications, 438(1):331–341, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005551.

Myskova:2012:IMP

[Mys12] Helena Mysková. Interval max-plus systems of linear equations. Linear Algebra and its Applications, 437(8): 1992–2000, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003539.

Meng:2010:OPB

[MZ10] Lingsheng Meng and Bing Zheng. The optimal perturbation bounds of the Moore– Penrose inverse under the Frobenius norm. Linear Algebra and its Applications, 432(4):956–963, February 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcellan:2011:DPS

[MZ11] F. Marcellán and S. M. Zagorodnyuk. Density of polynomials in some L^{2} spaces on radial rays in the complex plane. Linear Algebra and its Applications, 435(1):128–146, July 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marcaida:2012:HBO

[MZ12a] S. Marcaida and I. Zaballa. On a homeomorphism between orbit spaces of linear systems and matrix polynomials. Linear Algebra and its Applications, 436(6):

1664–1682, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002254.

Matthies:2012:SSS

[MZ12b] Hermann G. Matthies and Elmar Zander. Solving stochastic systems with low-rank tensor compression. Linear Algebra and its Applications, 436(10):3819–3838, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003223.

Mo:2012:MSS

[MZ12c] Qun Mo and Xiaosheng Zhuang. Matrix splitting with symmetry and dyadic framelet filter banks over algebraic number fields. Linear Algebra and its Applications, 437(10):2650–2679, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005034.

Ma:2013:ESC

[MZ13] Chao Ma and Xingzhi Zhan. Extremal sparsity of the companion matrix of a polynomial. Linear Algebra and its Applications, 438(1): 621–625, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Ma:2014:IIZ

[MZ14] Chao Ma and Xingzhi Zhan. Inverse invariant zero-nonzero patterns. Linear Algebra and its Applications, 443(??): 184–190, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007088.

Nunez-Alarcon:2013:GOC

[NA13] Daniel Núñez-Alarcón. On the growth of the optimal constants of the multilinear Bohnenblust–Hille inequality. Linear Algebra and its Applications, 439(8): 2494–2499, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004163.

Najafi:2013:SRK

[Naj13] Hamed Najafi. Some results on Kwong functions and related inequalities. Linear Algebra and its Applications, 439(9):2634–2641, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004175.

Nakatsukasa:2010:ARW

[Nak10]  Yuji Nakatsukasa.  Abso-

lute and relative Weyl theorems for generalized eigenvalue problems. Linear Algebra and its Applications, 432 (1):242–248, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakatsukasa:2012:TRC

[Nak12] Yuji Nakatsukasa. The tan θ theorem with relaxed conditions. Linear Algebra and its Applications, 436(5): 1528–1534, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006227.

Nakamura:2013:BPF

[Nak13] Noboru Nakamura. Barbour path functions and related operator means. Linear Algebra and its Applications, 439 (8):2434–2441, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004345.

Natkaniec:2013:ASF

[Nat13] Tomasz Natkaniec. Algebrability of some families of Darboux-like functions. Linear Algebra and its Applications, 439(10):3256–3263, November 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005594.

Nakaoka:2014:SCM

[NCdS14] I. N. Nakaoka, E. L. Monte Carmelo, and O. J. N. T. N. dos Santos. Sharp covering of a module by cyclic submodules. Linear Algebra and its Applications, 458(??): 387–402, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400384X.

Nikolov:2013:RNG

[NCI13] Jovana Nikolov and Dragana S. Cvetković-Ilić. Rennd generalized inverses. Linear Algebra and its Applications, 439(10):2999–3007, November 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004990.

Neogy:2011:SMC

[ND11] S. K. Neogy and A. K. Das. On singular N_{0}-matrices and the class Q. Linear Algebra and its Applications, 434(3): 813–819, February 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nesetril:2013:NFV

[NdM13] J. Nesetril and P. Ossona de Mendez. A note on Fiedler value of classes with sublinear separators. Linear Algebra and its Applications, 439 (8):2216–2221, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300462X.

Nemirovsky:2013:TAM

[Nem13]  Danil Nemirovsky.  Ten-

sor approach to mixed highorder moments of absorbing Markov chains. Linear Algebra and its Applications, 438 (4):1900–1922, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006100.

Netzer:2010:SRN

[Net10] Tim Netzer. On semidefinite representations of nonclosed sets. Linear Algebra and its Applications, 432(12): 3072–3078, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neytcheva:2011:EES

[Ney11] Maya Neytcheva. On elementby-element Schur complement approximations. Linear Algebra and its Applications, 434 (11):2308–2324, June 1, 2011.

Niezgoda:2010:AOC

[Nie10] Marek Niezgoda. Accretive operators and Cassels inequality. Linear Algebra and its Applications, 433(1):136–142, July 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2011:MRC

[Nie11] Marek Niezgoda. Majorization and relative concavity. Linear Algebra and its Applications, 434(8):1968–1980, April 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2012:COG

[Nie12]  Marek Niezgoda.  Cone

orderings, group majorizations and similarly separable vectors. Linear Algebra and its Applications, 436(3): 579–594, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005258.

Niezgoda:2013:SSV

[Nie13] Marek Niezgoda. On sub- and superadditive vector-valued maps with applications to group majorization. Linear Algebra and its Applications, 438(11):4249–4259, June 1, 2013. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001122.

Nikiforov:2010:CZP

[Nik10a] Vladimir Nikiforov. A contribution to the Zarankiewicz problem. Linear Algebra and its Applications, 432(6):1405– 1411, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2010:SRG

[Nik10b] Vladimir Nikiforov. The spectral radius of graphs without paths and cycles of specified length. Linear Algebra and its Applications, 432(9):2243– 2256, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2011:SLS

[Nik11] Vladimir Nikiforov. On the sum of k largest singular values of graphs and matrices. Linear Algebra and its Applications, 435(10):2394–2401, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2013:IMF

[Nik13] V. Nikiforov. The influence of Miroslav Fiedler on spectral graph theory. Linear Algebra and its Applications, 439(4):818–821, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001930.

Nikiforov:2014:AMU

[Nik14] Vladimir Nikiforov. Analytic methods for uniform hypergraphs. Linear Algebra and its Applications, 457 (??):455–535, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002857.

Nitica:2010:SMM

[Nit10] V. Nitica. The structure of max–min hyperplanes. Linear Algebra and its Applications, 432(1):402–429, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ning:2013:SLS

[NLL13] Wenjie Ning, Hao Li, and Mei Lu. On the signless Laplacian spectral radius of irregular graphs. Linear Algebra and its Applications, 438(5):2280–2288, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007343.

Nazari:2014:IEP

[NM14] A. M. Nazari and F. Mahdinasab. Inverse eigenvalue problem of distance matrix via orthogonal matrix. Linear Algebra and its Applications, 450(??):202–216, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000846.

Nemeth:2010:HPI

[NN10] A. B. Németh and S. Z. Németh. How to project onto an isotone projection cone. Linear Algebra and its Applications, 433(1):41–51, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nemeth:2013:LLO

[NN13] A. B. Németh and S. Z. Németh. Lattice-like operations and isotone projection sets. Linear Algebra and its Applications, 439(10):2815– 2828, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005466.

Nelson:2014:DSS

[NNW14] Jelani Nelson, Huy L. Nguyẽn, and David P. Woodruff. On deterministic sketching and streaming for sparse recovery and norm estimation. Linear Algebra and its Applications, 441(??):152–167, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

Ning:2013:DSR

[NOL13] Wenjie Ning, Liangqi Ouyang, and Mei Lu. Distance spectral radius of trees with fixed number of pendent vertices. Linear Algebra and its Applications, 439(8):2240– 2249, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004291.

Nomura:2014:LPH

[Nom14] Kazumasa Nomura. Leonard pairs having LB–TD form. Linear Algebra and its Applications, 455(??):1–21, August 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002511.

Nordstrom:2011:CIM

[Nor11] Kenneth Nordström. Convexity of the inverse and Moore– Penrose inverse. Linear Algebra and its Applications, 434(6):1489–1512, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nordgren:2012:PSM

[Nor12] Ronald P. Nordgren. On properties of special magic square matrices. Linear Algebra and its Applications, 437 (8):2009–2025, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004223.

Nordgren:2014:SOS

[Nor14] Ronald P. Nordgren. Spectra of order-4 special magic squares. Linear Algebra and its Applications, 458 (??):80–98, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003401.

Neubauer:2010:OWD

[NP10] Michael G. Neubauer and Richard G. Pace. D-optimal (0, 1)-weighing designs for eight objects. Linear Algebra and its Applications, 432(10): 2634–2657, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nath:2012:DSR

[NP12] Milan Nath and Somnath Paul. On the distance spectral radius of bipartite graphs. Linear Algebra and its Applications, 436(5): 1285–1296, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Narayanan:2013:SAL

[NP13a] H. Narayanan and H. Priyadarshan. A subspace approach to linear dynamical systems. Linear Algebra and its Applications, 438 (9):3576–3599, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000402.

Neumann:2013:REP

[NP13b] M. Neumann and S. Pati. On reciprocal eigenvalue property of weighted trees. Linear Algebra and its Applications, 438(10):3817–3828, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006501.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Nath:2014:DLS

[NP14] Milan Nath and Somnath Paul. On the distance Laplacian spectra of graphs. Linear Algebra and its Applications, 460(??):97–110, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004753.

Neumann:2013:SCN

[NPP13] Michael Neumann, J. M. Peña, and Olga Pryporova. Some classes of nonsingular matrices and applications. Linear Algebra and its Applications, 438(4):1936– 1945, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007427.

Nazari:2010:CAN

[NR10] A. M. Nazari and D. Rajabi. Computational aspect to the nearest matrix with two prescribed eigenvalues. Linear Algebra and its Applications, 432(1):1–4, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Narayan:2011:RSF

[NRR^{+}11] Sivaram K. Narayan, Eileen L. Radzwion, Sara P. Rimer, Rachael L. Tomasino, Jennifer L. Wolfe, and Andrew M. Zimmer. Robustness and surgery of frames. Linear Algebra and its Applications, 434(8):1893–1901, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neuman:2012:IRR

[NRS12] A. Neuman, L. Reichel, and H. Sadok. Implementations of range restricted iterative methods for linear discrete ill-posed problems. Linear Algebra and its Applications, 436(10):3974–3990, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510004416.

Nastase:2011:MSF

[NS11a] Esmeralda L. Nāstase and Papa A. Sissokho. The minimum size of a finite subspace partition. Linear Algebra and its Applications, 435 (6):1213–1221, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neumann:2011:IMF

[NS11b] Michael Neumann and Nung-Sing Sze. On the inverse mean first passage matrix problem and the inverse M-matrix problem. Linear Algebra and its Applications, 434(7):1620– 1630, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nitica:2011:IVS

[NS11c] Viorel Nitica and Sergeĭ Sergeev. An interval version of separation by semispaces in max–min convexity. Linear Algebra and its Applications, 435(7):1637–1648, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nazari:2012:IEP

[NS12a] A. M. Nazari and F. Sherafat. On the inverse eigenvalue problem for nonnegative matrices of order two to five. Linear Algebra and its Applications, 436(7):1771–1790, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008597.

Nozaki:2012:GCS

[NS12b] Hiroshi Nozaki and Masashi Shinohara. A geometrical characterization of strongly regular graphs. Linear Algebra and its Applications, 437 (10):2587–2600, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005010.

Nozaki:2012:CSH

[NS12c] Hiroshi Nozaki and Sho Suda. A characterization of skew Hadamard matrices and doubly regular tournaments. Linear Algebra and its Applications, 437(3): 1050–1056, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002649.

Nair:2013:AMS

[NS13] Reshmi Nair and Bryan L. Shader. Acyclic matrices with a small number of distinct eigenvalues. Linear Algebra and its Applications, 438(10):4075–4089, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006532.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Ngo:2014:VCN

[NS14] Nham V. Ngo and Klemen Sivic. On varieties of commuting nilpotent matrices. Linear Algebra and its Applications, 452(??):237–262, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001761.

Nina:2013:JCF

[NSC13] Hans Nina, Ricardo L. Soto, and Domingos M. Cardoso. The Jordan canonical form for a class of weighted directed graphs. Linear Algebra and its Applications, 438(1):261–268, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005976.

Nakatsukasa:2013:MAG

[NSW13] Yuji Nakatsukasa, Naoki Saito, and Ernest Woei. Mysteries around the graph Laplacian eigenvalue 4. Linear Algebra and its Applications, 438(8):3231–3246, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008671.

Nomura:2010:STP

[NT10a] Kazumasa Nomura and Paul Terwilliger. On the shape of a tridiagonal pair. Linear Algebra and its Applications, 432(2–3):615–636, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2010:TPR

[NT10b] Kazumasa Nomura and Paul Terwilliger. Tridiagonal pairs of q-Racah type and the μconjecture. Linear Algebra and its Applications, 432(12): 3201–3209, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Napp:2011:LQC

[NT11a] D. Napp and H. L. Trentelman. Linear-quadratic control and quadratic differential forms for multidimensional behaviors. Linear Algebra and its Applications, 434 (1):117–130, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2011:TMN

[NT11b] Kazumasa Nomura and Paul Terwilliger. Tridiagonal matrices with nonnegative entries. Linear Algebra and its Applications, 434(12):2527– 2538, June 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Netzer:2012:PDR

[NT12a] Tim Netzer and Andreas Thom. Polynomials with and without determinantal representations. Linear Algebra and its Applications, 437(7): 1579–1595, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200331X.

Nomura:2012:KPL

[NT12b] Kazumasa Nomura and Paul Terwilliger. Krawtchouk polynomials, the Lie algebra ∫↕_{∈}, and Leonard pairs. Linear Algebra and its Applications, 437(1):345–375, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200136X.

Needell:2014:PGI

[NT14] Deanna Needell and Joel A. Tropp. Paved with good intentions: Analysis of a randomized block Kaczmarz method. Linear Algebra and its Applications, 441(??): 199–221, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000098.

Niendorf:2010:DHD

[NV10] V. Niendorf and H. Voss. Detecting hyperbolic and definite matrix polynomials. Linear Algebra and its Applications, 432(4):1017–1035, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Noutsos:2012:PFT

[NV12] Dimitrios Noutsos and Richard S. Varga. On the Perron– Frobenius theory for complex matrices. Linear Algebra and its Applications, 437(4):1071–1088, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200256X.

Nikiforov:2013:MNG

[NY13] Vladimir Nikiforov and Xiying Yuan. Maximum norms of graphs and matrices, and their complements. Linear Algebra and its Applications, 439(5):1538–1549, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Nikiforov:2014:MEP

[NY14] Vladimir Nikiforov and Xiying Yuan. More eigenvalue problems of Nordhaus-Gaddum type. Linear Algebra and its Applications, 451 (??):231–245, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001682.

ODorney:2014:MCT

[O’D14] Evan O’Dorney. Minimizing the Cayley transform of an orthogonal matrix by multiplying by signature matrices. Linear Algebra and its Applications, 448(??):97–103, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000615.

Oliveira:2010:BSG

[OdLdAK10] Carla Silva Oliveira, Leonardo Silva de Lima, Nair Maria Maia de Abreu, and Steve Kirkland. Bounds on the Q-spread of a graph. Linear Algebra and its Applications, 432(9):2342– 2351, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ogus:2013:EMW

[Ogu13] Arthur Ogus. The eigenvalues of matrices which commute with their derivative. Linear Algebra and its Applications, 438(12):4757–4759, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001596.

Dokovic:2012:GMT

[– Dok12] Dragomir Z. – Doković. Generalization of Mirsky’s theorem on diagonals and eigenvalues of matrices. Linear Algebra and its Applications, 437(10):2680–2682, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005046.

Ou:2014:EKR

[OLW14] Li Ou, Benjian Lv, and Kaishun Wang. The Erdős– Ko–Rado theorem for singular linear spaces. Linear Algebra and its Applications, 440(??):206–212, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006824.

Oseledets:2010:FOK

[OM10] Ivan Oseledets and Ekaterina Muravleva. Fast orthogonalization to the kernel of the discrete gradient operator with application to Stokes problem. Linear Algebra and its Applications, 432(6):1492– 1500, March 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oseledets:2012:RQM

[OM12] I. V. Oseledets and A. Yu. Mikhalev. Representation of quasiseparable matrices using excluded sums and equivalent charges. Linear Algebra and its Applications, 436(3):699–708, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005866.

Ogura:2013:GJS

[OM13] M. Ogura and C. F. Martin. Generalized joint spectral radius and stability of switching systems. Linear Algebra and its Applications, 439 (8):2222–2239, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004278.

Ogura:2014:LFJ

[OM14] M. Ogura and C. F. Martin. A limit formula for joint spectral radius with p-radius of probability distributions. Linear Algebra and its Applications, 458(??):605–625, October 1, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004030.

Omladic:2012:SAS

[OR12] Matjaz Omladic and Heydar Radjavi. Self-adjoint semigroups with nilpotent commutators. Linear Algebra and its Applications, 436(7): 2597–2603, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007385.

Oblak:2014:GAA

[OS14] Polona Oblak and Helena Smigoc. Graphs that allow all the eigenvalue multiplicities to be even. Linear Algebra and its Applications, 454(??):72–90, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514002316.

Olshevsky:2010:GM

[OSZ10] V. Olshevsky, G. Strang, and P. Zhlobich. Green’s matrices. Linear Algebra and its Applications, 432(1):218–241, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oseledets:2010:TCA

[OT10] Ivan Oseledets and Eugene Tyrtyshnikov. TT-cross approximation for multidimensional arrays. Linear Algebra and its Applications, 432(1): 70–88, January 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Osaka:2012:DPS

[OT12] Hiroyuki Osaka and Jun Tomiyama. Double piling structure of matrix monotone functions and of matrix convex functions II. Linear Algebra and its Applications, 437(3):735–748, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002054.

Olesky:2012:SPN

[OTdDv12] D. D. Olesky, M. J. Tsatsomeros, and P. den Driessche van. Sign patterns with a nest of positive principal minors. Linear Algebra and its Applications, 436(12): 4392–4399, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002953.

Otopal:2012:RKC

[Oto12] Nina Otopal. Restricted kernel canonical correlation analysis. Linear Algebra and its Applications, 437(1):1– 13, July 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001449.

Okoh:2010:RHC

[OZ10] Frank Okoh and Frank Zorzitto. The realization of hyperelliptic curves through endomorphisms of Kronecker modules. Linear Algebra and its Applications, 432(5):1189– 1217, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ozdemir:2013:SSS

[Özd13] Yunus Özdemir. Spaces of skew-symmetric matrices satisfying A^{3} = λA. Linear Algebra and its Applications, 438(3):1365–1371, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007173.

Pagacz:2012:WTD

[Pag12] Patryk Pagacz. On Woldtype decomposition. Linear Algebra and its Applications, 436(9):3065–3071, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006331.

Palfia:2013:WMM

[Pál13] Miklós Pálfia. Weighted matrix means and symmetrization procedures. Linear Algebra and its Applications, 438 (4):1746–1768, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100485X.

Paniello:2011:SMA

[Pan11] Irene Paniello. Stochastic matrices arising from genetic inheritance. Linear Algebra and its Applications, 434(3):791– 800, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2012:DMD

[Pan12a] Zhidong Pan. Derivable maps and derivational points. Linear Algebra and its Applications, 436(11):4251–4260, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512000912.

Pankov:2012:EGG

[Pan12b] Mark Pankov. Embeddings of Grassmann graphs. Linear Algebra and its Applications, 436(9):3413–3424, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007798.

Petrovic:2011:FRL

[PAS11a] Miroslav Petrović, Tatjana Aleksić, and Slobodan Simić. Further results on the least eigenvalue of connected graphs. Linear Algebra and its Applications, 435(9):2303–2313, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petrovic:2011:LEC

[PAS11b] Miroslav Petrović, Tatjana Aleksić, and Višnja Simić. On the least eigenvalue of cacti. Linear Algebra and its Applications, 435(10):2357–2364, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:2010:LBI

[Pat10] Thomas H. Pate. Lower bound inequalities for norms of symmetrized tensor powers. Linear Algebra and its Applications, 432(1):116–133, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:2012:HTM

[Pat12a] Thomas H. Pate. Hölder type matrix inequalities of Pate, Blakley, and Roy extended to the inner product of Frobenius. Linear Algebra and its Applications, 436(6):1763–1769, March 15, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006525.

Patricio:2012:MPI

[Pat12b] Pedro Patrício. The Moore– Penrose inverse of a companion matrix. Linear Algebra and its Applications, 437 (3):870–877, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002509.

Porras:2010:CCC

[PC10] José María Muñoz Porras and José Ignacio Iglesias Curto. Classification of convolutional codes. Linear Algebra and its Applications, 432(10):2701– 2725, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2012:SMS

[PCC12] Shaohua Pan, Yungyen Chiang, and Jein-Shan Chen. SOC-monotone and SOCconvex functions vs. matrixmonotone and matrix-convex functions. Linear Algebra and its Applications, 437(5):1264– 1284, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003060.

Patuzzi:2014:ISC

[PdFDV14] Laura Patuzzi, Maria Aguieiras A. de Freitas, and Renata R. Del-Vecchio. Indices for special classes of trees. Linear Algebra and its Applications, 442(??):106–114, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004588.

Potra:2013:CBA

[PE13] Florian A. Potra and Hans Engler. A characterization of the behavior of the Anderson acceleration on linear problems. Linear Algebra and its Applications, 438 (3):1002–1011, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006738.

Pellegrino:2014:SCA

[Pel14]  Daniel Pellegrino.  Sharp

coincidences for absolutely summing multilinear operators. Linear Algebra and its Applications, 440(??): 188–196, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006733.

Pena:2014:OTA

[Peñ14] J. M. Peña. An optimal test for almost strict total positivity. Linear Algebra and its Applications, 448 (??):274–284, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000603.

Peperko:2011:CGS

[Pep11] Aljoša Peperko. On the continuity of the generalized spectral radius in max algebra. Linear Algebra and its Applications, 435(4):902–907, August 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peperko:2012:BGJ

[Pep12] Aljosa Peperko. Bounds on the generalized and the joint spectral radius of Hadamard products of bounded sets of positive operators on sequence spaces. Linear Algebra and its Applications, 437(1):189–201, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001681.

Pereira:2011:BLR

[Per11] Rajesh Pereira. Bijective linear rank preservers for spaces of matrices over antinegative semirings. Linear Algebra and its Applications, 435(7):1666– 1671, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peretz:2012:CAS

[Per12] Y. Peretz. A characterization of all the static stabilizing controllers for LTI systems. Linear Algebra and its Applications, 437(2):525–548, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001978.

PereiradaSilvaeSilva:2013:GIE

[Per13] Diogo Diniz Pereira da Silva e Silva. On the graded identities for elementary gradings in matrix algebras over infinite fields. Linear Algebra and its Applications, 439 (5):1530–1537, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003091.

Peretz:2014:ASF

[Per14] Y. Peretz. On applications of Schauder’s fixed-point theorem for the solution of the non-symmetric algebraic Riccati equation. Linear Algebra and its Applications, 445 (??):29–55, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008203.

Petravchuk:2010:PCD

[Pet10] Anatoliy P. Petravchuk. On pairs of commuting derivations of the polynomial ring in one or two variables. Linear Algebra and its Applications, 433(3):574–579, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puchala:2011:PNR

[PGM^{+}11] Zbigniew Puchała, Piotr Gawron, Jarosław Adam Miszczak, Łukasz Skowronek, Man-Duen Choi, and Karol Życzkowski. Product numerical range in a space with tensor product structure. Linear Algebra and its Applications, 434(1):327–342, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Patricio:2012:DIP

[PH12] P. Patrício and R. E. Hartwig. The (2, 2, 0) Drazin inverse problem. Linear Algebra and its Applications, 437(11): 2755–2772, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005150.

Pokorny:2013:RQI

[PHS13] Milan Pokorný, Pavel Híc, and Dragan Stevanović. Remarks on Q-integral complete multipartite graphs. Linear Algebra and its Applications, 439(7):2029–2037, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003856.

Pan:2010:APM

[PIM^{+}10] Victor Y. Pan, Dmitriy Ivolgin, Brian Murphy, Rhys Eric Rosholt, Yuqing Tang, and Xiaodong Yan. Additive preconditioning for matrix computations. Linear Algebra and its Applications, 432(4):1070– 1089, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pinter:2011:ADT

[Pin11] Miklós Pintér. Algebraic duality theorems for infinite LP problems. Linear Algebra and its Applications, 434(3):688– 693, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pinkus:2012:BRM

[Pin12] A. Pinkus. On best rank n matrix approximations. Linear Algebra and its Applications, 437(9):2179–2199, November 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003680.

Pituk:2011:LBP

[Pit11] Mihály Pituk. A link between the Perron–Frobenius theorem and Perron’s theorem for difference equations. Linear Algebra and its Applications, 434(2):490–500, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Protasov:2013:LUB

[PJ13] V. Yu. Protasov and R. M. Jungers. Lower and upper bounds for the largest Lyapunov exponent of matrices. Linear Algebra and its Applications, 438(11): 4448–4468, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300089X.

Phat:2012:LAE

[PKR12] V. N. Phat, Y. Khongtham, and K. Ratchagit. LMI approach to exponential stability of linear systems with interval time-varying delays. Linear Algebra and its Applications, 436(1):243–251, January 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005283.

Pan:2012:SMQ

[PLL12] Xiang-Feng Pan, Xiuguo Liu, and Huiqing Liu. On the spectral moment of quasitrees. Linear Algebra and its Applications, 436(5):927– 934, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003405.

Papez:2014:DDA

[PLS14] J. Papez, J. Liesen, and Z. Strakos. Distribution of the discretization and algebraic error in numerical solution of partial differential equations. Linear Algebra and its Applications, 449(??):89–114, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000767.

Pastravanu:2010:DSI

[PM10] Octavian Pastravanu and Mihaela-Hanako Matcovschi. Diagonal stability of interval matrices and applications. Linear Algebra and its Applications, 433(8–10):1646–1658, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parraguez:2010:CVS

[PO10] Marcela Parraguez and Asuman Oktaç. Construction of the vector space concept from the viewpoint of APOS theory. Linear Algebra and its Applications, 432(8):2112–2124, April 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Park:2011:IWT

[PO11] Sungwoo Park and Dianne P. O’Leary. Implicitly-weighted total least squares. Linear Algebra and its Applications, 435(3):560–577, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Polderman:2012:TDD

[Pol12] J. W. Polderman. Timedomain description of behaviors over finite. Linear Algebra and its Applications, 436(5):1258–1266, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006173.

Poloni:2013:QVE

[Pol13] Federico Poloni. Quadratic vector equations. Linear Algebra and its Applications, 438 (4):1627–1644, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004484.

Poncet:2011:DTM

[Pon11] Paul Poncet. A decomposition theorem for maxitive measures. Linear Algebra and its Applications, 435(7):1672– 1680, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Popovych:2010:PSQ

[Pop10] Stanislav Popovych. Positive semidefinite quadratic forms on unitary matrices. Linear Algebra and its Applications, 433(1):164–171, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Popovici:2012:NEP

[Pop12] Dan Popovici. Norm equalities in pre-Hilbert C^{∗}- modules. Linear Algebra and its Applications, 436 (1):59–70, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004642.

Popov:2013:MSB

[Pop13a] Alexey I. Popov. On matrix semigroups bounded above and below. Linear Algebra and its Applications, 438 (11):4439–4447, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000980.

Popovych:2013:LOO

[Pop13b] Dmytro R. Popovych. Lieorthogonal operators. Linear Algebra and its Applications, 438(5):2090–2106, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007756.

Popovych:2014:CNU

[Pop14] Dmytro R. Popovych. Contractions with necessarily unbounded matrices. Linear Algebra and its Applications, 458(??):689–698, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003826.

Park:2011:CNS

[PP11a] Jin-Woo Park and Sung-Soo Pyo. Construction and nonnegativity of sign idempotent sign pattern matrices. Linear Algebra and its Applications, 435(11):2860–2869, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Poritz:2011:EPS

[PP11b] Alan B. Poritz and Jonathan A. Poritz. On entropy-preserving stochastic averages. Linear Algebra and its Applications, 434(6):1425–1443, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Piwowarczyk:2012:HPP

[PP12a] Kamila Piwowarczyk and Marek Ptak. On the hyperreflexivity of power partial isometries. Linear Algebra and its Applications, 437 (2):623–629, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001383.

Priyadarshan:2012:PCP

[PP12b] H. Priyadarshan and Harish K. Pillai. On “P ” property and the column-W property. Linear Algebra and its Applications, 436 (7):1969–1989, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007312.

Parton:2013:SMT

[PP13a] Maurizio Parton and Paolo Piccinni. Spheres with more than 7 vector fields: All the fault of Spin(9). Linear Algebra and its Applications, 438(3):1113–1131, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006684.

Puerta:2013:LSI

[PP13b] Ferran Puerta and Xavier Puerta. On the Lipschitz stability of (A, B)-invariant subspaces. Linear Algebra and its Applications, 438(1): 182–190, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006143.

Palfia:2014:WMO

[PP14] Miklós Pálfia and Dénes Petz. Weighted multivariable operator means of positive definite operators. Linear Algebra and its Applications, 463 (??):134–153, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005692.

Park:2013:MSI

[PPK13] Woongbae Park, Boram Park, and Suh-Ryung Kim. A matrix sequence {Γ(A^{m})}_{m=1}^{∞} might converge even if the matrix A is not primitive. Linear Algebra and its Applications, 438(5):2306–2319, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007227.

Popa:2012:KPI

[PPKR12] C. Popa, T. Preclik, H. Köstler, and U. Rüde. On Kaczmarz’s projection iteration as a direct solver for linear least squares problems. Linear Algebra and its Applications, 436(2):389–404, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511001297.

Paige:2014:AAP

[PPZ14] Christopher C. Paige, Ivo Panayotov, and Jens-Peter M. Zemke. An augmented analysis of the perturbed twosided Lanczos tridiagonalization process. Linear Algebra and its Applications, 447 (??):119–132, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003352.

Pan:2010:RPH

[PQ10] Victor Y. Pan and Guoliang Qian. Randomized preprocessing of homogeneous linear systems of equations. Linear Algebra and its Applications, 432(12):3272–3318, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2012:SLS

[PQ12] Victor Y. Pan and Guoliang Qian. Solving linear systems of equations with randomization, augmentation and aggregation. Linear Algebra and its Applications, 437 (12):2851–2876, December 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200506X.

Pan:2013:RPV

[PQZ13] Victor Y. Pan, Guoliang Qian, and Ai-Long Zheng. Randomized preprocessing versus pivoting. Linear Algebra and its Applications, 438 (4):1883–1899, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100187X.

Pan:2011:MCP

[PQZC11] Victor Y. Pan, Guoliang Qian, Ai-Long Zheng, and Zhao Chen. Matrix computations and polynomial root-finding with preprocessing. Linear Algebra and its Applications, 434(4):854–879, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Palacios:2010:SRH

[PR10] José Luis Palacios and José M. Renom. Sum rules for hitting times of Markov chains. Linear Algebra and its Applications, 433(2):491–497, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pereira:2013:WTS

[PR13a] Rajesh Pereira and Stephen Rush. Wielandt’s theorem, spectral sets and Banach algebras. Linear Algebra and its Applications, 439(4):852–855, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004673.

Perrucci:2013:ZNR

[PR13b] Daniel Perrucci and Marie-Françoise Roy. Zero-nonzero and real-nonreal sign determination. Linear Algebra and its Applications, 439(10):3016– 3030, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300565X.

Protasov:2010:WDS

[Pro10] V. Yu. Protasov. When do several linear operators share an invariant cone? Linear Algebra and its Applications, 433(4):781–789, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pivovarchik:2013:DNI

[PRT13] Vyacheslav Pivovarchik, Natalia Rozhenko, and Christiane Tretter. Dirichlet– Neumann inverse spectral problem for a star graph of Stieltjes strings. Linear Algebra and its Applications, 439 (8):2263–2292, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Popov:2011:NMS

[PRW11] Alexey I. Popov, Heydar Radjavi, and Peter Williamson. Nonnegative matrix semigroups with finite diagonals. Linear Algebra and its Applications, 434(6):1409– 1424, March 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pryporova:2010:CCD

[Pry10] Olga Pryporova. Complex D convergence and diagonal convergence of matrices. Linear Algebra and its Applications, 432(2–3):515–525, January 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pfaffel:2012:LSD

[PS12] Oliver Pfaffel and Eckhard Schlemm. Limiting spectral distribution of a new random matrix model with dependence across rows and columns. Linear Algebra and its Applications, 436(9): 2966–2979, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006240.

Petra:2014:ACR

[PS14a] Stefania Petra and Christoph Schnörr. Average case recovery analysis of tomographic compressive sensing. Linear Algebra and its Applications, 441(??):168–198, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004333.

Polak:2014:CSC

[PS14b] Agnieszka Polak and Daniel Simson. Coxeter spectral classification of almost TPcritical one-peak posets using symbolic and numeric computations. Linear Algebra and its Applications, 445(??):223–255, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008264.

Psarrakos:2012:DBP

[Psa12] Panayiotis J. Psarrakos. Distance bounds for prescribed multiple eigenvalues of matrix polynomials. Linear Algebra and its Applications, 436(11):4107–4119, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000493.

Pula:2011:MPT

[PSW11] Kyle Pula, Seok-Zun Song, and Ian M. Wanless. Minimum permanents on two faces of the polytope of doubly stochastic matrices. Linear Algebra and its Applications, 434(1):232–238, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Philipp:2013:NRN

[PT13a] Friedrich Philipp and Carsten Trunk. The numerical range of non-negative operators in Krein spaces. Linear Algebra and its Applications, 438(5):2542–2556, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007781.

Potts:2013:PEN

[PT13b] Daniel Potts and Manfred Tasche. Parameter estimation for nonincreasing exponential sums by Prony-like methods. Linear Algebra and its Applications, 439(4): 1024–1039, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007665.

Potts:2014:SPI

[PT14] Daniel Potts and Manfred Tasche. Sparse polynomial interpolation in Chebyshev bases. Linear Algebra and its Applications, 441 (??):61–87, January 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001079.

Possani:2010:UMT

[PTPL10] E. Possani, M. Trigueros, J. G. Preciado, and M. D. Lozano. Use of models in the teaching of linear algebra. Linear Algebra and its Applications, 432(8):2125–2140, April 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pulka:2011:MPE

[Puł11] Małgorzata Pułka. On the mixing property and the ergodic principle for nonhomogeneous Markov chains. Linear Algebra and its Applications, 434(6):1475–1488, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Protasov:2012:SNM

[PV12] V. Yu. Protasov and A. S. Voynov. Sets of nonnegative matrices without positive products. Linear Algebra and its Applications, 437 (3):749–765, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002005.

Wang:2014:GSS

[pWlW14] Xue ping Wang and Hui li Wang. The generators of the solution space for a system of inequalities. Linear Algebra and its Applications, 459(??):248–263, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004406.

Pang:2010:EOS

[PY10] Yongfeng Pang and Wei Yang. Elementary operators on strongly double triangle subspace lattice algebras. Linear Algebra and its Applications, 433(8–10):1678– 1685, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pei:2014:EAT

[PYZ14] Genhua Pei, Hongbo Yin, and Shunhua Zhang. Endomorphism algebras of tilting modules over m-replicated algebras. Linear Algebra and its Applications, 448 (??):292–298, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000214.

Pang:2011:CPP

[PZVJ11] Hong-Kui Pang, Ying-Ying Zhang, Seak-Weng Vong, and Xiao-Qing Jin. Circulant preconditioners for pricing options. Linear Algebra and its Applications, 434(11):2325– 2342, June 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qi:2011:CAL

[QCH11] Xiaofei Qi, Jianlian Cui, and Jinchuan Hou. Characterizing additive ξ-Lie derivations of prime algebras by ξ-Lie zero products. Linear Algebra and its Applications, 434(3):669– 682, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qi:2010:CDB

[QH10] Xiaofei Qi and Jinchuan Hou. Characterizations of derivations of Banach space nest algebras: All-derivable points. Linear Algebra and its Applications, 432(12):3183–3200, July 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qi:2013:CLD

[QH13] Xiaofei Qi and Jinchuan Hou. Characterization of Lie derivations on von Neumann algebras. Linear Algebra and its Applications, 438(1):533–548, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200643X.

Qi:2013:SNT

[Qi13] Liqun Qi. Symmetric nonnegative tensors and copositive tensors. Linear Algebra and its Applications, 439 (1):228–238, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002103.

Qi:2014:EOS

[QS14] Liqun Qi and Yisheng Song. An even order symmetric B tensor is positive definite. Linear Algebra and its Applications, 457(??):303– 312, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003176.

Qi:2014:RUH

[QSW14] Liqun Qi, Jia-Yu Shao, and Qun Wang. Regular uniform hypergraphs, s-cycles, spaths and their largest Laplacian H-eigenvalues. Linear Algebra and its Applications, 443(??):215–227, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007052.

Quarez:2010:SSA

[Qua10] Ronan Quarez. Sturm and Sylvester algorithms revisited via tridiagonal determinantal representations. Linear Algebra and its Applications, 433 (6):1082–1100, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Quarez:2012:SDR

[Qua12] Ronan Quarez. Symmetric determinantal representation of polynomials. Linear Algebra and its Applications, 436(9):3642–3660, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200050X.

Quinlan:2011:SMN

[Qui11] Rachel Quinlan. Spaces of matrices without non-zero eigenvalues in their field of definition, and a question of Szechtman. Linear Algebra and its Applications, 434(6):1580– 1587, March 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qiu:2012:CLC

[QY12] Wei Qiu and Weigen Yan. The coefficients of Laplacian characteristic polynomials of graphs. Linear Algebra and its Applications, 436 (7):2474–2479, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Rodriguez-Andrade:2011:ACD

[RAAGAVS11] M. A. Rodríguez-Andrade, G. Aragón-González, J. L. Aragón, and Luis Verde-Star. An algorithm for the Cartan– Dieudonné theorem on generalized scalar product spaces. Linear Algebra and its Applications, 434(5):1238–1254, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rada:2010:LBE

[Rad10] Juan Rada. Lower bounds for the energy of digraphs. Linear Algebra and its Applications, 432(9):2174–2180, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Radjabalipour:2013:RCF

[Rad13] Mehdi Radjabalipour. The rational canonical form via the splitting field. Linear Algebra and its Applications, 439 (8):2250–2255, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004692.

Rafiei:2014:LLV

[Raf14]  A. Rafiei.  Left-looking

version of AINV preconditioner with complete pivoting strategy. Linear Algebra and its Applications, 445 (??):103–126, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007957.

Rahmani:2013:ATM

[Rah13]  Mourad Rahmani.  The

Akiyama–Tanigawa matrix and related combinatorial identities. Linear Algebra and its Applications, 438(1): 219–230, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006131.

Rahimi-Alangi:2014:MLT

[RAY14] M. Rahimi-Alangi and Bamdad R. Yahaghi. On modules of linear transformations. Linear Algebra and its Applications, 445(??):127–137, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008008.

Rajkovic:2012:SOP

[RBP12] Predrag M. Rajković, Paul Barry, and Marko D. Petković. Sobolev orthogonal polynomials in computing of Hankel determinants. Linear Algebra and its Applications, 437 (10):2417–2428, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Rakic:2014:GMP

[RDD14] Dragan S. Rakić, Nebojsa C. Dincić, and Dragan S. Djordjević. Group, Moore–Penrose, core and dual core inverse in rings with involution. Linear Algebra and its Applications, 463(??):115–133, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005874.

Rande:2011:LPR

[RdSP11] Bernard Randé and Clément de Seguins Pazzis. The linear preservers of real diagonalizable matrices. Linear Algebra and its Applications, 435(6):1257–1266, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rezghi:2011:DTC

[RE11] Mansoor Rezghi and Lars Eldén. Diagonalization of tensors with circulant structure. Linear Algebra and its Applications, 435(3):422–447, August 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reff:2012:SPC

[Ref12] Nathan Reff. Spectral properties of complex unit gain graphs. Linear Algebra and its Applications, 436(9):

3165–3176, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951100718X.

Regalia:2013:MCA

[Reg13] Phillip A. Regalia. Monotonically convergent algorithms for symmetric tensor approximation. Linear Algebra and its Applications, 438(2): 875–890, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007300.

Rehkopf:2010:STD

[Reh10] Edward Rehkopf. Sharpening the triple diagonal form. Linear Algebra and its Applications, 433(2):342–347, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rehkopf:2011:CTD

[Reh11] Edward Rehkopf. Combining triple diagonal forms. Linear Algebra and its Applications, 434(3):785–790, February 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reichel:2011:RMM

[Rei11a]  Lothar Reichel.  Review

of “Matrices, Moments and Quadrature with Applications” by G. H. Golub and G. Meurant, Princeton University Press, 2010, xi + 363 pages. Linear Algebra and its Applications, 434(1):379– 380, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reis:2011:LEE

[Rei11b] Timo Reis. Lur’e equations and even matrix pencils. Linear Algebra and its Applications, 434(1):152–173, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rezaei:2013:CHG

[Rez13] H. Rezaei. On the convex hull generated by orbit of operators. Linear Algebra and its Applications, 438(11):4190–4203, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000943.

Rada:2013:EDH

[RGC13] Juan Rada, Ivan Gutman, and Roberto Cruz. The energy of directed hexagonal systems. Linear Algebra and its Applications, 439(7): 1825–1833, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300342X.

Rhee:2010:NFI

[Rhe10] Noah H. Rhee. Note on functional iteration technique for M / G / 1 type Markov chains. Linear Algebra and its Applications, 432(4):1042– 1048, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rhodes:2010:CPK

[Rho10] John A. Rhodes. A concise proof of Kruskal’s theorem on tensor decomposition. Linear Algebra and its Applications, 432(7):1818–1824, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rimas:2012:EEP

[Rim12] Jonas Rimas. Explicit expression for powers of tridiagonal 2-Toeplitz matrix of odd order. Linear Algebra and its Applications, 436 (9):3493–3506, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008615.

Robbiano:2010:IBL

[RJ10] María Robbiano and Raúl Jiménez. Improved bounds for the Laplacian energy of Bethe trees. Linear Algebra and its Applications, 432(9):2222– 2229, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-

Rojo:2011:LGC

[RJ11] Oscar Rojo and Raúl D. Jiménez. Line graph of combinations of generalized Bethe trees: Eigenvalues and energy. Linear Algebra and its Applications, 435(10):2402–2419, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rad:2011:LBA

[RJH11] Ali Ajdari Rad, Mahdi Jalili, and Martin Hasler. A lower bound for algebraic connectivity based on the connectiongraph-stability method. Linear Algebra and its Applications, 435(1):186–192, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rivero:2013:EKT

[RL13a] Abdon Eddy Choque Rivero and Andreas Lasarow. The Eneström–Kakeya theorem encounters the theory of orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 439 (5):1258–1285, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002528.

Rojo:2013:SUB

[RL13b] Oscar Rojo and Eber Lenes. A sharp upper bound on the incidence energy of graphs in terms of connectivity. Linear Algebra and its Applications, 438(3):1485–1493, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006775.

Rojo:2010:SCS

[RM10] Oscar Rojo and Luis Medina. Spectral characterization of some weighted rooted graphs with cliques. Linear Algebra and its Applications, 433(7):1388–1409, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rozikov:2014:CEA

[RM14] U. A. Rozikov and Sh. N. Murodov. Chain of evolution algebras of “chicken” population. Linear Algebra and its Applications, 450 (??):186–201, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400130X.

Rojo:2010:ACS

[RMAJ10] Oscar Rojo, Luis Medina, Nair Abreu, and Claudia Justel. On the algebraic connectivity of some caterpillars: a sharp upper bound and a total ordering. Linear Algebra and its Applications, 432(2– 3):586–605, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ramos:2014:ORM

[RMP14] C. Correia Ramos, Nuno Martins, and Paulo R. Pinto. Orbit representations from matrices. Linear Algebra and its Applications, 453 (??):44–58, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001955.

Ram:2011:PPP

[RMT11] Y. M. Ram, J. E. Mottershead, and M. G. Tehrani. Partial pole placement with time delay in structures using the receptance and the system matrices. Linear Algebra and its Applications, 434(7): 1689–1696, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodriguez-Ordonez:2010:HCB

[RO10] Hugo Rodríguez-Ordóñez. Homotopy classification of bilinear maps related to octonion polynomial multiplications. Linear Algebra and its Applications, 432(12):3117– 3131, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2011:RLP

[Rod11] Leiba Rodman. Remarks on Lipschitz properties of matrix groups actions. Linear Algebra and its Applications, 434(6):1513–1524, March 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2012:BRB

[Rod12a] Leiba Rodman. Book review: Matrix Completions, Moments, and Sums of Hermitian Squares by Mihály Bakonyi and Hugo J. Woerdeman, Princeton Series in Applied Mathematics, Princeton University Press (2011). xii + 518 pp., Hardback, ISBN 978-0-691-12889-4. Linear Algebra and its Applications, 436(7):2711–2715, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006471.

See [BW11].

Rodman:2012:LPS

[Rod12b] Leiba Rodman. Lipschitz properties of structure preserving matrix perturbations. Linear Algebra and its Applications, 437(7):1503–1537, October 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Rohn:2011:ACH

[Roh11] Jiri Rohn. An algorithm for computing the hull of the solution set of interval linear equations. Linear Algebra and its Applications, 435(2):193–201, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2011:LGE

[Roj11]  Oscar Rojo.  Line graph

eigenvalues and line energy of caterpillars. Linear Algebra and its Applications, 435(8): 2077–2086, October 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Romo:2014:CFP

[Rom14] Fernando Pablos Romo. Classification of finite potent endomorphisms. Linear Algebra and its Applications, 440(??):266–277, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006848.

Rosenthal:2012:WCB

[Ros12a] Daniel Rosenthal. Words containing a basis for the algebra of all matrices. Linear Algebra and its Applications, 436(7):2615–2617, April 1, 2012. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100632X.

Rosmanis:2012:FSP

[Ros12b] Ansis Rosmanis. Fixed space of positive trace-preserving super-operators. Linear Algebra and its Applications, 437(7):1704–1721, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003394.

Rost:2012:MRS

[Ros12c] Karla Rost. Matrix representations of split Bezoutians. Linear Algebra and its Applications, 436(10): 3904–3918, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510006166.

Rowlinson:2010:MET

[Row10] P. Rowlinson. On multiple eigenvalues of trees. Linear Algebra and its Applications, 432(11):3007–3011, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rowlinson:2011:EMG

[Row11] P. Rowlinson. On eigenvalue multiplicity and the girth of a graph. Linear Algebra and its Applications, 435 (10):2375–2381, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Row:2012:TCZ

[Row12a] Darren D. Row. A technique for computing the zero forcing number of a graph with a cut-vertex. Linear Algebra and its Applications, 436(12):4423–4432, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004174.

Rowlinson:2012:RSC

[Row12b] Peter Rowlinson. Regular star complements in strongly regular graphs. Linear Algebra and its Applications, 436(5):1482–1488, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006033.

Rowlinson:2014:EMC

[Row14a]  Peter Rowlinson.  Eigen-

value multiplicity in cubic graphs. Linear Algebra and its Applications, 444 (??):211–218, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300774X.

Rowlinson:2014:BGC

[Row14b] Peter Rowlinson. On bipartite graphs with complete bipartite star complements. Linear Algebra and its Applications, 458(??):149–160, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003760.

Rowlinson:2014:ISS

[Row14c] Peter Rowlinson. On independent star sets in finite graphs. Linear Algebra and its Applications, 442 (??):82–91, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003972.

Rowlinson:2014:SCC

[Row14d] Peter Rowlinson. Star complements and connectivity in finite graphs. Linear Algebra and its Applications, 442 (??):92–98, February 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004205.

Ransford:2011:PDD

[RR11] Thomas Ransford and Jérémie Rostand. Pseudospectra do not determine norm behavior, even for matrices with only simple eigenvalues. Linear Algebra and its Applications, 435(12):3024–3028, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reff:2012:OHA

[RR12] Nathan Reff and Lucas J. Rusnak. An oriented hypergraphic approach to algebraic graph theory. Linear Algebra and its Applications, 437 (9):2262–2270, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004466.

Rabe:2014:PPP

[RR14] H. Rabe and A. C. M. Ran. A peculiar permutation phenomenon arising from the singular vector entries of a special class of Toeplitz matrices. Linear Algebra and its Applications, 459(??):368–383, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004716.

Rakhimov:2012:DSC

[RRKK12] I. S. Rakhimov, I. M. Rikhsiboev, A. Kh. Khudoyberdiyev, and I. A. Karimjanov. Description of some classes of Leibniz algebras. Linear Algebra and its Applications, 437(9):2209–2227, November 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004491.

Rasajski:2011:MRC

[RRM11] M. Rašajski, Z. Radosavljević, and B. Mihailović. Maximal reflexive cacti with four cycles: The approach via Smith graphs. Linear Algebra and its Applications, 435 (10):2530–2543, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rendel:2013:ING

[RRZ13] Olaf Rendel, Anisa Rizvanolli, and Jens-Peter M. Zemke. IDR: a new generation of Krylov subspace methods? Linear Algebra and its Applications, 439(4): 1040–1061, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008269.

Rodman:2010:LTL

[RŠ10] Leiba Rodman and Peter Šemrl. A localization technique for linear preserver problems. Linear Algebra and its Applications, 433(11– 12):2257–2268, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roca:2012:PPC

[RS12a] A. Roca and F. C. Silva. Pencils with prescribed constant subpencils. Linear Algebra and its Applications, 436 (9):3315–3336, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007634.

Rodman:2012:CLI

[RS12b] Leiba Rodman and Ilya M. Spitkovsky. Compressions of linearly independent selfadjoint operators. Linear Algebra and its Applications, 436 (9):3757–3766, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007373.

Rybalkina:2012:TCC

[RS12c] Tetiana Rybalkina and Vladimir V. Sergeichuk. Topological classification of chains of linear mappings. Linear Algebra and its Applications, 437 (3):860–869, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002571.

Ren:2014:SAE

[RS14a] Guojing Ren and Yuming Shi. Self-adjoint extensions for discrete linear Hamiltonian systems. Linear Algebra and its Applications, 454 (??):1–48, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002304.

Roca:2014:PPC

[RS14b] A. Roca and F. C. Silva. Pencils with prescribed constant subpencils over arbitrary fields. Linear Algebra and its Applications, 445 (??):138–161, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008252.

Ruiz:2014:NSF

[RS14c] César Ruiz and Víctor M. Sánchez. Nonlinear subsets of function spaces and spaceability. Linear Algebra and its Applications, 463(??): 56–67, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005904.

Rodman:2010:CIC

[RSS10] Leiba Rodman, Hakan Seyalioglu, and Ilya M. Spitkovsky. On common invariant cones for families of matrices. Linear Algebra and its Applications, 432(4):911–926, February 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rowlinson:2010:SCR

[RTR10] Peter Rowlinson and Behruz Tayfeh-Rezaie. Star complements in regular graphs: Old and new results. Linear Algebra and its Applications, 432 (9):2230–2242, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ruan:2012:IGR

[Rua12] Hongshun Ruan. An improved Greub–Rheinboldt inequality. Linear Algebra and its Applications, 436(1): 126–134, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004782.

Rubei:2013:CSA

[Rub13] Elena Rubei. On completions of symmetric and antisymmetric block diagonal partial matrices. Linear Algebra and its Applications, 439 (10):2971–2979, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005478.

Rudolf:2012:RDK

[Rud12] Tina Rudolf. Reflexivity defect of kernels of the elementary operators of length 2. Linear Algebra and its Applications, 437(6):1366– 1379, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002996.

Rueda:2013:LSD

[Rue13]  Sonia L. Rueda.  Linear

sparse differential resultant formulas. Linear Algebra and its Applications, 438(11): 4296–4321, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000785.

Rukhin:2014:BES

[Ruk14] Andrew L. Rukhin. Bounds on elementary symmetric functions. Linear Algebra and its Applications, 448 (??):329–342, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000287.

Ragnarsson:2013:BTS

[RV13] Stefan Ragnarsson and Charles F. Van Loan. Block tensors and symmetric embeddings. Linear Algebra and its Applications, 438(2):853–874, January 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Ran:2013:DSE

[RvS13] André C. M. Ran and Jan H. van Schuppen. Distributed state estimation with communication of observations. Linear Algebra and its Applications, 439(3):600–612, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008245.

Roy:2010:EFD

[RW10] Sandip Roy and Yan Wan. An explicit formula for differences between Laplacianeigenvector components using coalesced graphs. Linear Algebra and its Applications, 433(7):1329–1335, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ran:2012:ERO

[RW12] André C. M. Ran and Michal Wojtylak. Eigenvalues of rank one perturbations of unstructured matrices. Linear Algebra and its Applications, 437(2):589–600, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200198X.

Radjavi:2012:IAS

[RY12a] Heydar Radjavi and Bamdad R. Yahaghi. On irreducible algebras spanned by triangularizable matrices. Linear Algebra and its Applications, 436(7):2001– 2007, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007579.

Robinson:2012:DNE

[RY12b] Stephen B. Robinson and Yilin Yang. Discrete nonlinear equations and the Fuci’k Spectrum. Linear Algebra and its Applications, 437(3): 917–931, August 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002479.

Simon:2010:MAC

[SA10] Dan Simon and Jeff Abell. A majorization algorithm for constrained correlation matrix approximation. Linear Algebra and its Applications, 432(5):1152–1164, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soleymani:2014:LPC

[SA14] M. Soleymani and A. Armandnejad. Linear preservers of circulant majorization on R^{n}. Linear Algebra and its Applications, 440(??):286–292, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006757.

Sadkane:2012:LRK

[Sad12] Miloud Sadkane. A low-rank Krylov squared Smith method for large-scale discrete-time Lyapunov equations. Linear Algebra and its Applications, 436(8):2807–2827, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005337.

Sagnol:2011:CSP

[Sag11] Guillaume Sagnol. A class of semidefinite programs with rank-one solutions. Linear Algebra and its Applications, 435(6):1446–1463, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sagnol:2013:SRR

[Sag13] Guillaume Sagnol. On the semidefinite representation of real functions applied to symmetric matrices. Linear Algebra and its Applications, 439 (10):2829–2843, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300520X.

Sahi:2010:NRN

[Sah10] Siddhartha Sahi. A note on the resolvent of a nonnegative matrix and its applications. Linear Algebra and its Applications, 432(10):2524–2528, May 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sanchez:2010:CCA

[San10] Luis A. Sanchez. Convex cones associated to generalized cones in R^{N}. Linear Algebra and its Applications, 433(11–12):2122–2138, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sano:2014:FDO

[San14] Takashi Sano. Fréchet derivatives for operator monotone functions. Linear Algebra and its Applications, 456(??): 88–92, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003601.

Sarkar:2014:WDD

[Sar14] Jaydeb Sarkar. Wold decomposition for doubly commuting isometries. Linear Algebra and its Applications, 445(??):289–301, March 15, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008094.

Sato:2011:NPL

[Sat11] Iwao Sato. New proofs for Levine’s theorems. Linear Algebra and its Applications, 435(5):943–952, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sato:2014:NPB

[Sat14]  Iwao Sato.  New proofs

of Bapat and Sivasubramanian’s theorems. Linear Algebra and its Applications, 448(??):1–10, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000391.

Savostyanov:2012:QRO

[Sav12] Dmitry Savostyanov. QTTrank-one vectors with QTTrank-one and full-rank Fourier images. Linear Algebra and its Applications, 436(9): 3215–3224, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100749X.

Savostyanov:2014:QMV

[Sav14] Dmitry V. Savostyanov. Quasioptimality of maximumvolume cross interpolation of tensors. Linear Algebra and its Applications, 458(??): 217–244, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003711.

Somasundaram:2011:SMM

[SB11] Kiran K. Somasundaram and John S. Baras. Solving multimetric network problems: an interplay between idempotent semiring rules. Linear Algebra and its Applications, 435(7): 1494–1512, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schulz-Baldes:2012:SIT

[SB12] Hermann Schulz-Baldes. Sturm intersection theory for periodic Jacobi matrices and linear Hamiltonian systems. Linear Algebra and its Applications, 436(3):498–515, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005155.

Sari:2011:RRM

[SBM11] Bilal Sari, Olivier Bachelier, and Driss Mehdi. Robust S-regularity of matrix pencils applied to the analysis of descriptor models. Linear Algebra and its Applications, 435(5):923–942, September 1,

Simic:2010:CGF

[SBMT10] Slobodan K. Simić, Francesco Belardo, Enzo Maria Li Marzi, and Dejan V. Tošić. Connected graphs of fixed order and size with maximal index: Some spectral bounds. Linear Algebra and its Applications, 432(9):2361–2372, April 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sburlati:2010:PFD

[Sbu10] Giovanni Sburlati. On prime factors of determinants of circulant matrices. Linear Algebra and its Applications, 432 (1):100–106, January 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stegeman:2010:SBR

[SC10] Alwin Stegeman and Pierre Comon. Subtracting a best rank-1 approximation may increase tensor rank. Linear Algebra and its Applications, 433(7):1276–1300, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Somasunderam:2012:ILS

[SC12] Naveen Somasunderam and Shivkumar Chandrasekaran. On the infinitesimal limits of the Schur complements of tridiagonal matrices. Linear Algebra and its Applications, 436(3):659–681, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005799.

Sorensen:2013:TDB

[SC13] Mikael Sørensen and Pierre Comon. Tensor decompositions with banded matrix factors. Linear Algebra and its Applications, 438(2): 919–941, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007452.

Schiebold:2010:CTD

[Sch10] Cornelia Schiebold. Cauchytype determinants and integrable systems. Linear Algebra and its Applications, 433 (2):447–475, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schomburg:2011:PCD

[Sch11] Bernd Schomburg. On the primitivity of Cayley digraphs. Linear Algebra and its Applications, 434(1):356– 360, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra-Capizzano:2011:ACS

[SCS11] Stefano Serra-Capizzano and Debora Sesana. Approximating classes of sequences: The Hermitian case. Linear Algebra and its Applications, 434 (4):1163–1170, February 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra-Capizzano:2010:EDP

[SCSS10] Stefano Serra-Capizzano, Debora Sesana, and Elizabeth Strouse. The eigenvalue distribution of products of Toeplitz matrices — Clustering and attraction. Linear Algebra and its Applications, 432(10): 2658–2678, May 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soto:2013:NMP

[SDNS13] Ricardo L. Soto, Roberto C. Díaz, Hans Nina, and Mario Salas. Nonnegative matrices with prescribed spectrum and elementary divisors. Linear Algebra and its Applications, 439(11):3591–3604, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005909.

Savas:2013:KTM

[SE13] Berkant Savas and Lars Eldén. Krylov-type methods for tensor computations I. Linear Algebra and its Applications, 438(2): 891–918, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511007877.

Seddighin:2011:SJA

[Sed11] Morteza Seddighin. Slant joint antieigenvalues and antieigenvectors of operators in normal subalgebras. Linear Algebra and its Applications, 434(5):1395–1408, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sego:2010:TSH

[Šeg10] Vedran Šego. Two-sided hyperbolic SVD. Linear Algebra and its Applications, 433(7): 1265–1275, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sego:2014:HSD

[Seg14] Vedran Sego. The hyperbolic Schur decomposition. Linear Algebra and its Applications, 440(??):90–110, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006721.

Seidel:2014:FTB

[Sei14] Markus Seidel. Fredholm theory for band-dominated and related operators: a survey. Linear Algebra and its Applications, 445(??): 373–394, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007970.

Gora:2010:SON

[Sem10] Semigroups Working Group at LAW’08 Kranjska Gora. Semigroups of operators with nonnegative diagonals. Linear Algebra and its Applications, 433(11–12):2080–2087, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seo:2013:GPS

[Seo13] Yuki Seo. Generalized Pólya– Szegő type inequalities for some non-commutative geometric means. Linear Algebra and its Applications, 438 (4):1711–1726, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007336.

Seo:2014:OPM

[Seo14] Yuki Seo. Operator power means due to Lawson–Lim– Pálfia for 1 < t < 2. Linear Algebra and its Applications, 459(??):342–356, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004601.

Sergeev:2011:MAA

[Ser11] Sergeĭ Sergeev. Max-algebraic attraction cones of nonnegative irreducible matrices. Linear Algebra and its Applications, 435(7):1736–1757, October 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sergeev:2013:FPS

[Ser13] Sergeĭ Sergeev. Fiedler–Pták scaling in max algebra. Linear Algebra and its Applications, 439(4):822–829, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512000778.

Seven:2010:QFM

[Sev10] Ahmet I. Seven. Quivers of finite mutation type and skewsymmetric matrices. Linear Algebra and its Applications, 433(6):1154–1169, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seven:2013:MCF

[Sev13] Ahmet I. Seven. Mutation classes of finite type cluster algebras with principal coefficients. Linear Algebra and its Applications, 438(12): 4584–4594, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001560.

Seven:2014:MGS

[Sev14]  Ahmet I. Seven.  Maxi-

mal green sequences of skewsymmetrizable 3 × 3 matrices. Linear Algebra and its Applications, 440(??): 125–130, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006381.

Shao:2010:PDH

[SH10] Junwei Shao and Xiaorong Hou. Positive definiteness of Hermitian interval matrices. Linear Algebra and its Applications, 432(4):970– 979, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2013:DIL

[SH13] Yunfeng Shi and Guolin Hou. On the Drazin inverse of the linear combinations of two idempotents in a complex Banach algebra. Linear Algebra and its Applications, 439 (11):3532–3540, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005879.

Shahryari:2010:RSP

[Sha10a] M. Shahryari. Relative symmetric polynomials. Linear Algebra and its Applications, 433(7):1410–1421, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shankar:2010:T

[Sha10b] Shiva Shankar. TOP. Linear Algebra and its Applications, 433(6):1077–1081, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sharifi:2011:POC

[Sha11] K. Sharifi. The product of operators with closed range in Hilbert C^{∗}-modules. Linear Algebra and its Applications, 435(5):1122–1130, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shahidi:2013:NSC

[Sha13a]  Farruh Shahidi.  Neces-

sary and sufficient conditions for doubly stochasticity of infinite-dimensional quadratic operators. Linear Algebra and its Applications, 438(1): 96–110, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006210.

Shao:2013:GPT

[Sha13b]  Jia-Yu Shao.  A general

product of tensors with applications. Linear Algebra and its Applications, 439(8): 2350–2366, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004618.

Shang:2014:ACM

[Sha14a] Yilun Shang. Average consensus in multi-agent systems with uncertain topologies and multiple time-varying delays. Linear Algebra and its Applications, 459(??): 411–429, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004698.

Shao:2014:FSM

[Sha14b] Meiyue Shao. On the finite section method for computing exponentials of doublyinfinite skew-Hermitian matrices. Linear Algebra and its Applications, 451(??):65– 96, June 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001657.

Shitov:2011:IGM

[Shi11] Yaroslav Shitov. Inequalities for Gondran–Minoux rank and idempotent semirings. Linear Algebra and its Applications, 435(7):1769– 1777, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shinohara:2012:TLM

[Shi12a] Hidehiro Shinohara. Thin Lehman matrices arising from finite groups. Linear Algebra and its Applications, 436(4):850–857, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001819.

Shitov:2012:KRT

[Shi12b] Yaroslav Shitov. On the Kapranov ranks of tropical matrices. Linear Algebra and its Applications, 436 (9):3247–3253, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007403.

Shitov:2012:TMS

[Shi12c] Yaroslav Shitov. On tropical matrices of small factor rank. Linear Algebra and its Applications, 437(11): 2727–2732, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004971.

Shitov:2013:CBM

[Shi13] Yaroslav Shitov. On the complexity of Boolean matrix ranks. Linear Algebra and its Applications, 439 (8):2500–2502, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004321.

Shparlinski:2010:SCQ

[Shp10]  I. E. Shparlinski.  Some

counting questions for matrices with restricted entries. Linear Algebra and its Applications, 432(1):155–160, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shen:2012:MRT

[SHS12] Xiaoling Shen, Yaoping Hou, and Li Sheng. On the minimum rank of the third power of a starlike tree. Linear Algebra and its Applications, 436(12):4503–4511, June 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100629X.

Shao:2010:LSB

[SHZ10] Jin-Liang Shao, Ting-Zhu Huang, and Guo-Feng Zhang. Linear system based approach for solving some related problems of M-matrices. Linear Algebra and its Applications, 432(1):327–337, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Simson:2010:IBF

[Sim10] Daniel Simson. Integral bilinear forms, Coxeter transformations and Coxeter polynomials of finite posets. Linear Algebra and its Applications, 433(4):699–717, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singer:2010:ETF

[Sin10a] Ivan Singer. Elementary topical functions on b-complete semimodules over b-complete idempotent semifields. Linear Algebra and its Applications, 433(11–12):2139–2146, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singh:2010:ETF

[Sin10b] Preeti Singh. Equiangular tight frames and signature sets in groups. Linear Algebra and its Applications, 433 (11–12):2208–2242, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sinkovic:2010:MNO

[Sin10c] John Sinkovic. Maximum nullity of outerplanar graphs and the path cover number. Linear Algebra and its Applications, 432(8):2052–2060, April 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sivic:2012:VCTb

[Siv12a] Klemen Sivic. On varieties of commuting triples II. Linear Algebra and its Applications, 437(2):461–489, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005970.

Sivic:2012:VCTa

[Siv12b] Klemen Sivic. On varieties of commuting triples III. Linear Algebra and its Applications, 437(2):393–460, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005982.

Sivakumar:2013:NGI

[Siv13] K. C. Sivakumar. Nonnegative generalized inverses and certain subclasses of singular Q-matrices. Linear Algebra and its Applications, 438(12):4701–4708, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001018.

Snobl:2010:CSL

[ŠK10] L. Šnobl and D. Karásek. Classification of solvable Lie algebras with a given nilradical by means of solvable extensions of its subalgebras. Linear Algebra and its Applications, 432(7):1836– 1850, March 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sinkovic:2012:MRO

[SK12] John Sinkovic and Mark Kempton. Minimum rank of outerplanar graphs. Linear Algebra and its Applications, 436(9):3701–3720, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000560.

Sharma:2013:RUB

[SK13] R. Sharma and R. Kumar. Remark on upper bounds for the spread of a matrix. Linear Algebra and its Applications, 438(11):4359–4362, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000803.

Seo:2014:CPR

[SK14] Jong-Hyeon Seo and Hyun-Min Kim. Convergence of pure and relaxed Newton methods for solving a matrix polynomial equation arising in stochastic models. Linear Algebra and its Applications, 440(??):34–49, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006782.

Skowronek:2011:CMC

[Sko11] Łukasz Skowronek. Cones with a mapping cone symmetry in the finite-dimensional case. Linear Algebra and its Applications, 435(2):361–370, July 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Skulj:2013:CID

[Sku13] Damjan Skulj. A classification of invariant distributions and convergence of imprecise Markov chains. Linear Algebra and its Applications, 439 (9):2542–2561, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004527.

Sun:2014:RGQ

[SL14] Xiu-Hong Sun and Yuan Li. The range of generalized quantum operations. Linear Algebra and its Applications, 452(??):120–129, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001852.

Slapnicar:2010:RMM

[Sla10] Ivan Slapničar. Review of: Matrix Mathematics: Theory, Facts, and Formulas by Dennis S. Bernstein. Linear Algebra and its Applications, 433 (11–12):2269–2271, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Slowik:2012:LCS

[Slo12a]  Roksana Slowik.  The

lower central series of subgroups of the Vershik–Kerov group. Linear Algebra and its Applications, 436(7):2299– 2310, April 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006318.

Slowik:2012:OPN

[Slo12b] Roksana Slowik. On one property of normal subgroups of UT_{∞}(R). Linear Algebra and its Applications, 437 (9):2300–2307, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512004314.

Slowik:2013:EIM

[Slo13] Roksana Slowik. Expressing infinite matrices as products of involutions. Linear Algebra and its Applications, 438(1):399–404, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005915.

Slowik:2014:PMF

[Slo14] Roksana Slowik. On products of matrices of a fixed order. Linear Algebra and its Applications, 446(??):104– 114, April 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008410.

Shpiz:2013:CES

[SLS13] Grigory B. Shpiz, Grigory L. Litvinov, and Sergei N. Sergeev. On common eigenvectors for semigroups of matrices in tropical and traditional linear algebra. Linear Algebra and its Applications, 439(6):1651–1656, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003170.

Sun:2010:GHO

[SM10a] Shanli Sun and Xuefeng Ma. Generalized Hermitian operators. Linear Algebra and its Applications, 433(4):737– 749, October 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2010:SJA

[SM10b] Shanli Sun and Xuefeng Ma. Solvable Jordan algebras of compact operators. Linear Algebra and its Applications, 432(5):1337–1347, February 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2012:LTD

[SM12] Shanli Sun and Xuefeng Ma. Lie triple derivations of nest algebras on Banach spaces. Linear Algebra and its Applications, 436 (9):3443–3462, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511008354.

Shaked-Monderer:2013:MAM

[SM13] Naomi Shaked-Monderer. Matrices attaining the minimum semidefinite rank of a chordal graph. Linear Algebra and its Applications, 438(10):3804–3816, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006082.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Stosic:2011:CSP

[SMC11] Marko Stošić, Manuel Marques, and João Paulo Costeira.

Convex solution of a permutation problem. Linear Algebra and its Applications, 434 (1):361–369, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singer:2012:TFS

[SN12] Ivan Singer and Viorel Nitica. Topical functions on semimodules and generalizations. Linear Algebra and its Applications, 437(10):2471– 2488, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004697.

Singer:2014:EVT

[SN14] Ivan Singer and Viorel Nitica. Extended-valued topical and anti-topical functions on semimodules. Linear Algebra and its Applications, 446(??):25–70, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008331.

Sowa:2013:FMD

[Sow13]  Artur Sowa.  Factorizing

matrices by Dirichlet multiplication. Linear Algebra and its Applications, 438(5): 2385–2393, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007355.

Sain:2013:ONA

[SP13] Debmalya Sain and Kallol Paul. Operator norm attainment and inner product spaces. Linear Algebra and its Applications, 439 (8):2448–2452, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300459X.

Spector:2011:CTZ

[Spe11] Oren Spector. A characterization of trace zero symmetric nonnegative 5 × 5 matrices. Linear Algebra and its Applications, 434(4):1000– 1017, February 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanimirovic:2012:SCI

[SPKS12] Predrag S. Stanimirović, Dimitrios Pappas, Vasilios N. Katsikis, and Ivan P. Stanimirović. Symbolic computation of A_{T,S}^{(2)}-inverses using QDR factorization. Linear Algebra and its Applications, 437(6):1317–1331, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003023.

Song:2014:IFD

[SQ14] Yisheng Song and Liqun Qi. Infinite and finite dimensional Hilbert tensors. Linear Algebra and its Applications, 451(??):1–14, June 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001670.

Sou:2012:GMA

[SR12] Kin Cheong Sou and Anders Rantzer. On generalized matrix approximation problem in the spectral norm. Linear Algebra and its Applications, 436(7):2331–2341, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007038.

Serrano:2013:DFP

[SR13a] Daniel Hernández Serrano and Fernando Pablos Romo. Determinants of finite potent endomorphisms, symbols and reciprocity laws. Linear Algebra and its Applications, 439(1):239–261, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300164X.

Serrano-Rodriguez:2013:ASM

[SR13b] Diana Marcela Serrano-Rodríguez. Absolutely γ-summing multilinear operators. Linear Algebra and its Applications, 439 (12):4110–4118, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006174.

Serrano-Rodriguez:2013:ICF

[SR13c] Diana Marcela Serrano-Rodríguez. Improving the closed formula for subpolynomial constants in the multilinear Bohnenblust–Hille inequalities. Linear Algebra and its Applications, 438(7):3124– 3138, April 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200835X.

Sra:2013:EEC

[Sra13] Suvrit Sra. Explicit eigenvalues of certain scaled trigonometric matrices. Linear Algebra and its Applications, 438(1):173–181, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005654.

So:2010:ATK

[SRdAG10] Wasin So, María Robbiano, Nair Maria Maia de Abreu, and Ivan Gutman. Applications of a theorem by Ky Fan in the theory of graph energy. Linear Algebra and its Applications, 432(9):2163– 2169, April 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Srinivasan:2013:REL

[Sri13] Murali K. Srinivasan. Radial eigenvectors of the Laplacian of the nonbinary hypercube. Linear Algebra and its Applications, 438(5): 2557–2560, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200729X.

Saez-Schwedt:2010:CAR

[SS10a] A. Sáez-Schwedt. Cyclic accessibility of reachable states characterizes von Neumann regular rings. Linear Algebra and its Applications, 433(6): 1187–1193, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sander:2010:EDP

[SS10b] J. W. Sander and T. Sander. On the eigenvalues of distance powers of circuits. Linear Algebra and its Applications, 432(12):3132–3140, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schaffrin:2010:TLS

[SS10c] Burkhard Schaffrin and Kyle Snow. Total Least-Squares regularization of Tykhonov type and an ancient racetrack in Corinth. Linear Algebra and its Applications, 432(8): 2061–2076, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Simoncini:2010:FVO

[SS10d] Valeria Simoncini and Daniel B. Szyld. On the field of values of oblique projections. Linear Algebra and its Applications, 433(4):810–818, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanimirovic:2010:ILC

[SS10e] Predrag Stanimirović and Stefan Stanimirović. Inverting linear combinations of identity and generalized Catalan matrices. Linear Algebra and its Applications, 433(7):1472– 1480, December 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saleh:2011:EMC

[SS11a] Ossama A. Saleh and Ronald L. Smith. The elliptic matrix completion problem. Linear Algebra and its Applications, 434(8):1824–1835, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sander:2011:ICG

[SS11b] J. W. Sander and T. Sander. Integral circulant graphs of prime power order with maximal energy. Linear Algebra and its Applications, 435 (12):3212–3232, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shang:2011:FZM

[SS11c] Ying Shang and Michael K. Sain. Fixed zeros in the model matching problem for systems over semirings. Linear Algebra and its Applications, 434(1): 18–43, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2011:SAE

[SS11d] Yuming Shi and Huaqing Sun. Self-adjoint extensions for second-order symmetric linear difference equations. Linear Algebra and its Applications, 434(4):903–930, February 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stampach:2011:EPP

[ŠŠ11e] F. Štampach and P. Šťovíček. On the eigenvalue problem for a particular class of finite Jacobi matrices. Linear Algebra and its Applications, 434 (5):1336–1353, March 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sato:2012:EFG

[SS12a] Iwao Sato and Seiken Saito. The edge L-function of a graph. Linear Algebra and its Applications, 436(5):1376– 1384, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005994.

Savchuk:2012:PAM

[SS12b] Yurii Savchuk and Konrad Schmüdgen. Positivstellensätze for algebras of matrices. Linear Algebra and its Applications, 436(3):758– 788, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005751.

Schwetlick:2012:NRF

[SS12c] Hubert Schwetlick and Kathrin Schreiber. Nonlinear Rayleigh functionals. Linear Algebra and its Applications, 436(10): 3991–4016, May 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379510003447.

Sato:2013:GBZ

[SS13a] Iwao Sato and Seiken Saito. A generalized Bartholdi zeta function for a regular covering of a bipartite graph. Linear Algebra and its Applications, 438(3):1025–1056, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006726.

Sesana:2013:SAI

[SS13b] Debora Sesana and Valeria Simoncini. Spectral analysis of inexact constraint preconditioning for symmetric saddle point matrices. Linear Algebra and its Applications, 438(6):2683–2700, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008270.

Setyadi:2013:EGS

[SS13c] A. Setyadi and C. K. Storm. Enumeration of graphs with the same Ihara zeta function. Linear Algebra and its Applications, 438(1):564– 572, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005599.

Sherman:2013:CNM

[SS13d] Michael D. Sherman and Ronald L. Smith. Completely normal matrices. Linear Algebra and its Applications, 439(7):2114–2122, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003911.

Sherman:2013:PNM

[SS13e] Michael D. Sherman and Ronald L. Smith. Principally normal matrices. Linear Algebra and its Applications, 438(5):2617–2627, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007276.

Stampach:2013:CFJ

[SS13f] F. Stampach and P. Stóvícek. The characteristic function for Jacobi matrices with applications. Linear Algebra and its Applications, 438(11): 4130–4155, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000839.

Sen:2014:GIE

[SS14] Mausumi Sen and Debashish Sharma. Generalized inverse eigenvalue problem for matrices whose graph is a path. Linear Algebra and its Applications, 446(??):224– 236, April 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Shan:2010:EGT

[SSGL10] Hai-Ying Shan, Jia-Yu Shao, Fei Gong, and Yue Liu. An edge grafting theorem on the energy of unicyclic and bipartite graphs. Linear Algebra and its Applications, 433(3): 547–556, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sumi:2013:TRT

[SSM13] Toshio Sumi, Toshio Sakata, and Mitsuhiro Miyazaki. Typical ranks for m×n×(m−1)n tensors with m ≤ n. Linear Algebra and its Applications, 438(2):953–958, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005842.

Shader:2014:NPM

[SSMS14] Bryan Shader, Naomi Shaked-Monderer, and Daniel B. Szyld. Nearly positive matrices. Linear Algebra and its Applications, 449(??):520– 544, May 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000561.

Shi:2013:SPS

[SSR13] Yuming Shi, Chunmei Shao, and Guojing Ren. Spectral properties of self-adjoint subspaces. Linear Algebra and its Applications, 438(1): 191–218, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006106.

Saez-Schwedt:2013:FCL

[SSS13] Andrés Sáez-Schwedt and Wiland Schmale. Feedback classification of linear systems over von Neumann regular rings. Linear Algebra and its Applications, 438(4):1852– 1862, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004575.

Seto:2014:GMR

[SST14] Michio Seto, Sho Suda, and Tetsuji Taniguchi. Gram matrices of reproducing kernel Hilbert spaces over graphs. Linear Algebra and its Applications, 445(??):56–68, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007994.

Shao:2013:SPD

[SSZ13] Jia-Yu Shao, Hai-Ying Shan, and Li Zhang. On some properties of the determinants of tensors. Linear Algebra and its Applications, 439 (10):3057–3069, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005132.

Saito:2010:DWT

[ST10a] Kichi-Suke Saito and Masaru Tominaga. A Dunkl–Williams type inequality for absolute value operators. Linear Algebra and its Applications, 432 (12):3258–3264, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saldanha:2010:CPQ

[ST10b] Nicolau C. Saldanha and Carlos Tomei. Cut-and-paste of quadriculated disks and arithmetic properties of the adjacency matrix. Linear Algebra and its Applications, 432 (9):2423–2437, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shieh:2012:CBS

[ST12] Min-Zheng Shieh and Shi-Chun Tsai. Computing the ball size of frequency permutations under Chebyshev distance. Linear Algebra and its Applications, 437 (1):324–332, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001590.

Son:2013:SCR

[ST13] Nguyen Khoa Son and Do Duc Thuan. The structured controllability radii of higher order systems. Linear Algebra and its Applications, 438(6): 2701–2716, March 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008026.

Stanic:2012:SGW

[Sta12] Zoran Stanić. Some graphs whose second largest eigenvalue does not exceed \sqrt{2}. Linear Algebra and its Applications, 437(7):1812–1820, October 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512003321.

Stanford:2014:UOS

[Sta14] Robert E. Stanford. On the uniqueness of optimal strategies in symmetric matrix games. Linear Algebra and its Applications, 452 (??):192–201, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001712.

Stevanovic:2010:RAC

[Ste10] Dragan Stevanović. Resolution of AutoGraphiX conjectures relating the index and matching number of graphs. Linear Algebra and its Applications, 433(8–10):1674–1677, December 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stevanovic:2011:TSC

[Ste11] Dragan Stevanović. Two spectral characterizations of regular, bipartite graphs with five eigenvalues. Linear Algebra and its Applications, 435 (10):2612–2625, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2013:GUH

[Ste13] Michael Stewart. A generalized unitary Hessenberg matrix. Linear Algebra and its Applications, 439(12):3807– 3821, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006277.

Storm:2011:SPG

[Sto11] Christopher Storm. Some properties of graphs determined by edge zeta functions. Linear Algebra and its Applications, 434(5):1285–1294, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stoll:2012:KSA

[Sto12] Martin Stoll. A Krylov– Schur approach to the truncated SVD. Linear Algebra and its Applications, 436(8): 2795–2806, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005349.

Stupariu:2012:FWQ

[Stu12] Mihai-Sorin Stupariu. Filtrations, weights and quiver problems. Linear Algebra and its Applications, 436(3): 648–658, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511005763.

Stuart:2013:SFM

[Stu13]  Jeffrey Stuart.  Special

families of matrices — a talk in honor of Miroslav Fiedler. Linear Algebra and its Applications, 439(4): 830–835, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200078X.

Sun:2013:ANJ

[Sun13] Xiaosong Sun. On additivenilpotency of Jacobian matrices of polynomial maps. Linear Algebra and its Applications, 439(12):3746–3751, December 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006332.

Suzuki:2013:SLC

[Suz13] Akito Suzuki. Spectrum of the Laplacian on a covering graph with pendant edges I: the one-dimensional lattice and beyond. Linear Algebra and its Applications, 439 (11):3464–3489, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005739.

Soto:2011:NPS

[SV11] M. J. Soto and J. L. Vicente. The Newton Procedure for several variables. Linear Algebra and its Applications, 435(2):255–269, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stacho:2013:HIS

[SV13] L. L. Stachó and R. Vajda. Hermite interpolation sequences over fields. Linear Algebra and its Applications, 439(1):66–77, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001638.

Sinkovic:2011:MSR

[SvdH11] John Sinkovic and Hein van der Holst. The minimum semidefinite rank of the complement of partial ktrees. Linear Algebra and its Applications, 434(6):1468– 1474, March 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seeger:2011:IEP

[SVP11] Alberto Seeger and José Vicente-Pérez. Inverse eigenvalue problems for linear complementarity systems. Linear Algebra and its Applications, 435(12):3029–3044, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sergeev:2011:BSS

[SW11] Sergeĭ Sergeev and Edouard Wagneur. Basic solutions of systems with two max-linear inequalities. Linear Algebra and its Applications, 435(7): 1758–1768, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2013:LIE

[SW13] Lingsheng Shi and Hui Wang. The Laplacian incidence energy of graphs. Linear Algebra and its Applications, 439 (12):4056–4062, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006629.

Shmueli:2012:UKM

[SWA12] Yaniv Shmueli, Guy Wolf, and Amir Averbuch. Updating kernel methods in spectral decomposition by affinity perturbations. Linear Algebra and its Applications, 437 (6):1356–1365, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003114.

Schlote:2013:HMT

[SWBS13] A. Schlote, F. Wirth, A. Berman, and R. Shorten. On the higher moments of TCP. Linear Algebra and its Applications, 439(4):899–913, August 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006088.

Song:2013:NUB

[SWT13] Haizhou Song, Qiufen Wang, and Lulu Tian. New upper bounds on the spectral radius of trees with the given number of vertices and maximum degree. Linear Algebra and its Applications, 439 (9):2527–2541, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004771.

Samoilenko:2012:KTU

[SY12] Yurii Samoilenko and Kostyantyn Yusenko. Kleiner’s theorem for unitary representations of posets. Linear Algebra and its Applications, 437 (2):581–588, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002029.

Shen:2014:SCN

[SYH14] Shu-Qian Shen, Juan Yu, and Ting-Zhu Huang. Some classes of nonsingular matrices with applications to localize the real eigenvalues of real matrices. Linear Algebra and its Applications, 447 (??):74–87, April 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001067.

Salce:2014:PEI

[SZ14] Luigi Salce and Paolo Zanardo. Products of elementary and idempotent matrices over integral domains. Linear Algebra and its Applications, 452(??):130–152, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Szechtman:2014:ENF

[Sze14a] Fernando Szechtman. Equivalence and normal forms of bilinear forms. Linear Algebra and its Applications, 443(??):245–259, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007076.

Szechtman:2014:MRH

[Sze14b] Fernando Szechtman. Modular representations of Heisenberg algebras. Linear Algebra and its Applications, 457(??):49–60, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400295X.

Szollosi:2013:CHM

[Szö13]  Ferenc Szöllősi.  Com-

plex Hadamard matrices and equiangular tight frames. Linear Algebra and its Applications, 438(4):1962–1967, February 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004460.

Tadej:2012:DKP

[Tad12] Wojciech Tadej. Defect of a Kronecker product of unitary matrices. Linear Algebra and its Applications, 436 (7):1924–1959, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006379.

Tam:2012:BRF

[Tam12] Tin-Yau Tam. Book review on Fuzhen Zhang’s Matrix Theory: Basic Results and Techniques (Universitext), second ed., xvii + 399 pp., Springer (2011) [Paperback], ISBN 978-1-4614-1098-0, e-ISBN 978-1-4614-1099-7. Linear Algebra and its Applications, 437(6):1426–1427, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002960.

Tan:2010:WTG

[Tan10a] Shang-Wang Tan. On the weighted trees with given degree sequence and positive weight set. Linear Algebra and its Applications, 433(2):380– 389, August 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2010:NMA

[Tan10b] Yi-Jia Tan. On nilpotency of matrices over antirings. Linear Algebra and its Applications, 433(8–10):1541–1554, December 15, 2010. CO-

Tan:2011:NSM

[Tan11] Yi-Jia Tan. On nilpotent subsemigroups of the matrix semigroup over an antiring. Linear Algebra and its Applications, 435(9):2247–2258, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2014:BSC

[Tan14a] Yi-Jia Tan. Bases in semimodules over commutative semirings. Linear Algebra and its Applications, 443(??): 139–152, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007234.

Tan:2014:IPS

[Tan14b] Yi-Jia Tan. Inner products on semimodules over a commutative semiring. Linear Algebra and its Applications, 460(??):151–173, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004935.

Tao:2011:NOS

[Tao11] Jiyuan Tao. A note on the Ostrowski–Schneider type inertia theorem in Euclidean Jordan algebras. Linear Algebra and its Applications, 434 (8):1902–1909, April 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tao:2013:CQP

[Tao13]  Jiyuan Tao.  On the

completely-Q property for linear transformations on Euclidean Jordan algebras. Linear Algebra and its Applications, 438(5):2054–2071, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007471.

Tian:2013:UBE

[TC13] Gui-Xian Tian and Shu-Yu Cui. On upper bounds for the energy of digraphs. Linear Algebra and its Applications, 438(12):4742–4749, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001626.

Tie:2011:RIH

[TCL11] Lin Tie, Kai-Yuan Cai, and Yan Lin. Rearrangement inequalities for Hermitian matrices. Linear Algebra and its Applications, 434(2):443– 456, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tudisco:2011:PAP

[TD11] Francesco Tudisco and Carmine Di Fiore. A preconditioning approach to the PageRank computation problem. Linear Algebra and its Applications, 435(9):2222–2246, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Talebi:2013:BBM

[TD13] Gholamreza Talebi and Mohammad Ali Dehghan. The below boundedness of matrix operators on the special Lorentz sequence spaces. Linear Algebra and its Applications, 439(8):2411–2421, October 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004199.

Tudisco:2013:ORM

[TDT13] F. Tudisco, C. Di Fiore, and E. E. Tyrtyshnikov. Optimal rank matrix algebras preconditioners. Linear Algebra and its Applications, 438(1):405–427, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006015.

Teranishi:2011:SLS

[Ter11] Yasuo Teranishi. Subgraphs and the Laplacian spectrum of a graph. Linear Algebra and its Applications, 435(5):1029– 1033, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Terwilliger:2013:FDI

[Ter13]  Paul Terwilliger.  Finite-

dimensional irreducible U_{q}(∫↕_{∈})- modules from the equitable point of view. Linear Algebra and its Applications, 439 (2):358–400, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002176.

Terwilliger:2014:BAF

[Ter14] Paul Terwilliger. Billiard Arrays and finite-dimensional irreducible U_{q}(sl_{2})-modules. Linear Algebra and its Applications, 461(??):211–270, November 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005242.

Tan:2010:VEI

[TF10] Ying-Ying Tan and Yi-Zheng Fan. The vertex (edge) independence number, vertex (edge) cover number and the least eigenvalue of a graph. Linear Algebra and its Applications, 433(4):790–795, October 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tao:2013:MEI

[TG13] J. Tao and M. Seetharama Gowda. Minimum eigenvalue inequalities for Ztransformations on proper and symmetric cones. Linear Algebra and its Applications, 438(8):3476–3489, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000438.

Tisseur:2011:DQM

[TGM11] Françoise Tisseur, Seamus D. Garvey, and Christopher Munro. Deflating quadratic matrix polynomials with structure preserving transformations. Linear Algebra and its Applications, 435(3):464– 479, August 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tao:2014:BNP

[TGS14] Jiyuan Tao, M. Seetharama Gowda, and Roman Sznajder. On the block norm-P property. Linear Algebra and its Applications, 459(??): 121–135, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004212.

Tang:2010:IGI

[TH10] Zikai Tang and Yaoping Hou. The integral graphs with index 3 and exactly two main eigenvalues. Linear Algebra and its Applications, 433(5): 984–993, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takane:2011:NFC

[TH11] Yoshio Takane and Michael A. Hunter. A new family of constrained principal component analysis (CPCA). Linear Algebra and its Applications, 434(12):2539–2555, June 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

terHorst:2013:EAE

[tHR13] Sanne ter Horst and André C. M. Ran. Equivalence after extension and matricial coupling coincide with Schur coupling, on separable Hilbert spaces. Linear Algebra and its Applications, 439(3):793–805, August 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001948.

Tian:2010:EII

[Tia10] Yongge Tian. Equalities and inequalities for inertias of Hermitian matrices with applications. Linear Algebra and its Applications, 433(1):263–296, July 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2011:SEO

[Tia11a] Gui-Xian Tian. On the skew energy of orientations of hypercubes. Linear Algebra and its Applications, 435(9):2140– 2149, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2011:MMR

[Tia11b] Yongge Tian. Maximization and minimization of the rank and inertia of the Hermitian matrix expression A − BX − (BX)^{∗} with applications. Linear Algebra and its Applications, 434(10):2109–2139, May 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2012:FCE

[Tia12]  Yongge Tian.  Formulas

for calculating the extremum ranks and inertias of a fourterm quadratic matrix-valued function and their applications. Linear Algebra and its Applications, 437(3):835– 859, August 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002522.

Tifenbach:2011:SSD

[Tif11] R. M. Tifenbach. Strongly self-dual graphs. Linear Algebra and its Applications, 435 (12):3151–3167, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Timotin:2014:SCR

[Tim14] Dan Timotin. Schur coupling and related equivalence relations for operators on a Hilbert space. Linear Algebra and its Applications, 452 (??):106–119, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001827.

Tao:2014:SMI

[TKLX14] J. Tao, Lingchen Kong, Ziyan Luo, and Naihua Xiu. Some majorization inequalities in Euclidean Jordan algebras. Linear Algebra and its Applications, 461(??):92– 122, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004996.

Tan:2013:DAP

[TL13a] Qianrong Tan and Mao Li. Divisibility among power GCD matrices and among power LCM matrices on finitely many coprime divisor chains.

Linear Algebra and its Applications, 438(3):1454–1466, February 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200660X.

Tan:2013:NIC

[TL13b] Xuezhong Tan and Bolian Liu. The nullity of k − 1cyclic graphs. Linear Algebra and its Applications, 438 (7):3144–3153, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008403.

Li:2010:SNB

[tLyLWqW10] Yao tang Li, Yan yan Li, Rui-Wu Wang, and Ya qiang Wang. Some new bounds on eigenvalues of the Hadamard product and the Fan product of matrices. Linear Algebra and its Applications, 432(2– 3):536–545, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tebbens:2014:IGC

[TMSS14] J. Duintjer Tebbens, G. Meurant, H. Sadok, and Z. Strakos. On investigating GMRES convergence using unitary matrices. Linear Algebra and its Applications, 450 (??):83–107, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001116.

Tao:2011:SCM

[TmYsH11] Min Tao, Xiao ming Yuan, and Bing sheng He. Solving a class of matrix minimization problems by linear variational inequality approaches. Linear Algebra and its Applications, 434(11):2343–2352, June 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Theja:2014:NGM

[TN14] Siva Theja and H. Narayanan. On the notion of generalized minor in topological network theory and matroids. Linear Algebra and its Applications, 458(??):1–46, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003735.

Tayebi:2012:GRF

[TNP12] Akbar Tayebi, Mohammad Shahbazi Nia, and Esmaeil Peyghan. On generalized m-th root Finsler metrics. Linear Algebra and its Applications, 437 (2):675–683, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001802.

Trigueros:2013:UEM

[TP13] Mari’a Trigueros and Edgar Possani. Using an economics model for teaching linear algebra. Linear Algebra and its Applications, 438(4):1779– 1792, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511003053.

Turkmen:2012:SIM

[TPZ12] Ramazan Turkmen, Vehbi E. Paksoy, and Fuzhen Zhang. Some inequalities of majorization type. Linear Algebra and its Applications, 437(6):1305– 1316, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002972.

Traldi:2012:LAL

[Tra12] Lorenzo Traldi. On the linear algebra of local complementation. Linear Algebra and its Applications, 436(5): 1072–1089, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004988.

Tran:2013:PRC

[Tra13] Ngoc Mai Tran. Pairwise ranking: Choice of method can produce arbitrarily different rank order. Linear Algebra and its Applications, 438(3):1012–1024, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006520.

Trench:2010:CPM

[Tre10] William F. Trench. Characterization and properties of matrices with k-involutory symmetries II. Linear Algebra and its Applications, 432(11): 2782–2797, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2011:APL

[Tre11] William F. Trench. Asymptotic preconditioning of linear homogeneous systems of differential equations. Linear Algebra and its Applications, 434(7):1631–1637, April 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2012:CPC

[Tre12] William F. Trench. Characterization and properties of (R, S_{σ})-commutative matrices. Linear Algebra and its Applications, 436(11): 4261–4278, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200122X.

Torres:2012:TMC

[TS12] Maria M. Torres and Pedro C. Silva. Tensors, matchings and codes. Linear Algebra and its Applications, 436(6):1606–1622, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511001170.

Tsai:2011:NRWb

[Tsa11] Ming Cheng Tsai. Numerical ranges of weighted shift matrices with periodic weights. Linear Algebra and its Applications, 435(9):2296–2302, November 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tsatsomeros:2012:MMM

[Tsa12]  Michael Tsatsomeros.  In

memoriam: Michael (Miki) Neumann (1946–2011). Linear Algebra and its Applications, 436(1):1–11, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004794.

Tam:2010:DPS

[TT10] Tin-Yau Tam and Mary Clair Thompson. Determinant and Pfaffian of sum of skew symmetric matrices. Linear Algebra and its Applications, 433 (2):412–423, August 1, 2010.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tomeo:2011:TAS

[TT11] V. Tomeo and E. Torrano. Two applications of the subnormality of the Hessenberg matrix related to general orthogonal polynomials. Linear Algebra and its Applications, 435(9):2314–2320, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tam:2012:DSO

[TT12] Tin-Yau Tam and Mary Clair Thompson. Determinants of sum of orbits under compact Lie group. Linear Algebra and its Applications, 436(6):1644–1650, March 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004691.

Taslaman:2013:TMP

[TTZ13] Leo Taslaman, Françoise Tisseur, and Ion Zaballa. Triangularizing matrix polynomials. Linear Algebra and its Applications, 439(7): 1679–1699, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003182.

Tam:2010:RSL

[TW10a] Bit-Shun Tam and Shu-Hui Wu. On the reduced signless Laplacian spectrum of a degree maximal graph. Linear Algebra and its Applications, 432(7):1734–1756, March 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Teufl:2010:DIL

[TW10b] Elmar Teufl and Stephan Wagner. Determinant identities for Laplace matrices. Linear Algebra and its Applications, 432(1):441–457, January 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2011:CEM

[TW11a] Yongge Tian and Hongxing Wang. Characterizations of EP matrices and weighted-EP matrices. Linear Algebra and its Applications, 434(5):1295– 1318, March 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tsai:2011:NRWa

[TW11b] Ming Cheng Tsai and Pei Yuan Wu. Numerical ranges of weighted shift matrices. Linear Algebra and its Applications, 435(2):243–254, July 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Telloni:2014:SFC

[TW14] Agnese Ilaria Telloni and Gerald Williams. Smith forms of circulant polynomial matrices. Linear Algebra and its Applications, 458(??):559–572, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004017.

Tang:2012:GLF

[TX12] Peipei Tang and Aimin Xu. Generalized Leibniz functional matrices and divided difference form of the Lagrange–Bürmann formula. Linear Algebra and its Applications, 436(3):618–630, February 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005714.

Takane:2012:TEM

[TZ12a] Yoshio Takane and Lixing Zhou. On two expressions of the MLE for a special case of the extended growth curve models. Linear Algebra and its Applications, 436 (7):2567–2577, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006483.

Tang:2012:SCG

[TZ12b] Gaohua Tang and Yiqiang Zhou. Strong cleanness of generalized matrix rings over a local ring. Linear Algebra and its Applications, 437 (10):2546–2559, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004983.

Tang:2013:CFM

[TZ13a] Gaohua Tang and Yiqiang Zhou. A class of formal matrix rings. Linear Algebra and its Applications, 438(12):4672–4688, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300150X.

Tang:2013:WEL

[TZ13b] Gaohua Tang and Yiqiang Zhou. When is every linear transformation a sum of two commuting invertible ones? Linear Algebra and its Applications, 439 (11):3615–3619, December 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300606X.

Uchiyama:2010:MSO

[Uch10] Mitsuru Uchiyama. Majorization and some operator monotone functions. Linear Algebra and its Applications, 432(8): 1867–1872, April 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhlig:2013:CGC

[Uhl13] Frank Uhlig. On computing the generalized Crawford number of a matrix. Linear Algebra and its Applications, 438(4):1923–1935, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511004733.

Uschmajew:2013:GAU

[UV13] André Uschmajew and Bart Vandereycken. The geometry of algorithms using hierarchical tensors. Linear Algebra and its Applications, 439(1):133–166, July 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002115.

Urenda:2011:MAM

[UW11] Julio C. Urenda and Piotr J. Wojciechowski. Matrix algebras with Multiplicative Decomposition Property. Linear Algebra and its Applications, 434(4):931–943, February 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Urschel:2014:SBG

[UZ14] John C. Urschel and Ludmil T. Zikatanov. Spectral bisection of graphs and connectedness. Linear Algebra and its Applications, 449(??):1–16, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000743.

Vandebril:2010:TMU

[Van10] Raf Vandebril. On tridiagonal matrices unitarily equivalent to normal matrices. Linear Algebra and its Applications, 432(12):3079–3099, July 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vassiliou:2014:SMM

[Vas14] P.-C. G. Vassiliou. Semi-Markov migration process in a stochastic market in credit risk. Linear Algebra and its Applications, 450 (??):13–43, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001293.

Vinjamoor:2010:IFI

[VB10] Harsh Vinjamoor and Madhu N. Belur. Impulse free interconnection of dynamical systems. Linear Algebra and its Applications, 432(2–3):637– 660, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vonBelow:2013:SHS

[vBM13] Joachim von Below and Delio Mugnolo. The spectrum of the Hilbert space valued second derivative with general self-adjoint boundary conditions. Linear Algebra and its Applications, 439(7):1792–1814, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003376.

vanDam:2014:LSE

[vDF14] E. R. van Dam and M. A. Fiol. The Laplacian spectral excess theorem for distanceregular graphs. Linear Algebra and its Applications, 458(??):245–250, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003656.

vanderHolst:2013:IIP

[vdH13] Hein van der Holst. The inverse inertia problem for the complements of partial k-trees. Linear Algebra and its Applications, 439(7): 2167–2175, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

vanderHolst:2014:ISG

[vdH14] Hein van der Holst. The inertia sets of graphs with a 2-separation. Linear Algebra and its Applications, 463(??):95–114, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005989.

vanDam:2011:GWN

[vDO11] E. R. van Dam and G. R. Omidi. Graphs whose normalized Laplacian has three eigenvalues. Linear Algebra and its Applications, 435 (10):2560–2569, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vinagre:2013:MLE

[VDVJT13] Cybele T. M. Vinagre, Renata R. Del-Vecchio, Dagoberto A. R. Justo, and Vilmar Trevisan. Maximum Laplacian energy among threshold graphs. Linear Algebra and its Applications, 439 (5):1479–1495, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003066.

vanderWoude:2014:SPP

[vdW14] Jacob van der Woude. A straightforward proof of the polynomial factorization of a positive semi-definite polynomial matrix. Linear Algebra and its Applications, 456 (??):214–220, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000809.

Vijayakumar:2014:FLG

[Vij14] G. R. Vijayakumar. From finite line graphs to infinite derived signed graphs. Linear Algebra and its Applications, 453(??):84–98, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001918.

Vladimir:2012:ESS

[Vla12] Balan Vladimir. Erratum to “Spectra of symmetric tensors and m-root Finsler models” [Linear Algebra Appl. 436(5) (2012) 1061–1071]. Linear Algebra and its Applications, 437(4):1182, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002728.

See [Bal12b].

Voynov:2013:SPP

[Voy13] Andrey Voynov. Shortest positive products of nonnegative matrices. Linear Algebra and its Applications, 439 (6):1627–1634, September 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003017.

Varga:2012:ANM

[VR12] Richard S. Varga and Alessandro Rizzo. An application of nonnegative matrices to the synchronization of chaotic oscillators. Linear Algebra and its Applications, 436(2): 265–275, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951000354X.

Valcher:2010:AEC

[VS10] Maria Elena Valcher and Paolo Santesso. Asymptotic exponential cones of Metzler matrices and their use in the solution of an algebraic problem. Linear Algebra and its Applications, 432(4):980– 1006, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verde-Star:2011:ITM

[VS11]  Luis Verde-Star.  Infinite

triangular matrices, q-Pascal matrices, and determinantal representations. Linear Algebra and its Applications, 434 (1):307–318, January 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verde-Star:2013:CCC

[VS13] Luis Verde-Star. Characterization and construction of classical orthogonal polynomials using a matrix approach. Linear Algebra and its Applications, 438(9): 3635–3648, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000761.

Dooren:2014:SEP

[VS14a] Paul Van Dooren and Gilbert Strang. Symmetric elimination without pivoting. Linear Algebra and its Applications, 452(??):40–45, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001700.

Verde-Star:2014:RCD

[VS14b] Luis Verde-Star. Recurrence coefficients and difference equations of classical discrete orthogonal and q-orthogonal polynomial sequences. Linear Algebra and its Applications, 440(??):293–306, January 1, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006964.

Vsemirnov:2012:ERC

[Vse12] Maxim Vsemirnov. On the evaluation of R. Chapman’s ‘evil determinant’. Linear Algebra and its Applications, 436(11):4101–4106, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006239.

Vandanjav:2014:NRS

[VU14] Adiyasuren Vandanjav and Batzorig Undrakh. On the numerical range of some weighted shift matrices and operators. Linear Algebra and its Applications, 449 (??):76–88, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000858.

Vulanovic:2012:SFD

[Vul12] Relja Vulanović. Stability of a finite-difference discretization of a singular perturbation problem. Linear Algebra and its Applications, 436(2):326–334, January 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511002692.

Vassiliou:2013:ABS

[VV13] P.-C. G. Vassiliou and Aglaia Vasileiou. Asymptotic behaviour of the survival probabilities in an inhomogeneous semi-Markov model for the migration process in credit risk. Linear Algebra and its Applications, 438 (7):2880–2903, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951200849X.

Vale:2010:SGF

[VW10a] Richard Vale and Shayne Waldron. The symmetry group of a finite frame. Linear Algebra and its Applications, 433(1): 248–262, July 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vartziotis:2010:CPS

[VW10b] Dimitris Vartziotis and Joachim Wipper. Characteristic parameter sets and limits of circulant Hermitian polygon transformations. Linear Algebra and its Applications, 433(5):945–955, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wenzel:2010:ICI

[WA10] David Wenzel and Koenraad M. R. Audenaert. Impressions of convexity: an illustration for commutator bounds. Linear Algebra and its Applications, 433(11–12):1726–1759, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wada:2014:SWC

[Wad14] Shuhei Wada. Some ways of constructing Furuta-type inequalities. Linear Algebra and its Applications, 457 (??):276–286, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003243.

Wagneur:2011:WET

[Wag11]  Edouard Wagneur.  The

Whitney embedding theorem for tropical torsion modules: Classification of tropical modules. Linear Algebra and its Applications, 435(7):1786– 1795, October 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wilson:2013:EEB

[WAH13] Richard C. Wilson, Furqan Aziz, and Edwin R. Hancock. Eigenfunctions of the edge-based Laplacian on a graph. Linear Algebra and its Applications, 438(11): 4183–4189, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513000566.

Waldron:2011:FVS

[Wal11a] Shayne Waldron. Frames for vector spaces and affine spaces. Linear Algebra and its Applications, 435(1):77– 94, July 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Walker:2011:RAA

[Wal11b] Ronald A. Walker. Regarding an adaptive algorithm for testing multivariate linear dependence. Linear Algebra and its Applications, 434(2):605– 613, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Walker:2014:NGB

[Wal14] Stephen G. Walker. A note on geometric bounds for eigenvalues. Linear Algebra and its Applications, 457(??):400– 407, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003607.

Wang:2011:AMM

[Wan11a] Yu Wang. Additivity of multiplicative maps on triangular rings. Linear Algebra and its Applications, 434(3):625– 635, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-

Wang:2011:HPM

[Wan11b]  Zeng-Qi Wang.  on hy-

brid preconditioning methods for large sparse saddle-point problems. Linear Algebra and its Applications, 434(11): 2353–2366, June 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2014:DAS

[Wan14a] Weiping Wang. A determinantal approach to Sheffer sequences. Linear Algebra and its Applications, 463 (??):228–254, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005941.

Wang:2014:LDU

[Wan14b] Yu Wang. Lie n-derivations of unital algebras with idempotents. Linear Algebra and its Applications, 458(??): 512–525, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400398X.

Waterhouse:2013:BMT

[Wat13] William C. Waterhouse. Bhattacharyya’s matrix theorem. Linear Algebra and its Applications, 438(1):7– 9, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005071.

Wang:2011:NSL

[WB11] Jianfeng Wang and Francesco Belardo. A note on the signless Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 435 (10):2585–2590, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2012:SCG

[WB12] JianFeng Wang and F. Belardo. Spectral characterizations of graphs with small spectral radius. Linear Algebra and its Applications, 437 (10):2408–2416, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004764.

Wang:2013:GWL

[WBHM13] Jianfeng Wang, Francesco Belardo, QiongXiang Huang, and Enzo M. Li Marzi. On graphs whose Laplacian index does not exceed 4.5. Linear Algebra and its Applications, 438(4):1541–1550, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Wang:2013:GET

[WBWH13] JianFeng Wang, Francesco Belardo, Wei Wang, and QiongXiang Huang. On graphs with exactly three Qeigenvalues at least two. Linear Algebra and its Applications, 438(7):2861–2879, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512008397.

Wu:2011:SIP

[WC11] Deyu Wu and Alatancang Chen. Spectral inclusion properties of the numerical range in a space with an indefinite metric. Linear Algebra and its Applications, 435 (5):1131–1136, September 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2012:PDI

[WC12] Zhou Wang and Jianlong Chen. Pseudo Drazin inverses in associative rings and Banach algebras. Linear Algebra and its Applications, 437 (6):1332–1345, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003151.

Weng:2013:SLS

[WCKL13] Peter Chang-Yi Weng, Eric King-Wah Chu, Yueh-Cheng Kuo, and Wen-Wei Lin. Solving large-scale nonlinear matrix equations by doubling. Linear Algebra and its Applications, 439(4):914– 932, August 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006180.

Wang:2010:PSS

[WD10] Qing-Wen Wang and Chang-Zhou Dong. Positive solutions to a system of adjointable operator equations over Hilbert C^{∗}-modules. Linear Algebra and its Applications, 433(7): 1481–1489, December 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Weiner:2010:OSM

[Wei10] Mihály Weiner. On orthogonal systems of matrix algebras. Linear Algebra and its Applications, 433(3):520– 533, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Weis:2011:QCS

[Wei11] Stephan Weis. Quantum convex support. Linear Algebra and its Applications, 435 (12):3168–3188, December 15,

Wei:2013:UID

[Wei13a] Guangsheng Wei. The uniqueness for inverse discrete transmission eigenvalue problems. Linear Algebra and its Applications, 439(12):3699–3712, December 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006617.

Wei:2013:JMI

[Wei13b] Ying Wei. A Jacobi matrix inverse eigenvalue problem with mixed data. Linear Algebra and its Applications, 439(10):2774–2783, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513004953.

Wang:2010:LEG

[WF10] Yi Wang and Yi-Zheng Fan. The least eigenvalue of a graph with cut vertices. Linear Algebra and its Applications, 433(1):19–27, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2012:LES

[WF12] Yi Wang and Yi-Zheng Fan. The least eigenvalue of signless Laplacian of graphs under perturbation. Linear Algebra and its Applications, 436(7):2084–2092, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006276.

Wang:2014:SLG

[WF14] Long Wang and Yi-Zheng Fan. The signature of line graphs and power trees. Linear Algebra and its Applications, 448(??):264–273, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951400038X.

Wang:2011:BMD

[WFM11] Long Wang, Yizheng Fan, and Xiaobin Ma. On bilinear maps determined by rank one matrices with some applications. Linear Algebra and its Applications, 434(5):1354–1361, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2012:SEC

[WGL12] Dengyin Wang, Hui Ge, and Xiaowei Li. Solvable extensions of a class of nilpotent linear Lie algebras. Linear Algebra and its Applications, 437(1):14–25, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Huang:2012:CLT

[wH12] Hau wen Huang. The classification of Leonard triples of QRacah type. Linear Algebra and its Applications, 436(5):1442–1472, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006161.

Huang:2013:TLT

[wH13] Hau wen Huang. Two-lit trees for lit-only σ-game. Linear Algebra and its Applications, 438(3):1057–1066, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200657X.

Wikramaratna:2011:CIM

[Wik11] Roy S. Wikramaratna. The centro-invertible matrix: a new type of matrix arising in pseudo-random number generation. Linear Algebra and its Applications, 434(1):144– 151, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [Wik12].

Wikramaratna:2012:CCI

[Wik12] Roy S. Wikramaratna. Corrigendum to “The centroinvertible matrix: a new type of matrix arising in pseudorandom number generation” [Linear Algebra Appl. (2011) 144–151]. Linear Algebra and its Applications, 437(6): 1428, September 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003874.

See [Wik11].

Wildstrom:2011:DRL

[Wil11] D. Jacob Wildstrom. Dynamic resource location with tropical algebra. Linear Algebra and its Applications, 435(7):1796–1811, October 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wilson:2013:OAG

[Wil13] James B. Wilson. Optimal algorithms of Gram– Schmidt type. Linear Algebra and its Applications, 438(12):4573–4583, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513001572.

Williams:2014:SFA

[Wil14] Gerald Williams. Smith forms for adjacency matrices of circulant graphs. Linear Algebra and its Applications, 443(??):21–33, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

Wu:2012:NAI

[WJ12] Xiaoqian Wu and Erxiong Jiang. A new algorithm on the inverse eigenvalue problem for double dimensional Jacobi matrices. Linear Algebra and its Applications, 437(7):1760–1770, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003618.

Weiguo:2013:FIM

[WJT13] Li Weiguo, Li Juan, and Qiao Tiantian. A family of iterative methods for computing Moore–Penrose inverse of a matrix. Linear Algebra and its Applications, 438 (1):47–56, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005708.

Wang:2010:GGD

[WKV10] H. Wang, R. E. Kooij, and P. Van Mieghem. Graphs with given diameter maximizing the algebraic connectivity. Linear Algebra and its Applications, 433(11–12):1889– 1908, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2010:PBI

[WL10a] Musheng Wei and Sitao Ling. On the perturbation bounds of g-inverses and oblique projections. Linear Algebra and its Applications, 433(11– 12):1778–1792, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2010:LOS

[WL10b] Yaping Wu and Huiqing Liu. Lexicographical ordering by spectral moments of trees with a prescribed diameter. Linear Algebra and its Applications, 433(11–12):1707– 1713, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2012:LEL

[WL12] Weizhong Wang and Yanfeng Luo. On Laplacian-energy-like invariant of a graph. Linear Algebra and its Applications, 437(2):713–721, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512001942.

Wang:2011:IED

[WLG11] Dengyin Wang, Xiaowei Li, and Hui Ge. Idempotent elements determined matrix algebras. Linear Algebra and its Applications, 435(11): 2889–2895, December 1, 2011.

Wang:2012:MDA

[WLG12] Dengyin Wang, Xiaowei Li, and Hui Ge. Maps determined by action on identityproduct elements. Linear Algebra and its Applications, 436(1):112–119, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004800.

Wang:2011:ASC

[WLGM11] Kaishun Wang, Fenggao Li, Jun Guo, and Jianmin Ma. Association schemes coming from minimal flats in classical polar spaces. Linear Algebra and its Applications, 435(1): 163–174, July 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2010:BSQ

[WLHL10] Shiying Wang, Jing Li, Wei Han, and Shangwei Lin. The base sets of quasi-primitive zero-symmetric sign pattern matrices with zero trace. Linear Algebra and its Applications, 433(3):595–605, September 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wen:2014:FPF

[WLL14] Ching-Feng Wen, Chia-Cheng Liu, and Yung-Yih Lur. Fixed points of functions with maxweighted quasi-arithmetic mean operator. Linear Algebra and its Applications, 445(??):302–315, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008045.

Wang:2011:GDT

[WLLX11] Wei Wang, Feng Li, Hongliang Lu, and Zongben Xu. Graphs determined by their generalized characteristic polynomials. Linear Algebra and its Applications, 434(5):1378– 1387, March 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2013:BRG

[WML13] Wei Wang, Lihuan Mao, and Hongliang Lu. On biregular graphs determined by their generalized characteristic polynomials. Linear Algebra and its Applications, 438(7):3076–3084, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512008452.

Wong:2014:GAZ

[WMZ14] Dein Wong, Xiaobin Ma, and Jinming Zhou. The group of automorphisms of a zero-divisor graph based on rank one upper triangular matrices. Linear Algebra and its Applications, 460 (??):242–258, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004923.

Wicker:2013:NDM

[WNM13] Nicolas Wicker, Canh Hao Nguyen, and Hiroshi Mamitsuka. A new dissimilarity measure for comparing labeled graphs. Linear Algebra and its Applications, 438(5):2331–2338, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007318.

Wodka:2014:SSF

[Wód14] Julia Wódka. Subsets of some families of real functions and their algebrability. Linear Algebra and its Applications, 459(??):454–464, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004649.

Wojcik:2014:MAT

[Wój14a] Pawel Wójcik. On mappings approximately transferring relations in finite-dimensional normed spaces. Linear Algebra and its Applications, 460(??):125–135, November 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004972.

Wojcik:2014:SRO

[Wój14b] Pawel Wójcik. On some restrictions of an operator to an invariant subspace. Linear Algebra and its Applications, 450(??):1–6, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514001281.

Wolf:2012:GFM

[Wol12] Reinhard Wolf. On the gap of finite metric spaces of pnegative type. Linear Algebra and its Applications, 436(5):1246–1257, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511006148.

Worawannotai:2013:DPG

[Wor13] Chalermpong Worawannotai. Dual polar graphs, the quantum algebra U_{q}(∫↕_{∈}), and Leonard systems of dual q-Krawtchouk type. Linear Algebra and its Applications, 438(1):443–497, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Woracek:2014:RKA

[Wor14] Harald Woracek. Reproducing kernel almost Pontryagin spaces. Linear Algebra and its Applications, 461 (??):271–317, November 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514005011.

Wang:2012:LGL

[WS12] Jianfeng Wang and Shuning Shi. The line graphs of lollipop graphs are determined by their spectra. Linear Algebra and its Applications, 436(7):2630–2637, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951100721X.

See comments [WY14a].

Wang:2012:OTA

[WT12] Xing-Ke Wang and Shang-Wang Tan. Ordering trees by algebraic connectivity. Linear Algebra and its Applications, 436(9):3684–3691, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000596.

Tan:2013:LCG

[wTlW13] Shang wang Tan and Qi long Wang. On the Laplacian coefficients of graphs under some transformations. Linear Algebra and its Applications, 439 (10):2746–2761, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513005077.

Tan:2012:LCT

[wTmS12] Shang wang Tan and Tian mei Song. On the Laplacian coefficients of trees with a perfect matching. Linear Algebra and its Applications, 436(3):595–617, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005829.

Wu:2010:MWS

[Wu10a] Honglin Wu. On the 0–1 matrices whose squares are 0– 1 matrices. Linear Algebra and its Applications, 432(11): 2909–2924, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2010:GPB

[Wu10b] Huazhang Wu. Generalized polynomial Bezoutian with respect to a Jacobson chain basis over an arbitrary field. Linear Algebra and its Applications, 432(12):3351–3360, July 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2013:CAI

[Wu13a] Junhua Wu. Conics arising from internal points and their binary codes. Linear Algebra and its Applications, 439(2):422–434, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002474.

Wu:2013:QFL

[Wu13b] Mingzhong Wu. Quasi R_{n} filiform Lie algebras. Linear Algebra and its Applications, 439(5):1203–1220, September 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002206.

Wulling:2013:DIS

[Wül13]  Wolfgang Wülling.  The

Drazin inverse of a singular, unreduced tridiagonal matrix. Linear Algebra and its Applications, 439(10):2736– 2745, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951300551X.

Wang:2010:CHS

[WW10] Qing-Wen Wang and Zhong-Cheng Wu. Common Hermitian solutions to some operator equations on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 432(12): 3159–3171, July 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2013:DNR

[WW13a] Kuo-Zhong Wang and Pei Yuan Wu. Diagonals and numerical ranges of weighted shift matrices. Linear Algebra and its Applications, 438(1): 514–532, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006064.

Wang:2013:JHU

[WW13b] Yao Wang and Yu Wang. Jordan homomorphisms of upper triangular matrix rings. Linear Algebra and its Applications, 439(12):4063–4069, December 15, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006058.

See correction [DWW14].

Wang:2013:MLD

[WW13c] Yao Wang and Yu Wang. Multiplicative Lie n-derivations of generalized matrix algebras. Linear Algebra and its Applications, 438(5):2599– 2616, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007859.

Wasson:2013:NDS

[WW13d] Ryan D. Wasson and Hugo J. Woerdeman. The normal defect of some classes of matrices. Linear Algebra and its Applications, 438(8): 3530–3546, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000529.

Wang:2013:DTA

[WWD13] Yao Wang, Yu Wang, and Yiqiu Du. n-Derivations of triangular algebras. Linear Algebra and its Applications, 439(2):463–471, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002279.

Wang:2010:CNPb

[WWG10] Kuo-Zhong Wang, Pei Yuan Wu, and Hwa-Long Gau. Crawford numbers of powers of a matrix. Linear Algebra and its Applications, 433 (11–12):2243–2254, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2011:HDT

[WX11] Feng Wei and Zhankui Xiao. Higher derivations of triangular algebras and its generalizations. Linear Algebra and its Applications, 435(5):1034– 1054, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2014:GME

[WX14] Wen-Huan Wang and Wei-Wei Xu. Graphs with the maximal Estrada indices. Linear Algebra and its Applications, 446(??):314–328, April 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008422.

Wang:2010:SLB

[WXH10a] Lin Wang, Mao-Zhi Xu, and Ting-Zhu Huang. Some lower bounds for the spectral radius of matrices using traces. Linear Algebra and its Applications, 432(4):1007–1016, February 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2010:CNPa

[WXH10b] Yingnan Wang, Naihua Xiu, and Jiye Han. On cone of nonsymmetric positive semidefinite matrices. Linear Algebra and its Applications, 433 (4):718–736, October 1, 2010.

Xu:2014:MGM

[wXL14] Wei wei Xu and Hao Liu. A modified general modulusbased matrix splitting method for linear complementarity problems of H-matrices. Linear Algebra and its Applications, 458(??):626–637, October 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003887.

See corrigendum [wXZ19].

Xu:2019:CSM

[wXZ19] Wei wei Xu and Lei Zhu. Corrigendum to “A modified general modulus-based matrix splitting method for linear complementarity problems of H-matrices” [Linear Algebra Appl. 458 (2014) 626–637]. Linear Algebra and its Applications, 561(??): 204–206, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379518304191.

See [wXL14].

Williams:2013:IMB

[WY13] Joseph J. Williams and Qiang Ye. Infinite matrices bounded on weighted ℓ^{1} spaces. Linear Algebra and its Applications, 438(12):4689–4700, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513001547.

Wang:2014:CSL

[WY14a] JianFeng Wang and Juan Yan. Comments to “The line graphs of lollipop graphs are determined by their spectra”. Linear Algebra and its Applications, 440(??): 342–344, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006642.

See [WS12].

Wu:2014:GAK

[WY14b] Wenming Wu and Wei Yuan. On generators of abelian Kadison–Singer algebras in matrix algebras. Linear Algebra and its Applications, 440(??):197–205, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006654.

Wu:2012:CSA

[WZ12] Yuezhu Wu and Linsheng Zhu. Center of Schrödinger algebra and annihilators of Verma modules for Schrödinger algebra. Linear Algebra and its Applications, 437 (1):184–188, July 1, 2012.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000791.

See corrigendum [WZ14c].

Wang:2013:TEM

[WZ13a] Bin Wang and Xinyun Zhu. On the traces of elements of modular group. Linear Algebra and its Applications, 438(1):604–608, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005848.

Wang:2013:PEA

[WZ13b] Xing Tao Wang and Lei Zhang. Partial eigenvalue assignment of high order systems with time delay. Linear Algebra and its Applications, 438(5):2174–2187, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007215.

Wang:2013:DSR

[WZ13c] Yanna Wang and Bo Zhou. On distance spectral radius of graphs. Linear Algebra and its Applications, 438(8): 3490–3503, April 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000116.

Wang:2013:HPP

[WZ13d] Yi Wang and Bin Zhang. Hadamard powers of polynomials with only real zeros. Linear Algebra and its Applications, 439(10):3173– 3176, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005119.

Wu:2013:SWM

[WZ13e] Yuezhu Wu and Linsheng Zhu. Simple weight modules for Schrödinger algebra. Linear Algebra and its Applications, 438(1):559– 563, January 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512005885.

Wang:2014:LCA

[WZ14a] Yi Wang and Zhi-Hai Zhang. Log-convexity of Aigner– Catalan–Riordan numbers. Linear Algebra and its Applications, 463(??):45–55, December 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005928.

Wu:2014:ESS

[WZ14b] Gang Wu and Lu Zhang. On expansion of search subspaces for large non-Hermitian eigenproblems. Linear Algebra and its Applications, 454(??): 107–129, August 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002456.

Wu:2014:CSS

[WZ14c] Yuezhu Wu and Linsheng Zhu. Corrigendum to “Center of Schrödinger algebra and annihilators of Verma modules for Schrödinger algebra” [Linear Algebra Appl. 437

(1) (2012) 184–188]. Linear Algebra and its Applications, 445(??):372, March 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007015.

See [WZ12].

Wu:2012:UBS

[WZL12] Junliang Wu, Pingping Zhang, and Wenshi Liao. Upper bounds for the spread of a matrix. Linear Algebra and its Applications, 437(11): 2813–2822, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005113.

Wong:2013:CLG

[WZL13] Dein Wong, Min Zhu, and Wenping Lv. A characterization of long graphs of arbitrary rank. Linear Algebra and its Applications, 438 (3):1347–1355, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006829.

Wang:2014:BOG

[WZY14] Jianfeng Wang, Lu Zhao, and Chengfu Ye. Bicyclic oriented graphs with the second largest skew-energy. Linear Algebra and its Applications, 459(??):43–57, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514004224.

Xie:2013:EAT

[XC13] Jinshan Xie and An Chang. On the Z-eigenvalues of the adjacency tensors for uniform hypergraphs. Linear Algebra and its Applications, 439 (8):2195–2204, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004679.

Xu:2013:RWM

[XCS13] Qingxiang Xu, Yonghao Chen, and Chuanning Song. Representations for weighted Moore–Penrose inverses of partitioned adjointable operators. Linear Algebra and its Applications, 438 (1):10–30, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200568X.

Cai:2011:AMP

[xCwXL11] Li xia Cai, Wei wei Xu, and Wen Li. Additive and multiplicative perturbation bounds for the Moore–Penrose inverse. Linear Algebra and its Applications, 434(2):480– 489, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xie:2012:CNN

[XD12] Huiqing Xie and Hua Dai. On condition numbers of a nondefective multiple eigenvalue of a nonsymmetric matrix pencil. Linear Algebra and its Applications, 437(7): 1628–1640, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003357.

Xu:2010:SLO

[XDFL10] Xiao-Ming Xu, Hong-Ke Du, Xiaochun Fang, and Yuan Li. The supremum of linear operators for the *-order. Linear Algebra and its Applications, 433(11–12):2198–2207, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xue:2011:FIS

[XE11] Fei Xue and Howard C. Elman. Fast inexact subspace iteration for generalized eigenvalue problems with spectral transformation. Linear Algebra and its Applications, 435 (3):601–622, August 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2011:EET

[XF11] Kexiang Xu and Lihua Feng. Extremal energies of trees with a given domination number. Linear Algebra and its Applications, 435(10):2382– 2393, November 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2013:OGW

[XG13] Guang-Hui Xu and Shi-Cai Gong. On oriented graphs whose skew spectral radii do not exceed 2. Linear Algebra and its Applications, 439 (10):2878–2887, November 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513005193.

Xie:2010:P

[XJ10] Baohua Xie and Yueping Jiang. PU. Linear Algebra and its Applications, 433 (11–12):2168–2177, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2013:CSN

[XLG^{+}13] Wei-Ru Xu, Ying-Jie Lei, Xian-Ming Gu, Yong Lu, and Yan-Ru Niu. Comment on “A note on the inverse eigenvalue problem for symmetric doubly stochastic matrices”. Linear Algebra and its Applications, 439(8):2256–2262, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003881.

See [HP04, Fan10b].

Xu:2011:LSQ

[XM11] Ying Xu and Jixiang Meng. The Laplacian spread of quasi-tree graphs. Linear Algebra and its Applications, 435(1):60–66, July 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2014:RGI

[XSH14] Qingxiang Xu, Chuanning Song, and Lili He. Representations for the group inverse of anti-triangular block operator matrices. Linear Algebra and its Applications, 443(??):191–203, February 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513007155.

Xue:2013:CAR

[XSS13] Wenjuan Xue, Chungen Shen, and Wenqiong Shao. Convergence analysis of a regularized interior point algorithm for the barrier problems with singular solutions. Linear Algebra and its Applications, 438(12):4654–4671, June 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300133X.

Xu:2013:SCQ

[XSZ13] Qingxiang Xu, Chuanning Song, and Li Zhang. Solvability of certain quadratic operator equations and representations of Drazin inverses. Linear Algebra and its Applications, 439(2):291–309, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513002188.

Xu:2011:FMK

[Xu11] Hongguo Xu. Functions of a matrix and Krylov matrices. Linear Algebra and its Applications, 434(1):185–200, January 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2012:RSK

[Xu12] Yinghong Xu. The reconstruction of a special kind of periodic Jacobi matrices. Linear Algebra and its Applications, 436(9):3618–3633, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000043.

Xu:2014:MFT

[Xu14] Bangteng Xu. Multidimensional Fourier transforms and nonlinear functions on finite groups. Linear Algebra and its Applications, 452 (??):89–105, July 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001773.

Xiao:2010:CMG

[XW10a] Zhankui Xiao and Feng Wei. Commuting mappings of generalized matrix algebras. Linear Algebra and its Applications, 433(11–12):2178–2197, December 30, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xiao:2010:JHD

[XW10b] Zhankui Xiao and Feng Wei. Jordan higher derivations on triangular algebras. Linear Algebra and its Applications, 432(10):2615–2622, May 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2011:LSG

[XW11] Li-Wen Xu and Song-Gui Wang. Linear sufficiency in a general growth curve model. Linear Algebra and its Applications, 434(2):593–604, January 15, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xiao:2012:LTD

[XW12] Zhankui Xiao and Feng Wei. Lie triple derivations of triangular algebras. Linear Algebra and its Applications, 437 (5):1234–1249, September 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512002911.

Xu:2013:ACG

[XWD13] Genjiu Xu, Wenna Wang, and Hua Dong. Axiomatization for the center-of-gravity of imputation set value. Linear Algebra and its Applications, 439(8):2205–2215, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004254.

Xu:2012:ECD

[XWS12] Qingxiang Xu, Yimin Wei, and Chuanning Song. Explicit characterization of the Drazin index. Linear Algebra and its Applications, 436 (7):2273–2298, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007415.

Xu:2012:TLC

[XX12] Chuang Xu and Runzhang Xu. Tripotency of a linear combination of two involutory matrices and a tripotent matrix that mutually commute. Linear Algebra and its Applications, 437(9): 2091–2109, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004399.

See corrigendum [Kis15].

Xiao:2014:SBI

[XX14] Mingqing Xiao and Jianhong Xu. Sharp bounds of the inverse matrices resulted from five-point stencil in solving Poisson equations. Linear Algebra and its Applications, 444(??):231–245, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007635.

Xu:2011:ADC

[XY11] Jia Xu and Yong Yao. An algorithm for determining copositive matrices. Linear Algebra and its Applications, 435(11):2784–2792, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xing:2013:LEL

[XZ13a] Rundan Xing and Bo Zhou. On least eigenvalues and least eigenvectors of real symmetric matrices and graphs. Linear Algebra and its Applications, 438(5):2378–2384, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512007495.

Xing:2013:DDS

[XZ13b] Rundan Xing and Bo Zhou. On the distance and distance signless Laplacian spectral radii of bicyclic graphs. Linear Algebra and its Applications, 439(12):3955–3963, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006241.

Xing:2013:LEC

[XZ13c] Rundan Xing and Bo Zhou. On the least eigenvalue of cacti with pendant vertices. Linear Algebra and its Applications, 438(5):2256– 2273, March 1, 2013. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Xing:2014:SBS

[XZ14] Rundan Xing and Bo Zhou. Sharp bounds for the spectral radius of nonnegative matrices. Linear Algebra and its Applications, 449 (??):194–209, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000986.

Xing:2014:EGT

[XZD14] Rundan Xing, Bo Zhou, and Fengming Dong. The effect of a graft transformation on distance spectral radius. Linear Algebra and its Applications, 457(??):261–275, September 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003152.

Yamazaki:2013:EPA

[Yam13] T. Yamazaki. An elementary proof of arithmeticgeometric mean inequality of the weighted Riemannian mean of positive definite matrices. Linear Algebra and its Applications, 438(4):1564– 1569, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007865.

Yanagi:2010:URG

[Yan10a]  Kenjiro Yanagi.  Uncer-

tainty relation on generalized Wigner–Yanase–Dyson skew information. Linear Algebra and its Applications, 433(8– 10):1524–1532, December 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2010:ACF

[Yan10b] Shihai Yang. Algebraic convergence of finitely generated Kleinian groups in all dimensions. Linear Algebra and its Applications, 432(5):1147– 1151, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yan:2011:NJC

[Yan11] Dan Yan. A note on the Jacobian Conjecture. Linear Algebra and its Applications, 435(9):2110–2113, November 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yan:2014:NRM

[Yan14]  Zi-Zong Yan.  New rep-

resentations of the Moore– Penrose inverse of 2 × 2 block matrices. Linear Algebra and its Applications, 456(??):3–15, September 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006246.

Ye:2011:SI

[Ye11] Ke Ye. The stabilizer of immanants. Linear Algebra and its Applications, 435(5):1085– 1098, September 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ye:2010:MSL

[YFW10] Miao-Lin Ye, Yi-Zheng Fan, and Hai-Feng Wang. Maximizing signless Laplacian or adjacency spectral radius of graphs subject to fixed connectivity. Linear Algebra and its Applications, 433(6):1180– 1186, November 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2013:BGS

[YFW13] Guihai Yu, Lihua Feng, and Qingwen Wang. Bicyclic graphs with small positive index of inertia. Linear Algebra and its Applications, 438(5):2036–2045, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007458.

Yu:2012:CSR

[YHH12] Ber-Lin Yu, Ting-Zhu Huang, and Hong-Bo Hua. Critical sets of refined inertias for irreducible zero-nonzero patterns of orders 2 and 3. Linear Algebra and its Applications, 437(2):490–498, July 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512002017.

Yuan:2014:CPD

[YHY14] Pingzhi Yuan, Zilong He, and Lihua You. A conjecture on the primitive degree of tensors. Linear Algebra and its Applications, 450 (??):175–185, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514001360.

Yanagihara:2012:BCA

[YiKIS12] Hirokazu Yanagihara, Ken ichi Kamo, Shinpei Imori, and Kenichi Satoh. Biascorrected AIC for selecting variables in multinomial logistic regression models. Linear Algebra and its Applications, 436(11):4329–4341, June 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512000821.

Yuan:2012:EKI

[YJ12] Jiangtao Yuan and Guoxing Ji. Extensions of Kadison’s inequality on positive linear maps. Linear Algebra and its Applications, 436(3): 747–752, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

You:2010:MLS

[YL10] Zhifu You and Bolian Liu. The minimum Laplacian spread of unicyclic graphs. Linear Algebra and its Applications, 432(2–3):499–504, January 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2012:MKS

[YM12] Chao Yang and Juan C. Meza. Minimizing the Kohn– Sham total energy for periodic systems. Linear Algebra and its Applications, 436(8):2764–2779, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005362.

Peng:2011:CIE

[yPjXL11] Xiang yang Peng, Hui jun Xiong, and Wei Liu. The constrained inverse eigenvalue problem and its approximation for normal matrices. Linear Algebra and its Applications, 435(12):3115–3123, December 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2013:SBS

[YS13] Lihua You and Jian Shen. A survey on bases of sign pattern matrices. Linear Algebra and its Applications, 439 (2):346–357, July 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513002152.

Shu:2011:BSS

[ySpW11] Qian yu Shu and Xue ping Wang. Bases in semilinear spaces over zerosumfree semirings. Linear Algebra and its Applications, 435(11): 2681–2692, December 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shu:2012:SOV

[ySpW12] Qian yu Shu and Xue ping Wang. Standard orthogonal vectors in semilinear spaces and their applications. Linear Algebra and its Applications, 437(11):2733–2754, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512005137.

Shu:2014:CBS

[ySpW14] Qian yu Shu and Xue ping Wang. The cardinality of bases in semilinear spaces over commutative semirings. Linear Algebra and its Applications, 459(??):83–100, October 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004182.

Yan:2013:LTS

[YT13a] Dan Yan and Guoping Tang. The linear triangularizability of some Keller maps. Linear Algebra and its Applications, 438(9):3649–3653, May 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513000049.

Yan:2013:ADS

[YT13b] Wen Yan and Tin-Yau Tam. Anti-diagonals of symmetric and skew symmetric matrices with prescribed eigenvalues. Linear Algebra and its Applications, 438(3):1446– 1453, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006623.

Yu:2013:LDE

[Yu13]  Guanglong Yu.  On the

least distance eigenvalue of a graph. Linear Algebra and its Applications, 439(8): 2428–2433, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951300428X.

Yuan:2014:MIF

[Yua14] Xiying Yuan. Maxima of the Q-index: forbidden odd cycles. Linear Algebra and its Applications, 458(??): 207–216, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514003796.

See corrigendum [Yua15].

Yuan:2015:CSQ

[Yua15] Xiying Yuan. Corrigendum to “Maxima of the Q-index: forbidden odd cycles” [Linear Algebra Appl. 458 (2014) 207–216]. Linear Algebra and its Applications, 465(??): 426–429, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514006119.

See [Yua14].

Yang:2010:SNP

[YW10] Xiaodan Yang and Yuwen Wang. Some new perturbation theorems for generalized inverses of linear operators in Banach spaces. Linear Algebra and its Applications, 433 (11–12):1939–1949, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2011:PNP

[YW11] Lihua You and Yuhan Wu. Primitive non-powerful symmetric loop-free signed digraphs with given base and minimum number of arcs. Linear Algebra and its Applications, 434(5):1215–1227, March 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2011:SBS

[YWS11a] Guanglong Yu, Yarong Wu, and Jinlong Shu. Sharp bounds on the signless Laplacian spectral radii of graphs. Linear Algebra and its Applications, 434(3):683–687, February 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2011:SLS

[YWS11b] Guanglong Yu, Yarong Wu, and Jinlong Shu. Signless Laplacian spectral radii of graphs with given chromatic number. Linear Algebra and its Applications, 435(8):1813– 1822, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yuan:2013:LPI

[YWX13] Shi-Fang Yuan, Qing-Wen Wang, and Zhi-Ping Xiong. Linear parameterized inverse eigenvalue problem of bisymmetric matrices. Linear Algebra and its Applications, 439(7):1990–2007, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003686.

Yang:2013:EM

[YX13] Shangjun Yang and Changqing Xu. On extreme U_{1} matrices. Linear Algebra and its Applications, 438(10): 3905–3912, May 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008585.

Special issue in honor of Abraham Berman, Moshe Goldberg, and Raphael Loewy.

Yang:2014:CSL

[YY14a] Jieshan Yang and Lihua You. On a conjecture for the signless Laplacian eigenvalues. Linear Algebra and its Applications, 446(??):115– 132, April 1, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513008409.

Yuan:2014:ST

[YY14b] Pingzhi Yuan and Lihua You. On the similarity of tensors. Linear Algebra and its Applications, 458(??): 534–541, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Yuan:2014:SRT

[YY14c] Pingzhi Yuan and Lihua You. Some remarks on P , P_{0}, B and B_{0} tensors. Linear Algebra and its Applications, 459(??):511–521, October 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004947.

Yu:2010:NLD

[YZ10] Weiyan Yu and Jianhua Zhang. Nonlinear Lie derivations of triangular algebras. Linear Algebra and its Applications, 432(11):2953–2960, June 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ye:2011:IIS

[YZ11] Qiang Ye and Ping Zhang. Inexact inverse subspace iteration for generalized eigenvalue problems. Linear Algebra and its Applications, 434(7):1697– 1715, April 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ye:2012:EBA

[YZ12a] Qiang Ye and Weifeng Zhi. Eigenvalue bounds for an alignment matrix in manifold learning. Linear Algebra and its Applications, 436(8):2944–2962, April 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005192.

Yu:2012:NLD

[YZ12b] Weiyan Yu and Jianhua Zhang. Nonlinear *-Lie derivations on factor von Neumann algebras. Linear Algebra and its Applications, 437 (8):1979–1991, October 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004235.

Yin:2013:RDT

[YZ13] Hongbo Yin and Shunhua Zhang. Representation dimensions of triangular matrix algebras. Linear Algebra and its Applications, 438(5):2004–2017, March 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512007823.

Yu:2014:IWU

[YZF14] Guihai Yu, Xiao-Dong Zhang, and Lihua Feng. The inertia of weighted unicyclic graphs. Linear Algebra and its Applications, 448(??):130–152, May 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

Zhou:2012:SBM

[ZBW12] Jiang Zhou, Changjiang Bu, and Yimin Wei. Some block matrices with signed Drazin inverses. Linear Algebra and its Applications, 437(7): 1779–1792, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003527.

Zhang:2014:RNG

[ZC14] Fuji Zhang and Zhibo Chen. Ramsey numbers, graph eigenvalues, and a conjecture of Cao and Yuan. Linear Algebra and its Applications, 458(??):526–533, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514004029.

Zhuang:2012:JLG

[ZCC12] Guifen Zhuang, Jianlong Chen, and Jian Cui. Jacobson’s lemma for the generalized Drazin inverse. Linear Algebra and its Applications, 436(3):742–746, February 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511005891.

Zappavigna:2012:ENN

[ZCKS12] A. Zappavigna, P. Colaneri, S. Kirkland, and R. Shorten. Essentially negative news about positive systems. Linear Algebra and its Applications, 436(9):3425–3442, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511008573.

Zhou:2013:CAL

[ZCQ13] Guanglu Zhou, Louis Caccetta, and Liqun Qi. Convergence of an algorithm for the largest singular value of a nonnegative rectangular tensor. Linear Algebra and its Applications, 438(2): 959–968, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004885.

Zhou:2013:SNB

[ZCWZ13] Duanmei Zhou, Guoliang Chen, Guoxing Wu, and Xiangyun Zhang. On some new bounds for eigenvalues of the Hadamard product and the Fan product of matrices. Linear Algebra and its Applications, 438 (3):1415–1426, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006787.

Zhang:2011:MPP

[ZH11a] Wen Zhang and Jinchuan Hou. Maps preserving peripheral spectrum of Jordan semitriple products of operators. Linear Algebra and its Applications, 435(6):1326–1335, September 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2011:PCI

[ZH11b] Sheng-Wei Zhou and Ting-Zhu Huang. On Perron complements of inverse N_{0}- matrices. Linear Algebra and its Applications, 434(9):2081– 2088, May 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zou:2012:SIU

[ZH12] Limin Zou and Chuanjiang He. On some inequalities for unitarily invariant norms and singular values. Linear Algebra and its Applications, 436(9):3354–3361, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007737.

Zhang:2012:BRB

[Zha12a] Fuzhen Zhang. Book review: Inequalities: Theory of Majorization and Its Applications (Springer Series in Statistics) by Albert W.

Marshall, Ingram Olkin and Barry C. Arnold, 2nd edition, Springer (2011) xxvii + 909 pp, Hardback, ISBN 978-0-387-40087-7; e-ISBN 978-0-387-68276-1. Linear Algebra and its Applications, 436(5):1535–1540, March 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006045.

See [MOA11].

Zhang:2012:DSR

[Zha12b] Xiaoling Zhang. On the distance spectral radius of some graphs. Linear Algebra and its Applications, 437(7):1930–1941, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003126.

Zhao:2012:OPI

[Zha12c]  Jingang Zhao.  An open

problem on inverse matrices from industrial organization, and a partial solution. Linear Algebra and its Applications, 437(1):294–306, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001632.

Zhao:2012:ATM

[Zha12d] Yun-Bin Zhao. An approximation theory of matrix rank minimization and its application to quadratic equations. Linear Algebra and its Applications, 437(1):77–93, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200167X.

Zhang:2013:AAP

[Zha13] Fuzhen Zhang. An analytic approach to a permanent conjecture. Linear Algebra and its Applications, 438 (4):1570–1579, February 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007105.

Zhang:2014:IED

[Zha14a] Xiaoling Zhang. The inertia and energy of distance matrices of complete kpartite graphs. Linear Algebra and its Applications, 450 (??):108–120, June 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514000974.

Zhang:2014:EMI

[Zha14b] Yun Zhang. Eigenvalue majorization inequalities for positive semidefinite block matrices and their blocks. Linear Algebra and its Applications, 446(??):216–223, April 1, 2014. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513008306.

Zhang:2014:SOC

[Zha14c] Zhihua Zhang. Some operator convex functions of several variables. Linear Algebra and its Applications, 463 (??):1–9, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379514005436.

Zhang:2013:SSL

[ZHG13] Jing-Ming Zhang, Ting-Zhu Huang, and Ji-Ming Guo. The smallest signless Laplacian spectral radius of graphs with a given clique number. Linear Algebra and its Applications, 439(9):2562–2576, November 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513004783.

Zhou:2010:SLS

[Zho10] Bo Zhou. Signless Laplacian spectral radius and Hamiltonicity. Linear Algebra and its Applications, 432(2–3): 566–570, January 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2011:LMD

[Zho11] Jiren Zhou. Linear mappings derivable at some nontrivial elements. Linear Algebra and its Applications, 435(8):1972– 1986, October 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhong:2012:SSP

[Zho12] Jin Zhong. Structural and sparsity properties of symmetric 7-matrices. Linear Algebra and its Applications, 436(7):2425–2439, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007683.

Zou:2013:IAV

[ZHQ13] Limin Zou, Chuanjiang He, and Shahid Qaisar. Inequalities for absolute value operators. Linear Algebra and its Applications, 438(1): 436–442, January 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006490.

Zhang:2014:CDN

[ZHQ14] Yanfang Zhang, Jinchuan Hou, and Xiaofei Qi. Characterizing derivations for any nest algebras on Banach spaces by their behaviors at an injective operator. Linear Algebra and its Applications, 449(??):312–333, May 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S002437951400086X.

Zhu:2010:SLS

[Zhu10a] Bao-Xuan Zhu. On the signless Laplacian spectral radius of graphs with cut vertices. Linear Algebra and its Applications, 433(5):928–933, October 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2010:UBL

[Zhu10b] Dongmei Zhu. On upper bounds for Laplacian graph eigenvalues. Linear Algebra and its Applications, 432(11): 2764–2772, June 1, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2011:BEG

[Zhu11a] Bao-Xuan Zhu. Bounds on the eigenvalues of graphs with cut vertices or edges. Linear Algebra and its Applications, 434(9):2030–2041, May 1, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2011:CSM

[Zhu11b] Jiandong Zhu. On consensus speed of multi-agent systems with double-integrator dynamics. Linear Algebra and its Applications, 434(1):294– 306, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2012:LEG

[Zhu12a] Bao-Xuan Zhu. The least eigenvalue of a graph with a given domination number. Linear Algebra and its Applications, 437(11):2713– 2718, December 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004375.

Zhu:2012:MET

[Zhu12b] Jianming Zhu. Minimal energies of trees with given parameters. Linear Algebra and its Applications, 436 (9):3120–3131, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006926.

Zhu:2012:OUG

[Zhu12c] Jianming Zhu. Oriented unicyclic graphs with the first ⌊(n − 9)/2⌋ largest skew energies. Linear Algebra and its Applications, 437(10):2630– 2649, November 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512004995.

Zhang:2013:UMU

[ZHZF13] Ling Zhang, Ting-Zhu Huang, Qing-Fang Zhu, and Lili Feng. Uni-mode uniqueness conditions for CANDECOMP/ PARAFAC decomposition of n-way arrays with linearly dependent loadings. Linear Algebra and its Applications, 439(7):1918–1928, October 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513003650.

Zimmer:2013:NLB

[Zim13] Andrew M. Zimmer. A new lower bound for the positive semidefinite minimum rank of a graph. Linear Algebra and its Applications, 438 (3):1095–1112, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379512006647.

Zivkovic:2012:EFC

[Ziv12] Miodrag Zivković. Extremal families containing no two sets and their union. Linear Algebra and its Applications, 436(4):845–849, February 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379511002977.

Zou:2010:EES

[ZJ10] Limin Zou and Youyi Jiang. Estimation of the eigenvalues and the smallest singular value of matrices. Linear Algebra and its Applications, 433(6):1203–1211, November 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2014:MEG

[ZK14] Jianbin Zhang and Haibin Kan. On the minimal energy of graphs. Linear Algebra and its Applications, 453 (??):141–153, July 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514002237.

Zhou:2011:BSL

[ZL11] Yunkai Zhou and Ren-Cang Li. Bounding the spectrum of large Hermitian matrices. Linear Algebra and its Applications, 435(3):480–493, August 1, 2011. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2011:TSM

[ZLD11] Bin Zhou, James Lam, and Guang-Ren Duan. Toward solution of matrix equation X = Af(X)B + C. Linear Algebra and its Applications, 435(6):1370–1398, September 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2014:PRP

[ZLH^{+}14] Ling Zhang, Zhongshan Li, Ting-Zhu Huang, Qing-Fang Zhu, Jian Hua, and Lihua Zhang. Periodic, reducible, powerful ray pattern matrices. Linear Algebra and its Applications, 444(??):81–88, March 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007040.

Zhang:2012:EET

[ZLL12] Dongye Zhang, Zhiping Lin, and Yongzhi Liu. On eigenvalues and equivalent transformation of trigonometric matrices. Linear Algebra and its Applications, 436 (1):71–78, January 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511004745.

Zhai:2012:SUB

[ZLW12] Mingqing Zhai, Huiqiu Lin, and Bing Wang. Sharp upper bounds on the second largest eigenvalues of connected graphs. Linear Algebra and its Applications, 437(1):236–241, July 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001322.

Zhou:2014:LSS

[ZSWB14] Jiang Zhou, Lizhu Sun, Wenzhe Wang, and Changjiang Bu. Line star sets for Laplacian eigenvalues. Linear Algebra and its Applications, 440(??):164–176, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007027.

Zuo:2010:NDS

[Zuo10] Kezheng Zuo. Nonsingularity of the difference and the sum of two idempotent matrices. Linear Algebra and its Applications, 433(2):476–482, August 1, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhai:2012:PCS

[ZW12a] Mingqing Zhai and Bing Wang. Proof of a conjecture on the spectral radius of C_{4}-free graphs. Linear Algebra and its Applications, 437(7):1641–1647, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003552.

Zhang:2012:DIL

[ZW12b] Shifang Zhang and Junde Wu. The Drazin inverse of the linear combinations of two idempotents in the Banach algebra. Linear Algebra and its Applications, 436 (9):3132–3138, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007191.

Zhao:2013:ICP

[ZWL13] Guopeng Zhao, Ligong Wang, and Ke Li. Q-integral complete r-partite graphs. Linear Algebra and its Applications, 438(3):1067–1077, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002437951200674X.

Zhang:2010:CSP

[ZWZ10] Shifang Zhang, Zhenying Wu, and Huaijie Zhong. Continuous spectrum, point spectrum and residual spectrum of operator matrices. Linear Algebra and its Applications, 433(3): 653–661, September 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zuo:2013:CCM

[ZXX13] Guoxin Zuo, Mingyuan Xia, and Tianbing Xia. Constructions of composite T - matrices. Linear Algebra and its Applications, 438(3): 1223–1228, February 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512006635.

Zhu:2010:MMS

[ZXZ10] Jun Zhu, Changping Xiong, and Hong Zhu. Multiplicative mappings at some points on matrix algebras. Linear Algebra and its Applications, 433(5):914–927, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2012:SPG

[ZY12a] Jin Zhang and Jikun Yi. A simple proof of the generalized Craig–Sakamoto theorem. Linear Algebra and its Applications, 437(3):781– 782, August 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512001991.

Zhaoa:2012:SAS

[ZY12b] Jinling Zhaoa and Qingzhi Yang. Several acceleration schemes for solving the multiple-sets split feasibility problem. Linear Algebra and its Applications, 437(7): 1648–1657, October 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379512003801.

Zhu:2014:CHO

[ZY14] Jiandong Zhu and Lijun Yuan. Consensus of highorder multi-agent systems with switching topologies. Linear Algebra and its Applications, 443(??):105–119, February 15, 2014. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513007143.

Zhang:2010:DSA

[ZYL10] Zhengzhi Zhang, Zhenghong Yang, and Cheng Li. Displacement structure approach to q-adic polynomial-Vandermonde and related matrices. Linear Algebra and its Applications, 433(2):319–332, August 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhao:2010:JAD

[ZZ10] Sha Zhao and Jun Zhu. Jordan all-derivable points in the algebra of all upper triangular matrices. Linear Algebra and its Applications, 433 (11–12):1922–1938, December 30, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zeng:2011:JHA

[ZZ11] Hongyan Zeng and Jun Zhu. Jordan higher all-derivable points on nontrivial nest algebras. Linear Algebra and its Applications, 434(2):463– 474, January 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhao:2012:JHA

[ZZ12] Jinping Zhao and Jun Zhu. Jordan higher all-derivable points in triangular algebras. Linear Algebra and its Applications, 436(9):3072– 3086, May 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511006938.

Zhang:2013:SLC

[ZZ13] Jie Zhang and Xiao-Dong Zhang. The signless Laplacian coefficients and incidence energy of bicyclic graphs. Linear Algebra and its Applications, 439(12):3859–3869, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513006605.

Zhu:2014:CTA

[ZZC14] Huihui Zhu, Xiaoxiang Zhang, and Jianlong Chen. Centralizers and their applications to generalized inverses. Linear Algebra and its Applications, 458(??):291–300, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379514003802.

Zhang:2013:MPI

[ZZCW13] Xiaoxiang Zhang, Shuangshuang Zhang, Jianlong Chen, and Long Wang. Moore– Penrose invertibility of differences and products of projections in rings with involution. Linear Algebra and its Applications, 439(12):4101– 4109, December 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379513006666.

Zhang:2011:EIT

[ZZL11] Jianbin Zhang, Bo Zhou, and Jianping Li. On Estrada index of trees. Linear Algebra and its Applications, 434(1):215– 223, January 1, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2010:ADP

[ZZW10] Lin Zhang, Jun Zhu, and Junde Wu. All-derivable points in nest algebras. Linear Algebra and its Applications, 433(1):91–100, July 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).